A Complete Bibliography of Publications in Journal of Computational and Applied Mathematics (2015–2019)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

17 February 2023 Version 1.14

Title word cross-reference

(2 + 1) [TW19]. (A, B) [CL18b]. (C, α, ρ, d) [MLS15]. (C, D) [CL18b]. (p, ν) [DP19a]. (p, q) [DP19b, KL17]. (ψ, ϕ) [RMMRC15]. 0 [PBS15]. 0 < p ≤ 1 [ZY19b]. 1 [CGSS17, CZ16, Guo17, JS17, ML17a, PBS15, WD19, XZN^{+}19, ZJG^{+}15]. 1 + \sqrt{2} [AR15c]. 10 [GK18b, GK19a]. 14 [MLSK16]. 2 [BBG19b, CJ19, CROC16, DAM15b, GC15, GLL19, GRS19, HMMV17, KR16, KV16, KV17, LW15, LW17b, LSH^{+}19, PT17a, QDBG18, ST16a, SJ17, SCG18, WW16a, WD19, Wu17]. 2.5 [DFGS16]. 2n [DXML19]. 2n − 1 [Ahn19]. 2 × 2 [EHTL19]. 3 [DMAS^{+}16, EKKT17, GD17, ID16, MS19a, RABB19, RdSLL16, SDH17, SW17d, TBAH16, VS15b, ZLS^{+}15, Zha19a, ZWZ^{+}18]. 3/2 [CP15]. 4 [ABRS17, ADD^{+}16, BAAA^{+}15, BL17, KBL19, ZMTT19]. 6 [BK19a]. ^{(}R) [MPM^{+}17]. _{N} [VFGPG^{+}18]. A [SW17a, ZCC15]. α [CFG^{+}19, SEV^{+}18]. α − ψφ [SHS^{+}15]. αβ [ZHL15]. AXB = C [LWJ15]. AXD − BCX = E [CL18b]. B [GWLL18, NS16]. BQ [NS16]. C^{0} [BySZ19, CG17b, DFS15, Tho16, YBLH17]. C^{1} [AHBC15, BDIR19a, BDIR19b, BAHC17, BK19a, GS18a, GS19, GK18b, HL18a, KB19, ZWL18]. C^{2} [BIR17a, BC16, GK19a, HL18a, KV18a, KP19]. C^{k,γ}(Ω) [Sai19]. c_{0} [DHAM17]. D [MG15c, KKV19]. Δ [MB15b]. ℓ^{1} [LZ18b]. ℓ_{1} [DHAM17, ZPL17, ZPL18]. F [FZW19, AY18, ABT16]. F_{1}(·) [DP19a]. G [Sun16, YL19a, AA18]. G^{1} [Hu16, LZ18a, MB15a, ZwKN15]. G^{2} [Ahn19, BF18, KV18a, Lu15c, LJ16]. G^{3} [LAHB15]. G^{n} [AH18]. Γ [GLZ18]. ΓQR [GLZ18]. GCROT(m, k) [MLJ16]. H [AR15b, CL19d, HM19b, KR18, KLL18, LA15b, LLV18a, ZW16]. H(div) [JM15, CDF^{+}16]. H^{+} [HZ18a]. H_{+} [LY18]. H_{0}^{1}(Ω) ↪ L^{p}(Ω) [TSMO17]. hp [BS16]. I [LP15a, TLTJ18, WL17c]. INAR(1) [QLZ19]. J [SB19]. k [ABHVM16, AAB^{+}18b, CCIL18, DTKP17, FPGK19, Kam17, MT17a, MM17, RIK16, RMVA17, Wan17a]. KAK = A^{s+1} [LRT17]. L [BK16, LaT15, LaTG17]. l^{1} [EW18]. L^{2} [BW19, FS19a, JM15, Men17, EW18]. l^{2}/l^{p} [ZY19b]. L^{∞} [LZC15]. L^{q} [NP19]. l_{0} [YSL17]. l_{1} [PKMM19, ZLAA15]. l_{2} [PKMM19]. l_{p} [CPL^{+}19a]. λ [DP16, RM19]. LDU [dBGMUA18]. LU [FNV17]. M [DLLB17, JMSL19, LLL19b, LHZ15, ZS18b, CSB18, DG16b]. M_{1} [CFSS18b]. R [KMB^{+}18]. R^{3} [DD16b, LD15]. R^{n} [VS18]. R^{s} [NN17]. Z_{2} [RFO17]. O(1/\sqrt{k}) [YSL17]. μ [JSC18]. N [BBCP16, BHMT16, WLM16, Ahn19, CCIL18, CSE^{+}19, Ery19, FPGK19, FS19a, Kam17, KPS16, MM17, RIK16, RM19, WH16, Zha18]. ω [DP16]. P [MRMA15, SVAKGR19b, ZCC15, CKL18, DWXO15, Miy18, Yao18]. p(x) [CZ17a]. p + 3 [CJSCT18]. P_{0} [WYW15]. P_{k}Q_{l} [Arm18]. p ≥ 5 [CJSCT18]. ϕ [FZW19]. ψ [KK19c]. q [BDA17, HWC16a, LJX17, Nat19]. Q_{k} [ZM16a]. Q_{k−1} [ZM16a]. QR [APS17a, JWZC18, SY19]. R [LaT15, SY18, TBS18, XLS^{+}19]. R^{3} [dFLM18]. S [GWLL18, DSZ15]. S^{1} [MGMDGVR17]. σ [ÁRS16]. T [BR16, BCDwK17, BLMS19, LCC^{+}15, MM15a, ZDLD15]. θ [GG15, JN18, LY15a, Tan18, YL19a]. U [CQW17]. US [CQW17]. v [KK19c]. ε [XF16]. φ [KS17c]. W [GPHAPR17]. X + AX̄B = C [LH19]. X − AX̂B = CY + R [SC18]. X − AXB = CY + R [SC18]. x^{−1/2} [HMMV17]. X^{p} = A + M(B + X^{−1})^{−1}M^{∗} [MK17]. Z [CS19, KLL18, LLV18a, LMUZ18].

[?]/Kuo:2018:CMC. -adaptive [XLS^{+}19]. -analogue [KL17]. -bases [JSC18]. -BEM [BS16]. -Bézier [HWC16a, LJX17, LA15b]. -body [BHMT16]. -contractions [SHS^{+}15]. -contractive [RMMRC15]. -control [FZW19]. -convexity [MLS15]. -coupled [FZW19]. -D [CROC16, LW15, XZN^{+}19]. -dimensional [WW16a]. -divergence [AY18]. -eigenpairs [CQW17, KLL18, LLV18a]. -eigenvalue [LLL19b]. -eigenvector [CS19]. -elliptic [CL19d]. -EM [LY15a, Tan18]. -equivariant [RFO17]. -expectation [Sun16]. -extension [DP19a, DP19b]. -fixed [KS17c]. -fold [WH16]. -fractional [AA18]. -framework [YL19a]. -functions [AR15b]. -generalized [DG16b]. -Hessenberg [SB19]. -homotopic [BDA17]. -hypergeometric [SCG18]. -inverse [ML17a]. -iterated [KR16]. -Lanczos [GLZ18]. -Laplacian [CZ17a, CKL18, DWXO15, Yao18]. -localized [LaT15, LaTG17]. -matrices [LY18]. -matrix [DLLB17, HZ18a, LHZ15, ZS18b]. -Meixner [MB15b]. -meshes [BR16, BLMS19, ZDLD15]. -method [YL19a]. -methods [GPHAPR17]. -Nekrasov [GWLL18]. -node [MLSK16]. -nonbacktracking [CFG^{+}19]. -norm [CPL^{+}19a, TBS18]. -optimal [MG15c]. -out-of- [CCIL18, Ery19, FPGK19, Kam17, MM17, RIK16]. -overload [ÁRS16]. -parameter [RABB19]. -periodic [SY18]. -phase [HMMV17]. -point [ABRS17, DXML19, KBL19, ZMTT19]. -projection [JM15]. -record [Wan17a]. -regularized [LZ18b]. -section [KPS16]. -simplices [KKV19]. -smooth [KV18a]. -space [ADD^{+}16, BAAA^{+}15]. -species [CSE^{+}19]. -spline [LCC^{+}15]. -splines [BCDwK17]. -split [BK19a, GK18b]. -splits [GK19a]. -stability [MM15a]. -stable [BK16, SW17a, SVAKGR19b]. -step [ABHVM16, AAB^{+}18b, RMVA17]. -step-ahead [ZJG^{+}15]. -strict [DTKP17]. -SVR [XF16]. -symmetries [RM19]. -tensor [ZW16]. -tensors [HM19b]. -th [Miy18, Zha18]. -th-order [BBCP16]. -times [SDH17]. -timescales [Nat19]. -tridiagonal [MT17a]. -type [JMSL19, LMUZ18, WLM16]. -wave [ABT16]. -widths [FS19a].

2 [SPG^{+}16b]. 2015 [HLV16]. 2017 [CPRS19, DLY18]. 260 [Beh16]. 266 [ÁD15]. 271 [CDKN15]. 278 [DNL16]. 280 [AB16]. 283 [BP17a]. 2D [GCTGC16].

305 [SW16a]. 318 [LGB18]. 319 [MHY21, ZPL18]. 325 [JSW18b]. 336 [CPL^{+}19b]. 340 [Chr19]. 343 [GMSR21]. 346 [HWXC19b]. 354 [GAA21]. 358 [BKMO20]. 3rd [LPW^{+}19].

4th [AN17b].

A-posteriori [LC17b]. a.s [ABT17]. Abel [GV17, PSK16]. ability [ICMPG16]. absence [KElY16]. absolute [AHM18, IIA15, LLW18, ME18, ZL18a]. absorbing [AMP18, HZ17a]. Accelerated [And17, NRV18, CTZK18, LC18, LSZB17, MS19a, Ou18, RP16b, Wan18c, ZHWS16, ZSW16, ZSZ17]. Accelerating [AdGvG19, EW18, HVV15, HH19b, Wan18f, Wan18g]. Acceleration [GG17a, HSS19, GHV19, SX18, Zha15a]. acceptable [WXY17]. accessibility [HVM18]. account [SS18b]. accounts [XDZZ19]. accretive [KKP^{+}19]. Accuracy [LW15, LRKV16, LW17b, ASKG18, Dos19, FFKP17, HMDSR19, MPM^{+}17, MJS19, SRH17, TYS17, VF17, Zak19, ZHC19].

Accuracy-Conserving [LRKV16]. Accurate [Hua19, MM19, Moh17, PD17, AIS^{+}17b, CL19a, CXZ19, DIS^{+}18, DIP^{+}19, DMS19, JB18, KT15, KK17, MV16c, OOO15, Sey18, SMS18]. ACOMEN [SDG15]. Acoustic [TW19, CROC16, DMCC^{+}18, GM16a, GDCK19, LLL19a, SFGJR^{+}16]. acoustics [CD16a, GYX17, KM18, Kre19, Tun18]. across [BN19]. acting [Gor16]. actions [Lu18]. active [GDC15b]. activity [FCJ16]. actuarial [DSDV15, PT17b]. Adaptation [ABT16, BBG19b]. adapted [HWH19b, IDGH18]. Adaptive [AHP19, Arà19, BBD16, CS18a, CL19d, DKL^{+}18, HWH19a, HCSC16, Kaw17, KRR18, LC15a, LAVMA18, MRVV16, MD18, PB15, SRBD18, WCS17, XZ18, ZNG15, AN17a, AN19, And16, And17, BTKM15, BJK19, Bie17, BG15, BN15, CPZ19, GKQW17, GZDC19, Gia18, GD17, JN18, KW16, KCX17, LWW17, LW15, LL16a, LW17b, LC15b, LWZ17, LN17, LN19, LSP18, MAK19, MHAA15, Mat18, MM15c, NAK17, Ou18, PSW18, Pol16, Prz19, RVC19, SGL17, SYZ19, SN19, WLLC18, WXJ19, XLS^{+}19, YN18, xZWlC15]. adaptivity [CLP16]. Add [HY19a]. Add-sub [HY19a]. addends [ZF15]. Addendum [CDKN15]. Additive [LN19, AY19, JC15, LWZ^{+}15, ZZHS17]. adequate [FH16]. ADI [DM17b, HHEG17, ZZW16]. Adiabatic [LTY17]. adjoint [AZ18, CMRP17, HL15, JZ18]. adjoints [Jan17a]. adjusted [TLTJ18]. ADMM [CLZL18, LSH18a, WS17]. Adomian [NAM18]. adsorption [CFMN15]. Advanced [ZDFF18, GC18, SDG15]. advances [KKS16]. advection [CDP17, CM16, CLA19, CEJ18, CNQR^{+}18, Dun17, GHM18, GPHAPR15, GPHAPR17, LCL16, MDAP19, SB16]. advection-diffusion [CLA19, CNQR^{+}18]. advection-reaction [SB16]. aerodynamic [VG15]. aerodynamics [GGC^{+}18]. aesthetic [LX16c]. AFC [SAT15]. AFC-stabilized [SAT15]. affected [VFN18]. Affine [CPL^{+}18, CPL^{+}19b, BW17, Fer15, HLBP15, HJS19, HZ17c, RR16]. Agarwal [AAKT19, DLY18]. age [SS18d, Wan18h]. agent [MO18a]. aggregate [NCJ16]. Aggregated [TK15]. aggregation [AL18, AGM19, SKBT16, pWmGkY^{+}19, XZ18]. agricultural [NY19]. ahead [ZJG^{+}15]. AHP [BDGIPG15]. AIDS [KJ19]. air [AVGCMVM17, CFMN15, SPG^{+}16a, SPG^{+}16b]. air-water [CFMN15]. aircraft [OLRPdCJ19, Xu17]. airfoil [FFP16]. airfoils [DS19]. Airy [Kan18, Xia19]. algebra [RANG15]. Algebraic [ELL^{+}19, LZA19, TDA18, AMRX17, ALV19, ATV15, BPD15b, wCW15a, CH19c, CGvA^{+}15, CPRS19, Els15, FwC17, GMG19, GLSW16, GLSW18, HMT^{+}17, HP17, KL15, LDDF18, LSW19, LC18, LSH^{+}18b, Mär17, Miy17, Miy19, NSW15, PB15, Pis15, SH18, Wan18b, WGZ15, WZ18, pWmGkY^{+}19, YFwC19, ZN19]. algebras [CNT17, CRR19, FFN^{+}17, LSS19]. Algorithm [PBS15, ACDT19, And16, AGR^{+}15b, AJY18, BM17a, BW16, BW17, BGNGR^{+}17, BPP17, CZ19a, CM16, CDZ16, Che16b, CSXY19, wCW15a, CSB18, CPL^{+}19a, DIS^{+}18, DIP^{+}19, DGH15, DY15, DGN19, EK19, FZ15, FP16a, FWW^{+}18, GH16a, GH19a, GC15, GGBTP17, GH17a, GW17, GJSC18, GHL18, Haj19, HLG^{+}17, HS19, HX15, HFMM16, HZ19, HZ17c, JHJB17, JCC15, JSL15, JWZC18, JGTB15, KSDM16, KDW17a, KKW16, KCX17, KLC17, KMLB17, KHS^{+}18, LX15a, LQZ16, LL16a, LRL^{+}18, LDL^{+}19, LZ16c, LX16b, LF18, LOAGBR15, LJ16, MCLW16, MT16a, MR19, MG15c, NSM16, NY19, NGGFM17, OZ18, OGR19, PSW18, PW17, PST18, RDRM15, RVC19, RMK19, Sad19, SGK16, Sar19, SW16a, SW16b, Sey18, SHW19, SNY19, SQRE16, SCP19, SBCJ18, TV18, TC19, VRBS19, WGLV16, WS17, WSL17, WZBM18, WXJ19, WCW19, WLC^{+}18, pWmGkY^{+}19, Xu15, Xu17, XY18a, XY18b]. algorithm [YSCY17, YZ15c, YWY19, ZPL17, ZPL18, ZCFW17, ZHS18, ZCC15, ZL17a, ZLAA15, dBAAA15, dBGMUA18]. Algorithmic [CBP18, APS15a, Els15, HWH19b, WL17c]. Algorithms [KKP^{+}19, LRT17, AIS^{+}17b, And17, AJ18, CQW17, CYY^{+}16, CJSCT18, DSA15, DK16, DFS15, GLdCJS17, GLZ18, HZ18a, JL17, JCH17, KPS16, Kow17, LY15b, LWZZ16, LZW18, LH18, LSH18a, LLHO18, LW19b, Lu15a, LC16, MVQDdR^{+}18, MPM^{+}17, NC16, OGR17a, OGR17b, LGB18, RF18, SGK16, SJ15, SRH17, SM17, TH18, Wu16b, YZZL19, Yao18, Yin19, ZDGM17]. alignment [WXJ19]. Allee [BYYV16]. Allen [JK18, LHY19, SJL16a]. allocation [ÁRS16, CDRY18, Iid15]. allowing [HHZ18]. alloy [ID18]. Almost [APS15a, CW15a, LWZ19a, Tan18, APS15b, APS17b, APS17a, APS19a, APS19b, EC18, KP19, Mil15]. along [GKN18]. Alpert [Sar19]. Alpha [DNK19, NKD^{+}19]. Alternating [LDG19, LQ19, CLZS19, CSXY19, HH19b, MV16a, Sad19, YSL17, ZZZ15, GC15, LLL19a]. Alternating-Direction-Implicit [GC15]. Alternation [Har18]. alternative [EDO19, MGMDGVR17, dBAAA15]. ambiguity [LWZ19b]. American [CEJ16, CEJV16, FNV17, JHK17, JYK16, KDB17, LDK17, MNFS18, ZY18]. AMF [GPHAPR15]. AMGe [lCCVV18]. Ampère [YX19]. analogue [Had15, KL17]. Analysis [AM19, ÁD15, ACCF19, BGI18, BCVV18, DMR19, ES18, DD18b, FM17, FMMQ19, GRS19, GG17b, Güc17, HJL18, JY19, Kad17, EP18, LX16a, MHV17, Mir15, OZ17, SJ18, VG17, WLW19, YMGA19, AALMLM19, AAB18a, AG16a, AJR15, AK18, AW19a, AZ17, AMZ19, AG19, AJ17, AFM19, BM17a, BM18, BJM16, BDA17, BBGG19, Ber15, BLM16, BEMA16, CDP19, CD16a, CRR16, CLA19, CV17, wCW15a, CST^{+}16, CEJV16, CFMN15, CNQR^{+}18, CPRS19, DDP16, DVS15, DMMSK17, DTKP17, DGSM15, DM17a, DLM17, DF16b, DSN17, EJJ15, FE16, FSNS16, Fer15, GHM18, GZD^{+}19, GMSR18, GMSR21, GMW18, GAA16, GM19, GHL19, GSRJ18, HH19a, HLG^{+}17, HH16, HS18c, HM19a, HH18d, ID18, ID16, Jin16, JLCZ17, KS18, KV19a, KDB18, KMD^{+}16, KJ19, KElY16, KPW17, KK15, KAS18, Lad15, LD19a, LD19b]. analysis [LA15a, LW15, LZC15, LHY16, LSZ16, LSLZ17, LYS19, LWJ15, LHZ15, LWS16, LWTQ17, LSBL18a, LSBL18b, Luo18, MJ18a, MJ18b, MMO15, MS17a, MD18, MYM16, Mar19b, Mej17, MLSK16, MS17e, Ngo19b, NTPVNTDT15, OE19, Par15, PCJ18b, PI19, QDY^{+}19, RNM16, RCN17, RQ16, RF18, SS18c, SNY19, SP17, SM15a, Sto18, SSZL16, SW17d, TA17, TDA18, Tho16, UG18, Ver16, VF17, WCS17, WCSL15, WW16a, WG18, WXY17, vtWvdHvdVV16, XGH18, XWWI18, XY19, XY18a, XY18b, XLS^{+}19, XHH^{+}19, YZPL18, YSG18, YX19, YZZ19, YAA19, YJBM19, ZNG15, yZ17, ZCZM17, ZNS18, ZL19, ZP19, ZS18c, ZWWW19, ZY19b, ZLAA15, ZC16b, ZN19, dSBV16, ÁD14]. Analytic [Hac18, KT15, AGR15a, GLR16, GD18, KSDM16, KSW19, RP16b, TBS18, XC18]. Analytical [AFdOdS15, BBB17c, FNdO15, MMS19, STZ16, TGSS17, DMCC^{+}18, HA18, HK18, KDW17b, KDW17a, LOAGBR15, SLA19, SPM15, WLM16, You19]. Analytically [ZL15c]. analyze [GH17b]. analyzing [HMKCR19]. angiogenesis [ZKRL15]. angle [AM17a, KKV19]. angles [SM15b]. angular [DMAS^{+}16, LP19]. Anisotropic [BBG19b, NLG15, CSS19, CV19, CRSV19, DKL^{+}18, Gia18, RP15]. anisotropic-adaptive [Gia18]. Annual [LPW^{+}19]. annuity [FJD17]. annular [BBQ18, SLH15]. anomalies [ACMT17a, OCU17]. Anomalous [KLM18, CCW17]. answers [Mär17]. Ant [Xu17, FP16a]. ant-local [FP16a]. anti [Bou18, GdLT17, PR15]. anti-diffusion [GdLT17]. anti-diffusive [Bou18]. anti-Gauss [PR15]. antithetics [God16]. Antwerp [SPG^{+}16a]. any [CT15]. AOR [Had15, ZWT19]. aortic [LS17a]. aperture [MA19a, PGD^{+}18]. appearing [FM19b]. Appell [DP19a, KR16]. Appl [ÁD15, AB16, Beh16, BP17a, BKMO20, Chr19, CDKN15, CPL^{+}19b, DNL16, GMSR21, GAA21, JSW18b, MHY21, LGB18, RdSMO18, SW16a, TGJ19, ZPL18]. Application [CDH^{+}17, DI15, GM16a, GTC16, GRVC15, GKPS19, KHS^{+}18, KBL19, ML17b, MLW19, MS17c, SPG^{+}16a, SB18, Xu15, ZST15, AdHR17, AHKT15, ABB^{+}15, AB16, BCFN16, BP17a, BDG15, BDSB16, CFSS18a, CZ15, CT16, CDZ16, CKB17, CTAT19, EBJS19, FGF19, GMSRP16, GAA19, GAA21, Güc17, HH19a, HX15, HHEG17, HXC19, HJL18, Kam16a, KKP^{+}19, LaTG17, LMUZ18, LZL^{+}18a, LLV18b, Luo18, ML18, Mar19a, MMRS15, MT19, NP19, Pap16a, PAC15, PBMJC18, PN16, QLT^{+}17, RVA15, Sai19, SHS^{+}15, SE18, ST16b, STS15, SCP19, TJA18, Uda18, WZ16, WS18, Xie15, XY16, Yao18, YL18d, ZZ19, ZCC15, ZZ18, SM16, VZDV18]. Applications [BBB17b, DLY18, IAAW18, SMS18, ADG19, AJR15, ABEHV16, AMS17, AFdOdS15, BDF15, BCNP15, CFR18, CMM17, CYY^{+}16, CLW17, CLA19, CAO^{+}19, CNQRR17, Dzh18, ELL^{+}19, FZW19, FLN16, FM19b, GT17, Hag15, Hai17, HWZ19, JCC15, JZ18, JSC18, JKL^{+}15, KGL19, KDB18, KKW16, KLK16a, KT15, KS17c, KCNH16, LaT15, LSS19, Lu15b, MA19b, MG17, NKD^{+}19, NW19, NN19, OAS19, PT17b, QL18, RRM18, RdSMO17, RdSMO18, TLTJ18, VDS15, WSW15, WGLV16, Xia19, XHH^{+}19, Yan17, YCF18, YZZL19, YMGA19, ZDGM17]. Applied [AJ18, BEJR17, Bog16, DVJ18, LD19a, LD19b, LZS15, MG15b, PG16, SPM15, VG17, Wan18f, ZHH^{+}19]. Applying [Mit15b, FFKP17, MH16a]. approach [ALS19, AW19a, BM18, BD19, BB18, BKLR16, BS19, BLV19, BBG19b, BS15, BAI17, COP17, CCJ16, CCJ18, CP18, CEJ16, CH19c, Cou19, DWS^{+}18, DF16b, DT17, Dzh16, DD17, ERPD18, FE16, FM18b, FJLB18, FLMS16, GLR16, GGG19, GYS17, GX17, Gun18, GSRJ18, HYF18, HYQ^{+}18, HCA^{+}18, HJ17a, HAH17, HK18, HY19b, HKZZ15, IANS15, JK19, KMRS16, KC17, KGES19, KDW17b, KV18b, KS17c, LDK17, MDT16, MD19, MD18, Mar19b, Mat18, MS15a, MB17, NCQ^{+}18, NMKK19, OD17, PR16, PGRC18a, RM17, RMVA17, RW16, RMA19, SA17, SK15, SV16b, SRA^{+}15, TLL16, VG17, VZDV18, VARC17, WZBM18, WLZS18, XLS^{+}19, YGS18, ZCFW17, Zha17a, ZS17c]. approaches [ABB18, CM19b, GT17, HA18, HA19, LWW17, SZ18]. Approaching [AGM19]. approximants [BRZ15, LP15a, Tra15, dlCYC15]. Approximate [FFP16, GLTW15, MA18, WMI16, AD18, AvG16, BWS17, CPB19, EHTL19, GLP16, GT17, HHKR15, LZ18c, LDW18, LL18, MDG17a, MDGVR17, Miy15, Prz19, Qui15, RdSMO17, RdSMO18, YCX19, ZST15, ZWGL18]. approximately [HAH17]. Approximating [AALMLM18, JGAFAL19, AH18, CT17b, HG17, HK19, OZ18, ZMTT19]. Approximation [DLS17, FM18b, GR19, KPRA19, KtTBK18, MOY18, MH16a, RRW17, AKR18, Ahn19, AEKGP19, AT17, AM19, BDFM19, BHRR18, BW19, BLP17, BCDwK17, CZ19a, Che16a, lCCVV18, CPZ19, CKB17, Cro15, CC15, DVA19, DP17b, DM17a, Din16, DSA16, EL19, FP18, FJLB18, Fuk15, GS18b, GH17a, GH19b, GS15b, GO15, GG17b, Han15, HS18a, HMMV17, HS18b, HS19, HWT15, Hu16, HXC19, IAAW18, JW17, Kaz18, KCX17, KT15, LP15b, LL16b, LCL16, LZL16, LWXF18, LZ15, LLW17, LLG^{+}18, LX16c, Luo18, MJ18a, MRVV16, MOPP18, Mil16, MKKV19, Mir15, MR15, MP17b, NT15, NSYX17, NS18, NH19, PPM16, PCSR19, PI19, Prz15, RS18, RR16, SS19a, Sey18, SSW19, SWL15, ST16b, Skr17, SI16, VŽ19, VFGPG^{+}18, WGLV16, Wol17, WAGVRB19, Yam17, YH18a, Zag18, Zha15b, ZqZwZ15, ZJG^{+}15, Zha19b]. approximation [ZZHS17, Zyg17]. Approximations [YZY17, AH16c, AC17b, BK15a, CFR18, CLC16, CZ16, DBX19, DIP^{+}19, El 17, FJD17, GD16, Gia18, JM15, LY15a, LC15a, LZC15, LRKV16, LLY19, LFKG16, MPH18, Mir16, MG17, NCJ16, NR19, OP19, PBP16, QWS15, RABB19, SW17c, SH15, Tho16, YZZL19, ZWH15a]. AQMEII [SPG^{+}16b]. AQMEII-2 [SPG^{+}16b]. aquifer [LCX18]. arbitrage [Rom18]. arbitrarily [NR15]. Arbitrary [BVZ17, AM17a, CDE18, CT17b, CJSCT18, DFS15, FP18, Hu16, HLMS19, KTT16, RVZ15, ZML16]. arbitrary-dimensional [CDE18]. arc [Zyg17]. architectures [GDC15a, LRL^{+}18]. arcs [AH18, VŽ19, Zag18]. Area [MN19b, Zag18]. Area-preserving [MN19b]. areas [FAVGC^{+}15]. arguments [LL15, LSL17, LSL18, Mil16, Zha16]. ARIMA [OLRPdCJ19]. arising [ALS19, BKMO19, BKMO20, Ceb17, ID18, KDB17, LLSZ17, MFWK18, RMA19, SR16a, Sch18, yZLX15]. arithmetic [HOM17]. ARKN [LSW19]. Arnoldi [AHJ17, GW17, HLLY19, JL15, RS18]. Arrival [GMSR18, GMSR21]. Arrow [CHS17]. art [LSS19]. artificial [AB18a, NGGFM17, TDA18]. Asian [HOM17, MZ16, Sho18, dSBV16]. Askey [Luo18]. Aspects [CBP18, TAT18].

Assessing [BBM15]. Assessment [BG15, WL17c, Wu18, WH19]. asset [FEC18, PX17]. asset-liability [PX17]. assimilation [JBFS15, NPR18, RS16]. assistance [MHY17, MHY21]. assisted [KLS19, NRV18]. associated [BHN15b, CL18b, CGPR16, DXML19, HMH18, Hua19, KSR15, LRT17, MDT16, Mej17, MVPN15]. associative [LSS19]. assumption [BS17c, Rom18]. assumptions [LLY19]. astrophysics [RMA19]. asymmetric [ALMLM17, HMKCR19, Pap16a]. Asymptotes [BPD15a]. Asymptotic [FL16, GQM16, GWY17, KS18, KDB17, Lan18, LL15, MMMMB16, MMZ19, OLPT16, PHCG16, RP17, ÁAM15, BP17b, DIJ16, LCX18, MFX15, MCLW16, MYM16, NSYX17, PA19, Prz15, YZJC15]. Asymptotical [Ogu17]. Asymptotically [AN17b]. Asymptotics [FF15, GZH19]. atomization [SRRJ16]. attempt [DBdLdM^{+}19]. attraction [HH18a]. Attractivity [AL16, FL23, LNP17]. Attractors [GN18]. Attribute [CMRP17]. Augmented [PKMM19, FLZ19, GxLlW15, HWH19b, Mia18, XYZ15]. August [WWST18]. automatic [SV19]. automorphisms [PDB15]. automotive [DMCC^{+}18, GCTGC16]. autonomous [BBCP16, BBC^{+}18, CCJ18, CNQRR18, Das18, GLSW16]. autoregressive [FF15, HYF18, ZLPF19]. Auxiliary [EHV19, MRMA17]. availability [CCIL18]. AVaR [Ugu17]. average [CL18a, HWT15, Xie15, YZZL19, ZqZwZ15]. averaged [CH17a, DRS16, DRST19, JRT17]. averaging [BL17, CSW19, DS17b, GMCD15]. Avoiding [AMCR19, CR17]. aware [SZM^{+}18]. axis [CM18b]. axisymmetric [KH18, Zha19b]. Azzalini [DPN15].

B [AZ18, BDGO15, FWW^{+}18, GK18b, HH15, Kre19, KB16, LLG^{+}18, LM17b, LZ19, Maz18, SuSR16, US19, YW15a, ZTGZ18, ZD19]. B-spline [AZ18, FWW^{+}18, GK18b, Kre19, KB16, LLG^{+}18, LM17b, LZ19, Maz18, YW15a, ZTGZ18, ZD19]. B-Splines [HH15, BDGO15, SuSR16, US19]. B.C [GRVC15]. background [CNJM18, ZB18]. backscattering [BTKM15]. backup [Nke18a]. Backward [APS17b, CZ19a, CW15a, FP16b, HK07, HRO16, KH18, Liu17, MKKV19, Reu16a, SWZW15, Sun16, TGJ19, THL^{+}17, WWZ15]. Baja [TLL16]. Bakhvalov [BZ16a]. Bakhvalov-type [BZ16a]. balance [AL18, BP17b, SKBT16]. balanced [AHT18, CL19b, KD19, LSG18]. balancing [NC16]. Baleanu [DLY18]. ball [FWW^{+}18, NP15]. ballistic [KLM18]. BAM [AYSE19]. Banach [AMB19, ABEHV16, AM15b, BKLR15, CHVRT15, Fer15, HSS19, KGMS18, MP17a, WWH19, Yao16]. Band [Elo19, LC17a]. bandlimited [Che16b]. bands [ZZW18]. Bandwidth [HB16, Iid15]. bang [BAO^{+}16]. bang-singular-bang [BAO^{+}16]. banks [UPD^{+}15]. bar [AMM19]. Barenblatt [GKBG17]. Barrier [HLL15, SCG18, BRA16, DYD19, GB18, KYY15, MB17, SM18]. Barycentric [LHPZ18, AHP19, CH19b, VS18]. Based [Lu18, ABG17, AL18, AH19, AGM19, AdHR17, APYW19, And17, ABP15, ACEJ18, AALZ15, AHKT15, BRW16, BM18, BBH^{+}17, BSLL18, BDA17, BD19, BGII15a, BGII15b, BIR17a, BIJM^{+}19, BOF^{+}16, BKLR16, BBG19b, BZ19b, CCY18, CFSS18a, CG17a, CCIL18, CDF^{+}16, CLZL18, CDZ16, CCT19, CH19a, CDKN14, CDKN15, CDHP17, CPL^{+}19a, DDG^{+}16, DF15, DP17b, DIP^{+}19, DAM15a, DGH15, DWS^{+}18, DLM16b, DGM16a, DY19, DS17b, Els15, FMC^{+}18, FP18, GLR16, GC15, GK19a, GX17, GGC15, Gun18, GUB15, Haj19, HM16, HS18a, HMMV17, HRS16, HWT15, HL18b, HSAA19, HR16, HZ19, HXC19, HG19, HOM17, HZ17b, HH19b, Iid15, ID18, IAMCI18, JC18, JMSL19, JL19b, JA17, KSDM16, KL17, KV15, KBL19, KK17, KB16, LBL15, LAVMA18, LQZ16, LWZZ16, LLCG16, LY18, LXD^{+}18, LSG18, LSH18a, LHLL18, LSDY19, LCZ18, LWC17, LSBL18a, LSL15]. based [LYG15, MTCC15, MDT16, MT16a, MD18, MP15, MB15a, MAS19, Mu19b, NCJ16, NS19, NN18a, NSM16, NGGFM17, NDH16, NC16, OH18, OD17, OG16, PKMM19, PD19, PG19, PBMJC18, PBJCM19, PDO19, QLZ19, QWZ18, RNPK15, RDRM15, RM17, RCPA15, RF18, RRS19, Rót15, SE16, Sar19, SV19, SPCSA16, SZZ^{+}16, SQRE16, SP17, SJL16b, SM15a, SV16a, SW17c, SX18, TXFH18, THC^{+}18, VFG18, WCSL15, WSL17, WWM17, WCW18, WZS19, Wu16a, WL16, Wu18, WH19, pWmGkY^{+}19, Xie15, XHZ17, XF16, Xu17, YY16, YL18a, YHZ18, YZZL19, YN18, YC19, YP17, YLZG18, ZNG15, ZW16, ZG17, ZS18b, ZJQW18, ZLPF19, ZL15b, ZHL15, ZZ15, vSV15, vV16, BKK18, DMP19, DVJ18, LFKG16, NGGFM16]. bases [CDF^{+}16, Che18, GS15a, JSC18, Maz18, Pap16b]. basic [PST18]. basins [HH18a]. Basis [NRV18, AD18, BM17a, BDSB16, CH18, CL15, GK18b, HW18a, Jeb19, JL16b, KV18a, KDB18, Lu15b, LX17, MAK19, PPM16, TLFC18, UPD^{+}15, WWM17, YBC^{+}16, Zha17a, ZHL15]. basket [ST16b, SX18]. batch [WXJ19]. bathtub [DNK19]. bathtub-shaped [DNK19]. batteries [FFKP17, VMM18]. Baxter [ZCD17]. Baxter-like [ZCD17]. Bay [GRVC15]. Bayes [TK15, WS16]. Bayesian [ABB18, BMLR15, CG19, GMSRP16, GT17, Gun19, HYF18, JO17, LSZB17, NMKK19, SMTdF18, SM19b, ZNS18]. BCR [Haj19]. BDDC [KCX17, PSW18]. BDF [Jan17a]. BDF2 [CLC18a]. BDF2-FEM [CLC18a]. BE [ZCFW17]. beam [CLS18, RL18]. beams [AG16a, Mar19b, Sto18]. beams/rings [AG16a]. bed [MFK^{+}15, OSC15]. bee [NGGFM17]. behavior [ALS19, Ahn15, DQIM18, DdSL16, FJD17, LZ17b, MMMMB16, SAA18, YZJC15]. behaviors [You19]. Behaviour [AHS19, GMW18]. behind [GKN18]. being [SSCM18]. Bell [Orr19]. BEM [ADG19, AMP15, BS16, GM16a, ZLS^{+}15]. BEMs [CFSS18a]. benchmark [ZG17]. benchmarks [LW15]. benefit [ZJQW18]. benefits [FJD17]. Benjamin [DAM15b, Güc17]. Berenger [HJL18]. Bernoulli [Baz15, CCNQ^{+}17, MLW19, Qi19, ROB17]. Bernstein [Beh16, BM18, BDSB16, BWS17, EDO19, GS15a, HS17b, JBT16, KL17, Lu15b, LX17, ZHL15]. Bernstein-like [ZHL15]. Besov [ABH19]. Bessel [Ceb17, CY18b, CJV17, LZZL17, Xia19, XM16, ZuI17]. best [BIR17a, HMMV17, LLW17, TSMO17]. Bethe [GLZ18]. Betti [LDDF18]. between [BPD15b, CL18b, CJT17, CBMPP15, FWW^{+}18, HLV^{+}15, JC18, MSZ18, OSMZ18, PDB15, QL19, Tru19]. Bézier [Ahn19, BDSB16, CFJMVG16, GS15a, GW16, HWC16a, HG17, HFMM16, LA15a, LA15b, LZ18a, Lu15a, LX17, LJX17, SR15, SJ15, WZH19, YW15b, KL17, Hu16, LJ16]. Bezout [WGZ15]. Bezoutians [DTB17]. BFGS [And17, And18, YN18, YSW^{+}18]. Bhattacharya [MDGVR17, MDAP19]. Bi [RFO17, CD16b, CMTV17, OGR17a, LGB18, ZZG^{+}19]. Bi-center [RFO17]. bi-diagonal [ZZG^{+}19]. bi-object [CD16b]. bi-parametric [CMTV17, OGR17a, LGB18]. biangular [ZwKN15]. biased [GK19b, LSZ16]. BiCGCR2 [GHC16]. bicubic [ZDLD15]. bidiagonalization [BEJR17, WGLV16]. Bidomain [OS15, CLW17, SZ16]. bifurcation [SE18, You19]. big [HLLY19, Lu18]. biharmonic [CG17b, HH15]. bilevel [ESAE18]. bilinear [ABB17, ABG17, DLMZ19, KV18a, NSYX17]. bilinear-like [KV18a]. binary [DY19, OGR17b, ZCS^{+}19]. binary-tree [ZCS^{+}19]. binomial [COGD19, Ery16, WZS19, WS16]. biochemical [Ver16]. biodegradation [AAB18a]. bioheat [RdSLL16]. Biological [SS18d, JCC18, SR16b]. biomass [NGGL^{+}19]. biomolecular [DXZ18]. Biorthogonal [AJR15, AHL17, MT19]. Biot [CCX18, GLMT16, HRGZ17]. bipolar [CLM19]. biquadratic [HH15, ZDLD15]. Birational [NSW15, PDB15]. Birkhoff [CL15, DD16a]. birth [You19, vD15]. birth-death [vD15]. bistability [TLE19]. Bivariate [AGR^{+}15b, Guo18, GH17a, GKM18, GUB15, Har18, JZS16, LWZ19a, LLG^{+}18, QL18, QWZ18, RN17, SKBT16, TJA18, ZWL18]. Black [KV18b, JHK17, KDB17, KV16, KV17]. bladder [GMSRP16]. Blaschke [NT15]. Blind [DTB17, LLD18]. Bloch [LMM19, QLT^{+}17]. Block [BCMP16, HS17b, JM18, JL17, XHZ17, ZS18b, ATG18, AALZ15, AKM18, Beh16, BBC17, BJK19, Byl18, CLZL18, CLZS19, CLW17, DB15, FV17, GDCK19, GW19, HL15, KBG17, LZ18c, LZZ19, LCX18, MLJ16, MAS19, NE17, NZL17, RMVA17, SHCJ19, WS17, XMW^{+}15, yZLX15, ZZG^{+}19, ZZZ15, xZWlC15, XY19]. block-centered [LCX18]. block-pulse [MAS19, XY19]. block-Toeplitz [NZL17]. Block-tridiagonal [BCMP16]. block-wise [GDCK19]. blocks [NZL17, NR15]. blood [KAM16b]. blow [Mus17, TMS^{+}17, YZZ19]. blow-up [Mus17, TMS^{+}17, YZZ19]. blur [Lu16]. BMU [JC18]. BNS [VL19]. Board [Ano18b, Ano18c, Ano18d, Ano18e, Ano18a, Ano18f, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18g, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r]. bodies [OGR19, VMR17]. body [AMP15, BHMT16, OCU17, OCU18, RL18, SKB15]. Boltzmann [COdSR16, VX19, XY16]. Bona [DAM15b, Güc17]. Bond [NM17, FNV17, Wu19]. bond-pricing [Wu19]. Bone [JC18, BCS17, CCD16, DMAS^{+}16, JC19]. Boolean [ADMV19, LZA19, LZL^{+}18b]. bordered [JSL15, MT17a]. Borsa [DI15]. Bose [Bad18]. both [WD19]. bound [Akh19, HZ17c, LXL15, LXL16, LSLZ17, Miy15, NTPVNTDT15, Tru19, WWZ15, WGZ15]. boundaries [CLZW19, LRKV16, LZ18a, LLZ18, XDW18]. Boundary [BKMS18a, BGS19, ZM15a, ZHC19, AAA15, ABH19, ALS19, AS18, AB18a, AL19a, ABL^{+}19, AK18, AZ18, AT17, AMP18, AMCR19, AR19, AD18, AC17b, AP18, BR15a, BBQ18, BKLR15, Bog16, BySZ19, CHX18, CHR^{+}15, CT17a, CS18b, CV19, CGSS17, CN16, DI19, DWXO15, DGM16a, DYZ18, DM16b, DD18b, Dzh16, Dzh18, EJJ15, FLZ19, FGF19, GHB17, GDQ17, GD17, GS15b, GS15c, GHZ^{+}15, GZZZ16, HHLS18, HA19, HZ17a, IN16, JAN17b, JLLZ16, JLCZ17, KM18, KCK18, KDW17a, KS16, KKK16, KtTBK18, LD19a, LD19b, LBL15, LSZ16, LC17a, LW17a, LLSZ17, LX16b, LM17b, MD19, Mat18, NXLX18, PW16, PRP19, PTV17, PCJ17, PCJ18a, PCJ18b, PCSR19, QM15, QDBG18, RC19, RVC19, RW16, RT18a, RMK19, SAA18, SDH17, SYZ19, SGG19, SS19b, Slo15, SLH15, SVL19, Sta18, SI16, TV16b]. boundary [VCL^{+}19, Vis15, WXW15, WPA^{+}18, WZS18a, WW16b, vtWvdHvdVV16, Xia17, YW15a, ZL15a, ZH16, ZLZ17, vV16, BKMS18b]. boundary-value [DM16b, KtTBK18]. Bounded [AAC17, AKHSU18, BDFM19, MDG17b, MDG17a, MDGVR17, TSMO17]. Boundedness [GDM16, MSH15, Wan17b].

Bounds [LDDF18, BLMS19, CJP17, DS15b, GWLL18, GYS17, GAB17, IN16, KVO18, LLV18a, Mir16, TBS18, Vej18]. Boussinesq [OZ17]. box [GW16, RKB16, XY16]. BPX [OS15]. BQN [HLBP15]. branch [Kow17]. branched [BFMP19]. Bratu [JM18, TBR16, TP19]. Bratu-type [TP19]. breakdowns [SB19]. Brezzi [LSD15]. brick [MLSK16]. bridge [EM15, SM16]. Brinkman [KSW19, AVB17, CJCDLL19, ES18, GLS15, HWZ18a, KLM17, LT18, WZZ16]. Brownian [Lev18, Sho18]. Broyden [Fan15]. Broyden-like [Fan15]. BSDEs [DL16b]. bubble [ZNG15]. bubble-type [ZNG15]. Buchberger [BW17]. Buckley [Che19a]. Buckling [Sto18, MM15b]. Building [NR15, SS18b]. bulk [Rät15]. bulk-surface [Rät15]. bundle [Gru16]. Burgers’ [CZ19b, Sey18, AHL17, BK19c, DAM15b, GHB17, MDG17a, MDGVR17, MS18, RCBG19, SEÖ16]. buried [BTKM15, YL18b]. burst [VOB16]. burst-correcting [VOB16]. business [HEA18]. BVP [FM18b]. by-claims [FL16].

cables [LMS17]. CAGD [KL17]. Cahn [Che19b, CFWW19, CSE^{+}19, DF15, JK18, LHY19, Son16, SJL16a, WCW19, XZ19, YZH18]. Calculating [HLLS18, QDBG18]. calculation [ML18, OT17]. calculations [CDH^{+}17, LC17a]. calculus [Abd15, AM19, DJAMJ17, FLN16, IAAW18, Uda18]. Calderon [YOK15]. calibration [GRVC15, GB18, MRMA15, MRMA17]. California [TLL16]. call [CEJ16]. CAM [HWXC19b]. Camassa [YFS18]. Can [DMS19, DD16c]. cancer [Mer19, SN19]. cancerous [KCDACC17]. canonical [CGMGV15, CSXY19, DLM16b]. Cantor [Man15]. capacity [ASV17, BBG^{+}19a]. capillarity [CM19a]. capturing [Pol16]. Caputo [CHX18, FJLB18, GS15b, LCL16, SF18, YMGA19]. Caputo-type [LCL16]. car [BM19b]. carbonation [PCJ18a, PCJ18b]. carcinoma [GMSRP16]. cardiac [AdQdS16, COdSR16]. cardiovascular [FNdO15]. Carlo [ABT17, BH17b, CFMÖ19, Dan17, God16, HWZ18a, HWH19a, Kaw17, MOY18, Nak15, SD15, SSD16, SMTdF18, SM19b, SX18, XW18]. Cartesian [ML16]. cascadic [SSW19, SWH16]. case [AHKT15, COGD19, CCNQ^{+}16, CCJ18, GO15, IN16, MOY18, PN16, dR16, dSBV16]. case-study [CCNQ^{+}16]. cases [CFG^{+}19, DLM16a]. cast [NGGA^{+}18]. Casteljau [SJ15, KMLB17]. Casteljau-type [SJ15]. Cauchy [AB18a, BLM16, CNQR^{+}18, DJC15, HS18b, HX15, HMH18, KR17, KW16, LZZL17, LMA18, Men17, TN19, YCX19]. cavitation [ML19a, SJRR17, SL18]. cavity [LW15]. Cayley [BCT16]. CBD [GD16]. CDO [OG16]. celestial [OGR19]. cell [AALMLM19, ALMLM17, SN19, YGC19, ZH16, ZBL^{+}16]. cells [JC18]. Cellular [ALS19]. censored [FPGK19, Gun18, HG19, Wan18e, Wu18, WH19, YZPL18, ZSW16, ZSZ17, ZNS18]. censoring [ABB18, BSLL18, BSLL19, NAK17, Wan15b, WS16, WL17c, WLC^{+}18, ZG19a]. center [EHVM18, RFO17]. centered [KD17, LCX18]. centerline [NGGA^{+}18]. Centers [BL17, GLV15, AIL17, GV17, IL15, LI15]. Central [GS15b, LPW^{+}19, CL19b, QCSG18]. centralities [FR18a]. centrality [CFG^{+}19]. centroidal [WKXT15]. Certain [AQCJ17, DM16b, KR18, NT15]. Cesàro [DMMSK17]. CEV [ALS18, JYK16, LRZ15, Mel19, Sta19, WRZ18, YL18d, YL19b, ZYY19]. CFEM [ZCH18]. CGRS [GC18]. Chaikin [DY19]. Chaikin-based [DY19]. chain [CH19a, CNQRR17, MT19]. chain-based [CH19a]. chains [HLLS18]. Challenges [YOK15, RS19]. chance [OAN17]. change [LC16, Mit15b, SVV18, WZS19, WAGVRB19]. change-point [LC16]. change-points [LC16]. changeable [SYT16]. changed [Sho18]. changes [NCQ^{+}18]. channel [EKKT17, Liu16, WS19]. channels [ABL^{+}19, Xin16]. chaos [LHM16, Luo18, OKGB19]. chaotification [ZZ18]. Characteristic [DLMZ19, GDC15b, GC17, JLCZ17, YB17, YSL^{+}18, ZYZ18]. characteristics [FW18, LD19a, LD19b, MT16b]. characterization [APS15a, APS15b, KSR15]. characterizations [La19]. Characterizing [BPD15b]. charge [BIJM^{+}19]. cheapest [KYS16]. Chebyshev [ABK19, AGM15, CRS^{+}18, CSTU17, DDM15, DRST19, GLR16, GOGM17, HXX16, HB16, LL16b, LSH^{+}19, MA16, MD18, NSM16, NT15, NDH16, NH19, PD17, YJBM19, ZM15b, dR16]. Chebyshevian [Maz18]. Chemical [BCMP16, CSE^{+}19, LK18, PCJ18b, RKV19, SR16a]. chemistry [LK18]. chemotaxis [AA17, Mer19, RKS16, ZZLY16]. chemotaxis-haptotaxis [Mer19]. chemotherapy [KCDACC17]. Chen [Wu18]. Cherednik [Mej17]. chikungunya [VZDV18]. China [ML17b, MLW19, WWST18]. choice [HHS^{+}15, LF19, PB15, SWZW15, VAR15, YL18a]. cholera [LPST17]. Choosing [CGT18]. Choquet [Nak15]. CIR [AHT18, NM17, Ruj16, Sta19, Wu19]. CIR/CEV [Sta19]. Circle [Ahn19, BD16, BCD15, BCA18, BCCB15, CBMO15]. circles [LAHB15]. circuits [Sch18]. Circulant [Hon19]. Circular [Zag18, Zyg17, AH18, Bej16, CKKJ15, FdGKZ17, VŽ19]. cited [DLY18]. CK [LSXL19]. claim [EG17, FY18, GZH19, RVZ15]. claims [BRA16, CLPZ17, DT18, FL16, NY19, RVZ15]. Clarke [HND18]. Class [FMC^{+}18, ABG17, AFBJGJ17, AGR15a, AG16b, AH15, BLXD18, BS18, BHN15a, BAO^{+}16, BG17, CFR18, CDRY18, CN16, DVA19, FCJ16, FR18b, GH16a, GMG19, HXX16, HK07, IÇ18, JMSL19, JGAFAL19, KK19a, KCK18, KS17a, LD19a, LD19b, LY18, LH18, LDG19, LDL^{+}19, LSH^{+}19, LN17, LNP17, MFX15, MT18, Men17, Mil15, Mok15, NBAK19, Nat19, NH19, NN18b, OM17a, PHD17, Ple17, Ros17, SHW19, SJ17, SVAKGR19a, SW17b, SLW17, TGJ19, VDT19, WNML16, WZL17, WL19a, WZWJ19, WL17a, XY16, YP18, YZZL19, Yao18, YJBM19, ZS18a, ZSL15, ZSL18, ZWT19, ZZZ15, ZM16b, ZZ15]. Classes [Mel19, Baz15, HZ18a, LFCW19, LZ16b, LDZ18, LPT17, Meh19, RFO17, RN17, Yao18, ZWX18]. classic [OCU17]. Classical [ABB18, CRGB17, RCBG17, BDGO15, CLPZ17, MDG17a, RVA15, RCBG19, dR16]. classification [BD19, CFJMVG16, LC16, RDRM15, SBP17, WCS17]. Classifier [TK15]. Clement [OV17]. Clenshaw [Wan18b]. climate [SPG^{+}16a]. close [OGR17b]. Closed [LaTG17, BN19, HZ18b, LWC17, RZ19b, Ruj16]. closed-form [HZ18b, LWC17, RZ19b, Ruj16]. closely [GC18]. closures [MS18]. clothoids [BF18]. cloud [CCZ^{+}18, SK18]. clouds [QDG^{+}15]. Clough [BI17, KB19]. cluster [GHA^{+}15]. Clustered [Hai17]. clustering [BCC^{+}18, EWR18, JWQ18, SBP17, VFN18]. clusters [DS15a]. CMAQ [SPG^{+}16b]. CMC [YZY17]. CMM [GHL18]. CMMSE [VACR15, VAMGR17, VAMDGR19]. CMRH [ATG18]. CMV [BDG15]. CMV-like [BDG15]. CNLS [RWW17]. co [GR16, MLV16]. co-creation [GR16]. co-infection [MLV16]. coalescence [BV16a]. coalition [MTMZ19]. coarse [lCCVV18, CDKN14, CDKN15, SPM15]. coarser [LW15]. Coast [TLL16]. coastlines [RTV15]. coated [GDC15b]. coaxial [LMS17]. code [GGK18]. Coded [MA19a, PGD^{+}18]. codes [VOB16]. coefficient [Abd18, Bay15, CL18a, FF15, GYL16, JJX17, LLL19a, LSXL19, LYG15, PCSR19, SJL16b, SZJ19, SYW19, WZS^{+}18b, YY16, ZLPF19]. coefficients [BS17b, BWS17, CTZK18, CZLV17, CLC18b, CL19d, DIJ17, FYY17, GC15, GCY18, KCX17, LMM19, LP15b, LWC17, MR19, MT17b, ML19b, MSV18, MHC16, Prz15, SW16a, SW16b, Sun19, WG18, Xu18, ZJ19, ZWH15b]. coercive [DMS19]. coexistence [SS18d]. Cohen [LJ15]. coherence [LC19]. coherent [BO16, FOK15]. Coincidence [BP17a, PAC15, dHKdHMM15]. cointegrated [CW15b]. cold [FOK15, NCPD19]. coli [BYWPA19]. collage [BKLR16, LCZ18]. collage-based [BKLR16]. Collective [DT18, Lu18]. Collocation [PSK16, SPK18, AKHSU18, AST19, AT17, CW16b, CZLV17, DP17a, DSOBK16, Elg19, FMR15, GAA16, HMT^{+}17, HR15, HXX16, HNT19, KPT15, LHPZ18, LLHO18, LFWC19, LL17, MMO15, MHL16, Moh19, Mok15, NN18a, Nem15, NH19, PSS17, PRP19, PTV17, Pis15, SR16b, SGG19, SLH15, TJB17, WWM17, WW19a, WCS18, YHZ18, Zak19, ZWGL18]. Colony [Xu17, NGGFM17]. color [PBJCM19]. coloring [FP16a]. colour [PKMM19, PBMJC18]. Column [HVV15]. combination [HHEG17, LL16a, Zha17b]. Combined [APS19a, MG15c, TAT18, BCS17, DD16a, HSWH15, Mit15b, NH19, XDW18, YSL^{+}18, ZZLY16, ZZZ^{+}17, ZMTT19, ZDF^{+}17]. combining [Elg19, HWH19b, JGTB15, JWQ18]. comment [FL23]. Comments [SF18, TGJ19]. commodity [GVHMR17, GVHMR18, ST16b]. common [BWS17, Gun19]. communication [KBT15]. commutative [GLMY18, LW19b]. commutativity [CRR19]. commutator [Tho16]. Comonotonic [FJD17]. Comp [GMSR21, MHY21]. Compact [CROC16, BK16, CRCM16, DP19c, JGGS16, JS19b, KK19c, LLL19a, MT17b, MS19a, WLW19, WSL19, Wu17, ZZW16]. compactly [HW18a]. companies [ZS17c]. Comparative [CN18, DSN17, KPS17, SR16a, dSBV16, CFSS18b, LLSM15]. Comparing [APS19b]. Comparison [BS16, Das15, HZ18a, JCH17, MSH15, RF18, Su18, BCC^{+}18, CJT17, EPC17, KVO18, NGGA^{+}18, RFLM16, SVW19]. Comparisons [YH18b, GGC15]. compatible [LA15a]. compensated [LSL18]. competing [Wan18e, WS16, ZSW16]. competition [CM18b]. competitive [PBS15, Wan18h]. complement [BO18, NZL17]. complementarity [AHM18, DLLB17, Fan15, GWLL18, HZ18a, HWCC15, HWT15, LY18, WYW15, WL16, YZZL19, ZqZwZ15, ZL15b]. complementary [HWH19b]. Complete [DD16a, EJJ15, Fuk15, PW15a, RMMRC15]. Completely [QL19]. completion [BDGIPG15]. Complex [ASKG18, NS18, AY19, APYW19, AKM18, BN19, BBB17c, CQW17, CM18a, CFG^{+}19, Elg19, GGK18, HWXC19a, HWXC19b, Kow17, LRRV18, LH18, LH19, NZ15, PM15, QL18, RWW17, SZZ^{+}16, SLQ16, TJSW19, WW16a, WZL17, WCW18, WL17b, YH18b, ZZZ15]. Complex-step [NS18]. Complex-Step-Finite-Difference [ASKG18]. Complexity [DGSM15, ACDT18, CT15, GM17b, YSL17]. compliance [FS15, HRS16, VMR17]. component [DS15a, DKLP15, FOK15, GVHMR18, HHKR15, Kav16, KS16, Vab18, YFS18]. component-by-component [DKLP15]. components [CJ18, LSBL18b, MM17, RIK16, SM16]. componentwise [DZ18]. composite [DCNY18, EL19, LWZ17, MCLW16, YZ15a]. composites [CJP17, GK18a]. Compound [Ery16, CLHMT15, GLP16, WH16, WZS19]. compounding [SBC16]. Compressed [CG17a, AALZ15, CL19a]. compressible [CFS19, CZ16, DBdLdM^{+}19, HK19, IDGH18, LRX16, SSG16, ZCH18]. Compression [BDG15]. compressive [MA19a]. Comput [ÁD15, AB16, Beh16, BP17a, BKMO20, Chr19, CDKN15, CPL^{+}19b, DNL16, GAA21, JSW18b, LGB18, RdSMO18, SW16a, TGJ19, ZPL18]. Computation [CPOG18, DS19, FFN^{+}17, KK19a, KL15, Xie15, ADG17, BBG^{+}19a, BWS17, CDDD19, CDR18, DIS^{+}18, DI19, Fuk15, JR15, KF16, LDZ18, ML19a, MM19, MS19b, Miy17, Miy18, Miy19, Mol15, Pap16a, PP15, PN16, PKR18, QWZ16, SS16a, SIAJ^{+}19, SS16b, SM17, SL18, TDA18, XM16]. Computational [AG16a, BdSFL16, BK19d, DNY17, DMCC^{+}18, KAM16b, SDG15, TJCC17, WWST18, WL17c, AA18, ALMLM16, BG17, BLV19, CNQRR17, DCNY18, DGN19, GM17b, JTCC18, KS17a, KV18b, MH16b, SK15, TAT18, TJCC16, VFG18, VAMGR17, WLC^{+}18, YGS18, VACR15]. Computationally [PS18]. computations [ÁD14, ÁD15, ACEJ18, OV17, Tan17, ZDL16]. compute [ACDT15, BCMP16, GGBTP17, MG15c]. computed [MA19a]. computer [GM16a, KLS19, MHY17, MHY21, PN16]. Computing [ABP15, CBP18, CCNQ^{+}17, CNQRR18, DD18a, EG17, KW16, Mia18, NS17, NR19, PCJ18a, SS18b, SSH16, AIS^{+}17b, BCJV18, BCVV18, CQW17, CM18b, CHYZ19, CV17, Chr18, Chr19, CEJV16, CNQR^{+}18, DIP^{+}19, DB15, Elo19, GD18, GXZ15, GW17, GJSC18, GM16b, HS17a, HRS16, HAH17, Kam17, KLL18, LXL15, ML17a, PCJ18b, RSID16, Sad19, SNY19, WO19, WHS17, Xie19]. concentration [TNTL18]. concentrations [NGGFM17, SRA^{+}15]. Concept [CYY^{+}16, Ahn15, CMRP17, FZW19]. Concomitants [DG16b]. concrete [PCJ18a, PCJ18b]. condensates [Bad18]. condensers [BBG^{+}19a]. Condition [Dia18, AM17a, AR19, BySZ19, CW15a, DWQ16, DZ18, DLHY15, DD18b, EHVM18, Fer15, GS15c, HMS19, HZ17a, JZ18, KB17, KM18, KKV19, KS17c, LT18, Li18, SLH15, WP16a, YW16]. Conditional [CEJ18, XW18, GSM18, GM17a, Hag15, Ruj16, Ugu18, ZLPF19]. conditioned [KR18, ML18]. Conditioning [AK18]. conditions [AB18a, AL19a, ABL^{+}19, AK18, AMP18, AK17a, AH16c, AEHVM18, AG19, BKMS18a, BKMS18b, BHN15a, CJCDLL15, CJVW15, CXW19, DGM16a, Dos19, HMD19, IN16, JLCZ17, KDW17a, KS16, KGMS18, LD19a, LD19b, LX16b, Mol15, PW16, QM15, RMMRC15, SGG19, SVL19, VCL^{+}19, WPA^{+}18, WZS18a, WZS19, WW16b, Xu18, Zho16]. conduction [MCLW16, QDBG18, VDS15]. conductive [GRS19]. conductivity [CMHHS17, KV19b, dAFvdMR19]. cone [BDF15, Fer15, HWCC15, NS19, NTPVNTDT15]. Conference [DNY17, WWST18, CPRS19, BLLM^{+}15]. confidence [YS19, ZZW18]. configuration [CCIL18]. configurations [ADMV19, GH16a, GH19a, Uda18]. conformable [Abd15, Chu15, KK19b]. conformal [DS19, Moo17]. conforming [OZ17]. congruence [OSMZ18]. Conic [Chr19, Chr18, DLMZ19, HG17, Hu16, NR15]. conjugacy [DLHY15, GKM18]. Conjugate [GW19, JL17, And16, And17, DLHY15, DZM16, Fat16, GC18, GHC16, JJ19, LSDY19, LH19, LS18a, Miy19, Ou18, VL19, YN18, YZ15c, YWY19, ZEBM17]. connected [BD16, FJ19, LZZL17, XZXQ18]. connecting [BBB17a]. connection [AGR^{+}15b, CBMPP15, MP15, PDB15]. conquer [LRL^{+}18]. Conservation [BMG^{+}19, RCBG19, BSS15, BM19a, CRGB17, HWC16b, HW18b, JN18, LBL15, LSXL19, LV16b]. Conservative [BAK18, JSW17, JSW18b, CW16a, CZ17b, CFS19, CVW17, DGM19, GdLT17, MM15c, PG16, WHZ16, WD18, WLW19, WD19, WSL19, WW19a, YB17, ZWW^{+}18, ZKRL15]. conservative/dissipative [MM15c]. Conserving [LRKV16, BGS19, FS19b, WCW15b]. considering [BE19]. consistence [Col16]. consistency [GWT18, Tho16]. Consistent [BDGIPG15, BCC^{+}18, CW15b, CSE^{+}19, GSVSM16, Han19, Kad17, KW18, LRZ15, MDG17a, MPH18, OD17, OGR17b, OGR19]. consisting [BE15]. consolidation [CCX18, HRGZ17]. constant [CH19b, GZH19, GB18, Men17, MS19b, Mil15, Mil16, SFH^{+}16, TSMO17, Wan18c, Wan18d, YFwC19, Zha16, vSV15]. constant-stress [Wan18c, Wan18d]. constants [BYW18]. constitutive [AM16, KMD^{+}16]. constrained [AFN17, BBH^{+}17, BCFN16, BKLR16, BySZ19, CD16b, DDM15, DP17b, EWR18, GM17a, Haj19, HHS^{+}15, HMMV17, HYZ16, HY19b, LLW17, LW19b, MHAA15, NS19, NRV18, PLP17, RKB16, SV19, SCP19, XCG19, Yua15, ZXQH17, ZL17a, ZZW18]. Constraint [CYY^{+}16, BLXD18, JS19a, LMUZ18, OAN17, YZW15, Zho16]. constraints [BE19, Dia18, FZ15, GW16, HZ17c, Iid15, JCC15, JMN^{+}16, KCX17, MYG18, PX17, WP17, XDZZ19, ZqZwZ15, ZG17, ZX16]. construct [MRMA15, MRMA17, OGR17b, OGR19]. Constructing [CEJ16, EK19, GKN18, Maz18, WZH19, DD17]. Construction [AFH16, Abd16, DGM19, KBL19, LDZ17, PP19, TYMG17, YW15a, BIR17a, DKLP15, KMRS16, KV18a, MHY17, MHY21, MII16, SS18b, XMW^{+}15, XD17, ZHL15]. Consumer [ÁRS16]. consumption [ML17b, MYG18, TF18, WS18]. consumption-investment [MYG18]. Contact [KSD17, ACCF19, BS16, BOF^{+}16, DKP^{+}19, FS15, Gor16, HH19a, HMS19, HR16, SKB15, VMR17, XHH^{+}19]. contaminant [LYS19]. content [NGGFM17]. context [AMRM17]. Continuation [Elg19, KLL18, BC16, BMM17, BZ19b, XC18]. continue [AV16]. continuing [JGGS16]. continuity [LZ18a]. Continuous [HRS16, HLZ19, SCP19, BFMP19, BS17c, CGMGV15, CY18a, DGM19, DGN19, FR18a, GC18, GHV19, HP17, IAMCI18, KK17, LL15, LI15, LSL17, LSL18, MFK^{+}15, MGMDGVR17, NGGA^{+}18, OCU17, OCU18, Yao16, ZCS^{+}19, ZWH15b, dFLM18]. continuous-discrete [KK17]. Continuous-time [SCP19, HP17, KK17]. continuously [Xu18]. Continuum [VCL^{+}19, LWV18]. contour [Bej16, GG17a, VK16]. contraction [DTKP17]. contractions [SHS^{+}15]. contractive [KS17c, RMMRC15]. contracts [ZY19a]. contrast [AEKGP19, AdQdS16, GK18a, PHCG16, RP15, YOK15]. Control [LZL^{+}18b, PB15, Rót15, ABH19, ABN^{+}18, APYW19, AFN17, AL19c, BYW18, BGMS16, BYWPA19, BGNGR^{+}17, BySZ19, CHS^{+}18, CG17b, CSB18, El 17, ERPD18, FZW19, FAVGC^{+}15, GCY18, HNT19, HC15, HYZ16, JCC18, JAN17b, JZ18, KCDACC17, LE18, La19, LPST17, LLX17, LX16b, LDW17, LW19b, MO18a, MLV16, Mat18, MO18b, MS15a, MSY16, MSY17, OKGB19, Ple17, Prz19, RRW17, Ros19, SYZ19, SZZ^{+}16, SSCM18, Sun16, hShX17, TF18, Ugu18, WZZ18, YT16, ZYXT16, ZXQH17, ZWY16, Zho16, ZC16b]. controllability [GT19, LDW18]. Controlled [Ugu17]. controller [ZWWW19]. controllers [DMM18]. controls [HND18, WKBB17]. convection [ALC19, AW19b, BVZ17, BK15b, BK19d, CFX17, CHS^{+}18, CJ19, GLL19, GS15b, GO17, HSWH15, Jeo15, Jin16, LC17b, Liu18, MPH18, NSV18, PDO19, PCSR19, RR16, SYW19, WZS18a, WAGVRB19, ZCZM17, ZP19, ZL17c, iW16]. convection-diffusion [BK19d, CJ19, GO17, Liu18, MPH18, NSV18, ZCZM17]. convection-dominated [HSWH15, Jin16]. convective [KtTBK18]. Convergence [AHT18, AMB19, BCD15, BHRR18, CGPR17, DMMSK17, DIJ16, GM19, HHLS18, HLM18, HMT15, JC15, LZ17a, LZ17b, LHZ15, LSL17, LSL18, MJ18a, MMO15, Mao16, MLSK16, Mil15, MS17e, PI19, SYZ19, WZ17, YAA19, YJBM19, yZ17, ZWX18, iW16, Ais17a, ÁD14, ÁD15, AAB^{+}18b, ACDT18, AL19b, AR15c, AM15b, AM15a, AGM15, ACMT17b, AG19, AALZ15, AKM18, BM17a, BM18, CV17, CT15, CHVRT15, CJSCT18, DLW16b, DTKP17, DLM17, DIJ17, EL19, EHV19, Fer15, GDM16, GMdV19, GAA16, Guo17, HVMR19, HK07, HH18d, JYK16, KB17, KHLY17, KAS18, KGMS18, LP15a, LX18, LW19a, LJ15, LY18, LZW18, LWZ19a, LYS19, LM15, LLY19, LN19, MG15a, MA16, MS17a, MD18, MBM17, MPM^{+}17, ML19b, Mok15, NN17, NSV18, Ou18, Pis15, RP16b, RC19, RW16, SPK18, SNY19, SH18, SH19]. convergence [SBCJ18, TGJ19, Tan18, WW16a, Wan18b, XY18a, XY18b, YSG18, Yao16, YH18a, dlCYC15, YSW^{+}18, YWY19, ZQJY18, ZSL18, ZCC15, ZC16a, ZJ17b]. Convergent [CLZL18, AW19b, AK17c, Bay15, BTKM15, Che16a, EC18, JGTB15, LLW18, MDG17a, MR19, SJRR17, STS15, ZL18a]. convergent-divergent [SJRR17]. converging [RMA19]. Convex [HLMS19, Iid15, AAKT19, AHS19, BCFN16, CLZL18, CLZS19, CK16, CPL^{+}18, CPL^{+}19b, CPL^{+}19a, HAH17, KKP^{+}19, LV16a, LXL16, LDL^{+}19, MA19b, MTMZ19, NS19, Nie15, SCP19, TSMO17, WS17, WWH19, ZML16]. Convexity [Han15, GS18a, GS19, KElY16, MLS15]. Convexity-preserving [Han15]. convolution [AL16, BL16, DVS15, FL23, LA15a, SWZW15]. convolutions [GWT18, JB18]. Conway [GH17b]. cooperative [CZ15]. coordinate [BAO^{+}16, EK19]. coordinates [AHS19, MCLW16, VS18, pWmGkY^{+}19, ZwKN15]. Copula [Xie15, BSLL18, Rom18, TJA18]. Copula-based [Xie15]. copulas [Ahn15, BKMO19, BKMO20, DGM16b, GSM19, JN19, LCA17, RN17]. Cordero [Wan18f]. Cordero-Torregrosa [Wan18f]. core [yZyMxG^{+}16]. corner [CL19a, RR18]. Cornu [CC15]. Corput [Xia19]. Corput-type [Xia19]. Corrected [TBAH16, ALC19, AA17]. correcting [VOB16]. correction [CLZL18, EP18, Mia18, QM15, WXJ19]. Corrective [GLLS16]. Corrector [Elg19, PS15, YLL15]. correlated [KB17, Liu16, STZ16, WS19]. correlation [KV16, LWLL18, TEG15]. corresponding [PMN19, SvG15]. Corrigendum [ÁD15, AB16, BP17a, Chr19, CPL^{+}19b, GMSR21, LGB18, RdSMO18, SW16a, ZPL18]. corrosion [PCJ18b]. corrugated [LMS17]. cosine [CYY15b, DIP^{+}19, HOM17, SIAJ^{+}19]. cost [ALMLM16, BE19, CD15, SAC17, SS18b, WCS17]. costs [TF18, Ugu17, dFG17]. Coulomb [VMR17]. count [AFBJGJ17, QLZ19]. countermonotonicity [LCA17]. counting [BDAG15]. coupled [AMZ19, BM19b, BK19c, CGPR17, Con18, DAAF16, FZW19, FNdO15, HRO16, Kad17, LLCG16, LGLC18, LGYC19, Lin16, LSH^{+}18b, LWC17, LC17b, LCX18, RC19, SSD16, SW17d, VG17, XDY16]. Coupling [LWLL18, AMP15, BHN15a, BV16c, ERR18, GC18, HS18c, LWW17, MPM^{+}17]. Covariance [GYLQ16, Mar19a]. covariates [GMSRP16, ZZW18]. covolume [ZCL16]. CP [TPC17]. CPR [WLLC18]. CPR-type [WLLC18]. crack [LWW17]. Craig [iW16]. Crank [GSRJ18, JKL^{+}15, LMW18, LGYC19, Moh19, RZ19a, SW17d, ZZW16]. creation [GR16]. credit [MX16, WZBM18]. crisis [Nke18a]. criteria [HM19b, IABIPG17, IAMCI18, Ugu17]. criterion [BFLP15, LRZ15]. critical [NCF^{+}19]. criticality [SOMF19]. cross [Arm18, Luo18, MRVV16, SFH^{+}16]. cross-diffusion [SFH^{+}16]. cross-grid [Arm18]. crossflow [LZRH19]. cryosurgery [KS15b]. crystal [Bot16]. crystallizer [MFK^{+}15]. crystallography [PGD^{+}18]. crystals [AK17b, LC17a]. CS [NTPVNTDT15]. CT [HLG^{+}17]. CTIT [OM17b]. cubature [Ber15, Pap16a, Pap16b]. cubatures [NP15]. Cubed [Cro15, PC18b]. Cubed-Sphere [Cro15]. Cubic [GS19, HG18, AHBC15, ÁAM15, AIL17, Bal17, BC16, Han15, HZ17c, KB19, LI15, MG15b]. Cucker [GMW18]. CUDA [GDC15a]. cumulative [LC19, PT17b]. curl [CL19d]. currency [XY15, ZYL18]. Current [VAMGR17, AL19c]. Curtis [Wan18b]. curvature [LX16c, WZH19]. Curve [LXD^{+}18, ZHL15, BRW16, CYY^{+}16, LA15a, LA15b, SW16a, SW16b, TXFH18, YB17]. curved [AG16a, CDF^{+}16, LLZ18, Sto18, VRBS19]. curves [AH18, Ahn19, ACDT15, ALV19, ADD^{+}16, ATV15, BAAA^{+}15, BDSB16, BLV19, BPD15a, BPD15b, BZ16b, EL19, FW18, FGS19, FWW^{+}18, GS15a, GW16, HWC16a, HG17, Han19, Hu16, IANS15, JSC18, KL17, LA15a, LA15b, LFCW19, Lu15a, LJ16, LX16c, LX17, LJX17, LZ19, MPS19, MB15a, Nie15, PDB15, Rót15, SR15, SPD15, SJ15, VŽ19, WZH19, WGZ15, WZ18, YW15a, ZZ16, ZWY16]. curvilinear [MCLW16, VS15b]. cuts [Kow17]. CVaR [BBB17a, BBB17b, DLS17, GSM19]. CWENO [HW18b]. cyanobacteria [NGGFM17]. cyanotoxin [NGGFM17]. cycle [NGGFM16]. cycles [BL17, LP18, LI15, dFLM18]. Cyclic [BAO^{+}16, AALZ15, CTZK18, SHS^{+}15]. cylindrical [VG15].

D [AMM19, BBG19b, CGSS17, CJ19, CROC16, CZ16, DAM15b, DMAS^{+}16, DFGS16, EKKT17, GC15, GLL19, GRS19, GD17, Guo17, ID16, KV16, KV17, LW15, LW17b, MS19a, PT17a, QDBG18, RdSLL16, ST16a, SJ17, SW17d, TBAH16, TW19, VS15b, WD19, Wu17, XZN^{+}19, ZLS^{+}15, Zha19a, ZWZ^{+}18]. D-bar [AMM19]. D. [DLY18]. DAEs [JS17, Ran15, Sch18]. Dai [JJ19]. Dalian [WWST18]. damage [CNJM18, DJAMJ17]. damped [DCNY18, MG15b]. Damping [KTT16, BM19a, CRR16, EGK17]. damping/driving [BM19a]. Darbo [CT16]. Darcy [AB16, AFM19, ABB^{+}15, CGPR17, Chi17, ES18, ELRR17, ERR18, KLM17, KSW19, LLCG16, LGLC18, LGYC19, Ngo19b, ZH15]. Data [KElY16, TK15, YS19, AFBJGJ17, AdHR17, AG18, APYW19, BL18, BIR^{+}17b, BBQ18, BK16, BL16, CG17a, CG19, DD16a, FP18, GLdCJS17, GHL18, Gun18, GH17b, HHS^{+}15, HMKCR19, HG19, HZ17b, HLLY19, IANS15, JAN17b, JBFS15, KP15, LAHB15, LSZB17, Lu18, MTV17, MAK19, Mar18, MP15, NPR18, NGGFM16, Nie15, QWZ18, RVC19, RS16, Su16, VFN18, WSL17, Wan18e, WSG^{+}19, YZPL18, YC19, ZSW16, ZSZ17, ZNS18]. data-driven [CG17a, CG19]. database [PR16]. dated [CP15]. Daubechies [PM15]. Davey [Mus17]. Davidson [HMP15, JL15, Mia18]. DC [Mud19]. DDM [ZL17c]. death [vD15]. deblurring [DGH15, EW18, HH19b]. debt [Nke18a]. decay [FdOP17]. Decision [FLMS16, Ugu17, BDGIPG15]. decomposed [JGGS16, JS19b]. Decomposition [FMC^{+}18, AM17b, ALC19, APS17a, BV18, BW16, CSXY19, CT17b, DLM16b, DFS15, GHM18, GKQW17, GMCD15, GC17, GK18a, HA19, HWZ19, HVM18, HZ17b, KVO18, KK15, KNP^{+}17, LDK17, LD15, NAM18, OD17, RL18, SMS18, hShX17, TDA18, TPC17, Vab18, WCW18, ZY18]. decompositions [CRSV19]. deconvolution [DTB17, LLD18]. decoupled [AA17, ZHS18]. decoupling [Chi17, ZH15]. decreases [IGRL17]. decreasing [DKLP15]. deep [AG16a, LWLL18]. Default [JA17, JYP18, JSW18a]. defaultable [ZC19]. Defect [AHKT15, QM15]. Defect-based [AHKT15]. defect-correction [QM15]. deficient [GW19]. defined [BZ16b]. definite [AN17b, MK17, Miy19, WXZ19]. definiteness [CQW16, HQ18, ZW16]. definition [OCU17, OCU18, YMGA19]. definitions [SF18]. deflated [MLJ16, NF17, SHCJ19, xZWlC15]. deflating [FwC17]. Deflation [DVJ18, vSV15]. Deformation [HFMM16, HRS16]. deformities [DMAS^{+}16]. Degenerate [Mol15, BK19d, CKKJ15, Ple17]. Degree [SYT16, AM17a, Ahn19, ACMT17a, BL17, BDSB16, GS18a, Hu16, LHJR15, LX17, WWL15, ZG19c, dR16]. degrees [QL19]. Delannoy [QCSG18]. Delaunay [GLTW15, WKXT15]. Delay [HKZZ15, ALS18, AHT18, Akh19, AY19, ACAC17, AAC17, CPB19, CW15a, CWH18, DP17a, EC18, ERPD18, GZH19, GH16c, GQM16, HND18, HNT19, JY17, KEB19, KPW17, KAA16, Lan18, LXW19, LW19a, LZA19, LZY15, LZ16b, LMA18, MS15a, Mil15, OM17a, PHD17, SB16, SYD15, SYD17, Sta19, Ver16, Wan17b, WG18, WZ17, XH19, YAA19, ZSL15, Zha16, ZLF18, ZSL18, ZWH15a]. delay-claim [GZH19]. delay-dependent [CWH18]. Delay-partitioning [HKZZ15]. delayed [DT16, LRRV18, LZL^{+}18b, LDZ18, ML17b, Mar18, Mit15b, You19]. delays [AG18, APYW19, AYSE19, CCD16, CWH18, FE16, GYS17, GH16c, JCC17, JCC18, LJ15, MHL16, Nat19, yZ17, ZWWW19]. delivery [FNdO15, RFLM16]. Delphi [Lu18]. Delta [KV18b]. demand [ASV17, BLH^{+}17, EPC17, WS18]. denoising [DGH15, HS18a, MDH19]. denominator [JR15]. dense [GK18a]. densities [ÁAM15, DSA16]. density [CRG19, CNQRR18, GWT18, GK19b, NCJ16, NL17, You19]. dental [BOF^{+}16]. dependence [Ahn15, BS17c, CLHMT15, GZH19, GMSR18, GMSR21]. dependences [Wu19]. dependent [Abd18, AL19b, AR19, BBCK19, BE15, BM19a, CL18a, CGPR17, CS18b, CRG19, CJ19, CWH18, CH17b, FYY17, FL16, FY18, GHM18, GTC16, GO15, Guo17, ICMPG16, JAN17b, JYK17, JN19, JLCZ17, KMD^{+}16, KV19b, LMM19, LXW19, LKN17, LAVMA18, LHJ16, LZA19, MDG17b, OZ17, Par15, Prz19, SJ17, Slo15, SYW19, VFFGV16, WW16b, XHH^{+}19, YZJC15, YT16, YZY17, You19, ZQJY18, ZL17c, dlHV16]. depletion [MV16c]. deposit [AO18]. deposition [ELRR17, ERR18]. depth [EM15]. derivative [Abd16, AEHVM18, BMM17, BGMM18, CHX18, Chu15, CSB18, EBJS19, EHVM18, Fan15, FJLB18, GVHMR18, GS15b, HG18, HVMR19, HZ17c, JLLZ16, Kam16a, KLM18, KMB^{+}18, LRKV16, LF18, MG17, Moh17, NHA17, NAH18, NS18, RS16, SM19a, SO16, YMGA19, ZJG^{+}15, ZHH^{+}19, iW16]. derivative-free [Fan15, HVMR19, HZ17c, LF18, RS16]. derivatives [ÁNPPCSA15, AEP15, BGMS16, BCD15, BZ19b, CLX17, CT15, CJT17, DA19, Goć15, GO15, KVO18, LN16, LCL16, LX15b, Mir16, OP19, SVAKGR19b, SSG16, Xie19, dSBV16]. deriving [dR16]. Descent [MHMP16, BAO^{+}16, DLHY15, EJJ16, GS18b, LZW18, LF18, SCP19]. described [BGMS16, PBP16]. describing [SR16b]. description [Rót15, SF18, UPD^{+}15]. descriptor [ABG17, CCRN19]. design [CCIL18, HEA18, LZL^{+}18b, PGD^{+}18, XCG19, YW15b, ZWWW19]. Designing [AO18, NR15]. designs [DGN19, MG15c]. Deslauriers [Vis19]. detailed [HYQ^{+}18]. Detecting [LC16, PW15a, SR15, YL18b, LXZ17, NCQ^{+}18]. detection [CBC^{+}19, CNJM18, GJ17, JMSL19, EP18, RCPA15, RRS19]. deteriorated [CDW15]. determinantal [PP18a]. determinants [MT15]. Determination [HHOT17, LMM19, Slo15, CRR19, HRR15, PRRY18, PP18b]. Determining [MRMA15, VL15]. Deterministic [El 17, CDRY18, MDVM17]. develop [LOAGBR15]. developable [LZ18a]. Developing [CY18a, DXZ18, YOK15, ZL18a]. development [RMVA17]. developments [LLT16]. deviation [LX16c]. devices [DMCC^{+}18]. Devising [SS16c]. dewetting [Zha19b]. DG [LRKV16, HT18, LHLL18, vSV15]. DGMRES [SS18a]. diagonal [BBC17, FM19b, SHC^{+}19, WLZ17, ZZG^{+}19]. diagonalizable [LXD^{+}18]. Diagonally [HHR16, GLLS16, WKBB17]. diagram [AGR15a]. dichotomy [XqYL^{+}15]. dielectric [PSS17]. diesel [SRRJ16, SJRR17]. Difference [ASKG18, Dos19, LV16b, AAM18, ÁNPPCSA15, ACAC17, BS18, CRCM16, CLX17, CJLT15, CFWW19, CEJV16, CEJ18, CROC16, CP16, Das18, DA19, DP19c, FGF19, GDCK19, GUB^{+}17, GS15b, HCL17, HHEG17, Jeb19, JK18, Kha17, KV16, KKK16, KAA16, LBL15, LSZ16, LCX18, LW19c, MDG17a, MT17b, MVdRE^{+}17, MVQDdR^{+}18, MHV17, MT18, MMZ19, QL19, SAA18, SBU^{+}17, Sha15, SYD17, TYS17, TBR16, UGG^{+}19, WHZ16, WLW19, WD19, WSL19, Wu16a, Wu17, Yos17, dSBV16]. difference-based [Wu16a]. differences [AHP19, KD17, ZD19]. Differencing [AAKW16, AK19, BK15c, BK19c]. different [BE15, GLdCJS17, NADK18, RF18, ZTGZ18]. differentiable [Yao16]. Differential [ADD^{+}16, BBB17a, KR16, AÖ17, ABN^{+}18, AH16a, ALS19, Akh19, AM16, AA18, ABP15, AJ17, AJY18, BM17a, BM17b, BJ16, BBCP16, BO18, BDA17, BGII15a, BGII15b, BDIR19a, BT17, BPP17, Beh16, BS17b, BL17, BGL17, BS19, BS15, BR15b, BCVV18, CCNQ^{+}16, CCNQ^{+}17, CCJ18, CPB19, CW15a, CHR^{+}15, CWH18, Con18, CJV17, CNQRR18, DBX19, DA19, DG16a, DFLM19, DIJ16, DT17, Dzh16, Dzh18, Elg19, Els15, EC18, EBJS19, FCJ16, FLJT18, FM18b, FLN16, GZD^{+}19, GUB^{+}17, GMG19, GV17, GN18, GLSW16, GLSW18, GS15c, GR19, GH16c, GQM16, Guo17, GLMY18, GL19, Guo18, HHR16, HMD19, HMT^{+}17, HH16, HCA^{+}18, HSAA19, HLM18, HB16, HIS16, IÇ18, JY17, JC15, JMS17, KB17, KP15, KS17a, KFT17, KRR18, KKS16, KDW17b, KDW17a, KMV16, KHLY17, KPT15, KEB19]. differential [KBG17, KAA16, KW18, KPS17, Lad15, Lan18, LX18, LXW19, LW19a, LW17a, LXD^{+}18, LCZ18, LL15, LZY15, LM15, LZ16b, LTY17, LDW17, LDW18, LLHO18, LLY19, LFWC19, LNP17, LSL17, LSL18, Luo18, MDT16, MDG17b, MG15a, MMS19, Mao15, MYM16, Mao16, Mär17, MS19a, MOPP18, Mil15, Mil16, MM15c, MC18, Moh19, Mok15, Mol15, MJS19, NE17, NSM16, NY19, NT15, NSW15, NDH16, NH19, OM17a, PRP19, Pis17, ROB17, RRM18, RANG15, Reu16a, RM19, SF18, SAA18, Sak17, Sar19, SM19a, SB18, SPK18, SAT15, SYD15, TJB17, TMS^{+}17, TYS17, TGSS17, WWM17, Wan17b, WG18, WKBB17, WZ17, WL17a, WLM16, XY19, XZXQ18, XH19, Yam17, Yan17, YL19a, YZZ19, YAA19, YJBM19, ZS18a, ZSL15, Zha16, ZT18, ZSL18, ZH18, ZJ17b, ZG19b, ZWH15a, ZWX18, dFLM18, dlHV16]. differential-algebraic [HMT^{+}17, Mär17]. differential-difference [SAA18]. differentially [DLW16a]. differentiation [BZ16b, Els15, LS18b, ZJG^{+}15, ZWGL18]. difficult [SHC^{+}19, TPC17]. diffuse [DFRR15, XD17]. Diffusion [AFdOdS15, ALS18, ALC19, APYW19, AMCR19, AK19, AW19b, AAKW16, AR19, BKMS18a, BS18, BVZ17, BK15a, BK15b, BB18, BPP17, BK15c, BK16, Bog19, BK19d, CCW17, CDP17, CD16a, CM16, CW16a, CLC16, CFX17, CHS^{+}18, CLC18b, CLA19, CRG19, CPZ19, CKB17, CJ18, CJ19, CEJ18, CNQR^{+}18, DP16, Dan17, DVA19, DSA15, Din16, DAAF16, FEC18, GHM18, GC15, GZH19, GZD^{+}19, GLL19, GDC15b, GVMR16, GVHMR17, GVMR19, GPHAPR15, GPHAPR17, GO15, GO17, GHZ^{+}15, GG15, GdLT17, HCL17, HSWH15, HK07, HYZ16, Jeo15, JSW18a, Jin16, KLM18, KV19b, KKW16, KK15, Lee18, LX15a, LCL16, Lin16, LC17b, Liu18, LF19, MX16, MDAP19, MVM19, MC19, MPH18, MVC19, ML16, MS15a, MKKV19, MFLL18, MR15, Mud19, NSV18, Nke18b, PHD17, PCJ17, Prz19, RP16a, RC19, Rät15, RR16, STZ16]. diffusion [SFH^{+}16, SR16c, SS16c, SLA19, SWZW15, SZJ19, SYW19, TGJ19, TYMG17, TC19, THL^{+}17, TNTL18, VFFGV16, VFGPG^{+}17, WWZ15, Wan18h, WRZ18, WZS18a, WO19, WZ16, XY15, YS19, ZCZM17, ZZG^{+}19, ZP19, ZKRL15, ZL17c, iW16, vSV15, BKMS18b]. diffusion-reaction [GPHAPR17]. diffusion-wave [BB18, CLC18b, DSA15, SLA19]. diffusions [AK16, YYL15]. diffusive [Bou18, MDG17b]. digital [God16, KGL19]. Dimension [AvG16, AdGvG19, BLMS19, KV18a, Bac18, BVZ17, CNJM18, Dan17, MRMA17, NSV18, SMTdF18, XW18]. dimensional [AN19, AK19, BM17b, BM18, BBQ18, BSS15, BG17, BEHL17, CDF^{+}16, CTZK18, CQW16, Che16a, CL19b, CPZ19, CDE18, CJ18, CRR19, CNQR^{+}18, DA19, EJJ15, FSNS16, FP18, FY18, GH19a, GRVC15, GLMT16, GDQ17, GKLS18, GWY17, HH16, HMDSR19, HMH18, JMSL19, KGES19, KV19b, KBL19, KNP^{+}17, LP15b, LMW18, LLL19a, LLG19, LHPZ18, MS17a, NE17, OD17, PSS17, STZ16, SE16, SAC17, SLA19, SM19b, uIuD18, SJL16b, SZ18, SO16, TH19, WW16a, XHZ17, Zha15b, ZZW16, ZJ17a, Zha19a, iW16]. dimensionality [YZ15b]. Dimensions [ZDLD15, HM16, HHKR15, RKS16, RCBG19, TKMT17, ZWL18]. DIMSIMs [IJ18]. Dirac [WW16b]. Direct [Che19b, LXZ18, MHAA15, Mol15, QWZ16, AEKGP19, FNV17, HM16, MFX15, WHWY18]. directed [DMR19]. Direction [GC15, Han19, LLL19a, BI17, BDIR19a, CLZS19, HH19b, TLL16]. Direction-consistent [Han19]. Directional [KGL19, GZD^{+}19]. directions [HSS19, LDZ17, VL19]. directly [HIS16]. Dirichlet [CDKN15, ACEJ18, CJCDLL15, CL19a, CG17b, CDKN14, GS15c, KM18, SGG19, WZS18a, WO19]. Dirichlet-to-Neumann [CDKN15, CDKN14, KM18, ACEJ18]. disc [CKKJ15]. discontinuities [MAF17]. discontinuity [HK19, PCSR19]. Discontinuous [CDE18, FLN16, SLS18, YZ16, ACLY19, AK17b, BJ16, Bac18, BBD16, CL19b, CFSS18b, DLM17, DYZ18, FS19b, Gia18, Goć15, GP16, GTX^{+}19, HL18b, HCSC16, KM18, KCK18, KCX17, LRKV16, LSG18, LGYC19, LLSM15, Lin16, LI15, MTS17, SA17, SVV18, SF16, TX19, VS15b, WZS18a, XW18, XH19, YP18, ZZLY16, ZM16a, ZZZ^{+}17, ZCS^{+}19, ZP19, ZC18, ZG19c, dFLM18]. discounted [YL18a]. discrepancy [BDF15, HPT19]. Discrete [AAM18, DNL15, EAKS18, FR18a, NW19, PCSR19, SB16, YZ15b, AY19, ÁNPPCSA15, AW19a, BEJR17, CYY15a, wCW15a, DJAMJ17, DD18a, DI19, El 17, EG17, FH16, FH18, GHL19, HRS16, HMDSR19, HRR15, HC15, HZ17a, HOM17, HJ17b, Jan17a, EP18, KPW17, KK17, LDZ17, LWL17, LC18, LWS16, Liu16, LZY17, MT16a, MR19, MO18a, MMMMB16, MP17a, Miy17, Miy19, MSZ18, Skr17, SM18, SO16, TB18, Tho16, WPA^{+}18, XMW^{+}15, YZY17, Yan17, YCW17, ZYXT16, ZXQH17, ZHY^{+}17, ZHH^{+}19, DNL16]. Discrete-fracture-model [EAKS18]. discrete-time [AY19, AW19a, wCW15a, LC18, LWS16, Liu16, MO18a, MP17a, Miy17, Miy19, YZY17, ZYXT16, ZXQH17, ZHH^{+}19]. discretely [HLL15]. discretisation [SSKV17]. discretised [Kaz18, KRR18]. Discretization [AK17c, BK15a, BAK18, CM19a, CH17b, FP16b, GDCK19, GRS19, GG15, JZS16, LX15a, LSP18, MM15c, MFWK18, SM19a, VS15a, Wan18a, XDY16, ZJG^{+}15]. discretizations [BVZ17, CW17, HB16, Moo17, PSW18, RNPK15]. discretize [LW19b]. discretize-then-optimize [LW19b]. discretized [CHS15, GS18b, MS19a, yZyMxG^{+}16]. discriminative [LWLL18]. disease [CNQRR17, FH18]. disjoint [FFP16]. disk [MS19b]. dispersal [ZLF18]. Dispersion [CDP19, NS16, TNTL18, Wu17]. dispersive [ÁD17, DLM17, JB18, LHLL18, MFWK18, NZ15]. displacement [AG16a, LRX16, YSL^{+}18, ZCH18, ZZZ^{+}17]. displacement-function [AG16a]. disposal [LYS19]. dissipative [AY19, MM15c, MFWK18, WW19a]. dissipative-dispersive [MFWK18].

Dissipativity [WL17a]. dissolution

[BBPR15]. distance

[BPD15b, DI19, RFLM16, UG18, WCS17]. distinct [AAKW16, IAAW18]. distinctive [BS17a]. distortion [WP17]. Distributed [Ple17, CRR16, CHS^{+}18, CWH18, LRL^{+}18, MR15, PHD17, ZC16b]. Distribution [AMRM17, ÁRS16, (Ba18, COGD19, CCC19, CY18a, CCRN19, CLPZ17, DNK19, DG16b, EDO19, Ery16, GH17b, GUB15, HMKCR19, HG19, JGAFAL19, KV19b, KKW16, LM15, MDVM17, Mar19a, MRMA15, MRMA17, NC19, NADK18, NZ16, OK16, RABB19, SKC^{+}19, VZDV18, Wan15b, Wan17a, Wan18c, Wan18h, WZS19, WS16, WL17c, WLC^{+}18, WZ18, Wu18, WH19, ZH16, ZG19a]. distribution-free [CCC19]. distributions [ARCA^{+}19, AR15b, BSLL19, CDN19, Col16, CAO^{+}19, DD18a, EG17, Hui16, MVPN15, NKD^{+}19, OAS19, SBC16, SRA^{+}15, TJA18, TLTJ18, Wan18e]. disturbances [ZWWW19]. div [KR18]. divergence [AY18, Cro15, YP18]. divergent [SJRR17]. divide [LRL^{+}18]. divide-and-conquer [LRL^{+}18]. divided [ZD19]. dividends [TB18, YCW17]. division [ALMLM17, SHW19]. divisor [BWS17]. Domain [RL18, ALC19, AKHSU18, AM15b, AMS17, AEHVM18, BV18, BIJM^{+}19, BBQ18, Bot16, CLC16, Cou19, DLM17, DB15, Elg19, GHM18, GMCD15, GC17, GK18a, HA19, JL16a, KK15, LHLL18, MCGNBD17, SM19a, SM15a, SVL19, SMS18, Sun19, TSMO17, TDA18, Vab18, VG15, VS15b, ZCS^{+}19, ZY18]. domains [BD16, CJCDLL15, CLV16, CVW17, CLVW18, DAM15b, Elg19, FJ19, FSNS16, FMP^{+}15, HZ17a, KV19a, QDY^{+}19, SLW17, VCL^{+}19, WHWY18, XDW18, ZWGL18]. dominance [OLPT16]. dominant [GLLS16, ZP19]. dominated [HSWH15, Jeo15, Jin16]. dominating [BGMS16]. double [And18, FM19a, HWXC19a, HWXC19b, JK19, Qi19, SM18, YS19, ZS17a]. Doubling [wCW15a]. Doubly [KW18, DL16b, RW16]. Douglas [HOS16]. down [DNK19]. downdating [SVGM17]. DPCG [DVJ18]. Dragomir [AAKT19]. dramatically [BMLR15]. drawing [ZF15]. Drazin [SS16b]. DRBEM [Mer19]. drift [AH15, BHN15a, HHR16, IÇ18, KHLY17]. drift-flux [BHN15a]. drift-implicit [HHR16]. driven [AK16, CG17a, CG19, LW15, LSL18, YCF18, YS19]. driving [BM19a]. drop [BV16a, BG15]. Drude [NZ15, SLS18]. drug [FNdO15]. DSF [MT19]. DSG3 [NTPVNTDT15]. DtN [GYX17]. Dual [SBCJ18, dMPF15, AMRM17, CRCM16, CDZ16, DXML19, KYY15, LWV18, PST18, RW15, TBY17]. dual-continuum [LWV18]. dual-phase-lagging [CRCM16]. duality [HR16, MLS15, ZNS18]. duality-based [HR16]. Dubuc [Vis19]. Dubuc-Deslauriers [Vis19]. due [AVGCMVM17]. Dunkl [GB16]. dwarfism [AALMLM19]. dyadic [ZS18c]. Dynamic [AMM19, CG19, CW15b, DLY18, XGH18, dMPF15, BG15, CM19a, CFMN15, DKP^{+}19, DQIM18, DCNY18, FJD17, FS15, GYLQ16, HH19a, JHJB17, JYP18, JQWW16, ME18, Mar19b, OAN17, SBP17, SSCM18, Sun16, TEG15, VFN18, WCS17, ZG17, ZWWW19]. Dynamical [ALS19, BHMT16, ADMV19, AY19, APYW19, BJK19, CJT17, CP16, DMM18, KVO18, KBT15, Pul17, SCMS18, SM16, TJSW19, Xu15]. dynamically [CPZ19, MDG17a]. Dynamics [CCT19, HVMR19, KSD17, LZX19, ACLM18, ACMT17b, Bad18, BJK15, EGK17, ESSE18, GOGM17, GKN18, GKM18, Hei15, HMT15, LS17a, LKN17, LMUZ18, MG15b, MA16, Nat19, Ogu17, PC17, PC18a, Yua15, ZCST18, ZHH^{+}19].

Ebola [LRRV18]. economic [SGK16]. economical [SSW19]. economics [BGNGR^{+}17, VAMGR17]. economy [SSCM18]. eddy [AL19c]. Eden [ADMV19]. Edge [RCPA15, RRS19, BCS17, GJ17, KPS16]. Editorial [Ano18b, Ano18c, Ano18d, Ano18e, CBP18, DLY18, LPW^{+}19, ZDFF18, Ano18a, Ano18f, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18g, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r]. Effect [GCTGC16, TLL16, BYYV16, JBFS15, JL19b]. Effective [LHJR15, QDBG18]. Effectiveness [BYWPA19, CNJM18]. Effects [KCDACC17, ABL^{+}19, BD19, HJ17b, Lu18, NCF^{+}19, SH15]. efficiency [ALMLM16, CCC19, HMT15, KElY16, PP19, SS18b, SRH17]. Efficient [AIS^{+}17b, ACEJ18, BBCP16, BZ16b, CFSS18a, CDR18, CC15, DM16a, DGN19, HJ17a, HOM17, HRO16, Jan17a, JN19, KS17b, OG16, Pap16a, Prz19, RANG15, SS16a, SW17a, WLLC18, XMW^{+}15, XM16, XLS^{+}19, ZuI17, ZDL16, AFH16, ABHVM16, AAB^{+}18b, AR15c, BGMM18, BAMM19, BSGM19, BFMP19, BLM16, CHX18, CGSS17, CJ19, CGvA^{+}15, CJSCT18, DIS^{+}18, DIP^{+}19, DLM16a, Fat16, GB18, Haj19, JK19, KR18, LQZ16, LRL^{+}18, LZ16c, LL18, LSL15, MD19, MMD15, NBAK19, OAN17, PC18b, RVC19, Sar19, SS16c, SAH15, SKBT16, SK15, TP19, WXJ19, WCW19, Wat16, XqYL^{+}15, Xu17, YCX19, YOK15, YP17, dBGMUA18]. efficiently [CJ18]. eigendata [YB15]. eigendecomposition [BW16]. eigenfunction [LQZ16]. Eigenmaps [YZ15b]. eigenpairs [AvG16, CQW17, KLL18, LLV18a, Mia18, RS18, Xie19]. eigenspaces [Tru19]. eigenvalue [Ais17a, Bog16, CDH^{+}17, CHYZ19, CDR18, DZ18, GLT16, GLZ18, Haj19, HMP15, JL15, LKS17, LZL16, LLL19b, LWZZ19, LX15b, MJ18a, MT16a, MT19, OV17, PMN19, QWS15, Reu16b, Reu17, SvG15, VK16, WD16, WP16a, Xie19, YBLH17, YSG18, YB15, Yin19]. eigenvalues [Elo19, KLK16a, LX15b, SY18, Vej18]. eigenvector [CS19, FR18a]. eigenvector-like [FR18a]. eigenvectors [BB16, SvG15]. eight [MA16, PS15]. eight-step [PS15]. Eighth [HH18a, BGMM18, CMQT16, GKN18, LKN17, MBM17]. eighth-order [BGMM18, CMQT16, GKN18, LKN17]. eikonal [CV19]. Einstein [Bad18]. either [RRM19]. elastic [AG16a, CLS18, CLZW19, CY15, Don18, ESI^{+}19, RL18, TLLL19, VMR17, XH15, Zha19a, ZBL^{+}16]. elasticity [CZ17b, CL19c, DGM16a, FCM19, GKPS19, HQ18, LSD15, ML19a, PHCG16, QWZ16, SL18, WXW15, WWWZ16, Yi19]. elasto [HR16]. elasto-plastic [HR16]. elastodynamic [Ceb17]. elastodynamics [ZYF19]. elastoplasticity [CH18]. electric [TKMT17]. electrical [AMM19, CMHHS17, Sch18, dAFvdMR19]. electrocardiology [OS15]. electrochemical [HHS^{+}15]. electrodynamics [VF17]. electromagnetic [BN19, BFLP15, CMM17, DLM17, LMM19, PSS17, WSG^{+}19]. electromagnetics [vtWvdHvdVV16]. electromagnetism [Rad17]. electroneutrality [ABL^{+}19]. electrophysiology [COdSR16]. electrostatic [MXW17]. electrostatics [DXZ18]. Element [BBG19b, BV16b, BV16c, GLS15, GTC16, LS17b, MWY16b, PG16, RP15, SVL19, ZWW^{+}18, ACLY19, AHMY19, AFM19, BV18, BK15a, BTKM15, BCS17, Bou18, BH17b, BN15, CCZZ16, CM19a, CDF^{+}16, CL19a, CHS15, CZ17b, CFX17, CCX18, CZ19b, CK16, CY15, CL19d, CL19c, Cou19, DKP^{+}19, DS19, DAM15b, DA19, DXZ18, DGM19, ES18, DFS15, DD18b, Dun17, DAAF16, EJJ15, ESI^{+}19, FLN16, FCM19, GCZ17, GZD^{+}19, GDQ17, GD17, GG15, GM19, GHL19, GSRJ18, HSWH15, HH16, HS18c, HM19a, HHKR15, HC15, HRGZ17, HHZ18, HMY19, JM15, JL16a, JL16b, JLLZ16, JLCZ17, Ku16, LS17a, LC15a, LKS17, LHK19, LA15c, LL16b, LHY16, LLCG16, LMW18, LCH18, LGLC18, LGYC19, LLG19, LSYY19, LYS19, LHY19, LLSM15, Lin16, LLC18, LLY19, LFKG16, MM15b, MWY15b, MWY15a, MWY16a, MY18, Mu19a, NSYX17, OZ17, QM15, QDBG18, RP16a, RR16, RR18, SKB15, SSW19]. element [SYZ19, SAT15, SSZL16, Sun19, SH15, SI16, TH19, Tho16, TZZ18, VFGPG^{+}18, WWWZ16, WWZZ16, Wan18a, WCW19, WZWJ19, WD16, WAGVRB19, vtWvdHvdVV16, XDW18, YZ16, YCF18, YSL^{+}18, ZNG15, Zha15b, ZZZ^{+}17, ZHY^{+}17, ZCZM17, ZZZ18, ZYZ18, ZL18b, ZL19, ZHC19, ZCS^{+}19, ZYF19, ZYL19, ZG19c]. element-free [DAM15b]. element-upwind [HSWH15]. element/finite [DA19]. element/volume [AM16]. elementary [MVC19, Mel19, ZS18c]. Elements [MS15b, AM17a, Arm18, BI17, CFJMVG16, Che18, DLM^{+}19, GK19a, HLZ19, KYS16, LSLZ17, LN17, LN19, MLSK16, XH15, ZLS^{+}15, ZM16a, ZML16, ZG19c]. elevation [SYT16]. eliminating [ARTY19]. elimination [CH19a, ML17a]. ellipse [UG18]. ellipsoid [JCH17, UG18]. Elliptic [FMC^{+}18, Bog19, BW19, BLP17, BK19b, BySZ19, BH17b, BN15, CWXW17, CTZK18, CZ15, CZLV17, CH19a, CLV16, CL19d, DYZ18, FLZ19, Fuk15, GUB^{+}17, GFPGF19, HM16, HH16, HLZ19, JL16b, KCX17, LX16a, LCH18, LCZ18, LW19b, LL17, MJ18a, Mat18, MS19a, MZ19, Mu19a, Mu19b, Mus17, OCU17, PW16, SSW19, SYZ19, Sun19, SI16, VARC17, WHWY18, XDW18, XZN^{+}19, YP18, Yan17, YGC19, ZNG15, ZCL16, ZG19c]. Elsasser [LHK19]. embedded [AK15, Elg19, PS15]. embedding [TSMO17]. Emden [RMA19, SGG19, SK15]. Empirical [ZLPF19, Nke18b, YS19]. enabling [QDG^{+}15]. encoding [MJ18b]. encryption [SE18]. end [BGNGR^{+}17]. endpoint [Wan18b]. energetic [ADG19, HL18b]. Energy [BGS19, CH17b, Son16, Yos17, CFWW19, Don18, FS19b, LD19a, LD19b, LMW18, LHY19, Lu15c, MDVR19, SS18b, SEV^{+}18, WCW15b, XDY16, YZH18, Yao18, BGI18]. Energy-conserving [BGS19, FS19b, WCW15b]. energy-preserving [LD19a, LD19b, LMW18]. energy-stable [SEV^{+}18]. engine [TBAH16]. Engineering [SDG15, VACR15, VAMGR17, ZDGM17]. engines [OLRPdCJ19]. Enhanced [AW19a, OD17, GSM18, PS18]. enhancement [LZL^{+}18a]. entire [SH19]. entries [BCC^{+}18, YB15]. entropy [DSDV15, JD16, KPP19, MJ18b, PT17b, RKS16]. envelope [BRW16, YP17]. envelopment [KElY16]. environment [KJ19, NS18, Nke18b]. environmental [RTV15]. environments [FLMS16]. epidemic [FH16, KJ19, KPW17, LPST17, LZX19, ZLF18]. epidemics [BJK15]. equality [BLXD18, Dia18, LMUZ18, Mar19a]. Equation [DWS^{+}18, AVB17, AHM18, AKHSU18, AMP18, AL19b, AHL17, AR19, AK17b, AFN17, Bac18, BBPS17, BBCK19, BGMS16, BT17, Bay15, BBQ18, BLM16, BKN16, Bou18, BGS19, BMG^{+}19, CLS18, CLC18a, COP17, CJCDLL15, CRGB17, CR17, CYY15a, CDW15, CCW17, CCNQ^{+}16, CCNQ^{+}17, CRR16, CL19a, CPB19, CJLT15, CHR^{+}15, CLC16, CQ17, CL18b, Che19b, CXZ19, CZ19b, CFWW19, CRG19, CV19, CJV17, CNQR^{+}18, CH17b, DJC15, DP17a, DA19, DLM16a, DT17, Din16, DMS19, DB15, Dos19, DT16, Dun17, ESSE18, Els15, FM19a, FLJT18, FYY17, FFP16, FJLB18, FS19b, GLS15, GC15, GCZ17, GCY18, GHM17, GHB17, GR16, GT19, GHL19, HCL17, HA18, HKAR18, HSWH15, HL18b, HMH18, HWH19a, HZ17a, HLZ19, ID18, JM18, Jeb19, JHK17, JL16a, JK18, KV19a, KR17, KDB17, KH18, KV19b, KV17, KS16, KS17b, KAA16]. equation [KAS18, LP19, LHJ16, LMW18, LLL19a, LSDY19, LH19, LHY19, LS18a, LZ15, LZZL17, Liu18, LW19c, LZRH19, MDG17a, MDVM17, MDGVR17, MDG19, MM18, MK17, Mia18, MS17c, Miy19, Moh19, Mol15, MFWK18, OSMZ18, PD17, PPM16, PA19, PT17a, PST18, RTV15, RQ16, RCBG17, RCBG19, Ros17, SR16a, SEV^{+}18, SOMF19, SB18, SEÖ16, Sey18, SLA19, SSZ19, SWH16, SJ17, SJ18, SVAKGR19b, SKBT16, SS19b, SFGJR^{+}16, Son16, SYW19, SO16, TJB17, Tho16, TKMT17, TNTL18, UG18, VFFGV16, VFGPG^{+}17, WCW15b, WWZ15, WW16a, WHZ16, WSL17, WD18, WLW19, WD19, WSL19, WCW19, WR16, WCS18, Wol17, Wu17, XH15, Xia17, XY16, XZ19, XD17, Yak19, YZH18, YCX19, YX19, YAA19, YB17, YFS18, YH18b, ZL18a, ZM15a, ZJ17a, Zha18, ZH18, Zha19a, Zha17b, ZZG^{+}19, ZCD17]. equation-based [Els15]. Equations [BCMP16, AFH16, AÖ17, ABN^{+}18, AH16a, AKR18, AL18, ALC19, Akh19, AM16, AN17a, AN19, AST19, ABK19, AL16, AK19, ABHVM16, ABBB17, AR15a, AR16, ACAC17, ACDT19, AZ17, AW19b, ABT16, AAC17, AJ17, AJY18, AD18, AHKT15, AHK19, AKM18, BM17a, BM17b, BM18, BJM16, BH17a, BJ16, BBCP16, BBC^{+}18, BO18, BBB17a, BDA17, BK15b, BB18, Baz15, BPP17, Beh16, BMM17, BAK18, BS17b, BKLR16, BG17, BGL17, BHRR18, BS19, Bie17, Bog15, BLP17, BS15, Bot16, BAI17, BK19c, BH17b, BR15b, BK19d, BCVV18, CCJ18, CMHHS17, CGPR17, CW15a, CHS15, CT16, CW16a, CFX17, CLW17, CHS17, CZLV17, CLC18b, CL19b, CLA19, CK16, CS18b, wCW15a, CN18, CPZ19, Con18, CMQT16, CLM19, CNQRR18, DDR19, DKL^{+}18, DBX19, DF15, DNL15, DNL16, DHAM17, Das18]. equations [DL16a, DSA15, DAM15b, DG16a, DFLM19, DIJ16, DY15, DYZ18, DD18b, DLM^{+}19, Dzh16, Dzh18, Elg19, Els15, EC18, EBJS19, ERR18, EHTL19, EE16, FwC17, FNZ18, FL23, FSNS16, FLN16, GZD^{+}19, GLL19, GDCK19, GGG19, GUB^{+}17, GMG19, GV17, GYS17, GAA16, GKN16, GLSW16, GLSW18, GR19, GHZ^{+}15, GG15, Güc17, GM17b, GGC15, GH16c, GGRN17, GQM16, Guo17, GLMY18, GL19, HMD19, HMT^{+}17, HHOT17, HP17, HR15, HSAR18, HXX16, HH16, HM19a, HCA^{+}18, HHEG17, HH18a, HS17b, HSAA19, HZ19, HLM18, HMY19, HB16, HW17, HIS16, IIA15, IN16, IÇ18, JHJB17, JAN17b, JBT16, JB18, JY17, JL19a, JZ18, JGTB15, JC15, JMS17, JD16, JLCZ17, JS19b, KB17, KP15, KS17a, KGES19, KRR18, KMB15, KKS16, KR16, Kha17, KDW17b, KDW17a, KYS16, KLC17, KMV16, KHLY17, KV18b, KPT15]. equations [KEB19, KBG17, KR18, KW18, KB16, KPS17, KtTBK18, KMB^{+}18, KK19c, Lad15, Lan18, LX18, LXW19, LW19a, LLT16, LC15a, Lee18, LHK19, LaT15, La19, LP15b, LCL16, LW17a, LWL17, LH18, LSG18, LHLL18, LDG19, LLG19, LLW18, LZ16a, LYG17, Lin16, LC18, LCZ18, LL15, LZY15, LM15, LZ16b, LTY17, LDW17, LHPZ18, LDW18, LF18, LLC18, LQ19, LSXL19, LLY19, LFWC19, LL17, LNP17, LSL17, LSL18, Luo18, LYG15, LM18, MFX15, MDT16, MDG17b, MDAP19, MVM19, MG15a, MMO15, MS17a, MD18, MMS19, MN17, MN19a, ME18, Mao15, MYM16, Mao16, MT17b, MT17a, Mär17, MHV17, MPH18, MV16a, MS19a, MII16, ML19b, MOPP18, Men17, Mic18, MMZ19, Mil15, Mil16, MT16b, MV16b, MSV18, MV18, MHL16, MH16a, MH16b, MA18, MSH18, MAS19, Miy17, MM15c, MTKO17, MC18, Mok15, MR15, MJS19, MWY15b]. equations [MTS17, MS15b, NDN19, NE17, NBAK19, NN18a, Nem15, NSM16, NT15, NSW15, Ngo19a, NCPD19, NAM18, OM17a, PS18, PSK16, PMN19, PBP16, PW16, Par15, PRP19, PB15, PHD17, Pis15, Pis17, Ple17, QLT^{+}17, QM15, ROB17, RRM18, RVA15, RM17, RWW17, RZ19a, RC19, Reu16a, RKS16, RM19, SF18, SE16, SAC17, SLH18, SAA18, Sak17, SHS^{+}15, Sar19, SM19a, SDH17, SR16c, Sha15, SPK18, SS18c, SLS18, SJ17, SK16, SGG19, uIuD18, SK15, SAT15, SYD15, SYD17, SJL16a, SC18, SZJ19, SRH17, SSG16, SH18, SSZL16, SLW17, TMS^{+}17, TYS17, TH19, TX19, TV18, TZZ18, TP19, TGSS17, TN19, Vab18, VG17, VX19, VL15, WWZZ16, WWM17, Wan17b, Wan18a, WG18, WKBB17, WZ17, WL17a, WA17, WLM16, Wu19, XHZ17, XY19, XDY16, XF16, XH19, Yam17, Yam19, YY16]. equations [Yan17, YHZ18, YL19a, YZZ19, Yao18, YTSZ15, YJBM19, YFwC19, YZ15c, YH18b, ZCST18, ZS18a, ZSL15, yZLX15, ZM15b, ZZW16, ZH16, Zha16, yZ17, ZCFW17, ZZZ18, ZT18, ZSL18, ZX19, ZP19, ZL17c, ZJ17b, ZG19b, ZZ15, ZN19, ZWH15a, ZWX18, dlHV16, iW16]. equations- [wCW15a]. equiareality [YZ15a]. equidistant [AS19b, CH19b]. equidistributing [HH16]. equilibria [RANG15, Uda18]. Equilibrium [XY15, ABT17, AG16a, BHMT16, DMMSK17, DK16, FZ15, Man15, MP18, NDH16, NH19, OGR17b, OGR19, SJRR17, Sha15]. EQUIP [BGI18]. equity [PZ18, ZY19a]. equity-linked [PZ18, ZY19a]. Equivalence [HLBP15, GKBG17, MP15, MSZ18, ZS18a]. equivalences [HJS19]. equivalent [HJL18, NP19]. equivariant [RFO17]. Erlang [BRA16, GHL18]. Erlangized [DYD19]. erosion [AFdOdS15, ZFK18]. Errata [MHY21]. Erratum [BKMO20, DNL16, JSW18b]. Error [AZ17, ALMLM16, AJ17, BJM16, BS17c, CM19a, DDP16, DVS15, DAM15a, DM17a, DA19, GWLL18, ID18, KVO18, Lad15, Luo18, MAK19, Mir16, SSZL16, TX19, VS15a, XDY16, ZHS18, ZLAA15, ZC16b, AA16, AHMY19, Akh19, AK18, AW19a, AHKT15, AHK19, AK17c, BYW18, BBGG19, BLM16, CLC18a, CZ19a, wCW15a, CST^{+}16, Con18, DAM15b, DS15b, DGM16a, EJJ15, FSNS16, GZD^{+}19, GYX17, GPHAPR15, GP16, GGRN17, GHL19, GSRJ18, HOS16, HH16, IN16, Jan17a, JBFS15, JMSL19, KS18, KP15, KMB15, KH18, Ku16, LC17b, MTCC15, MM15a, Mat18, MOY18, Miy15, Mu19a, Mu19b, OOO15, PBP16, Par15, Prz15, SF16, SW17d, Tho16, US19, Wan15a, WWZ15, WHZ16, WKBB17, XY19, YBLH17, YGC19, ZCZM17, ZLZ17, ZP19]. Error-adaptive [MAK19]. error-free [OOO15]. errors [EJJ16, GH16b, LX15b]. erythropoiesis [ACLM18]. Escherichia [BYWPA19]. ESERK5 [MVK19]. Essential [Mär17, Wu16b]. essentially [JN18]. established [CNQRR17]. estate [HW18a]. Estimate [CZ15, AJ18, CZ19a, DAM15a, DAM15b, DA19, GVHMR17, Mu19a, Par15, WHZ16]. estimates [AA16, AMRM17, AK17c, BS17c, CLC18a, CM19a, FwC17, FM19b, GYX17, GPHAPR15, GP16, GGRN17, KC17, KP15, KMB15, KH18, Ku16, MPH18, Mu19b, NMKK19, NW19, PBP16, PGRC18b, TX19, US19, VS15a, XDY16, YBLH17, ZLZ17, ZHS18]. Estimating [ZS19, CL18b, EJJ16, KElY16, KK17, LYG15, VZDV18, WSL17, YY16]. Estimation [GVMR16, NAK17, Wan18c, Wan18d, AG18, AJY18, BSLL18, BSLL19, CCY18, CCZ^{+}18, Con18, CAO^{+}19, DLM16b, EDO19, FR19, GLTW15, GK19b, HZ17b, IANS15, JJX17, LC17b, LSZB17, MTCC15, Mar18, MM15a, NADK18, SF16, TJA18, Wan15a, Wan15b, WS19, WP16a, WS16, Xu15, YW16, YC19]. estimations [DWQ16, Moh17, YGC19]. estimator [AHMY19, DGM16a, GYQS15, RW15, Wu16a]. estimators [AHKT15, AHK19, GWT18, Jan17a, KD19, MRMA15, MRMA17]. ethanol [DQIM18]. Euclidean [ADD^{+}16, BAAA^{+}15]. Euler [TGJ19, CW15a, Dub16, DAAF16, FGS19, HK07, HLM18, KHLY17, LYG17, Mao15, Mao16, Mil16, MSH18, OM17a, RT18b, SRH17, YZZ19, YH18a, yZLX15, ZSL18]. Eulerian [BSS15, LS17a]. European [RZ19b, DS17a, HZ18b, HT18, JSW18a, MV16d, NP19, SPG^{+}16b]. evaluate [SSCM18]. Evaluation [CNJM18, IGRL17, AL18, BEHL17, Ceb17, CY18b, DM16a, EJJ15, ET19, HX15, HZ17a, MIT15a, NGGA^{+}18, RNM16, RANG15, SIAJ^{+}19, Sho18, WLC^{+}18, ZLS^{+}15, ZCS^{+}19]. evaluations [EBJS19]. even [TDA18]. Eventual [AKHSU18]. everywhere [KP19]. evidence [SM19b]. evolution [ABN^{+}18, AHK19, BYWPA19, BHRR18, CXZ19, HND18, JC19, KSW19, LLT16, LLHO18, LQ19, NY19, Ple17, SJ17]. Evolutionary [KS17a, PBS15, AAC17, HH19a, ICMPG16, MM15c, Ngo19b, SM19a, SW16a, SW16b]. evolving [SAT15]. evolving-in-time [SAT15]. Exact [KKK16, NCF^{+}19, RdSMO17, RdSMO18, Zha18, AR15b, CCIL18, CD16b, GL19, HZ17a, LY15a, LXW19, LL16b, LSH18a, LSXL19, MDVR19, NL17, NC19, RCBG19, Rót15, ZSL15]. Example [BYWPA19]. examples [Yam19]. excess [ALS18, QLZ19]. excess-of-loss [ALS18]. exchange [PG19, TBAH16, ZS17c]. exchangeability [BKMO19, BKMO20]. exchangeable [BE15, EG17]. excited [KTT16]. executing [DS15a]. exhaust [DMCC^{+}18]. Existence [BYYV16, BT17, CLS18, CFG^{+}19, FMC^{+}18, HSAR18, LZY15, Ros17, FM19a, HKAR18, LSW19, MTKO17, SLW17]. expanded [CW16a, LLG19, YP18]. expansion [GLP16, KDB17, LHM16, ZS19, ZY19a]. expansions [EW18, HOM17, KSDM16, LQZ16, LWC17, NSYX17, PHCG16]. expectation [Sun16]. expectations [Hag15]. expected [PX17, WS16, YCW17]. expenditure [YL19b]. experiment [SR16a]. experimental [BTKM15, DGN19, NGGFM17]. experiments [Goć15, NGGFM16, TYQ^{+}19, TN19]. expiry [KDB17]. Explicit [AHBC15, HK18, Hu16, Lu15a, MS18, MF16, ALC19, ACLM18, BKK18, BCMP16, DBdLdM^{+}19, DIJ16, FR18b, GPHAPR17, GQM16, HMD19, HPW16, IDGH18, JS17, JK18, LSW19, MVK19, MP18, Sta19, SV16a, WLZ15, WZS18a, WLM16, YGC19]. explicit-implicit [ALC19, BKK18]. exploiting [lCCVV18]. exploration [JGGS16]. exponent [BGMS16, SV16a]. Exponential [AAKW16, BM19a, BHU18, DT16, LY15a, LJ15, MP17a, WW19a, YL19a, ZSL15, ZLF18, ZWH15a, AMP18, AMCR19, AK19, AHS19, BZ19a, BR15a, BK15c, Bot16, BK19c, CZ19a, CW15a, CY18a, DEY19, DLS17, DIS^{+}18, DMR19, FdOP17, Gun19, GUB15, JY19, KV17, Lan18, LPT17, LYG15, MDGVR17, Mil15, NS17, NN17, Nie15, OK16, OAS19, RSID16, Sha15, SJL16b, WCW15b, Wan18d, WZ16, ZNS18]. exponential-Lévy [DLS17]. Exponentially [CDP17, MR19, CL19a, FR18b, Mok15]. exponentially-fitted [FR18b]. exponentials [DSA16, GAA19, GAA21]. exponentiated [HG19]. exposure [PC18a]. expressions [JN19, RdSMO17, RdSMO18]. Extended [AY19, AM15a, MS17b, NP15, WGLV16, ZL17b, GC18, HWZ18b, KD19, KK17, OK16, QDY^{+}19, Ruj16, WL19a]. Extending [AM15b, AEHVM18, BDGO15, JMN^{+}16]. extension [DP19a, DP19b, DPN15, GHC16, KBG17, NADK18, NY19, dFG19]. extensions [OV17, ST16a]. extensive [ZX16]. external [EGK17, SS18b, ZWWW19]. Extinction [DV16, VM15]. Extinction-time [DV16]. extraction [DS15a]. extragradient [TH18]. extraneous [GKN18, LKN17]. extrapolated [BK15c, MVQDdR^{+}18, MVK19]. Extrapolation [ZDF^{+}17, BR15a, Tan17, ZHWS16]. Extremal [DWXO15]. extreme [Ahn19, ID16]. extremum [WZS^{+}18b].

F [DLY18]. fabric [DWS^{+}18]. Fabrizio [FJLB18, YMGA19]. facilities [FLMS16]. factor [AL19a, ACLM18, GVMR16, HT18]. Factorization [LSS19, SVGM17, Byl18, HY19a, LHJR15, SY18, SY19, WCW18, YLZG18, ZST15, yZyMxG^{+}16, dBAAA15]. factorizations [Li18]. factors [BR15b, DB15, KL15, WW16a]. fading [Liu16, WS19]. Faedo [CPB19]. fail [LO19]. fails [RVA15]. failure [GMSR18, GMSR21, NMKK19]. Fair [KP19]. faithful [CNT17]. Falkner [RMVA17]. Falkner-type [RMVA17]. falsi [SV19]. families [CCT19, DD17, HMT15, MBM17, SPCSA16, Zha18]. Family [YL19b, ARTY19, ACMT17b, BMM17, BAMM19, BSGM19, BSS15, CCTV17, CMTV17, CGT18, Das18, DG16b, DLM^{+}19, GKN18, GKM18, HH18a, HH18b, HVMR19, KDB18, KMB^{+}18, LKN17, LX16a, MLSK16, NKD^{+}19, OK16, OCU17, OGR17a, LGB18, OCU18, PP18b, SBC16, SS18d, Wan15b, WWL15, ZJ17a, ZNS18, Zha19a, ZS17b]. far [WSG^{+}19]. far-field [WSG^{+}19]. FARGO [LL16a]. Farlie [DG16b]. Fast [CDDD19, DMP19, FM19b, HL15, HZ17a, KV18b, Miy17, Miy18, SIAJ^{+}19, TB18, WKXT15, WC17, WCLL18, WZS^{+}18b, ACDT19, DGH15, DKLP15, DMS19, EK19, EL19, Fuk15, GH19a, GGC^{+}18, JCH17, LFWC19, LL17, LK18, MT17a, NSM16, NAM18, PW17, PST18, RMA19, SAH15, STS15, ZX19]. fast-converging [RMA19]. FDE [ZZ19]. FDEs [AJ18]. FDM [Mer19]. FDTD [JB18, NZ15]. FDV [Par15]. FE [ZCFW17]. feasibility [LDL^{+}19, MMRS15]. feasible [TBY17]. FEAST [Yin19]. feature [LZ19, NLG15, TYQ^{+}19, WZBM18]. features [LAVMA18]. featuring [RTV15]. feedback [YT16, ZWWW19]. Fejér [AAKT19]. Feller [THL^{+}17]. FEM [AG16b, AMP15, BOF^{+}16, CLC18a, GM16a, GYX17, ML19a, SL18]. FEM-based [BOF^{+}16]. FEMs [KP15]. Fermi [LZ15, PD17]. ferrite [LHY16]. FETI [BKK18]. FETI-based [BKK18]. fever [BdSFL16]. FFT [AL18, GKQW17]. FFT-based [AL18]. FGM [TJA18]. Fictitious [Sun19, Cou19, ZG19c]. field [AG16a, CS17, DLY18, LWW17, LAVMA18, MS15a, TKMT17, WSG^{+}19]. fields [Che16a, DD16b, FFN^{+}17, HLMS19]. fifth [CHVRT15, HWC16b, MVK19]. fifth-order [CHVRT15, HWC16b, MVK19]. figures [OGR17b, OGR19]. filled [ZXQH17]. filling [DFGS16, MPS19, YSCY17]. film [AN17a, AN19]. films [BG15]. Filon [LTY17, MFX15]. Filon-type [LTY17]. filter [HP15, LWLL18, LLD18, SRBD18, SZM^{+}18, UPD^{+}15, WZBM18]. filtered [Kaz18]. filtering [DLM16b, Liu16]. Filters [CRSV19, KK17, LLT16, LAVMA18, LRKV16]. final [CL18a]. finance [DKV16, HHEG17, LQZ16]. financial [AdHR17, BDF15, DSDV15, DF16b, Hai17, Nke18a, WSW15, YL18d]. find [DD16c]. finders [GKN18, GKM18, LKN17, PN16]. Finding [SAH15, BIR^{+}17b, BW17, CN18, CT15, FWW^{+}18, GZDC19, Kow17, LV16a, MMD15, PP18a, PP19, RP16b, XYZ15, Yao16, Yao18, ZCST18]. fine [ZFK18]. fingering [ZBL^{+}16]. Finite [ASKG18, ABL^{+}19, Bou18, BV16b, BV16c, CWXW17, CEJV16, ESI^{+}19, GLS15, GTC16, JS19a, LS17b, LV16b, MC19, MWY16b, MS15b, PG16, RP15, SVL19, SH15, SI16, WCW19, ZWW^{+}18, AHP19, AALMLM18, AM17a, ACLY19, AHMY19, AM16, AdHR17, ACAC17, AW19b, ABT16, Arm18, AFM19, BV18, BSLL19, BS18, BI17, BK15a, BTKM15, BSS15, BCS17, BH17b, BN15, CCZZ16, CM19a, CDF^{+}16, CL19a, CPB19, CZ17b, CFX17, CCX18, CZ19b, CFWW19, CK16, CL19d, CL19c, CRR19, CEJ18, CROC16, Cou19, CP16, DKP^{+}19, DF15, DMMSK17, DA19, DXZ18, DGM19, ES18, DFS15, DD18b, Dun17, DAAF16, DP19c, EG17, FFN^{+}17, FLN16, FGF19, FCM19, GCZ17, GZD^{+}19, GLL19, GDCK19, GUB^{+}17, GG15, GM19, GHL19, GSRJ18, HS19, HSWH15, HH16, HS18c, HM19a, HHKR15, HHEG17, HC15, HRGZ17, HMY19, HLZ19, IGRL17, JM15, Jeb19]. finite [JL16a, JK18, JL16b, JLLZ16, JLCZ17, KD17, KYS16, Ku16, KAA16, KBL19, LS17a, LC15a, LBL15, LKS17, LHK19, LZC15, LA15c, LHY16, LSZ16, LLCG16, LCH18, LGLC18, LGYC19, LLG19, LSYY19, LHY19, LLSM15, Lin16, LWZ17, LCX18, LLC18, LLY19, LW19c, LMA18, LN17, LN19, LFKG16, MDG17a, MVdRE^{+}17, MVQDdR^{+}18, MM15b, Moh17, MWY15b, MWY15a, MWY16a, MY18, Mu19a, NSYX17, OZ17, QM15, RNPK15, RP16a, RR16, RR18, SBU^{+}17, SS19a, SKB15, SSW19, SYZ19, SKBT16, SAT15, SSZL16, Sun19, TYS17, TH19, TBR16, Tho16, TZZ18, Tra15, UGG^{+}19, VFGPG^{+}18, VM15, WWWZ16, WWZZ16, Wan18a, WLW19, WD19, WSL19, WZWJ19, WL19b, WD16, WAGVRB19, Wu17, XH15, Xin16, XDW18, YZ16, YZY17, YCF18, Yos17, YSL^{+}18, ZNG15, Zha15b, ZM16a, ZZZ^{+}17, ZHY^{+}17, ZCZM17, ZZZ18, ZL18b, ZL19, ZKRL15, ZML16, ZG19c, dSBV16]. finite-difference [KAA16, MDG17a]. finite-dimensional [KBL19]. finite-element [AHMY19, AM16, HM19a]. finite-element/volume [AM16]. finite-part [LWZ17]. finite-time [YZY17]. finiteness [BPD15b]. FIR [UPD^{+}15]. First [STZ16, SYD15, CQ17, CNQRR18, DKL^{+}18, DBdLdM^{+}19, DLM16a, FF15, GOGM17, GLSW16, GLSW18, GGRN17, Guo17, Kha17, KKW16, Ku16, MTS17, NAM18, OP19, OGR17b, Par15, RRM19, SK16, VARC17, YP17, ZJG^{+}15]. first- [RRM19]. first-order [CNQRR18, FF15, GLSW16, GLSW18, GGRN17, Guo17, Ku16, YP17, ZJG^{+}15]. Fisher [BK19c, CRG19, MDG17a, MDGVR17, MDG19, RCBG17, RCBG19, Sta18]. fit [CDN19]. fitness [GLdCJS17]. fitted [CDP17, DP16, FYY17, FR18b, KMRS16, LWDW18, TFS17, VAR15, YAA19]. Fitting [GHL18, TK16, DT16, LXD^{+}18, LLG^{+}18, LZ19, VARC17, ZTGZ18]. Fixed [SHS^{+}15, AMB19, BGL17, BP17a, CT16, DTKP17, FZW19, GKN18, Iid15, KEB19, KW18, KS17c, LKN17, NCF^{+}19, PAC15, RMMRC15, dSBV16]. fixed-stepsize [KW18]. fixing [CEJ16, PCJ17]. flame [CD16a]. flapping [PST18]. Fletcher [JJ19]. Flexible [SHCJ19, YW16, xZWlC15]. floating [JHK17]. flooding [hShX17]. Flow [YWW19, AN17a, AN19, AC17a, AVGCMVM17, AdQdS16, ABB^{+}15, AB16, BM19b, BBPR15, CYY15a, CM19a, CZ17b, CM18b, CFS19, CSW19, CLZW19, CLVW18, CZ16, DLW16a, DS19, DSOBK16, EKKT17, EAKS18, ELRR17, GCTGC16, HCSC16, JO17, Kau16, KAM16b, LWW17, LW17b, LLSZ17, MFWK18, OSC15, PBP16, PST18, PG16, RNPK15, RP15, SJRR17, SVW19, SLH15, TLL16, VCC^{+}19, WCSL15, WZZ16, ZHY^{+}17, ZWW^{+}18, ZQJY18]. flowfield [Par15]. flowfield-dependent [Par15]. flowgraph [GMSRP16]. flows [AEKGP19, CVW17, ES18, GN18, GT17, IDGH18, KLM17, LW15, LHM16, LFKG16, SVL19, Xin16]. Fluid [BG15, CSW19, CZ16, ELRR17, Hei15, LS17a, LY15b, LLZ16, LWV18, LZS15, PBP16, PST18, SH15, Wat16, Yam19, ZHS18]. fluid-fluid [ZHS18]. fluid-structure [LY15b, SH15]. fluidized [MFK^{+}15, OSC15]. Flux [MSV18, Vej18, Abd18, BHN15a, BIJM^{+}19, KV16, KtTBK18, PT17a, TNTL18, ZH16, TBAH16]. Flux-Corrected-Transport [TBAH16]. fluxes [DGM19]. fly [CZ19a]. Fokker [FJLB18, HL18b, LP19, LP15b, SRA^{+}15, Wol17, YHZ18]. fold [WH16]. following [HR16, Wu18]. FOM [LH18, ZHWS16]. force [EGK17, GZH19, LHJ16]. forced [AKHSU18]. forces [BM19a, Cou19, Gor16]. Forchheimer [AVB17, AC17a, AFM19, KSW19]. Forcing [HRRC19, AÖ17, ZH18]. Forecast [NGGL^{+}19]. forecasting [BLH^{+}17]. foreign [ZS17c]. Foreword [BLLM^{+}15]. form [CFJMVG16, CGMGV15, DDP16, HZ18b, ID18, JB18, KP19, LA15b, LaTG17, LCC^{+}15, LWC17, NN18a, PRRY18, PSW18, RVA15, RZ19b, Ruj16, SB19, SSG16, TV18, YP18, dR16]. formalism [SO16]. format [GH19b]. formation [MV16c]. forms [Guo18, SuSR16, ZW16, dR16]. formula [CCIL18, Das18, Don18, Dub16, HZ18b, Kam16a, MB15b, RZ19b, Ruj16, RT18b, SE16, ST16b, SH19, WLZ15, WLZ17]. formulae [AN17b, Ber15, SH18]. formulas [BDGO15, BCCB15, CBMPP15, GPHAPR15]. formulation [AdQdS16, AKM18, BBGG19, CWXW17, CW16a, DJAMJ17, DYZ18, GMdV19, HL18b, JHPT15, KK19c, LHLL18, LZA19, MV16b, MSV18, MV18, MWY16a, NCPD19, WWWZ16, Zha17b]. formulations [BEMA16, DMS19, GKQW17, LSYY19]. Fornberg [BAI17, CRGB17]. Forts [YZ18]. Forward [ZS17a, HRO16, Sun16]. forward-backward [HRO16, Sun16]. foundation [ZP16]. four [ARTY19, CQW16, LZ18a, SVAKGR19b, ZHH^{+}19, ZHL15]. four-point [ARTY19]. four-step [SVAKGR19b, ZHH^{+}19]. Fourier [AK19, Bad18, CYY15b, DKV16, Elg19, GKQW17, HM19a, HOM17, HRO16, LDK17, Lee18, LZZL17, PW17, WSG^{+}19, WP16b]. Fourier-Continuation [Elg19]. Fourier-cosine [CYY15b, HOM17]. Fourier-splitting [Bad18]. fourth [BDA17, BMM17, BK16, CQW16, CFWW19, CV17, CMTV17, Dos19, HMT15, HIS16, LLL19b, MR19, MMZ19, PW16, RT18b, WD19, WSL19, WKBB17, ZJ17a, Zha19a]. fourth-order [BDA17, BK16, CV17, HIS16, LLL19b, PW16, WD19, WSL19, WKBB17, ZJ17a, Zha19a]. Fox [Meh19]. fractal [Bal17, CNJM18]. fraction [Ceb17, CPL^{+}18, CPL^{+}19b]. Fractional [Chu15, DQIM18, DLY18, MRS17, OM17b, TJSW19, Abd15, AAM18, AÖ17, ABN^{+}18, AAKT19, AA18, AM19, AK19, AZ17, AJ17, AJ18, AJY18, AR19, BKMS18a, BKMS18b, BDA17, BT17, BB18, BS17b, BS19, BLP17, Bou18, BCVV18, CHX18, CPB19, CJLT15, CW16a, CLC16, CLC18b, CLA19, CT17a, CKB17, DMM18, DBX19, Das18, DAM15a, DSA15, DA19, DG16a, DWXO15, Din16, EBJS19, FLJT18, FM18b, FJLB18, FMR15, GDM16, GZD^{+}19, GMG19, GB16, GMW18, GN18, GS15b, GHZ^{+}15, GG15, HCL17, HND18, HA18, HCA^{+}18, HMDSR19, HSAA19, HNT19, IAAW18, Kam16a, KCK18, KFT17, KLM18, KK19b, KJ19, KV19b, KPT15, KPS17, KAS18, LRRV18, Lee18, LW17a, LZL^{+}18a, LDW18, LLY19, LFWC19, LF19, LMA18, LNP17, MDAP19, MVM19, MHAA15, MMS19, MO18a, Mar19b, MO18b, MNFS18, MDH19, Mic18, MKKV19, MA18]. fractional [MAS19, MLS15, MC18, Mok15, NE17, NSM16, OKGB19, PD17, PRP19, PTV17, PHD17, PC17, QLT^{+}17, ROB17, RWW17, RZ19a, Reu16b, Reu17, SF18, SB18, SPK18, SLA19, SWZW15, SJ17, SJ18, Sho18, SS19b, SZJ19, SLQ16, SYW19, SI16, TJB17, TYS17, TAT18, THL^{+}17, TNTL18, VDT19, WWZ15, WHZ16, WPA^{+}18, WCW19, bWzZjH16, XY19, YGS18, YHZ18, YMGA19, YJBM19, Zak19, ZT18, ZH18, ZZG^{+}19, ZZ18]. fractional-in-space [Lee18]. fractional-order [LRRV18, LZL^{+}18a, MO18b, MDH19, MC18, OKGB19, ROB17, bWzZjH16, YMGA19]. fractional-step [AZ17]. fracture [EAKS18, LWW17, ZHY^{+}17]. fractured [CSS19, EAKS18, LCX18, VCC^{+}19, ZHY^{+}17, ZWW^{+}18]. frame [AMRM17, CG17a, DWS^{+}18, HH19b, LSL15]. frame-based [HH19b]. frames [FGS19, Kre19, Vis19]. framework [CFR18, CMHHS17, DS17a, HT18, HWH19b, OM17b, PX17, PG19, RS16, TYQ^{+}19, VŽ19, WH16, WLM16, XZXQ18, YL19a]. FRAP [MP15]. Fredholm [Beh16, DNL16, GAA16, XHZ17, BM18, BLM16, CHR^{+}15, DJC15, DNL15, Dzh16, Dzh18, FSNS16, HXX16, HS17b, JD16, KS17a, KGES19, LHPZ18, MMO15, MN17, MN19a, Mol15, NN18a, Nem15, PMN19, PPM16, Reu16a, uIuD18, WW16a, XY19, XF16, YJBM19]. free [ABH19, Ahn15, AMP18, ACDT18, BMM17, BGMM18, BL16, CCC19, DAM15b, Fan15, HHLS18, HVMR19, HZ17c, KP19, KMB^{+}18, KK19c, LF18, ML19a, OOO15, PCJ17, PCJ18b, RTV15, RS16, TGSS17, WWWZ16, ZHC19, ZC16a]. free-form [KP19]. free-surface [RTV15]. Frenet [DWS^{+}18]. frequency [BBD16, BHU18, CCY18, JHPT15, LMM19, MCGNBD17, PS15, VAR15, WWM17, WSG^{+}19, YC19]. frequency-dependent [LMM19]. frequent [MAF17]. Frequentist [SMTdF18]. Frequentist-Bayesian [SMTdF18]. friction [QM15, VMR17]. frictional [BS16]. friendly [HS17a]. front [CEJ16, FM17, PCJ17, SF16]. front-fixing [CEJ16, PCJ17]. frozen [AG19, GS18b]. full [CEJ18, GRS19, LDZ17, VS15a, dBAAA15]. full-pivoting [dBAAA15]. Fully [CYY15a, CFS19, GGRN17, GHL19, HC15, Kaz18, LSH^{+}18b, Pap16b, SK18, YP18]. fully-coupled [LSH^{+}18b]. fully-discrete [GHL19]. Function [NRV18, AQCJ17, AG16a, AMRM17, BSLL18, Bal17, BCJV18, CJ16, CP18, CSTU17, CNQRR18, CD15, CPL^{+}18, CPL^{+}19b, DP19a, DP19b, DNK19, Dia18, DV15, EDO19, FZW19, Fuk15, GD18, Gun18, GYLQ16, Jeb19, KD19, KK19a, KYY15, LKN17, LZS15, LYG15, MS17a, MRMA15, MRMA17, Mel19, MV16b, MLS15, MS17d, Moh17, PA19, STS15, WYW15, XLS^{+}19, YBC^{+}16, ZXQH17, ZS19, vV16]. function-based [vV16]. function-finite [Jeb19]. function-to-function [LKN17]. function-vorticity [MV16b]. Functional [Mat18, AL16, BJM16, BS15, CT16, FL23, GLR16, GYS17, LW17a, LNP17, MR19, Ros17, WL17a, Yao16, YGC19, ZJ17b]. functional-discrete [MR19]. functional-integro-differential [WL17a]. Functions [FLM17, AHM18, AAKT19, ADMV19, AIS^{+}17b, AR15b, AD18, (Ba18, Beh16, BS17a, BDFM19, BHN15b, CBC^{+}19, CGMGV15, Ceb17, CT16, CCT19, CT15, DDG^{+}16, GLdCJS17, GH17a, GKM18, GC18, GAB17, GWT18, HS17a, HS19, HX15, HS17b, Hon19, IANS15, IAAW18, KDB18, KL17, Kow17, LKN17, MAK19, Mar19a, MHY17, MHY21, MOY18, MA19b, Meh19, MT18, MH16a, MA18, MAS19, Mir16, MG17, MMRS15, MXW17, PM15, PD17, QL19, RP16b, dHKdHMM15, Rom18, SCP19, SH19, TLFC18, TBS18, UPD^{+}15, WGHR17, WCLL18, XW18, XHZ17, XY19, XD17, YSW^{+}18, YWY19, YCW17, ZCST18, Zha17a, ZHL15, ZD19]. fund [JYP18, JQWW16, Nke18b]. fundamental [AS19b, Bog16, BDGO15]. Further [SLQ16, Vig17, You19]. fusion [MDH19]. futures [GVHMR17]. fuzziness [ET15]. Fuzzy [TK15, YH18a, ABN^{+}18, AH19, AdHR17, Che19a, CLM19, GX17, Kha17, LC16, MDT16, MGMGJLRB17, SE16, SBP17, WH16, ZS18a, ZWWW19].

G [NHA17]. G-symplectic [NHA17]. Gabor [OZ18]. GAITA [ZPL18, ZPL17]. Galerkin [AHMY19, AST19, AD18, AK17b, BJ16, Bac18, BBD16, BKLR15, BKLR16, BW19, CPB19, CHR^{+}15, CCX18, Che19b, CL19b, CZ19b, CFSS18b, CW17, DKL^{+}18, DAM15b, DGM19, DLM17, DYZ18, DAAF16, FLN16, FS19b, GCZ17, GGG19, Gia18, GP16, GTX^{+}19, HWZ18a, HL18b, HCSC16, HMY19, HB16, KM18, KCK18, KCX17, LX15a, LRKV16, LMW18, LSG18, LCH18, LGLC18, LLSM15, Lin16, LLC18, LW19c, Men17, Mol17, MWY15b, MWY15a, MWY16a, MWY16b, MZ19, Mu19a, Mu19b, PMN19, PBP16, Sar19, SA17, SSW19, SLS18, SZ18, SVV18, SF16, TX19, TZZ18, VG17, VS15b, WWWZ16, WWZZ16, WZZ16, Wan18a, WZS18a, WZWJ19, XH19, YP18, YZ16, YL18c, Yi19, Zha15b, ZZLY16, ZZZ^{+}17, ZL18b, ZL19, ZP19, ZC18, ZG19c]. Galerkin-collocation [AST19]. Galerkin-front [SF16]. Galue [KK19a]. game [PW15b]. games [MTMZ19, NDH16, NH19]. gamma [BCJV18, CCRN19, FPGK19, Mar19a, NZ16, SKC^{+}19, SBC16]. Gander [PP18b]. gap [LC17a]. Garden [ADMV19]. Garden-of-Eden [ADMV19]. gas [ML17b, MFWK18, TV16a, TBAH16, WS18]. gas-exchange [TBAH16]. Gauss [ZPL18, CDZ16, CBMPP15, DRS16, FEC18, JR15, LLCG16, ML17a, NS16, PM15, PR15, SRBD18, ZPL17]. Gauss-Hermite [FEC18]. Gauss-type [CBMPP15, PM15]. Gauss4 [Wu16b]. Gaussian [AHBC15, BC16, BAHC17, DRST19, Jeb19, KB19, MIT15a, NP15, PBMJC18]. Gautschi [XDY16]. Gautschi-type [XDY16]. gear [XGH18]. GEM [GHL18]. GEM-CMM [GHL18]. General [DDP16, Din16, DD17, EHTL19, NBAK19, YZW15, ZHH^{+}19, AFH16, Abd16, BM18, BGII15a, BGMM18, BCS17, BBM15, CFR18, CJVW15, DLLB17, DMMSK17, DTKP17, Dzh18, FSNS16, GLSW18, Kaw17, LCH18, LGLC18, MT17a, MS17e, MT17c, Mol15, NMKK19, NHA17, PRRY18, RMVA17, VŽ19, WWH19, Xia17, YZPL18, YPD16, ZqZwZ15, dR16]. general-purpose [BBM15]. generalised [ADG17, CC15]. generalization [AY18, CLZS19, CNQRR17, KKV19]. Generalizations [CT16]. Generalized [FMC^{+}18, FCM19, GYL16, JRT17, JS17, KK19b, MIT15a, NJCS19, OAS19, Orr19, PSK16, PR15, ST16a, SPCSA16, SVV18, ZTGZ18, ZWT19, ZG19a, AQCJ17, AGR15a, AH16c, AG19, AFM19, BBGG19, BS15, BR16, BCDwK17, BMG^{+}19, BCVV18, CRGB17, CL19a, CL18b, CM18a, CT17a, CS18b, CRG19, CL19d, CL19c, CCRN19, CGN18, DF16a, DAM15b, DM17a, DRS16, DRST19, DG16b, DB15, FZ15, FS19b, GUB^{+}17, GLV15, GLLS16, Gun18, GYLQ16, HH18c, HCC18, HWXC19a, HWXC19b, IAAW18, JBT16, JSL15, Kam17, KK19a, KV15, LZW18, LLW18, LZ17b, LZ18c, LZ16c, LSBL18a, LSXL19, LOAGBR15, LYG15, ML17a, MT16a, MC18, Mus17, NM17, OK16, PKR18, PT17b, ROB17, RCN17, RCBG17, RT18b, SKC^{+}19, Sai16, SBU^{+}17, SEV^{+}18, SAH15, SPK18, SBC16, SLQ16, UGG^{+}19, WS17, WD19, WP16a, YL18c, ZLPF19, BV16b, BV16c, GLS15]. Generalized [GTC16, LS17b, MWY16b, PG16, RP15, SVL19]. generalized- [SEV^{+}18]. Generalizing [Con18, OCU17]. generated [DGM16b, MT18, WO19]. generating [Han19, ZJK16]. generation [ÁD17, SBP17, WKXT15, ZNG15]. generator [CAO^{+}19]. generators [Vig17]. generic [CDE18, TV18, SO16]. genetic [GLdCJS17, NF17, WZBM18]. geodesic [DS17b, YW15a, YW15b]. geomechanics [LWW17]. Geometric [CFJMVG16, FW18, KLM17, LA15a, MXW17, NCQ^{+}18, CPRS19, DWS^{+}18, DFGS16, Gru16, Lev18, MJ18b, MN19b, OAS19, PA19, QG17, QL18, RDRM15, RTV15, Sho18, WH16, ZTGZ18]. Geometrical [OCU17, OCU18]. geometrically [Sto18, Zag18]. geometries [KHS^{+}18, MTMZ19, VRBS19]. Geometry [ZSZ17, ADD^{+}16, VFFGV16, Xin16, ZLS^{+}15, Zha19b]. geotechnics [CH18]. Gerber [WXY17, ZS19]. Gersgorin [LZL16]. GFDM [UGG^{+}19]. ghost [VRBS19]. Gibbs [ARTY19, BCA18, DF16b]. Gilman [GKBG17]. GIMLS [DM17a]. given [HBH17]. Gl-LSMR [MT17c]. Glass [ESSE18]. Global [AC17a, BEJR17, CST^{+}16, LH18, MT17c, ZL17a, ZEBM17, AHJ17, Akh19, BEJR16, Con18, DGM16b, GC18, Jan17a, KFT17, LZW18, LH19, MTKO17, NF17, NRV18, Ou18, PDO19, SPG^{+}16a, SH18, SH19, WLZS18, WKBB17, YSW^{+}18, YWY19]. global-in-time [MTKO17]. Global-local [AC17a]. Globally [BTKM15, Bay15, DIJ17, LZ16c, ZWH15b]. glucose [DDG^{+}16]. GMAW [JL19b]. GMLS [Mir16]. GMRES [BRS15, CSB18, GJSC18, LH18, SHCJ19, xZWlC15, HLV^{+}15]. GMRES-Power [GJSC18]. GMsFEM [SVL19, CG19, CLP16, CLV16, CL16, MJ18a, VMM18]. Goal [BN15, CLP16, YL18a, Jan17a, LWZ19b]. Goal-based [YL18a]. Goal-oriented [BN15, CLP16, Jan17a]. goal-reaching [LWZ19b]. Goerisch [Vej18]. Golub [BEJR17]. Gompertz [HMKCR19, WS16, WH19]. Good [God15, CJT17]. goodness [CDN19]. Google [SvG15]. Gordon [WLW19, AMP18, Bac18, BBCK19, EGK17, KS17b, TN19, XDY16, YTSZ15]. Goursat [BGMS16]. governed [CG17b, GCY18].

GPU [DS15a]. GPUs [COdSR16, DF16a, RdSLL16]. graded [CXZ19]. Gradient [ABH19, JL17, SBCJ18, And16, And17, DdSL16, DSOBK16, DLHY15, DZM16, Fat16, FMMQ19, GC18, GM17a, GW19, JJ19, LZW18, LSDY19, LS18a, LL18, MDH19, MSZ18, MHMP16, NS19, Ou18, SCP19, YN18, YZ15c, YWY19, ZEBM17]. gradient-based [NS19]. gradient-like [SCP19]. Gradual [IAMCI18]. Gram [Elg19, Lu15b]. granular [KSD17, SODBK19]. Graph [PBJCM19, FFN^{+}17, FP16a, HLLY19]. Graph-based [PBJCM19]. graphics [PN16]. graphs [CGPR16, DEY19, PBMJC18, XZ18]. Grassi [OKGB19]. gravity [Sai16]. greater [LSW19]. greatest [BWS17]. Greedy [GH19b, KC19, ZLAA15]. Greek [KV18b]. Greeks [HL15, KC17]. green [FAVGC^{+}15, ACCF19, XD17, vV16]. Grey [WS18, ML17b, MLW19]. Grey-related [WS18]. grid [Arm18, AB18b, CZ17b, CY15, HSWH15, LC15b, LS18b, SRBD18, SPM15, TDA18, ZCH18, ZH15]. grids [AL18, BG15, GDCK19, KD17, KK19c, MHV17, MPH18, NSYX17, PR16, PDO19, ZCL16]. Gronwall [AAM18]. Grossberg [LJ15]. ground [KLS19]. groundwater [YBC^{+}16]. group [DLW16b]. growing [ZJ19]. growth [ACLM18, HMD19, NGGFM16, ZHC19]. GSS [HH18d, SBP17]. guaranteed [FJD17, ZY19a]. guess [KLC17]. Gumbel [DG16b]. gust [SH15].

H [SNY19]. H-spectral [SNY19]. H7 [BYWPA19]. Haar [BM17b, SK16, SGG19]. Hadamard [AAKT19, AAKT19, LWZ17, MA19b, WPA^{+}18]. Hagstrom [AMP18]. Hagstrom-Warburton [AMP18]. Halanay [LWTQ17]. Halanay-type [LWTQ17]. half [GH16a, HG19, LWZ^{+}15, NDN19, Tun18, WPA^{+}18]. half-line [NDN19, WPA^{+}18]. half-plane [GH16a, Tun18]. half-quadratic [LWZ^{+}15]. halfspace [DFT18]. Halley [ABK19, AGM15, CRS^{+}18, CSTU17, GD18, MA16]. Halley-type [AGM15]. Halton [HPT19]. Hamilton [LQ19]. Hamiltonian [BGS19, DZ18, LWL17, MPS19, NHA17, SW17c, ZZHS17]. Hammerstein [DNL16, AST19, DNL15, GAA16, HR15, MN17, MN19a, SE16, SR16a, WYLD19, XHZ17]. hand [ATG18, HHOT17, KH18, MLJ16, MT17c, REBF16, xZWlC15]. Handling [WSL17]. hanging [HHZ18]. Hankel [CQW16, Pul17]. happiness [SCMS18]. haptotaxis [Mer19]. hard [GH19a, HWP^{+}15, YSL17]. hardening [CS17]. Harmonic [JL15, Mej17, ALV19, AC17b, AFN17, GM16a, GH16b, NML18, Wan18a, YH18b]. Hartley [Mej17]. harvesting [Wan18h]. Hashin [CJP17]. Hausdorff [BPD15b]. having [ABK19, PCSR19, RIK16]. Hawkes [MX16]. hazard [BD19, DNK19]. HB [DP19b]. HCIR [Jeb19]. HDG [CHS^{+}18, Jin16, LX16a, SSZ19, ZC16b]. healing [BCS17]. heat [Abd18, CCJ16, CRCM16, FAVGC^{+}15, GRS19, KH18, MCLW16, PT17a, SSZ19, SVV18, VDS15, YSCY17]. heated [DLW16a]. heating [CS17, NGGL^{+}19]. Heaviside [WGHR17]. heavy [FF15]. heavy-tailed [FF15]. hedging [CP15, DKV16, DLS17, PZ18, XW18, ZYL18, dSBV16]. Heine [MB15b]. Heisenberg [DLW16b, GB16]. Hele [ZBL^{+}16]. Helmert [NS16]. Helmholtz [CDKN15, ACEJ18, AKM18, BBD16, CL19a, CQ17, CDKN14, DMS19, GT19, HMH18, JS19b, KR17, KPRA19, LZZL17, TKMT17, Wu17, YBLH17, YH18b]. Helsinki [SPG^{+}16a]. hemivariational [HH19a, XHH^{+}19]. hepatic [DDG^{+}16]. herd [Ahn15]. Hermite [AAKT19, BI17, BCD15, BCA18, BF18, CH19b, CH19c, DIP^{+}19, DD16a, DD17, FEC18, GS18a, GS19, HG18, JY19, LP15a, LAHB15, Lu15c, MN19b, MA19b, MS17b, RdV19, SRBD18, ZWGL18, ZwKN15]. Hermitian [Cro15, GLT16, GHC16, Hon19, LLW17, Miy19, SWL15, Tru19, WXZ19, Yin19]. heroin [LZX19]. Hessenberg [GHY^{+}18, MTT15, SB19]. Hessian [YZ15b]. Hestenes [DLHY15]. Heston [RZ19b, DI15, FM19a, GB18, HZ18b, MNFS18, dFG19]. heterogeneities [QDBG18]. Heterogeneous [GCY18, GK18a, AC17a, AEK19, ABB^{+}15, AB16, CL16, CVW17, CL19d, CDKN14, CDKN15, Don18, GHM18, GM16a, GC15, GT17, HCSC16, KLM17, LRRV18, LT18, MSV18, PG16, SRA^{+}15, ZCFW17, ZWW^{+}18]. heuristic [LZ18b, SW16a, SW16b]. hexagonal [VFFGV16]. HH [AY18]. Hierarchical [CDF^{+}16, BH17b, HM16, SA17, VFN18, WYLD19, ZDLD15, ZJK16]. High [AL19a, AMRX17, ACAC17, CBP18, CJLT15, DP17a, DM17b, DP19c, FP16b, GM16a, GKLS18, Gri19, GM17b, GTX^{+}19, HHEG17, JB18, LCL16, LSG18, LS18b, LZ19, MJS19, RSID16, TH19, Xin16, ZG19c, Abd16, AEKGP19, AGM15, AJY18, BM17a, BJ16, BS17c, CDF^{+}16, DF16a, DSA15, DGM19, Elg19, FR18b, GH19a, GK18a, GGC15, HMDSR19, JHPT15, JMSL19, KSDM16, LP15b, LSZ16, LLL19a, LHPZ18, MVQDdR^{+}18, MCGNBD17, MS19a, NZ15, OD17, PHCG16, RP16a, RP15, Reu17, RT18a, SEÖ16, SMTdF18, SJL16b, WLW19, YOK15, YC19, YFwC19, Zak19, ZST15, ZLS^{+}15, ZYL19]. high-contrast [PHCG16, RP15]. high-dimensional [LP15b, LHPZ18, SJL16b]. high-frequency [JHPT15, MCGNBD17]. High-order [ACAC17, DP17a, DM17b, DP19c, GM16a, Gri19, GM17b, GTX^{+}19, HHEG17, JB18, LCL16, LS18b, Abd16, AJY18, BJ16, DSA15, Elg19, FR18b, GGC15, KSDM16, MVQDdR^{+}18, Reu17, RT18a, WLW19, ZLS^{+}15]. High-Performance [CBP18, DF16a]. High-quality [LZ19]. high-rank [YFwC19]. Higher [MVdRE^{+}17, ACMT17b, BZ19b, HMT^{+}17, KV19b, KMB^{+}18, MTV17, MG15a, NGGL^{+}19, WWL15, Yam17, ZLZ17, ZQJY18]. higher-index [HMT^{+}17]. Higher-order [MVdRE^{+}17, ZLZ17]. Highly [BSGM19, CJSCT18, BBB17c, CL19a, CY18b, CL16, FYY17, GCTGC16, GCY18, GT17, HXX16, JGGS16, JS19b, KLM17, KCX17, KT15, LT18, LTY17, ML18, Mil15, MSV18, Mol17, uIZ15, ZX19, DLY18]. highly-accurate [CL19a]. Hilbert [AAA15, AA16, ABH19, AP18, BJM16, BD16, BHRR18, CDHP17, DK16, HS19, MV16a, OZ18, SAA18, Sak17, WGHR17, XM16]. Hilbert-Besov [ABH19]. Hildreth [JCC15]. Hilfer [HND18, Kam16a]. Hill [BBPS17, DDG^{+}16]. Hilliard [Che19b, CFWW19, CSE^{+}19, DF15, Son16, WCW19, XZ19, YZH18]. Hilliard-like [XZ19]. Hilliard-reaction [DF15]. historical [KV19b]. history [XHH^{+}19]. history-dependent [XHH^{+}19]. HIV [KJ19, MLV16, PC17, PC18a]. HIV/AIDS [KJ19]. HMM [CYY15a]. Hocking [Gru16]. hodograph [BLV19, FW18, KLK16b, LAHB15, LFCW19, ZWY16]. Hohenberg [SEV^{+}18]. Holder [Sai19]. hole [YSCY17]. hole-filling [YSCY17]. Holling [YT16]. Holm [YFS18]. Holomorphic [BHN15b]. homogeneous [CCNQ^{+}16, CNQRR18, JW17, MDVR19, QDBG18, SJRR17, VCL^{+}19, Wol17, ZW16]. Homogenization [CJCDLL15, CSW19, CJCDLL19, TV16a, VFGPG^{+}18]. homotopic [BDA17]. Homotopy [YMGA19, BC16, BZ19b, CHYZ19, EWR18, GGC15, HSAR18]. honor [DLY18]. Hopf [Gru16]. Hopfield [LDZ18]. horizon [BGNGR^{+}17, CPOG18, JS19a, Ugu18]. horn [GM16a]. host [TLLL19]. household [YL19b]. Householder [LWZZ16]. hp [LSP18]. HPC [GDC15a]. Hsieh [BI17]. HSS [RWW17, Yan17, ZZ15]. Huber [Bie17]. Hull [FNV17]. human [AMM19, BOF^{+}16, SCMS18]. humid [HWP^{+}15]. hump [Sad19]. humps [NS17]. Hurwicz [CHS17]. Hybrid [BLH^{+}17, MS17a, NF17, SBCJ18, AM16, ABB18, BSLL18, Beh16, CLX17, CCC19, DMMSK17, FYY17, FP16a, GC18, GJ17, HS17b, HCSC16, JK18, KMRS16, LLZ16, LRL^{+}18, LWDW18, Lu16, MAS19, NAK17, Ngo19a, NGGFM16, NGGFM17, OLRPdCJ19, QDG^{+}15, SGK16, SX18, WZBM18, WS16, XqYL^{+}15, ZSW16, ZCFW17, ZS17b]. hybridizable [DYZ18, KCK18]. Hybridized [Jeo15, MWY16a, SA17]. hydrodynamic [GRVC15, Gor16]. hydrostatic [LSG18]. Hyers [Xia17]. Hyper [PP15, WZ16, ZZW18]. hyper-exponential [WZ16]. hyper-plane [ZZW18]. Hyper-power [PP15]. Hyperbolic [SuSR16, BGMS16, BP17b, CTAT19, ELL^{+}19, HMDSR19, HWH19a, HWC16b, HW18b, JN18, KP15, LD19a, LD19b, Luo18, MDG19, MDVR19, MY18, OCU18, Par15, Rót15, SS17, WCW15b, YZH18]. hyperelastic [LY15b]. Hyperelliptic [ATV15]. hyperexponential [Hac18]. hyperfunction [OH18]. hypergeometric [AQCJ17, DP19b, DGSM15, SCG18]. hyperpower [SSG18]. hypersingular [NAM18, SS18c]. hypersurfaces [ADD^{+}16, BAAA^{+}15, LDDF18]. hypertension [TS17]. hyperthermia [RdSLL16].

i.i.d [OLPT16]. IB [CLZW19]. IBVPs [AFGT18]. ideal [JZS16]. identical [DD18a]. Identification [HBH17, JAN17b, LFCW19, SS19b, SS17, SZJ19, SYW19, VDS15, Abd18, KMD^{+}16, PT17a]. Identifying [AR19, HM19b, WSG^{+}19, ZW16]. identities [MVC19]. IDR [DSZ15]. IGA [CFSS18a, Mat18]. IGA-BEMs [CFSS18a]. II [BV16c, Gun18, HJ17b, LD19b, NAK17, YZPL18, YT16, ZG19a]. III [LCL16, AHKT15]. ILC [LDW18]. ill [BEJR17, DdSL16, HRR15, HJ17b, KR18, MRVV16, MP15, MHMP16, PB15, XZXQ18]. ill-conditioned [KR18]. ill-posed [BEJR17, DdSL16, HRR15, HJ17b, MRVV16, MP15, MHMP16, PB15, XZXQ18]. illustration [TDA18]. ILU [CMM17]. Image [WGHR17, BEJR16, CFR18, CG17a, DF16a, DGH15, DTB17, GDC15a, HS18a, HH19b, KKP^{+}19, LAVMA18, LLD18, LZL^{+}18a, LSL15, MDH19, PKMM19, SE18, SAH17]. imagery [CCRN19]. images [HLG^{+}17, NCQ^{+}18, PBMJC18, PBJCM19, SAH17, WXJ19]. imaginary [GKN18]. imaging [BTKM15, QLT^{+}17]. IMEX [BP17b, CDP17, Kad17, KS18, LLT16, MV16d, Son16]. Immersed [CCZZ16, LKS17, ACLY19, GKPS19, MZ19, YZ16, ZG19c]. immersed-isogeometric [GKPS19]. immune [BdSFL16, LRRV18]. Impact [ZS17c, CD15, PC18a]. Impacts [SPG^{+}16a, BG15]. IMPEC [CFS19]. impedance [AMM19, HHS^{+}15]. Implementation [CDE18, SLH15, GKPS19, HHKR15, KBG17, MMPGCSP17, SM16]. implementations [BW17]. implemented [VMR17]. Implicit [ABB17, ZJ19, AH16a, ALC19, ADD^{+}16, BKK18, BL16, CCW17, DBdLdM^{+}19, DLM17, HHR16, IDGH18, IÇ18, KW18, LZS15, MS18, MP18, Mil15, QDY^{+}19, SGL17, SAT15, SW17b, VL15, WD18, WZS18a, WKBB17, XZ19, YZZ19, ZS18a, ZST15, ZJ17b, ZDF^{+}17, GC15, LLL19a]. implicit-explicit [DBdLdM^{+}19, IDGH18, WZS18a]. implicitly [BZ16b, NML18]. implicitly-defined [BZ16b]. implied [OT17]. implosions [VG15]. Importance [RIK16, ABT17, Kaw17, MM17]. impression [DI15]. improve [ICMPG16, OGR17a, LGB18]. Improved [FE16, FPGK19, JJ19, MCGNBD17, NZL17, Ran15, Zho16, AS18, AJ17, BFLP15, CSTU17, FP16a, HX15, HZ19, LZS15, LOAGBR15, OGR19, SJ17, SKBT16, ZCC15]. Improvement [FFKP17, JL19b, OOO15, CCC19, CY18a, Wu16b]. Improvements [GHA^{+}15, AB18b, HB16]. Improving [AMS17, HVM18, dMPF15, CDH^{+}17, CNQRR17]. impulse [El 17]. impulses [LDW18]. impulsive [HND18, HCA^{+}18, LJ15, LDW17, ZSL15, Zha16, Zha18]. impurity [EGK17]. imputation [GLdCJS17]. in-situ [AAB18a]. inactivity [BO16]. incidence [ZM15a]. including [EAKS18, KR18, LWW17, OCU17]. inclusion [DTKP17, Fer15, LZL16, LLL19b, MMD15, TSMO17]. inclusions [BP17a, HND18, PAC15]. income [MLW19, dSBV16]. incomplete [BDGIPG15, LHJR15]. incompressible [BAK18, CHS15, CZ17b, CHS17, CLZW19, Hei15, KtTBK18, LW15, LW17b, ML19a, MV18, Ngo19a, PST18, WZZ16, Wat16, ZZZ18]. incorporating [DMCC^{+}18]. increasing [Nie15, LRKV16]. incremental [HRS16]. indefinite [HS17a, WL17b, dBAAA15]. independence [DM16b, JGAFAL19, ZJK16]. independent [GR19, NL17]. indeterminate [BHN15b]. Index [GSM18, Ahn15, CZ15, CY18a, Che19a, GYL16, HMT^{+}17, HW18a, JS17, Pis15, Rad17, SS18a, SCMS18, SJL16b, WL17c, WLC^{+}18, Wu18, WH19]. index- [JS17]. index-3 [Pis15]. indexing [GSM18]. indices [LN16]. indifference [LN16]. Induced [AvG16, AdGvG19]. Induction [CS17]. inductive [DS17b]. industrial [LN16]. industry [EPC17]. inequalities [AM17b, AÖ17, AAKT19, DMMSK17, MA19b, TH18, XHH^{+}19]. inequality [AAM18, CT17a, El 17, FE16, HH19a, LA15c, LWS16, LWTQ17, NS19, Qi19]. Inertial [TH18, CST^{+}16]. Inexact [AEP15, BBM15, HWH19b, LSH18a, MP17b, Pol16, RWW17, YSG18, YWY19]. inf [LT18]. inf-sup [LT18]. Infeasible [YZZL19, YLL15]. infeasible-interior-point [YLL15]. infection [LRRV18, MLV16]. Inference [AS19a, Gun18, PD19, Wan17a, Wan18e, Gun19, HG19, ZSW16, ZG19a, ZLPF19]. inferential [ABB18]. infiltration [LWV18]. Infinite [KNP^{+}17, BGNGR^{+}17, DHAM17, DFT18, FCJ16, HSAR18, Kan18, MTT15, SLW17, Ugu18, Zha18]. Infinite-dimensional [KNP^{+}17]. infinitely [MOY18, VL19]. inflated [QLZ19]. Influence [MPM^{+}17, Sai16, BOF^{+}16, CJP17, GR16, SLH15, SSCM18]. influenced [BYWPA19]. influenza [KPW17]. Information [WWST18, AY19, AdGvG19, MJ18b, MDH19, MP17b, RDRM15, WZS^{+}18b]. Infotaxis [EKKT17]. infrastructure [SPG^{+}16b]. ingredients [GDC15b]. inherent [Mär17]. inhomogeneous [AL19a, HMH18, JHK17, TNTL18, XD17]. INID [(Ba18]. initial [AAM18, CLX17, KP15, KLC17, LC15b, LWC17, PRP19, PTV17, RMVA17, RMA19, SK15, VARC17, vV16]. initial-boundary [vV16]. initial-value [RMVA17, VARC17]. initialization [LZS15, SF18]. Injecting [ICMPG16]. injection [SRRJ16]. Injectivity [ZZ16]. injectors [SJRR17]. inner [SJRR17]. innovations [FF15, HYF18, QLZ19]. Inout [GW17]. inpainting [WGHR17]. input [BJK19, BCNP15, Luo18, VZDV18]. input-output [VZDV18]. inputs [JL16b]. Inspired [PBS15]. instabilities [BLMS19, Rät15]. instability [Gru16, IGRL17]. insurance [DT18, MSY16, MSY17, NY19]. insurances [AO18]. insurer [HOTZ18, LRZ15, WRZ18, ZJQW18, ZC19]. integer [ZL17a]. integrable [MT16a]. Integral [QL18, Qi18, AFH16, AL16, AD18, BM18, BJM16, BH17a, Baz15, BG17, BP17a, BLM16, Bie17, CCJ16, CT16, DJC15, DDR19, DP17a, DNL15, DNL16, DHAM17, DL16a, DVS15, DLM16a, DYZ18, Don18, Elg19, EJJ15, EHTL19, FL23, FSNS16, GYS17, GD17, GAA16, GKN16, HHLS18, HR15, HSAR18, HKAR18, HXX16, HS17b, HSAA19, JYK17, JD16, KGES19, KMB15, KR16, LYG17, LS18a, LHPZ18, MFX15, MMO15, MS17a, MD18, MN17, MN19a, MII16, MA19b, ML19b, MT18, Men17, Mic18, MHL16, MH16a, MH16b, MS17c, MA18, MSH18, MAS19, Mit15b, MTS17, NDN19, NN18a, Nem15, NAM18, OD17, PSK16, PMN19, PAC15, PPM16, PA19, Pis15, PST18, Ros17, SE16, SAC17, SLH18, SR16a, SHS^{+}15, SDH17, SS18c, SJ17, SK16, uIuD18, WW16a, WPA^{+}18, WR16, WCS18, XHZ17]. integral [XY19, XF16, Yak19, YCX19, ZM15a, ZM15b, yZ17, ZX19, Zho16, ZN19, dAFvdMR19]. integral-algebraic [Pis15, ZN19]. integral-difference [MT18]. integrals [AM17b, AQCJ17, AAKT19, AL18, BEHL17, Ceb17, Fuk15, HS17a, HS18b, HX15, IGRL17, KK19a, KL15, Kan18, KW16, LWZ17, ML18, MIT15a, Mol17, Nak15, Orr19, PM15, uIZ15, SLQ16, VK16, Wan18b, ZuI17, ZLS^{+}15, ZCS^{+}19]. Integrated [HLLY19, Lev18]. integrating [AMCR19, CR17]. integration [Abd16, AL19a, ALMLM16, ACLM18, BBCP16, BHU18, BR15b, DKLP15, EK19, GG17a, God15, God16, GPHAPR15, GPHAPR17, GKLS18, Kan18, KS17b, KW18, KNP^{+}17, MG15a, MF16, NAH18, NV19, NN17, OH18, OGR17a, LGB18, RT18b, SS19a, TGSS17, VS18, Vis15, WCLL18]. integrations [LLCG16]. integrator [AMP18, JM18, MDG17b, SOMF19, XDY16, Zha17b]. integrators [BBPS17, BBC^{+}18, BBCK19, BM19a, JSW17, JSW18b, LWL17, WLM16]. integro [ALS19, BM17a, BM17b, BDA17, Beh16, BGL17, BS19, Bog15, BS15, CHR^{+}15, DBX19, DA19, DG16a, Dzh16, Dzh18, FCJ16, GMG19, GS15c, HSAA19, KP15, KS17a, KDW17b, KDW17a, KAA16, KPS17, Moh19, Mok15, Mol15, NE17, NSM16, Pis17, Reu16a, Sar19, SPK18, TJB17, WZ17, WL17a, YZZ19, YAA19, YJBM19, ZS18a]. integro-differential [ALS19, BM17a, BM17b, BDA17, Beh16, BGL17, BS19, BS15, CHR^{+}15, DBX19, DG16a, Dzh16, Dzh18, FCJ16, GMG19, GS15c, HSAA19, KP15, KS17a, KDW17b, KDW17a, KAA16, KPS17, Moh19, Mok15, Mol15, NE17, NSM16, Pis17, Reu16a, Sar19, SPK18, TJB17, WZ17, YZZ19, YAA19, YJBM19, ZS18a]. integro-parabolic [Bog15]. integro-partial [DA19]. integrodifferential [VDS15]. Intel [GGK18]. intelligence [KS17a]. intensity [DB15, PG19]. inter [GMSR18, GMSR21, PD19, RVZ15]. inter-claim [RVZ15]. inter-failure [GMSR18, GMSR21]. inter-record [PD19]. interaction [BBB17c, LY15b, MCGNBD17, VG15, ZHS18]. interactions [JC18, SH15, VS15b]. InterCriteria [RF18]. interest [BZ19a, GZH19, GVMR16, GVMR19, HZ18b, ID16, PX17, RZ19b, ZS17a, dFG19]. Interface [ZG19b, ACLY19, CCZZ16, CWXW17, CS18a, CFMN15, DYZ18, FLZ19, GSRJ18, JCC15, KS16, MZ19, Mu19a, Mu19b, SZ18, Sun19, TLLL19, WXW15, XY16, XZN^{+}19, YZ16, Zha19b, ZG19c]. interfaces [HBH17, NZ15, XD17]. interfacial [XZ19]. interior [BySZ19, CG17b, CGSS17, HK19, HL18b, HZ17c, JL15, KYY15, LLHO18, QC18, Qui15, RC19, SZ18, SM15b, TBY17, YLL15, YZZL19]. interior-point [YZZL19]. interlaced [God15]. interleaved [LHJR15]. intermittent [IABIPG17, IAMCI18]. internal [Pol16]. Internality [DRST19]. International [DNY17, WWST18]. interphase [QDBG18]. interpolants [BIR17a, BCD15, CDDD19, CH19b, ELL^{+}19, SST15]. Interpolating [CJ16, ABRS17, Bej16, DDR19, DAM15b, DM17a, ZMTT19]. Interpolation [BF18, CDHP17, GK19a, ABB17, Arà19, Bal17, BL18, BGII15a, BI17, BDIR19a, BDIR19b, BIJM^{+}19, BCA18, CFSS18a, Che18, CL15, DMP19, DD16a, DXML19, DD17, FSNS16, FP18, GS19, GH16b, HG18, Han19, Har18, HW18a, ID18, JZS16, KMLB17, KB16, LHPZ18, Lu15c, MPM^{+}17, MN19b, Moh17, NJCS19, NY19, QWZ18, RRS19, SJ17, SJ18, TXFH18, WCLL18, WWL15, ZwKN15]. interpolations [BLV19, ZYY19]. interpolatory [GMdV19]. interpretation [GSVSM16]. intersection [ADD^{+}16, BAAA^{+}15, DD16b, DD16c, FWW^{+}18]. intersections [BHN15a]. interval [AG18, AHL17, BSLL19, FE16, HKZZ15, Kan18, KS15a, LaT15, LaTG17, La19, LXL15, LXL16, LWZ17, MHC16, Wan15b, bWzZjH16, WL17c, WLC^{+}18, Wu18, WH19, Zha18]. intervals [FFP16, GMdV19, LLL19b, Moh17, YS19]. intimate [LE18]. Introduction [ZDL16, ZDFF18]. intrusion [ZYZ18]. intrusive [CH18]. invariance [ADG19, Rom18]. Invariant [BGI18, FwC17, Fer15, GGBTP17, VFN18, YZH18]. invariants [AK17a, FFN^{+}17]. invasion [Mer19, SN19]. inventory [EPC17]. Inverse [BW17, HVV15, LHJ16, MHC16, Su16, WW16b, Aks17, Bay15, BFLP15, BKLR15, BKLR16, DNK19, DFRR15, FLJT18, FM19b, GHA^{+}15, GLT16, Haj19, JO17, LRT17, LH16, LSZ16, LLZ18, LCZ18, Liu17, Liu18, LF19, ML17a, MM19, MP15, NAK17, RVC19, SNY19, VF17, WLZ15, WWH19, WD16, XCG19, YB15, ZWGL18]. inverses [HK18, IAAW18, ML17a, MTT15, MT15, PP15, PKR18, SS16b, SAH15, SSH16]. Inversion [AGR^{+}15b, FM18a, HHS^{+}15, JYK16, NPR18]. invert [HLLY19, JL15]. Investment [TF18, ALS18, GWY17, HYZ16, HOTZ18, LRZ15, MYG18, Mud19, Nke18a, RP17, WRZ18, YX19, YL19b, ZC19, dFG17]. Investment-consumption [TF18]. investments [FY18]. Involutions [AH16b]. Involutive [RM19]. involving [AQCJ17, AKR18, BLP17, Ceb17, CGN18, DVA19, KK19a, PM15, SCMS18]. ion [ABL^{+}19, FFKP17, VMM18]. IPDG [CYY15a, CGPR17]. IPG [YBLH17]. irreducible [CHYZ19, SNY19]. irregular [JK19, Prz15, RTV15, Xin16]. isentropic [BHN15a]. iSIRA [HLLY19]. island [FAVGC^{+}15, NC16]. island-based [NC16]. iso [SL18]. iso-parametric [SL18]. isochronous [AIL17, GLV15, IL15, LI15]. Isogeometric [QDY^{+}19, QDBG18, CDP19, CPRS19, GKPS19, GDQ17, GD17, KV18a, KV19a, SM15a, XLS^{+}19]. isospectral [Kau16]. isosurface [DS15a]. isothermal [BBPR15, OSC15]. isotopisms [FFN^{+}17]. isotropic [LWZ^{+}15]. Issue [DLY18, VAMDGR19]. issues [AB18a]. Istanbul [DI15]. Itô [AK16, BJ16, HMD19, MSH18, MAS19, SLH18]. Itô-Volterra [MSH18]. Italy [CPRS19]. iterated [KR16]. iterates [Bog15]. Iteration [FLM17, AM15a, BL16, CT15, EL19, GXZ15, HH18d, HWXC19a, HWXC19b, LY18, Mar19b, MM15a, MV16a, NH19, PLP17, RWW17, SNY19, SH19, Wan15a, WZL17, WWH19, WXZ19, WHS17, WL16, Xu15, YZW15, Yan17, YSG18, YZ18, ZZ19, ZZZ15, ZZ15]. Iterations [GLLS16, KMB^{+}18, SSH16, SSG18]. Iterative [CQW17, CW17, CPL^{+}19a, DG16a, HM19b, KKW16, LWW17, LDL^{+}19, Sak17, ZZZ18, Ais17a, ABHVM16, AAB^{+}18b, ACDT18, AP18, BCT16, BH17a, BBH^{+}17, BW16, BSGM19, BBC17, CCTV17, CFS19, CCT19, CJT15, CMQT16, CJT17, CMTV17, CT17b, CGT18, CJSCT18, DTKP17, DLM16b, DD17, EL19, GHM17, GD18, GFPGF19, GM16b, Had15, HH18a, HVMR19, HSAA19, HMT15, JMNB18, LLD18, LHZ15, LD15, LDW17, LWC17, LLG^{+}18, LJ16, Mic18, PKR18, PI19, RMK19, SLH18, SR16c, SH18, TBR16, TP19, WA17, Xia17, XY16, Xie19, XC18, YSL17, ZCST18, ZPL17, ZPL18, yZLX15, ZW16, ZTGZ18, Zho16]. iteratively [LH16, LSH18a]. IVPs [FR18b, Goć15, PS15, SVAKGR19a].

J [ÁD15, AB16, Beh16, BP17a, BKMO20, Chr19, CDKN15, CPL^{+}19b, DNL16, DLY18, GMSR21, GAA21, JSW18b, MHY21, LGB18, RdSMO18, SW16a, TGJ19]. J. [DLY18, ZPL18]. Jacobi [GMG19, BDSB16, HMP15, JL15, LQ19, LX17, MMMMB16, Mia18, SAC17, SB18, Su16, Zak19]. Jacobian [ACDT18, TGSS17, WYW15]. Jacobian-free [ACDT18]. Jacobian-vector [TGSS17]. Joint [MM17, CLPZ17, EG17, KC19, PW15a, SGK16]. Jordan [ML17a]. Joule [CS17]. judgments [BDGIPG15]. jump [ALS18, AHT18, AG18, APYW19, Dan17, DP19c, FEC18, GVMR16, GVHMR17, GVMR19, Gri19, HK19, HK07, HYZ16, JSW18a, KKW16, MX16, MS15a, MFLL18, Mud19, Nke18b, Prz19, RP16a, Sta19, Sun19, TGJ19, TYMG17, WC17, WRZ18, WW16b, WZ16, XY15, YYL15]. jump-diffusion [ALS18, Dan17, GVMR16, GVHMR17, GVMR19, HK07, HYZ16, KKW16, MS15a, MFLL18, Nke18b, Prz19, TGJ19, XY15]. jump-SDEs [Gri19]. jump-to-default [JSW18a]. jumping [AYSE19]. jumps [CXW19, DFLM19, DL16b, GKLS18, HOM17, MYM16, ST16b, ZS17a, ZC19].

Kahan [BEJR17]. Kalman [DLM16b, KK17]. Kantorovich [AM15a, AH16c]. Kawahara [MM18]. KdV [WD18, BMG^{+}19, LSXL19, WD19]. Kelvin [PBP16]. Kernel [OD17, AAA15, AA16, AVB17, AAM18, AP18, BM18, BJM16, BH17a, BDA17, BLM16, CXZ19, CDHP17, DJC15, DVS15, HSAA19, JW17, KMB15, KV15, LW17a, LF19, MMO15, ML19b, Moh19, NGGFM17, PMN19, SAA18, Sak17, VFG18, VDS15, WGHR17, WR16, XZN^{+}19, YCX19, ZCS^{+}19]. kernels [Bie17, DL16a, GKN16, HXX16, Mol15, MTS17, NSM16, NAM18, SDH17, SS17, ZX19]. kind [BM18, BH17a, BLM16, DJC15, DLM16a, DRST19, GV17, Men17, Mic18, MTS17, NSM16, NAM18, PPM16, SK16, ZM15b, dR16]. kinds [YH18b]. kinetic [MP18, RKS16]. King [AR15c, CCTV17, MBM17]. kinks [EGK17, GKLS18]. Kirchhoff [MM15b]. Klein [KS17b, AMP18, BBCK19, TN19, WLW19, XDY16, YTSZ15]. Klein/Sine [YTSZ15]. Klein/Sine-Gordon [YTSZ15]. Knapsack [PBS15]. Knights [GGK18]. knock [LDK17]. knock-out [LDK17]. knot [AHBC15, BAHC17]. knowledge [ICMPG16]. Koiter [BEMA16]. Kolmogorov [MDG19, Wat16]. König [LP18, PP18a]. König-like [PP18a]. Korteweg [AKHSU18, BGS19, FS19b]. Kou [CGT18]. Krawczyk [HZ19]. Krawtchouk [AGR^{+}15b, RCPA15]. Krízek [KKV19]. Krylov [ABG17, Bai15, Bot16, BHU18, BL16, CGvA^{+}15, GG17a, KBG17, Mia18, REBF16]. Kumaraswamy [Wan17a, Wan18e]. Kuramoto [MFWK18]. kurtosis [BOF^{+}16]. Kutta [AK15, ZM15b, AH15, AK17a, BYW18, BP17b, CWH18, DT16, GPHAPR15, GPHAPR17, GQM16, HHR16, HIS16, KS18, LL15, MVK19, MAF17, Son16, SV16a, TFS17, WZS18a, ZST15, Zha16, ZZHS17, ZDF^{+}17].

L [DB15]. L-shaped [DB15]. L1 [THC^{+}18]. lag [Sai16, SVAKGR19b]. lagging [CRCM16]. Lagrange [AB18a, DLM^{+}19, Har18, KBL19, LJ15, WWM17]. Lagrangian [BSS15, CH17a, FP16b, Guo17, HWH19b, PKMM19, SK18]. Laguerre [AQCJ17, LZ15, ST16a, ZWGL18, ZS19, ZY19a]. laminar [LLSZ17, MFWK18]. Lanczos [BEJR16, GLZ18, Haj19, LH19, NML18, WGLV16]. Landau [Wol17]. Landing [GGK18]. Lane [RMA19, SGG19, SK15]. Laplace [CS18a, DB15, Dos19, FM18a, JMSL19, JYK16, MVC19, OM17b]. Laplace-interface [CS18a]. Laplacian [CZ17a, CKL18, DWXO15, HLLS18, HLLY19, VDT19, Yao18]. Large [wCW15a, YFwC19, And16, AvG16, AMP15, Bai15, BCC^{+}18, BEJR17, DSN17, DY15, FM19b, GH19a, HLG^{+}17, HJ17b, HWZ18b, JK19, KSD17, LSH^{+}18b, LF18, LL18, Per16, SM16, WLLC18, WP16a, YZ15c, ZDL16, yZyMxG^{+}16]. large-body [AMP15]. Large-scale [wCW15a, YFwC19, And16, DSN17, DY15, HLG^{+}17, HWZ18b, KSD17, LSH^{+}18b, LF18, SM16, WLLC18, WP16a, YZ15c, ZDL16]. largest [CHYZ19, CDR18]. latency [PC17]. Lattice [COdSR16, DKLP15, EK19, God15, Uda18]. lattices [CMRP17]. Lavrentiev [XC18]. law [CJCDLL19, CS17, CS18b]. laws [BSS15, BM19a, BP17b, BMG^{+}19, CRGB17, GAA19, GAA21, HWC16b, HW18b, JN18, Kaw17, LSXL19, LV16b, RCBG19]. Lax [LSZ16]. Layer [MT16b, ABT16, DBdLdM^{+}19, LLHO18, LZRH19, PGRC18a, Qui15, SAA18, SBU^{+}17, ZFK18]. layered [YL18b]. layers [BR15a, CGSS17, Pol16, RC19]. layout [JCC15]. lead [DMS19]. leading [GW19, MJ18b]. leap [RKV19]. Leapfrog [JKL^{+}15, LX16b]. Learning [CFMÖ19, TW17, HYQ^{+}18, LWLL18, LDW17, NGGA^{+}18, NGGL^{+}19]. Least [HMT^{+}17, AHMY19, AS18, CQ17, CSXY19, DF15, DDM15, DM17a, Dia18, DLM16b, HLV^{+}15, HHS^{+}15, JL16a, JL17, Ku16, LC15a, LHK19, Lin16, LWJ15, LLG^{+}18, Mir15, MJS19, Nie15, PLP17, SY19, TW17, WS18, XCG19, YW16, Yua15, ZPL17, ZPL18, dBGMUA18]. Least-squares [HMT^{+}17, AHMY19, DDM15, Ku16, LC15a, Lin16, MJS19, Nie15, XCG19, Yua15]. Lebesgue [CH19b, MS19b]. Leffler [AAM18, Hui16, IAAW18]. left [LSS19]. leg [MSH15, WZ17, WL17a]. Legendre [DNL16, AST19, Che19b, DNL15, EBJS19, HC17, HNT19, MN17, MN19a, MC18, Moh19, Mok15, NN18a, Nem15, PMN19, PPM16, SR16b, SM18, YTSZ15, YJBM19, yZ17]. Lehmann [Vej18]. Leland [KV15]. lemma [ZS18a]. lemmas [Xia19]. length [MVQDdR^{+}18]. lengths [LDW17]. Leray [XWWI18]. level [AB18b, CMM17, CZLV17, CGvA^{+}15, Dan17, GdLT17, GUB15, HWZ18a, HHZ18, KC17, LWW17, LX15a, LA15c, LCZ18, LZS15, Ngo19a, QM15, WD18, YZPL18, vSV15]. level-set [LWW17]. levels [WXY17]. Levenberg [AR15a, AR16, IIA15, LM18]. Levin [ML18, Mol17]. Levy [RdSMO18, NP19, WH16, BZ19a, CCY18, DLS17, DS17a, DYD19, FCJ16, Hac18, Hai17, HXC19, JA17, RdSMO17, YC19, ZS19]. Li [FFKP17, VMM18]. Li-ion [FFKP17, VMM18]. liability [DT18, PX17]. Liapunov [PB15]. library [GDC15a]. lid [ABT16, LW15]. lid-driven [LW15]. Lie [CNT17, GKBG17, YX19]. Lienard [KAS18]. life [AALMLM18, BO16, GAB17, Kav16, LSZB17, OLRPdCJ19, Wan18c, ZSW16]. life-span [AALMLM18]. lifetime [WL17c, WLC^{+}18, Wu18, WH19]. lifting [NLG15]. light [VS15b]. light-matter [VS15b]. like [AMB19, AG19, BDG15, DD16c, Fan15, FR18a, GKM18, GxLlW15, HKAR18, KV18a, KLC17, LZ16a, PR16, PP18a, Skr17, SCP19, XZ19, ZZ19, ZCD17, ZHL15]. likelihood [GYQS15, LC16, RABB19, YS19, ZLPF19]. likelihood-based [ZLPF19]. Limit [LA15b, LI15, dFLM18, AMP15, BL17, CFG^{+}19, HRS16, KS17b, LSLZ17, NTPVNTDT15, Su18, SM15b]. limited [KV16, RKB16, VL19, ZZG^{+}19]. limited-memory [VL19, ZZG^{+}19]. limiting [CFSS18b, FCM19]. Lindley [COGD19, DNK19, NZ16, PD19]. Lindsay [ACCF19]. line [AR16, CD16b, Fan15, GB16, HWZ18b, JJ19, NDN19, WPA^{+}18, XqYL^{+}15, YWY19]. lineal [MGMGJLRB17]. Linear [AH16a, BLV19, BV16b, Mar18, YZH18, AFH16, Abd16, AHP19, ATG18, BBCP16, BBC^{+}18, Bai15, BCFN16, BGNGR^{+}17, BFLP15, Baz15, BWS17, BCVV18, CFR18, CJVW15, CCJ18, CGMM18, CZ17b, CLW17, CM18a, CL19c, CNQRR18, CN16, DLLB17, Das18, DdSL16, DG16a, Dia18, DIJ17, DGM16a, DSZ15, Dzh16, Dzh18, Elg19, FE16, FLN16, FGF19, GWLL18, GLMT16, GUB^{+}17, GH19b, GKPS19, GFPGF19, GHC16, GHY^{+}18, GH16c, GYQS15, GYLQ16, GSRJ18, HZ18a, HMT^{+}17, HOS16, HMMV17, HLLS18, HZ17a, HZ17b, Hua19, HWXC19a, HWXC19b, JMN^{+}16, JSL15, JZ18, JK19, JGTB15, JD16, Kad17, Kha17, KV17, KV18b, KPT15, KR18, KBL19, Lad15, LaT15, LaTG17, La19, LWXF18, LMUZ18, LYG17, LHZ15, LZ16c, Liu16, LHPZ18, LLV18b, LLY19, LM17b, LYG15, MDG17b, MT17b, MT17a, MP17a, ML19b, MLJ16, MB17, MS17c, MA18, Miy15, MSZ18]. linear [MT17c, MHC16, MSH15, MY18, NHA17, NW19, OD17, OK16, OAS19, Par15, PB15, PDB15, PBGG17, PHD17, PHCG16, Pul17, QWZ16, RWW17, REBF16, Reu16a, SAC17, SHW19, SW17c, Tho16, TBAH16, TBY17, TN19, UGG^{+}19, Vis15, VX19, WZL17, WLZ17, WD18, WXZ19, WA17, Wu16a, WL16, WL17b, YW16, Yan17, Yi19, ZZ19, ZSL15, yZLX15, Zha16, ZZZ15, ZL15b, xZWlC15, ZS17b, ZN19, ZJK16, dAFvdMR19, dBGMUA18, dFLM18]. linearity [LO19]. Linearization [JC15, JMS17, CCNQ^{+}16, RNPK15, SHW19]. Linearized [HA19, RZ19a, BHN15a, CLZS19, MDG19]. linearly [ÁNPPCSA15, SGL17, YZZ19, ZST15]. lines [BK19c, DM17a, WHWY18]. link [SvG15]. linked [PZ18, ZY19a]. Liouville [AÖ17, CKB17, JLLZ16, MR19, SF18]. lipid [Hei15]. Lipschitz [AH16c, BS17b, CXW19, DBX19, DIJ17, EHVM18, HMD19, IÇ18, KB17, KGMS18, Yao16, ZWH15b]. liquid [AK17b, BG15, MFWK18, RANG15]. liquidity [CPOG18, MV16d, PX17]. Liu [Wu16a]. liver [KHS^{+}18]. living [SR16b]. Ljusternik [Yao18]. LN [AH18]. load [HRS16]. loads [SV16b]. Lobatto [CBMPP15, MVPN15]. Lobatto-type [MVPN15]. Local [AGM15, AK17b, DLW16b, HM19a, JHPT15, JC15, JMS17, Jin16, Ku16, SNY19, AK18, AGM19, AC17a, AG16b, AAB^{+}18b, AR19, AHKT15, AHK19, Bac18, BT17, BM19a, CYY^{+}16, CVW17, DXML19, Fer15, FP16a, GJ17, GTX^{+}19, HVMR19, JM15, KB17, KSDM16, KtTBK18, LAVMA18, LLCG16, MA16, MKKV19, MFLL18, PBMJC18, ST16b, TX19, VX19, WZS^{+}18b, WLZS18, WZS18a, WKBB17, YGS18, YH18a, ZNG15, ZC16a, ZC18, VCL^{+}19]. local/global [WLZS18]. localization [PP16]. localizations [RW15]. localized [LaT15, LaTG17, La19, LWV18]. locally [AS19a, BK15c, CRR16, DGM19, DLM17, HMD19, SSKV17, Yao16, ZJ17a, ZWW^{+}18, Zha19a]. location [FAVGC^{+}15, Gun19, Qui15]. loci [TBS18]. locking [ML19a, WWWZ16]. locking-free [ML19a, WWWZ16]. LOD [BK15c]. Log [Orr19, LX16c]. log-aesthetic [LX16c].

Log-sine [Orr19]. Logarithmic [AV16, AT17, AD18, HS18b, KL15, MB17, Qi18, QL19, WCLL18]. logistic [FLMS16, HG19, PCJ17]. Lomax [ABB18, CAO^{+}19]. London [SPG^{+}16a]. Long [GGK18, Ngo19a, BBB17c, BGNGR^{+}17, CP15, DAM15b, DMAS^{+}16, NV19, RFLM16, WSL19]. long-bone [DMAS^{+}16]. long-short [BBB17c]. long-term [BGNGR^{+}17]. Long-time [GGK18]. long-wave [DAM15b]. Longest [KPS16]. Longest-edge [KPS16]. longitudinal [HZ17b]. Lookback [FM19a, JHK17]. loop [NDH16]. loss [ALS18, DT18, KD19, LE18, LN16, SJL16b]. loss-based [SJL16b]. losses [NCJ16]. lot [SGK16]. Love [LS18a]. Low [ACDT18, AB18b, CT15, LSH18a, SAC17, Arm18, BLXD18, BBD16, BDF15, BL16, CCY18, CGMM18, CPZ19, DY15, GH17a, GH19b, HS18a, HLZ19, MOPP18, MS19b, SRRJ16, SHC^{+}19, SZ18, WGLV16]. Low-complexity [ACDT18, CT15]. low-dimensional [CPZ19]. low-discrepancy [BDF15]. low-frequency [BBD16, CCY18]. low-order [Arm18]. Low-rank [AB18b, BLXD18, BL16, CGMM18, GH17a, GH19b, HS18a, MOPP18, SHC^{+}19]. Lower [LSLZ17, CBC^{+}19, Tru19, Vej18, Wan15b, WP17, ZZG^{+}19]. lowest [Yi19]. lowest-order [Yi19]. loyalty [GR16]. LROM [SMS18]. LS [HLBP15]. LSMR [MT17c, MT17c]. LU [LZ18c]. Lupas [HWC16a, KL17, LJX17]. Lyapunov [MHY21, AÖ17, CT17a, MHY17, OSMZ18, PS18, SV16a].

M [DLY18]. MacCormack [Ngo19b]. Mach [IDGH18]. Mach-sensitive [IDGH18]. machine [NGGA^{+}18, NGGL^{+}19, WS18]. machines [RRW17]. Mackey [ESSE18]. Maclaurin [Dub16, RT18b]. macro [BK19a, GK19a, KB19]. macro-elements [GK19a]. macro-triangles [BK19a, KB19]. macroscopic [BM19b]. Madrid [SPG^{+}16a]. magnetic [COP17, CS17, QLT^{+}17]. magnetized [LHY16]. magneto [Sai16]. magneto-thermoelastic [Sai16]. magnetohydrodynamic [KLC17, LHK19]. magnetohydrodynamics [AMRX17, AZ17]. magnetostatic [MTCC15]. Mahony [DAM15b, Güc17]. majorant [AH16c, Fer15]. majorization [KCNH16]. makes [JK19]. making [BDGIPG15]. Malliavin [Yam17]. malware [MVdRE^{+}17]. management [AVGCMVM15, CY18a, Nke18a, Nke18b, PX17]. Mandelbrot [KPP19]. Manifold [DS17b, CST^{+}16, ZNS18]. Manifold-valued [DS17b]. manifolds [CDF^{+}16, CRS^{+}18, KMLB17]. Mann [MM15a, Wan15a]. mantle [PDO19]. manufacturing [CCC19]. many [QG17, REBF16, VL19]. map [ACEJ18, Bej16, DLMZ19, GKM18, OKGB19, SE18, LZY17]. Mapped [HW18b]. mapping [DS19, XLS^{+}19]. mappings [BP17a, DTKP17, Iid15, KLK16b, PAC15, Ugu18]. Marching [LLZ18, vtWvdHvdVV16]. marching-on-in-time [vtWvdHvdVV16]. marginal [Ahn15]. mark [DV15]. market [CPOG18, DSDV15, RdSMO17, RdSMO18, ZS17c]. markets [ST16b, YL18d]. Markov [BZ19a, CH19a, CNQRR17, DS17a, Ery16, LY15a, RP17, Ugu17, YCW17, ZS17c]. Markov-modulated [BZ19a, DS17a, RP17]. Markov-switching [ZS17c]. Markovian [GMSR21, AG18, APYW19, AYSE19, CXW19, GMSR18, OM17a, WW19b, YYL15]. Markovian-jumping [AYSE19]. Marotto [SE18]. Marquardt [AR15a, AR16, IIA15, LM18]. Marsaglia [Vig17]. Marshall [BO16, DGM16b, NKD^{+}19, OK16, OE19]. Martensen [DS15b]. martingale [DSDV15, NP19]. Maruyama [CW15a, GQM16, HLM18, KHLY17, LYG17, Mao15, Mao16, Mil16, OM17a, ZSL18]. mask [XMW^{+}15]. masked [LSZB17]. mass [BBCK19, CFS19, DKP^{+}19, GMCD15, PG16, YB17, ZL17c]. mass-conservative [CFS19]. mass-preserving [ZL17c]. Master [BCMP16]. match [LAHB15]. Matched [WXW15, FLZ19, SBU^{+}17]. matching [DFT18, SODBK19, WYLD19]. material [HBH17, KSD17, NZ15, SODBK19, YOK15]. materials [Don18, DCNY18, HWP^{+}15, MCLW16, YSCY17]. Math [ÁD15, AB16, Beh16, BP17a, BKMO20, Chr19, CDKN15, CPL^{+}19b, DNL16, GMSR21, GAA21, JSW18b, MHY21, LGB18, RdSMO18, SW16a, TGJ19]. Math. [ZPL18]. Mathematical [ABL^{+}19, DNY17, JTCC18, KDB18, KPW17, LE18, LHY16, MMPGCSP17, TJCC16, TJCC17, Ver16, VACR15, CCIL18, CCD16, DJAMJ17, DMAS^{+}16, FZ15, MVC19, SH15, TLE19, ZqZwZ15]. matrices [ADG19, APS15a, APS15b, APS17b, APS17a, APS19a, APS19b, AvG16, AK17c, AB18b, BCC^{+}18, BDG15, BB16, CDR18, Elo19, FV17, FM19b, GWLL18, GLLS16, HWZ19, Hon19, HK18, Hua19, HLLY19, JMN^{+}16, LWZZ16, LY18, Li18, LLW17, MM19, Mar19a, MTT15, MT15, MS17c, OV17, SS18a, SY18, Su16, XCG19, ZWGL18, ZCD17, dBAAA15]. Matrix [BL16, BCCB15, CM19b, DFS15, SVGM17, AIS^{+}17b, BBPS17, BWS17, CZ19a, CCJ18, CNT17, CSTU17, CT17b, CPL^{+}18, CPL^{+}19b, DLLB17, DIS^{+}18, DIP^{+}19, DMR19, DY15, Don18, DI15, FE16, GLT16, GXZ15, GYLQ16, Had15, HZ18a, HY19a, HCC18, JZ18, KLK16a, LDZ17, LY18, LH18, Li18, LH19, LWJ15, LHZ15, LLW17, LS18b, Lu15b, MT16a, MT17a, MOPP18, MK17, Miy18, MT19, NE17, NS17, NS16, OV17, OOO15, PS18, RSID16, SS16a, Sad19, SB19, SvG15, SS19a, SIAJ^{+}19, SWL15, SLH15, STS15, SC18, SSH16, TGSS17, Tru19, WLZ15, WNML16, WHS17, WL16, YN18, YBD16, Yua15, ZST15, ZS18b, ZCD17]. Matrix-free [BL16, TGSS17]. matter [VS15b]. max [CLM19, LaT15, LaTG17]. max-plus [LaTG17]. max-product [CLM19]. max-separable [LaT15]. maximal [AKR18, WZS^{+}18b]. maximally [GW19]. maximization [Sad19]. Maximizing [LWZ19b]. Maximum [RABB19, Sun16, AK17c, BFLP15, GYQS15, JD16, KKV19, LC17b, LC16, MYG18, MS15a, WZ18]. maxterm [ADMV19]. Maxwell [GH17b, ABBB17, Bot16, CMHHS17, CS18b, DLM^{+}19, JB18, KS17b, LHLL18, NCPD19, SLS18, Wan18a, YH18b, ZCFW17]. MCC [LRX16]. MCGM [Abd18]. MDS [VOB16]. Mead [SGK16]. Mean [BO16, CW15b, CJV17, LZ16b, LDZ18, WW19b, ZB18, ABT17, AH16a, BE19, BCVV18, CCJ18, DMMSK17, DSOBK16, GQM16, GWT18, GUB15, HOM17, ID16, Kav16, LRZ15, LC16, MS15a, QG17, QL18, SMTdF18, ZWH15a]. mean-field [MS15a]. Mean-risk [ZB18]. mean-square [ABT17, AH16a, GQM16]. mean-VaR [BE19]. Mean-variance [CW15b, WW19b, LRZ15]. meander [PR16]. meander-like [PR16]. means [FFN^{+}17, GLdCJS17, MGMGJLRB17, NGGL^{+}19, Slo15]. Measure [Rom18, DSDV15, HSAR18, LSL18, MJ18b, NP19, Ros17, Sai19]. Measure-invariance [Rom18]. measurement [CS18b, Liu16, SS19b, VFN18]. measurements [BTKM15, CL18a, Slo15, WS19]. measures [BYWPA19, CPOG18, DRST19, FJD17, JR15, Man15, MM17, SPCSA16, SM19b, WP17]. mechanical [DCNY18, KTT16, SV16b]. Mechanics [CBP18, Chu15, XHH^{+}19, Yam19, ZX16]. mechanism [ESAE18]. mechanisms [CMRP17, SS18d]. media [AAB18a, AEKGP19, AdQdS16, AFM19, CYY15a, CM19a, CFS19, CSW19, CSS19, CL16, DLM17, DFT18, DVJ18, ES18, EAKS18, ESI^{+}19, GM16a, GT17, HCSC16, KLM17, LT18, LRX16, LWV18, LHLL18, LCX18, MDVR19, PSS17, PG16, RNPK15, RP15, SRA^{+}15, VCC^{+}19, WAGVRB19, XD17, ZWW^{+}18]. medial [CM18b]. medical [AFdOdS15, NCQ^{+}18]. medium [BFLP15, CJCDLL19, CV19, LMM19, Sai16, YL18b]. Meeting [LPW^{+}19]. Mehler [MB15b]. Mehrotra [YLL15]. Mehrotra-type [YLL15]. Meixner [MB15b]. Mellin [DL16a]. melts [AM16]. member [YBD16]. members [CGT18]. membrane [ABL^{+}19]. membranes [Hei15]. memories [ARCA^{+}19]. memory [CCTV17, CCT19, CJT17, FdOP17, HL15, LRL^{+}18, NBAK19, RKB16, SS17, VDS15, VL19, Wan18g, ZZG^{+}19]. memoryless [And17, YN18]. mental [XDZZ19]. Menten [TK16]. Merging [GW16, HM16, KHS^{+}18, Lu15a, LJ16]. Merton [JA17, WC17]. Merton-type [JA17]. Mesh [JWQ18, AN17a, AN19, BS18, BI17, BCS17, BBG19b, BZ16a, DM16b, GLTW15, KHS^{+}18, LBL15, LW15, LL16a, LW17b, MZ16, MS15b, SN19, WKXT15, ZNG15]. mesh-independence [DM16b]. meshes [BDIR19a, BR16, BLMS19, CDF^{+}16, CL19a, Che16a, CXZ19, CDE18, Das15, HWZ18a, HH16, LW15, LRKV16, LGLC18, ML16, Mu19a, Mu19b, SLS18, SM15b, VFFGV16, WL19b, ZDLD15, ZKRL15, ZML16]. meshfree [SK18, TKH16]. meshing [SL18]. Meshless [uIuD18, WW16a, AD18, DAM15a, FSNS16, ID18, SJ17, SJ18]. metabolism [DDG^{+}16]. metaheuristic [RF18]. metaheuristics [AGM19, SQRE16]. metamaterial [TW19]. metamaterials [BN19, SLS18]. metastasis [JC18, JC19]. meteorological [QDG^{+}15]. meteorology [SPG^{+}16a]. Method [BBQ18, HVV15, LD19a, LD19b, Sch18, SVL19, YMGA19, AAA15, AA16, AVB17, AHP19, ABH19, AHJ17, ASKG18, ACLY19, AHMY19, AHT18, AS18, AG16a, Aks17, AKHSU18, AM16, ACDT15, AN17a, AN19, AZ18, AMB19, ABHVM16, ABEHV16, AAB^{+}18b, AR15a, AR16, ATG18, AZ17, And18, AW19b, ALMLM17, AS19b, AM15b, AH16c, AMS17, ACMT17a, AEHVM18, AEP15, AvG16, AdGvG19, AP18, BRW16, BJM16, BH17a, BJ16, Bac18, BLXD18, BYW18, BDA17, BGII15a, BIJM^{+}19, BBD16, BCFN16, Bay15, BFLP15, BMM17, BKK18, BEJR16, BJK19, BKLR16, BG17, BLM16, Bie17, Bog19, Bog16, BW19, Bou18, BWS17, BK19c, BH17b, COdSR16, CR17, CWXW17, CCW17, CRS^{+}18, CL19a, CH17a, CHX18, CH18, CLZS19, CYY15b, CHR^{+}15, CD16b, CW16b, CZ17b, CQ17, CFX17, CHS17, CZLV17, CM18b, CHS^{+}18, CCX18, Che19b]. method [CFS19, CHYZ19, CL19b, CXZ19, CZ19b, CLZW19, Chi17, CK16, CV17, CG17b, CPZ19, CL19c, CJ19, CGvA^{+}15, CHVRT15, CT17b, CN16, Cou19, CPL^{+}19a, DF16a, DKP^{+}19, DLW16b, Dan17, DF15, DL16a, DP17b, DMMSK17, DAM15a, DM17a, DFLM19, DLM17, DLM16a, DFT18, DIJ16, DIJ17, DI19, ES18, Din16, DPN15, DLHY15, DYZ18, DCNY18, DB15, Dos19, DM16b, DD18b, DFRR15, DD16c, DAAF16, ESAE18, Elg19, EJJ16, EC18, EDO19, EE16, EHVM18, EHV19, Fan15, FNZ18, FSNS16, Fat16, FLZ19, Fer15, FGF19, FMR15, FS19b, FCM19, FR19, GZD^{+}19, GLL19, GOGM17, GZDC19, GHA^{+}15, GUB^{+}17, GCY18, GS18b, GMG19, Gia18, GC18, GD18, GM17a, GDQ17, GD17, GK18a, GAA16, GKN16, GLSW16, GFPGF19, GHC16, GHY^{+}18, Güc17, GM19, GdLT17, GM16b, GxLlW15, GJ17, GLMY18, GTX^{+}19, GHL19, GSRJ18, GHV19]. method [Had15, HLV^{+}15, HMS19, HOS16, HWCC15, HHLS18, HWZ18a, HMMV17, HS17a, HR15, HRS16, HSAR18, HKAR18, HWT15, HSWH15, HXX16, HHKR15, Hei15, HL18b, HVM18, HSAA19, HR16, HLL15, HNT19, HCSC16, HL15, HWC16b, HRGZ17, HLM18, HWH19b, HMY19, HP15, HS17c, HW17, HH18c, HWXC19a, HLLY19, HH19b, HWXC19b, HIS16, ID18, IIA15, JMNB18, JL16a, JO17, JLLZ16, JD16, JLCZ17, JGGS16, JS19b, Kad17, KS18, Kam17, KB17, KM18, KCK18, KYY15, KMB15, KV15, KBG17, KAA16, KGMS18, KNP^{+}17, LS17a, Lad15, Lan18, LX18, LXW19, LBL15, LKS17, Lee18, LHK19, LW15, LXL15, LRX16, LLCG16, LW17b, LC17a, LLX17, LW17a, LMW18, LLD18, LHLL18, LGLC18, LDG19, LLL19a, LGYC19, LLG19, LSYY19, LSDY19, LWZZ19, LH19, LHY19, LLW18, LZ17b, LYG17, Lin16, LCZ18, LC15b, LD15, LZS15, LZ16b, LZZL17, LWC17, LXZ17, Liu17, LCX18, LDW18, LP18]. method [LLC18, Liu18, LFWC19, LF19, LM17b, LN19, LL17, Lu16, LSL17, LSL18, LM18, MTV17, MFX15, ML17a, ML18, MDG17a, MDGVR17, MDAP19, MG15a, MG15b, MR19, MHAA15, MS17a, Mao15, Mao16, MBM17, MVK19, Mar19b, MV16a, MS19a, ML19b, MLJ16, Men17, Mic18, MM15b, MMD15, MH16b, Mit15b, MC18, MT17c, Mol15, Mol17, MFLL18, MWY15b, MWY15a, MWY16a, MY18, MZ19, NE17, NN18a, Nem15, Ngo19a, Ngo19b, NTPVNTDT15, NZ15, NGGFM17, NDH16, NH19, NML18, NF17, NAM18, OM17a, OZ17, OM17b, PKMM19, PSK16, PS15, Par15, PSS17, PLP17, PP18a, PCJ17, Pis17, PI19, QWS15, QM15, QC18, QDBG18, Qui15, RKB16, RRM19, RVA15, RM17, RC19, RS18, RKV19, Reu16a, Reu16b, Reu17, RT18a, RR18, STZ16, SAA18, SR16b, Sak17, SBU^{+}17, Sar19, SEV^{+}18, SB18, SV19]. method [SA17, SD15, SSD16, SR16c, SSW19, SAH15, SPK18, SYZ19, SWH16, SVAKGR19b, SK16, SGG19, Skr17, SLH15, SM18, STS15, SAT15, SCP19, SJL16a, SVV18, SM16, SN19, SSZL16, SX18, Sun19, SHCJ19, SO16, SPM15, TJB17, TA17, TYMG17, Tan17, TH19, TX19, TBR16, TLFC18, TKH16, TP19, TBY17, UGG^{+}19, VFG18, VCL^{+}19, Vej18, VFGPG^{+}18, VFN18, VL19, WYW15, WXW15, WWZ15, WMI16, WWWZ16, WW16a, WZZ16, WZL17, Wan18f, Wan18g, WO19, WSG^{+}19, WWH19, WYLD19, WL19b, WXZ19, Wat16, WCS18, WA17, vtWvdHvdVV16, WL17b, WHWY18, pWmGkY^{+}19, XYZ15, XY16, Xie19, XC18, XMW^{+}15, XM16, XF16, Xu17, XDW18, XZN^{+}19, YZW15, YLL15, YW15c, YY16, YSCY17, YBLH17, YSL17, Yan17, YL18c, YSG18, YZH18, YCX19, YL19a, YZZ19, Yao16, YN18, YAA19, YZ15b, Yi19, Yos17, YJBM19, YSW^{+}18, YSL^{+}18, ZL18a, ZCH18, ZZ19, ZLS^{+}15, ZL15a]. method [ZM15b, ZqZwZ15, ZHWS16, ZZLY16, yZ17, ZZZ^{+}17, ZXQH17, ZHY^{+}17, Zha17a, ZS18b, ZTGZ18, ZZZ18, ZYZ18, ZQJY18, ZSL18, ZL18b, ZH18, ZL19, ZHC19, ZCS^{+}19, ZWT19, ZYY19, ZCL16, ZC16a, ZP19, ZZZ15, ZKRL15, ZL15b, ZC16b, ZC18, ZEBM17, ZH15, yZyMxG^{+}16, dR16, dSBV16, dFG17, BV16b, BV16c, GLS15, GTC16, MWY16b, PG16, RP15, ZWW^{+}18]. methodological [FLMS16]. methodologies [CNQRR17]. methodology [SMS18]. Methods [BGS19, HP17, SDG15, TJCC17, VACR15, ZDFF18, AFH16, Abd16, ABT17, AB18a, AL19a, Ais17a, AST19, AT17, AMCR19, ABK19, ÁD14, ÁD15, AK19, AK17a, AL19b, ABT16, AR15c, AGM15, ACMT17b, AG19, AEK19, AK16, AHKT15, AHK19, AL19c, BCT16, Bad18, BV18, Bai15, BBH^{+}17, BD16, BAMM19, BSGM19, BTKM15, BRS15, BSS15, BBM15, BySZ19, BGI18, BCCB15, BN15, CCTV17, CCZZ16, CJVW15, CDP17, CHS15, CLA19, CN18, CY15, CL19d, CDE18, CEJV16, CWH18, Con18, CJT15, CMQT16, CJT17, CMTV17, CGT18, CPRS19, DDP16, DP16, DKL^{+}18, DP17a, DNL15, DNL16, DdSL16, DXZ18, DGM19, DMCC^{+}18, DZM16, DT16, DD17, EWR18, FYY17, FP16b, FGF19, FR18b, GHM18, GCZ17, GLMT16, GHM17, GG17a, GPHAPR17, GT19, GHZ^{+}15, GP16, GQM16, Guo17, HHR16, HMD19, HM19a, HH18a]. methods [HH18b, HMP15, HWH19a, HPW16, HC15, HHZ18, HH18d, HMT15, HRO16, IÇ18, Jan17a, JM15, JS17, JL19a, JL15, JL16b, JJ19, Jin16, JTCC18, KMRS16, KKS16, KEB19, KVO18, KS16, Ku16, KW18, KK15, KB16, KLL18, LW19a, LLT16, LA15c, LX16a, LY18, LSG18, LWXF18, LWDW18, LSW19, LSH^{+}19, LLSM15, LS18a, LL15, LHZ15, LTY17, LHPZ18, LDZ18, LL18, LQ19, LV16b, Lu18, MZ16, MA16, MMO15, MN17, MN19a, MPM^{+}17, MP15, MII16, Mil15, MT16b, MHL16, MSZ18, MS17e, MAF17, Moh19, MSH15, Mu19a, MHMP16, NDN19, Nak15, NBAK19, NHA17, NAH18, NADK18, Ou18, PS18, PMN19, PW16, PBJCM19, PP15, PKR18, PP19, Pol16, RNM16, RRM18, RMVA17, RWW17, RP16b, Ran15, REBF16, RRS19, SW17a, SKB15, SGL17, SS16c, SEÖ16, SLS18, SVAKGR19a, uIuD18, SW17b, Son16, SV16a]. methods [SM17, Su18, SK18, SW17c, TYS17, TYQ^{+}19, TB18, TJCC16, VAR15, VARC17, WWM17, Wan18a, WZS18a, WW19a, WZWJ19, WR16, WZ17, WL17a, WL16, XW18, Xu18, XH19, YP18, YYL15, YZ16, YHZ18, YCF18, YTSZ15, YP17, YH18b, ZCST18, Zak19, ZuI17, ZST15, ZM15a, Zha15a, yZLX15, Zha16, ZCZM17, ZX19, ZYF19, ZYL19, ZY18, ZM16b, ZJ17b, ZZHS17, ZZ15, ZG19c, ZDL16, ZDF^{+}17, LS17b]. Metric [BDFM19, HKAR18, FZW19, FZ15, KS17c, dHKdHMM15, RMMRC15, TBS18]. Metric-like [HKAR18]. metrics [CFMÖ19]. metropolitan [FAVGC^{+}15]. Mexico [GRVC15]. Meyer [KR17]. MFDSA [DLY18]. MFE [LRX16]. MFEM [ZCH18]. MFS [RVC19]. MHAM [YMGA19]. MHD [LSH^{+}18b]. MHSS [ZZ19]. MHSS-like [ZZ19]. MIB [FLZ19, CWXW17]. MIC [WZS^{+}18b]. Michaelis [TK16]. microchannels [KAM16b]. micromagnetics [AMP15]. micromagnetism [AEK19]. micropolar [CZ16]. microscale [CRCM16]. Microscopic [BM19b]. midpoint [CCW17]. Milan [SPG^{+}16a]. mildly [KMV16, XZXQ18]. Miller [OKGB19]. Milstein [GLMY18, ZWX18]. Mimetic [LV16b, BAK18, SFGJR^{+}16]. Mindlin [Che18, NTPVNTDT15]. mineral [BBPR15]. Minimal [BYW18, CNT17, LZL16, Prz15, AM17a, AKR18, DSDV15, FGS19, GM17b, HG18, KP15, LCA17, LH19, LLY19, NP19, Pap16b, dBGMUA18]. minimal-twist [FGS19]. minimal/maximal [AKR18]. minimax [Fuk15, GGG19, MLS15, Yao16, Yao18]. minimising [SSKV17]. minimization [CZ17a, CPL^{+}18, CPL^{+}19b, CPL^{+}19a, DLS17, KKP^{+}19, LWZ^{+}15, LZ18b, MHMP16, WS17, YSL17, YGC19, ZY19b]. minimizations [BLXD18]. Minimizing [ZYXT16, FGS19, Wu17, ZXQH17]. Minimum [DD16b, SWL15, LHJR15, Lu15c, MYG18, RKS16, SS18b]. minimum-entropy [RKS16]. Minkowski [BLV19, DV15, KLK16b]. MINRES [HJ17b]. minterm [ADMV19]. Miranda [NW19]. miscible [LRX16, ZCH18, ZZZ^{+}17]. Missing [GLdCJS17, WSL17]. misspecified [EP18]. Mittag [AAM18, Hui16, IAAW18]. Mittag-Leffler [AAM18, Hui16, IAAW18]. mixability [PW15a]. Mixed [AFM19, CLV16, CL16, HS18c, MNFS18, SKB15, YSL^{+}18, ZYF19, AÖ17, AG16b, Arm18, BGMS16, BS16, BK15a, BKK18, BEMA16, CLS18, CM19a, CCJ16, CHS15, CW16a, CL19c, Cou19, ET19, FM18b, GSM18, GSM19, GYS17, HR15, HM19a, HC15, HCSC16, HLZ19, IN16, KP15, KDW17a, LLG19, LSYY19, LYS19, LLSM15, Mol15, MWY16a, NM17, RFLM16, SM19a, Sho18, YCF18, ZM16a, ZHY^{+}17, ZYZ18, ZCL16, ZL17a, ZY19b, ZH15, iW16]. mixed-hybrid [HCSC16]. Mixing [DEY19]. mixture [BSLL19, GHL18]. mixtures [BRA16]. Möbius [GKM18, YZ15a]. mock [DDM15]. mock-Chebyshev [DDM15]. mode [HHKR15, SODBK19]. Model [CTAT19, KVO18, Per16, SSKV17, WYLD19, ALS18, AALMLM19, AHJ17, AHT18, AAB18a, ABB17, ABG17, ABL^{+}19, AH19, AdHR17, AC17a, ÁAM15, ALMLM17, ACLM18, ABB18, AFM19, BE19, BSLL18, BSLL19, BYYV16, BFMP19, Ber15, BCS17, BEMA16, BKN16, BRA16, BLH^{+}17, BBPR15, CM19a, CCC19, CP15, CCD16, Chi17, CFSS18b, CKL18, CVW17, CLVW18, CSE^{+}19, CFMN15, CNQRR17, CD15, CLPZ17, DDG^{+}16, DJAMJ17, DBdLdM^{+}19, DMAS^{+}16, DFGS16, DYD19, DI15, EGK17, EAKS18, ERR18, EG17, ET19, FNV17, FM19a, FM17, FNdO15, FLMS16, FF15, FL16, FY18, FCM19, FR19, GZH19, GRVC15, GKBG17, GLMT16, GKPS19, GMCD15, GC17, GLP16, GHL18, Gun18, GWY17, GB18, Hac18, HVV15, HYF18, HS18a, HZ18b, HWP^{+}15, HRGZ17, HYZ16, HJL18, JCC18, JBFS15, JC19, JMSL19, JO17, JSW18a, JL19b, JYK16, JSW17]. model [JSW18b, JA17, KMD^{+}16, KJ19, KKW16, KV15, KV17, KSW19, KPW17, KAS18, LRRV18, LV16a, LPST17, LRZ15, LHY16, LWV18, LZY17, LM17a, LCX18, LZX19, LPT17, LFKG16, ML17b, MLW19, MDVM17, MDG19, MDVR19, MFK^{+}15, ME18, MVdRE^{+}17, MVQDdR^{+}18, MCGNBD17, MNFS18, Mel19, Mer19, MFLL18, MMPGCSP17, Mud19, NM17, NMKK19, NS16, NGGFM16, NN18b, OAN17, OLRPdCJ19, OS15, OE19, PD19, PBP16, PBMJC18, PC17, PCJ17, QLZ19, RdSLL16, RZ19b, RL18, SR16b, Sai16, SJRR17, SCMS18, SSZ19, SSCM18, SP17, SJL16b, SCG18, SZ16, Sta18, SV16b, TBAH16, TK16, TLE19, VS15a, VMM18, VZDV18, WH16, WC17, Wan18d, WRZ18, WW19b, WD16, WS16, Wu16a, WS18, Wu19, XWWI18, XY15, Xu18, YZPL18, YZJC15, YT16, YZY17, YL18a, You19, Yua15, YL18d, YL19b, YCW17, ZB18, Zha15b, ZZLY16, ZSW16, ZSZ17, ZHY^{+}17]. model [ZLF18, ZYL18, ZS19, ZYY19, ZLPF19, Zha19b, ZH15, dAFvdMR19, dFG19]. Modeling [DSA16, GMSR18, GMSR21, JCC17, JCC19, MLV16, QLZ19, TJCC17, AMZ19, BRW16, BV16a, BdSFL16, Bot16, BHU18, BDF15, CROC16, DWS^{+}18, DSN17, EM15, ESI^{+}19, GR16, JC18, JTCC18, KS15b, LS17a, LE18, MAK19, MXW17, OSC15, Sch18, SH15, TJCC16, Xu17, ZHY^{+}17, ZHC19]. Modelling [BN19, DNY17, Tun18, AK17b, DMCC^{+}18, MX16, OD17, RdSMO17, RdSMO18, TBAH16]. Models [ZDFF18, AALMLM18, AA17, AG16b, ABP15, ALMLM16, BHN15a, BZ19a, BAMM19, BSGM19, BM19b, BKMO19, BKMO20, BCNP15, CRCM16, CEJ18, CLHMT15, Dan17, DLS17, DT17, DV16, DM17b, DP19c, FEC18, FH16, FH18, FFP16, FFKP17, GM16a, GMSRP16, GMSR18, GMSR21, GD16, GMW18, GVMR16, GVHMR17, GVMR19, GYQS15, GYLQ16, GYL16, GUB15, HT18, HXC19, HZ17b, JMSL19, JJX17, KLM17, KSR15, EP18, KS16, KK17, LY15b, LLZ16, LXD^{+}18, LYG15, NCF^{+}19, OAN17, RTV15, RANG15, RKS16, SSD16, SS18d, Sta19, SMS18, SN19, SPG^{+}16b, TYMG17, TC19, TW19, VZDV18, VX19, WS17, WAGVRB19, YY16]. Modern [DLY18]. modification [MH16a, NAM18]. Modified [BDSB16, GKN16, KPT15, LL16a, MFX15, iW16, AR15a, AR16, BW19, CCW17, CL19b, CPL^{+}19a, DL16a, DRST19, DLHY15, DZM16, EE16, FZW19, FNZ18, GKN18, GKM18, GxLlW15, HWC16b, HS17c, HH18c, KPRA19, KMLB17, KAS18, Lan18, LX18, LW19a, LSH^{+}19, LF19, LX17, MDGVR17, MM18, Mej17, MWY15b, TZZ18, XC18, YW15c, YMGA19, YSW^{+}18, ZL17c]. modified-upwind [ZL17c]. modulated [BZ19a, DS17a, RP17]. Modulus [BBH^{+}17, LY18, SPCSA16, WL16, ZL15b]. Modulus-based [BBH^{+}17, LY18, WL16, ZL15b]. MOL [DM17a]. mold [DFGS16]. molecular [MXW17]. Möller [BW17]. moment [BHN15b, GL19, KF16, MG17, NCJ16, PA19, RKS16, ZLPF19]. moment-recovered [MG17]. moments [DD18a, Hag15, JN19, Lev18, MS17d, Ruj16]. momentum [AVB17, GHM17, MPM^{+}17]. Monge [YX19]. monitor [XLS^{+}19]. monitored [HLL15]. mono [HL18b]. mono-energetic [HL18b]. monochromatic [BHU18]. Monodromy [AGR15a]. Monolithic [BZ19b]. monoliths [DMCC^{+}18]. monomial [CL15, JR15]. Monotone [Bog15, AK17c, CLS18, CW15a, CW17, GJ17, HMS19, MM18, SDH17, SLW17, TH19, WZH19]. monotone-type [CW15a]. monotonic [MTMZ19, QL19]. monotonicity [GS18a, GS19, MVM19, MSH15]. monotony [BCS17]. Monte [SD15, SSD16, ABT17, BH17b, CFMÖ19, Dan17, God16, HWZ18a, HWH19a, Kaw17, MOY18, Nak15, SMTdF18, SM19b, SX18, XW18]. Monte-Carlo [Dan17]. Moore [MM19]. Moran [You19]. Moreau [YP17]. Morgenstern [DG16b]. morphing [ZF15]. Morphological [TLLL19, LAVMA18]. Morrison [WLZ17]. Morse [CZ15]. mortality [GD16]. mortar [PSW18, YCF18]. Moser [AM15b]. most [BM18, CGT18]. motion [Lev18, OGR17a, LGB18, PBP16, Sho18]. motions [FGS19]. Motivations [Bai15]. motor [DT18]. movement [SK18]. Moving [MZ16, VFFGV16, ACLY19, AN17a, AN19, ABT16, BS18, DM17a, GO17, LBL15, LLSZ17, LC15b, Mir15, PCJ18a, SV16b, SN19, TW17]. moving-mesh [BS18]. MP [SMS18]. MP-LROM [SMS18]. MPI [GDC15a]. MR [HJ17b]. MR-II [HJ17b]. MSEIR [MVdRE^{+}17]. mud [ZFK18]. Multi [AK16, HWZ18a, IABIPG17, LSH^{+}19, Moo17, OSC15, QLT^{+}17, VCL^{+}19, VG15, yZyMxG^{+}16, AK19, AGM15, AP18, BE19, BB18, BKK18, BJK19, BP17a, BDSB16, BHU18, CG17a, CCIL18, CLZL18, CLC18b, CLA19, CH19a, CGvA^{+}15, CMRP17, Dan17, DS19, DSA15, DFRR15, EAKS18, ESI^{+}19, FJ19, FLJT18, FFKP17, HYQ^{+}18, IDGH18, IAMCI18, KV19a, KS16, LSBL18a, LLV18b, LNP17, LX17, MO18a, NDN19, PMN19, PAC15, RdV19, RANG15, Sar19, SLA19, WS17, WWM17, WSG^{+}19, XMW^{+}15, XqYL^{+}15, YZPL18, YMGA19]. multi-adjoint [CMRP17]. multi-agent [MO18a]. multi-block [CLZL18, WS17, XMW^{+}15]. multi-component [KS16]. Multi-conformal-symplectic [Moo17]. Multi-Continuum [VCL^{+}19]. Multi-core [yZyMxG^{+}16]. Multi-criteria [IABIPG17, IAMCI18]. multi-degree [BDSB16, LX17]. multi-dimensional [AK19]. Multi-domain [VG15]. multi-element [DS19]. multi-elimination [CH19a]. multi-frequency [BHU18, WWM17, WSG^{+}19]. multi-Galerkin [PMN19]. multi-input [BJK19]. Multi-level [HWZ18a, CGvA^{+}15, Dan17, YZPL18]. multi-linear [LLV18b]. multi-objective [CCIL18, XqYL^{+}15]. multi-output [BJK19]. multi-parameter [RANG15]. multi-patch [FJ19, KV19a]. Multi-phase [OSC15]. multi-physics [ESI^{+}19]. multi-point [AP18]. multi-point-parametric [AGM15]. multi-product [HYQ^{+}18]. multi-scale [CG17a, DFRR15, EAKS18, FFKP17, IDGH18]. multi-state [LSBL18a]. Multi-step [AK16]. Multi-term [QLT^{+}17, BB18, CLC18b, CLA19, DSA15, FLJT18, LNP17, SLA19]. multi-time-step [BKK18]. multi-valued [BP17a, PAC15]. Multi-variable [LSH^{+}19]. multi-wavelets [RdV19, Sar19]. multicomponent [BSLL18, CFS19]. multicontinua [VCC^{+}19]. multidelay [ZZW16]. Multidimensional [RMMRC15, ACCF19, BK15c, SYD17, WCS18, Yam17]. multiframe [HS18a]. multifrontal [yZyMxG^{+}16]. Multigrid [KLM17, lCCVV18, LP19, LLX17, LSH^{+}18b, LX16b, MS19a, RTV15, SSW19, SWH16, SPM15, TC19, pWmGkY^{+}19]. multilayer [MC19]. Multilevel [GT17, Chi17, Dos19, LXZ17]. multimemetic [NC16]. multinomial [HXC19, Zha17a]. multiparticle [CDH^{+}17]. multiphase [BHN15a, ZWW^{+}18]. Multiple [CM19b, CBMO15, HAH17, AAA15, ABK19, ATG18, GH16a, GH19a, GKN18, GKM18, GGK18, GYQS15, HH18b, HHZ18, JHPT15, LAVMA18, LDZ17, LLSZ17, LSBL18b, MLJ16, MSY16, MSY17, Nie15, PLP17, PN16, PP18b, PP19, SB16, SLQ16, VOB16, WZBM18, XH15, XGH18, Xie19, Yao18, ZCST18, xZWlC15, ZWZ^{+}18]. multiple-zero [GKN18, GKM18]. Multiplex [PGRC18b, PGRC18a]. multiplication [OOO15, SS19a]. multiplicative [CDZ16, GVHMR18, JMS17, LZ16c, Lu16]. multiplicity [PN16, TLE19]. multiplier [AB18a, HWH19b]. multipliers [CLZS19, OZ18]. multiply [BD16, FJ19]. multiply-connected [FJ19]. multipoint [DD17, PP19, Reu16a, dlCYC15, YPD16]. multipole [ZX19]. multiquadric [WWL15, ZYY19]. multiresolution [AJR15, SRH17]. multiresolutions [KBL19]. Multiscale [BV16b, BV16c, CLVW18, GLS15, GHM18, GTC16, JO17, LS17b, LWV18, MWY16b, PPM16, PG16, RP15, SVL19, VMM18, AEKGP19, AEK19, CYY15a, CHR^{+}15, CVW17, CPZ19, CL19d, CL19c, ESI^{+}19, FCM19, FR19, GCY18, GDC15b, JL16b, KGL19, LA15c, Liu18, NPR18, SVV18, SPM15, TH19, YSCY17, ZCFW17, ZHY^{+}17, ZZZ18, Zha17b, ZWW^{+}18]. multisplitting [yZLX15]. multistep [HSAA19, LYS19, MAF17, MSH15, yZ17]. Multivariance [OD17]. Multivariate [DS17a, LCA17, NN17, AL18, BW17, COGD19, CM19b, CL15, Kaw17, KSR15, KNP^{+}17, MLW19, Qi18, TLTJ18, WWL15, Wu19, ZW16]. multiwavelet [PPM16, SM18]. multiwavelets [AHL17]. multiwise [Lu15a, LJ16]. Müntz [EBJS19, GMG19, HNT19, NN18a]. mutative [WZS19]. mutual [ZJQW18]. mutually [ZTGZ18]. myocardium [AdQdS16, AMZ19].

nabla [Das18]. Naive [TK15]. nano [ALS19]. nano-structures [ALS19]. nanoparticles [CSS19, EAKS18, RdSLL16]. Nanoscale [ALS19, VS15b]. Nash [NDH16, NH19]. natural [FP18, LS17a, ML17b, WAGVRB19, WS18]. Navier [BAK18, CJCDLL19, CDW15, CGPR17, CHS15, CHS17, Chi17, CK16, GM17b, GGC15, GGRN17, HMY19, JLCZ17, KtTBK18, KK19c, LA15c, LZ16a, LHM16, LLC18, MT16b, MV16b, MV18, Ngo19a, QM15, SSG16, ZZZ18, ZL19]. NAVIMA [AGR^{+}15b]. near [BIR17a, BLH^{+}17, FFP16, Gor16, KDB17, LRKV16, SB19, SOMF19]. near-best [BIR17a]. near-contact [Gor16]. Nearest [KLK16a]. nearly [CZ17b, EJJ15, FdGKZ17, ML19a, ZLS^{+}15, ZCS^{+}19]. negation [CLM19]. Negative [Ahn15, APS15a, CCY18, COGD19, DYD19, Ery16, JS19a, LV16a, PW17, SR16c, WZS19, Yao18, YC19]. negatively [HHS^{+}15]. negativity [AHT18]. Neighborhood [SQRE16, CCZ^{+}18, YLL15]. neighbors [FP18]. neighbouring [Qi19]. Nekrasov [GWLL18]. Nelder [SGK16]. Nernst [ABL^{+}19, HS18c, SSZL16]. nested [GSVSM16, HWZ18a]. nets [God16]. Network [SSG18, HJ17a, IAMCI18, KBT15, NS19, NN18b, NN19, Ogu17, SPCSA16, ALS19]. networked [WS19]. networks [AG18, AY19, APYW19, AYSE19, AVGCMVM17, ASV17, BMLR15, CFG^{+}19, DMR19, FR18a, LRRV18, LJ15, LZA19, LWS16, LWTQ17, LZL^{+}18b, LDZ18, SZZ^{+}16, TJSW19, bWzZjH16]. Neumann [CDKN15, ACEJ18, BySZ19, CDKN14, GHB17, JHK17, KM18, RVC19, SGG19, SLH15]. Neural [SSG18, AG18, AYSE19, HJ17a, LJ15, LMUZ18, LWS16, LWTQ17, LDZ18, NS19, NN18b, NN19, bWzZjH16]. neural-dynamics [LMUZ18]. neuro [SBP17]. neuro-fuzzy [SBP17]. Neuroelectric [PP16]. neuron [KGL19]. neuronal [BCNP15]. neutral [AYSE19, CPB19, GVMR16, GVHMR17, GVMR19, GH16c, JY17, LY15a, Lan18, LW19a, LZ16b, Mil15, OM17a, Reu16a, Tan18, Wan17b, WZ17, ZJ19, ZWH15a]. neutral-type [AYSE19, ZJ19]. neutron [VFFGV16, VFGPG^{+}17, VFGPG^{+}18]. Newton [AGR15a, AM15a, AH16c, AMS17, AEHVM18, AG19, BCFN16, CZ17a, Chu15, CJT15, DLW16b, DM16b, EHVM18, EHV19, FNZ18, Fer15, GHA^{+}15, GKN18, GKM18, GM17a, GHV19, HLV^{+}15, HR16, KLC17, LSDY19, LLW18, MG15b, NF17, RKB16, RVA15, RM17, SVW19, XYZ15, Xu15, ZY18, ZL15b]. Newton-like [AG19, GKM18]. Newton-type [CJT15, GKN18]. Newtonian [LLZ16]. Nicolson [GSRJ18, JKL^{+}15, LMW18, LGYC19, Moh19, RZ19a, SW17d, ZZW16]. Nieto [DLY18]. NIG [JA17]. Nikishin [LP15a]. nilpotent [AGR15a]. NIRK [KK17]. NIRK-based [KK17]. NLMC [VCL^{+}19]. NMF [TYQ^{+}19]. no [CT15, Rom18, ZP16]. no-arbitrage [Rom18]. Nodal [ZC18, KLS19]. node [MLSK16]. nodes [CH19b, GGK18, HHZ18, PM15]. noise [CDZ16, GHB17, GLMY18, JC15, JMS17, KB17, Liu16, Lu16, PBMJC18, Yam19, ZZHS17]. noises [WS19]. noisy [NS18, ZLAA15]. Non [BAAA^{+}15, BKMO19, BKMO20, CP16, HHS^{+}15, LW19b, ZD19, AHT18, ABRS17, ADD^{+}16, AG16b, AR19, BBCP16, BBC^{+}18, BS16, BS17b, BWS17, BBPR15, CJCDLL19, CFR18, CCJ18, CH18, Che16b, CXW19, CK16, CNQRR18, CPL^{+}18, CPL^{+}19b, CPL^{+}19a, DBX19, Das18, DD18a, DAM15b, DIJ17, EG17, GLMT16, GUB^{+}17, GKN16, GHC16, HMS19, HLMS19, IÇ18, JMN^{+}16, JY19, JN18, Kad17, KD17, KH18, KHLY17, KV17, KV18b, KS17b, LV16a, LLZ16, LGYC19, MTV17, MDVR19, MM18, MVQDdR^{+}18, MHV17, MPH18, MP18, Mol15, NSYX17, OSC15, PHD17, PW17, PST18, Qi19, SuSR16, TBAH16, UGG^{+}19, VCL^{+}19, WCLL18, WW19b, WXZ19, WW16b, WA17, Yin19, Zak19, ZX16, ZKRL15, ZWH15b, VCL^{+}19]. non-autonomous [BBCP16, BBC^{+}18, CCJ18, CNQRR18, Das18]. non-bandlimited [Che16b]. Non-commutative [LW19b]. non-convex [CK16, CPL^{+}18, CPL^{+}19b, CPL^{+}19a, HLMS19]. non-degenerate [Mol15]. non-equilibrium [MP18]. Non-exchangeability [BKMO19, BKMO20]. non-extensive [ZX16]. non-globally [DIJ17, ZWH15b]. non-Hermitian [GHC16, WXZ19, Yin19]. non-homogeneous [MDVR19, VCL^{+}19]. non-identical [DD18a]. non-intrusive [CH18]. non-isothermal [BBPR15, OSC15]. non-linear [BWS17, CFR18, GLMT16, GUB^{+}17, Kad17, KV17, KV18b, PHD17, TBAH16, UGG^{+}19]. non-Lipschitz [BS17b, CXW19, DBX19, IÇ18]. non-local [AG16b, AR19, VCL^{+}19]. non-Markovian [WW19b]. non-monotone [HMS19, MM18]. non-negative [LV16a, PW17]. Non-negatively [HHS^{+}15]. non-negativity [AHT18]. non-Newtonian [LLZ16]. non-oscillatory [JN18]. non-overlapping [ZKRL15]. non-periodic [CJCDLL19]. Non-polynomial [ZD19, WCLL18]. non-rectangular [DAM15b]. non-regular [KHLY17]. non-relativistic [KS17b]. non-ruin [EG17]. non-selfadjoint [WW16b]. non-smooth [GKN16, MTV17, Zak19]. non-square [JMN^{+}16, WA17]. non-stabilized [BS16]. non-standard [MVQDdR^{+}18]. non-stationary [ABRS17, JY19, LGYC19, SuSR16]. Non-transversal [BAAA^{+}15, ADD^{+}16]. non-uniform [KD17, MHV17, MPH18, NSYX17]. non-viscous [PST18]. non-zero [KH18, Qi19]. nonautonomous [MT16a, MP17a]. nonbacktracking [CFG^{+}19]. noncompactness [HSAR18, Ros17, Sai19]. Nonconforming [LS17b, HRGZ17, KYS16, ZCZM17, ZML16]. nonconstant [Wan18c, Wan18d]. nonconvex [BP17a, LSL15, PAC15, SBCJ18, YSW^{+}18, YWY19]. Nondegenerate [GV17]. Nonexistence [KFT17]. nonexpansive [Iid15]. nonhomogeneous [RVZ15]. nonignorably [WSL17]. noninstantaneous [LDW18]. Nonlinear [BO18, BV16c, lCCVV18, ELRR17, HYF18, MFK^{+}15, VK16, ZBL^{+}16, AAA15, AÖ17, AA18, ABK19, ÁD17, ABHVM16, ARTY19, AR15a, AR16, ACDT18, ACDT19, AEP15, AAC17, AJ17, AJ18, AD18, AK17b, AHKT15, AHK19, BM17b, BM18, BS18, BBB17c, BT17, Baz15, BPP17, Beh16, BMM17, BAMM19, BSGM19, Bie17, Bog15, Bog19, BK19b, CLS18, CLC18a, CR17, CCW17, CRR16, CD16b, CL19b, CS17, CS18b, CN18, CMQT16, CJSCT18, CTAT19, DDR19, DP17a, Das15, DSN17, DWXO15, DY19, DY15, ESAE18, ERR18, EHTL19, Fan15, FNZ18, FLN16, FGF19, FCM19, GM17a, GAA16, GS15c, GGC15, Guo17, HWCC15, HA18, HA19, HSAR18, HKAR18, HSWH15, HH18a, HSAA19, HK07, HZ19, HP15, HW17, HMT15, HKZZ15, IÇ18, JGTB15, JJX17, JLCZ17, KS17a, KYY15, KRR18, KV19b, KMV16, KBG17, KKK16, KMB^{+}18]. nonlinear [LH16, LY18, LMW18, LMUZ18, LL15, LM15, LHPZ18, LF18, Liu18, LM17b, MLW19, ML19a, MDG17b, MDAP19, MD19, MMO15, MS17a, MD18, MYM16, MK17, Mil15, MH16a, MH16b, MA18, MSH18, MAS19, Mol15, MRS17, MJS19, NDN19, NE17, NBAK19, NW19, NDH16, NN18b, NXLX18, PTV17, PT17a, PCJ17, QWS15, QC18, Qui15, RRM18, RVA15, RM17, RZ19a, RdSLL16, RQ16, Ros19, RW16, RMK19, SE16, SAC17, SLH18, SY19, SK16, Sto18, SL18, SLW17, SW17d, SH15, TGJ19, TH19, THL^{+}17, VZDV18, VDT19, VX19, WYW15, WCW15b, Wan17b, WWH19, WCS18, WL17a, XGH18, XY16, XHZ17, XY19, XDY16, XH19, YP18, YT16, YY16, YGS18, YZZ19, YAA19, YTSZ15, YZ15c, ZCST18, ZL15a, ZH18, ZHH^{+}19, ZC16a, Zha17b, ZJ17b, ZZ15]. Nonlocal [VCC^{+}19, WPA^{+}18, ABN^{+}18, AK18, Dos19, DAAF16, GG17b, PW16, TV16b, VS15a, VX19, ZLF18]. nonmonotone [DK16, Fan15, HW17, HWZ18b]. Nonnegative [BL18, CHYZ19, CS19, KLL18, SNY19, SVGM17]. nonnegativity [RR16]. nonoverlapping [ZCL16]. Nonparametric [CCY18, YC19, LSZB17]. nonpolynomial [FMR15]. Nonreflecting [LD19b]. nonsingular [GW19, MT15]. Nonsmooth [KSD17, BK16, HW17, MHMP16, ZL15b]. nonstandard [MVdRE^{+}17]. nonstationary [BK15a, HH19b, Ngo19a]. nonsymmetric [ATG18, CDW15, CGMM18, DSZ15, FwC17, GHY^{+}18, HS17c, HH18c, JK19]. nonuniform [KK19c, LRKV16]. Nordsieck [DDP16]. norm [AK17c, CPL^{+}19a, DI15, JL16a, LC17b, LF18, Pul17, SWL15, TBS18, YLL15]. Normal [CCZ^{+}18, CCRN19, FS15, KSR15, NL17, NLG15, TV18, VMR17]. normal-compliance [VMR17]. Normality [CBMO15]. normalized [Beh16, HLLS18]. Northern [NGGFM17]. Note [LX17, AL19c, (Ba18, Chr18, Chr19, CLHMT15, FV17, LT18, LZY17, MSY17, MM15c, Ou18, PP18a, RZ19b, Tru19, Wan15a, WHS17]. notice [Beh16, GAA21, HWXC19b]. notion [MII16]. Novel [AJ18, BBCK19, AJY18, BE19, BLXD18, CCC19, FLJT18, LZC15, LWS16, ML17b, MLW19, MBM17, MVC19, NS19, NN18b, NN19, NAM18, Reu16b, RW16, TYQ^{+}19, VFN18, WLZS18, Wan18g, ZL15a, ZZ18]. nozzle [SJRR17]. nozzles [SJRR17]. nuclear [LYS19]. Null [CBC^{+}19, JGAFAL19]. number [BRA16, BS17c, CLPZ17, HLLS18, HMH18, JZ18, LDDF18, WZZ18, WGZ15, WZ18, YW16]. numbers [AdHR17, DWQ16, Dia18, DZ18, GX17, LDDF18, Li18, QG17, QCSG18, Qi19, YH18b]. Numeric [MTS17]. Numerical [AALMLM19, AKR18, AAB18a, AS18, Akh19, AM16, AT17, AVGCMVM17, ÁD17, ACLM18, BM17a, BM17b, BD16, BJK15, BV16a, BHRR18, BR15b, CD16a, CRR16, Ceb17, CW16a, CY18b, CLA19, CXW19, CSS19, CV19, CKB17, CEJV16, CFMN15, CNQR^{+}18, DP16, DBdLdM^{+}19, EM15, FEC18, FJLB18, GLMT16, GS18b, GMG19, GHB17, GMW18, Goć15, GAA16, GO15, GGC15, HH19a, HR15, HLL15, JL19a, JLCZ17, JS19b, Kan18, KF16, KS15b, LP15b, LMS17, LM15, MOPP18, MAS19, MTKO17, MR15, Mus17, NN18a, Nem15, NXLX18, OP19, OH18, PW16, PCJ18b, ROB17, RQ16, SE16, SRRJ16, SODBK19, SS19a, SEÖ16, SS18c, SPD15, SM17, TJB17, TMS^{+}17, TPC17, TKH16, Wol17, XHZ17, XWWI18, XY19, XF16, XHH^{+}19, YHZ18, YZZ19, YFS18, ZDGM17, ZDFF18, dlHV16, AVB17, Abd16, AM19, AK15]. numerical [ALMLM16, ALMLM17, ASV17, AHL17, AJY18, BM18, BBCP16, BBD16, BK19a, BW17, Bay15, Baz15, BGL17, BFMP19, Bog19, BS15, BKN16, BZ16b, BGS19, CRCM16, CHX18, CLC18b, CXZ19, CGSS17, CJ19, Col16, CEJ16, CZ16, DVA19, DAM15a, DSA15, DSN17, EK19, EC18, FLJT18, FLN16, FNdO15, GCTGC16, God15, GYS17, Gri19, GFPGF19, GJ17, GL19, HK19, HKAR18, HHKR15, Hei15, HRO16, IÇ18, JBFS15, Jeo15, JL19b, KB17, KLM17, KKS16, KV15, KV17, KW18, KB16, KPS17, KAS18, LS17a, LD19a, LD19b, LXW19, LLSZ17, LYS19, LDZ18, MTV17, MDVM17, MDG19, MD18, MYM16, MT17a, Mic18, Mil15, MIT15a, MV16c, Mud19, NV19, NSM16, NDH16, PCJ17, PCJ18a, PST18, Qui15, RC19, RdSLL16, Reu17, RW16, RT18a, STZ16, SAC17, SS18a, SLH18, SR16a, SOMF19]. numerical [SDH17, SR16c, SS16c, SAH15, SLA19, SVAKGR19a, SVAKGR19b, SM18, STS15, Sta18, Sta19, Su18, TV16a, TYMG17, TSMO17, TDA18, TV16b, TLL16, TF18, TN19, VM15, WCW15b, WC17, Wan17b, WG18, Wat16, XZ19, YYL15, YAA19, ZuI17, ZSL15, ZLS^{+}15, yZLX15, ZJG^{+}15, Zha19b, ZJ17b, ZJ19, ZDL16]. Numerically [CSTU17, BL16]. numerics [GLR16]. NURBS [LCC^{+}15, YZ15a, ZZ16]. nutrient [AJ18]. Nyström [AK15, DDP16, DL16a, JM18].

O157 [BYWPA19]. obesity [KCDACC17]. object [CD16b, LWLL18]. objective [CCIL18, XqYL^{+}15]. objects [BTKM15, YOK15]. oblique [BG15, Per16, Yin19]. Obrechkoff [SVAKGR19b]. observation [CCY18]. observer [DLM16b]. obstacle [AKR18, GH16a, GG17b, GM19]. obstacles [DL16b, VG15]. obtain [BIJM^{+}19, RRM19]. occlusal [BOF^{+}16]. Occupation [WZ16]. occurrences [EG17]. ocean [LXZ17, LXZ18]. odd [CAO^{+}19, TDA18]. odd-even [TDA18]. ODE [Mär17, MV16d]. ODEs [Abd16, AHP19, AK16, RRM19, RMVA17, Ran15, VARC17]. Off [SHC^{+}19]. Off-diagonal [SHC^{+}19]. offline [GH16a]. offline/online [GH16a]. offsets [ACDT15]. Ohm [CS18b]. oil [YSL^{+}18, ZS17c]. oil-water [YSL^{+}18]. Old [RS19, Mär17]. Olkin [BO16, DGM16b, NKD^{+}19, OK16, OE19]. On-the-fly [CZ19a]. One [BD19, CJT15, DB15, AHK19, Bac18, CL19b, CJ18, CNQR^{+}18, Dzh16, EWR18, GLMT16, GGW16, HHR16, HH16, HT18, KEB19, LC18, MSH15, NSV18, NCF^{+}19, Prz15, QLZ19, RKS16, SM19b, WCW18, WZ17, WL17a, YLL15, ZJ17a, Zha19a]. One-block [DB15]. one-dimensional [CL19b, CNQR^{+}18, GLMT16, HH16, SM19b, ZJ17a, Zha19a]. one-factor [HT18]. one-leg [MSH15, WZ17, WL17a]. one-norm [YLL15]. One-point [CJT15, Prz15]. one-shot [GGW16]. one-step [AHK19]. One-way [BD19, NCF^{+}19]. ones [QLZ19]. online [GH16a, LWLL18, Lu18, YCF18]. onset [Mit15b]. onto [BG15, JCH17]. open [IANS15, KPS16, LSS19, NDH16]. open-loop [NDH16]. OpenDA [GRVC15]. OpenFOAM [MPM^{+}17]. operating [Che19a]. Operational [Pis17, BS19, BS15, Iid15, MD18, MS17c, NE17]. operations [CY18a, DXML19, KS15b]. Operator [KV15, LK18, Rad17, SZ16, ZZLY16, AA18, CLS18, CM19b, CG17b, CDKN14, CDKN15, EE16, FFKP17, HZ19, Lee18, MV16a, Mej17, MV16b, MMRS15, SS19a, SPM15, WW16b, XZXQ18, Zha17b]. operator-splitting [MV16b]. operators [AKR18, AK18, AG16b, BLP17, CGN18, DD16a, EW18, GMW18, HM16, HVM18, KK19b, KKP^{+}19, KT15, MT18, OD17, OZ18, SB16, WWL15]. opinion [MGMDGVR17]. opposite [JSL15]. opposite-bordered [JSL15]. optical [BHU18, DFRR15, TLE19, WCSL15]. Optimal [ALS18, DMM18, ERPD18, FAVGC^{+}15, GLL19, GGRN17, HG17, HYZ16, KP15, LLY19, MO18a, MRMA17, MS15a, MSY16, MP17b, NSV18, Nke18a, Nke18b, PX17, PBP16, PZ18, Pis15, hShX17, SW17d, WWZ15, Wan15b, Wan18h, WRZ18, WS19, ZCST18, ZJQW18, ABH19, ABN^{+}18, ABK19, AN17b, AFN17, AL19c, BYW18, BGMS16, BGMM18, BAMM19, BySZ19, CLC18a, CM16, CDRY18, CD15, DVA19, DI19, DGN19, DLM^{+}19, Dun17, DD17, GCY18, GKN18, HYQ^{+}18, HND18, HNT19, HOTZ18, HY19b, JW17, KD19, KV17, KCDACC17, LKN17, LPST17, LXL15, LQZ16, LXL16, LLX17, LX16b, LL18, LW19b, LN17, LN19, MLV16, MBM17, MG15c, MO18b, MSY17, Mud19, RRW17, Ros19, SE16, SWH16, SF16, Ugu18, VŽ19, WP17, YX19, YL19b, ZYXT16, ZXQH17, ZC19, Zho16, ZC16b, dFG17]. optimal-optimal [SF16]. Optimality [MLS15, SYZ19]. optimisation [NF17]. optimization [And16, And18, AEP15, BE19, BBB17b, Ber15, CFMÖ19, CDD^{+}15, CD16b, DZM16, ESAE18, Fat16, GLT16, GC18, GSM19, GGW16, HSS19, HEA18, HFMM16, HZ17c, HWZ18b, Iid15, IABIPG17, IAMCI18, ICMPG16, JGTB15, KS15b, KTT16, LV16a, LaTG17, LMUZ18, LSDY19, LZS15, LL18, Lu16, MA19a, MMRS15, NRV18, PW15b, QC18, RKB16, Ros19, SV19, SHW19, SQRE16, VL19, WS18, XqYL^{+}15, XY18a, XY18b, YLL15, ZZ19, ZHH^{+}19, ZL17a, ZEBM17]. optimize [LW19b, SGK16, SW16a, SW16b].

Optimized [Xu18, YW15b, AMM19, AK15, CDP19, NGGFM16]. Optimizing [PB15, SM15a, YZ15a]. optimum [BDA17, MRMA15]. Option [BZ19a, HC17, VFG18, CEJV16, DS17a, DM17b, DP19c, FCJ16, FM19a, Hac18, HLL15, JHK17, JYK16, KC17, KDB17, KDB18, ST16b, SM18, SCG18, TYMG17, WH16, XW18, ZYY19, dSBV16]. options [CEJ16, FEC18, FNV17, GB18, HZ18b, HW18a, HT18, HXC19, HOM17, JYK17, JSW18a, LDK17, LM17a, MZ16, MNFS18, MFLL18, MV16d, NP19, PG19, RP16a, RZ19b, Sho18, SX18, TEG15, TB18, WC17, WZ16, XY15, ZS17a, Zha17a, ZYL18, ZY18]. orbital [OGR17a, LGB18, PS15]. Order [AK17a, BE15, CJVW15, Abd16, AH16a, AL19a, AMRX17, ALC19, AZ18, AMCR19, ACAC17, AH15, ACDT18, ALMLM17, ATV15, AR15c, AGM15, ACMT17b, Arm18, AFGT18, AJ18, AJY18, AN17b, BM17a, BJ16, BBCP16, BBC^{+}18, BDA17, BMM17, BGMM18, BAMM19, BK16, BW19, BMG^{+}19, CR17, CRCM16, CDF^{+}16, CJLT15, CHR^{+}15, CQW16, CW16b, CQ17, CLA19, CXZ19, CFWW19, CV17, CLV16, CT15, CHVRT15, CMQT16, CMTV17, CT17b, CJSCT18, CNQRR18, CTAT19, DMM18, DKL^{+}18, DP17a, DVA19, DD18a, DSA15, DGM19, DSN17, DG16b, DCNY18, Dos19, DFS15, DM17b, DP19c, Elg19, EC18, FP16b, FR18b, FF15, FR19, GDM16, GM16a, GH19a, GGBTP17, GUB^{+}17, GKN18, GKM18, GMG19, GLSW16, GLSW18, Gri19, GM17b, GGC15, GGRN17, Guo17, GTX^{+}19, HHR16, HWCC15, HCL17, HHEG17, HMDSR19, HH18a, HSAA19, HWC16b, HLZ19]. order [HMT15, HIS16, JB18, KP15, KSDM16, KD17, Kha17, KJ19, Ku16, KMB^{+}18, LRRV18, LKN17, Lee18, LX16a, LCL16, LSZ16, LW17a, LSG18, LCH18, LHLL18, LWZ19a, LLL19a, LSW19, LLL19b, LHY19, LZL^{+}18a, LZY15, LTY17, LS18b, LM17b, MTV17, MCLW16, MG15a, MA16, MR19, MMS19, MBM17, MVdRE^{+}17, MVQDdR^{+}18, MVK19, MHV17, MO18b, MS19a, MDH19, ML19b, MMZ19, MA18, MAS19, MC18, MR15, Mu19a, Mu19b, NE17, NS19, NTPVNTDT15, NZ15, OGR17b, OGR19, OKGB19, PW16, PD17, Par15, Per16, PHD17, PC17, ROB17, RP16a, RRM19, RMVA17, Reu16b, Reu17, RT18a, RM19, RT18b, SEÖ16, SSW19, SSZ19, SVAKGR19a, SP17, SYD15, SJL16a, SMS18, SI16, TYS17, TAT18, UGG^{+}19, WCW15b, WWM17, WZS18a, WLW19, WD19, WSL19, WKBB17, bWzZjH16, XWWI18, Xin16, Yam17, YZH18, YMGA19, Yi19, YP17, Zak19, ZST15]. order [ZLS^{+}15, ZJG^{+}15, ZJ17a, ZLZ17, Zha18, ZT18, Zha19a, ZYL19]. ordered [RMMRC15]. ordering [(Ba18, LHJR15]. orders [CN18, MFX15, TFS17]. ordinary [BJ16, BO18, BR15b, Con18, HIS16, KW18, LTY17, MG15a, NSW15, PRP19, RM19, TMS^{+}17, WKBB17]. orientation [LAVMA18, TXFH18]. oriented [BN15, CLP16, Jan17a]. originated [SuSR16]. Ornstein [BCNP15, WSW15]. orography [TLL16]. Orthogonal [DV15, IANS15, Man15, PR16, ÁNPPCSA15, ÁAM15, CBMO15, DGSM15, FMP^{+}15, HSS19, JMNB18, KVO18, MMMMB16, MB15b, hShX17, WCW18, WYLD19]. orthogonality [HZ17b, YY16]. orthogonality-projection [YY16]. orthonormal [JBT16, Pap16b, UPD^{+}15]. OrthoQuad [BLLM^{+}15]. orthotropic [GC15]. oscillating [KCX17, SS16c, SVAKGR19a]. Oscillation [WG18, BCA18, HG18, LZL^{+}18a, TDA18]. oscillations [AKHSU18]. oscillator [LZL^{+}18a, Ver16]. oscillatory [BR15b, CY18b, FYY17, FR18b, GCY18, HS17a, HS19, HX15, HXX16, JGGS16, JS19b, JN18, KL15, Kan18, LWDW18, LZY15, LTY17, ML18, Mol17, PS15, uIZ15, uIuD18, WWM17, XM16, ZuI17, ZX19]. Oseen [ZP16]. ossification [FM17]. ostectomy [DMAS^{+}16]. Ostrowski [Wan18g]. Ostrowski-type [Wan18g]. outer [PP15, PKR18, SSH16]. outliers [TJA18]. output [BJK19, LZL^{+}18b, RNM16, VZDV18, ZWWW19]. overdetermination [DVS15]. overlapping [GMCD15, GC17, KK15, Vab18, ZKRL15]. overload [ÁRS16]. overrelaxation [AALZ15]. oversampling [ZHY^{+}17]. overspecified [JAN17b]. overview [ABBB17, GOGM17]. oxidation [YBC^{+}16].

P [DLY18, SVAKGR19a]. P-stable [SVAKGR19a]. Pachpatte [AAKT19]. Padé [AT17, BRZ15, Din16, LP15a, SW17c, Tra15, dlCYC15]. Padé-type [BRZ15]. PageRank [GXZ15, GW17, GJSC18, PGRC18b, PGRC18a, SHC^{+}19, Tan17, WHS17, ZHWS16]. pair [AK15]. paired [HQ18]. pairs [CW15b, RM19, SS16a, Tru19, TFS17]. pairwise [BCC^{+}18]. pantograph [GL19, JBT16, ROB17]. paper [EPC17]. papers [VAMDGR19]. parabolic [Abd18, ABH19, ACAC17, AFGT18, Bay15, BKK18, Bog15, BLP17, BK19d, CGSS17, CJ18, CJ19, DVS15, GCY18, GS15c, GS15d, GM19, GSRJ18, HHOT17, HHEG17, HC15, JAN17b, LLX17, LX16b, LSP18, MDAP19, MVM19, MHV17, MSV18, MAS19, MTKO17, PT17a, SGL17, Slo15, SZ18, TH19, UGG^{+}19, Vab18, VS15a, VDS15, VDT19, YZ16, YX19, YZZ19, ZZW16, MV16d]. parabolic-ODE [MV16d]. Paraexp [KBG17]. Parallel [AG16b, CTZK18, GYQS15, HHKR15, LSP18, ZYZ18, ADMV19, BO18, COdSR16, DS15a, DF16a, HLG^{+}17, KBG17, LSZB17, LN17, LN19, MF16, SM16, SM17, WLLC18, yZLX15, ZS17b, yZyMxG^{+}16]. parallel-in-time [BO18]. parallelization [CSS19]. Parameter [CLX17, DVA19, NMKK19, PRRY18, RC19, TJA18, APYW19, And18, AALZ15, Bog19, CEJV16, CD15, DLM16b, FR19, GTC16, GMdV19, GS15d, Gun19, HG17, HHS^{+}15, HMP15, HRR15, KMD^{+}16, LZ18b, LS18a, LF19, MTV17, OP19, OV17, PB15, PCSR19, RABB19, RANG15, SWZW15, THC^{+}18, Wan18c, Wan18g, WW16b, Xu15, XZ19, Zha15a, Zha18]. parameter-dependent [GTC16]. Parameter-uniform [CLX17, RC19]. parameterization [FJ19, SM15a, WLZS18, XMW^{+}15, YLZG18]. parameterizations [KV18a]. Parameterized [FMC^{+}18, LZ18c, Das15, JL16b, LZ17b, SP17]. parameters [AMP18, AMS17, BOF^{+}16, BZ16a, CNQRR17, GB18, JAN17b, Mar18, NAK17, OZ17, VZDV18, Wan18d, XGH18, Zho16]. Parametric [SMS18, ZH16, AGM15, CMTV17, LCZ18, LZ19, OGR17a, LGB18, RRM19, SPD15, SL18, TA17, TXFH18, VŽ19, WZ18, Zag18]. parametrically [KP19]. parametrizations [SW16a, SW16b]. parametrized [AH16b, CTAT19]. Parareal [Wu16b]. Pareto [CDN19, Gun18]. Parisian [CLPZ17]. Part [AHKT15, AK17c, CGMM18, IGRL17, LWZ17, SLQ16, LD19a, LD19b]. Partial [RKS16, WCW18, AH19, BM17b, BDA17, BPP17, BRZ15, DA19, DT17, Elg19, FCJ16, FLJT18, FLN16, GZD^{+}19, GUB^{+}17, GLSW16, GLSW18, GR19, Guo17, KB17, KKS16, KMV16, KBG17, Lad15, LCZ18, LLY19, LFWC19, LX15b, MDT16, MDG17b, MS19a, MM15c, Moh19, SF18, SAT15, SYD15, Su16, TYS17, TGSS17, WLM16, XYZ15, YB15, ZH18, ZG19b, dlHV16]. Partial-moment [RKS16]. partial-robust-ridge-based [AH19]. partially [HZ17b, JJX17, LLL19b, LYG15, RWW17, RMMRC15, WYW15, Wu16a, YY16, ZSW16, ZSL18]. participation [Lu18]. particle [TKH16, XqYL^{+}15, XY18a, XY18b]. particles [Gor16, VRBS19]. particular [TLFC18]. partition [DMP19, DP17b, FP16a, ZS18c]. Partitioned [LY15b, WMI16]. partitioning [HKZZ15, pWmGkY^{+}19]. partitions [BIR17a, DI19, GSVSM16, WO19]. partner [LE18]. party [DT18, MGMDGVR17]. Pascal [KF16]. passage [KKW16, STZ16]. Passivity [APYW19]. Passivity-based [APYW19]. passport [KC17]. past [DS19]. patch [FJ19, HS18a, KV18a, KV19a, LSYY19, Uda18, WZZ18]. patch-based [HS18a]. patchy [KJ19]. Path [GH16c, HR16, JYK17, Prz19]. path-dependent [JYK17, Prz19]. path-following [HR16]. Path-regularization [GH16c]. pathogen [BYWPA19]. pathological [LS17a]. paths [HAH17]. Pathwise [KB17, KRR18]. pattern [BIR^{+}17b, SJRR17]. patterns [SFH^{+}16, SS16c]. PC [AJ17]. PDE [CH19a, FR19, GTC16, LW19b, Pol16, SV19]. PDE-constrained [LW19b]. PDEs [BM19a, CTZK18, CHS^{+}18, GPHAPR15, GPHAPR17, GFPGF19, GGW16, HM16, JL16b, LX16b, LV16b, Moo17, NW19, RRM18, UGG^{+}19, YGS18, yZyMxG^{+}16]. Péclet [BS17c]. pedestrian [BM19b]. pediatric [TS17]. peer [HPW16, KW18, SW17a, SGL17, SW17b, WKBB17]. penalization [LWZ19b]. penalized [BEMA16]. Penalty [SCP19, BySZ19, CD16b, CG17b, CP18, CPL^{+}18, CPL^{+}19b, ESAE18, HMS19, HWCC15, HWT15, HL18b, HCSC16, HP15, JMSL19, KYY15, MRS17, SZ18, WWH19, YCW17, ZC16a]. penalty-based [HWT15, JMSL19]. penalty-free [ZC16a]. pencils [MT16a]. penetrable [LXZ18, YL18b, YOK15]. Peng [KS16]. Penrose [MM19]. pension [Mud19, Nke18b]. pentadiagonal [MT15, WLZ15]. pentagonal [HLLS18]. percentile [ZZW18]. Perception [TS17]. perfect [UPD^{+}15]. perfect-reconstruction [UPD^{+}15]. perfectly [SBU^{+}17]. perforated [CJCDLL15, CLV16, CVW17, CLVW18, SVL19, VCL^{+}19]. perforations [VCL^{+}19]. Performance [Ber15, CBP18, HLG^{+}17, LSH^{+}18b, RNM16, CY18a, Che19a, DF16a, DMCC^{+}18, NGGA^{+}18, RSID16, TW17, WL17c, WLC^{+}18, Wu18, WH19, YSCY17, dMPF15]. perfusion [AdQdS16, CL18a]. period [BE19]. Periodic [LD19a, ÁD17, AK15, AL19c, BCNP15, CJCDLL19, CCZZ16, DCNY18, FM18b, FS19a, HK18, LCC^{+}15, MCLW16, SY18, SM19a, SM17, VAR15, YSCY17]. periodicity [AKHSU18]. periodogram [HHS^{+}15]. permanganate [YBC^{+}16]. Perry [And17, Ou18]. Persistence [JC19]. personalized [PGRC18b]. perspectives [MLV16]. Perturbation [LWJ15, Wu19, DVA19, ERPD18, GGC15, HSAR18, ZEBM17]. perturbations [BB16, HKZZ15, LP15a, LI15]. perturbed [AS18, AZ18, BR15a, BZ16a, CLX17, CW16b, CGSS17, CJ18, CJ19, Das15, DVA19, ESSE18, EC18, GO15, GO17, HCA^{+}18, HMH18, KAA16, KK15, Lin16, LC15b, LC17b, LM17a, LLHO18, LM17b, MTV17, OP19, PCSR19, Qui15, RC19, SAA18, SR16c, Sha15, YAA19, ZC18, dR16]. Petrov [DKL^{+}18, Zha15b, ZG19c]. Petrovski [MDG19]. Petviashvili [ÁD15, ÁD14]. Phase [ARCA^{+}19, ÁAM15, IL15, TFS17, YWW19, AG16b, CJP17, CM19a, CRCM16, CSW19, Cou19, ES18, EAKS18, Gru16, HMMV17, HCSC16, KS16, LWW17, LZW18, Mit15b, MV16c, OSC15, PGD^{+}18, PG16, RNPK15, RP15, Sai16, SVAKGR19b, SVV18, Tho16, WAGVRB19, YSL^{+}18]. phase-field [LWW17]. Phase-fitted [TFS17]. phase-lag [SVAKGR19b]. Phase-type [ARCA^{+}19]. phenomena [TKH16]. phenomenon [ARTY19, HWP^{+}15, KD17, SR16a, TDA18]. Phi [GGK18]. photonic [Bot16, LC17a]. PHT [QDY^{+}19]. PHT-splines [QDY^{+}19]. physical [Mar18]. Physics [CBP18, ESI^{+}19, HVV15, WCSL15]. physics-based [WCSL15]. phytoplankton [AJ18]. PI [TJSW19]. Picard [MDT16]. PID [DMM18]. pieces [LZ18a]. Piecewise [LWXF18, Men17, BCFN16, CGMGV15, CS18a, Fuk15, GLR16, GB18, HH15, JD16, LDDF18, LL15, LSL17, LSL18, Maz18, Mil16, MA18, MGMGJLRB17, Pis15, WGZ15, WZ18, Zha16, dFLM18, vSV15]. piecewise-linear [dFLM18]. piecewise-smooth [GLR16]. piers [EM15]. Pin [VFGPG^{+}18]. Pin-wise [VFGPG^{+}18]. pinning [TJSW19]. pipe [BHN15a, LLSZ17]. pipe-to-pipe [BHN15a]. pipeline [HYQ^{+}18]. pipelines [DS15a]. Piper [SST15]. Piscounov [MDG19]. Pitkäranta [LSD15]. pivoting [APS19b, HY19a, dBAAA15]. pixel [MA19a]. plan [HEA18, Mud19]. Planar [Lu15c, ALV19, AC17b, BLMS19, CL19c, FJ19, FW18, KV19a, LFCW19]. Planck [Wol17, ABL^{+}19, FJLB18, HS18c, HL18b, LP19, LP15b, SRA^{+}15, SSZL16, YHZ18]. Planck/Stokes [HS18c]. plane [ACDT15, DS19, DGM16a, GH16a, GZD^{+}19, KM18, PDB15, SPD15, Tun18, TW19, YH18b, ZZW18]. planetary [LOAGBR15, NV19, XGH18]. planning [ASV17, DMAS^{+}16, RFLM16]. plans [FPGK19]. plant [JCC17, JCC18, JCC19]. plasma [NCPD19]. plastic [HR16]. plasticity [DJAMJ17, HRS16]. plate [Bej16, Che18, FdGKZ17]. Plateau [HL18a]. platensis [NGGFM16]. plates [MM15b, NTPVNTDT15, SM17]. Platonic [ADG19]. players [PW15b]. plus [LaTG17, WLZ17]. PML [HJL18]. PN [Gia18]. POD [DVJ18, LHLL18, LFKG16, SSZ19, WMI16].

POD-based [DVJ18, LFKG16]. Poincaré [HM16, Tun18]. Point [BDIR19a, DSOBK16, SK18, UG18, WHZ16, ABRS17, ARTY19, AHS19, Arà19, AGM15, ABB^{+}15, AB16, AP18, BR15a, BBC17, BGL17, CJP17, CDW15, CCZ^{+}18, CHX18, CT16, CJT15, DTKP17, DXML19, EHV19, FZW19, FSNS16, FGF19, GLTW15, GS15b, GW19, HW18a, HXC19, HS17c, HH18c, HH18d, HCC18, Iid15, ID18, KYY15, KKK16, KBL19, KS17c, LZ16a, LZ17b, LZ18c, LZZ19, LM17b, LC16, LZ19, Miy15, MF16, Prz15, QC18, RCN17, dHKdHMM15, RMMRC15, Rót15, RT18a, SHS^{+}15, SJ17, SJ18, SRH17, TBY17, WNML16, WZS19, WL19a, YZW15, YLL15, YZZL19, Zha15a, ZMTT19, ZWT19, ZM16b, ZL17b, ZG19b, ZwKN15]. Point-to-ellipse [UG18]. point-to-ellipsoid [UG18]. point-value [SRH17]. Point-wise [WHZ16]. points [Ahn19, AMB19, AS19b, AEHVM18, BP17a, DKLP15, EK19, EHVM18, GKN18, Har18, LKN17, LC16, MRMA15, MS19b, PAC15, uIZ15, WZ18, Yao16]. Pointwise [GK19b, MP17b]. Poisson [ABL^{+}19, BBQ18, CJCDLL15, CY18a, DFLM19, GLP16, GH17b, HS18c, IN16, JL16a, LSL18, MF16, QLZ19, SSZL16, TLTJ18, VX19, XY16, XD17, YB17]. Poissonian [HH19b]. Polak [YZ15c, YWY19, ZEBM17]. polar [CT17b, YLZG18]. poles [AAKW16, IANS15, IAAW18]. policies [EPC17]. policy [NH19]. policyholder [FJD17]. pollutant [LWC17]. pollution [AVGCMVM17, LWC17]. Polyak [YZ15c, YWY19]. polygon [CK16]. polygonal [BBG^{+}19a, Mu19a, Mu19b, WP16b]. polygons [AHS19, BDAG15, MS15b, ZWY16, Zyg17]. polyhedral [JL16a, PDO19]. polyhedral-based [PDO19]. polylinear [CP18]. polymer [AM16]. Polynomial [BIR^{+}17b, GZZZ16, LXW19, WZWJ19, AH18, BW17, BL17, BDG15, BLV19, CGMGV15, CV17, CH19c, DM16a, DSA16, FFP16, GS18a, GS19, GKM18, GV17, God15, GLP16, GR19, GM16b, GJ17, HMD19, HMMV17, HJS19, HW18a, HMP15, HB16, IANS15, JZ18, Kaz18, LDDF18, LP15b, Liu17, LX16c, Luo18, ML17b, MR19, MMD15, MSH18, MWY15a, NR19, PP18a, Pis15, PI19, RRM18, RRM19, SR15, SIAJ^{+}19, TXFH18, TV18, VŽ19, VZDV18, WCLL18, WWL15, ZD19]. polynomially [AH16b]. polynomials [AQCJ17, ALV19, ÁNPPCSA15, ÁAM15, AGR^{+}15b, ADG17, BM18, Baz15, Beh16, BWS17, CGN18, CBMO15, DGSM15, DV15, EDO19, EBJS19, FMP^{+}15, GZDC19, HS17b, HC17, JBT16, JY19, KS15a, KR16, KSR15, LV16a, LSS19, LSH^{+}19, LP18, MG15b, MMMMB16, Man15, MB15b, MT19, NN18a, NSM16, Orr19, PP18b, Qi18, QWZ18, RCPA15, ST16a, SAC17, WWM17, YZ18, Zag18, vD15]. polytopal [BAK18, CDE18]. polytope [YBD16]. polytopes [HAH17]. population [AALMLM18, AL18, ALMLM16, ALMLM17, BYYV16, BJK15, DV16, PCJ17, SKBT16, SS18d, Wan18d, WZBM18]. populations [Gun19]. pore [VMM18]. pore-scale [VMM18]. poro [ESI^{+}19]. poro-elastic [ESI^{+}19]. poroelasticity [BV16b, BV16c, HRGZ17]. porosity [KSW19]. porous [AAB18a, AEKGP19, AdQdS16, AFM19, CJCDLL19, CYY15a, CM19a, CFS19, CSW19, CSS19, DVJ18, ES18, EAKS18, FMMQ19, GT17, KLM17, LRX16, LLSZ17, LWV18, PG16, RNPK15, VCC^{+}19, WAGVRB19, YSCY17, ZWW^{+}18]. portable [GGC^{+}18]. Portfolio [ZG17, ZX16, BE19, JS19a, KV17, OAN17, WW19b, XDZZ19, YL18a, ZB18]. portraits [IL15]. posed [BEJR17, DdSL16, HRR15, HJ17b, MRVV16, MP15, MHMP16, PB15, XZXQ18]. posedness [AMZ19, DBX19, LMA18]. positioning [VRBS19]. Positive [CQW16, DP17b, HQ18, CLS18, CEJ16, CS19, KV16, MDG17a, MDGVR17, MM19, Maz18, MK17, Miy19, Nie15, PCJ18a, QG17, SLW17, WXZ19, ZW16, Zha18]. positively [BCFN16]. positivity [CL19b, CEJ18, KRR18, MDG19, Sta19, SN19, TYMG17, ZZLY16]. positivity-preserving [CL19b, CEJ18, MDG19, ZZLY16]. possibility [GX17]. possibly [FF15]. post [TX19]. post-processing [TX19]. posted [XDW18]. posterior [AR15b, Mu19a]. posteriori [AHMY19, BBGG19, BEMA16, BBG19b, Das15, DGM16a, GP16, GHL19, GSRJ18, HH16, LC17b, LF19, MTCC15, Mu19b, SWZW15, YGC19, ZCZM17, ZP19]. potential [DS19, GDQ17, GD17, RCBG17, dFG17]. Pouzet [ZM15b]. Powell [BI17, BK19a, SM15a]. Power [GAA19, GAA21, BCDwK17, BCVV18, DNK19, GB16, HWCC15, LWZZ19, NKD^{+}19, PD19, PG19, PP15, RN17, SVW19, Sho18, SBC16, GJSC18]. powers [BLP17, PRP19, Sad19]. practical [BE19, Das18]. pre [PC18a]. pre-exposure [PC18a]. precipitation [BBPR15]. Precise [Fuk15]. precision [NR15]. Precompact [SBCJ18]. Preconditioned [LS18a, AB18b, CGvA^{+}15, LHZ15, LFWC19, RCN17]. preconditioner [AFN17, CDW15, CMM17, CTZK18, CLW17, CM18a, CH19a, DLLB17, LP19, LZZ19, LN17, SHC^{+}19, WLLC18, ZZG^{+}19]. Preconditioners [ABB^{+}15, AB16, CGMM18, GC15, AKM18, BBC17, BFMP19, GHZ^{+}15, Hon19, HCC18, LZ16a, LZ18c, LSH^{+}18b, NZL17, OS15, VFGPG^{+}17, WNML16, WL19a, YOK15, ZL17b]. Preconditioning [WLZ17, AK17c, AL19c, BK19b, LHJR15, NSV18, YZW15, vSV15]. predator [YT16].

Predecessors [ADMV19]. predict [ML17b, MLW19]. predicting [HWP^{+}15, NGGFM16, NGGFM17, WS18]. prediction [CLZL18, GMSRP16, JA17, NGGA^{+}18, RABB19]. prediction-correction-based [CLZL18]. Predictor [Elg19, BZ19b, PS15, YLL15]. predictors [AH19]. predictors-responses [AH19]. Preface [APW^{+}17, CPS16, DGL16, HLV16, WWST18]. preference [YL18a]. preintegration [GKLS18]. premium [WP17]. PrEP [PC18a]. Prescribed [LP18, KLK16a, YB15]. presence [BV16a, ET15, MAF17, Sey18, TJA18, TDA18]. preservation [ABRS17]. preserves [BS18]. Preserving [ACDT18, BCS17, BGI18, RR16, AHT18, AK17a, Bal17, BM19a, BWS17, CL19b, CEJ18, CP16, GS18a, GS19, GZZZ16, GLZ18, Han15, HPW16, IJ18, JWZC18, KLK16b, LD19a, LD19b, LWZZ16, LMW18, MDG19, MDAP19, MN19b, NSW15, Sta18, Sta19, SN19, TYMG17, WZWJ19, Yos17, Zag18, ZZLY16, ZKRL15, ZL17c]. pressure [DD18b, LZRH19, MPM^{+}17, SRRJ16, ZM16a]. pressure-velocity [MPM^{+}17]. preventive [BYWPA19]. prey [YT16]. price [ÁRS16, ASV17, DT17, WZ16]. prices [DQIM18, ST16b]. Pricing [CP15, HW18a, JYK17, JSW18a, JQWW16, KC17, LM17a, NP19, PG19, BZ19a, CEJV16, DS17a, DM17b, DP19c, FCJ16, FNV17, GVHMR18, GB18, Hac18, HZ18b, HC17, HLL15, HT18, HXC19, HOM17, JYP18, KDB17, KDB18, LN16, MZ16, MNFS18, MFLL18, NM17, RP16a, RZ19b, SM18, SCG18, SX18, TYMG17, TEG15, VFG18, WH16, Wu19, XW18, ZS17a, ZYY19, ZL15c, dSBV16]. Primal [CDZ16, SBCJ18, KYY15, KCX17, TBY17, WWWZ16]. Primal-Dual [SBCJ18, CDZ16, KYY15, TBY17]. primary [SRRJ16]. principal [BB16, CYY^{+}16, HS18b, HX15, KW16, Miy18]. principle [DM16b, GB16, MS15a, Sun16]. principles [Mej17]. priori [AW19a, Das15, GYX17, Mu19a, ZP19]. priors [AMM19]. Probabilistic [CCNQ^{+}16, CDRY18, Ber15, CCNQ^{+}17, HWP^{+}15]. probabilities [ABP15, CYY15b, Tra15, YZJC15]. Probability [AM17b, EWR18, AALMLM18, CNQRR18, DSA16, EG17, FM18a, FL16, GLP16, GWY17, Kaw17, LWZ19b, RVZ15, SRA^{+}15, YZY17]. Probability-one [EWR18]. Problem [BF18, PBS15, ALS18, ACLY19, ALS19, AB18a, Aks17, AFBJGJ17, AGM19, ACCF19, AMZ19, BGMS16, BBD16, BBGG19, Bay15, BFLP15, BR15a, BCA18, Bog19, BZ16a, BySZ19, BHMT16, CZ17a, CL19a, CHX18, CDD^{+}15, CCX18, CT17a, CKL18, CG17b, CLV16, CKB17, Cou19, CPL^{+}18, CPL^{+}19b, DLLB17, DVS15, DWQ16, Dia18, DGM16a, DFRR15, ESAE18, ESI^{+}19, FLJT18, FLZ19, FS15, FP16a, GM16a, GCY18, GYX17, GLT16, GO17, GS15c, GS15d, GG17b, GP16, GM19, HZ18a, HH19a, HMS19, HWZ18a, HMH18, HOTZ18, ID18, ICMPG16, JM15, JHK17, KM18, KR17, KFT17, KLC17, KV16, LT18, LRT17, LKS17, LaTG17, LSD15, LZC15, LA15c, LLCG16, LLSZ17, LZW18, LGLC18, LGYC19, LSYY19, LDL^{+}19, LWZZ19, LHM16, LCZ18, LWJ15, LZZL17, LLW17, Liu18, LF19, LX15b, MYG18, MP15, MVC19, MM15b, MKKV19, MV16c, MMRS15, Ngo19b, OP19]. problem [OCU17, OCU18, OZ17, PMN19, PT17a, PGD^{+}18, PCSR19, Qui15, RVC19, RCN17, RFO17, RMK19, RR16, RR18, SGK16, SWL15, SYZ19, SWZW15, SQRE16, Slo15, SVV18, Su16, TBR16, TV16b, TF18, TN19, THL^{+}17, VDS15, VMR17, VM15, WWZ15, WPA^{+}18, Wat16, WD16, WW16b, Xie19, Xu17, XCG19, Yak19, YW16, YBLH17, Yan17, YSG18, Yua15, YSL^{+}18, ZCH18, ZM15a, ZZZ^{+}17, ZHS18, ZYZ18, ZQJY18, ZL18b, ZHC19, ZX16, ZC16b, dFG17]. problem-dependent [ICMPG16].

Problems [FMC^{+}18, AAA15, AAM18, ABH19, AKR18, AS18, Ais17a, AZ18, AA18, AT17, AM19, AMCR19, AW19a, AK15, ACEJ18, AGR^{+}15b, AC17b, ABB^{+}15, AB16, AFN17, AL19c, AP18, BS18, BYW18, BD16, BS16, BJK15, BK15a, BAO^{+}16, BGNGR^{+}17, BBC17, BKK18, BEJR17, BKLR15, BKLR16, Bog16, BW19, BK19b, BV16b, BV16c, BHN15b, BN15, BS17c, CDW15, CCZZ16, CWXW17, CDP17, CCJ16, Ceb17, CLX17, CM16, CH18, CS18a, CHR^{+}15, CW16b, CDRY18, CH19a, CLVW18, CL19d, CV19, CGSS17, CP18, CDKN14, CDKN15, CW17, CGvA^{+}15, CN16, CTAT19, CL15, DKP^{+}19, Das15, DVA19, DdSL16, DTKP17, DFT18, DZ18, DVJ18, DWXO15, DK16, DYZ18, DCNY18, DM16b, DD16c, Dzh16, Dzh18, El 17, FCJ16, Fan15, FYY17, Fer15, FGF19, FMR15, GHM18, GWLL18, GHA^{+}15, GDQ17, GD17, GN18, GS15b]. problems [GO15, GW17, GJSC18, GLZ18, GSRJ18, HLBP15, Haj19, HOS16, HWCC15, HL18a, HBH17, HA19, HWT15, HK07, HR16, HMP15, HRR15, HNT19, HC15, HWH19b, HS17c, HJ17b, HH18c, HH18d, HCC18, Iid15, Jeo15, JHPT15, JL17, JL15, JO17, Jin16, JLLZ16, JGGS16, KCK18, KF16, KCX17, KKP^{+}19, KPS16, KMLB17, KT15, Kre19, KKK16, KR18, KK15, KtTBK18, LY15b, LBL15, LHJ16, LX15a, LX16a, LQZ16, LH16, LLX17, LW17a, LY18, LLZ18, LCH18, LZ18b, LWDW18, LZ16a, LZ17b, LZ18c, LZZ19, LC15b, LC17b, Liu17, LLHO18, LW19b, LMA18, LM17b, LSP18, MTV17, MCLW16, MJ18a, MTCC15, MRVV16, MD19, MR19, MHAA15, Mat18, Mit15b, MY18, MZ19, Mu19a, Mu19b, MT19, MHMP16, NHA17, NS19, NT15, NSV18, NN18b, NN19, NXLX18, PS15, PRRY18, PRP19, PTV17, PLP17, PCJ18a, PCJ18b, Ple17, QWS15, RMVA17, RS19]. problems [RP15, Reu16a, Reu16b, Reu17, RW16, RT18a, RMA19, SBU^{+}17, SKB15, SGL17, SVW19, SSW19, SHC^{+}19, SHW19, SY19, SS17, SK15, SZ18, Sun19, SI16, TGJ19, TAT18, VK16, VCC^{+}19, VCL^{+}19, VDT19, VAR15, VARC17, VF17, Vis15, WYW15, WXW15, WWWZ16, WNML16, WZS18a, WL19a, WWH19, WP16a, WL16, WHWY18, XY16, XZXQ18, XDW18, Xu18, XZN^{+}19, YP18, YZW15, YZ16, YZZL19, YB15, YGC19, Yin19, Zak19, ZZ19, ZNG15, ZL15a, Zha15a, ZHWS16, ZYXT16, ZXQH17, ZCZM17, ZLZ17, ZL19, ZWT19, ZYL19, ZCL16, ZL15b, ZM16b, ZL17b, Zho16, ZC18, ZG19c, ZP16, vSV15, vV16]. procedure [CCT19, EL19, WL17c]. procedures [JK19]. Proceedings [BLLM^{+}15]. process [ALS18, BCNP15, CCY18, CCC19, CY18a, DAM15a, JA17, NM17, RP17, Ruj16, STZ16, WRZ18, YS19, YC19, ZG17]. Processes [GMSR18, GMSR21, AS19a, BGMS16, FCJ16, GVMR16, GVHMR17, GVMR19, Hai17, HVMR19, Hui16, MX16, ML16, NP19, PA19, RP16a, STZ16, WSW15, WZ16, Ugu17]. processing [GDC15a, MMPGCSP17, TX19, ZCC15]. produced [Mia18]. product [BFLP15, Ceb17, CS18a, CLM19, DFS15, FMP^{+}15, Hag15, HYQ^{+}18, HOTZ18, Hua19, LZZ19, RT18b]. Products [Mar19a, OD17, NT15, PZ18, TGSS17, VS18, WL17c, WLC^{+}18, Wu18, WH19]. profile [LWC17]. program [LXL15, ZqZwZ15, ZL17a]. Programming [dMPF15, BCFN16, CCIL18, CLZL18, CLZS19, CDE18, CP18, HP15, KYY15, LXL16, LZ16c, MB17, MLS15, MHC16, NTPVNTDT15, NN18b, NN19, Sun16, TBY17, ZC16a]. programs [FZ15]. Progressive [FMC^{+}18, LLG^{+}18, EL19, NAK17, Wan15b, WL17c, WLC^{+}18, Wu18, WH19]. progressive-iteration [EL19]. progressively [ABB18, BSLL18, BSLL19, Gun18, Wan18e, WS16, YZPL18, ZSW16, ZG19a]. project [FZ15]. Projected [FwC17, HMD19]. Projection [DK16, NDN19, Ais17a, AALZ15, BW19, DNL15, DNL16, DMMSK17, GD16, GKN16, JMNB18, JM15, JCH17, LDG19, MN17, MN19a, PS18, US19, bWzZjH16, YY16, Yin19, ZS18b, ZQJY18]. projections [GH16b, Per16, PBGG17, SS16a]. Projective [HJS19, BW17, MG15a]. projectors [DDR19]. Prony [Skr17]. Prony-like [Skr17]. propagation [BN19, CROC16, DLM17, FH16, FH18, FM17, GM16a, GTX^{+}19, JCC17, JCC18, JCC19, KSW19, LMS17, MVdRE^{+}17, Per16, SODBK19]. propensity [Lu18]. Proper [FMC^{+}18, KVO18, SPD15, hShX17]. Properties [CAO^{+}19, Lu15b, NADK18, NZ16, AH16a, BRS15, BP17b, CL19b, CJ16, lCCVV18, CP16, DP19a, DP19b, DXML19, FH18, HPT19, HCC18, KPS16, LV16b, Meh19, MMZ19, MSH15, Ogu17, QCSG18, TBS18, WNML16, Yos17, dR16, vD15, COGD19]. property [CQW16, DM16a, QDBG18, SM15b, Yam19]. prophylaxis [PC18a]. Proportional [MTMZ19, CSB18, HYZ16, TF18, ZC19]. proportional-derivative [CSB18]. protection [JYP18]. protections [JQWW16]. prove [HKAR18]. Proximal [MMRS15, YSL17]. PSAI [JK19]. pseudo [ACAC17, DTKP17, DXZ18, LZ18b, Su16, XZXQ18, yZ17]. pseudo-contraction [DTKP17]. pseudo-differential [XZXQ18]. pseudo-heuristic [LZ18b]. pseudo-Jacobi [Su16]. pseudo-parabolic [ACAC17]. pseudo-spectral [yZ17]. pseudo-time [DXZ18]. pseudodifferential [LL17]. pseudomonotone [TH18]. pseudopotential [SSD16]. pseudospectra [FV17]. pseudospectral [CLC16, Elg19, LZ15, MHAA15, NDH16, SLH15]. pseudospectrum [NR19]. pseudostress [BBGG19]. psi [QL19]. PSO [NGGFM16]. PSPIKE [ZS17b]. PSS [CDW15]. PU [MFLL18]. pulsating [GCTGC16]. Pulse [HS17b, XHZ17, Beh16, GO17, Gru16, MAS19, NE17, SOMF19, YT16, XY19]. pure [DGM16a, RFLM16]. purely [GKN18, Mus17]. Purkinje [AMZ19]. Purkinje/myocardium [AMZ19]. purpose [BBM15]. pursuit [KC19, WYLD19]. put [CEJV16, JYK16]. PyFly [GGC^{+}18]. Pythagorean [BLV19, FW18, KLK16b, LAHB15, LFCW19, ZWY16]. Pythagorean-hodograph [FW18, KLK16b].

QN [HLBP15]. QP [JGTB15, VMR17]. QR [HZ17b]. QUadratic [BGI18, BCT16, PBS15, AK17a, BCFN16, BK19a, CCJ18, DKV16, GK18b, Haj19, HKAR18, HWH19b, KYY15, KS15a, LA15a, LXL15, LWZ^{+}15, LXL16, LFWC19, LX15b, MA18, QWS15, SST15, SL18, WL19b, Xie19, YSG18, YZ18, ZL18a, ZL17a, ZWL18]. quadratically [Abd16, ZL17a]. quadratization [YZH18]. quadrature [AHBC15, AN17b, BGII15a, BGII15b, BC16, BAHC17, BK19a, BCCB15, CFSS18a, CBMPP15, DRS16, DRST19, EJJ16, FEC18, JR15, JRT17, KW16, KB19, MDT16, MIT15a, PM15, PC18b, PR15, Pul17, SE16, uIZ15, TB18, Wan18b]. quadrature-based [MDT16]. quadratures [CDP19, LL16b, MVPN15]. quadrics [CFJMVG16]. quadrilateral [AM17a, Che18, LSLZ17, YW15a, YW15b, ZCL16, ZML16]. Qualitative [FH18]. qualitatively [FH16]. quality [GLTW15, LZ19, SPG^{+}16a, SPG^{+}16b]. quantification [BK19d, CTAT19]. quantile [MS17d, Xie15]. quantitative [Mej17]. Quanto [TEG15]. Quantum [CRR19, GS15a, PBS15, BDGO15]. quartic [AIL17, BIR17a, BDIR19b, GK19a, HG17, IL15]. Quasi [BDIR19b, CZ17a, FP18, God16, LSDY19, MOY18, Nak15, AS19b, BGII15a, BIR17a, BDIR19a, BIJM^{+}19, BGNGR^{+}17, CFSS18a, CMHHS17, CBMPP15, DDR19, DCNY18, ELL^{+}19, El 17, FNZ18, GCTGC16, HLV^{+}15, HL18a, Iid15, JW17, KW18, KB16, LH19, QC18, RKB16, SST15, SYZ19, TBAH16, WWL15, XW18, ZYY19]. quasi- [TBAH16]. quasi-2D [GCTGC16]. quasi-consistent [KW18]. quasi-equidistant [AS19b]. quasi-interpolants [BIR17a, ELL^{+}19, SST15]. quasi-interpolating [DDR19]. Quasi-interpolation [BDIR19b, FP18, BGII15a, BDIR19a, BIJM^{+}19, CFSS18a, KB16, WWL15]. quasi-interpolations [ZYY19]. quasi-linear [BGNGR^{+}17]. quasi-minimal [LH19]. Quasi-Monte [God16, MOY18, Nak15, XW18]. Quasi-Newton [CZ17a, LSDY19, FNZ18, HLV^{+}15, RKB16]. quasi-nonexpansive [Iid15]. quasi-optimal [JW17]. quasi-optimality [SYZ19]. quasi-periodic [DCNY18]. quasi-Plateau [HL18a]. quasi-static [CMHHS17]. quasi-tangential [QC18]. quasi-variational [El 17]. quasilinear [MHV17, MPH18, Pol16]. Quasistatic [VMR17]. quaternion [JWZC18, LWZZ16, LWZZ19, TXFH18]. quaternionic [ST16a]. query [MMRS15]. question [DV15, Mär17]. questionnaires [GLdCJS17]. questions [LSS19]. queues [TA17]. Quintic [LM17b, LX16c, BAHC17, FW18, GLV15, Lu15c, TXFH18]. quintic-polynomial-based [TXFH18]. Quintin [GRVC15]. quota [ZJQW18]. quota-share [ZJQW18]. quotient [MMRS15, Yos17].

R. [DLY18]. raceway [NGGFM16]. Rachford [HOS16]. radar [BTKM15]. Radau [MVPN15]. Radial [NRV18, YBC^{+}16, Zha17a, AD18, FSNS16, GCTGC16, HW18a, ID18, Jeb19, KDB18, MAK19, SDH17, SJ17, SJ18, SVAKGR19b]. radial-polynomial [HW18a]. Radially [Bej16]. Radiation [FdGKZ17, CFSS18b, LP19]. radiative [CH17b, GRS19, Gia18, SWH16]. radiative-conductive [GRS19]. radius [SNY19]. Radon [GH16b]. railway [MCGNBD17]. RAMs [BIJM^{+}19]. Random [NRV18, YBD16, AK16, BD19, (Ba18, BCJV18, BCVV18, CJCDLL15, CCJ16, CCNQ^{+}17, CCJ18, CZLV17, CNQR^{+}18, CDHP17, DEY19, FF15, GC18, GAB17, JN19, LSL18, Luo18, MM17, NL17, NC19, NCF^{+}19, OLPT16, PP16, Pul17, RIK16, RdSMO17, RdSMO18, YL19b, ZLPF19, ZWZ^{+}18]. randomized [YCW17]. Randomizing [CNQRR17]. randomly [LDW17]. Range [SHW19, BRS15, SS18a, ZZW18, ZP16]. Rank [LLW17, AGM19, AB18b, BLXD18, BDG15, BL16, CGMM18, CPL^{+}18, CPL^{+}19b, DY15, GH17a, GH19b, GW19, HS18a, Hua19, LDZ17, LSH18a, MOPP18, NS16, SWL15, SHC^{+}19, WGLV16, WCW18, WXJ19, YFwC19, ZCD17]. rank-correction [WXJ19]. rank-deficient [GW19]. rank-one [WCW18]. rank-structured [BDG15, Hua19]. rank-two [ZCD17]. ranking [GX17]. rankings [CGPR16]. Raphson [SVW19]. Rapid [BEHL17, PKR18, Ngo19b]. rapidly [BR15b]. Rate [dlCYC15, AL19b, EL19, GVMR16, GVMR19, HZ18b, HLM18, ID16, JC15, KHLY17, LZ17a, LLZ16, LWZ19a, Mok15, NMKK19, NN17, RZ19b, Tan18, Wan18b, You19, ZCC15, ZS17c]. rate-type [LLZ16]. rates [BZ19a, HK07, LX18, LW19a, LLY19, MYG18, Mao16, PX17, TGJ19, ZS17a, dFG19]. ratio [GSM19, PP19, Qi19]. Rational [FGS19, GLSW18, LLHO18, AHP19, AHJ17, ACDT15, Arà19, Bal17, BHN15b, CFJMVG16, Ceb17, CW16b, CH19b, DMP19, DLMZ19, DD17, EDO19, Fuk15, GS19, GG17a, HJS19, IANS15, IAAW18, JSC18, LP15a, LA15b, MMPGCSP17, NSW15, PD17, SJ15]. rationality [CEJV16]. ratios [LKN17, Mar19a, SHW19]. ray [PGD^{+}18]. Rayleigh [WLC^{+}18, ZG19a]. Razumikhin [GL19]. Razumikhin-type [GL19]. RBF [DMP19, DP17b, FP18, MFLL18, RRS19]. RBF-based [DMP19, DP17b]. RBF-PU [MFLL18]. RCLL [DL16b]. RCP [SPG^{+}16a]. re [FLM17, LZS15]. re-initialization [LZS15]. re-visited [FLM17]. reachable [RRW17]. reaching [LWZ19b]. reaction [APYW19, AMCR19, AK19, AAKW16, BK15c, BK16, Bog19, CFX17, CJ18, CEJ18, DP16, DF15, DVA19, DAM15a, DAAF16, GHM18, GPHAPR15, GPHAPR17, GO15, HSWH15, KK15, Lee18, Lin16, LK18, MR15, RC19, Rät15, RR16, SR16c, SS16c, SJ18, SB16, SZJ19, ZKRL15]. reaction-diffusion [Bog19, CJ18, Lee18]. reaction-subdiffusion [DAM15a, SJ18]. reactions [CSE^{+}19]. Reactive [CSE^{+}19, OSC15]. reactors [OSC15]. Real [LWZZ16, MG15b, AdHR17, AAKW16, AKM18, BLH^{+}17, CNQRR17, GOGM17, GB16, HW18a, IAAW18, JWZC18, KS15a, LDDF18, LSS19, PW17, SH18, SH19, TBS18, WG18, ZS18c]. real-time [BLH^{+}17]. real-world [AdHR17]. realistic [XDZZ19]. Realizable [OZ18]. realization [BCMP16]. realizations [Bai15]. rebates [LDK17]. reciprocal [QG17]. recognition [LZ19]. Recombined [HXC19]. reconstruct [RVC19]. reconstructing [Che16b]. Reconstruction [CL18a, CS18b, GS15d, Hag15, Mok15, TKMT17, WP16b, ABK19, CGvA^{+}15, DVS15, GLTW15, GSRJ18, KGL19, LSG18, LSYY19, LD15, SAH17, THC^{+}18, UPD^{+}15, ZCC15, ZLAA15]. reconstructions [AMM19, Vej18]. record [PD19, Wan17a]. records [RABB19]. recovered [MG17]. Recovering [KV19b, dAFvdMR19, BBQ18, LWC17]. Recovery [FM18a, MS17d, TNTL18, Zak19, ZY19b, DGH15, GZZZ16, JC19, KC19, LSH18a, LC19, LXZ18, Skr17, WYLD19, WZWJ19, YP17]. recruitment [DT16]. recruitment/renewal [DT16]. rectangular [CZ17b, DAM15b, HC15]. recurrence [CHVRT15, CBMO15, REBF16]. Recursive [ADG17, Liu16, Nat19, JB18, QWZ18, RMA19, SY18, WS19]. recursively [pWmGkY^{+}19]. rediscovery [GD18]. redistribution [DKP^{+}19]. Reduced [JL16b, BMLR15, CH18, DKLP15, DSN17, EK19, LHLL18, SSZ19, SJL16a, SMS18, XWWI18]. reduced-order [DSN17, LHLL18]. Reduction [AvG16, AHJ17, ABB17, ABG17, AMCR19, AC17a, BDSB16, CR17, CTZK18, CST^{+}16, CVW17, CLVW18, CMRP17, CD15, CTAT19, Dan17, FR19, HL15, JO17, KVO18, LWV18, LFKG16, LX17, MFK^{+}15, MWY15a, Per16, Rad17, SB19, SP17, SV16b, VMM18, XW18, YZ15b, AdGvG19]. Reductions [CRG19, RRM18, RRM19]. reductive [SZM^{+}18]. Reeves [JJ19]. refinability [CPRS19]. refinable [CJ16]. refined [JL15, NML18]. refinement [DG16a, JMNB18, KS15a, KHS^{+}18, LW15, LL16a, LW17b, MII16, MS15b, SM15b]. refinements [MPS19]. refining [FwC17]. Reflected [DL16b, HMKCR19]. reflection [CJ16]. reflexive [BKLR15, GLT16]. regime [BBD16, DF16b, DT17, Gor16, JQWW16, KKW16, KV17, KS17b, LZ17a, MZ16, MGMDGVR17, RP16a, Su18, WW19b]. regime-switching [KKW16, KV17, LZ17a, WW19b]. region [ESAE18, HP15, RKB16, RS16]. regions [AC17b, CDF^{+}16, MFWK18, PR16, SBU^{+}17]. registration [TC19]. regression [AH19, AdHR17, BLH^{+}17, CCC19, GYQS15, LSH^{+}19, MG17, SZM^{+}18, TW17, Xie15, YZPL18, ZPL17, ZPL18]. Regula [SV19]. regular [APS15b, APS17b, APS17a, APS19a, APS19b, KHLY17, SS16a, Tru19, Vis19, WZZ18]. regularity [CBC^{+}19, GH16b, LLY19, SZ18]. Regularization [HRR15, KR17, KH18, BBH^{+}17, CPL^{+}19a, DWQ16, GH16c, HZ17c, IGRL17, LWZ^{+}15, LH16, LF19, LSL15, MRS17, PRRY18, PB15, SV19, SWZW15, THC^{+}18, WWZ15, WWH19, XC18, XZXQ18, YW15c, ZqZwZ15]. Regularized [LLD18, ABB^{+}15, AB16, Che16b, CSXY19, CP18, DAM15b, LZ18b, LLG^{+}18, TW17, WSL19, ZPL17, ZPL18]. regularizer [GS18b]. regularizing [DdSL16, HJ17b]. regulation [You19]. reinforced [RL18]. reinsurance [ALS18, HYZ16, HOTZ18, HY19b, LRZ15, WP17, ZJQW18, ZC19]. reinsurance-investment [LRZ15]. reinsurer [HOTZ18, LRZ15, WRZ18, ZJQW18, WP17]. reinsurers [MSY16, MSY17]. related [ÁNPPCSA15, BKLR16, BCA18, CDD^{+}15, GGG19, Hui16, ML17a, Meh19, PS15, SB19, WS18, Yak19]. relation [CLM19, SvG15, Sun16]. relations [CHVRT15, CBMO15]. relationship [OSMZ18]. Relative [SZM^{+}18]. relativistic [KS17b]. Relaxation [CKL18, AL19b, DF16a, JMN^{+}16, Lad15, LZL^{+}18a, YZH18]. relaxed [CDW15]. release [YBC^{+}16]. Reliability [BSLL18, BSLL19, ET19, FOK15, GMSR18, GMSR21, Kam17, LSZB17, OE19, YZPL18, BBM15, ET15, Ery19, FPGK19, Gun18, LSBL18a, LSBL18b, MM17, RdSMO17, RdSMO18, ZG19a]. Reliable [Gia18, CEJ16, HRS16, MT17a, WR16]. remaining [GUB15, OLRPdCJ19]. remark [ZT18]. remarks [LJX17]. remediation [YBC^{+}16]. remeshing [NLG15]. remodeling [CCD16]. removal [CDZ16, Lu16]. removals [WS16]. renewable [BGNGR^{+}17]. renewal [BRA16, CLHMT15, DT16, FL16, YZJC15]. reparametrization [SPD15]. Repeated [DXML19, CN18, VL19]. Replacing [FGF19]. replenishment [EPC17]. replicating [CDD^{+}15]. represent [Kre19]. Representation [HH15, OD17, COGD19, Kam16a, PP18a, TYQ^{+}19, ZH16]. representations [BRZ15, CNT17, KGL19, MT17b, QL18, Qi18]. represented [MGMGJLRB17]. Reprint [BKMS18b, XY18b]. Reproducing [AVB17, SAA18, AAA15, AA16, AP18, BJM16, BH17a, CDHP17, JY19, JW17, KMB15, LW17a, ML19b, Sak17, VFG18, WGHR17, XZN^{+}19]. reproduction [CH19c, DM16a, WWL15]. research [XY19]. researchers [DLY18]. reserving [DT18]. reservoir [LZRH19, NGGFM17, hShX17, WLLC18]. reservoirs [ZHY^{+}17]. reset [BIJM^{+}19]. Residual [MTCC15, YCF18, BO16, BBG19b, DGM16a, GAB17, GHC16, HHS^{+}15, HW17, HLLY19, JL15, LH19, PT17b, RW15, YSG18]. Residual-based [MTCC15, DGM16a]. Resistive [BIJM^{+}19, ARCA^{+}19, LSH^{+}18b]. resolution [GCTGC16, HWH19b]. Resonance [EGK17, ACEJ18, QLT^{+}17]. resonant [NCPD19]. resonating [YOK15]. resources [BGNGR^{+}17]. respect [LLW17]. response [ÁRS16, BdSFL16, BHU18, LRRV18, SH15, ZP16]. responses [AH19, SM17]. Restarted [GHY^{+}18, AvG16, NML18]. restarters [MAF17]. restarting [CSB18, xZWlC15]. Restoration [AEP15, BBM15, BEJR16, CFR18, CG17a, KKP^{+}19, LSL15, PKMM19]. restoring [HLG^{+}17]. restrained [CS17]. Restricted [BCNP15, Wu16a, BRS15, ZZW18]. restriction [Xia17]. Restrictions [ADG19, HB16, ZLPF19]. results [AM15a, CGN18, FZW19, GH16b, GAB17, GWY17, HHKR15, KS17c, KCNH16, LXL16, LSS19, LLW18, MN19a, MT15, NSV18, Pis15, RP17, SKC^{+}19, STZ16, SHS^{+}15, SLQ16]. Resuspension [ZFK18]. retardation [SYD15]. retarded [BEHL17]. retinal [CNJM18]. Retraction [Beh16, GAA21, HWXC19b]. retrieval [LZW18, PGD^{+}18, YWW19]. return [ID16, YZJC15]. returns [FL16, GWY17, KElY16]. revenues [RVZ15]. reversed [BD19]. reversed-hazard-based [BD19]. reversible [AHK19, CSE^{+}19]. reversion [HOM17, ID16]. review [CDN19, OT17]. Revised [FYY17]. Revisit [CKKJ15]. Revisited [AIL17, BN15, DS19, Mir15, SZ16]. revolution [VL15]. reweighted [LSH18a, ZLAA15]. Rhythm [LZL^{+}18a]. Ribière [YZ15c, YWY19, ZEBM17]. Riccati [CCNQ^{+}16, wCW15a, FwC17, HP17, LC18, Miy17, Miy19, Sak17, YFwC19]. rich [SQRE16]. Richards [CYY15a]. Richardson [BR15a, ZDF^{+}17]. Ricker [You19]. ridge [AH19]. Riemann [AÖ17, BD16, JLLZ16, SF18]. Riemannian [CRS^{+}18]. Riesz [CJLT15, THL^{+}17]. right [ATG18, HHOT17, KH18, LWZZ19, MLJ16, MT17c, REBF16, xZWlC15]. right-hand [ATG18, HHOT17, MLJ16, MT17c, REBF16, xZWlC15]. rigid [ABT16]. rigid-lid [ABT16]. rigidity [MXW17]. Rigorous [GLR16]. rings [AG16a, NV19]. Risk [ZX16, ALS18, BBB17b, BRA16, CCY18, CPOG18, CLHMT15, CLPZ17, DYD19, EG17, FJD17, FL16, FY18, GZH19, GSM18, GVMR16, GVHMR17, GVMR19, GWY17, HYZ16, JYP18, LZY17, MX16, MSY16, MSY17, RP17, SSKV17, Ugu18, WP17, WRZ18, YZJC15, YZY17, YCW17, ZB18, ZG17, ZNS18, ZS19, ZS17c, FM18a]. risk-minimising [SSKV17]. risk-neutral [GVMR16, GVHMR17, GVMR19]. risks [DSDV15, Nke18a, Wan18e, WS16, ZSW16, ZC19]. risky [FY18]. Ritz [FJLB18]. RK [WLM16]. RLW [WD18]. Robin [LX16b, SYZ19, SGG19, SVL19]. Robinson [KS16]. robot [LMUZ18]. Robust [BV18, GSM19, HOTZ18, JJX17, LYG15, SJL16b, Ugu18, ZC19, AH19, CFX17, CC15, Fer15, LWZ19b, MTV17, RNPK15, SOMF19, TC19, yZyMxG^{+}16]. Robustness [DKV16, GM16b]. rock [KMD^{+}16, KEB19]. rod [WD16]. Rodrigues [FGS19]. role [GSVSM16]. Root [KS15a, BW17, CT15, GZDC19, Kow17, LKN17, Miy18, PN16, PP19, RP16b, YPD16]. root-finders [PN16]. root-finding [BW17, CT15, RP16b]. root-ratio [PP19]. root-solvers [YPD16]. rootfinding [BDG15]. roots [BGMM18, CN18, HH18b, PN16, SH18, ZCST18]. Rosenau [WD18, WD19]. Rosenau-KdV [WD19]. Rosenau-KdV-RLW [WD18]. Rosenbrock [Ran15]. Rosenbrock-Wanner [Ran15]. rotating [Bad18, DLW16a, OGR19]. rotation [FGS19]. rotation-minimizing [FGS19]. rotationally [Pap16b]. rough [XDW18, ZM15a]. Rounding [LX15b]. Route [RFLM16]. routing [SQRE16]. row [LDZ17, dlCYC15, ZS18b, JS17]. row-rank [LDZ17]. ROW-type [JS17]. RSAI [JK19]. RTE [THC^{+}18]. RTE-based [THC^{+}18]. ruin [BRA16, CYY15b, CLPZ17, EG17, FM18a, FL16, GLP16, GWY17, RVZ15, Tra15, YZJC15, YZY17]. rule [LZ18b, LWZ17, LF19, PM15, PC18b, RT18b, SBP17, VS18, ZJG^{+}15]. rules [AHBC15, BC16, BAHC17, CFSS18a, CBC^{+}19, DRS16, DRST19, God15, JR15, JRT17, MIT15a, OLPT16, Pap16a, Pap16b, PR15, Pul17, uIZ15]. run [KSDM16, OE19]. run-up [KSDM16]. Runge [GPHAPR15, ZM15b, AH15, AK15, AK17a, BYW18, BP17b, CWH18, DT16, GPHAPR17, GQM16, HHR16, HIS16, KS18, LL15, MVK19, MAF17, Son16, SV16a, TFS17, WZS18a, ZST15, Zha16, ZZHS17, ZDF^{+}17].

S [KEB19, ZWWW19]. S-ROCK [KEB19]. Sabin [BI17, BK19a, SM15a]. saddle [ABB^{+}15, AB16, BBC17, CDW15, GW19, HXC19, HS17c, HH18c, HH18d, HCC18, LZ16a, LZ17b, LZ18c, LZZ19, Miy15, RCN17, WNML16, WL19a, YZW15, Yao16, Zha15a, ZWT19, ZM16b, ZL17b]. saddle-point [GW19, HXC19, LZ17b, YZW15]. Sakurai [PI19]. sales [HEA18]. saliency [JWQ18]. Salpeter [GLZ18]. salt [KMD^{+}16]. saltwater [ZYZ18]. same [LA15b]. Sample [SP17, AFBJGJ17, GK19b, HWT15, NCF^{+}19, WP16a, Wu18, WH19, ZqZwZ15]. Sampled [AG18, APYW19]. Sampled-data [AG18, APYW19]. samples [GLTW15, WP16b]. Sampling [JW17, ABT17, BFLP15, Che16b, DF16b, Kaw17, LWXF18, LXZ17, LZ19, PP16, YZZL19]. Samuelson [LPT17]. San [GRVC15]. SAR [CCRN19]. SAT [GDCK19]. satisfaction [GR16]. SBP [GDCK19]. Scalable [SM16, vSV15, BV18, LLZ16]. scalar [Had15, HLMS19, RVA15, SRH17]. scale [And16, BHRR18, BH17b, CG17a, CKKJ15, wCW15a, DSN17, DY15, DCNY18, DFRR15, EAKS18, FFKP17, FM19b, GC17, HLG^{+}17, HJ17b, HWZ18b, HWXC19a, HWXC19b, IDGH18, JHJB17, KElY16, KSD17, LSH^{+}18b, LF18, LL18, PM15, Per16, RKV19, SA17, SM16, VMM18, VFN18, WLLC18, WP16a, YFwC19, YZ15c, Zha17b, ZDL16]. scale-invariant [VFN18]. scale-SOR [HWXC19b, HWXC19a]. scaled [And18, FM18a, OP19, SSH16, SSG18]. scales [FR18a, LPT17, Ogu17, XH15]. scaling [And17, BS18, HZ17c, KLM18, YN18]. scanning [LDZ17]. scatter [SGK16]. Scattered [QWZ18, BIR^{+}17b, DD16a, FP18]. scatterer [LD15]. scatterers [LXZ18, YL18b]. scattering [DFT18, GH16a, GH19a, JHPT15, KM18, Kre19, LP19, LLZ18, ZM15a, ZCFW17]. scenarios [LE18, SPG^{+}16a]. schedules [IABIPG17, KCDACC17]. Scheduling [QDG^{+}15, HYQ^{+}18, Xu17]. scheme [ALC19, AA17, ABRS17, AAKW16, AJ17, AD18, BS18, BGMM18, BCS17, BK15c, BK16, CRCM16, CLX17, CS18a, CW15a, CDZ16, CLC18b, CFWW19, CGSS17, Col16, CEJ18, CP16, DTKP17, DBdLdM^{+}19, DSOBK16, DY19, DL16b, Dun17, DM17b, DP19c, FP18, GCTGC16, GRS19, GGRN17, HK19, HCL17, HW18b, IDGH18, JK18, JY17, JMS17, JN18, KHLY17, KV16, KSW19, KKK16, KBL19, KAS18, LSD15, LCH18, LSP18, MB15a, MV16b, NAK17, RNPK15, RP16a, RZ19a, SKBT16, SFGJR^{+}16, SYD17, Sta18, Sta19, SF16, Tan18, TZZ18, TF18, Vab18, WCW15b, WHZ16, WWZZ16, WD18, WLW19, WD19, Wu17, Yam17, YB17, YGC19, ZW16, ZMTT19, ZL17c, ZZ18, iW16]. schemes [AH16a, ARTY19, AH15, ACDT18, AK17b, BVZ17, BBC17, BKLR15, BP17b, CM19b, CJLT15, CFSS18b, CH19c, CSTU17, CDHP17, DM16a, DXML19, DS17b, GGG19, Gri19, GG15, HHEG17, HMDSR19, JB18, Jeo15, JY19, JC15, KBL19, KPS17, LSZ16, LLZ18, LW19c, MG15a, MC19, MVdRE^{+}17, MVQDdR^{+}18, Mat18, MHV17, MP18, MS17b, MV18, MV16d, NR15, SuSR16, SM15b, SW17d, Wan18a, WSL19, Xin16, XZ19, YZH18, Yos17, ZZW16, ZJ17a, ZWGL18, Zha19a, ZWH15b, ZWX18]. Schmidt [OZ18]. Schnirelman [Yao18]. Scholes [JHK17, KDB17, KV16, KV17, KV18b]. Schröder [PP18a, SH18, SH19]. Schrödinger [AL19b, AHKT15, CLC18a, COP17, CR17, HA18, HZ17a, LMW18, RZ19a, SVAKGR19b, SLW17, SW17d, TX19, VG17, WHZ16, WLW19, Zha17b]. Schultz [PKR18]. Schur [BO18, NZL17]. Schwarz [AL19b, LN17, LN19, VFGPG^{+}17, Xu18]. Schwarzschild [TW19]. SCI [GLLS16]. Science [DNY17, VACR15, WWST18, PT17b, VAMGR17]. Scientific [ZDFF18, ZDL16, ZDGM17]. score [LYG15]. scoring [WZBM18]. scour [EM15]. SCR [HJ17a]. scramblings [Vig17]. screening [PC18a]. SDA [DdSL16]. SDC [DdSL16]. SDDEs [LY15a, Tan18, ZWH15b]. SDE [AH15]. SDE’s [MP17b, ABT17, AK17a, CXW19, DIJ17, Gri19, HRO16, Prz15, Prz19, STS15]. Search [NRV18, SQRE16, AGM19, AR16, CD16b, EK19, Fan15, FP16a, GC18, HSS19, HWZ18b, ICMPG16, JJ19, SGK16, WZS^{+}18b, XqYL^{+}15, YBD16, YWY19]. search-based [AGM19]. searching [BHMT16]. seasonal [GVHMR18]. seasonality [JCC19]. Secant [AM15b, KGMS18]. Secant-type [KGMS18]. Second [ALC19, CPRS19, DCNY18, MCLW16, TYS17, Abd16, AH15, ALMLM17, AEHVM18, AFGT18, BH17a, BBC^{+}18, BMM17, BLM16, BW19, CRCM16, CHR^{+}15, CXZ19, CLV16, DJC15, DRST19, EC18, EHVM18, GUB^{+}17, GV17, HWCC15, HCL17, HLZ19, JB18, KP15, KD17, KMB^{+}18, Lee18, LX16a, LCH18, LHY19, LZY15, LTY17, LM17b, MHV17, Men17, Mic18, Mu19a, Mu19b, NHA17, NAH18, NS19, NSM16, NTPVNTDT15, OGR19, PPM16, PA19, RRM19, RMVA17, Reu16b, SvG15, SSW19, SVAKGR19a, UGG^{+}19, WCW15b, WWM17, YZH18, ZM15b, dR16]. second-derivative [Abd16]. Second-order [DCNY18, MCLW16, ALMLM17, BBC^{+}18, CHR^{+}15, CXZ19, HWCC15, HCL17, JB18, Lee18, LCH18, LTY17, NS19, NTPVNTDT15, RRM19, RMVA17, SVAKGR19a, WWM17]. section [KPS16, LPW^{+}19]. sections [HG17, Hu16]. sector [IAMCI18, Zag18]. security [Nke18a]. sediments [ZFK18].

Segmentation [FJ19, DF16a, HMMV17, HLMS19, JWQ18, ZWZ^{+}18]. segments [MGMGJLRB17]. segregated [GGRN17]. segregation [NGGA^{+}18]. Seidel [ZPL18, CDZ16, ZPL17]. SEIR [MVQDdR^{+}18]. Selected [VAMDGR19]. selection [IAMCI18, JS19a, LZ18b, OAN17, SP17, SJL16b, SV16a, WZBM18, WW19b, XDZZ19, ZB18, ZG17, ZX16]. self [AZ18, And17, BPP17, HYQ^{+}18, JZ18, Kad17, NC16, OD17, Wan18g, YN18, GLLS16]. self-accelerating [Wan18g]. self-adjoint [AZ18, JZ18]. self-balancing [NC16]. self-consistent [Kad17, OD17]. Self-Corrective [GLLS16]. self-diffusion [BPP17]. self-learning [HYQ^{+}18]. self-scaling [And17, YN18]. selfadjoint [WW16b]. SEM [PDO19]. Semi [HH18d, IÇ18, Vis19, AH16a, BCFN16, CKKJ15, CQW16, DI19, Fer15, FP16b, Guo17, HZ17a, JS17, KDW17b, KDW17a, KSW19, Lad15, LCC^{+}15, LWL17, LOAGBR15, LM18, PS15, Tho16, TN19, WLM16, XZ19, Yao16, YCW17, dR16]. semi-analytic [KSW19]. semi-analytical [KDW17b, KDW17a, LOAGBR15, WLM16]. semi-circular [CKKJ15]. semi-classical [dR16]. Semi-convergence [HH18d]. semi-definiteness [CQW16]. semi-differentiable [Yao16]. semi-discrete [DI19, HZ17a, LWL17, Tho16]. semi-embedded [PS15]. semi-explicit [JS17]. Semi-implicit [IÇ18, AH16a, XZ19]. semi-Lagrangian [FP16b, Guo17]. semi-linear [Lad15, TN19]. semi-local [Fer15]. semi-Markov [YCW17]. semi-NURBS [LCC^{+}15]. semi-smooth [BCFN16]. semi-symmetric [LM18]. Semidefinite [YB15, KYY15, NN18b, NN19, TBY17, ZC16a]. semidiscrete [LLY19, LMA18, PBP16]. semilinear [AFGT18, CZLV17, DVS15, DM16b, GS15d, MVM19, MTKO17, Ple17, RC19, Slo15, VDS15, ZZW16, vV16]. Semilocal [CHVRT15, KGMS18, EHV19]. seminorm [CSXY19]. Semiparametric [HYF18, JMSL19]. Semismooth [ZY18, HR16]. sense [Col16]. sensing [AALZ15, CG17a]. sensitive [IDGH18, NLG15]. Sensitivity [KMD^{+}16, BBB17b, SD15, SP17]. separable [CLZL18, CLZS19, LaT15, SY19, WS17]. separation [CL18b, GTC16, SA17, ZH18]. septic [LFCW19, ZWY16]. Sequence [AH18, DHAM17, HPT19]. sequences [AHBC15, ÁNPPCSA15, BE15, BDF15, SY18, dlCYC15]. sequential [HSS19, NPR18]. series [ABB17, AFGT18, BCVV18, DQIM18, DF16b, GLR16, MAK19, MJ18b, MGMGJLRB17, QLZ19, RN17, SBC16, WCS17, ZS19, ZY19a]. SERK2v3 [KMV16]. set [BDFM19, GdLT17, LWW17, LaTG17, LZL16, LCZ18, LZS15]. set-valued [BDFM19]. sets [BE15, CGPR16, GGBTP17, Iid15, Man15, MS19b, RRW17]. setting [AW19a, HRS16]. seven [HWZ19]. seventh [BMG^{+}19]. Several [LL18, GH16c, Gun19, JCH17, LSW19, Meh19, MT17c, NGGA^{+}18]. SGBEM [VMR17]. shale [TV16a]. shallow [ABT16, CL19b, JL19a, LSG18, LFKG16, Xin16]. shallow-water [ABT16]. Shape [ABRS17, Bal17, HFMM16, SSG16, Uda18, BDG15, CFMÖ19, LA15a, LZS15, Wan18c, ZWZ^{+}18]. shaped [DNK19, DB15, WHWY18]. shapes [MM18, WP16b]. share [ZJQW18]. Sharp [PGRC18b, TSMO17, Miy15, Zha19b]. sharp-interface [Zha19b]. sharpening [PBJCM19]. Shaw [ZBL^{+}16]. Sheffer [CGN18]. shell [BEMA16, JL19a]. shells [Kaz18]. Shepard [CDDD19, DD16a]. Sherman [WLZ17]. shift [CM18a, HS17c, HH18c, HLLY19, JL15, NML18, RCN17, SR16c, WL19a, ZL17b]. shift-invert [HLLY19, JL15]. shift-splitting [CM18a, HS17c, HH18c, RCN17, WL19a, ZL17b]. shifted [CCZ^{+}18, DSZ15, GHY^{+}18, Gun19, JBT16, KBL19, LH19, SHCJ19]. shifted-exponential [Gun19]. shifts [LC16]. Shiu [WXY17, ZS19]. shock [BKMO19, BKMO20, Col16, DGM16b, ET19, OE19, VG15]. shocks [BO16, MV16d]. shooting [AAA15, FGF19, HAH17, SM16]. short [BBB17c, REBF16, SOMF19]. shortest [HAH17]. shot [GGW16]. Shtrikman [CJP17]. SI [HLV16]. SIAC [LRKV16]. SIAM [LPW^{+}19, HLV16]. SIAM-2015 [HLV16]. side [EM15, HHOT17, KH18, MPH18]. sided [LFWC19, WXY17]. sides [ATG18, MLJ16, MT17c, REBF16, xZWlC15]. sign [APS15b, APS17b, APS17a, APS19a, APS19b, CSTU17, GS18a, GS19, STS15]. Signal [LC19, Skr17, LMS17, ZCC15]. signals [AS19a, BCNP15, Che16b, JW17, TYQ^{+}19]. Signature [Sch18, LSBL18a, LSBL18b]. significance [GSVSM16]. Signorini [GP16]. silencers [DSOBK16, SODBK19]. similarities [ALV19, HLV^{+}15]. Simple [KR18, MV16a, BGMM18, BDAG15, JLLZ16, LKN17, LX15b, MY18, Mud19, Xie19, LZ16a]. Simple-iteration [MV16a]. SIMPLE-like [LZ16a]. simple-root [LKN17]. simpler [xZWlC15]. simplest [MT16b]. simplexes [VS18]. simplices [KKV19]. simplicial [SM15b]. Simplified [LW19c, ML19b, Gia18, WWZ15, XZN^{+}19]. simply [MM15b]. simulate [Jeb19]. simulating [GTX^{+}19, KS16, KSW19]. Simulation [AdQdS16, SPG^{+}16b, AMRX17, AVGCMVM17, AG16b, CT16, CW16a, CL16, DBdLdM^{+}19, DXZ18, DIJ17, DFGS16, GCTGC16, Hei15, HL15, JL19b, Kaw17, KSD17, LHY16, LWV18, LYS19, LDZ18, QDG^{+}15, RNPK15, RQ16, dHKdHMM15, SRRJ16, TKH16, VMM18, WCW19, YBC^{+}16, ZWW^{+}18]. Simulations [LE18, BG15, COdSR16, CSS19, CY15, GTC16, GGK18, NZ16, PDO19, RNM16, RdSLL16, VG15, WLLC18, YZY17, ZBL^{+}16, dlHV16]. simulator [GGC^{+}18]. Simultaneous [Abd18, GGW16, MDH19, ZZW18, HWZ19, PBJCM19, PP18a, PP18b, PI19, RP16b]. simultaneously [CV17, Xie19]. Sinc [HR15, MMO15, AKHSU18, Moh17]. Sinc-collocation [HR15, MMO15]. sine [Bac18, EGK17, LDK17, Orr19]. sine-Gordon [Bac18, EGK17, YTSZ15]. sines [Tru19]. single [AG16b, CS18b, FOK15, GVMR16, GGW16, HH18b, MA19a, RP15, SJL16b, WXZ19]. single-factor [GVMR16]. single-index [SJL16b]. single-phase [AG16b, RP15]. single-step [GGW16, WXZ19]. singly [WKBB17]. Singular [Chr19, AT17, BDA17, BAO^{+}16, BS19, Bie17, CXZ19, CK16, Chr18, Col16, DA19, DLM16a, EJJ15, ERPD18, HH18d, KL15, KKK16, LC17a, LLSZ17, LZ17b, LM17a, MMO15, MN19a, Men17, MIT15a, Moh19, NS16, NSM16, NXLX18, PMN19, PPM16, Pis17, RW16, RT18a, RMK19, SK15, TF18, WR16, WCS18, WZ18, XYZ15, YZW15, YCX19, ZLS^{+}15, Zha18, ZCS^{+}19]. Singularities [Gor16, ACDT15, AT17, CL19a, Kan18, Wan18b, WCLL18]. singularity [HS18b, RR18, ZG19b]. singularly [AS18, AZ18, BR15a, BZ16a, CLX17, CW16b, CGSS17, CJ18, CJ19, Das15, DVA19, ESSE18, EC18, GO15, GO17, KAA16, KK15, Lin16, LC15b, LC17b, LLHO18, LM17b, MTV17, OP19, PCSR19, Qui15, RC19, SAA18, SR16c, YAA19, ZC18]. SIP [vSV15]. SIR [WSL17]. sites [LWXF18]. siting [FLMS16]. Sitter [TN19]. situ [AAB18a]. Sivashinsky [ID18, MFWK18]. six [MA16]. sixteenth [BAMM19, MBM17]. sixteenth-order [MBM17]. sixth [CT17b, GKM18, HMT15]. sixth-order [CT17b, GKM18]. size [ABL^{+}19, ALMLM16, CMRP17, MV16a, NZL17, Prz19, SV16a, Wan18h]. size-structured [ALMLM16]. sizes [FY18, NCF^{+}19]. sizing [SGK16]. Skeletal [CDE18]. Skew [HMKCR19, WSW15, CGMM18, LLW17, LS18b, NL17, SWL15]. Skew-Reflected-Gompertz [HMKCR19]. skew-symmetric [CGMM18, LS18b]. skewness [BOF^{+}16]. skinning [LCC^{+}15]. slabs [NGGA^{+}18]. sliding [DD16b]. slip [JLCZ17]. slit [DB15]. Slow [RKV19, YBC^{+}16]. slow-release [YBC^{+}16]. Slow-scale [RKV19]. Smale [GMW18]. Small [WP16a, DVA19, LS18a, NZL17, Sey18]. Small-sample [WP16a]. smaller [NS16]. Smallest [Chr19, Chr18]. SMART [CPRS19]. Smooth [TXFH18, YY16, BCFN16, GLR16, GKN16, HWT15, KV18a, LYG15, MTV17, MOY18, NR15, Zag18, Zak19]. Smooth-threshold [YY16, LYG15]. Smoothing [GKLS18, PTV17, AHM18, Fan15, PBMJC18, PBJCM19, WYW15, YS19]. smoothly [VRBS19]. Smoothness [DS15b, KP15, LRKV16]. Smoothness-Increasing [LRKV16]. SMRPI [FSNS16]. snapshots [KVO18, WMI16]. Sneyd [iW16]. Sobolev [BGMS16, BS17b, FMP^{+}15, GCZ17, KB16, LLG19, MMMMB16, Mir16, MB15b]. Sobolev-type [BS17b]. Soccer [PW15b]. social [Ogu17, SCMS18]. soft [JCC15, JMN^{+}16, ZM15a]. soil [dAFvdMR19]. solely [Slo15]. solid [SM15b, Zha19b]. solid-state [Zha19b]. solidification [ID18]. solids [ADG19]. Solitary [HA18, MM18]. solitons [CZ15]. soluble [BV16a]. solute [TNTL18]. Solution [Beh16, BGL17, BS19, MS19a, AVB17, AHT18, AKR18, AS18, Aks17, AT17, AK15, AFGT18, AD18, BM17a, BM18, BBD16, BJK15, BT17, BK15b, BFLP15, Baz15, BFMP19, BS15, Bot16, BGS19, CCNQ^{+}16, CL19a, CV19, CEJ16, CGvA^{+}15, CJV17, DP16, DDR19, DdSL16, DAM15a, DA19, Dia18, DMS19, EHTL19, FM19a, GMG19, GHB17, GAA16, GO15, GLSW16, HZ18a, HVV15, HK19, HR15, HSAR18, HKAR18, HH16, HMDSR19, IN16, JSL15, KVO18, KT15, LXW19, LY15b, LaTG17, LS18a, LZRH19, MDVM17, MHAA15, MMS19, MLJ16, MK17, Mer19, MKKV19, MH16a, MA18, MAS19, MV16c, Miy19, MTKO17, Mok15, Mud19, MTS17, NN18a, Nem15, NDH16, NXLX18, OP19, PD17, PPM16, PA19, Prz19, ROB17, RW16, SE16, SAC17, SLH18, SR16b, SDH17]. solution [SEÖ16, SS18c, SLA19, SVAKGR19a, SVAKGR19b, TJB17, TV16b, VF17, VM15, XHZ17, XF16, YCX19, yZLX15, dBGMUA18]. Solutions [NT15, SC18, AAA15, ÁD17, AS19b, AHL17, AAC17, BM17b, BBB17c, Bie17, Bog16, BCMP16, BHMT16, CLS18, CRGB17, CCNQ^{+}17, CRCM16, CLX17, CPB19, CXW19, CDRY18, CK16, Col16, CD15, DJC15, Das18, DD16b, DWXO15, Dzh18, FFP16, FLN16, GH19b, GLSW18, GR19, GL19, HP17, HA18, Hua19, KFT17, KDB17, Kha17, KLS19, Kre19, LDDF18, LY15a, La19, LRKV16, LLSZ17, LZY15, LSXL19, LNP17, MDG17a, MDGVR17, MDVR19, MVM19, MYM16, MM18, MT17b, MPH18, Mel19, MMZ19, Mil16, MSH18, Miy15, Miy17, MJS19, NSW15, PS15, PCJ18a, Pis15, RCBG19, Ros17, SR16a, SWL15, SLA19, SVAKGR19a, SZ18, SLW17, TMS^{+}17, TLFC18, TLE19, VDT19, Wan17b, WG18, Wu19, XYZ15, YHZ18, Yao18, YFS18, Zak19, ZSL15, ZLZ17, Zha18, ZJ19, ZG19b]. solvabilities [LaT15]. Solvability [AB18a, CS17, DHAM17, HND18, CLM19]. solvable [CNT17]. solve [ACDT19, DFT18, Dzh16, Haj19, HSAR18, HKAR18, KGES19, KDW17b, KDW17a, KMLB17, LLV18b, ME18, MII16, NAM18, PB15, PLP17, RM17, VFFGV16]. solver [AEKGP19, Che19b, FNV17, HM16, HMMV17, LLZ16, MT17a, Ngo19b, RTV15, VS15b]. solvers [BO18, BV18, lCCVV18, GGC15, KR18, SVW19, YPD16, ZS17b]. Solving [AHM18, BCVV18, CCJ16, CCJ18, CHS17, CJ18, CNQR^{+}18, FCJ16, GUB^{+}17, Goć15, HS17b, HIS16, JBT16, JD16, KV19a, KMV16, PRP19, PBS15, SY19, UGG^{+}19, ZCD17, AA18, ATG18, ACDT18, BJM16, BH17a, BMM17, BSGM19, BG17, BLM16, Bog15, CLZS19, CP18, CMQT16, CJSCT18, Cou19, DSA15, DAM15b, DK16, DSZ15, Dzh18, Elg19, EE16, FLZ19, FR18b, GYS17, GHZ^{+}15, GHC16, GHY^{+}18, HWH19a, HNT19, HWH19b, HWXC19a, HWXC19b, IIA15, JHJB17, JS17, KS17a, Kau16, KMB15, KKP^{+}19, LRT17, LDL^{+}19, LHY19, Lin16, LC18, LZ16c, LZZL17, Liu17, LHPZ18, LDW18, LF18, LM17b, LM18, MD19, MP15, MV16a, MH16b, MS17c, MAS19, MC18, MT17c, Mol15, NE17, NBAK19, NS19, NN18b, NN19, OAN17, PGD^{+}18, Pul17, RVA15, RMVA17, Reu16b, Reu17, RT18a, SAA18, Sar19, SB18, SJ18, SQRE16]. solving [SK15, SRH17, SVV18, TBR16, TZZ18, TP19, VARC17, WWZZ16, WZZ16, WR16, WKBB17, WA17, WLM16, XY16, YGS18, YHZ18, YB17, ZL18a, ZHWS16, ZWT19, ZL15b]. Some [AA16, BRS15, CHS15, CGN18, DZM16, FLJT18, FZW19, GAB17, KMB15, KCNH16, LXL16, LLV18a, LJX17, MDVR19, MBM17, MT15, MII16, QCSG18, SKC^{+}19, ÁD17, Baz15, BK19d, Ceb17, EG17, FH16, FH18, GUB15, KB16, LQZ16, Li18, LSS19, LPT17, LL17, MDG17b, MVM19, NH19, NN19, RABB19, RFO17, SY18, SVAKGR19b, SK16, SM15b, TAT18, TPC17, Yos17, ZDF^{+}17, dR16]. SOR [HWXC19b, GxLlW15, HWXC19a, MSZ18]. SOR-like [GxLlW15]. SOR-type [MSZ18]. sound [AO18, GH19a, SODBK19, ZM15a]. sound-hard [GH19a]. source [AS19b, AR19, CCW17, CS18b, GS15d, JAN17b, LWC17, Liu17, PP16, PCSR19, SS19b, Slo15, THL^{+}17]. sources [BHU18, LXZ17, WSG^{+}19]. SP [VFGPG^{+}18]. Space [AM15b, SGL17, AAA15, ADG19, ALC19, ADD^{+}16, AK19, AP18, BJM16, Bac18, BBCK19, BAAA^{+}15, BVZ17, BPD15a, BCMP16, BMLR15, Bou18, CL18a, CJLT15, CLA19, CRG19, CDKN14, CDKN15, DA19, Din16, DLM16b, DK16, FH16, GZD^{+}19, GG15, HA18, JW17, KV19b, EP18, Lee18, LHJ16, LLY19, LFWC19, LSP18, MPS19, MV16a, MGMDGVR17, MKKV19, RWW17, RZ19a, SAA18, Sai19, SWZW15, TYS17, THL^{+}17, UPD^{+}15, WGHR17, XDY16, Yao16, ZYL19, ZZG^{+}19]. space- [BBCK19]. space-dependent [CL18a, LHJ16]. space-filling [MPS19]. space-fractional [AK19, Bou18, CJLT15, KV19b, LLY19, LFWC19, SWZW15, TYS17]. Space-time [SGL17, ADG19, BVZ17, FH16, GZD^{+}19, GG15, LSP18, ZYL19]. Spaces [BR16, AA16, ABH19, AMB19, ABEHV16, BH17a, BGMS16, BKLR15, BHRR18, BLMS19, CDF^{+}16, lCCVV18, CHVRT15, CDHP17, DHAM17, DKLP15, FZW19, Fer15, God15, HKAR18, HSS19, KV18a, KGMS18, KS17c, MOY18, MP18, MP17a, dHKdHMM15, RMMRC15, Sak17, WWH19, XHZ17, ZDLD15, ZWL18]. spacetime [TN19]. SPAI [JK19]. Spain [NGGFM17]. spamming [SvG15]. span [AALMLM18]. spanning [HLLS18]. Sparse [XH15, YWW19, Bai15, BZ16b, JK19, KC19, PW17, SRBD18, WP16b, YP17, ZS17b]. sparse-grid [SRBD18]. Spatial [BDF15, CZ15, GFPGF19, MB15a, PP16, SM19a]. spatial-stepping [GFPGF19]. spatially [Abd18, Wol17]. spatio [Abd18]. spatio-temporal [Abd18]. Spatiotemporal [NY19]. Special [DLY18, VAMDGR19, BCFN16, HHKR15, HIS16, MT18, NN18a, UPD^{+}15]. specializations [SS16b]. species [CSE^{+}19, SR16b, SRA^{+}15, Wan18h]. specific [HWH19b]. Specifying [KElY16]. Spectral [BFMP19, CLC16, HCC18, LWL17, LV16b, RCN17, Vis15, WNML16, YTSZ15, vD15, AK19, AdGvG19, CW16b, Che19b, CY15, CN16, DNL15, DNL16, EE16, FSNS16, HNT19, HW17, JWQ18, KT15, Lee18, LL16b, LMW18, LLHO18, MN17, MN19a, MS17e, Moh19, NH19, SS16a, SR16b, SB18, SWL15, SNY19, SJ17, SJ18, Su16, TJB17, WW16b, WCS18, XDY16, YHZ18, YJBM19, Zak19, yZ17, ZHC19, ZYL19, dFG17]. Spectrally [DYD19, CCY18, YC19]. spectroscopy [HHS^{+}15]. spectrum [HLLS18, PGRC18a]. speed [ID16]. SPH [VRBS19]. Sphere [Cro15, AS19b, DLW16a, MF16, PC18b]. spherical [BL18, BDF15, Cro15, HL18b, Kaz18, VX19, WLZS18]. Spiral [DWS^{+}18, CC15]. spirals [AV16, LAHB15]. Spirulina [NGGFM16]. Spline [DDR19, HHZ18, AZ18, BGII15a, BGII15b, BI17, BIJM^{+}19, BLMS19, CFSS18a, Che18, DIS^{+}18, ELL^{+}19, FWW^{+}18, GK18b, Guo18, KV18a, KPT15, Kre19, KB16, LAHB15, LCC^{+}15, LLG^{+}18, LFWC19, LM17b, LZ19, Maz18, PTV17, SST15, SYT16, TXFH18, XMW^{+}15, YW15a, ZDLD15, ZTGZ18, ZWL18, ZD19]. spline-based [BGII15b]. Splines [HH15, AHBC15, BL18, BIR17a, BDIR19b, BC16, BAHC17, Bej16, BR16, BCDwK17, BDGO15, DS15b, FJ19, GS18a, GH17a, GKPS19, GK18b, Han15, KPRA19, RdV19, SuSR16, SM15a, US19, QDY^{+}19]. split [AH15, BK19a, DMMSK17, DTKP17, GK18b, IÇ18, LSL17, LSL18, RKV19, SST15]. split-drift [AH15]. split-step [IÇ18, LSL17, LSL18, RKV19]. splits [GK19a]. Splitting [MV18, AMRX17, AMP18, AMCR19, AHKT15, BH17a, Bad18, BPP17, COP17, CLW17, CM18a, CJ18, DBdLdM^{+}19, DGH15, EAKS18, FFKP17, GHM17, GXZ15, Güc17, HOS16, HS17c, HH18c, HCC18, KV15, Lee18, LY18, LLV18b, LK18, MP17a, MV16b, OOO15, RCN17, SEÖ16, SZ16, SX18, WNML16, WL19a, WHS17, WL16, WL17b, ZZLY16, ZS18b, ZZZ15, ZM16b, ZL17b, ZL17c]. splittings [AK17c]. spray [SRRJ16]. Spurious [DJC15]. SQP [HP15]. square [ABT17, AH16a, BCVV18, CCJ18, CJV17, DB15, GQM16, JMN^{+}16, LZ16b, LDZ18, LLG^{+}18, TW17, WA17, Zho16, ZWH15a]. squared [SJL16b]. squares [AHMY19, AS18, CQW16, CQ17, CSXY19, DF15, DDM15, DM17a, Dia18, DLM16b, HLV^{+}15, HMT^{+}17, HHS^{+}15, JL16a, JL17, Ku16, LC15a, LHK19, Lin16, LWJ15, Mir15, MJS19, Nie15, PLP17, SY19, WS18, XCG19, YW16, Yua15, ZPL17, ZPL18, dBGMUA18]. Srivastava [DP19b]. SSP [Son16]. stabilities [JS19b]. Stability [CCTV17, CMQT16, CMTV17, DLW16a, KJ19, LD19a, LD19b, LSZ16, LWTQ17, Mar19b, MP18, Ngo19b, OM17a, Par15, SBU^{+}17, SFH^{+}16, Sha15, US19, VDT19, WAGVRB19, vtWvdHvdVV16, YYL15, You19, Zha16, ZQJY18, ZWWW19, ZDF^{+}17, AH16a, AYSE19, APS17b, AJ17, BYYV16, CW15a, CEJ18, CWH18, DIJ16, FE16, GM16b, GQM16, Guo17, GL19, HPW16, HLM18, IJ18, JKL^{+}15, KRR18, LY15a, Lan18, LXW19, LJ15, LL15, LZ16b, LWS16, LDZ18, LPT17, LSL17, LSL18, MYM16, MM15a, Mil15, MFWK18, Ngo19a, TLLL19, Xia17, YL19a, ZSL15, ZLF18, ZWH15a, ZWX18]. Stabilization [Arm18, PBGG17, JKL^{+}15, LSD15]. stabilize [GPHAPR17]. Stabilized [Pol16, XZ19, BS16, BBC17, DD18b, JLCZ17, LLCG16, LFKG16, MVK19, QM15, TV18, SAT15]. stabilizing [HP17, Miy17]. Stable [Che16a, CL15, EBJS19, KYS16, ZM16a, Abd16, ABT17, ARTY19, BK15b, BK16, CJLT15, CZ17b, CFWW19, CSTU17, CGT18, DMP19, DXZ18, HZ17a, KSR15, KS16, LLZ16, LHY19, NC19, NCPD19, PCJ18a, RdSMO17, RdSMO18, SEV^{+}18, SW17a, SVAKGR19a, SVAKGR19b, WD19, YZH18, YBD16, ZL18a, ZL18b, ZWL18]. Stackelberg [AVGCMVM15]. stages [LSW19]. Staggered [CY15, GDCK19, KCX17, ZP19]. Staggered-grid [CY15]. standard [CP16, GPHAPR17, KW16, LA15b, MVQDdR^{+}18]. standby [FOK15]. star [HPT19, WHWY18]. star-shaped [WHWY18]. STARR [GSM19]. Starting [EHVM18, AEHVM18, ZS17a]. State [LSS19, AG18, BGNGR^{+}17, BCMP16, BySZ19, DLM16b, GC15, EP18, KLS19, KS16, LZA19, LSBL18a, QDBG18, RNM16, SYD15, UPD^{+}15, YT16, Zha19b]. state-constrained [BySZ19]. state-dependent [LZA19, YT16]. state-space [BCMP16, UPD^{+}15]. States [LPW^{+}19, SFH^{+}16, Sha15]. static [AG18, CMHHS17, CDD^{+}15, LQ19]. stationary [ABRS17, AS19a, BYYV16, CJCDLL19, CCIL18, JY19, LGYC19, LM15, SuSR16, uIZ15, TZZ18, WWZZ16, ZL19]. statistic [Mar19a]. Statistical [HG19, PD19, VF17, ZSW16, ZG19a, LO19, RDRM15, WP16a, YW16, ZNS18, ZX16]. statistics [BE15, CJP17, DD18a, DG16b, JGAFAL19, LWZ19a]. status [KElY16]. steady [BGNGR^{+}17, BKN16, CDW15, CHS15, CHS17, GC15, LW15, LW17b, LZ16a, QDBG18, RNM16, SFH^{+}16]. steady-state [GC15, RNM16]. steel [CS17, NGGA^{+}18]. steepest [EJJ16, GS18b]. steered [ZF15]. Stefan [Mit15b, MV16c, RVC19, VM15]. Steffensen [ABEHV16, HVM18, NBAK19]. Stein [LH19]. Stein-conjugate [LH19]. Steklov [AC17b, DIJ16, DIJ17, HM16]. Step [ASKG18, YMGA19, ABHVM16, AAB^{+}18b, AZ17, AK16, AHK19, BKK18, DDG^{+}16, FYY17, FR18b, GXZ15, GGW16, IÇ18, KMRS16, KW18, LWDW18, LSZB17, LSL17, LSL18, MVQDdR^{+}18, MV16a, NS18, PS15, Prz19, RMVA17, RKV19, SW17a, SVAKGR19a, SVAKGR19b, SV16a, SV16b, WXZ19, WKBB17, WHS17, WZ16, ZJG^{+}15, ZSW16, ZHH^{+}19]. step-size [Prz19, SV16a]. step-stress [LSZB17, ZSW16]. stepping [CM16, Con18, GFPGF19, GM17b, vV16]. steps [ABEHV16, AR15a, AR16, HWXC19a, HWXC19b]. stepsize [KW18]. stepsizes [LL18]. Stewartson [Gru16, Mus17]. stiction [HWP^{+}15]. Stiefel [DLHY15, KMLB17]. Stieltjes [MVPN15, QG17]. stiff [AFH16, Abd16, AHP19, AH15, HHR16, HOS16, KMV16, KW18, MG15a, MD19, Ran15, RKV19, SW17b, WKBB17]. stiffness [CKL18]. Stirling [AMB19]. Stirling-like [AMB19]. Stochastic [AYSE19, GD16, ML16, SW17c, ZZHS17, dMPF15, AH16a, Akh19, AH15, ABP15, AK17a, BJ16, BYW18, BZ19a, BS17b, BHRR18, BH17b, CPB19, CW15a, CZLV17, CPZ19, CPOG18, DBX19, DFLM19, DIJ16, DV16, DI15, DM17b, DP19c, EPC17, FL16, GH16a, GVMR19, GQM16, GWY17, GLMY18, GL19, GB18, HHR16, HYF18, HMD19, HWZ18a, HWT15, HZ18b, HT18, HLM18, Hui16, IÇ18, JC19, JY17, JC15, JMS17, KB17, KRR18, KHLY17, KEB19, KK17, Lan18, LX18, LXW19, LW19a, LWXF18, LHM16, LYG17, LM15, LZ16b, LDW17, LM17a, LDZ18, LZX19, LSL17, LSL18, MTCC15, MDVM17, Mao15, MYM16, Mao16, MS15a, Mil15, Mil16, MT16b, MH16b, MSH18, MAS19, NH19, OM17a, OLPT16, PX17, RP17, RKV19, RZ19b, SLH18, SCMS18, ST16b, SSCM18, SCG18, Sun16, TJB17, WS19, XY15, Yam17, YZJC15, YL19a]. stochastic [YZZL19, ZF15, ZqZwZ15, ZS17a, ZSL18, ZJ17b, ZJ19, ZL15c, ZEBM17, ZWH15a, ZWX18, dFG19]. stochastically [Sha15]. stock [ID16, RdSMO17, RdSMO18, ZS17c]. Stokes [ACLY19, AB18a, AFN17, BBGG19, BAK18, CJCDLL19, CDW15, CGPR17, CHS15, CZ17b, CHS17, Chi17, CK16, CVW17, Cou19, DD18b, ERR18, GM17b, GGC15, GGRN17, HK19, HS18c, HM19a, HMY19, JM15, JLCZ17, KYS16, KLC17, KtTBK18, KK19c, LC15a, LZC15, LA15c, LLCG16, LGLC18, LGYC19, LSYY19, LZ16a, LHM16, LLC18, LW19c, MT16b, MV16b, MV18, MWY15b, Ngo19a, Ngo19b, QM15, RR18, SSG16, Sun19, TZZ18, WWZZ16, ZZZ18, ZL18b, ZL19, ZH15]. Stokes-like [KLC17]. Stokes/Darcy [Chi17]. Stokes/elliptic [Sun19]. stopping [MDVM17]. strain [FMMQ19, FCM19, Lu15c]. strain-limiting [FCM19]. Strang [CR17, GHZ^{+}15, Güc17]. Strang-type [GHZ^{+}15]. strategies [APS19b, AVGCMVM15, CDD^{+}15, CFSS18b, CW15b, DKV16, DSN17, NC16, PZ18, QDG^{+}15, SSKV17, WLLC18, WRZ18]. strategy [ASV17, CD16b, CSB18, DYD19, LRZ15, Mud19, NML18, SWZW15, SV16a, SL18, WP17, YL19b, ZC19]. stratified [LXZ17, LXZ18]. Stratonovich [AK17a, MH16a, MS17c]. stream [MV16b]. streaming [MAK19]. strength [BSLL18, BSLL19, ET15, GUB15, LSBL18a, LSBL18b]. Stress [ET15, LSBL18a, LSBL18b, BSLL18, BSLL19, DB15, LSZB17, MV18, Wan18c, Wan18d, YZPL18, ZSW16]. Stress-strength [LSBL18a, LSBL18b, BSLL18, BSLL19]. stressed [JL19a]. stresses [QWZ16]. stretched [AHBC15]. strict [DTKP17]. strictly [APS15a, APS15b, APS17b, APS17a, APS19a, APS19b, MM19]. strike [JHK17]. strip [BBPR15]. stroke [CNQRR17]. Strong [DIJ17, HK07, HPW16, IJ18, LX18, LW19a, LYG17, TGJ19, ZSL18, ZJ17b, BYYV16, HHR16, HM19b, JJ19, MVPN15, ZZHS17]. strongly [BK19b, BK19d, CW17, KRR18, LC17b, RMK19, TH18]. Structural [GR16, Lu18, MX16, ZN19, JA17, Yua15]. Structure [BMLR15, GLZ18, ÁAM15, BBB17c, BWS17, DY15, HOS16, JWZC18, LY15b, LWZZ16, LZA19, MDAP19, MTMZ19, SZM^{+}18, SH15, Yos17]. structure-aware [SZM^{+}18]. Structure-preserving [GLZ18, BWS17, JWZC18, LWZZ16, MDAP19, Yos17]. Structured [DWQ16, Li18, ALMLM16, BLXD18, BDG15, DZ18, FV17, Hua19, JZ18, NR19, SS18d, ZYL19, yZyMxG^{+}16]. structures [ALS19, BN19, CCZZ16, GZH19, TLL16, XCG19]. Struve [KK19a]. student [GR16]. studies [ABL^{+}19, Nke18b]. Study [ALMLM16, BCT16, VX19, AM16, AAB^{+}18b, AFdOdS15, BDA17, BOF^{+}16, BKN16, BHMT16, CCNQ^{+}16, CFSS18b, CN18, Col16, CMQT16, FNdO15, GGG19, GKM18, GJ17, JS19b, KLM17, KAM16b, KV17, KLS19, KPS17, LLSZ17, LMS17, LLSM15, Mus17, OLRPdCJ19, OGR17a, LGB18, SJRR17, SD15]. Studying [NC16, ARCA^{+}19]. Sturm [CKB17, MR19]. sub [AB18a, BMLR15, CCW17, HY19a]. sub-Cauchy [AB18a]. sub-diffusion [CCW17]. sub-networks [BMLR15]. subdifferential [HND18]. subdiffusion [DAM15a, SJ18]. subdividing [BMLR15]. subdivision [ABRS17, ARTY19, CM19b, CH19c, CDHP17, CPRS19, DM16a, DXML19, DY19, DS17b, GZDC19, GMdV19, JY19, KBL19, MS17b, MB15a, NR15, RS19, SuSR16, ZCS^{+}19, ZMTT19]. subdomains [Vab18]. subject [HZ17c]. subjected [BO16]. Submatrix [XCG19]. submerged [FdGKZ17]. Subordination [BB18]. subproblem [JGTB15, QC18]. subsampled [Mar18]. subsonic [Su18]. subspace [ABG17, Bai15, Bot16, GG17a, HSS19, KC19, KBG17, Mia18, vSV15]. subspace-based [ABG17]. subspaces [FwC17]. substantial [HCL17]. substitute [PW15b]. substitution [Reu16a]. substochastic [BB16]. subsurface [CSS19, JO17]. successful [NGGFM16]. successive [AKR18, EK19, FGF19, QWS15, US19]. successive-projection [US19]. sufficient [DLHY15]. Sugiura [PI19]. suitable [WZS18a]. sum [CQW16, NL17, NC19]. sum-of-squares [CQW16]. Summation [ZF15, LWS16, RT18b]. sums [DSA16, JN19, OLPT16]. Sumudu [BS17a]. Sundman [LOAGBR15]. sup [LT18]. super [AW19b, CDD^{+}15]. super-convergent [AW19b]. super-replicating [CDD^{+}15]. superconductors [VS15a]. Superconvergence [Bac18, BW19, HC15, JM15, MN17, MN19a, MHL16, NSYX17, WL19b, XH19, YP18, LYG17, ZC18]. Superconvergent [SST15, AST19, DT17]. Superlinear [AKM18, ZC16a]. superlinearly [JGTB15, ZJ19]. SUPG [Jeo15]. supply [IABIPG17, IAMCI18]. support [MLS15, RRW17, WS18]. supported [HW18a, Man15, MM15b]. supraconvergence [KD17]. sure [CW15a, LWZ19a, Mil15, Tan18]. Surface [LCC^{+}15, YLZG18, AN17a, AN19, AS19b, BOF^{+}16, BKN16, Cou19, LXD^{+}18, LLG^{+}18, Rät15, RTV15, YW15a, YW15b, ZM15a]. surfaces [AH16b, DD16b, HH15, HJS19, HLMS19, JSC18, KP19, KPRA19, LZ18a, Rót15, SAT15, VL15, WZZ18, YZ15a]. surfactants [BV16a]. surrogates [NRV18]. Survey [JSC18, ABN^{+}18, BCCB15]. survival [AALMLM18, ABP15, GH17b, GUB15, LSBL18a, LSBL18b]. SV [dFG19]. SVM [NGGFM16, NGGFM17, OLRPdCJ19]. SVM-based [NGGFM16, NGGFM17]. SVR [XF16]. swaps [CP15, DT17, ZL15c]. swarm [XqYL^{+}15, XY18a, XY18b]. sweep [WL16]. Swift [SEV^{+}18]. switched [GGBTP17, PBGG17]. switching [CXW19, DF16b, DT17, JQWW16, KKW16, KV17, LY15a, LZ17a, MZ16, OM17a, RP16a, WW19b, YYL15, ZS17c]. Sylvester [CL18b, OSMZ18, PS18]. symbol [GMdV19]. Symbolic [SS16b, JSL15, MR19, PN16]. Symmetric [BW16, LWDW18, NAH18, SZ18, Ais17a, Bej16, BBG^{+}19a, CGMM18, CM18a, HMKCR19, HY19a, HJ17b, HQ18, HWXC19a, HWXC19b, LLL19b, LS18b, LF18, LM18, PS18, Pap16b, RWW17, WZL17, WCW18, WL17b, XCG19, YLL15, dBAAA15]. symmetrical [MII16]. symmetrically [AHBC15]. Symmetries [ALV19, CRGB17, CRG19, GKBG17, HJS19, LSXL19, RCBG17, RCBG19, RM19, SR15]. symmetrization [MP18]. Symmetrized [AHK19]. symmetry [BS18, MII16, VX19, YX19, ZYF19]. Symplectic [BBPS17, BBC^{+}18, BBCK19, LSW19, Moo17, NHA17, SW17c, ZZHS17]. Synchronization [SZZ^{+}16, AY19, APYW19, LWTQ17, TJSW19]. Synge [KKV19]. synthesis [HHKR15]. synthetic [OG16]. system [Abd18, AJ18, AKM18, BCFN16, BO16, CJCDLL19, CQ17, CLW17, CSE^{+}19, DHAM17, DVA19, DMMSK17, DG16a, DAAF16, ET19, Ery19, EHTL19, GUB15, HK19, HHLS18, HMMV17, HSAR18, HS17b, JGTB15, Kam17, Kav16, KBT15, Ku16, KTT16, LP15a, LaTG17, La19, LDZ17, LC15b, LC17b, LCX18, LSBL18a, LSBL18b, MT16a, MM17, MS17c, MAS19, MV16d, Mus17, Pis17, RIK16, RC19, Ros17, RFLM16, SFH^{+}16, Sha15, SS18c, Su18, Sun16, SW17d, VG17, WZBM18, bWzZjH16, WL17b, XGH18, XY19, Yan17, ZZG^{+}19, ZS17b, ZZ18]. Systems [DLY18, AH16a, ABB17, ABG17, AL19a, ADMV19, AGR15a, ÁD17, AR16, ATG18, AH15, ACDT18, ACDT19, AAKW16, Bai15, BBB17c, BPP17, BJK19, BL17, BGL17, BK15c, BK16, Bie17, BZ16b, CCIL18, CGMM18, CZ15, CDH^{+}17, CM18a, CJ18, CJ19, CWH18, CJSCT18, DP16, DMM18, DSN17, DLM16b, DSZ15, DFS15, ERPD18, FE16, FM18b, FPGK19, FOK15, GDM16, GLR16, GGBTP17, GGG19, GH19b, GRS19, GLV15, Goć15, GM17a, GLSW18, GW19, GHC16, GHY^{+}18, GxLlW15, HHR16, HMDSR19, HZ19, Hua19, HWXC19a, HWXC19b, HMT15, HKZZ15, IABIPG17, IIA15, JSL15, JK19, Kad17, KKS16, KDW17a, KVO18, KR18, LDDF18, LD19a, LD19b, LLT16, LXD^{+}18, LHZ15, Liu16, LSZB17, LLV18b, LK18, MDT16, MO18a, MO18b, MP18, MP17a, MS15a, MLJ16, Miy15, MSZ18, MT17c, NBAK19, Nat19, NZL17, OGR17b, PDB15]. systems [PBGG17, Pul17, RM17, RWW17, REBF16, Rät15, RKV19, RFO17, RFLM16, SS16c, SW17b, SM16, SV16b, SW17c, TV18, WZL17, WLZ17, WW19a, WYLD19, WS19, WXZ19, WA17, Xu15, Yam19, ZZ19, yZLX15, ZZZ15, ZWWW19, xZWlC15, ZZHS17, ZJ19, ZZ15, ZJK16, dBGMUA18, dFLM18, vSV15]. Szász [CGN18]. Szego [CBMPP15, JRT17].

T [OSMZ18, ZWWW19]. T-congruence [OSMZ18]. T-S [ZWWW19]. Tabu [SQRE16]. tailed [Che19a, FF15]. Takagi [WCW18]. taking [SS18b]. Talenti [NW19]. Tamed [JY17]. tangent [Han19]. tangential [BJK19, DD16c, QC18]. tapered [ZM15a]. target [KK17]. tau [RKV19, GMG19, GAA16, MC18, Pis17]. Tau-collocation [GAA16]. tau-leap [RKV19]. Taylor [AH16a, CZ19a, HMD19, SIAJ^{+}19, ZJG^{+}15]. Taylor-type [ZJG^{+}15]. TB [MLV16]. Tchebycheffian [BLMS19]. technical [CNQRR17]. technique [ABT17, ACAC17, CVW17, DAM15b, DA19, Els15, GTC16, GMCD15, GVHMR17, GdLT17, HHEG17, HWZ18b, ID18, IGRL17, Jeb19, LC18, RL18, SLH18, SBP17, US19, Wan18f]. techniques [AA18, CCNQ^{+}16, FLJT18, FFKP17, GG17a, GGC15, Hac18, HWH19b, NGGA^{+}18, NGGL^{+}19]. Technology [DNY17]. telecommunications [ASV17]. telegraph [CJLT15, HWH19a]. teleoperation [ERPD18]. temperature [CL18a, DSOBK16, OZ17]. temperature-dependent [OZ17]. tempered [AM19]. Template [FJ19, BRW16]. Temporal [AEK19, DLM17, MVM19, Abd18, FR18a, OGR17a, LGB18]. temporally [KV19b]. tension [AN17a, AN19, AZ18, Cou19]. tensor [AL18, BW16, CS18a, CHYZ19, CSXY19, DFS15, GH19b, HS18a, HFMM16, LZL16, LSH18a, LLV18a, LDG19, LLL19b, LLV18b, LM18, SNY19, XH15, ZW16]. tensors [CQW17, CQW16, CS19, HQ18, HM19b, KLL18, TPC17, WCW18]. term [AÖ17, BGNGR^{+}17, BB18, CCW17, CLC18b, CLA19, CS18b, CRG19, DSA15, FLJT18, HHOT17, HLZ19, KV17, LNP17, MRS17, NV19, PCSR19, QLT^{+}17, SLA19, WCW15b, WXJ19, YN18, ZH18, iW16]. terminal [FMR15, Zak19]. termination [WZS19, WS16]. terms [APYW19, IÇ18, JAN17b, YFwC19, YZ15c]. ternary [SuSR16, ZMTT19]. Territorial [HEA18]. test [FPGK19, Mar19a, OV17, SMTdF18, Wan18c, ZP16, BCT16]. testing [Che19a, JGAFAL19, Mar19a, Rom18, SM19b, WL17c, YZPL18]. tests [AFBJGJ17, CDN19, GSVSM16, LO19, LSZB17, ZSW16]. tetrahedral [BIR17a]. tetrahedron [MB15a]. tetrahedron-based [MB15a]. textiles [GDC15b, GMCD15]. texture [BOF^{+}16]. TFPL [SVAKGR19a]. th [BBCP16, Miy18, RM19, Zha18]. th-order [RM19]. THB [FJ19]. THB-splines [FJ19]. THBox [GKPS19]. THBox-splines [GKPS19]. their [AGR15a, DWQ16, Dzh18, GKN18, GT17, Hai17, HH18a, IAAW18, KSR15, KCNH16, LKN17, MPH18, MA19b, QL18, ST16a, SvG15, VL15, WSW15, YZZL19, YMGA19]. them [MII16]. theorem [CT16, CDH^{+}17, Ou18]. theorems [GL19, dHKdHMM15, RMMRC15]. Theoretical [Nke18b, ZP16]. theories [LOAGBR15, OGR19]. Theory [BCJV18, GAA19, GAA21, ACCF19, AH16c, BL17, FMMQ19, GX17, JSC18, MLV16, Mär17, OH18, OGR17b, Ros19, SV16a, Tho16, YX19]. thermal [KV19b, QDBG18]. thermally [JL19a]. thermo [DCNY18]. thermo-mechanical [DCNY18]. thermocapillary [JL19b]. thermodynamic [RANG15]. thermodynamically [CSE^{+}19]. thermodynamically-consistent [CSE^{+}19]. thermoelastic [Sai16]. thermoelasticity [FMMQ19]. thermoviscoelastic [ACCF19]. Theta [ZWH15b, LZ16b, LDZ18, LSL17, LSL18, ZWH15a, ZWX18]. theta-Milstein [ZWX18]. thin [AN17a, AN19, Bej16, BBPR15, ZFK18]. Third [ACMT17a, WZS18a, AZ18, ATV15, CW16b, DT18, GGBTP17, WKBB17, ZHC19]. third- [WKBB17]. Third-degree [ACMT17a]. third-order [AZ18, CW16b, GGBTP17]. Thomas [JSL15, LZ15, PD17]. Three [AR16, ADD^{+}16, AHS19, BI17, BDIR19a, CJP17, CTZK18, GH19a, GRVC15, GC17, GDQ17, HZ18a, HM16, HMH18, KC17, KKK16, KPS17, LLL19a, MGMDGVR17, SAC17, Sai16, TKMT17, WD18, YN18, YZ15c, YH18b, Zha15b, ZJ17a]. three-dimensional [CTZK18, GRVC15, LLL19a, SAC17, ZJ17a]. three-level [WD18]. three-party [MGMDGVR17]. three-phase-lag [Sai16]. three-point [AHS19, KKK16]. Three-steps [AR16]. three-term [YN18]. three-terms [YZ15c]. threshold [BJK15, JSW18a, LYG15, YY16]. thresholding [CPL^{+}19a, YSL17, ZPL17, ZPL18]. tight [CG17a, Vis19]. tightly [Kad17]. Tikhonov [BBH^{+}17, CP18, DWQ16, GS18b, MRS17, PRRY18, PB15, WWZ15, YW15c]. Time [AMP18, AAKW16, LRZ15, MGMGJLRB17, MFWK18, ADG19, ALC19, AG18, AY19, APYW19, AYSE19, ÁRS16, AK19, AW19a, AS19a, AL19b, AR19, AHK19, AFN17, AL19c, BBCK19, BO18, BVZ17, BK15a, BO16, BB18, BKK18, BK15c, BM19a, BEHL17, Bot16, BHU18, BK19c, BLH^{+}17, CL18a, CGPR17, CM16, CLC16, CLC18b, CLA19, CW15b, CS18b, wCW15a, CJ19, Con18, CLPZ17, DEY19, DSA15, DA19, DXZ18, DLM17, DF16b, Din16, DV16, EAKS18, ERPD18, EG17, EPC17, FLJT18, FH16, FE16, FPGK19, FP16b, FR18a, FY18, GHM18, GM16a, GZD^{+}19, GPHAPR15, GPHAPR17, GO15, GGK18, GG15, GM17b, Guo17, HCL17, HP17, HA18, HKZZ15, JHJB17, JAN17b, JLCZ17, KC17, Kav16, KMD^{+}16, KKW16, EP18, KBG17, KVO18, KS17b, KK17, KTT16, LXW19, LLT16, LJ15, LHLL18, LMUZ18, LC18, LWS16, Liu16]. time [LZY17, LF19, LSP18, ML17b, MAK19, MDG17b, MDVM17, MJ18b, MO18a, MO18b, MP17a, Miy17, Miy19, MTKO17, Ngo19a, Ogu17, OCU17, OCU18, Prz15, QLZ19, QLT^{+}17, RZ19a, SB18, SGL17, SLA19, SZZ^{+}16, SJ17, Sho18, Slo15, SAT15, SCP19, SYD15, SYD17, SZJ19, SYW19, Tra15, TGSS17, TNTL18, US19, VFFGV16, VS15b, VM15, WCS17, WWZ15, Wan18a, WS19, vtWvdHvdVV16, WS16, YZJC15, YZY17, YHZ18, YH18b, ZYXT16, ZXQH17, ZQJY18, ZH18, ZHH^{+}19, ZYL19, Zha17b, ZZG^{+}19, ZWWW19, ZL17c, dlHV16, vV16]. time-average [CL18a]. time-censored [FPGK19]. Time-consistent [LRZ15, CW15b]. time-correlated [Liu16, WS19]. time-delayed [ML17b]. time-dependent [AL19b, AR19, BBCK19, BM19a, CGPR17, FY18, GHM18, Guo17, JAN17b, JLCZ17, LXW19, MDG17b, SJ17, Slo15, SYW19, VFFGV16, YZJC15, ZQJY18, ZL17c, dlHV16]. time-domain [LHLL18, VS15b]. time-fractional [AR19, BB18, LF19, QLT^{+}17, SB18, SYW19, WWZ15]. time-harmonic [AFN17, GM16a, Wan18a, YH18b]. time-irregular [Prz15]. time-optimal [MO18b]. time-parallel [KBG17]. time-periodic [AL19c]. time-reversible [AHK19]. time-series [MAK19]. time-space [CLA19, Din16, HA18, RZ19a, ZZG^{+}19]. time-splitting [EAKS18]. time-stepping [CM16, Con18]. time-varying [AG18, AY19, APYW19, AS19a, EPC17, FE16, HKZZ15, LJ15, LMUZ18, SZZ^{+}16, ZHH^{+}19, ZWWW19]. times [GMSR18, GMSR21, GDC15b, GC17, LPT17, MOY18, PD19, STZ16, SDH17, VL19, WZ16]. timescales [Nat19]. timestepping [KRR18]. timing [LQZ16]. Timoshenko [RL18]. Titchmarsh [Yak19]. TKEMPR [OD17]. Tocher [BI17, KB19]. Toeplitz [Elo19, Hon19, NZL17, WLZ15, WLZ17, ZZG^{+}19, ZZ15]. Toeplitz-plus-diagonal [WLZ17]. together [SR16b]. tokamak [HVV15]. tolerance [La19]. tolerance-control [La19]. Tomographic [SAH17, CGvA^{+}15]. tomography [AMM19, AALZ15, DFRR15, LDZ17, MA19a]. tool [BCMP16, Rom18]. tools [VAMGR17]. topics [SB19]. topological [LZA19]. topology [LZS15]. TOPSIS [CCIL18]. toric [WZZ18]. Torii [PI19]. torrefaction [NGGL^{+}19]. Torregrosa [Wan18f]. Torres [DLY18]. total [LH16, LSL15]. totally [APS15a, MM19, Maz18]. tourist [MLW19]. toy [JSW17, JSW18b]. TR [Wu16b]. traces [JHPT15]. track [MCGNBD17]. tracking [GSM18, KK17, LWLL18, LHM16, LZL^{+}18b, SF16]. Tractability [HPT19]. traction [DGM16a]. Trade [ZYL18]. traded [DSDV15]. trading [CW15b]. traditional [Ran15]. traffic [AVGCMVM17, BM19b]. training [SAH17]. trait [BYWPA19]. tranches [OG16]. transaction [BE19, TF18, dFG17]. transfer [Abd18, CRCM16, CH17b, GRS19, Gia18, GMCD15, GYLQ16, SWH16, SYD17, YSCY17]. Transform [GFPGF19, CCJ16, CM18b, FM18a, GB16, HS19, LDK17, Mej17, OM17b, QG17, XYZ15, XM16]. transformation [BDSB16, CCNQ^{+}16, JK19, KK19c, LX17, OM17b, PTV17, YPD16]. transformation-free [KK19c]. transformations [CD15, GKBG17, LWZZ16, LSXL19, NSW15, OCU17, OGR17a, LGB18, OCU18, YZ15a]. transformed [DNK19, IJ18]. transforms [CY18b, DKV16, JYK17, MVC19, Xia19]. transient [AW19a, BK15b, GGC15, LZRH19]. transition [GN18, IAMCI18]. transmission [GYX17, MDVR19, Xu18, YBLH17]. transmutation [KT15]. transmuted [SKC^{+}19]. transport [BFMP19, BS17c, CSS19, CFSS18b, CLVW18, DKL^{+}18, DI19, EAKS18, FLMS16, LP19, TV16a, Tho16, VFGPG^{+}18, ZFK18, TBAH16]. transportation [RFLM16]. transported [SRA^{+}15]. transversal [ADD^{+}16, BAAA^{+}15]. trapezoidal [LWZ17]. Trasona [NGGFM17]. Traub [ACMT17a, PP18b]. traveling [ÁD14, ÁD15, ÁD17, ZLF18]. travelling [CRGB17]. treatment [Akh19, BS17a, BGL17, LPST17, LSZ16, MLV16, SB19, WZS18a]. tree [HXC19, Zha17a, ZCS^{+}19]. trees [HLLS18, LZ17a]. Trefftz [Liu17]. trends [SDG15, VACR15]. trial [LDW17]. triangle [AJR15, KF16, Pap16a, Pap16b]. triangles [AV16, BK19a, KB19]. triangular [BBC17, CDDD19, CNT17, HY19a, MPS19, NLG15, WL19b, ZZG^{+}19, ZM16b]. triangulated [HLMS19]. triangulation [WKXT15]. triangulations [BDIR19b, GK18b, HRRC19, ZWL18]. Tridiagonal [OV17, BCMP16, CDR18, HK18, JSL15, LRL^{+}18, MT16a, MT17a, SY18, OD17]. Trigonometric [WWM17, AIS^{+}17b, BS17a, LL17, NJCS19, Rót15, TLFC18]. trigonometric/hyperbolic [Rót15]. trigonometrically [DP16, FYY17, KMRS16, LWDW18, VAR15]. trigonometrically-fitted [LWDW18, VAR15]. triharmonic [KV19a]. triple [DP19b]. triples [WKBB17]. tripling [DY15]. trivariate [BIR17a, SST15]. Truncated [DRS16, DFLM19, DRST19, GLMY18, HLM18, Lan18, LX18, LW19a, Mao15, Mao16, Wan15b, ZSL18, ZJK16]. truncated/adjusted [TLTJ18]. truncation [BYW18, HB16]. trust [ESAE18, HP15, RKB16, RS16]. trust-region [ESAE18, HP15, RKB16]. Tucker [GH19b, LSH18a]. tumor [JC18, KCDACC17, TLLL19, ZHC19]. tumor-host [TLLL19]. turbine [GCTGC16]. turbulence [JSW17, JSW18b, MS18]. turbulent [EKKT17, GN18, MFWK18]. Turing [Rät15]. Turing-type [Rät15]. turning [BR15a]. TV [PKMM19, THC^{+}18]. Tweedie [KSR15]. twist [FGS19]. twisted [WS17]. Two [Che19a, DSA15, El 17, HA18, LA15c, OGR17b, OV17, PSS17, QM15, RRM18, Vab18, WR16, WL16, Yam19, Yao18, ZCH18, ZwKN15, AFBJGJ17, AN19, ABEHV16, AR15a, ABT16, AB18b, BM17b, BM18, BE15, BBQ18, BSS15, BG17, BCD15, BEHL17, BPD15b, BWS17, BRA16, BZ16a, BH17b, CJP17, CM19a, CDF^{+}16, Ceb17, CHX18, CMM17, CLZS19, Che16a, CLW17, CZLV17, CSW19, Cou19, DHAM17, DA19, DBdLdM^{+}19, DT18, DD16b, ES18, DCNY18, EAKS18, EJJ15, FEC18, FYY17, FSNS16, FGF19, FR18b, FY18, GS15b, GR19, GXZ15, GWY17, HSWH15, HHKR15, HMDSR19, HMP15, HCSC16, HB16, JKL^{+}15, KMRS16, KV18a, KGES19, KKP^{+}19, KS16, KW18, LA15a, LX15a, LLCG16, LMW18, LWDW18, LSW19, LLG19, LFCW19, LHZ15, LZ16b, LC17b, LDZ18, LFWC19, LM17b, LZRH19, MTV17, MS17a, MP15, MPH18, Mit15b, MV16c, NE17]. two [Ngo19a, OP19, OCU17, OCU18, PGRC18a, PCSR19, PG16, Qi19, RNPK15, RKS16, RCBG19, RT18a, STZ16, SE16, SW17a, SKB15, SLA19, SVAKGR19a, uIuD18, SZ18, SV16b, SO16, WKBB17, WHS17, WXY17, XHZ17, YL18b, YFS18, YSL^{+}18, ZZ19, ZZW16, Zha18, Zha17b, ZZZ15, ZCD17, ZWX18, ZH15, iW16, vSV15]. two-asset [FEC18]. two-barrier [BRA16]. two-block [CLZS19]. two-body [OCU17, OCU18, SKB15]. two-by-two [CLW17, ZZ19, ZZZ15]. Two-component [Vab18, YFS18]. Two-dimensional [PSS17, AN19, BM17b, BM18, BBQ18, BSS15, BEHL17, CDF^{+}16, Che16a, DA19, EJJ15, FSNS16, FY18, GWY17, HMDSR19, KGES19, LLG19, MS17a, NE17, STZ16, SE16, SLA19, uIuD18, SZ18, SO16, XHZ17, ZZW16, iW16]. Two-grid [ZCH18, HSWH15, ZH15]. two-layer [ABT16, DBdLdM^{+}19, LZRH19, PGRC18a]. two-layered [YL18b]. Two-level [LA15c, QM15, AB18b, CMM17, CZLV17, LX15a, Ngo19a, vSV15]. Two-parameter [OV17, HMP15, PCSR19]. two-patch [KV18a]. two-phase [CM19a, CSW19, Cou19, ES18, EAKS18, HCSC16, KS16, Mit15b, MV16c, PG16, RNPK15]. Two-point [ZwKN15, CJP17, CHX18, FGF19, GS15b, LM17b, RT18a]. two-sample [AFBJGJ17]. two-scale [BH17b, DCNY18, Zha17b]. two-side [MPH18]. two-sided [LFWC19, WXY17]. two-step [FYY17, FR18b, GXZ15, KW18, LWDW18, SW17a, SVAKGR19a, SV16b, WKBB17, WHS17]. Two-sweep [WL16]. Two-tailed [Che19a]. Type [TLTJ18, AAM18, ARCA^{+}19, AÖ17, AAKT19, AKHSU18, AYSE19, AST19, ABK19, ÁD14, ÁD15, AL16, ABEHV16, AR15c, AGM15, ABB18, BIR17a, BDIR19b, BO16, BS17b, Bog15, BAI17, BZ16a, BRZ15, CW15a, CJT15, CSTU17, CBMPP15, DP16, DL16a, DG16a, DAAF16, DGM16b, FL23, GS18b, GKN18, GMW18, GYS17, GHZ^{+}15, Güc17, Gun18, GL19, Haj19, HR15, HIS16, JM18, JS17, JMSL19, JW17, JA17, KK19a, KL15, KB16, KGMS18, LP15a, LLZ16, LA15c, LCL16, LMUZ18, LC18, LTY17, LWTQ17, MHAA15, MA19b, Mej17, MVPN15, MIT15a, MB15b, NBAK19, NP15, PM15, RMVA17, REBF16, Rät15, Reu16a, RMA19, Sha15, SJ15, SK15, Sta19, TC19, TP19, VFGPG^{+}17, WLLC18, Wan18g, WLM16, WL17c, WLC^{+}18, WH19, Xia19, XDY16, XZN^{+}19, Yak19, YLL15, YT16, ZuI17, ZNG15, ZJG^{+}15, ZG19a, ZJ19, VS15a]. Type [Wu18, YZPL18, ZSW16, MSZ18, NAK17]. type-1 [BDIR19b]. type-6 [BIR17a]. type-I [ABB18, VS15a, ZSW16]. type-II [ZG19a, YZPL18, NAK17]. types [DT18, LSBL18b, LSXL19, ML17a].

Uhlenbeck [BCNP15, WSW15]. Ulam [Xia17]. ultimate [GLP16, Wan17b]. ultra [FS19b]. ultra-weak [FS19b]. ultraspherical [HB16]. ultraspherical/Chebyshev [HB16]. unbounded [CCZZ16, CLC16, GM16a, HZ17a, Nat19, OM17a, Ugu17, ZM15a, ZWGL18]. Uncertain [XDZZ19, CD15, IN16, KV16, OAN17, WS19, ZB18]. uncertainties [APYW19, CJV17, HWP^{+}15, LWV18]. uncertainty [ASV17, BK19d, CNQRR18, CTAT19, GB16, Mej17, NP19, TA17, VFN18, YL18d]. Unconditional [CLC18a, JKL^{+}15, Ngo19a]. Unconditionally [KS16, CJLT15, LHY19, YZH18]. unconstrained [And16, And18, DZM16, Fat16, HWZ18b, HY19b, LLX17, LSDY19, LL18]. Uncorrelated [Nie15]. underdetermined [GHA^{+}15]. underlying [LKN17, Mär17]. unemployment [SSCM18]. unfitted [BV18]. Unified [AG19, GFPGF19, XD17, CLC18b, CJT15, RMVA17]. Uniform [AIL17, HS18b, HS19, Wan17b, AL18, BAHC17, Bog19, CLX17, DVA19, GDCK19, HMMV17, IL15, KD17, LT18, LI15, MHV17, MPH18, NSV18, NSYX17, RC19]. uniform-convergence [NSV18]. uniformization [ATV15]. Uniformly [YZJC15, EC18]. unifying [HY19b]. unilateral [HMS19]. Uniqueness [ZLZ17, CS19, FM19a, KLS19, SDH17, SLW17]. unit [BCA18, BCCB15, CBMO15, MS19b]. unitary [BDG15]. unity [DMP19, DP17b]. univariate [COGD19, GZDC19, Han15, OD17]. university [GR16]. unknown [AR19, GS15c, GS15d, PS15, PN16, SCP19]. unmixing [TYQ^{+}19]. unsaturated [CYY15a]. unstable [ABT17]. unsteady [CM16, GGW16, LCX18, MV16b, Yan17]. unstructured [Che16a, NR19, SLS18, WKXT15]. unsymmetric [AW19b, AvG16]. until [BRA16, CLPZ17]. update [And17, VL19]. Updating [HVV15, SVGM17, Yua15]. Updating/downdating [SVGM17]. upon [MD18, QWZ18]. upper [CNT17, GAB17, LXL15, LXL16, NTPVNTDT15, SB19, Tru19, WP17, WGZ15, Yao16]. upper-bound [NTPVNTDT15]. upper-triangular [CNT17]. Upscaling [VCL^{+}19, AEK19, CMHHS17, VCC^{+}19]. upside [DNK19]. upside-down [DNK19]. Upwind [CZ16, Jeo15, GLL19, HSWH15, LSZ16, LYS19, ZL17c]. upwind-biased [LSZ16]. urban [AVGCMVM17, BLH^{+}17, FAVGC^{+}15, FLMS16, SPG^{+}16a]. Urysohn [GKN16]. use [AS19b, DAM15b, ZS18c]. used [Mar19b]. useful [OLRPdCJ19]. user [HS17a, JCC15]. user-friendly [HS17a]. Using [BFLP15, Kre19, SJRR17, AHM18, Abd18, AG16a, AKHSU18, AM16, ALV19, AA18, AMM19, AdQdS16, AW19a, AS19a, AR15b, ABB18, AdGvG19, AKM18, BSLL19, Beh16, BDGIPG15, BGL17, BCD15, BS19, BS17c, COdSR16, CCNQ^{+}16, CCJ18, Che18, CDE18, Con18, CHVRT15, DS19, DSN17, DLM16b, DI15, DD17, EDO19, EW18, FEC18, FJLB18, GRVC15, GUB^{+}17, GH16b, GH17a, GMG19, God16, GSM18, GSM19, GD17, GGK18, GFPGF19, HVV15, HH15, HR15, HH16, Hu16, HIS16, IGRL17, Jan17a, JB18, JYK17, JN19, KV19a, KR17, KW16, KS15a, KMB15, KVO18, KtTBK18, LHK19, LW15, LCC^{+}15, LSYY19, LSH^{+}19, LHM16, LSBL18b, LOAGBR15, LC16, MAK19, MDH19, MT16b, MXW17, NN18a, Nem15, NTPVNTDT15, NSV18, NGGFM17, NLG15, Pap16b, PD17, PRP19, PBGG17, QDY^{+}19, Qui15, ROB17, RRW17]. using [RANG15, RT18b, SF18, SY19, SBP17, SLH15, SAH17, hShX17, SPG^{+}16b, TXFH18, TLFC18, Ugu18, UGG^{+}19, VFG18, VMM18, VCL^{+}19, VFGPG^{+}18, VL19, WGHR17, WLZ17, Wan18g, WYLD19, WLM16, Yam17, YZ15a, YZH18, YS19, YB17, YTSZ15, ZCST18, ZHY^{+}17, ZWZ^{+}18, ZG19c, Zyg17]. utilities [HOTZ18]. Utility [LN16, Iid15, PX17, dFG17]. utility-based [Iid15]. utilizing [AD18, HMMV17]. Uzawa [CHS15, LZ17b].

vacations [TA17]. Vademecum [FMC^{+}18]. validation [TMS^{+}17, GRVC15]. valuation [FEC18, HJ17a, OG16, WC17, XY15]. Value [BGS19, Chr19, ZX16, AAA15, AHM18, AAM18, ALS19, AS18, AZ18, AT17, AMCR19, AC17b, AP18, BKLR15, CLX17, CHX18, CHR^{+}15, CT17a, Chr18, CV19, CN16, DWXO15, DM16b, Dzh16, Dzh18, FGF19, FMR15, GSM18, GS15b, GHZ^{+}15, HS18b, HA19, HX15, IIA15, ID16, JLLZ16, KCK18, KW16, KKK16, KtTBK18, LXL15, LW17a, LLSZ17, LLW18, LC15b, LM17b, MD19, ME18, Mat18, NGGL^{+}19, NXLX18, PRP19, PTV17, PCSR19, RMVA17, RW16, RT18a, RMA19, RMK19, SK15, SRH17, SI16, TV16b, VARC17, Vis15, WPA^{+}18, ZL18a, Zak19, ZL15a, ZG17, ZLZ17, vV16, FM18a].

Value-at-Risk [ZX16, GSM18, ZG17]. valued [BDFM19, BP17a, DS17b, PAC15]. values [Arà19, HG17, LXL16, MTMZ19, NCF^{+}19, PD19, SLQ16, Wan17a]. Valuing [JHK17, ZY19a, FM19a, LDK17, Zha17a]. valves [LS17a]. vanilla [JSW18a]. vanished [SVAKGR19b]. vanishing [MHL16, XH19, yZ17]. vapour [RANG15]. vapour-liquid [RANG15]. VaR [BE19, BBB17a, BBB17b]. variability [ARCA^{+}19]. Variable [BK19b, MVQDdR^{+}18, SQRE16, ZH18, BGMS16, CTZK18, CLC18b, CLA19, DMM18, FJD17, GS18a, GTC16, GYS17, Gun18, LW17a, LLL19a, LSH^{+}19, LSXL19, MMS19, MT17b, ML19b, SYD15, SJL16b, TAT18, Xu18, XZ19]. variable-coefficient [LSXL19]. variable-order [CLA19, DMM18]. variable-separation [GTC16]. variables [BE15, (Ba18, DHAM17, GAB17, GR19, HYZ16, LHK19, LOAGBR15, NL17, NC19, OLPT16, RdSMO17, RdSMO18]. variance [CP15, CW15b, DT17, GAB17, LRZ15, LC16, WW19b]. variant [LH19, RS18, Wan18f]. variants [DRST19, SHCJ19]. variation [BDFM19, DD16b, Don18, LSL15, Par15, ZYXT16, ZXQH17, ZKRL15]. Variational [Aks17, BK15a, FMC^{+}18, RW15, AK17b, BKLR16, DMMSK17, DTKP17, El 17, GGG19, HH19a, HCA^{+}18, HMDSR19, LA15c, LH16, LWL17, MHAA15, Mar19b, NS19, PSW18, RS16, SSG16, SO16, TAT18, TH18, XHH^{+}19, ZZZ18]. variational-hemivariational [HH19a, XHH^{+}19]. variations [AM19]. variety [LMM19]. various [CN18, KS16]. varying [AG18, AY19, APYW19, AYSE19, AS19a, EPC17, FE16, GYL16, HKZZ15, JJX17, LJ15, LMUZ18, LDW17, LYG15, MMMMB16, SZZ^{+}16, SJL16b, YY16, ZHH^{+}19, ZWWW19]. varying-coefficient [LYG15, SJL16b]. vascularisation [KHS^{+}18]. vector [CZ15, Che16a, DD16b, LAVMA18, MD18, MRMA17, RVA15, RdV19, RRW17, TGSS17, WS18]. Vectorial [LWZ^{+}15]. Vectors [DVJ18, BAHC17, CJ16, Han19, PW17, SMTdF18]. vehicle [SQRE16]. Velocity [AW19a, BBGG19, DD18b, LHM16, MPM^{+}17]. velocity-pseudostress [BBGG19]. Ventcell [Xu18]. ventilation [AMM19]. verification [HZ19, LW17b, MTKO17, Wat16]. Verified [Miy19, HP17, Miy17, Miy18]. versions [GHM17]. versus [Mär17]. vertex [DEY19]. vertex-weighted [DEY19]. vertical [ZqZwZ15]. vertices [AHS19]. very [LS18a]. vessel [ABT16]. via [ALS19, AAKT19, AEK19, BC16, BL17, CLS18, CCZ^{+}18, CT16, CSW19, CPL^{+}18, CPL^{+}19b, DKV16, DP17b, DM17a, DMR19, DSA16, FJ19, FZW19, GKM18, GAA16, GHL18, HSAR18, HJ17a, HMDSR19, JD16, LDK17, LSH18a, LWTQ17, LX17, MHAA15, MA18, OZ18, PB15, Prz19, dHKdHMM15, Ros17, SAC17, XY19, YBC^{+}16, YWW19, dAFvdMR19]. vibrating [WD16, XCG19]. view [CJT15]. violence [LE18]. virial [CDH^{+}17]. virtual [ZYF19]. virus [JCC17, JCC18, JCC19, LRRV18]. visco [VMR17]. visco-elastic [VMR17]. viscoelastic [FS15, Mar19b, PBP16, ZQJY18, ZFK18]. viscoelasticity [AFdOdS15]. viscosity [Sey18]. viscous [Hei15, PST18, SLH15, Wat16, YZH18]. visibility [BDAG15]. visited [FLM17]. visual [LWLL18]. visualization [DS15a]. Vlasov [YB17]. VMS [LSH^{+}18b]. Voigt [PBP16]. volatility [CP15, DI15, DM17b, DP19c, GVMR19, GB18, HYF18, HT18, LM17a, MFLL18, NM17, OT17, ST16b, SCG18, XY15, YL18d, ZS17a, ZL15c]. voltage [BIJM^{+}19]. Volterra [FL23, AFH16, ABB17, AL16, BM17b, BG17, Bie17, Bog15, GYS17, GAA16, HS17b, HSAA19, KS17a, KMB15, LYG17, MFX15, MD18, MII16, ML19b, MHL16, MH16a, MH16b, MS17c, MSH18, MAS19, MTS17, NN18a, Nem15, Pis17, Reu16a, SLH18, Sar19, SK16, WR16, WCS18, XHZ17, XY19, XF16, YAA19, ZM15b, yZ17, ZX19]. Volume [BG15, HMMV17, AM16, AW19b, ABT16, BSS15, BCS17, CWXW17, CSW19, DF15, ES18, GLL19, GMCD15, HSWH15, KHS^{+}18, LZC15, LGYC19, LYS19, MC19, NSYX17, RNPK15, SKBT16, WL19b, Xin16, YSL^{+}18, ZKRL15]. volume-finite [BCS17, ES18]. Volume-of-Fluid [BG15]. volumes [Rót15]. volumetric [NCQ^{+}18, XMW^{+}15]. Voronoi [WKXT15]. vortex [Uda18, Xu17]. vortices [MF16]. vorticity [MV16b, SO16]. Vries [AKHSU18, BGS19, FS19b]. vs [FR18a]. vulnerable [JYK17, PG19].

wake [Xu17]. wake-vortex [Xu17]. walkers [ZWZ^{+}18]. walks [CRR19, SPCSA16]. wall [LLSZ17, SS18b]. Walsh [God15]. Wanner [Ran15]. Warburton [AMP18]. warping [WCS17]. Wasserstein [DI19]. waste [LYS19]. wasted [Kav16]. wastewater [AVGCMVM15]. water [ABT16, BKN16, BLH^{+}17, CL19b, CFMN15, FdGKZ17, IABIPG17, IAMCI18, LSG18, LFKG16, hShX17, Xin16, YSL^{+}18]. wave [ÁD14, ÁD15, ÁD17, ABT16, AK17b, BBB17c, BB18, CRGB17, CRR16, CLC18b, CY15, CROC16, DSA15, DAM15b, DLM17, GM16a, GDCK19, GHL19, HA18, HMH18, JGGS16, KM18, LHJ16, LWL17, LLL19a, Liu17, LXZ18, MM18, MS15b, Per16, RQ16, SBU^{+}17, SLA19, SS19b, SFGJR^{+}16, SRH17, VG15, WSL19, XH15, YL18b, YH18b, ZM15a, ZJ17a, Zha19a, Zha17b]. wave-penetrable [LXZ18, YL18b]. waveform [AL19b, Lad15]. waveguide [LXZ17, LXZ18]. waveguides [LLZ18]. wavelet [BM17a, BM17b, CRSV19, DFRR15, EW18, GWT18, KR17, Liu18, LSL15, MC18, NGGFM17, ROB17, SK16, SGG19, UPD^{+}15, Vis19, YL18c]. wavelet-Galerkin [YL18c]. Wavelets [ABBB17, CS18a, OG16, PM15, PSS17, RdV19, Sar19, YTSZ15]. wavelets-based [OG16]. waves [BN19, BYYV16, BKN16, CL16, FdGKZ17, FdOP17, PSS17, ZLF18, ZFK18]. way [BD19, NCF^{+}19]. Weak [BDAG15, CCX18, CWH18, GCZ17, KHLY17, Mu19b, AHMY19, AH16a, AH15, BW19, CZ19b, FS19b, HWZ18a, HMY19, ID18, JC15, JMS17, JSW17, JSW18b, KGMS18, LX15a, LCH18, LGLC18, LLSM15, LLC18, LW19c, MWY15b, MWY15a, MWY16a, MWY16b, MZ19, Mu19a, SSW19, TZZ18, WWWZ16, WWZZ16, WZZ16, Wan18a, WZWJ19, Wu19, Yam17, Yi19, ZL18b, ZL19, ZYF19]. Weaker [LLW18]. weakly [BDA17, BS19, Bie17, CXZ19, DA19, MMO15, MN19a, NSM16, PMN19, PPM16, SNY19, WR16, WCS18, ZZ15]. wealth [JS19a]. Weber [Yak19]. Weibull [GH17b, NAK17, NADK18, RABB19, Wan18c, WL17c, YZPL18, ZSW16]. weight [CCT19, GKM18, LA15b, LKN17, ZCST18]. Weighted [HWC16a, JL16a, RR18, DKP^{+}19, DEY19, DKLP15, GLL19, God15, HPT19, JN18, LJX17, MOY18, MP18, MM17, QDY^{+}19, RW15, Wan18h, Xie15, YW16, ZY19b]. weights [DKLP15, MM17, PM15, RIK16, Yam17, ZTGZ18]. Weiner [LLD18, LHM16]. Well [DBX19, LMA18, AMZ19, CL19b, LSG18, ML18, SSCM18, Zha15b, ZH16]. well-balanced [CL19b, LSG18]. well-being [SSCM18]. well-conditioned [ML18]. Well-posedness [DBX19, AMZ19]. Wendroff [LSZ16]. WENO [CLZW19, GJ17, JN18, Xin16]. WENO- [JN18]. WENOZ [HWC16b]. Werner [AR15c]. Werner-type [AR15c]. wetting [TKH16]. WG [CFX17]. wheelset [MCGNBD17]. wheelset-track [MCGNBD17]. where [NS16, RVA15]. which [PS15]. while [BHMT16]. White [FNV17]. Whitham [BKN16, BAI17, CRGB17]. Wick [LHM16]. Wick-stochastic [LHM16]. wide [FCJ16]. width [LWW17]. widths [FS19a]. Wiener [Luo18, STZ16]. Wigner [SD15, SSD16]. Wilbraham [BCA18]. Willmore [DD16c]. Willmore-like [DD16c]. wind [TLL16]. window [GB18]. wing [PST18]. Winslow [XLS^{+}19]. Wirtinger [YWW19]. wise [GDCK19, VFGPG^{+}18, WHZ16]. Wishart [NM17]. without [ABL^{+}19, AEP15, HP15]. Wolfe [JJ19]. Woodbury [WLZ17]. world [AdHR17]. wormhole [GTX^{+}19, KSW19]. Worsey [SST15]. Worst [IN16, MOY18]. woven [DWS^{+}18]. WR [Yan17]. WRF [SPG^{+}16b]. Wright [Meh19, Sta18]. written [LN16]. WZ [Byl18].

X [PGD^{+}18]. X-ray [PGD^{+}18]. Xeon [GGK18]. xoroshiro128 [LO19]. xorshift [Vig17]. xorshift1024 [LO19]. Xorshift1024* [LO19]. xorshift128 [LO19].

Y. [DLY18]. Yang [ZCD17]. yellow [BdSFL16]. Yuan [JJ19].

Zakharov [Su18]. Zernike [ADG17]. Zero [HRRC19, ÁAM15, GKN18, GKM18, GD18, KH18, Qi19, QLZ19, TLTJ18, TBS18]. zeroand-one [QLZ19]. zero-truncated [TLTJ18]. zerotruncated/adjusted [TLTJ18]. zeroing [ZHH^{+}19]. zeros [ABK19, CV17, KKP^{+}19, MMD15, MB15b, PP18a, PP18b, PP19, PI19, QLZ19, SH19]. Zhang [SSG18]. Zhou [DLY18]. Zipf [KPP19]. ZNN [ZJG^{+}15]. zones [FAVGC^{+}15]. zooplankton [AJ18].

References

Abbasbandy:2016:SEE

[AA16] Saeid Abbasbandy and Babak Azarnavid. Some error estimates for the reproducing kernel Hilbert spaces method. Journal of Computational and Applied Mathematics, 296 (??):789–797, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                                        http://www.sciencedirect.
                                                        com/science/article/
                                                        pii/S0377042715005464.

Akhmouch:2017:CDS

[AA17] M. Akhmouch and M. Benzakour Amine. A corrected decoupled scheme for chemotaxis models. Journal of Computational and Applied Mathematics, 323(??):36–52, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                                        www.sciencedirect.com/
                                                        science/article/pii/
                                                        S0377042717301565.

Alipour:2018:NCT

[AA18] Mohsen Alipour and Hamzeh Agahi. New computational techniques for solving nonlinear problems using gfractional differential operator. Journal of Computational and Applied Mathematics, 330(??):70–74, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730393X.

Abbasbandy:2015:SRK

[AAA15] Saeid Abbasbandy, Babak Azarnavid, and Mohammed S. Alhuthali. A shooting reproducing kernel Hilbert space method for multiple solutions of nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 279(??):293–305, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004932.

Agouzal:2018:NAS

[AAB18a] Abdellatif Agouzal, Karam Allali, and Siham Binna. Numerical analysis of insitu biodegradation model in porous media. Journal of Computational and Applied Mathematics, 344 (??):190–228, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830298X.

Amat:2018:LCS

[AAB^{+}18b] S. Amat, I. K. Argyros, S. Busquier, M. A. Hernández-Verón, and E. Martínez. On the local convergence study for an efficient k-step iterative method. Journal of Computational and Applied Mathematics, 343(??):753–761, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301079.

Ashyralyev:2017:BSD

[AAC17] Allaberen Ashyralyev, Deniz Agirseven, and Burcu Ceylan. Bounded solutions of delay nonlinear evolutionary equations. Journal of Computational and Applied Mathematics, 318(??):69– 78, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630588X.

Ahmad:2019:HHH

[AAKT19] Bashir Ahmad, Ahmed Alsaedi, Mokhtar Kirane, and Berikbol T. Torebek. Hermite–Hadamard, Hermite–Hadamard–Fejér, Dragomir–Agarwal and Pachpatte type inequalities for convex functions via new fractional integrals. Journal of Computational and Applied Mathematics, 353(??):120– 129, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307611.

Asante-Asamani:2016:RDP

[AAKW16] E. O. Asante-Asamani, A. Q. M. Khaliq, and B. A. Wade. A real distinct poles exponential time differencing scheme for reaction–diffusion systems. Journal of Computational and Applied Mathematics, 299(??):24– 34, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004690.

Abia:2018:ASP

[AALMLM18] L. M. Abia, O. Angulo, J. C. López-Marcos, and M. A. López-Marcos. Approximating the survival probability in finite life-span population models. Journal of Computational and Applied Mathematics, 330(??):783–793, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302406.

Abia:2019:NAC

[AALMLM19] L. M. Abia, O. Angulo, J. C. López-Marcos, and M. A. López-Marcos. Numerical analysis of a cell dwarfism model. Journal of Computational and Applied Mathematics, 349(??):82–92, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305892.

Arroyo:2015:CBC

[AALZ15] Fangjun Arroyo, Edward Arroyo, Xiezhang Li, and Jiehua Zhu. The convergence of the block cyclic projection with an overrelaxation parameter for compressed sensing based tomography. Journal of Computational and Applied Mathematics, 280(??):59–67, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005159.

Alvarez:2015:PSA

[ÁAM15] Gabriel Álvarez, Luis Martínez Alonso, and Elena Medina. Phase structure and asymptotic zero densities of orthogonal polynomials in the cubic model. Journal of Computational and Applied Mathematics, 284(??):10–25, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005330.

Abdeljawad:2018:DML

[AAM18] Thabet Abdeljawad and Qasem M. Al-Mdallal. Discrete Mittag-Leffler kernel type fractional difference initial value problems and Gronwall’s inequality. Journal of Computational and Applied Mathematics, 339(??):218–230, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305228.

Axelsson:2016:CPR

[AB16] Owe Axelsson and Radim Blaheta. Corrigendum to “Preconditioners for regularized saddle point problems with an application for heterogeneous Darcy flow problems” [J. Comput. Appl. Math. 280 (2015) 141– 157]. Journal of Computational and Applied Mathematics, 298(??):252–255, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005889. See

[ABB^{+}15].

Ahmed:2018:SCS

[AB18a] Elyes Ahmed and Amel Ben Abda. The sub-Cauchy–Stokes problem: Solvability issues and Lagrange multiplier methods with artificial boundary conditions. Journal of Computational and Applied Mathematics, 338(??):258–279, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300712.

Axelsson:2018:LRI

[AB18b] Owe Axelsson and Radim Blaheta. Low-rank improvements of two-level grid preconditioned matrices. Journal of Computational and Applied Mathematics, 340(??):432–442, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304582.

Axelsson:2015:PRS

[ABB^{+}15] Owe Axelsson, Radim Blaheta, Petr Byczanski, János Karátson, and Bashir Ahmad. Preconditioners for regularized saddle point problems with an application for heterogeneous Darcy flow problems. Journal of Computational and Applied Mathematics, 280(??):141–157, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005238. See

corrigendum [AB16].

Ahmad:2017:IVS

[ABB17] Mian Ilyas Ahmad, Ulrike Baur, and Peter Benner. Implicit Volterra series interpolation for model reduction of bilinear systems. Journal of Computational and Applied Mathematics, 316(??):15– 28, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304770.

Asl:2018:CBI

[ABB18] Mehri Noori Asl, Reza Arabi Belaghi, and Hossien Bevrani. Classical and Bayesian inferential approaches using Lomax model under progressively type-I hybrid censoring. Journal of Computational and Applied Mathematics, 343(??):397–412, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302097.

Amat:2017:WME

[ABBB17] Sergio Amat, Pedro J. Blázquez, Sonia Busquier, and Concepción Bermúdez. Wavelets for the Maxwell’s equations: an overview. Journal of Computational and Applied Mathematics, 321(??):555–565, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300730.

Abdeljawad:2015:CFC

[Abd15] Thabet Abdeljawad. On conformable fractional calculus. Journal of Computational and Applied Mathematics, 279(??):57– 66, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004622.

Abdi:2016:CHO

[Abd16] A. Abdi. Construction of high-order quadratically stable second-derivative general linear methods for the numerical integration of stiff ODEs. Journal of Computational and Applied Mathematics, 303(??):218–228, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630111X.

Abdelhamid:2018:SIS

[Abd18] Talaat Abdelhamid. Simultaneous identification of the spatio-temporal dependent heat transfer coefficient and spatially dependent heat flux using an MCGM in a parabolic system. Journal of Computational and Applied Mathematics, 328(??):164–176, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303394.

Amat:2016:STM

[ABEHV16] S. Amat, S. Busquier, J. A. Ezquerro, and M. A. Hernández-Verón. A Steffensen type method of two steps in Banach spaces with applications. Journal of Computational and Applied Mathematics, 291(??):317–331, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001910.

Ahmad:2017:KSB

[ABG17] Mian Ilyas Ahmad, Peter Benner, and Pawan Goyal. Krylov subspace-based model reduction for a class of bilinear descriptor systems. Journal of Computational and Applied Mathematics, 315(??):303–318, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305428.

Abdulla:2019:GMH

[ABH19] Ugur G. Abdulla, Vladislav Bukshtynov, and Ali Hagverdiyev. Gradient method in Hilbert-Besov spaces for the optimal control of parabolic free boundary problems.

Journal of Computational and Applied Mathematics, 346(??):84–109, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303959.

Amat:2016:ESI

[ABHVM16] S. Amat, C. Bermúdez, M. A. Hernández-Verón, and Eulalia Martínez. On an efficient k-step iterative method for nonlinear equations. Journal of Computational and Applied Mathematics, 302(??):258–271, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300437.

Alshomrani:2019:ORC

[ABK19] Ali Saleh Alshomrani, Ramandeep Behl, and V. Kanwar. An optimal reconstruction of Chebyshev–Halley type methods for nonlinear equations having multiple zeros. Journal of Computational and Applied Mathematics, 354(??):651– 662, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307702.

Aitbayev:2019:MSP

[ABL^{+}19] Rakhim Aitbayev, Peter W. Bates, Hong Lu, Lijun Zhang, and Mingji Zhang. Mathematical studies of Poisson–Nernst– Planck model for membrane channels: Finite ion size effects without electroneutrality boundary conditions. Journal of Computational and Applied Mathematics, 362 (??):510–527, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306459.

Agarwal:2018:SFF

[ABN^{+}18] Ravi P. Agarwal, Dumitru Baleanu, Juan J. Nieto, Delfim F. M. Torres, and Yong Zhou. A survey on fuzzy fractional differential and optimal control nonlocal evolution equations. Journal of Computational and Applied Mathematics, 339(??):3– 29, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304703.

Andreoli:2015:CSP

[ABP15] Alessandro Andreoli, Luca Vincenzo Ballestra, and Graziella Pacelli. Computing survival probabilities based on stochastic differential models. Journal of Computational and Applied Mathematics, 277(??):127–137, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003872.

Akram:2017:SPP

[ABRS17] Ghazala Akram, Khalida Bibi, Kashif Rehan, and Shahid S. Siddiqi. Shape preservation of 4-point interpolating non-stationary subdivision scheme. Journal of Computational and Applied Mathematics, 319(??):480–492, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730047X.

Ardakani:2016:AWF

[ABT16] H. Alemi Ardakani, T. J.

Bridges, and M. R. Turner. Adaptation of f-wave finite volume methods to the two-layer shallowwater equations in a moving vessel with a rigidlid. Journal of Computational and Applied Mathematics, 296(??):462–479, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004781.

Ableidinger:2017:IST

[ABT17] M. Ableidinger, E. Buckwar, and A. Thalhammer. An importance sampling technique in Monte Carlo methods for SDEs with a.s. stable and meansquare unstable equilibrium. Journal of Computational and Applied Mathematics, 316(??):3– 14, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304125.

Alotaibi:2017:GLM

[AC17a] Manal Alotaibi and Eric Chung. Global-local model reduction for heterogeneous Forchheimer flow. Journal of Computational and Applied Mathematics, 321(??):160–184, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042717300985.

Auchmuty:2017:SAH

[AC17b] Giles Auchmuty and Manki Cho. Steklov approximations of harmonic boundary value problems on planar regions. Journal of Computational and Applied Mathematics, 321 (??):302–313, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300997.

Amiraliyev:2017:HOF

[ACAC17] Gabil M. Amiraliyev, Erkan Cimen, Ilhame Amirali, and Musa Cakir. High-order finite difference technique for delay pseudo-parabolic equations. Journal of Computational and Applied Mathematics, 321(??):1–7, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300821.

Aouadi:2019:AMT

[ACCF19] M. Aouadi, M. Campo, M. I. M. Copetti, and J. R. Fernández. Analysis of a multidimensional thermoviscoelastic contact problem under the Green– Lindsay theory. Journal of Computational and Applied Mathematics, 345: 224–246, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Alcazar:2015:NMC

[ACDT15] Juan Gerardo Alcázar, Jorge Caravantes, and Gema M. Diaz-Toca. A new method to compute the singularities of offsets to rational plane curves. Journal of Computational and Applied Mathematics, 290(??):385–402, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003209.

Amiri:2018:POC

[ACDT18] A. R. Amiri, A. Cordero, M. T. Darvishi, and J. R. Torregrosa. Preserving the order of convergence: Low-complexity Jacobianfree iterative schemes for solving nonlinear systems. Journal of Computational and Applied Mathematics, 337(??):87–97, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042718300256.

Amiri:2019:FAS

[ACDT19] Abdolreza Amiri, Alicia Cordero, Mohammad Taghi Darvishi, and Juan R. Torregrosa. A fast algorithm to solve systems of nonlinear equations. Journal of Computational and Applied Mathematics, 354(??):242– 258, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302826.

Araujo-Cabarcas:2018:ERC

[ACEJ18] Juan Carlos Araujo-Cabarcas, Christian Engström, and Elias Jarlebring. Efficient resonance computations for Helmholtz problems based on a Dirichletto-Neumann map. Journal of Computational and Applied Mathematics, 330 (??):177–192, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042717304016.

Angulo:2018:NIE

[ACLM18] O. Angulo, F. Crauste, and J. C. López-Marcos. Numerical integration of an erythropoiesis model with explicit growth factor dynamics. Journal of Computational and Applied Mathematics, 330 (??):770–782, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042717300547.

Adjerid:2019:IDF

[ACLY19] Slimane Adjerid, Nabil Chaabane, Tao Lin, and Pengtao Yue. An immersed discontinuous finite element method for the Stokes problem with a moving interface. Journal of Computational and Applied Mathematics, 362 (??):540–559, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S037704271830459X.

Argyros:2017:TDA

[ACMT17a] Ioannis K. Argyros, Alicia Cordero, Á. Alberto Magreñán, and Juan R. Torregrosa. Third-degree anomalies of Traub’s method. Journal of Computational and Applied Mathematics, 309(??):511–521, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300425.

Argyros:2017:CHO

[ACMT17b] Ioannis K. Argyros, Alicia Cordero, Ángel Alberto Magreñán, and Juan Ramón Torregrosa. On the convergence of a higher order family of methods and its dynamics. Journal of Computational and Applied Mathematics, 309(??):542–562, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301959.

Alvarez:2014:PTM

[ÁD14] J. Álvarez and A. Durán.

Petviashvili type methods for traveling wave computations: I. Analysis of convergence. Journal of Computational and Applied Mathematics, 266(??):39–51, August 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714000363. See

corrigendum [ÁD15].

Alvarez:2015:CPT

[ÁD15] J. Álvarez and A. Durán.

Corrigendum to “Petviashvili type methods for traveling wave computations: I. Analysis of convergence” [J. Comput. Appl. Math. 266 (2014) 39–51]. Journal of Computational and Applied Mathematics, 277(??):215–216, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004002. See

[ÁD14].

Alvarez:2017:NGP

[ÁD17] J. Álvarez and A. Durán. Numerical generation of periodic traveling wave solutions of some nonlinear dispersive wave systems. Journal of Computational and Applied Mathematics, 316(??):29– 39, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630406X.

Assari:2018:MGS

[AD18] Pouria Assari and Mehdi Dehghan. A meshless Galerkin scheme for the approximate solution of nonlinear logarithmic boundary integral equations utilizing radial basis functions. Journal of Computational and Applied Mathematics, 333(??):362– 381, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305800.

Alessio:2016:DGN

[ADD^{+}16] O. Aléssio, M. Düldül, B. Uyar Düldül, Nassar H. Abdel-All, and Sayed Abdel-Naeim Badr. Differential geometry of non-transversal intersection curves of three implicit hypersurfaces in Euclidean 4-space. Journal of Computational and Applied Mathematics, 308(??):20–38, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630228X.

Area:2017:RCG

[ADG17] I. Area, Dimitar K. Dimitrov, and E. Godoy. Recursive computation of generalised Zernike polynomials. Journal of Computational and Applied Mathematics, 312(??):58– 64, March 1, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0377042715005622.

Aimi:2019:RMP

[ADG19] A. Aimi, M. Diligenti, and C. Guardasoni. Restrictions matrices for platonic solids invariance and applications to space-time energetic BEM. Journal of Computational and Applied Mathematics, 360(??):18–40, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271930158X.

Astudillo:2019:AID

[AdGvG19] R. Astudillo, J. M. de Gier, and M. B. van Gijzen. Accelerating the Induced Dimension Reduction method using spectral information. Journal of Computational and Applied Mathematics, 345: 33–47, January 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Alfonso:2017:FRM

[AdHR17] G. Alfonso, A. F. Roldán López de Hierro, and C. Roldán. A fuzzy regression model based on finite fuzzy numbers and its application to real-world financial data. Journal of Computational and Applied Mathematics, 318(??):47– 58, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305945.

Aledo:2019:PGE

[ADMV19] Juan A. Aledo, Luis G.

Diaz, Silvia Martinez, and Jose C. Valverde. Predecessors and Garden-of-Eden configurations in parallel dynamical systems on maxterm and minterm Boolean functions. Journal of Computational and Applied Mathematics, 348(??):26–33, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304928.

Alves:2016:SCP

[AdQdS16] J. R. Alves, R. A. B.

de Queiroz, and R. W. dos Santos. Simulation of cardiac perfusion by contrast in the myocardium using a formulation of flow in porous media. Journal of Computational and Applied Mathematics, 295(??):13–24, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002186.

Argyros:2018:EDS

[AEHVM18] I. K. Argyros, J. A. Ezquerro, M. A. Hernández-Verón, and Á.A. Magreñán. Extending the domain of starting points for Newton’s method under conditions on the second derivative. Journal of Computational and Applied Mathematics, 340(??):1–10, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300530.

Arjmand:2019:TUM

[AEK19] Doghonay Arjmand, Stefan Engblom, and Gunilla Kreiss. Temporal upscaling in micromagnetism via heterogeneous multiscale methods. Journal of Computational and Applied Mathematics, 345:99–113, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Akbari:2019:MDS

[AEKGP19] Hani Akbari, Allan P. Engsig-Karup, Victor Ginting, and Felipe Pereira. A multiscale direct solver for the approximation of flows in high contrast porous media. Journal of Computational and Applied Mathematics, 359(??):88–101, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301530.

Arouxet:2015:IRM

[AEP15] Ma. Belén Arouxét, Nélida E.

Echebest, and Elvio A. Pilotta. Inexact restoration method for nonlinear optimization without derivatives. Journal of Computational and Applied Mathematics, 290(??):26–43, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002733.

Alba-Fernandez:2017:CTT

[AFBJGJ17] M. V. Alba-Fernández, A. Batsidis, M. D. Jiménez-Gamero, and P. Jodrá. A class of tests for the twosample problem for count data. Journal of Computational and Applied Mathematics, 318(??):220– 229, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716304794.

Azhdari:2015:DVE

[AFdOdS15] Ebrahim Azhdari, José A. Ferreira, Paula de Oliveira, and Pascoal M. da Silva. Diffusion, viscoelasticity and erosion: Analytical study and medical applications. Journal of Computational and Applied Mathematics, 275(??):489–501, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714000533.

Aryana:2018:SSS

[AFGT18] S. Aryana, F. Furtado, V. Ginting, and P. Torsu. On series solution for second order semilinear parabolic IBVPs. Journal of Computational and Applied Mathematics, 330 (??):499–518, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717304132.

Abdi:2016:CEG

[AFH16] A. Abdi, S. Fazeli, and G. Hojjati. Construction of efficient general linear methods for stiff Volterra integral equations. Journal of Computational and Applied Mathematics, 292(??):417–429, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003908.

Audu:2019:MFE

[AFM19] J. D. Audu, F. Fairag, and K. Mustapha. Mixed finite element analysis for generalized Darcy–Forchheimer model in porous media. Journal of Computational and Applied Mathematics, 353(??):191– 203, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307349.

Axelsson:2017:POC

[AFN17] Owe Axelsson, Shiraz Farouq, and Maya Neytcheva. A preconditioner for optimal control problems, constrained by Stokes equation with a timeharmonic control. Journal of Computational and Applied Mathematics, 310(??):5–18, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716302631.

Ahmed:2016:CAE

[AG16a] S. Reaz Ahmed and Abhishek Kumar Ghosh. Computational analysis of elastic field of deep curved beams/rings using displacement-function equilibrium method. Journal of Computational and Applied Mathematics, 308 (??):464–487, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302217.

Alyoubi:2016:PMF

[AG16b] A. Alyoubi and M. Ganesh. Parallel mixed FEM simulation of a class of singlephase models with nonlocal operators. Journal of Computational and Applied Mathematics, 307(??):106–118, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301200.

Ali:2018:SDS

[AG18] M. Syed Ali and N. Gunasekaran. Sampled-data state estimation of Markovian jump static neural networks with interval time-varying delays. Journal of Computational and Applied Mathematics, 343(??):217–229, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302280.

Argyros:2019:UCA

[AG19] Ioannis K. Argyros and Santhosh George. Unified convergence analysis of frozen Newton-like methods under generalized conditions. Journal of Computational and Applied Mathematics, 347(??):95–107, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304874.

Argyros:2015:LCM

[AGM15] Ioannis K. Argyros, Santhosh George, and Á. Alberto Magreñán. Local convergence for multipoint-parametric Chebyshev– Halley-type methods of high convergence order. Journal of Computational and Applied Mathematics, 282(??):215–224, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042714005792.

Aledo:2019:ARA

[AGM19] Juan A. Aledo, Jose A. Gámez, and David Molina. Approaching the rank aggregation problem by local search-based metaheuristics. Journal of Computational and Applied Mathematics, 354(??):445– 456, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301602.

Algaba:2015:MCA

[AGR15a] A. Algaba, C. García, and M. Reyes. Monodromy of a class of analytic generalized nilpotent systems through their Newton diagram. Journal of Computational and Applied Mathematics, 287(??):78–87, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001648.

Area:2015:BKP

[AGR^{+}15b] I. Area, E. Godoy, J. Rodal, A. Ronveaux, and A. Zarzo. Bivariate Krawtchouk polynomials: Inversion and connection problems with the NAVIMA algorithm. Journal of Computational and Applied Mathematics, 284(??):50–57, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005019.

Amiri:2015:CWS

[AH15] Sadegh Amiri and S. Mohammad Hosseini. A class of weak second order split-drift stochastic Runge–Kutta schemes for stiff SDE systems. Journal of Computational and Applied Mathematics, 275(??):27–43, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003422.

Aghda:2016:LMS

[AH16a] A. S. Fatemion Aghda and S. Mohammad Hosseini. Linear meansquare stability properties of semi-implicit weak order 2.0 Taylor schemes for systems of stochastic differential equations. Journal of Computational and Applied Mathematics, 302(??):94–105, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300231.

Alcazar:2016:IPP

[AH16b] Juan Gerardo Alcázar and Carlos Hermoso. Involutions of polynomially parametrized surfaces. Journal of Computational and Applied Mathematics, 294(??):23–38, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004094.

Argyros:2016:MMT

[AH16c] Ioannis K. Argyros and Saïd Hilout. The majorant method in the theory of Newton–Kantorovich approximations and generalized Lipschitz conditions. Journal of Computational and Applied Mathematics, 291(??):332–347, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271400569X.

Ahn:2018:SGL

[AH18] Young Joon Ahn and Christoph Hoffmann. Sequence of G^{n} LN polynomial curves approximating circular arcs. Journal of Computational and Applied Mathematics, 341(??):117–126, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301912.

Akbari:2019:PRR

[AH19] Mohammad Ghasem Akbari and Gholamreza Hesamian. A partialrobust-ridge-based regression model with fuzzy predictors-responses. Journal of Computational and Applied Mathematics, 351 (??):290–301, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S037704271830671X.

Ait-Haddou:2015:EGQ

[AHBC15] Rachid Ait-Haddou, Michael Bartoň, and Victor Manuel Calo. Explicit Gaussian quadrature rules for C^{1} cubic splines with symmetrically stretched knot sequences. Journal of Computational and Applied Mathematics, 290 (??):543–552, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003301.

Abidi:2017:GRA

[AHJ17] O. Abidi, M. Hached, and K. Jbilou. A global rational Arnoldi method for model reduction. Journal of Computational and Applied Mathematics, 325(??):175–187, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730239X.

Auzinger:2019:SLE

[AHK19] Winfried Auzinger, Harald Hofstätter, and Othmar Koch. Symmetrized local error estimators for time-reversible onestep methods in nonlinear evolution equations. Journal of Computational and Applied Mathematics, 356(??):339–357, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300822.

Auzinger:2015:DBL

[AHKT15] Winfried Auzinger, Harald Hofstätter, Othmar Koch, and Mechthild Thalhammer. Defect-based local error estimators for splitting methods, with application to Schrödinger equations, part III: the nonlinear case. Journal of Computational and Applied Mathematics, 273(??):182–204, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002854.

Ashpazzadeh:2017:BMI

[AHL17] Elmira Ashpazzadeh, Bin Han, and Mehrdad Lakestani. Biorthogonal multiwavelets on the interval for numerical solutions of Burgers’ equation. Journal of Computational and Applied Mathematics, 317(??):510– 534, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305878.

Abdallah:2018:SAV

[AHM18] L. Abdallah, M. Haddou, and T. Migot. Solving absolute value equation using complementarity and smoothing functions. Journal of Computational and Applied Mathematics, 327(??):196–207, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303199.

Adler:2019:PEE

[AHMY19] James H. Adler, Xiaozhe Hu, Lin Mu, and Xiu Ye. An a posteriori error estimator for the weak Galerkin leastsquares finite-element method. Journal of Computational and Applied Mathematics, 362(??):383–399, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718306010.

Ahn:2015:NDC

[Ahn15] Jae Youn Ahn. Negative dependence concept in copulas and the marginal free herd behavior index. Journal of Computational and Applied Mathematics, 288(??):304–322, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002381.

Ahn:2019:CAG

[Ahn19] Young Joon Ahn. Circle approximation by G^{2} Bézier curves of degree n with 2n−1 extreme points. Journal of Computational and Applied Mathematics, 358(??):20–28, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301190.

Abdi:2019:ALB

[AHP19] A. Abdi, S. A. Hosseini, and H. Podhaisky. Adaptive linear barycentric rational finite differences method for stiff ODEs. Journal of Computational and Applied Mathematics, 357(??):204–214, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301165.

Anisimov:2019:BET

[AHS19] Dmitry Anisimov, Kai Hormann, and Teseo Schneider. Behaviour of exponential three-point coordinates at the vertices of convex polygons. Journal of Computational and Applied Mathematics, 350(??):114–129, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305995.

Aghda:2018:CNN

[AHT18] A. S. Fatemion Aghda, Seyed Mohammad Hosseini, and Mahdieh Tahmasebi. Convergence and non-negativity preserving of the solution of balanced method for the delay CIR model with jump. Journal of Computational and Applied Mathematics, 344(??):676–690, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303613.

Artes:2017:UIC

[AIL17] Joan C. Artés, Jackson Itikawa, and Jaume Llibre. Uniform isochronous cubic and quartic centers: Revisited. Journal of Computational and Applied Mathematics, 313(??):448–453, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304381.

Aishima:2017:CIP

[Ais17a] Kensuke Aishima. On convergence of iterative projection methods for symmetric eigenvalue problems. Journal of Computational and Applied Mathematics, 311(??):513–521, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304046.

Alonso:2017:EAA

[AIS^{+}17b] Pedro Alonso, Javier Ibáñez, Jorge Sastre, Jesús Peinado, and Emilio Defez. Efficient and accurate algorithms for computing matrix trigonometric functions. Journal of Computational and Applied Mathematics, 309(??):325–332, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302321.

Asl:2017:IPS

[AJ17] Mohammad Shahbazi Asl and Mohammad Javidi. An improved PC scheme for nonlinear fractional differential equations: Error and stability analysis. Journal of Computational and Applied Mathematics, 324(??):101–117, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301966.

Asl:2018:NAE

[AJ18] Mohammad Shahbazi Asl and Mohammad Javidi. Novel algorithms to estimate nonlinear FDEs: Applied to fractional order nutrient–phytoplankton– zooplankton system. Journal of Computational and Applied Mathematics, 339 (??):193–207, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717305319.

Ajmi:2015:BMA

[AJR15] Neyla Ajmi, Abdellatif Jouini, and Pierre Gilles Lemarié Rieusset. Biorthogonal multiresolution analysis on a triangle and applications. Journal of Computational and Applied Mathematics, 288(??):233–243, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002277.

Asl:2018:NHO

[AJY18] Mohammad Shahbazi Asl, Mohammad Javidi, and Yubin Yan. A novel highorder algorithm for the numerical estimation of fractional differential equations. Journal of Computational and Applied Mathematics, 342(??):180–201, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300153.

Anastassi:2015:OER

[AK15] Z. A. Anastassi and A. A.

Kosti. A 6(4) optimized embedded Runge–Kutta– Nyström pair for the numerical solution of periodic problems. Journal of Computational and Applied Mathematics, 275(??):311–320, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003355.

Asai:2016:MSM

[AK16] Y. Asai and P. E. Kloeden. Multi-step methods for random ODEs driven by Itô diffusions. Journal of Computational and Applied Mathematics, 294 (??):210–224, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042715004276.

Anmarkrud:2017:OCS

[AK17a] Sverre Anmarkrud and Anne Kværnø. Order conditions for stochastic Runge–Kutta methods preserving quadratic invariants of Stratonovich SDEs. Journal of Computational and Applied Mathematics, 316(??):40– 46, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304113.

Aursand:2017:LDG

[AK17b] P. Aursand and U. Koley. Local discontinuous Galerkin schemes for a nonlinear variational wave equation modelling liquid crystals. Journal of Computational and Applied Mathematics, 317(??):478– 499, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305994.

Axelsson:2017:DEE

[AK17c] Owe Axelsson and János Karátson. Discretization error estimates in maximum norm for convergent splittings of matrices with a monotone preconditioning part. Journal of Computational and Applied Mathematics, 310(??):155–164, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301492.

Aksoylu:2018:CEA

[AK18] Burak Aksoylu and Adem Kaya. Conditioning and error analysis of nonlocal operators with local boundary conditions. Journal of Computational and Applied Mathematics, 335(??):1– 19, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305836.

Alzahrani:2019:FSE

[AK19] S. S. Alzahrani and A. Q. M. Khaliq. Fourier spectral exponential time differencing methods for multi-dimensional spacefractional reaction–diffusion equations. Journal of Computational and Applied Mathematics, 361 (??):157–175, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL https://www.sciencedirect.
com/science/article/
pii/S0377042719301815.

Akhtari:2019:NTS

[Akh19] Bahar Akhtari. Numerical treatment of stochastic delay differential equations: a global error bound. Journal of Computational and Applied Mathematics, 361(??):249–270, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301979.

Al-Khaled:2018:EPF

[AKHSU18] Kamel Al-Khaled, Nicholas Haynes, William Schiesser, and Muhammad Usman. Eventual periodicity of the forced oscillations for a Korteweg–de Vries type equation on a bounded domain using a sinc collocation method. Journal of Computational and Applied Mathematics, 330 (??):417–428, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.

com/science/article/
pii/S0377042717304120.

Axelsson:2018:SCU

[AKM18] Owe Axelsson, János Karátson, and Fréderic Magoulès. Superlinear convergence using block preconditioners for the real system formulation of complex Helmholtz equations. Journal of Computational and Applied Mathematics, 340(??):424–431, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830058X.

Agnelli:2018:NAE

[AKR18] J. P. Agnelli, U. Kaufmann, and J. D. Rossi. Numerical approximation of equations involving minimal/maximal operators by successive solution of obstacle problems. Journal of Computational and Applied Mathematics, 342(??):133–146, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301845.

Aksoy:2017:VMS

[Aks17] N. Yildirim Aksoy. Variational method for the solution of an inverse problem. Journal of Computational and Applied Mathematics, 312(??):82–93, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000066.

Alvarez:2016:AFV

[AL16] Edgardo Alvarez and Carlos Lizama. Attractivity for functional Volterra integral equations of convolution type. Journal of Computational and Applied Mathematics, 301(??):230–240, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300280. See

comment [FL23].

Ahrens:2018:FBE

[AL18] Robin Ahrens and Sabine Le Borne. FFT-based evaluation of multivariate aggregation integrals in population balance equations on uniform tensor grids. Journal of Computational and Applied Mathematics, 338(??):280–297, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300840.

Ahmed:2019:HOI

[AL19a] Sameed Ahmed and Xinfeng Liu. High order integration factor methods for systems with inhomogeneous boundary conditions. Journal of Computational and Applied Mathematics, 348(??):89–102, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305235.

Antoine:2019:RCS

[AL19b] X. Antoine and E. Lorin.

On the rate of convergence of Schwarz waveform relaxation methods for the time-dependent Schrödinger equation. Journal of Computational and Applied Mathematics, 354 (??):15–30, July 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718307313.

Axelsson:2019:NPM

[AL19c] Owe Axelsson and Zhao-Zheng Liang. A note on preconditioning methods for time-periodic eddy current optimal control problems. Journal of Computational and Applied Mathematics, 352(??):262–277, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306757.

Akhavan:2019:SOT

[ALC19] Yousef Akhavan, Dong Liang, and Michael Chen. Second order in time and space corrected explicitimplicit domain decomposition scheme for convection– diffusion equations. Journal of Computational and Applied Mathematics, 357(??):38–55, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300901.

Angulo:2016:SEN

[ALMLM16] O. Angulo, J. C. López-Marcos, and M. A. López-Marcos. Study on the efficiency in the numerical integration of size-structured population models: Error and computational cost. Journal of Computational and Applied Mathematics, 291(??):391–401, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001752.

Angulo:2017:SON

[ALMLM17] O. Angulo, J. C. López-Marcos, and M. A. López-Marcos. A second-order numerical method for a cell population model with asymmetric division. Journal of Computational and Applied Mathematics, 309(??):522–531, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301212.

A:2018:OEL

[ALS18] Chunxiang A, Yongzeng Lai, and Yi Shao. Optimal excess-of-loss reinsurance and investment problem with delay and jump-diffusion risk process under the CEV model. Journal of Computational and Applied Mathematics, 342(??):317–336, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302152.

Agranovich:2019:DBI

[ALS19] G. Agranovich, E. Litsyn, and A. Slavova. Dynamical behavior of integrodifferential boundary value problem arising in nanostructures via Cellular Nanoscale Network approach. Journal of Computational and Applied Mathematics, 352(??):62– 71, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307143.

Alcazar:2019:SSP

[ALV19] Juan Gerardo Alcázar, Miroslav Lávicka, and Jan Vrsek. Symmetries and similarities of planar algebraic curves using harmonic polynomials. Journal of Computational and Applied Mathematics, 357(??):302–318, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301189.

Argyros:2015:ECR

[AM15a] Ioannis K. Argyros and Á. Alberto Magreñán. Extended convergence results for the Newton– Kantorovich iteration. Journal of Computational and Applied Mathematics, 286(??):54–67, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500148X.

Argyros:2015:ECD

[AM15b] Ioannis K. Argyros and Á. Alberto Magreñán. Extending the convergence domain of the secant and Moser method in Banach space. Journal of Computational and Applied Mathematics, 290 (??):114–124, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002782.

Al-Muslimawi:2016:NSD

[AM16] Alaa H. Al-Muslimawi.

Numerical study for differential constitutive equations with polymer melts by using a hybrid finiteelement/volume method. Journal of Computational and Applied Mathematics, 308(??):488–498, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716302771.

Acosta:2017:MAC

[AM17a] Gabriel Acosta and Gabriel Monzón. The minimal angle condition for quadrilateral finite elements of arbitrary degree. Journal of Computational and Applied Mathematics, 317(??):218– 234, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305830.

Agahi:2017:PID

[AM17b] Hamzeh Agahi and Radko Mesiar. Probability inequalities for decomposition integrals. Journal of Computational and Applied Mathematics, 315 (??):240–248, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716305477.

Almeida:2019:ANA

[AM19] Ricardo Almeida and M. Luísa Morgado. Analysis and numerical approximation of tempered fractional calculus of variations problems. Journal of Computational and Applied Mathematics, 361(??):1–12, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301906.

Alshomrani:2019:CSL

[AMB19] Ali Saleh Alshomrani, P. Maroju, and Ramandeep Behl. Convergence of a Stirling-like method for fixed points in Banach spaces. Journal of Computational and Applied Mathematics, 354(??):183– 190, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305880.

Alonso-Mallo:2019:AOR

[AMCR19] I. Alonso-Mallo, B. Cano, and N. Reguera. Avoiding order reduction when integrating reaction–diffusion boundary value problems with exponential splitting methods. Journal of Computational and Applied Mathematics, 357 (??):228–250, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042719300962.

Alsaker:2019:DOP

[AMM19] Melody Alsaker, Jennifer L. Mueller, and Rashmi Murthy. Dynamic optimized priors for D-bar reconstructions of human ventilation using electrical impedance tomography. Journal of Computational and Applied Mathematics, 362(??):276–294, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304655.

Aurada:2015:FBC

[AMP15] M. Aurada, J. M. Melenk, and D. Praetorius. FEM–BEM coupling for the large-body limit in micromagnetics. Journal of Computational and Applied Mathematics, 281(??):10– 31, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005214.

Alonso-Mallo:2018:TES

[AMP18] I. Alonso-Mallo and A. M. Portillo. Time exponential splitting integrator for the Klein–Gordon equation with free parameters in the hagstrom-warburton absorbing boundary conditions. Journal of Computational and Applied Mathematics, 333(??):185–199, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305472.

Arcos:2017:DFE

[AMRM17] A. Arcos, S. Martínez, M. Rueda, and H. Martínez. Distribution function estimates from dual frame context. Journal of Computational and Applied Mathematics, 318(??):242– 252, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304472.

Akbas:2017:HOA

[AMRX17] Mine Akbas, Muhammad Mohebujjaman, Leo G. Rebholz, and Mengying Xiao. High order algebraic splitting for magnetohydrodynamics simulation. Journal of Computational and Applied Mathematics, 321(??):128–142, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300869.

Argyros:2017:IDP

[AMS17] Ioannis K. Argyros, Ángel Alberto Magreñán, and Juan Antonio Sicilia. Improving the domain of parameters for Newton’s method with applications. Journal of Computational and Applied Mathematics, 318(??):124– 135, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305520.

Aouadi:2019:TMP

[AMZ19] S. Mani Aouadi, W. Mbarki, and N. Zemzemi. Towards the modeling of the Purkinje/myocardium coupled problem: a well-posedness analysis. Journal of Computational and Applied Mathematics, 351(??):136– 152, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306319.

Alharbi:2017:AMM

[AN17a] Abdulghani Alharbi and Shailesh Naire. An adaptive moving mesh method for thin film flow equations with surface tension. Journal of Computational and Applied Mathematics, 319(??):365–384, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300407.

Avdzhieva:2017:AOD

[AN17b] Ana Avdzhieva and Geno Nikolov. Asymptotically optimal definite quadrature formulae of 4th order. Journal of Computational and Applied Mathematics, 311(??):565–582, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630382X.

Alharbi:2019:AMM

[AN19] Abdulghani Alharbi and Shailesh Naire. An adaptive moving mesh method for two-dimensional thin film flow equations with surface tension. Journal of Computational and Applied Mathematics, 356(??):219–230, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042719300810.

Andrei:2016:ACG

[And16] Neculai Andrei. An adaptive conjugate gradient algorithm for large-scale unconstrained optimization. Journal of Computational and Applied Mathematics, 292(??):83–91, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003520.

Andrei:2017:AAP

[And17] Neculai Andrei. Accelerated adaptive Perry conjugate gradient algorithms based on the self-scaling memoryless BFGS update. Journal of Computational and Applied Mathematics, 325(??):149–164, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302236.

Andrei:2018:DPS

[And18] Neculai Andrei. A double parameter scaled BFGS method for unconstrained optimization. Journal of Computational and Applied Mathematics, 332 (??):26–44, April 2018.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717304995.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 333(??):ii, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306246.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Editorial

board. Journal of Computational and Applied Mathematics, 334(??):ii, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306507.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Editorial

board. Journal of Computational and Applied Mathematics, 335(??):ii, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718300396.

Anonymous:2018:EBd

[Ano18d]  Anonymous.  Editorial

board. Journal of Computational and Applied Mathematics, 336(??):ii, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300633.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Editorial

board. Journal of Computational and Applied Mathematics, 337(??):ii, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300931.

Anonymous:2018:EBf

[Ano18f]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 338(??):ii, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301274.

Anonymous:2018:EBl

[Ano18g]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 339(??):ii, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301419.

Anonymous:2018:EBg

[Ano18h]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 340(??):ii, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301535.

Anonymous:2018:EBh

[Ano18i]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 341(??):ii, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302577.

Anonymous:2018:EBi

[Ano18j]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 342(??):ii, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303108.

Anonymous:2018:EBj

[Ano18k]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 343(??):ii, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303431.

Anonymous:2018:EBk

[Ano18l]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 344(??):ii, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304126.

Anonymous:2019:EBa

[Ano19a]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 345(??):ii, January 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718304564.

Anonymous:2019:EBb

[Ano19b]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 346(??):ii, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305053.

Anonymous:2019:EBc

[Ano19c]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 347(??):ii, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305727.

Anonymous:2019:EBd

[Ano19d]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 348(??):ii, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306125.

Anonymous:2019:EBe

[Ano19e]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 349(??):ii, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306861.

Anonymous:2019:EBf

[Ano19f]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 350(??):ii, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306927.

Anonymous:2019:EBg

[Ano19g]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 351(??):ii, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307428.

Anonymous:2019:EBh

[Ano19h]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 352(??):ii, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042719300056.

Anonymous:2019:EBi

[Ano19i]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 353(??):ii, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042719300329.

Anonymous:2019:EBj

[Ano19j]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 354(??):ii, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042719300767.

Anonymous:2019:EBk

[Ano19k]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 355(??):ii, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301098.

Anonymous:2019:EBl

[Ano19l]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 356(??):ii, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301372.

Anonymous:2019:EBm

[Ano19m]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 357(??):ii, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301682.

Anonymous:2019:EBn

[Ano19n]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 358(??):ii, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301943.

Anonymous:2019:EBo

[Ano19o]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 359(??):ii, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S037704271930216X.

Anonymous:2019:EBp

[Ano19p]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 360(??):ii, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302341.

Anonymous:2019:EBq

[Ano19q]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 361(??):ii, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302705.

Anonymous:2019:EBr

[Ano19r]  Anonymous.  Editorial

Board. Journal of Computational and Applied Mathematics, 362(??):ii, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719303012.

Alvarez-Nodarse:2015:LRS

[ÁNPPCSA15] R. Álvarez-Nodarse, J. Petronilho, N. C. Pinzón-Cortés, and R. Sevinik-Adigüzel. On linearly related sequences of difference derivatives of discrete orthogonal polynomials. Journal of Computational and Applied Mathematics, 284(??):26–37, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400291X.

Agarwal:2017:LTI

[AÖ17] Ravi P. Agarwal and Abdullah Özbekler. Lyapunov type inequalities for mixed nonlinear Riemann– Liouville fractional differential equations with a forcing term. Journal of Computational and Applied Mathematics, 314 (??):69–78, April 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0377042716304988.

Assa:2018:DSD

[AO18] Hirbod Assa and Ramin Okhrati. Designing sound deposit insurances. Journal of Computational and Applied Mathematics, 327(??):226–242, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302960.

Azarnavid:2018:IRK

[AP18] Babak Azarnavid and Kourosh Parand. An iterative reproducing kernel method in Hilbert space for the multi-point boundary value problems. Journal of Computational and Applied Mathematics, 328(??):151–163, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303564.

Alonso:2015:AST

[APS15a] Pedro Alonso, J. M. Peña, and María Luisa Serrano. Almost strictly totally negative matrices: an algorithmic characterization. Journal of Computational and Applied Mathematics, 275(??):238–246, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003471.

Alonso:2015:CAS

[APS15b] Pedro Alonso, J. M. Peña, and María Luisa Serrano. On the characterization of almost strictly sign regular matrices. Journal of Computational and Applied Mathematics, 275(??):480–488, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714000600.

Alonso:2017:DAS

[APS17a] P. Alonso, J. M. Peña, and M. L. Serrano. QR decomposition of almost strictly sign regular matrices. Journal of Computational and Applied Mathematics, 318(??):646– 657, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005403.

Alonso:2017:BSA

[APS17b] P. Alonso, J. M. Peña, and M. L. Serrano. Backward stability with almost strictly sign regular matrices. Journal of Computational and Applied Mathematics, 322(??):71–80, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301322.

Alonso:2019:CMA

[APS19a] P. Alonso, J. M. Peña, and M. L. Serrano. Combined matrices of almost strictly sign regular matrices. Journal of Computational and Applied Mathematics, 354(??):144– 151, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305806.

Alonso:2019:CPS

[APS19b] P. Alonso, J. M. Peña, and M. L. Serrano. Comparing pivoting strategies for almost strictly sign regular matrices. Journal of Computational and Applied Mathematics, 354(??):96– 102, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718300864.

Arnold:2017:P

[APW^{+}17] Martin Arnold, Helmut Podhaisky, Rüdiger Weiner, Jason Frank, and Willem Hundsdorfer. Preface. Journal of Computational and Applied Mathematics, 316(??):1–2, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305404.

Ali:2019:PBS

[APYW19] M. Syed Ali, L. Palanisamy, J. Yogambigai, and Linshan Wang. Passivitybased synchronization of Markovian jump complex dynamical networks with time-varying delays, parameter uncertainties, reaction–diffusion terms, and sampled-data control. Journal of Computational and Applied Mathematics, 352(??):79–92, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306551.

Agarwal:2017:CII

[AQCJ17] Praveen Agarwal, Feng Qi, Mehar Chand, and Shilpi Jain. Certain integrals involving the generalized hypergeometric function and the Laguerre polynomials. Journal of Computational and Applied Mathematics, 313(??):307–317, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630454X.

Amini:2015:MTS

[AR15a] Keyvan Amini and Faramarz Rostami. A modified two steps Levenberg– Marquardt method for nonlinear equations. Journal of Computational and Applied Mathematics, 288(??):341–350, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002666.

Andrade:2015:EPD

[AR15b] J. A. A. Andrade and P. N. Rathie. On exact posterior distributions using H-functions. Journal of Computational and Applied Mathematics, 290(??):459–475, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003131.

Argyros:2015:CEK

[AR15c] Ioannis K. Argyros and Hongmin Ren. On the convergence of efficient King–Werner-type methods of order 1 + \sqrt{2}. Journal of Computational and Applied Mathematics, 285(??):169–180, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000916.

Amini:2016:TSM

[AR16] Keyvan Amini and Faramarz Rostami. Threesteps modified Levenberg– Marquardt method with a new line search for systems of nonlinear equations. Journal of Computational and Applied Mathematics, 300(??):30– 42, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715006251.

Asl:2019:IUT

[AR19] Nazdar Abdollahi Asl and Davood Rostamy. Identifying an unknown time-dependent boundary source in time-fractional diffusion equation with a non-local boundary condition. Journal of Computational and Applied Mathematics, 355(??):36–50, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300275.

Arandiga:2019:ARI

[Arà19] Francesc Aràndiga. Adaptive rational interpolation for point values. Journal of Computational and Applied Mathematics, 349(??):212–224, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305399.

Acal:2019:PTD

[ARCA^{+}19] C. Acal, J. E. Ruiz-Castro, A. M. Aguilera, F. Jiménez-Molinos, and J. B. Roldán. Phase-type distributions for studying variability in resistive memories. Journal of Computational and Applied Mathematics, 345: 23–32, January 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Armentano:2018:SLO

[Arm18] María Gabriela Armentano. Stabilization of low-order cross-grid P_{k}Q_{l} mixed finite elements. Journal of Computational and Applied Mathematics, 330(??):340–355, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717304193.

Alvarez:2016:CRP

[ÁRS16] Francisco Álvarez, José-Manuel Rey, and Raúl G. Sanchis. Consumer’s response to price distribution and σ-overload under time allocation. Journal of Computational and Applied Mathematics, 291(??):242–256, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001545.

Amat:2019:SFF

[ARTY19] Sergio Amat, Juan Ruiz, J. Carlos Trillo, and Dionisio F. Yáñez. On a stable family of four-point nonlinear subdivision schemes eliminating the Gibbs phenomenon. Journal of Computational and Applied Mathematics, 354(??):310– 325, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830253X.

Ahmadinia:2018:NSS

[AS18] M. Ahmadinia and Z. Safari. Numerical solution of singularly perturbed boundary value problems by improved least squares method. Journal of Computational and Applied Mathematics, 331(??):156–165, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304545.

Anderson:2019:ITV

[AS19a] Rachele Anderson and Maria Sandsten. Inference for time-varying signals using locally stationary processes. Journal of Computational and Applied Mathematics, 347(??):24–35, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304722.

Araujo:2019:UQE

[AS19b] António Araújo and Pedro Serranho. On the use of quasi-equidistant source points over the sphere surface for the method of fundamental solutions. Journal of Computational and Applied Mathematics, 359(??):55–68, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S037704271930144X.

Abreu:2018:ACS

[ASKG18] Rafael Abreu, Zeming Su, Jochen Kamm, and Jinghuai Gao. On the accuracy of the complexstep-finite-difference method. Journal of Computational and Applied Mathematics, 340(??):390–403, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301316.

Allouch:2019:LSG

[AST19] C. Allouch, D. Sbibih, and M. Tahrichi. Legendre superconvergent Galerkincollocation type methods for Hammerstein equations. Journal of Computational and Applied Mathematics, 353(??):253– 264, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307714.

Arregui:2017:NST

[ASV17] Iñigo Arregui, Beatriz Salvador, and Carlos Vázquez. A numerical strategy for telecommunications networks capacity planning under demand and price uncertainty. Journal of Computational and Applied Mathematics, 318(??):491– 503, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630036X.

Allouche:2017:NSS

[AT17] Hassane Allouche and Abderrahim Tazdayte. Numerical solution of singular boundary value problems with logarithmic singularities by Padé approximation and collocation methods. Journal of Computational and Applied Mathematics, 311(??):324–341, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303636.

Amini:2018:BCM

[ATG18] S. Amini, F. Toutounian, and M. Gachpazan. The block CMRH method for solving nonsymmetric linear systems with multiple right-hand sides. Journal of Computational and Applied Mathematics, 337(??):166–174, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300335.

Aptekarev:2015:HUA

[ATV15] A. I. Aptekarev, D. N. Toulyakov, and W. Van Assche. Hyperelliptic uniformization of algebraic curves of the third order. Journal of Computational and Applied Mathematics, 284(??):38–49, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005275.

Anatriello:2016:LSC

[AV16] Giuseppina Anatriello and Giovanni Vincenzi. Logarithmic spirals and continue triangles. Journal of Computational and Applied Mathematics, 296 (??):127–137, April 2016.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715004562.

Abbasbandy:2017:RKM

[AVB17] S. Abbasbandy, R. A. Van Gorder, and P. Bakhtiari. Reproducing kernel method for the numerical solution of the Brinkman– Forchheimer momentum equation. Journal of Computational and Applied Mathematics, 311(??):262– 271, February 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716303491.

Astudillo:2016:RID

[AvG16] R. Astudillo and M. B.

van Gijzen. A restarted induced dimension reduction method to approximate eigenpairs of large unsymmetric matrices. Journal of Computational and Applied Mathematics, 296(??):24– 35, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004665.

Alvarez-Vazquez:2015:SSW

[AVGCMVM15] L. J. Alvarez-Vázquez, N. García-Chan, A. Martínez, and M. E. Vázquez-Méndez. Stackelberg strategies for wastewater management. Journal of Computational and Applied Mathematics, 280(??):217–230, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271400541X.

Alvarez-Vazquez:2017:NSA

[AVGCMVM17] L. J. Alvarez-Vázquez, N. García-Chan, A. Martínez, and M. E. Vázquez-Méndez. Numerical simulation of air pollution due to traffic flow in urban networks. Journal of Computational and Applied Mathematics, 326(??):44–61, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302534.

Amanbek:2019:PEA

[AW19a] Yerlan Amanbek and Mary F. Wheeler. A priori error analysis for transient problems using enhanced velocity approach in the discretetime setting. Journal of Computational and Applied Mathematics, 361(??):459–471, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302432.

Angermann:2019:SCU

[AW19b] Lutz Angermann and Song Wang. A superconvergent unsymmetric finite volume method for convection–diffusion equations. Journal of Computational and Applied Mathematics, 358(??):179–189, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301426.

Agahi:2018:GHD

[AY18] Hamzeh Agahi and Milad Yadollahzadeh. A generalization of HH fdivergence. Journal of Computational and Applied Mathematics, 343 (??):309–317, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042718302632.

Ali:2019:EDS

[AY19] M. Syed Ali and J. Yogambigai. Extended dissipative synchronization of complex dynamical networks with additive time-varying delay and discrete-time information. Journal of Computational and Applied Mathematics, 348(??):328–341, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303558.

Ali:2019:SSN

[AYSE19] M. Syed Ali, J. Yogambigai, S. Saravanan, and S. Elakkia. Stochastic stability of neutraltype Markovian-jumping BAM neural networks with time varying delays. Journal of Computational and Applied Mathematics, 349(??):142–156, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305867.

An:2017:EAF

[AZ17] Rong An and Can Zhou. Error analysis of a fractionalstep method for magnetohydrodynamics equations. Journal of Computational and Applied Mathematics, 313(??):168–184, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304253.

Alinia:2018:NTB

[AZ18] N. Alinia and M. Zarebnia. A new tension Bspline method for thirdorder self-adjoint singularly perturbed boundary value problems. Journal of Computational and Applied Mathematics, 342(??):521–533, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301675.

Bayramoglu:2018:NOD

[(Ba18] Ismihan Bayramoglu (Bairamov).

A note on the ordering of distribution functions of INID random variables. Journal of Computational and Applied Mathematics, 343(??):49–54, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302231.

Badr:2015:NTI

[BAAA^{+}15] Sayed Abdel-Naeim Badr, Nassar H. Abdel-All, Osmar Aléssio, Mustafa Düldül, and Bahar Uyar Düldül. Non-transversal intersection curves of hypersurfaces in Euclidean 4space. Journal of Computational and Applied Mathematics, 288(??):81–98, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002071.

Baccouch:2018:SLD

[Bac18] Mahboub Baccouch. Superconvergence of the local discontinuous Galerkin method for the sine-Gordon equation in one space dimension. Journal of Computational and Applied Mathematics, 333 (??):292–313, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717305678.

Bader:2018:FSM

[Bad18] Philipp Bader. Fouriersplitting methods for the dynamics of rotating Bose–Einstein condensates. Journal of Computational and Applied Mathematics, 336(??):267– 280, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300062.

Barton:2017:GQR

[BAHC17] Michael Barton, Rachid Ait-Haddou, and Victor Manuel Calo. Gaussian quadrature rules for C^{1} quintic splines with uniform knot vectors. Journal of Computational and Applied Mathematics, 322(??):57–70, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300870.

Bai:2015:MRK

[Bai15] Zhong-Zhi Bai. Motivations and realizations of Krylov subspace methods for large sparse linear systems. Journal of Computational and Applied Mathematics, 283(??):71–78, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000370.

Boutarfa:2017:NAF

[BAI17] Bariza Boutarfa, Ali Akgül, and Mustafa Inc. New approach for the Fornberg– Whitham type equations. Journal of Computational and Applied Mathematics, 312(??):13–26, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715004689.

Beltman:2018:CPM

[BAK18] R. Beltman, M. J. H. Anthonissen, and B. Koren. Conservative polytopal mimetic discretization of the incompressible Navier–Stokes equations. Journal of Computational and Applied Mathematics, 340(??):443–473, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300785.

Balasubramani:2017:SPR

[Bal17] N. Balasubramani. Shape preserving rational cubic fractal interpolation function. Journal of Computational and Applied Mathematics, 319(??):277–295, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730033X.

Behl:2019:EOF

[BAMM19] Ramandeep Behl, S. Amat, Á.A. Magreñán, and S. S. Motsa. An efficient optimal family of sixteenth order methods for nonlinear models. Journal of Computational and Applied Mathematics, 354(??):271– 285, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300827.

Bayon:2016:CCD

[BAO^{+}16] L. Bayón, P. Fortuny Ayuso, J. A. Otero, P. M. Suárez, and C. Tasis. Cyclic coordinate descent in a class of bangsingular-bang problems. Journal of Computational and Applied Mathematics, 291(??):304–316, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001491.

Baysal:2015:GCN

[Bay15] Onur Baysal. A globally convergent numerical method for a coefficient inverse problem for a parabolic equation. Journal of Computational and Applied Mathematics, 289(??):153–172, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000990.

Bazm:2015:BPN

[Baz15] S. Bazm. Bernoulli polynomials for the numerical solution of some classes of linear and nonlinear integral equations. Journal of Computational and Applied Mathematics, 275(??):44–60, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003379.

Bourchtein:2016:PPE

[BB16] Ludmila Bourchtein and Andrei Bourchtein. On perturbations of principal eigenvectors of substochastic matrices. Journal of Computational and Applied Mathematics, 295(??):149–158, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042715000254.

Bazhlekova:2018:SAM

[BB18] Emilia Bazhlekova and Ivan Bazhlekov. Subordination approach to multi-term time-fractional diffusion-wave equations. Journal of Computational and Applied Mathematics, 339(??):179–192, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042717305630.

Balbas:2017:DEC

[BBB17a] Alejandro Balbás, Beatriz Balbás, and Raquel Balbás. Differential equations connecting VaR and CVaR. Journal of Computational and Applied Mathematics, 326 (??):247–267, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730290X.

Balbas:2017:VCS

[BBB17b] Alejandro Balbás, Beatriz Balbás, and Raquel Balbás. VaR as the CVaR sensitivity: Applications in risk optimization. Journal of Computational and Applied Mathematics, 309(??):175–185, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303089.

Baskonus:2017:ASN

[BBB17c] Haci Mehmet Baskonus, Hasan Bulut, and Fethi Bin Muhammad Belgacem. Analytical solutions for nonlinear longshort wave interaction systems with highly complex structure. Journal of Computational and Applied Mathematics, 312(??):257– 266, March 1, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716302692.

Beik:2017:BDB

[BBC17] Fatemeh Panjeh Ali Beik, Michele Benzi, and Sayyed-Hasan Azizi Chaparpordi. On block diagonal and block triangular iterative schemes and preconditioners for stabilized saddle point problems. Journal of Computational and Applied Mathematics, 326(??):15–30, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302455.

Bader:2018:SIS

[BBC^{+}18] Philipp Bader, Sergio Blanes, Fernando Casas, Nikita Kopylov, and Enrique Ponsoda. Symplectic integrators for second-order linear nonautonomous equations. Journal of Computational and Applied Mathematics, 330(??):909–919, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042717301516.

Bader:2019:NSI

[BBCK19] Philipp Bader, Sergio Blanes, Fernando Casas, and Nikita Kopylov. Novel symplectic integrators for the Klein–Gordon equation with space- and timedependent mass. Journal of Computational and Applied Mathematics, 350 (??):130–138, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718306186.

Bader:2016:ENI

[BBCP16] Philipp Bader, Sergio Blanes, Fernando Casas, and Enrique Ponsoda. Efficient numerical integration of N-th-order nonautonomous linear differential equations. Journal of Computational and Applied Mathematics, 291(??):380–390, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001399.

Barrios:2016:ANS

[BBD16] Tomás P. Barrios, Rommel Bustinza, and Víctor Domínguez. Adaptive numerical solution of a discontinuous Galerkin method for a Helmholtz problem in low-frequency regime. Journal of Computational and Applied Mathematics, 300(??):312– 340, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006366.

Bezrodnykh:2019:CCS

[BBG^{+}19a] Sergei Bezrodnykh, Andrei Bogatyrëv, Sergei Goreinov, Oleg Grigor’ev, Harri Hakula, and Matti Vuorinen. On capacity computation for symmetric polygonal condensers. Journal of Computational and Applied Mathematics, 361(??):271–282, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301554.

Boey:2019:ARB

[BBG19b] Edward Boey, Yves Bourgault, and Thierry Giordano. Anisotropic residual based a posteriori mesh adaptation in 2D: Element based approach. Journal of Computational and Applied Mathematics, 349(??):1–20, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305405.

Barrios:2019:PEA

[BBGG19] Tomás P. Barrios, Rommel Bustinza, Galina C. García, and María González. An a posteriori error analysis of a velocitypseudostress formulation of the generalized Stokes problem. Journal of Computational and Applied Mathematics, 357 (??):349–365, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300925.

Bai:2017:MBI

[BBH^{+}17] Zhong-Zhi Bai, Alessandro Buccini, Ken Hayami, Lothar Reichel, Jun-Feng Yin, and Ning Zheng. Modulus-based iterative methods for constrained Tikhonov regularization. Journal of Computational and Applied Mathematics, 319(??):1–13, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306331.

Birgin:2015:ARG

[BBM15] E. G. Birgin, L. F. Bueno, and J. M. Martínez. Assessing the reliability of general-purpose inexact restoration methods. Journal of Computational and Applied Mathematics, 282(??):1– 16, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005871.

Bringedal:2015:MNI

[BBPR15] Carina Bringedal, Inga Berre, Iuliu Sorin Pop, and Florin Adrian Radu. A model for non-isothermal flow and mineral precipitation and dissolution in a thin strip. Journal of Computational and Applied Mathematics, 289(??):346–355, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005585.

Bader:2017:SIM

[BBPS17] Philipp Bader, Sergio Blanes, Enrique Ponsoda, and Muaz Seydaoğlu. Symplectic integrators for the matrix Hill equation. Journal of Computational and Applied Mathematics, 316(??):47– 59, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630468X.

Bedin:2018:MRB

[BBQ18] L. Bedin, F. S. V.

Bazán, and J. R. Quiroz. Method for recovering boundary data in a twodimensional Poisson equation on annular domain. Journal of Computational and Applied Mathematics, 342(??):83–95, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301626.

Barton:2016:GQS

[BC16] Michael Bartoň and Victor Manuel Calo. Gaussian quadrature for splines via homotopy continuation: rules for C^{2} cubic splines. Journal of Computational and Applied Mathematics, 296(??):709–723, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004896.

Berriochoa:2018:GWO

[BCA18] E. Berriochoa, A. Cachafeiro, and J. M. García Amor. Gibbs–Wilbraham oscillation related to an Hermite interpolation problem on the unit circle. Journal of Computational and Applied Mathematics, 344(??):657–675, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303571.

Benitez:2018:CCE

[BCC^{+}18] J. Benítez, S. Carpitella, A. Certa, A. E. Ilaya-Ayza, and J. Izquierdo. Consistent clustering of entries in large pairwise comparison matrices. Journal of Computational and Applied Mathematics, 343(??):98–112, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830236X.

Bultheel:2015:MMQ

[BCCB15] Adhemar Bultheel, María José Cantero, and Ruymán Cruz-Barroso. Matrix methods for quadrature formulas on the unit circle. A survey. Journal of Computational and Applied Mathematics, 284(??):78–100, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004725.

Berriochoa:2015:CHI

[BCD15] E. Berriochoa, A. Cachafeiro, and J. Díaz. Convergence of Hermite interpolants on the circle using two derivatives. Journal of Computational and Applied Mathematics, 284(??):58–68, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004385.

Bracco:2017:APG

[BCDwK17] Cesare Bracco, Durkbin Cho, Catterina Dagnino, and Tae wan Kim. On the approximation power of generalized T -splines. Journal of Computational and Applied Mathematics, 311(??):423–438, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303302.

Barrios:2016:SSN

[BCFN16] J. G. Barrios, J. Y. Bello Cruz, O. P. Ferreira, and S. Z. Németh. A semismooth Newton method for a special piecewise linear system with application to positively constrained convex quadratic programming. Journal of Computational and Applied Mathematics, 301(??):91–100, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300206.

Braumann:2018:RGF

[BCJV18] C. A. Braumann, J.-C.

Cortés, L. Jódar, and L. Villafuerte. On the random gamma function: Theory and computing. Journal of Computational and Applied Mathematics, 335(??):142– 155, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306064.

Borri:2016:BTS

[BCMP16] A. Borri, F. Carravetta, G. Mavelli, and P. Palumbo. Block-tridiagonal statespace realization of chemical master equations: a tool to compute explicit solutions. Journal of Computational and Applied Mathematics, 296(??):410– 426, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005075.

Buonocore:2015:ROU

[BCNP15] A. Buonocore, L. Caputo, A. G. Nobile, and E. Pirozzi. Restricted Ornstein–Uhlenbeck process and applications in neuronal models with periodic input signals.

Journal of Computational and Applied Mathematics, 285(??):59–71, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000643.

Bessemoulin-Chatard:2017:PMC

[BCS17] Marianne Bessemoulin-Chatard and Mazen Saad. Preserving monotony of combined edge finite volumefinite element scheme for a bone healing model on general mesh. Journal of Computational and Applied Mathematics, 309(??):287–311, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303168.

Babajee:2016:SIM

[BCT16] D. K. R. Babajee, A. Cordero, and J. R. Torregrosa. Study of iterative methods through the Cayley Quadratic Test. Journal of Computational and Applied Mathematics, 291(??):358–369, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042714004270.

Burgos:2018:SRM

[BCVV18] C. Burgos, J.-C. Cortés, L. Villafuerte, and R. J. Villanueva. Solving random mean square fractional linear differential equations by generalized power series: Analysis and computing. Journal of Computational and Applied Mathematics, 339(??):94–110, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300104.

Balu:2016:NMR

[BD16] R. Balu and T. K. DeLillo.

Numerical methods for Riemann–Hilbert problems in multiply connected circle domains. Journal of Computational and Applied Mathematics, 307(??):248–261, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300450.

Baratnia:2019:OWC

[BD19] Mohammad Baratnia and Mahdi Doostparast. Oneway classification with random effects: a reversedhazard-based approach. Journal of Computational and Applied Mathematics, 349(??):60–69, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305752.

Baleanu:2017:NSW

[BDA17] Dumitru Baleanu, Rahmat Darzi, and Bahram Agheli. New study of weakly singular kernel fractional fourth-order partial integro-differential equations based on the optimum q-homotopic analysis method. Journal of Computational and Applied Mathematics, 320: 193–201, August 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Bygi:2015:WVC

[BDAG15] Mojtaba Nouri Bygi, Shervin Daneshpajouh, Sharareh Alipour, and Mohammad Ghodsi. Weak visibility counting in simple polygons. Journal of Computational and Applied Mathematics, 288(??):215–222, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002368.

Brauchart:2015:SLD

[BDF15] Johann S. Brauchart, Josef Dick, and Lou Fang. Spatial low-discrepancy sequences, spherical cone discrepancy, and applications in financial modeling. Journal of Computational and Applied Mathematics, 286(??):28–53, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500093X.

Berdysheva:2019:MAS

[BDFM19] Elena E. Berdysheva, Nira Dyn, Elza Farkhi, and Alona Mokhov. Metric approximation of set-valued functions of bounded variation. Journal of Computational and Applied Mathematics, 349(??):251–264, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305909.

Bevilacqua:2015:CUR

[BDG15] Roberto Bevilacqua, Gianna M. Del Corso, and Luca Gemignani. Compression of unitary rankstructured matrices to CMV-like shape with an application to polynomial rootfinding. Journal of Computational and Applied Mathematics, 278(??):326–335, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004300.

Benitez:2015:CCI

[BDGIPG15] J. Benítez, X. Delgado-Galván, J. Izquierdo, and R. Pérez-García. Consistent completion of incomplete judgments in decision making using AHP. Journal of Computational and Applied Mathematics, 290(??):412–422, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003179.

Budakci:2015:EFF

[BDGO15] Gülter Budakçi, Çetin Disibüyük, Ron Goldman, and Halil Oruç. Extending fundamental formulas from classical Bsplines to quantum Bsplines. Journal of Computational and Applied Mathematics, 282(??):17– 33, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005901.

Barrera:2019:PDQ

[BDIR19a] D. Barrera, C. Dagnino, M. J. Ibáñez, and S. Remogna. Point and differential C^{1} quasiinterpolation on three direction meshes. Journal of Computational and Applied Mathematics, 354 (??):373–389, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718305016.

Barrera:2019:QIC

[BDIR19b] D. Barrera, C. Dagnino, M. J. Ibáñez, and S. Remogna. Quasi-interpolation by C^{1} quartic splines on type-1 triangulations. Journal of Computational and Applied Mathematics, 349(??):225–238, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304825.

Bhrawy:2016:MJB

[BDSB16] A. H. Bhrawy, E. H.

Doha, M. A. Saker, and D. Baleanu. Modified Jacobi–Bernstein basis transformation and its application to multi-degree reduction of Bézier curves. Journal of Computational and Applied Mathematics, 302(??):369–384, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000133.

Bonin:2016:CMI

[BdSFL16] C. R. B. Bonin, R. W. dos Santos, G. C. Fernandes, and M. Lobosco. Computational modeling of the immune response to yellow fever. Journal of Computational and Applied Mathematics, 295(??):127–138, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000321.

Bayramoglu:2015:OSD

[BE15] Ismihan Bayramoğlu (Bairamov) and Serkan Eryilmaz. Order statistics of dependent sequences consisting of two different sets of exchangeable variables. Journal of Computational and Applied Mathematics, 286(??):1–6, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001223.

Babazadeh:2019:NMP

[BE19] Hossein Babazadeh and Akbar Esfahanipour. A novel multi period mean-VaR portfolio optimization model considering practical constraints and transaction cost. Journal of Computational and Applied Mathematics, 361(??):313–342, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718306472.

Behiry:2016:RNS

[Beh16] S. H. Behiry. Retraction notice to “Solution of nonlinear Fredholm integrodifferential equations using a hybrid of block pulse functions and normalized Bernstein polynomials” [J. Comput. Appl. Math. 260 (2014) 258– 265]. Journal of Computational and Applied Mathematics, 294(??):446, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715004112.

Bizzozero:2017:RET

[BEHL17] D. A. Bizzozero, J. A. Ellison, K. Heinemann, and S. R. Lau. Rapid evaluation of two-dimensional retarded time integrals. Journal of Computational and Applied Mathematics, 324(??):118–141, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301620.

Bejancu:2016:RST

[Bej16] Aurelian Bejancu. Radially symmetric thin plate splines interpolating a circular contour map. Journal of Computational and Applied Mathematics, 292(??):7–22, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003416.

Bentbib:2016:GLM

[BEJR16] A. H. Bentbib, M. El Guide, K. Jbilou, and L. Reichel. A global Lanczos method for image restoration. Journal of Computational and Applied Mathematics, 300 (??):233–244, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715006469.

Bentbib:2017:GGK

[BEJR17] A. H. Bentbib, M. El Guide, K. Jbilou, and L. Reichel. Global Golub– Kahan bidiagonalization applied to large discrete ill-posed problems. Journal of Computational and Applied Mathematics, 322(??):46–56, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301358.

Blouza:2016:PAP

[BEMA16] Adel Blouza, Linda El Alaoui, and Saloua Mani-Aouadi. A posteriori analysis of penalized and mixed formulations of Koiter’s shell model. Journal of Computational and Applied Mathematics, 296(??):138–155, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003696.

Bernardo:2015:PCF

[Ber15] Fernando P. Bernardo. Performance of cubature formulae in probabilistic model analysis and optimization. Journal of Computational and Applied Mathematics, 280(??):110–124, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005329.

Bertolazzi:2018:GHI

[BF18] Enrico Bertolazzi and Marco Frego. On the G^{2} Hermite interpolation problem with clothoids. Journal of Computational and Applied Mathematics, 341(??):99–116, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301924.

Bazan:2015:ULS

[BFLP15] Fermín S. V. Bazán, Juliano B. Francisco, Koung Hee Leem, and George Pelekanos. Using the linear sampling method and an improved maximum product criterion for the solution of the electromagnetic inverse medium problem. Journal of Computational and Applied Mathematics, 273(??):61–75, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002738.

Bergamaschi:2019:SPE

[BFMP19] L. Bergamaschi, E. Facca, Á. Martínez, and M. Putti. Spectral preconditioners for the efficient numerical solution of a continuous branched transport model. Journal of Computational and Applied Mathematics, 354(??):259– 270, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300517.

Brambilla:2015:ADA

[BG15] P. Brambilla and A. Guardone.

Assessment of dynamic adaptive grids in volumeof-fluid simulations of oblique drop impacts onto liquid films. Journal of Computational and Applied Mathematics, 281 (??):277–283, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S037704271400586X.

Berenguer:2017:CMS

[BG17] M. I. Berenguer and D. Gámez. A computational method for solving a class of two dimensional Volterra integral equations. Journal of Computational and Applied Mathematics, 318(??):403– 410, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630262X.

Brugnano:2018:AEQ

[BGI18] Luigi Brugnano, Gianmarco Gurioli, and Felice Iavernaro. Analysis of Energy and QUadratic Invariant Preserving (EQUIP) methods. Journal of Computational and Applied Mathematics, 335(??):51– 73, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717306040.

Barrera:2015:GSD

[BGII15a] D. Barrera, P. González, F. Ibáñez, and M. J. Ibáñez. A general spline differential quadrature method based on quasiinterpolation. Journal of Computational and Applied Mathematics, 275(??):465–479, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400082X.

Barrera:2015:SBD

[BGII15b] D. Barrera, P. González, F. Ibáñez, and M. J. Ibáñez. On splinebased differential quadrature. Journal of Computational and Applied Mathematics, 275(??):272–280, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003409.

Berenguer:2017:SSI

[BGL17] M. I. Berenguer, D. Gámez, and A. J. López Linares. Solution of systems of integro-differential equations using numerical treatment of fixed point. Journal of Computational and Applied Mathematics, 315(??):343–353, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S037704271630543X.

Behl:2018:OEG

[BGMM18] Ramandeep Behl, D. González, Prashanth Maroju, and S. S. Motsa. An optimal and efficient general eighth-order derivative free scheme for simple roots. Journal of Computational and Applied Mathematics, 330(??):666–675, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303874.

Bandaliyev:2016:OCP

[BGMS16] R. A. Bandaliyev, V. S. Guliyev, I. G. Mamedov, and A. B. Sadigov. The optimal control problem in the processes described by the Goursat problem for a hyperbolic equation in variable exponent Sobolev spaces with dominating mixed derivatives. Journal of Computational and Applied Mathematics, 305(??):11–17, October 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301510.

Bayon:2017:AQL

[BGNGR^{+}17] L. Bayón, P. J. García-Nieto, R. García-Rubio, J. A. Otero, P. M. Suárez, and C. Tasis. An algorithm for quasi-linear control problems in the economics of renewable resources: the steady state and end state for the infinite and long-term horizon. Journal of Computational and Applied Mathematics, 309(??):456–472, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301169.

Brugnano:2019:ECH

[BGS19] Luigi Brugnano, Gianmarco Gurioli, and Yajuan Sun. Energy-conserving Hamiltonian boundary value methods for the numerical solution of the Korteweg–de Vries equation. Journal of Computational and Applied Mathematics, 351(??):117–135, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306216.

Babolian:2017:SIM

[BH17a] Esmail Babolian and Danial Hamedzadeh. A splitting iterative method for solving second kind integral equations in reproducing kernel spaces. Journal of Computational and Applied Mathematics, 326(??):204–216, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302704.

Brown:2017:HFE

[BH17b] Donald L. Brown and Viet Ha Hoang. A hierarchical finite element Monte Carlo method for stochastic two-scale elliptic equations. Journal of Computational and Applied Mathematics, 323(??):16–35, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301590.

Budzko:2016:DSW

[BHMT16] Dzmitry A. Budzko, José L. Hueso, Eulalia Martínez, and Carles Teruel. Dynamical study while searching equilibrium solutions in N-body problem. Journal of Computational and Applied Mathematics, 297 (??):26–40, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/
pii/S0377042715005555.

Banda:2015:LCC

[BHN15a] Mapundi K. Banda, Michael Herty, and Jean Medard T. Ngnotchouye. On linearized coupling conditions for a class of isentropic multiphase driftflux models at pipe-to-pipe intersections. Journal of Computational and Applied Mathematics, 276 (??):81–97, March 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714003781.

Bultheel:2015:HFA

[BHN15b] Adhemar Bultheel, Erik Hendriksen, and Olav Njåstad. Holomorphic functions associated with indeterminate rational moment problems. Journal of Computational and Applied Mathematics, 284(??):101–114, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002726.

Bessaih:2018:NAS

[BHRR18] Hakima Bessaih, Erika Hausenblas, Tsiry Avisoa Randrianasolo, and Paul André Razafimandimby. Numerical approximation of stochastic evolution equations: Convergence in scale of Hilbert spaces. Journal of Computational and Applied Mathematics, 343(??):250–274, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302711.

Botchev:2018:EKT

[BHU18] M. A. Botchev, A. M. Hanse, and R. Uppu. Exponential Krylov time integration for modeling multi-frequency optical response with monochromatic sources. Journal of Computational and Applied Mathematics, 340(??):474–485, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306349.

Barrera:2017:HSI

[BI17] D. Barrera and M. J. Ibáñez. Hermite spline interpolation on a three direction mesh from Powell– Sabin and Hsieh–Clough– Tocher finite elements. Journal of Computational and Applied Mathematics, 318(??):565–579, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S037704271500624X.

Bieniasz:2017:AHM

[Bie17]  L. K. Bieniasz.  An

adaptive Huber method for nonlinear systems of Volterra integral equations with weakly singular kernels and solutions. Journal of Computational and Applied Mathematics, 323(??):136–146, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301887.

Barrera:2019:SQI

[BIJM^{+}19] D. Barrera, M. J. Ibáñez, F. Jiménez-Molinos, A. M. Roldán, and J. B. Roldán. A spline quasi-interpolation based method to obtain the reset voltage in resistive RAMs in the charge– flux domain. Journal of Computational and Applied Mathematics, 354 (??):326–333, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717306404.

Barrera:2017:CTN

[BIR17a] D. Barrera, M. J. Ibáñez, and S. Remogna. On the construction of trivariate near-best quasi-interpolants based on C^{2} quartic splines on type-6 tetrahedral partitions. Journal of Computational and Applied Mathematics, 311(??):252–261, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303508.

Barrera:2017:PPF

[BIR^{+}17b] D. Barrera, M. J. Ibáñez, A. M. Roldán, J. B. Roldán, and R. Yáñez. Polynomial pattern finding in scattered data. Journal of Computational and Applied Mathematics, 318 (??):107–116, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716305544.

Baccouch:2016:HOD

[BJ16] Mahboub Baccouch and Bryan Johnson. A high-order discontinuous Galerkin method for Itô stochastic ordinary differential equations. Journal of Computational and Applied Mathematics, 308 (??):138–165, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302680.

Bartoszewski:2015:NST

[BJK15] Z. Bartoszewski, Z. Jackiewicz, and Y. Kuang. Numerical solution of threshold problems in epidemics and population dynamics. Journal of Computational and Applied Mathematics, 279(??):40–56, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400466X.

Bentbib:2019:ABT

[BJK19] A. H. Bentbib, K. Jbilou, and Y. Kaouane. An adaptive block tangential method for multiinput multi-output dynamical systems. Journal of Computational and Applied Mathematics, 358(??):190–205, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301335.

Babolian:2016:EAR

[BJM16] Esmail Babolian, Shahnam Javadi, and Eslam Moradi. Error analysis of reproducing kernel Hilbert space method for solving functional integral equations. Journal of Computational and Applied Mathematics, 300(??):300– 311, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716000121.

Bause:2015:VTD

[BK15a] Markus Bause and Uwe Köcher. Variational time discretization for mixed finite element approximations of nonstationary diffusion problems. Journal of Computational and Applied Mathematics, 289(??):208–224, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000850.

Bayramov:2015:SST

[BK15b] N. R. Bayramov and J. K. Kraus. On the stable solution of transient convection–diffusion equations. Journal of Computational and Applied Mathematics, 280(??):275–293, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005391.

Bhatt:2015:LEE

[BK15c] H. P. Bhatt and A. Q. M.

Khaliq. The locally extrapolated exponential time differencing LOD scheme for multidimensional reaction–diffusion systems. Journal of Computational and Applied Mathematics, 285 (??):256–278, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000874.

Bhatt:2016:CFO

[BK16] H. P. Bhatt and A. Q. M.

Khaliq. A compact fourthorder L-stable scheme for reaction–diffusion systems with nonsmooth data. Journal of Computational and Applied Mathematics, 299(??):176– 193, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006068.

Barton:2019:NQC

[BK19a] Michael Barton and Jirí Kosinka. On numerical quadrature for C^{1} quadratic Powell–Sabin 6-split macro-triangles. Journal of Computational and Applied Mathematics, 349(??):239–250, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304783.

Borsos:2019:VPS

[BK19b] B. Borsos and J. Karátson. Variable preconditioning for strongly nonlinear elliptic problems. Journal of Computational and Applied Mathematics, 350 (??):155–164, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718306058.

Bratsos:2019:ETD

[BK19c] A. G. Bratsos and A. Q. M. Khaliq. An exponential time differencing method of lines for Burgers–Fisher and coupled Burgers equations. Journal of Computational and Applied Mathematics, 356(??):182–197, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300445.

Burger:2019:CUQ

[BK19d] Raimund Bürger and Ilja Kröker. Computational uncertainty quantification for some strongly degenerate parabolic convectiondiffusion equations. Journal of Computational and Applied Mathematics, 348 (??):490–508, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S037704271830551X.

Benes:2018:FBM

[BKK18] Michal Benes, Tomás Krejcí, and Jaroslav Kruis. A FETI-based mixed explicit-implicit multitime-step method for parabolic problems. Journal of Computational and Applied Mathematics, 333 (??):247–265, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717305502.

Berenguer:2015:GSI

[BKLR15] M. I. Berenguer, H. Kunze, D. La Torre, and M. Ruiz Galán. Galerkin schemes and inverse boundary value problems in reflexive Banach spaces. Journal of Computational and Applied Mathematics, 275(??):100–112, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003604.

Berenguer:2016:GMC

[BKLR16] M. I. Berenguer, H. Kunze, D. La Torre, and M. Ruiz Galán. Galerkin method for constrained variational equations and a collagebased approach to related inverse problems. Journal of Computational and Applied Mathematics, 292(??):67–75, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003386.

Bukovsek:2019:NEC

[BKMO19] Damjana Kokol Bukovsek, Tomaz Kosir, Blaz Mojskerc, and Matjaz Omladic. Non-exchangeability of copulas arising from shock models. Journal of Computational and Applied Mathematics, 358(??):61–83, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301049. See

erratum [BKMO20].

Bukovsek:2020:ENE

[BKMO20] Damjana Kokol Bukovsek, Tomaz Kosir, Blaz Mojskerc, and Matjaz Omladic. Erratum to “Nonexchangeability of copulas arising from shock models” [J. Comput. Appl. Math. 358 (2019) 61–83]. Journal of Computational and Applied Mathematics, 365(??):Article 112419, February 2020. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719304224. See

[BKMO19].

Baeumer:2018:BCF

[BKMS18a] Boris Baeumer, Mihály Kovács, Mark M. Meerschaert, and Harish Sankaranarayanan. Boundary conditions for fractional diffusion. Journal of Computational and Applied Mathematics, 336(??):408– 424, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300219. See

reprint [BKMS18b].

Baeumer:2018:RBC

[BKMS18b] Boris Baeumer, Mihály Kovács, Mark M. Meerschaert, and Harish Sankaranarayanan. Reprint of: Boundary conditions for fractional diffusion. Journal of Computational and Applied Mathematics, 339(??):414–430, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301365. See

[BKMS18a].

Borluk:2016:NSW

[BKN16] Handan Borluk, Henrik Kalisch, and David P. Nicholls. A numerical study of the Whitham equation as a model for steady surface water waves. Journal of Computational and Applied Mathematics, 296(??):293–302, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004707.

Breuer:2016:MFK

[BL16] Alex Breuer and Andrew Lumsdaine. Matrix-free Krylov iteration for implicit convolution of numerically low-rank data. Journal of Computational and Applied Mathematics, 308(??):98–116, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302229.

Benterki:2017:CLC

[BL17] Rebiha Benterki and Jaume Llibre. Centers and limit cycles of polynomial differential systems of degree 4 via averaging theory. Journal of Computational and Applied Mathematics, 313(??):273–283, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304162.

Baramidze:2018:NDI

[BL18] V. Baramidze and M. J.

Lai. Nonnegative data interpolation by spherical splines. Journal of Computational and Applied Mathematics, 342(??):463–477, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718302322.

Brentan:2017:HRM

[BLH^{+}17] Bruno M. Brentan, Edevar Luvizotto, Jr., Manuel Herrera, Joaquín Izquierdo, and Rafael Pérez-García. Hybrid regression model for near real-time urban water demand forecasting. Journal of Computational and Applied Mathematics, 309(??):532–541, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300565.

Bultheel:2015:FPO

[BLLM^{+}15] Adhemar Bultheel, Guillermo López-Lagomasino, Francisco Marcellán, Olav Njåstad, and Ramón Orive. Foreword to the Proceedings of the OrthoQuad 2014 Conference. Journal of Computational and Applied Mathematics, 284(??):1–9, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000473.

Beyrami:2016:NEM

[BLM16] Hossein Beyrami, Taher Lotfi, and Katayoun Mahdiani. A new efficient method with error analysis for solving the second kind Fredholm integral equation with Cauchy kernel. Journal of Computational and Applied Mathematics, 300(??):385– 399, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000224.

Bracco:2019:TSS

[BLMS19] Cesare Bracco, Tom Lyche, Carla Manni, and Hendrik Speleers. Tchebycheffian spline spaces over planar T -meshes: Dimension bounds and dimension instabilities. Journal of Computational and Applied Mathematics, 349(??):265–278, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306332.

Bonito:2017:APE

[BLP17] Andrea Bonito, Wenyu Lei, and Joseph E. Pasciak. The approximation of parabolic equations involving fractional powers of elliptic operators. Journal of Computational and Applied Mathematics, 315(??):32–48, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305052.

Bizzarri:2019:LCA

[BLV19] Michal Bizzarri, Miroslav Lávicka, and Jan Vrsek. Linear computational approach to interpolations with polynomial Minkowski Pythagorean hodograph curves. Journal of Computational and Applied Mathematics, 361(??):283–294, December 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042719302262.

Bai:2018:NMC

[BLXD18] Jianchao Bai, Jicheng Li, Fengmin Xu, and Pingfan Dai. A novel method for a class of structured low-rank minimizations with equality constraint. Journal of Computational and Applied Mathematics, 330(??):475–487, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304521.

Babaaghaie:2017:NSI

[BM17a] A. Babaaghaie and K. Maleknejad. Numerical solution of integro-differential equations of high order by wavelet basis, its algorithm and convergence analysis. Journal of Computational and Applied Mathematics, 325(??):125–133, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302066.

Babaaghaie:2017:NSN

[BM17b] A. Babaaghaie and K. Maleknejad. Numerical solutions of nonlinear two-dimensional partial Volterra integrodifferential equations by Haar wavelet. Journal of Computational and Applied Mathematics, 317 (??):643–651, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716306227.

Babaaghaie:2018:NAN

[BM18] A. Babaaghaie and K. Maleknejad. A new approach for numerical solution of two-dimensional nonlinear Fredholm integral equations in the most general kind of kernel, based on Bernstein polynomials and its convergence analysis. Journal of Computational and Applied Mathematics, 344(??):482–494, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303509.

Bhatt:2019:EIP

[BM19a] Ashish Bhatt and Brian E. Moore. Exponential integrators preserving local conservation laws of PDEs with timedependent damping/driving forces. Journal of Computational and Applied Mathematics, 352(??):341–351, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307283.

Borsche:2019:MMM

[BM19b] R. Borsche and A. Meurer. Microscopic and macroscopic models for coupled car traffic and pedestrian flow. Journal of Computational and Applied Mathematics, 348(??):356–382, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305247.

Bruzon:2019:CLG

[BMG^{+}19] M. S. Bruzón, A. P.

Márquez, T. M. Garrido, E. Recio, and R. de la Rosa. Conservation laws for a generalized seventh order KdV equation. Journal of Computational and Applied Mathematics, 354(??):682– 688, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830699X.

Bouhamed:2015:SSB

[BMLR15] Heni Bouhamed, Afif Masmoudi, Thierry Lecroq, and Ahmed Rebaï. Structure space of Bayesian networks is dramatically reduced by subdividing it in sub-networks. Journal of Computational and Applied Mathematics, 287(??):48–62, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001429.

Behl:2017:FSD

[BMM17] R. Behl, P. Maroju, and S. S. Motsa. A family of second derivative free fourth order continuation method for solving nonlinear equations. Journal of Computational and Applied Mathematics, 318(??):38– 46, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630601X.

Burg:2015:GOA

[BN15] Markus Bürg and Murtazo Nazarov. Goaloriented adaptive finite element methods for elliptic problems revisited. Journal of Computational and Applied Mathematics, 287(??):125–147, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001843.

Balbastre:2019:MPE

[BN19] J. V. Balbastre and L. Nuño. Modelling the propagation of electromagnetic waves across complex metamaterials in closed structures. Journal of Computational and Applied Mathematics, 352(??):40–49, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306691.

Bayramoglu:2016:MRL

[BO16] Ismihan Bayramoğlu (Bairamov)

and Murat Ozkut. Mean residual life and inactivity time of a coherent system subjected to Marshall–Olkin type shocks. Journal of Computational and Applied Mathematics, 298(??):190–200, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006214.

Badia:2018:NPT

[BO18] Santiago Badia and Marc Olm. Nonlinear parallelin-time Schur complement solvers for ordinary differential equations. Journal of Computational and Applied Mathematics, 344(??):794–806, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304648.

Bastos:2016:FBS

[BOF^{+}16] Flavia S. Bastos, Evelyn A. Oliveira, Leonardo G.

Fonseca, Sabrina M. Vargas, and Estevam B. Las Casas. A FEMbased study on the influence of skewness and kurtosis surface texture parameters in human dental occlusal contact. Journal of Computational and Applied Mathematics, 295(??):139–148, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000308.

Boglaev:2015:MIS

[Bog15] Igor Boglaev. Monotone iterates for solving nonlinear integro-parabolic equations of Volterra type. Journal of Computational and Applied Mathematics, 290(??):224–238, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003039.

Bogosel:2016:MFS

[Bog16] Beniamin Bogosel. The method of fundamental solutions applied to boundary eigenvalue problems. Journal of Computational and Applied Mathematics, 306(??):265–285, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630173X.

Boglaev:2019:PUN

[Bog19]  Igor Boglaev.  A pa-

rameter uniform numerical method for a nonlinear elliptic reaction-diffusion problem. Journal of Computational and Applied Mathematics, 350(??):178– 194, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306241.

Botchev:2016:KSE

[Bot16] Mikhail A. Botchev. Krylov subspace exponential time domain solution of Maxwell’s equations in photonic crystal modeling. Journal of Computational and Applied Mathematics, 293(??):20–34, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500240X.

Bouharguane:2018:FEM

[Bou18] Afaf Bouharguane. Finite element method for a space-fractional antidiffusive equation. Journal of Computational and Applied Mathematics, 328(??):497–507, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303576.

Berinde:2017:CCF

[BP17a] Vasile Berinde and Madalina Pacurar. Corrigendum to: “Coincidence and fixed points for multivalued mappings and its application to nonconvex integral inclusions” [J. Comput. Appl. Math. 283 (2015) 201– 217]. Journal of Computational and Applied Mathematics, 311(??):718–720, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303661. See

[PAC15].

Boscarino:2017:API

[BP17b] S. Boscarino and L. Pareschi. On the asymptotic properties of IMEX Runge– Kutta schemes for hyperbolic balance laws. Journal of Computational and Applied Mathematics, 316(??):60–73, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630396X.

Blasco:2015:ASC

[BPD15a] Angel Blasco and Sonia Pérez-Díaz. Asymptotes of space curves. Journal of Computational and Applied Mathematics, 278(??):231–247, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004592.

Blasco:2015:CFH

[BPD15b] Angel Blasco and Sonia Pérez-Díaz. Characterizing the finiteness of the Hausdorff distance between two algebraic curves. Journal of Computational and Applied Mathematics, 280(??):327–346, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005469.

Beauregard:2017:NSA

[BPP17] Matthew A. Beauregard, Joshua Padgett, and Rana Parshad. A nonlinear splitting algorithm for systems of partial differential equations with self-diffusion. Journal of Computational and Applied Mathematics, 321(??):8–25, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300845.

Becher:2015:RES

[BR15a] S. Becher and H.-G. Roos. Richardson extrapolation for a singularly perturbed turning point problem with exponential boundary layers. Journal of Computational and Applied Mathematics, 290(??):334–351, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003167.

Bunder:2015:NIO

[BR15b] J. E. Bunder and A. J. Roberts. Numerical integration of ordinary differential equations with rapidly oscillatory factors. Journal of Computational and Applied Mathematics, 282(??):54– 70, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005895.

Bracco:2016:SGS

[BR16] Cesare Bracco and Fabio Roman. Spaces of generalized splines over T - meshes. Journal of Computational and Applied Mathematics, 294(??):102– 123, March 1, 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715004136.

Boutsikas:2016:NCU

[BRA16] M. V. Boutsikas, A. C.

Rakitzis, and D. L. Antzoulakos. On the number of claims until ruin in a two-barrier renewal risk model with Erlang mixtures. Journal of Computational and Applied Mathematics, 294(??):124–137, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004203.

Bellalij:2015:SPR

[BRS15] M. Bellalij, L. Reichel, and H. Sadok. Some properties of range restricted GMRES methods. Journal of Computational and Applied Mathematics, 290(??):310–318, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003027.

Ba:2016:CMM

[BRW16] Wenlan Ba, Ning Ren, and Haihong Wu. A curve modeling method based on the envelope template. Journal of Computational and Applied Mathematics, 308(??):187–198, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302710.

Brezinski:2015:NRP

[BRZ15] Claude Brezinski and Michela Redivo-Zaglia. New representations of Padé, Padé-type, and partial Padé approximants. Journal of Computational and Applied Mathematics, 284(??):69–77, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003161.

Borhanifar:2015:NOA

[BS15] A. Borhanifar and Kh.

Sadri. A new operational approach for numerical solution of generalized functional integro-differential equations. Journal of Computational and Applied Mathematics, 279 (??):80–96, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714004403.

Banz:2016:CMB

[BS16] Lothar Banz and Ernst P.

Stephan. Comparison of mixed hp-BEM (stabilized and non-stabilized) for frictional contact problems. Journal of Computational and Applied Mathematics, 295(??):92–102, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000515.

Belgacem:2017:DST

[BS17a] Fethi Bin Muhammad Belgacem and Rathinavel Silambarasan. A distinctive Sumudu treatment of trigonometric functions. Journal of Computational and Applied Mathematics, 312(??):74–81, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716000029.

Benchaabane:2017:STF

[BS17b] Abbes Benchaabane and Rathinasamy Sakthivel. Sobolev-type fractional stochastic differential equations with non-Lipschitz coefficients. Journal of Computational and Applied Mathematics, 312 (??):65–73, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715006329.

Burman:2017:EET

[BS17c] Erik Burman and Isaac P. Santos. Error estimates for transport problems with high Péclet number using a continuous dependence assumption. Journal of Computational and Applied Mathematics, 309(??):267–286, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302965.

Baines:2018:MMF

[BS18] M. J. Baines and N. Sarahs. A moving-mesh finite difference scheme that preserves scaling symmetry for a class of nonlinear diffusion problems. Journal of Computational and Applied Mathematics, 340(??):380–389, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301195.

Biazar:2019:SWS

[BS19] Jafar Biazar and Khadijeh Sadri. Solution of weakly singular fractional integro-differential equations by using a new operational approach. Journal of Computational and Applied Mathematics, 352(??):453–477, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307337.

Behl:2019:HEF

[BSGM19] Ramandeep Behl, Í. Sarría, R. González, and Á.A. Magreñán. Highly efficient family of iterative methods for solving nonlinear models. Journal of Computational and Applied Mathematics, 346(??):110–132, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303947.

Bai:2018:REM

[BSLL18] Xuchao Bai, Yimin Shi, Yiming Liu, and Bin Liu. Reliability estimation of multicomponent stress-strength model based on copula function under progressively hybrid censoring. Journal of Computational and Applied Mathematics, 344 (??):100–114, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302693.

Bai:2019:RES

[BSLL19] Xuchao Bai, Yimin Shi, Yiming Liu, and Bin Liu. Reliability estimation of stress-strength model using finite mixture distributions under progressively interval censoring. Journal of Computational and Applied Mathematics, 348(??):509–524, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305685.

Benkhaldoun:2015:FFV

[BSS15] Fayssal Benkhaldoun, Saida Sari, and Mohammed Seaid. A family of finite volume Eulerian– Lagrangian methods for two-dimensional conservation laws. Journal of Computational and Applied Mathematics, 285 (??):181–202, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000849.

Bayour:2017:ESL

[BT17] Benaoumeur Bayour and Delfim F. M. Torres. Existence of solution to a local fractional nonlinear differential equation. Journal of Computational and Applied Mathematics, 312(??):127–133, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271600025X.

Beilina:2015:GCA

[BTKM15] Larisa Beilina, Nguyen Trung Thành, Michael V. Klibanov, and John Bondestam Malmberg. Globally convergent and adaptive finite element methods in imaging of buried objects from experimental backscattering radar measurements. Journal of Computational and Applied Mathematics, 289(??):371–391, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005342.

Bazhlekov:2016:NMD

[BV16a] I. Bazhlekov and D. Vasileva. Numerical modeling of drop coalescence in the presence of soluble surfactants. Journal of Computational and Applied Mathematics, 293(??):7–19, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002319.

Brown:2016:GMFa

[BV16b] Donald L. Brown and Maria Vasilyeva. A Generalized Multiscale Finite Element Method for poroelasticity problems. I: Linear problems. Journal of Computational and Applied Mathematics, 294(??):372–388, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004148.

Brown:2016:GMFb

[BV16c] Donald L. Brown and Maria Vasilyeva. A Generalized Multiscale Finite Element Method for poroelasticity problems. II: Nonlinear coupling. Journal of Computational and Applied Mathematics, 297(??):132–146, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500552X.

Badia:2018:RSD

[BV18] Santiago Badia and Francesc Verdugo. Robust and scalable domain decomposition solvers for unfitted finite element methods. Journal of Computational and Applied Mathematics, 344(??):740–759, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730465X.

Bank:2017:ADC

[BVZ17] Randolph E. Bank, Panayot S.

Vassilevski, and Ludmil T. Zikatanov. Arbitrary dimension convection– diffusion schemes for spacetime discretizations. Journal of Computational and Applied Mathematics, 310(??):19–31, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302023.

Batselier:2016:STD

[BW16] Kim Batselier and Ngai Wong. Symmetric tensor decomposition by an iterative eigendecomposition algorithm. Journal of Computational and Applied Mathematics, 308(??):69–82, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302588.

Batselier:2017:IMP

[BW17] Kim Batselier and Ngai Wong. Inverse multivariate polynomial rootfinding: numerical implementations of the affine and projective Buchberger– Möller algorithm. Journal of Computational and Applied Mathematics, 320: 15–29, August 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Bogrek:2019:SMW

[BW19] Betul Bogrek and Xiaoshen Wang. Superconvergence of a modified weak Galerkin approximation for second order elliptic problems by L^{2} projection method. Journal of Computational and Applied Mathematics, 346(??):53–62, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830387X.

Bourne:2017:NLS

[BWS17] Martin Bourne, Joab R.

Winkler, and Yi Su. A non-linear structurepreserving matrix method for the computation of the coefficients of an approximate greatest common divisor of two Bernstein polynomials. Journal of Computational and Applied Mathematics, 320(??):221–241, August 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042717300560.

Wu:2016:SFO

[bWzZjH16] Zeng bao Wu, Yun zhi Zou, and Nan jing Huang. A system of fractional-order interval projection neural networks. Journal of Computational and Applied Mathematics, 294(??):389– 402, March 1, 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715004598.

Bylina:2018:BWF

[Byl18]  Beata Bylina.  The

block WZ factorization. Journal of Computational and Applied Mathematics, 331(??):119–132, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304946.

Brenner:2019:CIP

[BySZ19] Susanne C. Brenner, Li yeng Sung, and Yi Zhang. C^{0} interior penalty methods for an elliptic stateconstrained optimal control problem with Neumann boundary condition. Journal of Computational and Applied Mathematics, 350(??):212–232, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042718306228.

Bakan:2018:MTE

[BYW18] Hacer Öz Bakan, Fikriye Yilmaz, and Gerhard-Wilhelm Weber. Minimal truncation error constants for Runge–Kutta method for stochastic optimal control problems. Journal of Computational and Applied Mathematics, 331(??):196–207, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305010.

Bani-Yaghoub:2019:ECP

[BYWPA19] Majid Bani-Yaghoub, Xueying Wang, Patrick O. Pithua, and Sharif S. Aly. Effectiveness of control and preventive measures influenced by pathogen trait evolution: Example of Escherichia coli O157:H7 . Journal of Computational and Applied Mathematics, 362 (??):366–382, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718305533.

Bani-Yaghoub:2016:ESS

[BYYV16] Majid Bani-Yaghoub, Guangming Yao, and Hristo Voulov. Existence and stability of stationary waves of a population model with strong Allee effect. Journal of Computational and Applied Mathematics, 307(??):385–393, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271500566X.

Brdar:2016:SPP

[BZ16a] Mirjana Brdar and Helena Zarin. A singularly perturbed problem with two parameters on a bakhvalov-type mesh. Journal of Computational and Applied Mathematics, 292(??):307–319, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003738.

Brown:2016:END

[BZ16b] David A. Brown and David W. Zingg. Efficient numerical differentiation of implicitly-defined curves for sparse systems. Journal of Computational and Applied Mathematics, 304(??):138–159, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301121.

Bao:2019:OPM

[BZ19a] Jiayong Bao and Yuexu Zhao. Option pricing in Markov-modulated exponential Lévy models with stochastic interest rates. Journal of Computational and Applied Mathematics, 357(??):146–160, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300603.

Brown:2019:MHC

[BZ19b] David A. Brown and David W. Zingg. Monolithic homotopy continuation with predictor based on higher derivatives. Journal of Computational and Applied Mathematics, 346(??):26–41, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303881.

Cordeiro:2019:OLG

[CAO^{+}19] Gauss M. Cordeiro, Ahmed Z. Afify, Edwin M. M. Ortega, Adriano K. Suzuki, and Mohamed E. Mead. The odd Lomax generator of distributions: Properties, estimation and applications. Journal of Computational and Applied Mathematics, 347(??):222–237, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304850.

Calabro:2019:NRD

[CBC^{+}19] F. Calabrò, D. Bravo, C. Carissimo, E. Di Fazio, A. Di Pasquale, A. A. M. O. Eldray, C. Fabrizi, J. G. S. Gerges, S. Palazzo, and J. F. F. T. Wassef. Null rules for the detection of lower regularity of functions. Journal of Computational and Applied Mathematics, 361(??):547–553, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719302304.

Cruz-Barroso:2015:MOP

[CBMO15] Ruymán Cruz-Barroso, Carlos Díaz Mendoza, and Ramón Orive. Multiple orthogonal polynomials on the unit circle. Normality and recurrence relations. Journal of Computational and Applied Mathematics, 284(??):115–132, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004750.

Cruz-Barroso:2015:CBS

[CBMPP15] Ruymán Cruz-Barroso, Carlos Díaz Mendoza, and Francisco Perdomo-Pío. A connection between Szegő–Lobatto and quasi Gauss-type quadrature formulas. Journal of Computational and Applied Mathematics, 284(??):133–143, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005007.

Calo:2018:EAA

[CBP18] Victor Calo, Santiago Badia, and Javier Principe. Editorial: Algorithmic aspects of high-performance computing for mechanics and physics. Journal of Computational and Applied Mathematics, 344(??):739, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304060.

Cross:2015:ERA

[CC15] Benjamin Cross and Robert J. Cripps. Efficient robust approximation of the generalised Cornu spiral. Journal of Computational and Applied Mathematics, 273(??):1–12, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002660.

Chakraborty:2019:NDF

[CCC19] Tanujit Chakraborty, Ashis Kumar Chakraborty, and Swarup Chattopadhyay. A novel distribution-free hybrid regression model for manufacturing process efficiency improvement. Journal of Computational and Applied Mathematics, 362(??):130–142, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271930247X.

Chen-Charpentier:2016:MMB

[CCD16] Benito M. Chen-Charpentier and Ibrahim Diakite. A mathematical model of bone remodeling with delays. Journal of Computational and Applied Mathematics, 291(??):76–84, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005044.

Carpitella:2018:IIS

[CCIL18] Silvia Carpitella, Antonella Certa, Joaquín Izquierdo, and Concetta Manuela La Fata. k-out-of-n systems: an exact formula for the stationary availability and multi-objective configuration design based on mathematical programming and TOPSIS. Journal of Computational and Applied Mathematics, 330(??):1007–1015, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300109.

Casaban:2016:SRM

[CCJ16] M.-C. Casabán, J.-C.

Cortés, and L. Jódar. Solving random mixed heat problems: a random integral transform approach. Journal of Computational and Applied Mathematics, 291(??):5–19, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004282.

Casaban:2018:SLQ

[CCJ18] M. C. Casabán, J. C. Cortés, and L. Jódar. Solving linear and quadratic random matrix differential equations using: a mean square approach. the nonautonomous case. Journal of Computational and Applied Mathematics, 330 (??):937–954, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S037704271630591X.

Casaban:2016:PSH

[CCNQ^{+}16] M.-C. Casabán, J.-C. Cortés, A. Navarro-Quiles, J.-V. Romero, M.-D. Roselló, and R.-J. Villanueva. Probabilistic solution of the homogeneous Riccati differential equation: a case-study by using linearization and transformation techniques. Journal of Computational and Applied Mathematics, 291(??):20–35, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400507X.

Casaban:2017:CPS

[CCNQ^{+}17] M.-C. Casabán, J.-C.

Cortés, A. Navarro-Quiles, J.-V. Romero, M.-D. Roselló, and R.-J. Villanueva. Computing probabilistic solutions of the Bernoulli random differential equation. Journal of Computational and Applied Mathematics, 309(??):396–407, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300814.

Cordeiro:2019:GGN

[CCRN19] Gauss M. Cordeiro, Renato J. Cintra, Leandro C. Rêgo, and Abraão D. C. Nascimento. The gamma generalized normal distribution: a descriptor of SAR imagery. Journal of Computational and Applied Mathematics, 347(??):257–272, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718304710.

Chicharro:2019:DIF

[CCT19] Francisco I. Chicharro, Alicia Cordero, and Juan R. Torregrosa. Dynamics of iterative families with memory based on weight functions procedure. Journal of Computational and Applied Mathematics, 354(??):286– 298, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300487.

Campos:2017:SKF

[CCTV17] Beatriz Campos, Alicia Cordero, Juan R. Torregrosa, and Pura Vindel. Stability of King’s family of iterative methods with memory. Journal of Computational and Applied Mathematics, 318(??):504– 514, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300140.

Cao:2017:IMM

[CCW17] Xuenian Cao, Xianxian Cao, and Liping Wen. The implicit midpoint method for the modified anomalous sub-diffusion equation with a nonlinear source term. Journal of Computational and Applied Mathematics, 318(??):199– 210, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305039.

Chen:2018:WGF

[CCX18] Yumei Chen, Gang Chen, and Xiaoping Xie. Weak Galerkin finite element method for Biot’s consolidation problem. Journal of Computational and Applied Mathematics, 330 (??):398–416, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042717304508.

Cai:2018:NES

[CCY18] Chunhao Cai, Nan Chen, and Honglong You. Nonparametric estimation for a spectrally negative Lévy risk process based on low-frequency observation. Journal of Computational and Applied Mathematics, 328(??):432–442, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042717303655.

Cao:2018:NES

[CCZ^{+}18] Junjie Cao, He Chen, Jie Zhang, Yujiao Li, Xiuping Liu, and Changqing Zou. Normal estimation via shifted neighborhood for point cloud. Journal of Computational and Applied Mathematics, 329(??):57–67, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301978.

Cao:2016:IFE

[CCZZ16] Yong Cao, Yuchuan Chu, Xiaoshi Zhang, and Xu Zhang. Immersed finite element methods for unbounded interface problems with periodic structures. Journal of Computational and Applied Mathematics, 307(??):72–81, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301935.

Crevecoeur:2015:CFT

[CD15] Guillaume Crevecoeur and Rob H. De Staelen. On cost function transformations for the reduction of uncertain model parameters’ impact towards the optimal solutions. Journal of Computational and Applied Mathematics, 289(??):392–399, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005093.

Carpes:2016:NAA

[CD16a] Charles Quevedo Carpes and Álvaro Luiz De Bortoli. Numerical analysis of the acoustics of a diffusion flame. Journal of Computational and Applied Mathematics, 295(??):115–126, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000345.

Chen:2016:LSE

[CD16b] Zhongwen Chen and Yu-Hong Dai. A line search exact penalty method with bi-object strategy for nonlinear constrained optimization. Journal of Computational and Applied Mathematics, 300(??):245– 258, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715006457.

Chen:2015:OPR

[CDD^{+}15] Xinliang Chen, Griselda Deelstra, Jan Dhaene, Daniël Linders, and Michèle Vanmaele. On an optimization problem related to static super-replicating strategies. Journal of Computational and Applied Mathematics, 278 (??):213–230, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042714004415.

Cavoretto:2019:FCT

[CDDD19] Roberto Cavoretto, Alessandra De Rossi, Francesco Dell’Accio, and Filomena Di Tommaso. Fast computation of triangular Shepard interpolants. Journal of Computational and Applied Mathematics, 354 (??):457–470, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718301584.

Cicuttin:2018:IDS

[CDE18] M. Cicuttin, D. A. Di Pietro, and A. Ern. Implementation of discontinuous skeletal methods on arbitrary-dimensional, polytopal meshes using generic programming. Journal of Computational and Applied Mathematics, 344(??):852–874, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730448X.

Castro:2016:HHO

[CDF^{+}16] Douglas A. Castro, Philippe R. B. Devloo, Agnaldo M. Farias, Sônia M. Gomes, and Omar Durán. Hierarchical high order finite element bases for h(div) spaces based on curved meshes for twodimensional regions or manifolds. Journal of Computational and Applied Mathematics, 301 (??):241–258, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042716300334.

Chen:2017:AVT

[CDH^{+}17] Goong Chen, Zhonghai Ding, Tingwen Huang, Lanxuan Yu, Viswanath Ramakrishna, and Joseph Ward. Application of the virial theorem for improving eigenvalue calculations of multiparticle systems. Journal of Computational and Applied Mathematics, 313(??):469–485, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304964.

Cotronei:2017:IRK

[CDHP17] Mariantonia Cotronei, Rosa Di Salvo, Matthias Holschneider, and Luigia Puccio. Interpolation in reproducing kernel Hilbert spaces based on random subdivision schemes. Journal of Computational and Applied Mathematics, 311(??):342–353, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303624.

Conen:2014:CSH

[CDKN14] Lea Conen, Victorita Dolean, Rolf Krause, and Frédéric Nataf. A coarse space for heterogeneous Helmholtz problems based on the Dirichlet-to-Neumann operator. Journal of Computational and Applied Mathematics, 271(??):83–99, December 1, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714001800. See

addendum [CDKN15].

Conen:2015:ACS

[CDKN15] Lea Conen, Victorita Dolean, Rolf Krause, and Frédéric Nataf. Addendum to “A coarse space for heterogeneous Helmholtz problems based on the Dirichlet-to-Neumann operator” [J. Comput. Appl. Math. 271 (2014) 83–99]. Journal of Computational and Applied Mathematics, 290(??):670–674, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002502. See

[CDKN14].

Chu:2019:RGF

[CDN19] J. Chu, O. Dickin, and S. Nadarajah. A review of goodness of fit tests for Pareto distributions. Journal of Computational and Applied Mathematics, 361(??):13–41, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302067.

Cardone:2017:EFI

[CDP17]  Angelamaria  Cardone,

Raffaele D’Ambrosio, and Beatrice Paternoster. Exponentially fitted IMEX methods for advection– diffusion problems. Journal of Computational and Applied Mathematics, 316(??):100–108, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303946.

Calo:2019:DOQ

[CDP19] Victor Calo, Quanling Deng, and Vladimir Puzyrev. Dispersion optimized quadratures for isogeometric analysis. Journal of Computational and Applied Mathematics, 355(??):283– 300, August 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/
                                        pii/S037704271930041X.

Coelho:2018:ECT

[CDR18] Diego F. G. Coelho, Vassil S. Dimitrov, and L. Rakai. Efficient computation of tridiagonal matrices largest eigenvalue. Journal of Computational and Applied Mathematics, 330(??):268–275, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303977.

Cheung:2018:PSC

[CDRY18] Ka Chun Cheung, Jan Dhaene, Yian Rong, and Sheung Chi Phillip Yam. Probabilistic solutions for a class of deterministic optimal allocation problems. Journal of Computational and Applied Mathematics, 336(??):394– 407, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300207.

Cao:2015:RDP

[CDW15] Yang Cao, Jun-Liang Dong, and Yu-Ming Wang. A relaxed deteriorated PSS preconditioner for nonsymmetric saddle point problems from the steady Navier–Stokes equation. Journal of Computational and Applied Mathematics, 273(??):41–60, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002714.

Chen:2016:PDA

[CDZ16] Dai-Qiang Chen, Xin-Peng Du, and Yan Zhou. Primal-dual algorithm based on Gauss–Seidel scheme with application to multiplicative noise removal. Journal of Computational and Applied Mathematics, 292(??):609–622, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002137.

Ceballos:2017:NEI

[Ceb17] Marcelo A. Ceballos. Numerical evaluation of integrals involving the product of two Bessel functions and a rational fraction arising in some elastodynamic problems. Journal of Computational and Applied Mathematics, 313(??):355–382, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304721.

Company:2016:CPR

[CEJ16] R. Company, V. N. Egorova, and L. Jódar. Constructing positive reliable numerical solution for American call options: a new front-fixing approach. Journal of Computational and Applied Mathematics, 291(??):422–431, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004099.

Company:2018:CFS

[CEJ18] Rafael Company, Vera N.

Egorova, and Lucas Jódar. Conditional full stability of positivity-preserving finite difference scheme for diffusion–advection– reaction models. Journal of Computational and Applied Mathematics, 341(??):157–168, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301109.

Company:2016:FDM

[CEJV16] Rafael Company, Vera Egorova, Lucas Jódar, and Carlos Vázquez. Finite difference methods for pricing American put option with rationality parameter: Numerical analysis and computing. Journal of Computational and Applied Mathematics, 304(??):1–17, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301091.

Criado:2019:NCC

[CFG^{+}19] Regino Criado, Julio Flores, Esther García, Alejandro J. García del Amo, Ángel Pérez, and Miguel Romance. On the αnonbacktracking centrality for complex networks: Existence and limit cases. Journal of Computational and Applied Mathematics, 350(??):35– 45, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306009.

Canton:2016:GEC

[CFJMVG16] A. Cantón, L. Fernández-Jambrina, E. Rosado María, and M. J. Vázquez-Gallo. Geometric elements and classification of quadrics in rational Bézier form. Journal of Computational and Applied Mathematics, 300(??):400– 419, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716000108.

Copetti:2015:NAA

[CFMN15] M. I. M. Copetti, J. R.

Fernández, M. C. Muñiz, and C. Núñez. Numerical analysis of an adsorption dynamic model at the air-water interface. Journal of Computational and Applied Mathematics, 281(??):82– 93, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005482.

Cellat:2019:LSM

[CFMÖ19] Serdar Cellat, Yu Fan, Washington Mio, and Giray Ökten. Learning shape metrics with Monte Carlo optimization. Journal of Computational and Applied Mathematics, 348(??):120–129, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305302.

Candela:2018:GFC

[CFR18] V. Candela, A. Falcó, and Pantaleón D. Romero. A general framework for a class of non-linear approximations with applications to image restoration. Journal of Computational and Applied Mathematics, 330(??):982–994, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301188.

Chen:2019:FMC

[CFS19] Huangxin Chen, Xiaolin Fan, and Shuyu Sun. A fully mass-conservative iterative IMPEC method for multicomponent compressible flow in porous media. Journal of Computational and Applied Mathematics, 362(??):1–21, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302468.

Calabro:2018:EQR

[CFSS18a] Francesco Calabrò, Antonella Falini, Maria Lucia Sampoli, and Alessandra Sestini. Efficient quadrature rules based on spline quasi-interpolation for application to IGA-BEMs. Journal of Computational and Applied Mathematics, 338(??):153–167, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042718300761.

Chidyagwai:2018:CSL

[CFSS18b] Prince Chidyagwai, Martin Frank, Florian Schneider, and Benjamin Seibold. A comparative study of limiting strategies in discontinuous Galerkin schemes for the M_{1} model of radiation transport. Journal of Computational and Applied Mathematics, 342(??):399–418, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301857.

Cheng:2019:ESF

[CFWW19] Kelong Cheng, Wenqiang Feng, Cheng Wang, and Steven M. Wise. An energy stable fourth order finite difference scheme for the Cahn–Hilliard equation. Journal of Computational and Applied Mathematics, 362 (??):574–595, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718303169.

Chen:2017:RWF

[CFX17] Gang Chen, Minfu Feng, and Xiaoping Xie. A robust WG finite element method for convection– diffusion–reaction equations. Journal of Computational and Applied Mathematics, 315(??):107–125, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305180.

Cao:2017:CSI

[CG17a] Chunhong Cao and Xieping Gao. Compressed sensing image restoration based on data-driven multiscale tight frame. Journal of Computational and Applied Mathematics, 309(??):622–629, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301315.

Chowdhury:2017:IPM

[CG17b] Sudipto Chowdhury and Thirupathi Gudi. A C^{0} interior penalty method for the Dirichlet control problem governed by biharmonic operator. Journal of Computational and Applied Mathematics, 317(??):290– 306, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305982.

Cheung:2019:DDD

[CG19] Siu Wun Cheung and Nilabja Guha. Dynamic datadriven Bayesian GMs-FEM. Journal of Computational and Applied Mathematics, 353(??):72– 85, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307350.

Caravantes:2015:CFC

[CGMGV15] Jorge Caravantes, M. Angeles Gomez-Molleda, and Laureano Gonzalez-Vega. A canonical form for the continuous piecewise polynomial functions. Journal of Computational and Applied Mathematics, 283(??):17–27, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005123.

Cerdan:2018:PNL

[CGMM18] J. Cerdán, D. Guerrero, J. Marín, and J. Mas. Preconditioners for nonsymmetric linear systems with lowrank skew-symmetric part. Journal of Computational and Applied Mathematics, 343(??):318–327, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302036.

Costabile:2018:SRG

[CGN18] Francesco Aldo Costabile, Maria Italia Gualtieri, and Anna Napoli. Some results on generalized Szász operators involving Sheffer polynomials. Journal of Computational and Applied Mathematics, 337(??):244–255, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300451.

Criado:2016:GAS

[CGPR16] Regino Criado, Esther García, Francisco Pedroche, and Miguel Romance. On graphs associated to sets of rankings. Journal of Computational and Applied Mathematics, 291(??):497–508, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001557.

Chaabane:2017:CIC

[CGPR17] Nabil Chaabane, Vivette Girault, Charles Puelz, and Beatrice Riviere. Convergence of IPDG for coupled time-dependent Navier–Stokes and Darcy equations. Journal of Computational and Applied Mathematics, 324 (??):25–48, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717301577.

Clavero:2017:ENS

[CGSS17] C. Clavero, J. L. Gracia, G. I. Shishkin, and L. P. Shishkina. An efficient numerical scheme for 1D parabolic singularly perturbed problems with an interior and boundary layers. Journal of Computational and Applied Mathematics, 318(??):634– 645, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005415.

Cordero:2018:CMS

[CGT18] Alicia Cordero, Lucia Guasp, and Juan R. Torregrosa. Choosing the most stable members of Kou’s family of iterative methods. Journal of Computational and Applied Mathematics, 330(??):759–769, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300705.

Cools:2015:MLP

[CGvA^{+}15] Siegfried Cools, Pieter Ghysels, Wim van Aarle, Jan Sijbers, and Wim Vanroose. A multi-level preconditioned Krylov method for the efficient solution of algebraic tomographic reconstruction problems. Journal of Computational and Applied Mathematics, 283(??):1–16, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000096.

Celledoni:2017:ALM

[CH17a] Elena Celledoni and Eirik Hoel Høiseth. The averaged Lagrangian method. Journal of Computational and Applied Mathematics, 316(??):161–174, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304769.

Czuprynski:2017:EDR

[CH17b] Kenneth Czuprynski and Weimin Han. Energy dependent radiative transfer equation and energy discretization. Journal of Computational and Applied Mathematics, 323(??):147–158, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301619.

Chakir:2018:NIR

[CH18] R. Chakir and J. K.

Hammond. A nonintrusive reduced basis method for elastoplasticity problems in geotechnics. Journal of Computational and Applied Mathematics, 337(??):1–17, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300128.

Chien:2019:MCB

[CH19a] Yu-Tse Chien and Feng-Nan Hwang. A Markov chain-based multi-elimination preconditioner for elliptic PDE problems. Journal of Computational and Applied Mathematics, 362 (??):116–129, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302559.

Cirillo:2019:LCB

[CH19b] Emiliano Cirillo and Kai Hormann. On the Lebesgue constant of barycentric rational Hermite interpolants at equidistant nodes. Journal of Computational and Applied Mathematics, 349 (??):292–301, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042718303637.

Conti:2019:AAP

[CH19c] Costanza Conti and Svenja Hüning. An algebraic approach to polynomial reproduction of Hermite subdivision schemes. Journal of Computational and Applied Mathematics, 349(??):302–315, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042718304862.

Chen:2016:SCA

[Che16a] Qingshan Chen. Stable and convergent approximation of two-dimensional vector fields on unstructured meshes. Journal of Computational and Applied Mathematics, 307(??):284–306, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300292.

Chen:2016:RSA

[Che16b] Weidong Chen. A regularized sampling algorithm for reconstructing non-bandlimited signals. Journal of Computational and Applied Mathematics, 301(??):259–270, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005865.

Chen:2018:QMP

[Che18] Juan Chen. The quadrilateral Mindlin plate elements using the spline interpolation bases. Journal of Computational and Applied Mathematics, 329(??):68–83, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302984.

Chen:2019:TTB

[Che19a] Kuen-Suan Chen. Twotailed Buckley fuzzy testing for operating performance index. Journal of Computational and Applied Mathematics, 361(??):55–63, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719302079.

Chen:2019:DSC

[Che19b] Lizhen Chen. Direct solver for the Cahn–Hilliard equation by Legendre– Galerkin spectral method. Journal of Computational and Applied Mathematics, 358(??):34–45, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271930127X.

Chidyagwai:2017:MDM

[Chi17] Prince Chidyagwai. A multilevel decoupling method for the Navier–Stokes/ Darcy model. Journal of Computational and Applied Mathematics, 325(??):74–96, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302157.

Chen:2015:MGM

[CHR^{+}15] Jian Chen, Yong Huang, Haiwu Rong, Tingting Wu, and Taishan Zeng. A multiscale Galerkin method for second-order boundary value problems of Fredholm integrodifferential equation. Journal of Computational and Applied Mathematics, 290 (??):633–640, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003428.

Chretien:2018:NCS

[Chr18] Stéphane Chrétien. A note on computing the smallest conic singular value. Journal of Computational and Applied Mathematics, 340(??):221–230, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718301146. See

corrigendum [Chr19].

Chretien:2019:CNC

[Chr19] Stéphane Chrétien. Corrigendum to “A note on computing the Smallest Conic Singular Value” [J. Comput. Appl. Math. 340 (2018) 221–230]. Journal of Computational and Applied Mathematics, 347(??):369, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304345. See

[Chr18].

Chen:2015:SUM

[CHS15] Puyin Chen, Jianguo Huang, and Huashan Sheng. Some Uzawa methods for steady incompressible Navier–Stokes equations discretized by mixed element methods. Journal of Computational and Applied Mathematics, 273(??):313–325, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002921.

Chen:2017:SSI

[CHS17] Puyin Chen, Jianguo Huang, and Huashan Sheng. Solving steady incompressible Navier– Stokes equations by the Arrow–Hurwicz method. Journal of Computational and Applied Mathematics, 311(??):100–114, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303296.

Chen:2018:HMD

[CHS^{+}18] Gang Chen, Weiwei Hu, Jiguang Shen, John R. Singler, Yangwen Zhang, and Xiaobo Zheng. An HDG method for distributed control of convection diffusion PDEs. Journal of Computational and Applied Mathematics, 343(??):643–661, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302991.

Chung:2015:FNM

[Chu15] Won Sang Chung. Fractional Newton mechanics with conformable fractional derivative. Journal of Computational and Applied Mathematics, 290 (??):150–158, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002824.

Cordero:2015:SCU

[CHVRT15] A. Cordero, M. A. Hernández-Verón, N. Romero, and J. R. Torregrosa. Semilocal convergence by using recurrence relations for a fifth-order method in Banach spaces. Journal of Computational and Applied Mathematics, 273(??):205–213, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002817.

Cen:2018:ENM

[CHX18] Zhongdi Cen, Jian Huang, and Aimin Xu. An efficient numerical method for a two-point boundary value problem with a Caputo fractional derivative. Journal of Computational and Applied Mathematics, 336(??):1– 7, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717306386.

Chen:2019:HMC

[CHYZ19] Liping Chen, Lixing Han, Hongxia Yin, and Liangmin Zhou. A homotopy method for computing the largest eigenvalue of an irreducible nonnegative tensor. Journal of Computational and Applied Mathematics, 355(??):174–181, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300172.

Choi:2016:IRF

[CJ16] Youngwoo Choi and Jaewon Jung. Interpolating refinable function vectors with reflection properties. Journal of Computational and Applied Mathematics, 294(??):297–308, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004239.

Clavero:2018:SEO

[CJ18] C. Clavero and J. C Jorge.

Solving efficiently one dimensional parabolic singularly perturbed reactiondiffusion systems: a splitting by components. Journal of Computational and Applied Mathematics, 344(??):1–14, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302905.

Clavero:2019:ENM

[CJ19] C. Clavero and J. C Jorge. An efficient numerical method for singularly perturbed time dependent parabolic 2D convection-diffusion systems. Journal of Computational and Applied Mathematics, 354(??):431– 444, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306411.

Calvo-Jurado:2015:HPE

[CJCDLL15]  Carmen  Calvo-Jurado,

Juan Casado-Díaz, and Manuel Luna-Laynez. Homogenization of the Poisson equation with Dirichlet conditions in random perforated domains. Journal of Computational and Applied Mathematics, 275(??):375–381, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042714003136.

Calvo-Jurado:2019:BLH

[CJCDLL19]  Carmen  Calvo-Jurado,

Juan Casado-Díaz, and Manuel Luna-Laynez. A Brinkman law in the homogenization of the stationary Navier–Stokes system in a non-periodic porous medium. Journal of Computational and Applied Mathematics, 354 (??):191–197, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718307027.

Chen:2015:HOU

[CJLT15] S. Chen, X. Jiang, F. Liu, and I. Turner. High order unconditionally stable difference schemes for the Riesz spacefractional telegraph equation. Journal of Computational and Applied Mathematics, 278(??):119–129, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400435X.

Calvo-Jurado:2017:ITP

[CJP17] Carmen Calvo-Jurado and William J. Parnell. The influence of two-point statistics on the Hashin– Shtrikman bounds for three phase composites. Journal of Computational and Applied Mathematics, 318(??):354–365, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042716304150.

Cordero:2018:HEI

[CJSCT18] Alicia Cordero, Cristina Jordán, Esther Sanabria-Codesal, and Juan R. Torregrosa. Highly efficient iterative algorithms for solving nonlinear systems with arbitrary order of convergence p + 3, p ≥

5. Journal of Computational and Applied Mathematics, 330(??):748–758, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300973.

Cordero:2015:OPN

[CJT15] Alicia Cordero, Cristina Jordán, and Juan R. Torregrosa. One-point Newton-type iterative methods: a unified point of view. Journal of Computational and Applied Mathematics, 275(??):366–374, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003288.

Cordero:2017:DCB

[CJT17] Alicia Cordero, Cristina Jordán, and Juan R. Torregrosa. A dynamical comparison between iterative methods with memory: Are the derivatives good for the memory? Journal of Computational and Applied Mathematics, 318(??):335– 347, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304186.

Cortes:2017:MSS

[CJV17] J.-C. Cortés, L. Jódar, and L. Villafuerte. Mean square solution of Bessel differential equation with uncertainties. Journal of Computational and Applied Mathematics, 309(??):383–395, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300139.

Cardone:2015:OCG

[CJVW15]  Angelamaria  Cardone,

Zdzislaw Jackiewicz, James H. Verner, and Bruno Welfert. Order conditions for general linear methods. Journal of Computational and Applied Mathematics, 290(??):44–64, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500268X.

Choi:2016:FEM

[CK16] Hyung Jun Choi and Jae Ryong Kweon. A finite element method for singular solutions of the Navier–Stokes equations on a non-convex polygon. Journal of Computational and Applied Mathematics, 292(??):342–362, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003684.

Ciesielski:2017:FSL

[CKB17] Mariusz Ciesielski, Małgorzata Klimek, and Tomasz Blaszczyk. The fractional Sturm–Liouville problem — numerical approximation and application in fractional diffusion. Journal of Computational and Applied Mathematics, 317(??):573– 588, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306240.

Chen:2015:RDS

[CKKJ15] Jeng-Tzong Chen, Shyh-Rong Kuo, Shing-Kai Kao, and Jie Jian. Revisit of a degenerate scale: a semi-circular disc. Journal of Computational and Applied Mathematics, 283(??):182–200, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000230.

Choi:2018:RML

[CKL18] Heesun Choi, Hongjoong Kim, and Marc Laforest. Relaxation model for the p-Laplacian problem with stiffness. Journal of Computational and Applied Mathematics, 344 (??):173–189, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302930.

Cui:2015:SMB

[CL15] Kai Cui and Na Lei.

Stable monomial basis for multivariate Birkhoff interpolation problems. Journal of Computational and Applied Mathematics, 277(??):162–170, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003914.

Chung:2016:MGSb

[CL16] Eric T. Chung and Wing Tat Leung. Mixed GMsFEM for the simulation of waves in highly heterogeneous media. Journal of Computational and Applied Mathematics, 306(??):69–86, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301650.

Cao:2018:RSD

[CL18a] K. Cao and D. Lesnic. Reconstruction of the space-dependent perfusion coefficient from final time or time-average temperature measurements. Journal of Computational and Applied Mathematics, 337(??):150–165, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300311.

Chen:2018:ESB

[CL18b] Xiao Shan Chen and Peng Lv. On estimating the separation between (A, B) and (C, D) associated with the generalized Sylvester equation AXD − BCX = E. Journal of Computational and Applied Mathematics, 330(??):128–140, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730376X.

Celiker:2019:HAF

[CL19a] Emine Celiker and Ping Lin. A highly-accurate finite element method with exponentially compressed meshes for the solution of the Dirichlet problem of the generalized Helmholtz equation with corner singularities. Journal of Computational and Applied Mathematics, 361(??):227–235, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301840.

Chen:2019:MCD

[CL19b] Aimin Chen and Maojun Li. A modified central discontinuous Galerkin method with positivitypreserving and well-balanced properties for the onedimensional nonlinear shallow water equations. Journal of Computational and Applied Mathematics, 345: 374–387, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Chung:2019:MGM

[CL19c] Eric T. Chung and Chak Shing Lee. A mixed generalized multiscale finite element method for planar linear elasticity. Journal of Computational and Applied Mathematics, 348(??):298– 313, March 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718305417.

Chung:2019:AGM

[CL19d] Eric T. Chung and Yanbo Li. Adaptive generalized multiscale finite element methods for H (curl)- elliptic problems with heterogeneous coefficients. Journal of Computational and Applied Mathematics, 345:357–373, January 2019. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic).

Chen:2019:NMA

[CLA19] Ruige Chen, Fawang Liu, and Vo Anh. Numerical methods and analysis for a multi-term timespace variable-order fractional advection-diffusion equations and applications. Journal of Computational and Applied Mathematics, 352(??):437–452, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307581.

Chen:2016:SPA

[CLC16] Hu Chen, Shujuan Lü, and Wenping Chen. Spectral and pseudospectral approximations for the time fractional diffusion equation on an unbounded domain. Journal of Computational and Applied Mathematics, 304(??):43–56, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301303.

Cai:2018:UOE

[CLC18a] Wentao Cai, Jian Li, and Zhangxin Chen. Unconditional optimal error estimates for BDF2-FEM for a nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 331(??):23– 41, March 15, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717304272.

Chen:2018:UNS

[CLC18b] Hu Chen, Shujuan Lü, and Wenping Chen. A unified numerical scheme for the multi-term time fractional diffusion and diffusion-wave equations with variable coefficients. Journal of Computational and Applied Mathematics, 330(??):380–397, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042717304284.

Cossette:2015:NCR

[CLHMT15] Hélène Cossette, Etienne Larrivée-Hardy, Etienne Marceau, and Julien Trufin. A note on compound renewal risk models with dependence. Journal of Computational and Applied Mathematics, 285(??):295–311, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001028.

Cornejo:2019:SBM

[CLM19] M. Eugenia Cornejo, David Lobo, and Jesús Medina. On the solvability of bipolar max-product fuzzy relation equations with the product negation. Journal of Computational and Applied Mathematics, 354(??):520– 532, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307015.

Chung:2016:GOA

[CLP16] Eric T. Chung, Wing Tat Leung, and Sara Pollock. Goal-oriented adaptivity for GMsFEM. Journal of Computational and Applied Mathematics, 296 (??):625–637, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S037704271500521X.

Czarna:2017:JDP

[CLPZ17] Irmina Czarna, Yanhong Li, Zbigniew Palmowski, and Chunming Zhao. The joint distribution of the Parisian ruin time and the number of claims until Parisian ruin in the classical risk model. Journal of Computational and Applied Mathematics, 313(??):499–514, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304745.

Cabrera:2018:EPS

[CLS18] I. J. Cabrera, B. López, and K. Sadarangani. Existence of positive solutions for the nonlinear elastic beam equation via a mixed monotone operator. Journal of Computational and Applied Mathematics, 327(??):306–313, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302029.

Chung:2016:MGSa

[CLV16] Eric T. Chung, Wing Tat Leung, and Maria Vasilyeva. Mixed GMsFEM for second order elliptic problem in perforated domains. Journal of Computational and Applied Mathematics, 304(??):84–99, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300942.

Chung:2018:MMR

[CLVW18] Eric T. Chung, Wing Tat Leung, Maria Vasilyeva, and Yating Wang. Multiscale model reduction for transport and flow problems in perforated domains. Journal of Computational and Applied Mathematics, 330(??):519–535, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304065.

Chen:2017:SPB

[CLW17] Hao Chen, Xiaolin Li, and Yan Wang. A splitting preconditioner for a block two-by-two linear system with applications to the bidomain equations. Journal of Computational and Applied Mathematics, 321(??):487–498, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730136X.

Cen:2017:PUH

[CLX17] Zhongdi Cen, Anbo Le, and Aimin Xu. Parameteruniform hybrid difference scheme for solutions and derivatives in singularly perturbed initial value problems. Journal of Computational and Applied Mathematics, 320: 176–192, August 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Chang:2018:CPC

[CLZL18] Xiaokai Chang, Sanyang Liu, Pengjun Zhao, and Xu Li. Convergent prediction-correction-based ADMM for multi-block separable convex programming. Journal of Computational and Applied Mathematics, 335(??):270– 288, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305939.

Chang:2019:GLA

[CLZS19] Xiaokai Chang, Sanyang Liu, Pengjun Zhao, and Dunjiang Song. A generalization of linearized alternating direction method of multipliers for solving two-block separable convex programming. Journal of Computational and Applied Mathematics, 357(??):251–272, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301013.

Cheng:2019:IWM

[CLZW19] Ziqiang Cheng, Yuan Liu, Mengping Zhang, and Jin Wang. IB–WENO method for incompressible flow with elastic boundaries. Journal of Computational and Applied Mathematics, 362(??):498–509, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718306356.

Chadha:2016:OTS

[CM16] Naresh M. Chadha and Niall Madden. An optimal time-stepping algorithm for unsteady advection–diffusion problems. Journal of Computational and Applied Mathematics, 294(??):57–77, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500391X.

Chen:2018:GSS

[CM18a] Cai-Rong Chen and Chang-Feng Ma. A generalized shift-splitting preconditioner for complex symmetric linear systems. Journal of Computational and Applied Mathematics, 344(??):691–700, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303650.

Chen:2018:CFM

[CM18b] Xiao-Diao Chen and Weiyin Ma. A competition flow method for computing medial axis transform. Journal of Computational and Applied Mathematics, 340(??):342–359, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301225.

Cao:2019:EEM

[CM19a] Xiulei Cao and Koondanibha Mitra. Error estimates for a mixed finite element discretization of a two-phase porous media flow model with dynamic capillarity. Journal of Computational and Applied Mathematics, 353 (??):164–178, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718307532.

Charina:2019:MMS

[CM19b] Maria Charina and Thomas Mejstrik. Multiple multivariate subdivision schemes: Matrix and operator approaches. Journal of Computational and Applied Mathematics, 349(??):279– 291, March 15, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718304904.

Caudillo-Mata:2017:FUE

[CMHHS17] L. A. Caudillo-Mata, E. Haber, L. J. Heagy, and C. Schwarzbach. A framework for the upscaling of the electrical conductivity in the quasistatic Maxwell’s equations. Journal of Computational and Applied Mathematics, 317(??):388– 402, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305933.

Cerdan:2017:TLI

[CMM17] J. Cerdán, J. Marín, and J. Mas. A two-level ILU preconditioner for electromagnetic applications. Journal of Computational and Applied Mathematics, 309(??):371–382, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301327.

Cordero:2016:SSE

[CMQT16] Alicia Cordero, Alberto Magreñán, Carlos Quemada, and Juan R. Torregrosa. Stability study of eighth-order iterative methods for solving nonlinear equations. Journal of Computational and Applied Mathematics, 291(??):348–357, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000187.

Cornejo:2017:ASR

[CMRP17] M. Eugenia Cornejo, Jesús Medina, and Eloísa Ramírez-Poussa. Attribute and size reduction mechanisms in multi-adjoint concept lattices. Journal of Computational and Applied Mathematics, 318(??):388– 402, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303314.

Cordero:2017:SFO

[CMTV17] Alicia Cordero, Javier G.

Maimó, Juan R. Torregrosa, and Maria P. Vassileva. Stability of a fourth order bi-parametric family of iterative methods. Journal of Computational and Applied Mathematics, 312(??):94–102, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000248.

Costabile:2016:NSM

[CN16] F. Costabile and A. Napoli.

A new spectral method for a class of linear boundary value problems. Journal of Computational and Applied Mathematics, 292(??):329–341, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003775.

Chun:2018:CSM

[CN18] Changbum Chun and Beny Neta. Comparative study of methods of various orders for finding repeated roots of nonlinear equations. Journal of Computational and Applied Mathematics, 340(??):11–42, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300803.

Colomer:2018:EFD

[CNJM18] Adrián Colomer, Valery Naranjo, Thomas Janvier, and Jose M. Mossi. Evaluation of fractal dimension effectiveness for damage detection in retinal background. Journal of Computational and Applied Mathematics, 337(??):341–353, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300268.

Cortes:2018:SRC

[CNQR^{+}18] J. C. Cortés, A. Navarro-Quiles, J. V. Romero, M. D. Roselló, and M. A. Sohaly. Solving the random Cauchy one-dimensional advection-diffusion equation: Numerical analysis and computing. Journal of Computational and Applied Mathematics, 330 (??):920–936, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042717300596.

Cortes:2017:RPM

[CNQRR17] J. C. Cortés, A. Navarro-Quiles, J. V. Romero, and M. D. Roselló. Randomizing the parameters of a Markov chain to model the stroke disease: a technical generalization of established computational methodologies towards improving real applications. Journal of Computational and Applied Mathematics, 324(??):225–240, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730211X.

Cortes:2018:CPD

[CNQRR18] J.-C. Cortés, A. Navarro-Quiles, J.-V. Romero, and M.-D. Roselló. Computing the probability density function of nonautonomous first-order linear homogeneous differential equations with uncertainty. Journal of Computational and Applied Mathematics, 337(??):190–208, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830044X.

Ceballos:2017:MFU

[CNT17] M. Ceballos, J. Núñez, and Á. F. Tenorio. Minimal faithful upper-triangular matrix representations for solvable Lie algebras. Journal of Computational and Applied Mathematics, 318(??):279– 292, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304356.

Campos:2016:LBM

[COdSR16] J. O. Campos, R. S. Oliveira, R. W. dos Santos, and B. M. Rocha. Lattice Boltzmann method for parallel simulations of cardiac electrophysiology using GPUs. Journal of Computational and Applied Mathematics, 295(??):70–82, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0377042715000692.

Calderin-Ojeda:2019:MNB

[COGD19] Enrique Calderín-Ojeda and Emilio Gómez-Déniz. The multivariate negative binomial–Lindley distribution. Properties and new representation for the univariate case. Journal of Computational and Applied Mathematics, 347(??):36–48, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304679.

Colombeau:2016:NSS

[Col16] M. Colombeau. A numerical scheme for singular shock solutions and a study of its consistence in the sense of distributions. Journal of Computational and Applied Mathematics, 294(??):78–92, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003817.

Constantinescu:2018:GGE

[Con18] Emil M. Constantinescu.

Generalizing global error estimation for ordinary differential equations by using coupled timestepping methods. Journal of Computational and Applied Mathematics, 332 (??):140–158, April 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717302480.

Caliari:2017:SAM

[COP17] M. Caliari, A. Ostermann, and C. Piazzola. A splitting approach for the magnetic Schrödinger equation. Journal of Computational and Applied Mathematics, 316(??):74–85, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304101.

Court:2019:FDA

[Cou19] Sébastien Court. A fictitious domain approach for a mixed finite element method solving the two-phase Stokes problem with surface tension forces. Journal of Computational and Applied Mathematics, 359(??):30–54, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/

science/article/pii/
S0377042719301542.

Chan:2015:PHL

[CP15] Leunglung Chan and Eckhard Platen. Pricing and hedging of long dated variance swaps under a 3/2 volatility model. Journal of Computational and Applied Mathematics, 278(??):181–196, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004427.

Cresson:2016:NSF

[CP16] Jacky Cresson and Frédéric Pierret. Non standard finite difference scheme preserving dynamical properties. Journal of Computational and Applied Mathematics, 303(??):15–30, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300541.

Clempner:2018:TRP

[CP18] Julio B. Clempner and Alexander S. Poznyak. A Tikhonov regularized penalty function approach for solving polylinear programming problems. Journal of Computational and Applied Mathematics, 328(??):267–286, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303837.

Chadha:2019:FGA

[CPB19] Alka Chadha, Dwijendra N. Pandey, and Dhirendra Bahuguna. Faedo– Galerkin approximate solutions of a neutral stochastic fractional differential equation with finite delay. Journal of Computational and Applied Mathematics, 347(??):238–256, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303339.

Cui:2018:AMR

[CPL^{+}18] Angang Cui, Jigen Peng, Haiyang Li, Chengyi Zhang, and Yongchao Yu. Affine matrix rank minimization problem via nonconvex fraction function penalty. Journal of Computational and Applied Mathematics, 336(??):353– 374, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042718300165. See

corrigendum [CPL^{+}19b].

Cui:2019:ITA

[CPL^{+}19a] Angang Cui, Jigen Peng, Haiyang Li, Meng Wen, and Junxiong Jia. Iterative thresholding algorithm based on nonconvex method for modified l_{p}-norm regularization minimization. Journal of Computational and Applied Mathematics, 347(??):173–180, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304989.

Cui:2019:CAM

[CPL^{+}19b] Angang Cui, Jigen Peng, Haiyang Li, Chengyi Zhang, and Youngchao Yu. Corrigendum to “Affine matrix rank minimization problem via non-convex fraction function penalty” [J. Comput. Appl. Math. 336 (2018) 353–374]. Journal of Computational and Applied Mathematics, 352(??):478–485, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/

                                        S0377042718305879. See

[CPL^{+}18].

Colldeforns-Papiol:2018:CMR

[CPOG18] Gemma Colldeforns-Papiol and Luis Ortiz-Gracia. Computation of market risk measures with stochastic liquidity horizon. Journal of Computational and Applied Mathematics, 342(??):431–450, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302188.

Cotronei:2019:SSC

[CPRS19] Mariantonia Cotronei, Francesca Pelosi, Lucia Romani, and Alessandra Sestini. SMART 2017 — second conference on subdivision, geometric and algebraic methods, isogeometric analysis and refinability in Italy. Journal of Computational and Applied Mathematics, 349(??):207–211, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306782.

Castillo:2016:P

[CPS16] José Castillo, Víctor Pereyra, and Godela Scherer. Preface. Journal of Computational and Applied Mathematics, 295(??):1, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004872.

Chung:2019:ADL

[CPZ19] Eric T. Chung, Sai-Mang Pun, and Zhiwen Zhang. An adaptive dynamically low-dimensional approximation method for multiscale stochastic diffusion equations. Journal of Computational and Applied Mathematics, 356(??):302–313, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300688.

Chen:2017:FOS

[CQ17] Huangxin Chen and Weifeng Qiu. A first order system least squares method for the Helmholtz equation. Journal of Computational and Applied Mathematics, 309(??):145–162, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302916.

Chen:2016:PSD

[CQW16] Yannan Chen, Liqun Qi, and Qun Wang. Positive semi-definiteness and sum-of-squares property of fourth order four dimensional Hankel tensors. Journal of Computational and Applied Mathematics, 302(??):356–368, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300668.

Che:2017:IAC

[CQW17] Maolin Che, Liqun Qi, and Yimin Wei. Iterative algorithms for computing US- and Ueigenpairs of complex tensors. Journal of Computational and Applied Mathematics, 317(??):547– 564, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630632X.

Cano:2017:AOR

[CR17] B. Cano and N. Reguera. Avoiding order reduction when integrating nonlinear Schrödinger equation with Strang method. Journal of Computational and Applied Mathematics, 316(??):86–99, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304538.

Castro:2016:CDS

[CRCM16] M. A. Castro, F. Rodríguez, J. Cabrera, and J. A. Martín. A compact difference scheme for numerical solutions of second order dual-phase-lagging models of microscale heat transfer. Journal of Computational and Applied Mathematics, 291(??):432–440, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004774.

Chulian:2019:RSG

[CRG19] S. Chulián, M. Rosa, and M. L. Gandarias. Reductions and symmetries for a generalized Fisher equation with a diffusion term dependent on density and space. Journal of Computational and Applied Mathematics, 354(??):689– 698, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718306988.

Camacho:2017:CST

[CRGB17] J. C. Camacho, M. Rosa, M. L. Gandarias, and M. S. Bruzón. Classical symmetries, travelling wave solutions and conservation laws of a generalized Fornberg–Whitham equation. Journal of Computational and Applied Mathematics, 318(??):149– 155, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305507.

Croisille:2015:HAS

[Cro15] Jean-Pierre Croisille. Hermitian approximation of the spherical divergence on the cubed-sphere. Journal of Computational and Applied Mathematics, 280(??):188–201, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005263.

Cordova:2016:CFD

[CROC16] L. J. Córdova, O. Rojas, B. Otero, and J. Castillo. Compact finite difference modeling of 2-D acoustic wave propagation. Journal of Computational and Applied Mathematics, 295(??):83–91, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000618.

Cavalcanti:2016:NAW

[CRR16] V. N. Domingos Cavalcanti, J. H. Rodrigues, and C. Rosier. Numerical analysis for the wave equation with locally nonlinear distributed damping. Journal of Computational and Applied Mathematics, 301(??):144–160, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630005X.

Combarro:2019:QWD

[CRR19] Elías F. Combarro, José Ranilla, and I. F. Rúa. Quantum walks for the determination of commutativity of finite dimensional algebras. Journal of Computational and Applied Mathematics, 354(??):496– 506, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718306368.

Castro:2018:CHM

[CRS^{+}18] R. A. Castro, J. C. Rodríguez, W. W. Sierra, G. L. Di Giorgi, and S. J. Gómez. Chebyshev– Halley’s method on Riemannian manifolds. Journal of Computational and Applied Mathematics, 336 (??):30–53, July 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717306398.

Cotronei:2019:FAW

[CRSV19] Mariantonia Cotronei, Milvia Rossini, Tomas Sauer, and Elena Volontè. Filters for anisotropic wavelet decompositions. Journal of Computational and Applied Mathematics, 349(??):316–330, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305600.

Chovan:2017:IHS

[CS17] Jaroslav Chovan and Marián Slodicka. Induction hardening of steel with restrained Joule heating and nonlinear law for magnetic induction field: Solvability. Journal of Computational and Applied Mathematics, 311(??):630–644, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303764.

Chegini:2018:APT

[CS18a] Nabi Chegini and Rob Stevenson. Adaptive piecewise tensor product wavelets scheme for Laplace-interface problems. Journal of Computational and Applied Mathematics, 336(??):72– 97, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306416.

Chovan:2018:RTD

[CS18b] Jaroslav Chovan and Marián Slodicka. Reconstruction of a time dependent source term from a single boundary measurement in Maxwell’s equations with nonlinear generalized Ohm’s law. Journal of Computational and Applied Mathematics, 334(??):58–76, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305745.

Cui:2019:UPZ

[CS19] Lu-Bin Cui and Yisheng Song. On the uniqueness of the positive Z-eigenvector for nonnegative tensors. Journal of Computational and Applied Mathematics, 352(??):72–78, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307222.

CuevasNunez:2018:PDC

[CSB18] Rolando Cuevas Núñez, Christian E. Schaerer, and Amit Bhaya. A proportional-derivative control strategy for restarting the GMRES(m) algorithm. Journal of Computational and Applied Mathematics, 337(??):209–224, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830030X.

Clavijo:2019:RSC

[CSE^{+}19] S. P. Clavijo, A. F. Sarmiento, L. F. R. Espath, L. Dalcin, A. M. A.

Cortes, and V. M. Calo. Reactive n-species Cahn– Hilliard system: a thermodynamicallyconsistent model for reversible chemical reactions. Journal of Computational and Applied Mathematics, 350(??):143–154, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306083.

Chen:2019:TNS

[CSS19] Meng-Huo Chen, Amgad Salama, and Shuyu Sun. The transport of nanoparticles in subsurface with fractured, anisotropic porous media: Numerical simulations and parallelization. Journal of Computational and Applied Mathematics, 350(??):1– 18, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305648.

Chung:2016:GEA

[CST^{+}16] Yu-Min Chung, Andrew Steyer, Michael Tubbs, Erik S. Van Vleck, and Mihir Vedantam. Global error analysis and inertial manifold reduction. Journal of Computational and Applied Mathematics, 307(??):204–215, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630070X.

Cordero:2017:NSI

[CSTU17] Alicia Cordero, F. Soleymani, Juan R. Torregrosa, and M. Zaka Ullah. Numerically stable improved Chebyshev–Halley type schemes for matrix sign function. Journal of Computational and Applied Mathematics, 318(??):189– 198, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305143.

Chen:2019:HTP

[CSW19] Jie Chen, Shuyu Sun, and Xiaoping Wang. Homogenization of two-phase fluid flow in porous media via volume averaging. Journal of Computational and Applied Mathematics, 353(??):265– 282, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307544.

Chen:2019:SRA

[CSXY19] Yannan Chen, Wenyu Sun, Min Xi, and Jinyun Yuan. A seminorm regularized alternating least squares algorithm for canonical tensor decomposition. Journal of Computational and Applied Mathematics, 347(??):296–313, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303236.

Cordero:2015:LCR

[CT15] Alicia Cordero and Juan R.

Torregrosa. Low-complexity root-finding iteration functions with no derivatives of any order of convergence. Journal of Computational and Applied Mathematics, 275(??):502–515, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714000521.

Chen:2016:GDF

[CT16] Jianhua Chen and Xianhua Tang. Generalizations of Darbo’s fixed point theorem via simulation functions with application to functional integral equations. Journal of Computational and Applied Mathematics, 296(??):564–575, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005129.

Chidouh:2017:GLI

[CT17a] Amar Chidouh and Delfim F. M. Torres. A generalized Lyapunov’s inequality for a fractional boundary value problem. Journal of Computational and Applied Mathematics, 312(??):192–197, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301625.

Cordero:2017:SOI

[CT17b] Alicia Cordero and Juan R. Torregrosa. A sixth-order iterative method for approximating the polar decomposition of an arbitrary matrix. Journal of Computational and Applied Mathematics, 318 (??):591–598, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715006019.

Crisovan:2019:MOR

[CTAT19] R. Crisovan, D. Torlo, R. Abgrall, and S. Tokareva. Model order reduction for parametrized nonlinear hyperbolic problems as an application to uncertainty quantification. Journal of Computational and Applied Mathematics, 348(??):466–489, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305636.

Chavez:2018:PAC

[CTZK18] Gustavo Chávez, George Turkiyyah, Stefano Zampini, and David Keyes. Parallel accelerated cyclic reduction preconditioner for three-dimensional elliptic PDEs with variable coefficients. Journal of Computational and Applied Mathematics, 344 (??):760–781, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305952.

Cholakov:2017:CAF

[CV17] Slav I. Cholakov and Maria T. Vasileva. A convergence analysis of a fourth-order method for computing all zeros of a polynomial simultaneously. Journal of Computational and Applied Mathematics, 321(??):270–283, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301036.

Churbanov:2019:NSB

[CV19] Alexander G. Churbanov and Petr N. Vabishchevich. Numerical solution of boundary value problems for the eikonal equation in an anisotropic medium. Journal of Computational and Applied Mathematics, 362(??):55–67, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302511.

Chung:2017:CLM

[CVW17] Eric T. Chung, Maria Vasilyeva, and Yating Wang. A conservative local multiscale model reduction technique for Stokes flows in heterogeneous perforated domains. Journal of Computational and Applied Mathematics, 321(??):389–405, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301139.

Chen:2015:ASE

[CW15a] Lin Chen and Fuke Wu.

Almost sure exponential stability of the backward Euler–Maruyama scheme for stochastic delay differential equations with monotone-type condition. Journal of Computational and Applied Mathematics, 282(??):44– 53, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005998.

Chiu:2015:DCP

[CW15b] Mei Choi Chiu and Hoi Ying Wong. Dynamic cointegrated pairs trading: Mean-variance time-consistent strategies. Journal of Computational and Applied Mathematics, 290 (??):516–534, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003234.

Chen:2016:NSC

[CW16a] Huanzhen Chen and Hong Wang. Numerical simulation for conservative fractional diffusion equations by an expanded mixed formulation. Journal of Computational and Applied Mathematics, 296(??):480– 498, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004744.

Chen:2016:RSC

[CW16b] Suqin Chen and Yingwei Wang. A rational spectral collocation method for third-order singularly perturbed problems. Journal of Computational and Applied Mathematics, 307(??):93–105, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301224.

Congreve:2017:IGD

[CW17] Scott Congreve and Thomas P. Wihler. Iterative Galerkin discretizations for strongly monotone problems. Journal of Computational and Applied Mathematics, 311(??):457–472, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303740.

Cong:2018:WDD

[CWH18] Yuhao Cong, Yanpei Wang, and Guangda Hu. Weak delay-dependent stability of Runge–Kutta methods for differential systems with distributed delays. Journal of Computational and Applied Mathematics, 342(??):70–82, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301882.

Cao:2017:FVF

[CWXW17] Yin Cao, Bao Wang, Kelin Xia, and Guowei Wei. Finite volume formulation of the MIB method for elliptic interface problems. Journal of Computational and Applied Mathematics, 321(??):60–77, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300857.

Chen:2019:NSS

[CXW19] Yingzi Chen, Aiguo Xiao, and Wansheng Wang. Numerical solutions of SDEs with Markovian switching and jumps under non-Lipschitz conditions. Journal of Computational and Applied Mathematics, 360(??):41–54, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301608.

Chen:2019:SOA

[CXZ19] Hongbin Chen, Da Xu, and Jun Zhou. A secondorder accurate numerical method with graded meshes for an evolution equation with a weakly singular kernel. Journal of Computational and Applied Mathematics, 356(??):152–163, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300470.

Chung:2015:SGS

[CY15] Eric T. Chung and Tang Fei Yu. Staggered-grid spectral element methods for elastic wave simulations. Journal of Computational and Applied Mathematics, 285(??):132–150, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000710.

Chen:2018:DPI

[CY18a] Kuen-Suan Chen and Chun-Ming Yang. Developing a performance index with a Poisson process and an exponential distribution for operations management and continuous improvement. Journal of Computational and Applied Mathematics, 343(??):737–747, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302127.

Chen:2018:NEH

[CY18b] Ruyun Chen and Gang Yang. Numerical evaluation of highly oscillatory Bessel transforms. Journal of Computational and Applied Mathematics, 342(??):16–24, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301894.

Cao:2015:FDI

[CYY15a] Haitao Cao, Tao Yu, and Xingye Yue. Fully discrete IPDG–HMM for multiscale Richards equation of unsaturated flow in porous media. Journal of Computational and Applied Mathematics, 290 (??):352–369, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003064.

Chau:2015:FCM

[CYY15b] K. W. Chau, S. C. P. Yam, and H. Yang. Fouriercosine method for ruin probabilities. Journal of Computational and Applied Mathematics, 281 (??):94–106, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042714005706.

Chen:2016:CLP

[CYY^{+}16] Dewang Chen, Jiateng Yin, Shiying Yang, Lingxi Li, and Peter Pudney. Constraint local principal curve: Concept, algorithms and applications. Journal of Computational and Applied Mathematics, 298(??):222–235, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005956.

Chen:2015:EMI

[CZ15] Xianjin Chen and Jianxin Zhou. Estimate of Morse index of cooperative elliptic systems and its application to spatial vector solitons. Journal of Computational and Applied Mathematics, 281(??):169– 181, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005251.

Crnjaric-Zic:2016:UNA

[CZ16] Nelida Crnjarić-Zic. Upwind numerical approximations of a compressible 1D micropolar fluid flow. Journal of Computational and Applied Mathematics, 303(??):81–92, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300693.

Caliari:2017:QNM

[CZ17a] M. Caliari and S. Zuccher. Quasi-Newton minimization for the p(x)- Laplacian problem. Journal of Computational and Applied Mathematics, 309(??):122–131, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302989.

Chen:2017:CSF

[CZ17b] Yuyan Chen and Shuo Zhang. A conservative stable finite element method for Stokes flow and nearly incompressible linear elasticity on rectangular grid. Journal of Computational and Applied Mathematics, 323(??):53–70, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301668.

Caliari:2019:FBE

[CZ19a] M. Caliari and F. Zivcovich. On-the-fly backward error estimate for matrix exponential approximation by Taylor algorithm. Journal of Computational and Applied Mathematics, 346 (??):532–548, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304680.

Chen:2019:WGF

[CZ19b] Yanli Chen and Tie Zhang.

A weak Galerkin finite element method for Burgers’ equation. Journal of Computational and Applied Mathematics, 348(??):103– 119, March 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718305314.

Chen:2017:TLS

[CZLV17] Luoping Chen, Bin Zheng, Guang Lin, and Nikolaos Voulgarakis. A twolevel stochastic collocation method for semilinear elliptic equations with random coefficients. Journal of Computational and Applied Mathematics, 315 (??):195–207, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716305192.

Dehghan:2019:EEF

[DA19] Mehdi Dehghan and Mostafa Abbaszadeh. Error estimate of finite element/finite difference technique for solution of two-dimensional weakly singular integropartial differential equation with space and time fractional derivatives. Journal of Computational and Applied Mathematics, 356(??):314–328, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718307593.

Duque:2016:EGF

[DAAF16] José C. M. Duque, Rui M. P. Almeida, Stanislav N. Antontsev, and Jorge Ferreira. The Euler–Galerkin finite element method for a nonlocal coupled system of reaction–diffusion type. Journal of Computational and Applied Mathematics, 296(??):116–126, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004719.

deAlba:2019:REC

[dAFvdMR19] P. Díaz de Alba, L. Fermo, C. van der Mee, and G. Rodriguez. Recovering the electrical conductivity of the soil via a linear integral model. Journal of Computational and Applied Mathematics, 352(??):132–145, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307246.

Dehghan:2015:EEN

[DAM15a] Mehdi Dehghan, Mostafa Abbaszadeh, and Akbar Mohebbi. Error estimate for the numerical solution of fractional reaction-subdiffusion process based on a meshless method. Journal of Computational and Applied Mathematics, 280(??):14– 36, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004993.

Dehghan:2015:UIE

[DAM15b] Mehdi Dehghan, Mostafa Abbaszadeh, and Akbar Mohebbi. The use of interpolating elementfree Galerkin technique for solving 2D generalized Benjamin–Bona– Mahony–Burgers and regularized long-wave equations on non-rectangular domains with error estimate. Journal of Computational and Applied Mathematics, 286(??):211–231, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001582.

Dang:2017:MLD

[Dan17] Duy-Minh Dang. A multilevel dimension reduction Monte-Carlo method for jump-diffusion models. Journal of Computational and Applied Mathematics, 324(??):49–71, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301693.

Das:2015:CPP

[Das15] Pratibhamoy Das. Comparison of a priori and a posteriori meshes for singularly perturbed nonlinear parameterized problems. Journal of Computational and Applied Mathematics, 290(??):16–25, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002538.

Dassios:2018:PFS

[Das18] Ioannis K. Dassios. A practical formula of solutions for a family of linear non-autonomous fractional nabla difference equations. Journal of Computational and Applied Mathematics, 339 (??):317–328, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717304612.

Dosiyev:2015:OBM

[DB15] A. A. Dosiyev and S. C.

Buranay. One-block method for computing the generalized stress intensity factors for Laplace’s equation on a square with a slit and on an Lshaped domain. Journal of Computational and Applied Mathematics, 289(??):400–411, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005081.

deBustos:2015:AFP

[dBAAA15] I. Fernández de Bustos, J. Agirrebeitia, G. Ajuria, and R. Ansola. An alternative full-pivoting algorithm for the factorization of indefinite symmetric matrices. Journal of Computational and Applied Mathematics, 274(??):44–57, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003100.

Demay:2019:NSC

[DBdLdM^{+}19] Charles Demay, Christian Bourdarias, Benoît de Laage de Meux, Stéphane Gerbi, and Jean-Marc Hérard. Numerical simulation of a compressible two-layer model: a first attempt with an implicitexplicit splitting scheme. Journal of Computational and Applied Mathematics, 346(??):357–377, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303790.

deBustos:2018:ELA

[dBGMUA18] I. Fernández de Bustos, V. García-Marina, G. Urkullu, and M. Abasolo. An efficient LDU algorithm for the minimal least squares solution of linear systems. Journal of Computational and Applied Mathematics, 344(??):346–355, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0377042718303145.

Dai:2019:WPE

[DBX19] Xinjie Dai, Weiping Bu, and Aiguo Xiao. Wellposedness and EM approximations for non-Lipschitz stochastic fractional integro-differential equations. Journal of Computational and Applied Mathematics, 356 (??):377–390, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL https://www.sciencedirect.
com/science/article/
pii/S0377042719300639.

Dong:2018:SOT

[DCNY18] Hao Dong, Junzhi Cui, Yufeng Nie, and Zihao Yang. Second-order two-scale computational method for damped dynamic thermo-mechanical problems of quasi-periodic composite materials. Journal of Computational and Applied Mathematics, 343(??):575–601, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302796.

DellAccio:2016:CHB

[DD16a] F. Dell’Accio and F. Di Tommaso. Complete Hermite–Birkhoff interpolation on scattered data by combined Shepard operators. Journal of Computational and Applied Mathematics, 300(??):192– 206, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006287.

Dieci:2016:MVS

[DD16b] Luca Dieci and Fabio Difonzo. Minimum variation solutions for sliding vector fields on the intersection of two surfaces in R^{3}. Journal of Computational and Applied Mathematics, 292(??):732–745, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000965.

Duldul:2016:CWF

[DD16c] Bahar Uyar Düldül and Mustafa Düldül. Can we find Willmore-like method for the tangential intersection problems? Journal of Computational and Applied Mathematics, 302(??):301–311, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300632.

Dzunic:2017:GAC

[DD17] Jovana Dzunić and Ivan Damnjanović. General approach to constructing optimal multipoint families of iterative methods using Hermite’s rational interpolation. Journal of Computational and Applied Mathematics, 321(??):261–269, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301048.

Davies:2018:CMD

[DD18a] Katherine Davies and Anna Dembińska. Computing moments of discrete order statistics from nonidentical distributions. Journal of Computational and Applied Mathematics, 328(??):340–354, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303588.

Du:2018:ASF

[DD18b] Shuaishuai Du and Huoyuan Duan. Analysis of a stabilized finite element method for Stokes equations of velocity boundary condition and of pressure boundary condition. Journal of Computational and Applied Mathematics, 337(??):290–318, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300475.

DAbbicco:2016:MHG

[DDG^{+}16] Marcello D’Abbicco, Nicoletta Del Buono, Patrizia Gena, Marco Berardi, Giuseppe Calamita, and Luciano Lopez. A model for the hepatic glucose metabolism based on Hill and step functions. Journal of Computational and Applied Mathematics, 292(??):746–759, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000485.

DeMarchi:2015:CMC

[DDM15] S. De Marchi, F. Dell’Accio, and M. Mazza. On the constrained mock-Chebyshev least-squares. Journal of Computational and Applied Mathematics, 280(??):94–109, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005111.

DAmbrosio:2016:GNM

[DDP16] Raffaele D’Ambrosio, Giuseppe De Martino, and Beatrice Paternoster. General Nyström methods in Nordsieck form: Error analysis. Journal of Computational and Applied Mathematics, 292(??):694–702, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002678.

Dagnino:2019:SQI

[DDR19] C. Dagnino, A. Dallefrate, and S. Remogna. Spline quasi-interpolating projectors for the solution of nonlinear integral equations. Journal of Computational and Applied Mathematics, 354(??):360– 372, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304151.

DeAsmundis:2016:RBS

[DdSL16] Roberta De Asmundis, Daniela di Serafino, and Germana Landi. On the regularizing behavior of the SDA and SDC gradient methods in the solution of linear ill-posed problems. Journal of Computational and Applied Mathematics, 302(??):81–93, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271600011X.

DeMuse:2019:MTV

[DEY19] Ryan DeMuse, Terry Easlick, and Mei Yin. Mixing time of vertexweighted exponential random graphs. Journal of Computational and Applied Mathematics, 362 (??):443–459, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718304643.

Dargaville:2015:LSB

[DF15] S. Dargaville and T. W. Farrell. A least squares based finite volume method for the Cahn–Hilliard and Cahn–Hilliard-reaction equations. Journal of Computational and Applied Mathematics, 273(??):225–244, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002933.

DAmbra:2016:PGR

[DF16a] Pasqua D’Ambra and Salvatore Filippone. A parallel generalized relaxation method for highperformance image segmentation on GPUs. Journal of Computational and Applied Mathematics, 293(??):35–44, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500254X.

DiPersio:2016:GSA

[DF16b] Luca Di Persio and Matteo Frigo. Gibbs sampling approach to regime switching analysis of financial time series. Journal of Computational and Applied Mathematics, 300(??):43– 55, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006226.

deFrutos:2017:SMO

[dFG17] Javier de Frutos and Víctor Gatón. A spectral method for an Optimal Investment problem with transaction costs under Potential Utility. Journal of Computational and Applied Mathematics, 319(??):262–276, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300341.

deFrutos:2019:EHS

[dFG19] Javier de Frutos and Víctor Gatón. An extension of Heston’s SV model to stochastic interest rates. Journal of Computational and Applied Mathematics, 354(??):174– 182, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305557.

Domenech:2016:TGM

[DFGS16] Luís Domenech, Antonio Falcó, Victor García, and Fernando Sánchez. Towards a 2.5D geometric model in mold filling simulation. Journal of Computational and Applied Mathematics, 291(??):183–196, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042715001132.

deFreitas:2018:LCC

[dFLM18] Bruno R. de Freitas, Jaume Llibre, and Joao C. Medrado. Limit cycles of continuous and discontinuous piecewise-linear differential systems in R^{3}. Journal of Computational and Applied Mathematics, 338(??):311–323, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300578.

Deng:2019:TEM

[DFLM19] Shounian Deng, Weiyin Fei, Wei Liu, and Xuerong Mao. The truncated EM method for stochastic differential equations with Poisson jumps. Journal of Computational and Applied Mathematics, 355(??):232–257, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300366.

Dubot:2015:WMS

[DFRR15] Fabien Dubot, Yann Favennec, Benoit Rousseau, and Daniel R. Rousse. A wavelet multi-scale method for the inverse problem of diffuse optical tomography. Journal of Computational and Applied Mathematics, 289(??):267–281, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000357.

Du:2015:MDA

[DFS15] Kui Du, Graeme Fairweather, and Weiwei Sun. Matrix decomposition algorithms for arbitrary order C^{0} tensor product finite element systems. Journal of Computational and Applied Mathematics, 275(??):162–182, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003720.

Dhia:2018:HMM

[DFT18] Anne-Sophie Bonnet-Ben Dhia, Sonia Fliss, and Antoine Tonnoir. The halfspace matching method: a new method to solve scattering problems in infinite media. Journal of Computational and Applied Mathematics, 338(??):44–68, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300505.

Deif:2016:IRS

[DG16a] Sarah A. Deif and Said R.

Grace. Iterative refinement for a system of linear integro-differential equations of fractional type. Journal of Computational and Applied Mathematics, 294(??):138–150, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500415X.

Domma:2016:CGO

[DG16b] Filippo Domma and Sabrina Giordano. Concomitants of m-generalized order statistics from generalized Farlie–Gumbel– Morgenstern distribution family. Journal of Computational and Applied Mathematics, 294(??):413–435, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004434.

Deng:2015:FIR

[DGH15] Liang-Jian Deng, Huiqing Guo, and Ting-Zhu Huang.

A fast image recovery algorithm based on splitting deblurring and denoising. Journal of Computational and Applied Mathematics, 287(??):88–97, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001880.

Dieci:2016:P

[DGL16] Luca Dieci, Nicola Guglielmi, and Luciano Lopez. Preface. Journal of Computational and Applied Mathematics, 292(??):637, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004215.

Dominguez:2016:RBP

[DGM16a] Carolina Domínguez, Gabriel N. Gatica, and Antonio Márquez. A residual-based a posteriori error estimator for the plane linear elasticity problem with pure traction boundary conditions. Journal of Computational and Applied Mathematics, 292(??):486–504, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042715003829.

Durante:2016:MOT

[DGM16b] Fabrizio Durante, Stéphane Girard, and Gildas Mazo. Marshall–Olkin type copulas generated by a global shock. Journal of Computational and Applied Mathematics, 296(??):638–648, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005324.

Deng:2019:CLC

[DGM19] Quanling Deng, Victor Ginting, and Bradley McCaskill. Construction of locally conservative fluxes for high order continuous Galerkin finite element methods. Journal of Computational and Applied Mathematics, 359(??):166–181, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301803.

Duan:2019:ECA

[DGN19] Jiangtao Duan, Wei Gao, and Hon Keung Tony Ng. Efficient computational algorithm for optimal continuous experimental designs. Journal of Computational and Applied Mathematics, 350(??):98–113, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305983.

Dehesa:2015:CAH

[DGSM15] J. S. Dehesa, A. Guerrero, and P. Sánchez-Moreno. Complexity analysis of hypergeometric orthogonal polynomials. Journal of Computational and Applied Mathematics, 284(??):144–154, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003707.

Das:2017:SIS

[DHAM17] Anupam Das, Bipan Hazarika, Reza Arab, and M. Mursaleen. Solvability of the infinite system of integral equations in two variables in the sequence spaces c_{0} and ℓ_{1}. Journal of Computational and Applied Mathematics, 326(??):183–192, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042717302881.

Roldan-Lopez-de-Hierro:2015:CPT

[dHKdHMM15] Antonio-Francisco Roldán-López de Hierro, Erdal Karapinar, Concepción Roldán-López de Hierro, and Juan Martínez-Moreno. Coincidence point theorems on metric spaces via simulation functions. Journal of Computational and Applied Mathematics, 275(??):345–355, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003306.

Duran:2015:AHS

[DI15] Ahmet Duran and Burhaneddin Izgi. Application of the Heston stochastic volatility model for Borsa Istanbul using impression matrix norm. Journal of Computational and Applied Mathematics, 281 (??):126–134, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714005767.

Dieci:2019:BMS

[DI19] Luca Dieci and J. D. Walsh III. The boundary method for semi-discrete optimal transport partitions and Wasserstein distance computation. Journal of Computational and Applied Mathematics, 353(??):318– 344, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307659.

Diao:2018:CNL

[Dia18] Huai-An Diao. Condition numbers for a linear function of the solution of the linear least squares problem with equality constraints. Journal of Computational and Applied Mathematics, 344 (??):640–656, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303273.

Diaz-Infante:2016:CAS

[DIJ16] Saúl Díaz-Infante and Silvia Jerez. Convergence and asymptotic stability of the explicit Steklov method for stochastic differential equations. Journal of Computational and Applied Mathematics, 291(??):36–47, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500028X.

Diaz-Infante:2017:LSM

[DIJ17] S. Díaz-Infante and S. Jerez.

The linear Steklov method for SDEs with nonglobally Lipschitz coefficients: Strong convergence and simulation. Journal of Computational and Applied Mathematics, 309(??):408–423, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301844.

Ding:2016:GPA

[Din16]  Hengfei Ding.  Gen-

eral Padé approximation method for time-space fractional diffusion equation. Journal of Computational and Applied Mathematics, 299(??):221– 228, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006032.

Defez:2019:EAA

[DIP^{+}19] Emilio Defez, Javier Ibáñez, Jesús Peinado, Jorge Sastre, and Pedro Alonso-Jordá. An efficient and accurate algorithm for computing the matrix cosine based on new Hermite approximations. Journal of Computational and Applied Mathematics, 348(??):1–13, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830534X.

Defez:2018:NEA

[DIS^{+}18] Emilio Defez, Javier Ibáñez, Jorge Sastre, Jesús Peinado, and Pedro Alonso. A new efficient and accurate spline algorithm for the matrix exponential computation. Journal of Computational and Applied Mathematics, 337(??):354–365, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305897.

Dassios:2017:MMP

[DJAMJ17] Ioannis Dassios, Andrey P. Jivkov, Andrew Abu-Muharib, and Peter James. A mathematical model for plasticity and damage: a discrete calculus formulation. Journal of Computational and Applied Mathematics, 312(??):27–38, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004240.

Dagan:2015:SSF

[DJC15] A. Dagan, T. Jeans, and R. H. Cooper. Spurious solutions in a Fredholm integral equation of the second kind with a Cauchy kernel. Journal of Computational and Applied Mathematics, 290(??):278–289, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003076.

Dinh:2016:PAS

[DK16] Bui Van Dinh and Do Sang Kim. Projection algorithms for solving nonmonotone equilibrium problems in Hilbert space. Journal of Computational and Applied Mathematics, 302(??):106–117, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300346.

Dahmen:2018:AAP

[DKL^{+}18] Wolfgang Dahmen, Gitta Kutyniok, Wang-Q Lim, Christoph Schwab, and Gerrit Welper. Adaptive anisotropic Petrov– Galerkin methods for first order transport equations. Journal of Computational and Applied Mathematics, 340(??):191–220, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830102X.

Dick:2015:RFC

[DKLP15] Josef Dick, Peter Kritzer, Gunther Leobacher, and Friedrich Pillichshammer. A reduced fast componentby-component construction of lattice points for integration in weighted spaces with fast decreasing weights. Journal of Computational and Applied Mathematics, 276 (??):1–15, March 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042714003744.

Dabaghi:2019:WFE

[DKP^{+}19] F. Dabaghi, P. Krejčí, A. Petrov, J. Pousin, and Y. Renard. A weighted finite element mass redistribution method for dynamic contact problems. Journal of Computational and Applied Mathematics, 345:338–356, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Daveloose:2016:RQH

[DKV16] Catherine Daveloose, Asma Khedher, and Michèle Vanmaele. Robustness of quadratic hedging strategies in finance via Fourier transforms. Journal of Computational and Applied Mathematics, 296 (??):56–88, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715004574.

DeBonis:2016:MNM

[DL16a] M. C. De Bonis and C. Laurita. A modified Nyström method for integral equations with Mellin type kernels. Journal of Computational and Applied Mathematics, 296(??):512–527, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005099.

Dumitrescu:2016:RSD

[DL16b] Roxana Dumitrescu and Céline Labart. Reflected scheme for doubly reflected BSDEs with jumps and RCLL obstacles. Journal of Computational and Applied Mathematics, 296(??):827–839, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005518.

Ysern:2015:RCR

[dlCYC15] B. de la Calle Ysern and J. Mínguez Ceniceros. Rate of convergence of row sequences of multipoint Padé approximants. Journal of Computational and Applied Mathematics, 284(??):155–170, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004531.

delaHoz:2016:NST

[dlHV16] Francisco de la Hoz and Fernando Vadillo. Numerical simulations of timedependent partial differential equations. Journal of Computational and Applied Mathematics, 295(??):175–184, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004452.

Dong:2015:MHS

[DLHY15] Xiao Liang Dong, Hong Wei Liu, Yu Bo He, and Xi Mei Yang. A modified Hestenes–Stiefel conjugate gradient method with sufficient descent condition and conjugacy condition. Journal of Computational and Applied Mathematics, 281(??):239– 249, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005378.

Dai:2017:GPL

[DLLB17] Ping-Fan Dai, Ji-Cheng Li, Yao-Tang Li, and Jianchao Bai. A general preconditioner for linear complementarity problem with an Mmatrix. Journal of Computational and Applied Mathematics, 317(??):100– 112, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305763.

Dezhbord:2016:NEM

[DLM16a] Atta Dezhbord, Taher Lotfi, and Katayoun Mahdiani. A new efficient method for cases of the singular integral equation of the first kind. Journal of Computational and Applied Mathematics, 296(??):156–169, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004811.

Ding:2016:KSF

[DLM16b] Feng Ding, Ximei Liu, and Xingyun Ma. Kalman state filtering based least squares iterative parameter estimation for observer canonical state space systems using decomposition. Journal of Computational and Applied Mathematics, 301(??):135–143, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630022X.

Descombes:2017:TCA

[DLM17] Stéphane Descombes, Stéphane Lanteri, and Ludovic Moya. Temporal convergence analysis of a locally implicit discontinuous Galerkin time domain method for electromagnetic wave propagation in dispersive media. Journal of Computational and Applied Mathematics, 316(??):122–132, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304587.

Duan:2019:FOL

[DLM^{+}19] Huoyuan Duan, Wei Liu, Junhua Ma, Roger C. E. Tan, and Shangyou Zhang. A family of optimal Lagrange elements for Maxwell’s equations. Journal of Computational and Applied Mathematics, 358(??):241–265, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301475.

Deng:2019:CCR

[DLMZ19] Chongyang Deng, Yajuan Li, Xiaowen Mu, and Yi Zhao. Characteristic conic of rational bilinear map. Journal of Computational and Applied Mathematics, 346(??):277–283, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304321.

Deaconu:2017:ACM

[DLS17] Madalina Deaconu, Antoine Lejay, and Khaled Salhi. Approximation of CVaR minimization for hedging under exponential-Lévy models. Journal of Computational and Applied Mathematics, 326 (??):171–182, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302418.

DAlessio:2016:SDH

[DLW16a] S. J. D. D’Alessio, N. Leung, and J. W. L. Wan. Stability of differentially heated flow from a rotating sphere. Journal of Computational and Applied Mathematics, 291(??):209–224, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001788.

Dali:2016:LCN

[DLW16b] Béchir Dali, Chong Li, and Jinhua Wang. Local convergence of Newton’s method on the Heisenberg group. Journal of Computational and Applied Mathematics, 300(??):217– 232, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006378.

Debbouche:2018:EMF

[DLY18] Amar Debbouche, Carlos Lizama, and Xiao-Jun Yang. Editorial: Modern Fractional Dynamic Systems and Applications, MFDSA 2017: Special issue in honor of: R. P. Agarwal, D. Baleanu, J. J. Nieto, D. F. M. Torres and Y. Zhou. Highly cited researchers in the field. Journal of Computational and Applied Mathematics, 339(??):1– 2, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301377.

Deng:2016:EES

[DM16a] Chongyang Deng and Weiyin Ma. Efficient evaluation of subdivision schemes with polynomial reproduction property. Journal of Computational and Applied Mathematics, 294(??):403–412, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004604.

Dratman:2016:NMM

[DM16b] Ezequiel Dratman and Guillermo Matera. Newton’s method and a mesh-independence principle for certain semilinear boundary-value problems. Journal of Computational and Applied Mathematics, 292(??):188–212, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003532.

Dehghan:2017:EAM

[DM17a] Mehdi Dehghan and Vahid Mohammadi. Error analysis of method of lines (MOL) via generalized interpolating moving least squares (GIMLS) approximation. Journal of Computational and Applied Mathematics, 321 (??):540–554, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301164.

During:2017:HOA

[DM17b] Bertram Düring and James Miles. High-order ADI scheme for option pricing in stochastic volatility models. Journal of Computational and Applied Mathematics, 316(??):109– 121, May 15, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716304678.

Domenech:2016:MMP

[DMAS^{+}16] L. Domenech, F. J.

Muñoz-Almaraz, C. I. Serra, C. Soler, and N. Montes. A 3D mathematical model for planning ostectomy on longbone angular deformities. Journal of Computational and Applied Mathematics, 291(??):58–65, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002265.

Denia:2018:CPA

[DMCC^{+}18] F. D. Denia, J. Martínez-Casas, J. Carballeira, E. Nadal, and F. J. Fuenmayor. Computational performance of analytical methods for the acoustic modelling of automotive exhaust devices incorporating monoliths. Journal of Computational and Applied Mathematics, 330(??):995–1006, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301292.

Dabiri:2018:OVO

[DMM18] A. Dabiri, B. P. Moghaddam, and J. A. Tenreiro Machado. Optimal variable-order fractional PID controllers for dynamical systems. Journal of Computational and Applied Mathematics, 339 (??):40–48, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718301080.

Deepho:2017:CAH

[DMMSK17] Jitsupa Deepho, Juan Martínez-Moreno, Kanokwan Sitthithakerngkiet, and Poom Kumam. Convergence analysis of hybrid projection with Cesàro mean method for the split equilibrium and general system of finite variational inequalities. Journal of Computational and Applied Mathematics, 318(??):658– 673, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500504X.

DeMarchi:2019:FSR

[DMP19] S. De Marchi, A. Martínez, and E. Perracchione. Fast and stable rational RBFbased partition of unity interpolation. Journal of Computational and Applied Mathematics, 349(??):331–343, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304400.

deMatos:2015:IPS

[dMPF15] Vitor L. de Matos, Andy B.

Philpott, and Erlon C. Finardi. Improving the performance of stochastic dual dynamic programming. Journal of Computational and Applied Mathematics, 290 (??):196–208, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002794.

DelaCruzCabrera:2019:ADN

[DMR19] Omar De la Cruz Cabrera, Mona Matar, and Lothar Reichel. Analysis of directed networks via the matrix exponential. Journal of Computational and Applied Mathematics, 355(??):182–192, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271930024X.

Diwan:2019:CCF

[DMS19] Ganesh C. Diwan, Andrea Moiola, and Euan A. Spence. Can coercive formulations lead to fast and accurate solution of the Helmholtz equation? Journal of Computational and Applied Mathematics, 352(??):110–131, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307258.

Dey:2019:APT

[DNK19] Sanku Dey, Mazen Nassar, and Devendra Kumar. Alpha power transformed inverse Lindley distribution: a distribution with an upside-down bathtub-shaped hazard function. Journal of Computational and Applied Mathematics, 348(??):130– 145, March 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718302176.

Das:2015:DLS

[DNL15] Payel Das, Gnaneshwar Nelakanti, and Guangqing Long. Discrete Legendre spectral projection methods for Fredholm– Hammerstein integral equations. Journal of Computational and Applied Mathematics, 278(??):293–305, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004580. See

erratum [DNL16].

Das:2016:EDL

[DNL16] Payel Das, Gnaneshwar Nelakanti, and Guangqing Long. Erratum to: Discrete Legendre spectral projection methods for Fredholm–Hammerstein integral equations [J. Comput. Appl. Math. 278 (2015) 293–305]. Journal of Computational and Applied Mathematics, 292(??):634–636, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004124. See

[DNL15].

Debbouche:2017:ICM

[DNY17] Amar Debbouche, Juan J. Nieto, and Anatoly Yagola. The International Conference: Mathematical and Computational Modelling in Science and Technology. Journal of Computational and Applied Mathematics, 312(??):1, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303739.

Dong:2018:NIF

[Don18] C. Y. Dong. A new integral formula for the variation of matrix elastic energy of heterogeneous materials. Journal of Computational and Applied Mathematics, 343(??):635–642, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830267X.

Dosiyev:2019:DMF

[Dos19] Adiguzel A. Dosiyev. Difference method of fourth order accuracy for the Laplace equation with multilevel nonlocal conditions. Journal of Computational and Applied Mathematics, 354(??):587– 596, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302413.

DAmbrosio:2016:NSR

[DP16] Raffaele D’Ambrosio and Beatrice Paternoster. Numerical solution of reaction– diffusion systems of λ– ω type by trigonometrically fitted methods. Journal of Computational and Applied Mathematics, 294 (??):436–445, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042715004197.

Darania:2017:HOC

[DP17a] P. Darania and S. Pishbin. High-order collocation methods for nonlinear delay integral equation. Journal of Computational and Applied Mathematics, 326(??):284–295, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302716.

DeRossi:2017:PCA

[DP17b] Alessandra De Rossi and Emma Perracchione. Positive constrained approximation via RBF-based partition of unity method. Journal of Computational and Applied Mathematics, 319(??):338–351, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300456.

Dar:2019:PPE

[DP19a] S. A. Dar and R. B. Paris. A (p, ν)-extension of the Appell function F_{1}(·) and its properties. Journal of Computational and Applied Mathematics, 358(??):12–19, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301207.

Dar:2019:PQE

[DP19b] S. A. Dar and R. B. Paris. A (p, q)-extension of Srivastava’s triple hypergeometric function HB and its properties. Journal of Computational and Applied Mathematics, 348 (??):237–245, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042718305326.

During:2019:HOC

[DP19c] Bertram Düring and Alexander Pitkin. High-order compact finite difference scheme for option pricing in stochastic volatility jump models. Journal of Computational and Applied Mathematics, 355(??):201–217, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300597.

Domma:2015:EAM

[DPN15] Filippo Domma, Bozidar V. Popović, and Saralees Nadarajah. An extension of Azzalini’s method. Journal of Computational and Applied Mathematics, 278(??):37– 47, April 15, 2015. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0377042714004142.

David:2018:FDB

[DQIM18] S. A. David, D. D. Quintino, C. M. C. Inacio, and J. A. T. Machado. Fractional dynamic behavior in ethanol prices series. Journal of Computational and Applied Mathematics, 339(??):85– 93, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300281.

daRocha:2016:GMD

[dR16] Zélia da Rocha. A general method for deriving some semi-classical properties of perturbed second degree forms: the case of the Chebyshev form of second kind. Journal of Computational and Applied Mathematics, 296(??):677–689, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005166.

Djukic:2016:TGA

[DRS16] Dusan Lj. Djukić, Lothar Reichel, and Miodrag M. Spalević. Truncated generalized averaged Gauss quadrature rules. Journal of Computational and Applied Mathematics, 308 (??):408–418, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302874.

Djukic:2019:IGA

[DRST19] Dušan Lj. Djukić, Lothar Reichel, Miodrag M. Spalević, and Jelena D. Tomanović. Internality of generalized averaged Gaussian quadrature rules and truncated variants for modified Chebyshev measures of the second kind. Journal of Computational and Applied Mathematics, 345:70–85, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

DAgostino:2015:PIE

[DS15a] Daniele D’Agostino and Frank J. Seinstra. A parallel isosurface extraction component for visualization pipelines executing on GPU clusters. Journal of Computational and Applied Mathematics, 273(??):383–393, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042714002647.

Demichelis:2015:SEB

[DS15b] Vittoria Demichelis and Matteo Sciarra. Smoothness and error bounds of Martensen splines. Journal of Computational and Applied Mathematics, 278(??):90–100, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004348.

Deelstra:2017:MEO

[DS17a] Griselda Deelstra and Matthieu Simon. Multivariate European option pricing in a Markovmodulated Lévy framework. Journal of Computational and Applied Mathematics, 317(??):171– 187, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305829.

Dyn:2017:MVS

[DS17b] Nira Dyn and Nir Sharon. Manifold-valued subdivision schemes based on geodesic inductive averaging. Journal of Computational and Applied Mathematics, 311(??):54–67, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630320X.

DeLillo:2019:CPP

[DS19] T. K. DeLillo and S. Sahraei.

Computation of plane potential flow past multielement airfoils using conformal mapping, revisited. Journal of Computational and Applied Mathematics, 362(??):246–261, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718305181.

Dehghan:2015:THO

[DSA15] Mehdi Dehghan, Mansour Safarpoor, and Mostafa Abbaszadeh. Two highorder numerical algorithms for solving the multi-term time fractional diffusion-wave equations. Journal of Computational and Applied Mathematics, 290(??):174–195, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002630.

Dumitrescu:2016:MPD

[DSA16] Bogdan Dumitrescu, Bogdan C. Sicleru, and Florin Avram. Modeling probability densities with sums of exponentials via polynomial approximation. Journal of Computational and Applied Mathematics, 292(??):513–525, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003945.

daSilva:2016:NFD

[dSBV16] Allan Jonathan da Silva, Jack Baczynski, and José V. M. Vicente. A new finite difference method for pricing and hedging fixed income derivatives: Comparative analysis and the case of an Asian option. Journal of Computational and Applied Mathematics, 297(??):98–116, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271500535X.

Dhaene:2015:MEM

[DSDV15] Jan Dhaene, Ben Stassen, Pierre Devolder, and Michel Vellekoop. The minimal entropy martingale measure in a market of traded financial and actuarial risks. Journal of Computational and Applied Mathematics, 282(??):111– 133, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005457.

Dimitriu:2017:CNA

[DSN17] Gabriel Dimitriu, Razvan Stefanescu, and Ionel M. Navon. Comparative numerical analysis using reduced-order modeling strategies for nonlinear large-scale systems. Journal of Computational and Applied Mathematics, 310(??):32–43, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303144.

Denia:2016:PCS

[DSOBK16] F. D. Denia, E. M.

Sánchez-Orgaz, L. Baeza, and R. Kirby. Point collocation scheme in silencers with temperature gradient and mean flow. Journal of Computational and Applied Mathematics, 291(??):127–141, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000680.

Du:2015:ISS

[DSZ15] Lei Du, Tomohiro Sogabe, and Shao-Liang Zhang. IDR(s) for solving shifted nonsymmetric linear systems. Journal of Computational and Applied Mathematics, 274(??):35–43, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003112.

Duan:2016:EFR

[DT16] Shengli Duan and Hongjiong Tian. Exponential fitting Runge–Kutta methods for the delayed recruitment/renewal equation. Journal of Computational and Applied Mathematics, 303(??):31–39, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300644.

Dilloo:2017:SPD

[DT17] Mehzabeen Jumanah Dilloo and Désiré Yannick Tangman. A superconvergent partial differential equation approach to price variance swaps under regime switching models. Journal of Computational and Applied Mathematics, 318(??):316– 334, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304216.

Denuit:2018:CLR

[DT18] Michel Denuit and Julien Trufin. Collective loss reserving with two types of claims in motor third party liability insurance. Journal of Computational and Applied Mathematics, 335 (??):168–184, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717306052.

Diaz-Toca:2017:BID

[DTB17] Gema M. Diaz-Toca and Skander Belhaj. Blind image deconvolution through Bezoutians. Journal of Computational and Applied Mathematics, 315 (??):98–106, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0377042716305106.

Deepho:2017:NGI

[DTKP17] Jitsupa Deepho, Phatiphat Thounthong, Poom Kumam, and Supak Phiangsungnoen. A new general iterative scheme for split variational inclusion and fixed point problems of kstrict pseudo-contraction mappings with convergence analysis. Journal of Computational and Applied Mathematics, 318 (??):293–306, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716304290.

Dubeau:2016:EMF

[Dub16]  François Dubeau.  On

Euler–Maclaurin formula. Journal of Computational and Applied Mathematics, 296(??):649–660, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715005336.

Dunca:2017:OFE

[Dun17] Argus A. Dunca. On an optimal finite element scheme for the advection equation. Journal of Computational and Applied Mathematics, 311(??):522–528, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303983.

Dresse:2015:OPM

[DV15] Zoé Dresse and Walter Van Assche. Orthogonal polynomials for Minkowski’s question mark function. Journal of Computational and Applied Mathematics, 284(??):171–183, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400332X.

Doubova:2016:ETS

[DV16] Anna Doubova and Fernando Vadillo. Extinctiontime for stochastic population models. Journal of Computational and Applied Mathematics, 295(??):159–169, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714006001.

Das:2019:PUO

[DVA19] Pratibhamoy Das and Jesus Vigo-Aguiar. Parameter uniform optimal order numerical approximation of a class of singularly perturbed system of reaction diffusion problems involving a small perturbation parameter. Journal of Computational and Applied Mathematics, 354(??):533– 544, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305861.

DiazCortes:2018:PBD

[DVJ18] G. B. Diaz Cortes, C. Vuik, and J. D. Jansen. On POD-based deflation vectors for DPCG applied to porous media problems. Journal of Computational and Applied Mathematics, 330(??):193–213, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303400.

DeStaelen:2015:EAR

[DVS15] R. H. De Staelen, K. Van Bockstal, and M. Slodicka. Error analysis in the reconstruction of a convolution kernel in a semilinear parabolic problem with integral overdetermination. Journal of Computational and Applied Mathematics, 275(??):382–391, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003070.

Diao:2016:SCN

[DWQ16] Huai-An Diao, Yimin Wei, and Sanzheng Qiao. Structured condition numbers of structured Tikhonov regularization problem and their estimations. Journal of Computational and Applied Mathematics, 308 (??):276–300, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302576.

Deng:2018:NGM

[DWS^{+}18] Daiguo Deng, Hefeng Wu, Peng Sun, Ruomei Wang, Zhuo Shi, and Xiaonan Luo. A new geometric modeling approach for woven fabric based on Frenet frame and spiral equation. Journal of Computational and Applied Mathematics, 329(??):84–94, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302297.

Ding:2015:ESN

[DWXO15] Youzheng Ding, Zhongli Wei, Jiafa Xu, and Donal O’Regan. Extremal solutions for nonlinear fractional boundary value problems with p-Laplacian. Journal of Computational and Applied Mathematics, 288 (??):151–158, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715002113.

Deng:2019:RLO

[DXML19] Chongyang Deng, Huixia Xu, Weiyin Ma, and Yajuan Li. Repeated local operations and associated interpolation properties of dual 2n-point subdivision schemes. Journal of Computational and Applied Mathematics, 349(??):344–353, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305818.

Deng:2018:DSF

[DXZ18] Weishan Deng, Jin Xu, and Shan Zhao. On developing stable finite element methods for pseudo-time simulation of biomolecular electrostatics. Journal of Computational and Applied Mathematics, 330 (??):456–474, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042717304211.

Dong:2015:TAL

[DY15] Ning Dong and Bo Yu.

On the tripling algorithm for large-scale nonlinear matrix equations with low rank structure. Journal of Computational and Applied Mathematics, 288(??):18–32, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001892.

Donat:2019:NCB

[DY19] Rosa Donat and Dionisio F. Yáñez. A nonlinear Chaikin-based binary subdivision scheme. Journal of Computational and Applied Mathematics, 349(??):379–389, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718305624.

Dong:2019:SNL

[DYD19] Hua Dong, Chuancun Yin, and Hongshuai Dai. Spectrally negative Lévy risk model under Erlangized barrier strategy. Journal of Computational and Applied Mathematics, 351(??):101–116, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306654.

Dong:2018:HDG

[DYZ18] Haixia Dong, Wenjun Ying, and Jiwei Zhang. A hybridizable discontinuous Galerkin method for elliptic interface problems in the formulation of boundary integral equations. Journal of Computational and Applied Mathematics, 344(??):624–639, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303054.

Diao:2018:SCC

[DZ18] Huai-An Diao and Jian Zhao. On structured componentwise condition numbers for Hamiltonian eigenvalue problems. Journal of Computational and Applied Mathematics, 335(??):74– 85, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730599X.

Dzhumabaev:2016:OAS

[Dzh16] D. S. Dzhumabaev. On one approach to solve the linear boundary value problems for Fredholm integro-differential equations. Journal of Computational and Applied Mathematics, 294(??):342–357, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004446.

Dzhumabaev:2018:NGS

[Dzh18] Dulat S. Dzhumabaev.

New general solutions to linear Fredholm integrodifferential equations and their applications on solving the boundary value problems. Journal of Computational and Applied Mathematics, 327(??):79– 108, January 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0377042717303102.

Du:2016:SMC

[DZM16] Xuewu Du, Peng Zhang, and Wenya Ma. Some modified conjugate gradient methods for unconstrained optimization. Journal of Computational and Applied Mathematics, 305(??):92–114, October 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301686.

El-Amin:2018:DFM

[EAKS18] Mohamed F. El-Amin, Jisheng Kou, and Shuyu Sun. Discrete-fracture-model of multi-scale timesplitting two-phase flow including nanoparticles transport in fractured porous media. Journal of Computational and Applied Mathematics, 333 (??):327–349, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S037704271730568X.

Erfani:2019:SEF

[EBJS19] S. Erfani, E. Babolian, S. Javadi, and M. Shamsi. Stable evaluations of fractional derivative of the Müntz–Legendre polynomials and application to fractional differential equations. Journal of Computational and Applied Mathematics, 348 (??):70–88, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S037704271830178X.

Erdogan:2018:UAS

[EC18] Fevzi Erdogan and Zhongdi Cen. A uniformly almost second order convergent numerical method for singularly perturbed delay differential equations. Journal of Computational and Applied Mathematics, 333(??):382–394, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305770.

Erdogan:2019:ADF

[EDO19] Mahmut Sami Erdogan, Çetin Disibüyük, and Özlem Ege Oruç. An alternative distribution function estimation method using rational Bernstein polynomials. Journal of Computational and Applied Mathematics, 353 (??):232–242, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718307647.

Esmaili:2016:MSM

[EE16] Sakine Esmaili and M. R. Eslahchi. A modified spectral method for solving operator equations. Journal of Computational and Applied Mathematics, 292(??):105–135, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271500326X.

Eryilmaz:2017:CFT

[EG17] Serkan Eryilmaz and Omer L. Gebizlioğlu. Computing finite time non-ruin probability and some joint distributions in discrete time risk model with exchangeable claim occurrences. Journal of Computational and Applied Mathematics, 313(??):235–242, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304459.

Ekomasov:2017:RDK

[EGK17] Evgenii G. Ekomasov, Azamat M. Gumerov, and Roman V. Kudryavtsev. Resonance dynamics of kinks in the sine-Gordon model with impurity, external force and damping. Journal of Computational and Applied Mathematics, 312(??):198–208, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301868.

Eshkuvatov:2019:GSN

[EHTL19] Z. K. Eshkuvatov, Hameed Husam Hameed, B. M. Taib, and N. M. A. Nik Long. General 2×2 system of nonlinear integral equations and its approximate solution. Journal of Computational and Applied Mathematics, 361(??):528–546, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302201.

Ezquerro:2019:APS

[EHV19] J. A. Ezquerro and M. A.

Hernández-Verón. Auxiliary point on the semilocal convergence of Newton’s method. Journal of Computational and Applied Mathematics, 354(??):198– 212, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301614.

Ezquerro:2018:SPN

[EHVM18] J. A. Ezquerro, M. A. Hernández-Verón, and Á. A. Magreñán. Starting points for Newton’s method under a center Lipschitz condition for the second derivative. Journal of Computational and Applied Mathematics, 330(??):721–731, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306239.

Elliott:2015:CEA

[EJJ15] David Elliott, Barbara M. Johnston, and Peter R. Johnston. A complete error analysis for the evaluation of a twodimensional nearly singular boundary element integral. Journal of Computational and Applied Mathematics, 279(??):261–276, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004944.

Elliott:2016:MSD

[EJJ16] David Elliott, Barbara M.

Johnston, and Peter R. Johnston. The method of steepest descent for estimating quadrature errors. Journal of Computational and Applied Mathematics, 303(??):93–104, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300759.

Ebert:2019:CLP

[EK19] Adrian Ebert and Peter Kritzer. Constructing lattice points for numerical integration by a reduced fast successive coordinate search algorithm. Journal of Computational and Applied Mathematics, 351(??):77–100, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830654X.

Eggels:2017:ITC

[EKKT17] A. W. Eggels, R. P. J.

Kunnen, B. Koren, and A. S. Tijsseling. Infotaxis in a turbulent 3D channel flow. Journal of Computational and Applied Mathematics, 310(??):44–58, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302746.

ElFarouq:2017:DIC

[El 17] Naïma El Farouq. Deterministic impulse control problems: Two discrete approximations of the quasi-variational inequality. Journal of Computational and Applied Mathematics, 309(??):200–218, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630303X.

Ebrahimi:2019:CIP

[EL19] A. Ebrahimi and G. B. Loghmani. A composite iterative procedure with fast convergence rate for the progressive-iteration approximation of curves. Journal of Computational and Applied Mathematics, 359(??):1–15, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301505.

Elgindy:2019:HOE

[Elg19] Kareem T. Elgindy. A high-order embedded domain method combining a predictor–corrector– Fourier-continuation–Gram method with an integral Fourier pseudospectral collocation method for solving linear partial differential equations in complex domains. Journal of Computational and Applied Mathematics, 361(??):372–395, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301578.

Eddargani:2019:AHS

[ELL^{+}19] S. Eddargani, A. Lamnii, M. Lamnii, D. Sbibih, and A. Zidna. Algebraic hyperbolic spline quasiinterpolants and applications. Journal of Computational and Applied Mathematics, 347(??):196–209, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304953.

Elouafi:2019:CEB

[Elo19] Mohamed Elouafi. On computing the eigenvalues of band Toeplitz matrices. Journal of Computational and Applied Mathematics, 357(??):12–25, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S037704271930086X.

Ervin:2017:NDF

[ELRR17] V. J. Ervin, Hyesuk Lee, and J. Ruiz-Ramírez. Nonlinear Darcy fluid flow with deposition. Journal of Computational and Applied Mathematics, 309(??):79–94, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302977.

Elsheikh:2015:EBA

[Els15]  Atiyah Elsheikh.  An

equation-based algorithmic differentiation technique for differential algebraic equations. Journal of Computational and Applied Mathematics, 281 (??):135–151, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714005822.

Ehteram:2015:NMS

[EM15] Mohammad Ehteram and Amin Mahdavi Meymand. Numerical modeling of scour depth at side piers of the bridge. Journal of Computational and Applied Mathematics, 280(??):68–79, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005184.

Kolei:2018:ADC

[EP18] Salima El Kolei and Frédéric Patras. Analysis, detection and correction of misspecified discrete time state space models. Journal of Computational and Applied Mathematics, 333(??):200–214, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305216.

Escuin:2017:CIR

[EPC17] David Escuín, Lorena Polo, and David Ciprés. On the comparison of inventory replenishment policies with time-varying stochastic demand for the paper industry. Journal of Computational and Applied Mathematics, 309(??):424–434, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301546.

Erfani:2018:OCT

[ERPD18] Ali Erfani, Sara Rezaei, Mehdi Pourseifi, and Hesameldien Derili. Optimal control in teleoperation systems with time delay: a singular perturbation approach. Journal of Computational and Applied Mathematics, 338(??):168–184, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300554.

Ervin:2018:DMC

[ERR18] V. J. Ervin and J. Ruiz-Ramírez. A deposition model coupling Stokes’ and Darcy’s equations with nonlinear deposition. Journal of Computational and Applied Mathematics, 340(??):151–172, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042718301006.

Eryilmaz:2016:CMN

[Ery16] Serkan Eryilmaz. Compound Markov negative binomial distribution. Journal of Computational and Applied Mathematics, 292(??):1–6, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003489.

Eryilmaz:2019:SR

[Ery19] Serkan Eryilmaz. (k_{1}, k_{2}, …, k_{m})- out-of-n system and its reliability. Journal of Computational and Applied Mathematics, 346 (??):591–598, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830462X.

Dine:2018:AFV

[ES18] Houssein Nasser El Dine and Mazen Saad. Analysis of a finite volumefinite element method for Darcy–Brinkman twophase flows in porous media. Journal of Computational and Applied Mathematics, 337(??):51–72, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300050.

El-Sobky:2018:PMT

[ESAE18] B. El-Sobky and Y. Abo-Elnaga. A penalty method with trust-region mechanism for nonlinear bilevel optimization problem. Journal of Computational and Applied Mathematics, 340(??):360–374, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301304.

Epov:2019:FEM

[ESI^{+}19] M. I. Epov, E. P. Shurina, N. B. Itkina, A. Y. Kutishcheva, and S. I. Markov. Finite element modeling of a multiphysics poro-elastic problem in multiscale media. Journal of Computational and Applied Mathematics, 352(??):1– 22, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305260.

El-Sayed:2018:DSP

[ESSE18] A. M. A. El-Sayed, S. M.

Salman, and N. A. Elabd. On the dynamics of the singularly perturbed Mackey–Glass equation. Journal of Computational and Applied Mathematics, 344(??):154–160, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302802.

Eryilmaz:2015:SSR

[ET15] Serkan Eryilmaz and G. Yazgi Tütüncü. Stress strength reliability in the presence of fuzziness. Journal of Computational and Applied Mathematics, 282(??):262– 267, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000163.

Eryilmaz:2019:RES

[ET19] Serkan Eryilmaz and Mustafa Tekin. Reliability evaluation of a system under a mixed shock model. Journal of Computational and Applied Mathematics, 352(??):255–261, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307362.

Escande:2018:ADU

[EW18] Paul Escande and Pierre Weiss. Accelerating l^{1}–l^{2} deblurring using wavelet expansions of operators. Journal of Computational and Applied Mathematics, 343(??):373–396, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302668.

Easterling:2018:POH

[EWR18] David R. Easterling, Layne T. Watson, and Naren Ramakrishnan. Probabilityone homotopy methods for constrained clustering. Journal of Computational and Applied Mathematics, 343(??):602–618, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302309.

Fan:2015:SBL

[Fan15] Bin Fan. A smoothing Broyden-like method with a nonmonotone derivativefree line search for nonlinear complementarity problems. Journal of Computational and Applied Mathematics, 290 (??):641–655, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003441.

Fatemi:2016:NEC

[Fat16] Masoud Fatemi. A new efficient conjugate gradient method for unconstrained optimization. Journal of Computational and Applied Mathematics, 300 (??):207–216, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715006470.

Fernandez:2015:OLG

[FAVGC^{+}15] F. J. Fernández, L. J.

Alvarez-Vázquez, N. García-Chan, A. Martínez, and M. E. Vázquez-Méndez. Optimal location of green zones in metropolitan areas to control the urban heat island. Journal of Computational and Applied Mathematics, 289(??):412–425, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004695.

Fakharany:2016:SPI

[FCJ16] M. Fakharany, R. Company, and L. Jódar. Solving partial integrodifferential option pricing problems for a wide class of infinite activity Lévy processes. Journal of Computational and Applied Mathematics, 296(??):739–752, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005373.

Fu:2019:GMF

[FCM19] Shubin Fu, Eric Chung, and Tina Mai. Generalized multiscale finite element method for a strain-limiting nonlinear elasticity model. Journal of Computational and Applied Mathematics, 359(??):153–165, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301785.

Farina:2017:RWW

[FdGKZ17] Leandro Farina, Rômulo L. da Gama, Sergey Korotov, and Juliana S. Ziebell. Radiation of water waves by a submerged nearly circular plate. Journal of Computational and Applied Mathematics, 310(??):165–173, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301819.

Ferreira:2017:EDW

[FdOP17] José A. Ferreira, Paula de Oliveira, and Gonçalo Pena. On the exponential decay of waves with memory. Journal of Computational and Applied Mathematics, 318(??):460– 478, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300383.

Farnam:2016:ILM

[FE16] Arash Farnam and Reza Mahboobi Esfanjani. Improved linear matrix inequality approach to stability analysis of linear systems with interval time-varying delays. Journal of Computational and Applied Mathematics, 294(??):49–56, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715003933.

Fakharany:2018:NVT

[FEC18] M. Fakharany, V. N. Egorova, and R. Company. Numerical valuation of two-asset options under jump diffusion models using Gauss-Hermite quadrature. Journal of Computational and Applied Mathematics, 330 (??):822–834, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717301553.

Ferreira:2015:RSL

[Fer15] O. P. Ferreira. A robust semi-local convergence analysis of Newton’s method for cone inclusion problems in Banach spaces under affine invariant majorant condition. Journal of Computational and Applied Mathematics, 279(??):318–335, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004981.

Fu:2015:ARC

[FF15] Ke-Ang Fu and Xiaoyong Fu. Asymptotics for the random coefficient first-order autoregressive model with possibly heavy-tailed innovations. Journal of Computational and Applied Mathematics, 285(??):116–124, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000904.

Farkas:2017:IAM

[FFKP17] Z. Farkas, I. Faragó, Á. Kriston, and A. Pfrang. Improvement of accuracy of multi-scale models of Li-ion batteries by applying operator splitting techniques. Journal of Computational and Applied Mathematics, 310(??):59–79, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630156X.

Falcon:2017:CIA

[FFN^{+}17] O. J. Falcón, R. M. Falcón, J. Núñez, A. M. Pacheco, and M. T. Villar. Computation of isotopisms of algebras over finite fields by means of graph invariants. Journal of Computational and Applied Mathematics, 318(??):307– 315, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304228.

Farina:2016:AEN

[FFP16] Leandro Farina, Marcos R. S. Ferreira, and Victor Péron. The airfoil equation on near disjoint intervals: Approximate models and polynomial solutions. Journal of Computational and Applied Mathematics, 298(??):97–104, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005786.

Filipov:2019:RFD

[FGF19] Stefan M. Filipov, Ivan D. Gospodinov, and István Faragó. Replacing the finite difference methods for nonlinear two-point boundary value problems by successive application of the linear shooting method. Journal of Computational and Applied Mathematics, 358(??):46–60, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/

science/article/pii/
S0377042719301232.

Farouki:2019:RMT

[FGS19] Rida T. Farouki, Carlotta Giannelli, and Alessandra Sestini. Rational minimaltwist motions on curves with rotation-minimizing Euler–Rodrigues frames. Journal of Computational and Applied Mathematics, 352(??):240–254, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307374.

Farago:2016:SQA

[FH16] István Faragó and Róbert Horváth. On some qualitatively adequate discrete space-time models of epidemic propagation. Journal of Computational and Applied Mathematics, 293(??):45–54, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001831.

Farago:2018:QPS

[FH18] István Faragó and Róbert Horváth. Qualitative properties of some discrete models of disease propagation. Journal of Computational and Applied Mathematics, 340(??):486–500, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304557.

Falini:2019:TSM

[FJ19] Antonella Falini and Bert Jüttler. THB-splines multi-patch parameterization for multiply-connected planar domains via template segmentation. Journal of Computational and Applied Mathematics, 349(??):390–402, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304886.

Feng:2017:CAR

[FJD17] Runhuan Feng, Xiaochen Jing, and Jan Dhaene. Comonotonic approximations of risk measures for variable annuity guaranteed benefits with dynamic policyholder behavior. Journal of Computational and Applied Mathematics, 311(??):272–292, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S037704271630334X.

Firoozjaee:2018:NAF

[FJLB18] M. A. Firoozjaee, H. Jafari, A. Lia, and D. Baleanu. Numerical approach of Fokker–Planck equation with Caputo–Fabrizio fractional derivative using Ritz approximation. Journal of Computational and Applied Mathematics, 339 (??):367–373, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717302674.

Fu:2016:ARP

[FL16] Ke-Ang Fu and Huijie Li.

Asymptotic ruin probability of a renewal risk model with dependent byclaims and stochastic returns. Journal of Computational and Applied Mathematics, 306(??):154–165, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301704.

Fang:2023:CAF

[FL23] Qingxiang Fang and Xiaoping Liu. A comment on “Attractivity for functional Volterra integral equations of convolution type”. Journal of Computational and Applied Mathematics, 425(??):??, June 2023. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042723000031. See

[AL16].

Fan:2018:SNN

[FLJT18] W. Fan, F. Liu, X. Jiang, and I. Turner. Some novel numerical techniques for an inverse problem of the multi-term time fractional partial differential equation. Journal of Computational and Applied Mathematics, 336(??):114– 126, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300025.

Farmer:2017:IFR

[FLM17] Michael Farmer, George Loizou, and Stephen Maybank. Iteration functions re-visited. Journal of Computational and Applied Mathematics, 311(??):484–496, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042716303818.

Fraile:2016:DMS

[FLMS16] Alberto Fraile, Emilio Larrodé, Á. Alberto Magreñán, and Juan Antonio Sicilia. Decision model for siting transport and logistic facilities in urban environments: a methodological approach. Journal of Computational and Applied Mathematics, 291(??):478–487, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005688.

Feng:2016:DGF

[FLN16] Xiaobing Feng, Thomas Lewis, and Michael Neilan. Discontinuous Galerkin finite element differential calculus and applications to numerical solutions of linear and nonlinear partial differential equations. Journal of Computational and Applied Mathematics, 299(??):68– 91, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005348.

Feng:2019:AMI

[FLZ19] Hongsong Feng, Guangqing Long, and Shan Zhao. An augmented matched interface and boundary (MIB) method for solving elliptic interface problem. Journal of Computational and Applied Mathematics, 361(??):426–443, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719302389.

Fernandez:2017:AMP

[FM17] J. R. Fernández and M. Masid. Analysis of a model for the propagation of the ossification front. Journal of Computational and Applied Mathematics, 318(??):624– 633, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005427.

Fadahunsi:2018:RRP

[FM18a] A. I. Fadahunsi and R. M. Mnatsakanov. Recovery of ruin probability and Value at Risk from the scaled Laplace transform inversion. Journal of Computational and Applied Mathematics, 342(??):249–262, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302061.

Feckan:2018:AAP

[FM18b] Michal Feckan and Kateryna Marynets. Approximation approach to periodic BVP for mixed fractional differential systems. Journal of Computational and Applied Mathematics, 339(??):208–217, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305290.

Fallah:2019:EUS

[FM19a] Somayeh Fallah and Farshid Mehrdoust. On the existence and uniqueness of the solution to the double Heston model equation and valuing lookback option. Journal of Computational and Applied Mathematics, 350(??):412–422, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306538.

Fika:2019:FED

[FM19b] Paraskevi Fika and Marilena Mitrouli. Fast estimates for the diagonal of the inverse of large scale matrices appearing in applications. Journal of Computational and Applied Mathematics, 355(??):91–105, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300238.

Falco:2018:EPV

[FMC^{+}18] A. Falcó, N. Montés, F. Chinesta, L. Hilario, and M. C. Mora. On the existence of a progressive variational vademecum based on the proper generalized decomposition for a class of elliptic parameterized problems. Journal of Computational and Applied Mathematics, 330(??):1093–1107, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303965.

Fernandez:2019:ASG

[FMMQ19] J. R. Fernández, A. Magaña, M. Masid, and R. Quintanilla. Analysis for the strain gradient theory of porous thermoelasticity. Journal of Computational and Applied Mathematics, 345:247–268, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Fernandez:2015:SOP

[FMP^{+}15] Lidia Fernández, Francisco Marcellán, Teresa E. Pérez, Miguel A. Piñar, and Yuan Xu. Sobolev orthogonal polynomials on product domains. Journal of Computational and Applied Mathematics, 284(??):202–215, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004117.

Ford:2015:NCM

[FMR15] N. J. Ford, M. L. Morgado, and M. Rebelo. A nonpolynomial collocation method for fractional terminal value problems. Journal of Computational and Applied Mathematics, 275(??):392–402, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002866.

Ferreira:2015:ANS

[FNdO15] J. A. Ferreira, J. Naghipoor, and Paula de Oliveira. Analytical and numerical study of a coupled cardiovascular drug delivery model. Journal of Computational and Applied Mathematics, 275(??):433–446, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002209.

Falco:2017:DSP

[FNV17] A. Falcó, Ll. Navarro, and C. Vázquez. A direct LU solver for pricing American bond options under Hull–White model. Journal of Computational and Applied Mathematics, 309(??):442–455, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302205.

Fang:2018:MQN

[FNZ18] Xiaowei Fang, Qin Ni, and Meilan Zeng. A modified quasi-Newton method for nonlinear equations. Journal of Computational and Applied Mathematics, 328(??):44–58, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303242.

Franko:2015:RCS

[FOK15] C. Franko, M. Ozkut, and C. Kan. Reliability of coherent systems with a single cold standby component. Journal of Computational and Applied Mathematics, 281(??):230– 238, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005858.

Fidanova:2016:IHA

[FP16a] Stefka Fidanova and Petrica Pop. An improved hybrid ant-local search algorithm for the partition graph coloring problem. Journal of Computational and Applied Mathematics, 293(??):55–61, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002496.

Filbet:2016:HOT

[FP16b] Francis Filbet and Charles Prouveur. High order time discretization for backward semi-Lagrangian methods. Journal of Computational and Applied Mathematics, 303 (??):171–188, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300036.

Feng:2018:QIS

[FP18] Renzhong Feng and Shun Peng. Quasi-interpolation scheme for arbitrary dimensional scattered data approximation based on natural neighbors and RBF interpolation. Journal of Computational and Applied Mathematics, 329(??):95–105, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300912.

Fernandez:2019:ITC

[FPGK19] Arturo J. Fernández, Carlos J. Pérez-González, and Akram Kohansal. Improved time-censored reliability test plans for kout-of-n gamma systems. Journal of Computational and Applied Mathematics, 361(??):42–54, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302092.

Flores:2018:ELC

[FR18a] Julio Flores and Miguel Romance. On eigenvectorlike centralities for temporal networks: Discrete vs. continuous time scales. Journal of Computational and Applied Mathematics, 330(??):1041–1051, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302558.

Franco:2018:CEH

[FR18b] J. M. Franco and L. Rández.

A class of explicit highorder exponentially-fitted two-step methods for solving oscillatory IVPs. Journal of Computational and Applied Mathematics, 342(??):210–224, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302073.

Fung:2019:MMM

[FR19] Samy Wu Fung and Lars Ruthotto. A multiscale method for model order reduction in PDE parameter estimation. Journal of Computational and Applied Mathematics, 350 (??):19–34, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718305946.

Fernandez:2015:DVC

[FS15] J. R. Fernández and D. Santamarina. A dynamic viscoelastic contact problem with normal compliance. Journal of Computational and Applied Mathematics, 276 (??):30–46, March 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714003732.

Floater:2019:PLN

[FS19a] Michael S. Floater and Espen Sande. On periodic L^{2} n-widths. Journal of Computational and Applied Mathematics, 349(??):403–409, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305454.

Fu:2019:ECU

[FS19b] Guosheng Fu and Chi-Wang Shu. An energyconserving ultra-weak discontinuous Galerkin method for the generalized Korteweg– de Vries equation. Journal of Computational and Applied Mathematics, 349(??):41–51, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305661.

Fatahi:2016:NSM

[FSNS16] Hedayat Fatahi, Jafar Saberi-Nadjafi, and Elyas Shivanian. A new spectral meshless radial point interpolation (SMRPI) method for the two-dimensional Fredholm integral equations on general domains with error analysis. Journal of Computational and Applied Mathematics, 294(??):196–209, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004252.

Fukushima:2015:PFC

[Fuk15] Toshio Fukushima. Precise and fast computation of complete elliptic integrals by piecewise minimax rational function approximation. Journal of Computational and Applied Mathematics, 282(??):71– 76, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000023.

Ferro:2017:NSP

[FV17] Richard Ferro and Jani A. Virtanen. A note on structured pseudospectra of block matrices. Journal of Computational and Applied Mathematics, 322(??):18–24, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301395.

Fang:2018:GCP

[FW18] Lincong Fang and Guozhao Wang. Geometric characteristics of planar quintic Pythagorean-hodograph curves. Journal of Computational and Applied Mathematics, 330(??):117–127, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730403X.

Fan:2017:PNA

[FwC17] Hung-Yuan Fan and Eric King wah Chu. Projected nonsymmetric algebraic Riccati equations and refining estimates of invariant and deflating subspaces. Journal of Computational and Applied Mathematics, 315(??):70– 86, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305076.

Fu:2018:AFI

[FWW^{+}18] Qian Fu, Zhongke Wu, Xingce Wang, Mingquan Zhou, and Jianmin Zheng. An algorithm for finding intersection between ball B-spline curves. Journal of Computational and Applied Mathematics, 327(??):260–273, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303151.

Fu:2018:TDR

[FY18] Ke-Ang Fu and Chenglong Yu. On a two-dimensional risk model with timedependent claim sizes and risky investments. Journal of Computational and Applied Mathematics, 344(??):367–380, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303200.

Fang:2017:RTF

[FYY17] Yonglei Fang, Yanping Yang, and Xiong You. Revised trigonometrically fitted two-step hybrid methods with equation dependent coefficients for highly oscillatory problems. Journal of Computational and Applied Mathematics, 318(??):266– 278, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304368.

Fang:2015:GPM

[FZ15] Minglei Fang and Zhibin Zhu. A generalized project metric algorithm for mathematical programs with equilibrium constraints. Journal of Computational and Applied Mathematics, 288(??):99–115, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042715002174.

Fan:2019:SCF

[FZW19] Yifan Fan, Chuanxi Zhu, and Zhaoqi Wu. Some ϕcoupled fixed point results via modified F -control function’s concept in metric spaces and its applications. Journal of Computational and Applied Mathematics, 349(??):70–81, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305582.

Gouyandeh:2016:NSN

[GAA16] Z. Gouyandeh, T. Allahviranloo, and A. Armand. Numerical solution of nonlinear Volterra– Fredholm–Hammerstein integral equations via Tau-collocation method with convergence analysis. Journal of Computational and Applied Mathematics, 308(??):435–446, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303004.

Gomez-Aguilar:2019:PEL

[GAA19] J. F. Gómez-Aguilar and Abdon Atangana. Power and exponentials laws: Theory and application. Journal of Computational and Applied Mathematics, 354(??):52–65, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S037704271930010X.

See retraction notice [GAA21].

Gomez-Aguilar:2021:RNP

[GAA21] J. F. Gómez-Aguilar and Abdon Atangana. Retraction notice to “Power and exponentials laws: Theory and application” [j. comput. appl. math. 354 (2019) 52–65]. Journal of Computational and Applied Mathematics, 394(??):??, October 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042721002314. See

[GAA19].

Goodarzi:2017:SRU

[GAB17] F. Goodarzi, M. Amini, and G. R. Mohtashami Borzadaran. Some results on upper bounds for the variance of functions of the residual life random variables. Journal of Computational and Applied Mathematics, 320:

30–42, August 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Ghazouani:2016:HUP

[GB16] Sami Ghazouani and Fethi Bouzeffour. Heisenberg uncertainty principle for a fractional power of the Dunkl transform on the real line. Journal of Computational and Applied Mathematics, 294(??):151–176, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003350.

Guterding:2018:HSV

[GB18] Daniel Guterding and Wolfram Boenkost. The Heston stochastic volatility model with piecewise constant parameters — efficient calibration and pricing of window barrier options. Journal of Computational and Applied Mathematics, 343(??):353–362, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302498.

Gao:2015:PBA

[GC15] Longfei Gao and Victor M. Calo. Preconditioners based on the Alternating-Direction-Implicit algorithm for the 2D steady-state diffusion equation with orthotropic heterogeneous coefficients. Journal of Computational and Applied Mathematics, 273(??):274–295, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002945.

Goessens:2017:CTT

[GC17] Tineke Goessens and Denis Constales. Characteristic times in a three scale model with overlapping domain decomposition. Journal of Computational and Applied Mathematics, 318(??):529– 537, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300103.

Gnandt:2018:CAH

[GC18] Christian Gnandt and Rainer Callies. CGRS — an advanced hybrid method for global optimization of continuous functions closely coupling extended random search and conjugate gradient method. Journal of Computational and Applied Mathematics, 333(??):99– 115, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305198.

Galindo:2016:ENS

[GCTGC16] J. Galindo, H. Climent, A. Tiseira, and L. M. García-Cuevas. Effect of the numerical scheme resolution on quasi-2d simulation of an automotive radial turbine under highly pulsating flow. Journal of Computational and Applied Mathematics, 291(??):112–126, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000953.

Ge:2018:HMM

[GCY18] Liang Ge, Yanzhen Chang, and Danping Yang. Heterogeneous multiscale method for optimal control problem governed by parabolic equation with highly oscillatory coefficients. Journal of Computational and Applied Mathematics, 328(??):116–131, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303291.

Gao:2017:WGF

[GCZ17] Fuzheng Gao, Jintao Cui, and Guoqun Zhao. Weak Galerkin finite element methods for Sobolev equation. Journal of Computational and Applied Mathematics, 317(??):188– 202, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305891.

Gbari:2016:SAC

[GD16] Samuel Gbari and Michel Denuit. Stochastic approximations in CBD mortality projection models. Journal of Computational and Applied Mathematics, 296(??):102–115, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004720.

Gong:2017:IBEb

[GD17] Y. P. Gong and C. Y. Dong. An isogeometric boundary element method using adaptive integral method for 3D potential problems. Journal of Computational and Applied Mathematics, 319(??):141–158, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730002X.

Gnang:2018:RHI

[GD18] Calvin Gnang and François Dubeau. On the rediscovery of Halley’s iterative method for computing the zero of an analytic function. Journal of Computational and Applied Mathematics, 335(??):129– 141, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306015.

Galizia:2015:MCL

[GDC15a] Antonella Galizia, Daniele D’Agostino, and Andrea Clematis. An MPI–CUDA library for image processing on HPC architectures. Journal of Computational and Applied Mathematics, 273(??):414–427, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002374.

Goessens:2015:CTM

[GDC15b] Tineke Goessens, Rob H.

De Staelen, and Denis Constales. Characteristic times for multiscale diffusion of active ingredients in coated textiles. Journal of Computational and Applied Mathematics, 289(??):426–432, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004610.

Gao:2019:SSF

[GDCK19] Longfei Gao, David C. Del Rey Fernández, Mark Carpenter, and David Keyes. SBP–SAT finite difference discretization of acoustic wave equations on staggered block-wise uniform grids. Journal of Computational and Applied Mathematics, 348(??):421–444, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305272.

Guermond:2017:CAD

[GdLT17] Jean-Luc Guermond, Manuel Quezada de Luna, and Travis Thompson. An conservative anti-diffusion technique for the level set method. Journal of Computational and Applied Mathematics, 321 (??):448–468, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730081X.

Gallegos:2016:BCF

[GDM16] Javier A. Gallegos and Manuel A. Duarte-Mermoud. Boundedness and convergence on fractional order systems. Journal of Computational and Applied Mathematics, 296(??):815– 826, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005531.

Gong:2017:IBEa

[GDQ17] Y. P. Gong, C. Y. Dong, and X. C. Qin. An isogeometric boundary element method for three dimensional potential problems. Journal of Computational and Applied Mathematics, 313(??):454–468, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304824.

Grylonakis:2019:ISS

[GFPGF19] E.-N. G. Grylonakis, C. K. Filelis-Papadopoulos, G. A. Gravvanis, and A. S. Fokas. An iterative spatial-stepping numerical method for linear elliptic PDEs using the unified transform. Journal of Computational and Applied Mathematics, 352(??):194–209, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307155.

Guan:2015:SFE

[GG15] Qingguang Guan and Max Gunzburger. θ schemes for finite element discretization of the space-time fractional diffusion equations. Journal of Computational and Applied Mathematics, 288(??):264–273, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002514.

Gockler:2017:ACI

[GG17a] T. Göckler and V. Grimm.

Acceleration of contour integration techniques by rational Krylov subspace methods. Journal of Computational and Applied Mathematics, 316(??):133– 142, May 15, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716304095.

Guan:2017:AAN

[GG17b] Qingguang Guan and Max Gunzburger. Analysis and approximation of a nonlocal obstacle problem. Journal of Computational and Applied Mathematics, 313(??):102–118, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304320.

Garcia-Gutierrez:2017:ACI

[GGBTP17] J. B. García-Gutiérrez, F. Benítez-Trujillo, and C. Pérez. An algorithm to compute invariant sets for third-order switched systems. Journal of Computational and Applied Mathematics, 318(??):348– 353, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716304174.

Guevel:2015:NCH

[GGC15] Y. Guevel, G. Girault, and J. M. Cadou. Numerical comparisons of highorder nonlinear solvers for the transient Navier– Stokes equations based on homotopy and perturbation techniques. Journal of Computational and Applied Mathematics, 289(??):356–370, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005500.

Garcia:2018:PFP

[GGC^{+}18] D. Garcia, M. Ghommem, N. Collier, B. O. N. Varga, and V. M. Calo. PyFly: a fast, portable aerodynamics simulator. Journal of Computational and Applied Mathematics, 344(??):875–903, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301237.

Garralda-Guillem:2019:MAS

[GGG19] A. I. Garralda-Guillem and M. Ruiz Galán. A minimax approach for the study of systems of variational equations and related Galerkin schemes. Journal of Computational and Applied Mathematics, 354(??):103– 111, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302772.

Graf:2018:LTS

[GGK18] Jonathan S. Graf, Matthias K.

Gobbert, and Samuel Khuvis. Long-time simulations with complex code using multiple nodes of Intel Xeon Phi Knights Landing. Journal of Computational and Applied Mathematics, 337(??):18–36, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300189.

Guillen-Gonzalez:2017:OFO

[GGRN17] F. Guillén-González and M. V. Redondo-Neble. Optimal first-order error estimates of a fully segregated scheme for the Navier–Stokes equations. Journal of Computational and Applied Mathematics, 321(??):348–370, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300900.

Gunther:2016:SSS

[GGW16] Stefanie Günther, Nicolas R. Gauger, and Qiqi Wang. Simultaneous single-step one-shot optimization with unsteady PDEs. Journal of Computational and Applied Mathematics, 294(??):12–22, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004057.

Ganesh:2016:OOA

[GH16a] M. Ganesh and S. C. Hawkins. An offline/online algorithm for a class of stochastic multiple obstacle scattering configurations in the half-plane. Journal of Computational and Applied Mathematics, 307(??):52–64, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301996.

Georgieva:2016:NRR

[GH16b] I. Georgieva and C. Hofreither. New results on regularity and errors of harmonic interpolation using Radon projections. Journal of Computational and Applied Mathematics, 293(??):73–81, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001430.

Guglielmi:2016:PRL

[GH16c] Nicola Guglielmi and Ernst Hairer. Pathregularization of linear neutral delay differential equations with several delays. Journal of Computational and Applied Mathematics, 292(??):785–794, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005846.

Georgieva:2017:ALR

[GH17a] I. Georgieva and C. Hofreither. An algorithm for low-rank approximation of bivariate functions using splines. Journal of Computational and Applied Mathematics, 310(??):80–91, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716301509.

Gupta:2017:WCM

[GH17b] Ramesh C. Gupta and Jie Huang. The Weibull– Conway–Maxwell–Poisson distribution to analyze survival data. Journal of Computational and Applied Mathematics, 311(??):171–182, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303077.

Ganesh:2019:FHO

[GH19a] M. Ganesh and S. C. Hawkins. A fast high order algorithm for multiple scattering from large sound-hard three dimensional configurations. Journal of Computational and Applied Mathematics, 362 (??):324–340, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306617.

Georgieva:2019:GLR

[GH19b] I. Georgieva and C. Hofreither. Greedy low-rank approximation in Tucker format of solutions of tensor linear systems.

Journal of Computational and Applied Mathematics, 358(??):206–220, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301219.

Gaudreau:2015:ICN

[GHA^{+}15] P. Gaudreau, K. Hayami, Y. Aoki, H. Safouhi, and A. Konagaya. Improvements to the cluster Newton method for underdetermined inverse problems. Journal of Computational and Applied Mathematics, 283(??):122–141, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000266.

Ghayebi:2017:NSB

[GHB17] Bakhtiyar Ghayebi, S. Mohammad Hosseini, and Dirk Blömker. Numerical solution of the Burgers equation with Neumann boundary noise. Journal of Computational and Applied Mathematics, 311(??):148–164, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S037704271630317X.

Gu:2016:BNE

[GHC16] Xian-Ming Gu, Ting-Zhu Huang, and Bruno Carpentieri. BiCGCR2: a new extension of conjugate residual method for solving non-Hermitian linear systems. Journal of Computational and Applied Mathematics, 305(??):115–128, October 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301595.

Gui:2018:FEM

[GHL18] Wenyong Gui, Rongtan Huang, and X. Sheldon Lin. Fitting the Erlang mixture model to data via a GEM-CMM algorithm. Journal of Computational and Applied Mathematics, 343(??):189–205, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302140.

Guo:2019:PEA

[GHL19] Yuling Guo, Jianguo Huang, and Junjiang Lai. A posteriori error analysis of a fully-discrete finite element method for the wave equation. Journal of Computational and Applied Mathematics, 356(??):231–247, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300627.

Geiser:2017:NVI

[GHM17] Jürgen Geiser, José L.

Hueso, and Eulalia Martínez. New versions of iterative splitting methods for the momentum equation. Journal of Computational and Applied Mathematics, 309(??):359–370, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302722.

Gander:2018:MAH

[GHM18] Martin J. Gander, Laurence Halpern, and Véronique Martin. Multiscale analysis of heterogeneous domain decomposition methods for time-dependent advection–reaction–diffusion problems. Journal of Computational and Applied Mathematics, 344 (??):904–924, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303352.

Gutierrez:2019:ACN

[GHV19] J. M. Gutiérrez and M.Á. Hernández-Verón. An acceleration of the continuous Newton’s method. Journal of Computational and Applied Mathematics, 354(??):213–220, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718301596.

Gu:2018:RHM

[GHY^{+}18] Xian-Ming Gu, Ting-Zhu Huang, Guojian Yin, Bruno Carpentieri, Chun Wen, and Lei Du. Restarted Hessenberg method for solving shifted nonsymmetric linear systems. Journal of Computational and Applied Mathematics, 331(??):166–177, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304880.

Gu:2015:STP

[GHZ^{+}15] Xian-Ming Gu, Ting-Zhu Huang, Xi-Le Zhao, Hou-Biao Li, and Liang Li. Strang-type preconditioners for solving fractional diffusion equations by boundary value methods. Journal of Computational and Applied Mathematics, 277(??):73–86, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003689.

Giani:2018:RAA

[Gia18]  Stefano Giani.  Reli-

able anisotropic-adaptive discontinuous Galerkin method for simplified PN approximations of radiative transfer. Journal of Computational and Applied Mathematics, 337(??):225–243, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300074.

Guo:2017:NSL

[GJ17] Jingyang Guo and Jae-Hun Jung. A numerical study of the local monotone polynomial edge detection for the hybrid WENO method. Journal of Computational and Applied Mathematics, 321(??):232–245, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300948.

Gu:2018:GPA

[GJSC18] Chuanqing Gu, Xianglong Jiang, Chenchen Shao, and Zhibing Chen. A GMRES-Power algorithm for computing PageRank problems. Journal of Computational and Applied Mathematics, 343(??):113–123, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301638.

Gorb:2018:HDD

[GK18a] Yuliya Gorb and Daria Kurzanova. Heterogeneous domain decomposition method for high contrast dense composites. Journal of Computational and Applied Mathematics, 337(??):135–149, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300293.

Groselj:2018:BSB

[GK18b] Jan Groselj and Marjeta Knez. A B-spline basis for C^{1} quadratic splines on triangulations with a 10-split. Journal of Computational and Applied Mathematics, 343(??):413–427, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302644.

Groselj:2019:ICQ

[GK19a] Jan Groselj and Marjeta Knez. Interpolation with C^{2} quartic macro-elements based on 10-splits. Journal of Computational and Applied Mathematics, 362 (??):143–160, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302651.

Guo:2019:PDE

[GK19b] Huijun Guo and Junke Kou. Pointwise density estimation for biased sample. Journal of Computational and Applied Mathematics, 361(??):444–458, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271930189X.

Garrido:2017:LSE

[GKBG17] T. M. Garrido, A. A. Kasatkin, M. S. Bruzón, and R. K. Gazizov. Lie symmetries and equivalence transformations for the Barenblatt–Gilman model. Journal of Computational and Applied Mathematics, 318(??):253– 258, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304435.

Griewank:2018:HDI

[GKLS18] Andreas Griewank, Frances Y. Kuo, Hernan Leövey, and Ian H. Sloan. High dimensional integration of kinks and jumps — smoothing by preintegration. Journal of Computational and Applied Mathematics, 344(??):259–274, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301778.

Geum:2018:SDM

[GKM18] Young Hee Geum, Young Ik Kim, and Á. Alberto Magreñán. A study of dynamics via Möbius conjugacy map on a family of sixth-order modified Newton-like multiple-zero finders with bivariate polynomial weight functions. Journal of Computational and Applied Mathematics, 344(??):608–623, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303583.

Grammont:2016:MPM

[GKN16] Laurence Grammont, Rekha P.

Kulkarni, and T. J. Nidhin. Modified projection method for Urysohn integral equations with nonsmooth kernels. Journal of Computational and Applied Mathematics, 294 (??):309–322, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042715004410.

Geum:2018:CFO

[GKN18] Young Hee Geum, Young Ik Kim, and Beny Neta. Constructing a family of optimal eighth-order modified Newton-type multiplezero finders along with the dynamics behind their purely imaginary extraneous fixed points. Journal of Computational and Applied Mathematics, 333 (??):131–156, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717305356.

Giannelli:2019:IIM

[GKPS19] Carlotta Giannelli, Tadej Kanduc, Francesca Pelosi, and Hendrik Speleers. An immersed-isogeometric model: Application to linear elasticity and implementation with THBoxsplines. Journal of Computational and Applied Mathematics, 349(??):410–423, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305788.

Gao:2017:FFA

[GKQW17] You Gao, Min Ku, Tao Qian, and Jianzhong Wang. FFT formulations of adaptive Fourier decomposition. Journal of Computational and Applied Mathematics, 324(??):204–215, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302005.

Guo:2019:RTT

[GL19] Ping Guo and Chong-Jun Li. Razumikhintype theorems on the moment stability of the exact and numerical solutions for the stochastic pantograph differential equations. Journal of Computational and Applied Mathematics, 355(??):77–90, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300202.

Galan:2017:MDI

[GLdCJS17] Celestino Ordóñez Galán, Fernando Sánchez Lasheras, Francisco Javier de Cos Juez, and Antonio Bernardo Sánchez. Missing data imputation of questionnaires by means of genetic algorithms with different fitness functions. Journal of Computational and Applied Mathematics, 311(??):704–717, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303727.

Gao:2019:OWU

[GLL19] Yulong Gao, Dong Liang, and Yonghai Li. Optimal weighted upwind finite volume method for convection–diffusion equations in 2D. Journal of Computational and Applied Mathematics, 359(??):73–87, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301438.

Guan:2016:SCI

[GLLS16] Jinrui Guan, Linzhang Lu, Ren-Cang Li, and Rongxia Shao. Self-Corrective Iterations (SCI) for generalized diagonally dominant matrices. Journal of Computational and Applied Mathematics, 302(??):285–300, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300681.

Gaspar:2016:NMO

[GLMT16] Francisco J. Gaspar, Francisco J. Lisbona, Piotr Matus, and Vo Thi Kim Tuyen. Numerical methods for a one-dimensional non-linear Biot’s model.

Journal of Computational and Applied Mathematics, 293(??):62–72, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001922.

Guo:2018:TMM

[GLMY18] Qian Guo, Wei Liu, Xuerong Mao, and Rongxian Yue. The truncated Milstein method for stochastic differential equations with commutative noise. Journal of Computational and Applied Mathematics, 338(??):298–310, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300359.

Goffard:2016:PEA

[GLP16]  Pierre-Olivier  Goffard,

Stéphane Loisel, and Denys Pommeret. A polynomial expansion to approximate the ultimate ruin probability in the compound Poisson ruin model. Journal of Computational and Applied Mathematics, 296(??):499–511, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003222.

Gameiro:2016:RNP

[GLR16] Marcio Gameiro, Jean-Philippe Lessard, and Yann Ricaud. Rigorous numerics for piecewisesmooth systems: a functional analytic approach based on Chebyshev series. Journal of Computational and Applied Mathematics, 292(??):654–673, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003106.

Galvis:2015:GMF

[GLS15] Juan Galvis, Guanglian Li, and Ke Shi. A Generalized Multiscale Finite Element Method for the Brinkman equation. Journal of Computational and Applied Mathematics, 280(??):294–309, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005172.

Grasegger:2016:SMA

[GLSW16] Georg Grasegger, Alberto Lastra, J. Rafael Sendra, and Franz Winkler. A solution method for autonomous first-order algebraic partial differential equations. Journal of Computational and Applied Mathematics, 300 (??):119–133, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715006421.

Grasegger:2018:RGS

[GLSW18] Georg Grasegger, Alberto Lastra, J. Rafael Sendra, and Franz Winkler. Rational general solutions of systems of first-order algebraic partial differential equations. Journal of Computational and Applied Mathematics, 331(??):88–103, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305009.

Gigola:2016:IEP

[GLT16] S. Gigola, L. Lebtahi, and N. Thome. The inverse eigenvalue problem for a Hermitian reflexive matrix and the optimization problem. Journal of Computational and Applied Mathematics, 291(??):449–457, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002058.

Gong:2015:ADM

[GLTW15] Wenyong Gong, Yong-Jin Liu, Kai Tang, and Tieru Wu. Approximate Delaunay mesh reconstruction and quality estimation from point samples. Journal of Computational and Applied Mathematics, 274(??):23–34, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003082.

Gine:2015:CIC

[GLV15] Jaume Giné, Jaume Llibre, and Claudia Valls. Centers and isochronous centers for generalized quintic systems. Journal of Computational and Applied Mathematics, 279(??):173–186, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004786.

Guo:2018:SPQ

[GLZ18] Zhen-Chen Guo, Tiexiang Li, and Ying-Ying Zhou.

Structure-preserving ΓQR and Γ-Lanczos algorithms for Bethe–Salpeter eigenvalue problems. Journal of Computational and Applied Mathematics, 341(??):12–30, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301456.

Ganesh:2016:HOF

[GM16a] M. Ganesh and C. Morgenstern. High-order FEM– BEM computer models for wave propagation in unbounded and heterogeneous media: Application to time-harmonic acoustic horn problem. Journal of Computational and Applied Mathematics, 307(??):183–203, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300711.

Guglielmi:2016:IMC

[GM16b] Nicola Guglielmi and Manuela Manetta. An iterative method for computing robustness of polynomial stability. Journal of Computational and Applied Mathematics, 292(??):638–653, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003349.

Goncalves:2017:NCG

[GM17a] Max L. N. Gonçalves and Jefferson G. Melo. A Newton conditional gradient method for constrained nonlinear systems. Journal of Computational and Applied Mathematics, 311(??):473–483, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303697.

Guermond:2017:HOT

[GM17b] Jean-Luc Guermond and Peter D. Minev. Highorder time stepping for the Navier–Stokes equations with minimal computational complexity. Journal of Computational and Applied Mathematics, 310(??):92–103, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302138.

Gudi:2019:CAF

[GM19] Thirupathi Gudi and Papri Majumder. Convergence analysis of finite element method for a parabolic obstacle problem. Journal of Computational and Applied Mathematics, 357(??):85–102, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300998.

Goessens:2015:VAO

[GMCD15] T. Goessens, B. Malengier, D. Constales, and R. H. De Staelen. A volume averaging and overlapping domain decomposition technique to model mass transfer in textiles. Journal of Computational and Applied Mathematics, 275(??):456–464, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714001071.

Gavhi-Molefe:2019:ISS

[GMdV19] Mpfareleni Rejoyce Gavhi-Molefe and Johan de Villiers. On interpolatory subdivision symbol formulation and parameter convergence intervals. Journal of Computational and Applied Mathematics, 349(??):354–365, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830356X.

Ghanbari:2019:NSC

[GMG19] F. Ghanbari, P. Mokhtary, and K. Ghanbari. Numerical solution of a class of fractional order integrodifferential algebraic equations using Müntz–Jacobi Tau method. Journal of Computational and Applied Mathematics, 362 (??):172–184, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302614.

Garcia-Mora:2018:MDI

[GMSR18] B. García-Mora, C. Santamaría, and G. Rubio. Modeling dependence in the inter-failure times. An analysis in reliability models by Markovian arrival processes. Journal of Computational and Applied Mathematics, 343(??):762–770, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306428. See

correction [GMSR21].

Garcia-Mora:2021:CMD

[GMSR21] B. García-Mora, C. Santamaría, and G. Rubio. Corrigendum to “Modeling dependence in the interfailure times. An analysis in Reliability models by Markovian Arrival Processes” [J. Comp. Appl. Math. 343, (2018), 762–770]. Journal of Computational and Applied Mathematics, 381 (??):Article 113085, January 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042720303769. See

[GMSR18].

Garcia-Mora:2016:BPF

[GMSRP16] B. García-Mora, C. Santamaría, G. Rubio, and J. L. Pontones. Bayesian prediction for flowgraph models with covariates. An application to bladder carcinoma. Journal of Computational and Applied Mathematics, 291(??):85–93, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001983.

Girejko:2018:NAB

[GMW18] Ewa Girejko, Dorota Mozyrska, and Malgorzata Wyrwas. Numerical analysis of behaviour of the Cucker–Smale type models with fractional operators. Journal of Computational and Applied Mathematics, 339(??):111–123, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717306337.

Goufo:2018:AFD

[GN18] Emile F. Doungmo Goufo and Juan J. Nieto. Attractors for fractional differential problems of transition to turbulent flows. Journal of Computational and Applied Mathematics, 339(??):329–342, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304156.

Gracia:2015:NAS

[GO15] J. L. Gracia and E. O’Riordan. Numerical approximation of solution derivatives in the case of singularly perturbed time dependent reaction–diffusion problems. Journal of Computational and Applied Mathematics, 273(??):13–24, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002684.

Gracia:2017:SPC

[GO17] J. L. Gracia and E. O’Riordan.

A singularly perturbed convection-diffusion problem with a moving pulse. Journal of Computational and Applied Mathematics, 321(??):371–388, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301127.

Gocwin:2015:SSI

[Goć15]  Maciej Goćwin.  Solv-

ing systems of IVPs with discontinuous derivatives — numerical experiments. Journal of Computational and Applied Mathematics, 290(??):476–499, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003295.

Goda:2015:GIP

[God15] Takashi Goda. Good interlaced polynomial lattice rules for numerical integration in weighted Walsh spaces. Journal of Computational and Applied Mathematics, 285(??):279–294, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001119.

Goda:2016:QMC

[God16] Takashi Goda. Quasi-Monte Carlo integration using digital nets with antithetics. Journal of Computational and Applied Mathematics, 304(??):26–42, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301042.

Garcia-Olivo:2017:FOR

[GOGM17] M. García-Olivo, José M. Gutiérrez, and Á. A. Magreñán. A first overview on the real dynamics of Chebyshev’s method. Journal of Computational and Applied Mathematics, 318(??):422– 432, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300966.

Gorb:2016:SHF

[Gor16]  Yuliya Gorb.  Singu-

larities of hydrodynamic forces acting on particles in the near-contact regime. Journal of Computational and Applied Mathematics, 307(??):82–92, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301881.

Gudi:2016:PEE

[GP16] Thirupathi Gudi and Kamana Porwal. A posteriori error estimates of discontinuous Galerkin methods for the Signorini problem. Journal of Computational and Applied Mathematics, 292(??):257–278, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003702.

Gonzalez-Pinto:2015:ARK

[GPHAPR15] S. Gonzalez-Pinto, D. Hernandez-Abreu, and S. Perez-Rodriguez. AMF–Runge– Kutta formulas and error estimates for the time integration of advection diffusion reaction PDEs. Journal of Computational and Applied Mathematics, 289(??):3–21, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002010.

Gonzalez-Pinto:2017:MSS

[GPHAPR17] S. González-Pinto, D. Hernández-Abreu, and S. Pérez-Rodríguez. W -methods to stabilize standard explicit Runge–Kutta methods in the time integration of advection– diffusion-reaction PDEs. Journal of Computational and Applied Mathematics, 316(??):143–160, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303958.

Guo:2016:AMS

[GQM16] Qian Guo, Mingming Qiu, and Taketomo Mitsui. Asymptotic meansquare stability of explicit Runge–Kutta Maruyama methods for stochastic delay differential equations. Journal of Computational and Applied Mathematics, 296(??):427–442, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005142.

Giner:2016:SEM

[GR16] Gabriela Ribes Giner and Agustín Peralt Rillo. Structural equation modeling of co-creation and its influence on the student’s satisfaction and loyalty towards university. Journal of Computational and Applied Mathematics, 291(??):257–263, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001144.

Groza:2019:ASP

[GR19] Ghiocel Groza and Mohsen Razzaghi. Approximation of solutions of polynomial partial differential equations in two independent variables. Journal of Computational and Applied Mathematics, 346(??):205–223, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S037704271830431X.

Grigo:2019:HON

[Gri19] Alexander Grigo. Highorder numerical schemes for jump-SDEs. Journal of Computational and Applied Mathematics, 354 (??):31–38, July 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718307490.

Ghattassi:2019:AFD

[GRS19] Mohamed Ghattassi, Jean Rodolphe Roche, and Didier Schmitt. Analysis of a full discretization scheme for 2D radiative-conductive heat transfer systems. Journal of Computational and Applied Mathematics, 346(??):1–17, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303807.

Grudzien:2016:IHS

[Gru16] Colin Grudzien. The instability of the Hocking– Stewartson pulse and its geometric phase in the Hopf bundle. Journal of Computational and Applied Mathematics, 307(??):162–169, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301066.

Garcia:2015:ATD

[GRVC15] Mariangel Garcia, Isabel Ramirez, Martin Verlaan, and Jose Castillo. Application of a three-dimensional hydrodynamic model for San Quintin Bay, B.C., Mexico. Validation and calibration using OpenDA. Journal of Computational and Applied Mathematics, 273(??):428–437, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002362.

Goldman:2015:QBB

[GS15a] Ron Goldman and Plamen Simeonov. Quantum Bernstein bases and quantum Bézier curves. Journal of Computational and Applied Mathematics, 288(??):284–303, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002459.

Gracia:2015:CDA

[GS15b] J. L. Gracia and M. Stynes. Central difference approximation of convection in Caputo fractional derivative two-point boundary value problems. Journal of Computational and Applied Mathematics, 273(??):103–115, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002702.

Grimmonprez:2015:NPI

[GS15c] Marijke Grimmonprez and Marián Slodicka. A nonlinear parabolic integrodifferential problem with an unknown Dirichlet boundary condition. Journal of Computational and Applied Mathematics, 275(??):421–432, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002210.

Grimmonprez:2015:RUS

[GS15d] Marijke Grimmonprez and Marián Slodicka. Reconstruction of an unknown source parameter in a semilinear parabolic problem. Journal of Computational and Applied Mathematics, 289(??):331–345, December 1, 2015. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714005834.

Gabrielides:2018:SMC

[GS18a] Nikolaos C. Gabrielides and Nickolas S. Sapidis. C^{1} sign, monotonicity and convexity preserving Hermite polynomial splines of variable degree. Journal of Computational and Applied Mathematics, 343(??):662–707, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301663.

George:2018:NAT

[GS18b] Santhosh George and M. Sabari. Numerical approximation of a Tikhonov type regularizer by a discretized frozen steepest descent method. Journal of Computational and Applied Mathematics, 330(??):488– 498, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0377042717304533.

Gabrielides:2019:CPC

[GS19] Nikolaos C. Gabrielides and Nickolas S. Sapidis. Cubic polynomial and cubic rational C^{1} sign, monotonicity and convexity preserving Hermite interpolation. Journal of Computational and Applied Mathematics, 357(??):184–203, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300974.

Goel:2018:ITE

[GSM18] Anubha Goel, Amita Sharma, and Aparna Mehra. Index tracking and enhanced indexing using mixed conditional valueat-risk. Journal of Computational and Applied Mathematics, 335(??):361– 380, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717306350.

Goel:2019:ROM

[GSM19] Anubha Goel, Amita Sharma, and Aparna Mehra. Robust optimization of mixed CVaR STARR ratio using copulas. Journal of Computational and Applied Mathematics, 347(??):62–83, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304746.

Gupta:2018:PEA

[GSRJ18] Jhuma Sen Gupta, Rajen Kumar Sinha, Gujji Murali Mohan Reddy, and Jinank Jain. A posteriori error analysis of the Crank–Nicolson finite element method for linear parabolic interface problems: a reconstruction approach. Journal of Computational and Applied Mathematics, 340(??):173–190, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301018.

Gibert:2016:RST

[GSVSM16] Karina Gibert, Beatriz Sevilla-Villanueva, and Miquel Sànchez-Marrè. The role of significance tests in consistent interpretation of nested partitions. Journal of Computational and Applied Mathematics, 292(??):623–633, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000436.

Guha:2017:MAB

[GT17] Nilabja Guha and Xiaosi Tan. Multilevel approximate Bayesian approaches for flows in highly heterogeneous porous media and their applications. Journal of Computational and Applied Mathematics, 317(??):700– 717, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304976.

Grote:2019:CMH

[GT19] Marcus J. Grote and Jet Hoe Tang. On controllability methods for the Helmholtz equation. Journal of Computational and Applied Mathematics, 358(??):306–326, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301414.

Gao:2016:AGM

[GTC16] Longfei Gao, Xiaosi Tan, and Eric T. Chung. Application of the Generalized Multiscale Finite Element Method in parameter-dependent PDE simulations with a variable-separation technique. Journal of Computational and Applied Mathematics, 300(??):183– 191, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006433.

Guo:2019:HOL

[GTX^{+}19] Hui Guo, Lulu Tian, Ziyao Xu, Yang Yang, and Ning Qi. Highorder local discontinuous Galerkin method for simulating wormhole propagation. Journal of Computational and Applied Mathematics, 350(??):247–261, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306289.

Gurler:2015:MRS

[GUB15] Selma Gurler, Burcu Hudaverdi Ucer, and Ismihan Bairamov. On the mean remaining strength at the system level for some bivariate survival models based on exponential distribution. Journal of Computational and Applied Mathematics, 290 (??):535–542, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003337.

Gavete:2017:SSO

[GUB^{+}17] L. Gavete, F. Ureña, J. J. Benito, A. García, M. Ureña, and E. Salete. Solving second order nonlinear elliptic partial differential equations using generalized finite difference method. Journal of Computational and Applied Mathematics, 318(??):378– 387, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303442.

Gucuyenen:2017:SSM

[Güc17] Nurcan Gücüyenen. Strang splitting method to Benjamin– Bona–Mahony type equations: Analysis and application. Journal of Computational and Applied Mathematics, 318(??):616– 623, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0377042715005609.

Gunasekera:2018:IRF

[Gun18] Sumith Gunasekera. Inference for the reliability function based on progressively type II censored data from the Pareto model: the generalized variable approach. Journal of Computational and Applied Mathematics, 343(??):275–288, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302218.

Gunasekera:2019:BIC

[Gun19] Sumith Gunasekera. Bayesian inference for the common location parameter of several shifted-exponential populations. Journal of Computational and Applied Mathematics, 353 (??):21–37, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718307301.

Guo:2017:SCS

[Guo17] Daniel X. Guo. On stability and convergence of semi-Lagrangian methods for the first-order timedependent nonlinear partial differential equations in 1D. Journal of Computational and Applied Mathematics, 324(??):72–84, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301929.

Guohui:2018:BSD

[Guo18] Zhao Guohui. Bivariate spline differential forms. Journal of Computational and Applied Mathematics, 343(??):748–752, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302462.

Gine:2017:NCA

[GV17] Jaume Giné and Claudia Valls. Nondegenerate centers for Abel polynomial differential equations of second kind. Journal of Computational and Applied Mathematics, 321(??):469–477, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301280.

Gomez-Valle:2017:NTE

[GVHMR17] L. Gómez-Valle, Z. Habibilashkary, and J. Martínez-Rodríguez. A new technique to estimate the risk-neutral processes in jump-diffusion commodity futures models. Journal of Computational and Applied Mathematics, 309(??):435–441, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006408.

Gomez-Valle:2018:MSC

[GVHMR18] L. Gómez-Valle, Z. Habibilashkary, and J. Martínez-Rodríguez. A multiplicative seasonal component in commodity derivative pricing. Journal of Computational and Applied Mathematics, 330(??):835–847, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302509.

Gomez-Valle:2016:ERN

[GVMR16] L. Gómez-Valle and J. Martínez-Rodríguez. Estimation of risk-neutral processes in single-factor jump-diffusion interest rate models. Journal of Computational and Applied Mathematics, 291(??):48–57, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001016.

Gomez-Valle:2019:RNS

[GVMR19] L. Gómez-Valle and J. Martínez-Rodríguez. The riskneutral stochastic volatility in interest rate models with jump-diffusion processes. Journal of Computational and Applied Mathematics, 347(??):49–61, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304758.

Gospodarczyk:2016:MBC

[GW16] Przemyslaw Gospodarczyk and Pawel Woźny. Merging of Bézier curves with box constraints. Journal of Computational and Applied Mathematics, 296 (??):265–274, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715005038.

Gu:2017:AIA

[GW17] Chuanqing Gu and Wenwen Wang. An Arnoldi– Inout algorithm for computing PageRank problems. Journal of Computational and Applied Mathematics, 309(??):219–229, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302606.

Greif:2019:CGN

[GW19] Chen Greif and Michael Wathen. Conjugate gradient for nonsingular saddle-point systems with a maximally rankdeficient leading block. Journal of Computational and Applied Mathematics, 358(??):1–11, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300895.

Gao:2018:EBL

[GWLL18] Lei Gao, Yaqiang Wang, Chaoqian Li, and Yaotang Li. Error bounds for linear complementarity problems of S-Nekrasov matrices and B–S–Nekrasov matrices. Journal of Computational and Applied Mathematics, 336(??):147– 159, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730660X.

Guo:2018:MCW

[GWT18] Huijun Guo, Jinru Wang, and Xinyan Tian. The mean consistency of wavelet estimators for convolutions of the density functions. Journal of Computational and Applied Mathematics, 343(??):1–11, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302401.

Guo:2017:ARR

[GWY17] Fenglong Guo, Dingcheng Wang, and Hailiang Yang. Asymptotic results for ruin probability in a twodimensional risk model with stochastic investment returns. Journal of Computational and Applied Mathematics, 325(??):198–221, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302273.

Gu:2017:NAR

[GX17] Qiupeng Gu and Zuxing Xuan. A new approach for ranking fuzzy numbers based on possibility theory. Journal of Computational and Applied Mathematics, 309(??):674–682, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302345.

Guo:2015:MSL

[GxLlW15] Peng Guo, Cui xia Li, and Shi liang Wu. A modified SOR-like method for the augmented systems. Journal of Computational and Applied Mathematics, 274(??):58–69, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003094.

Gu:2015:TSM

[GXZ15] Chuanqing Gu, Fei Xie, and Ke Zhang. A twostep matrix splitting iteration for computing PageRank. Journal of Computational and Applied Mathematics, 278(??):19–28, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004294.

Guo:2016:GVI

[GYL16] Chaohui Guo, Hu Yang, and Jing Lv. Generalized varying index coefficient models. Journal of Computational and Applied Mathematics, 300(??):1– 17, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005798.

Guo:2016:CMT

[GYLQ16] Guangbao Guo, Wenjie You, Lu Lin, and Guoqi Qian. Covariance matrix and transfer function of dynamic generalized linear models. Journal of Computational and Applied Mathematics, 296(??):613– 624, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005154.

Guo:2015:PML

[GYQS15] Guangbao Guo, Wenjie You, Guoqi Qian, and Wei Shao. Parallel maximum likelihood estimator for multiple linear regression models.

Journal of Computational and Applied Mathematics, 273(??):251–263, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002775.

Gokmen:2017:NAS

[GYS17] Elcin Gokmen, Gamze Yuksel, and Mehmet Sezer. A numerical approach for solving Volterra type functional integral equations with variable bounds and mixed delays. Journal of Computational and Applied Mathematics, 311(??):354–363, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303648.

Geng:2017:PEE

[GYX17] Hongrui Geng, Tao Yin, and Liwei Xu. A priori error estimates of the DtN– FEM for the transmission problem in acoustics. Journal of Computational and Applied Mathematics, 313(??):1–17, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716304241.

Gao:2019:FEM

[GZD^{+}19] Jing Gao, Meng Zhao, Ning Du, Xu Guo, Hong Wang, and Jiwei Zhang. A finite element method for space-time directional fractional diffusion partial differential equations in the plane and its error analysis. Journal of Computational and Applied Mathematics, 362 (??):354–365, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719302481.

Garcia-Zapata:2019:ASM

[GZDC19] Juan Luis García-Zapata, Juan Carlos Díaz Martín, and Álvaro Cortés Fácila. An adaptive subdivision method for root finding of univariate polynomials. Journal of Computational and Applied Mathematics, 352(??):146–164, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307131.

Gao:2019:ADC

[GZH19] Qingwu Gao, Jun Zhuang, and Zhongquan Huang. Asymptotics for a delayclaim risk model with diffusion, dependence structures and constant force of interest. Journal of Computational and Applied Mathematics, 353(??):219– 231, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307672.

Guo:2016:PPR

[GZZZ16] Hailong Guo, Zhimin Zhang, Ren Zhao, and Qingsong Zou. Polynomial preserving recovery on boundary. Journal of Computational and Applied Mathematics, 307(??):119–133, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301133.

Hashemi:2018:SWS

[HA18] M. S. Hashemi and Ali Akgül. Solitary wave solutions of time-space nonlinear fractional Schrödinger’s equation: Two analytical approaches. Journal of Computational and Applied Mathematics, 339 (??):147–160, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717305733.

Haynes:2019:LDD

[HA19] Ronald D. Haynes and Faysol Ahmed. Linearized domain decomposition approaches for nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 346(??):620–637, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302887.

Hackmann:2018:ATO

[Hac18] Daniel Hackmann. Analytic techniques for option pricing under a hyperexponential Lévy model. Journal of Computational and Applied Mathematics, 342(??):225–248, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302164.

Hadjidimos:2015:MAS

[Had15]  Apostolos  Hadjidimos.

The matrix analogue of the scalar AOR iterative method. Journal of Computational and Applied Mathematics, 288(??):366–378, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002447.

Hagwood:2015:RCE

[Hag15] Charles Hagwood. Reconstruction of conditional expectations from product moments with applications. Journal of Computational and Applied Mathematics, 276(??):129–142, March 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003756.

Hoai:2017:MSA

[HAH17] Tran Van Hoai, Phan Thanh An, and Nguyen Ngoc Hai. Multiple shooting approach for computing approximately shortest paths on convex polytopes. Journal of Computational and Applied Mathematics, 317(??):235– 246, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305155.

Hainaut:2017:CLP

[Hai17] Donatien Hainaut. Clustered Lévy processes and their financial applications. Journal of Computational and Applied Mathematics, 319(??):117–140, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300055.

Hajarian:2019:EAB

[Haj19] Masoud Hajarian. An efficient algorithm based on Lanczos type of BCR to solve constrained quadratic inverse eigenvalue problems. Journal of Computational and Applied Mathematics, 346(??):418–431, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830445X.

Han:2015:CPA

[Han15]  Xuli Han.  Convexity-

preserving approximation by univariate cubic splines. Journal of Computational and Applied Mathematics, 287(??):196–206, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002046.

Han:2019:DCT

[Han19]  Xuli Han.  Direction-

consistent tangent vectors for generating interpolation curves. Journal of Computational and Applied Mathematics, 346(??):237–246, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304266.

Harris:2018:APB

[Har18] Lawrence A. Harris. Alternation points and bivariate Lagrange interpolation. Journal of Computational and Applied Mathematics, 340(??):43–52, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300852.

Huang:2016:BTC

[HB16] Zhu Huang and John P.

Boyd. Bandwidth truncation for Chebyshev polynomial and ultraspherical/ Chebyshev Galerkin discretizations of differential equations: Restrictions and two improvements. Journal of Computational and Applied Mathematics, 302(??):340–355, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300279.

Haslinger:2017:IPG

[HBH17] J. Haslinger, R. Blaheta, and R. Hrtus. Identification problems with given material interfaces. Journal of Computational and Applied Mathematics, 310(??):129–142, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302953.

Hou:2015:SFD

[HC15] Tianliang Hou and Yanping Chen. Superconvergence of fully discrete rectangular mixed finite element methods of parabolic control problems. Journal of Computational and Applied Mathematics, 286(??):79–92, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005317.

Hok:2017:OPL

[HC17] Julien Hok and Tat Lung (Ron) Chan. Option pricing with Legendre polynomials. Journal of Computational and Applied Mathematics, 322(??):25–45, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301462.

Heidarkhani:2018:VAP

[HCA^{+}18] S. Heidarkhani, A. Cabada, G. A. Afrouzi, S. Moradi, and G. Caristi. A variational approach to perturbed impulsive fractional differential equations. Journal of Computational and Applied Mathematics, 341(??):42–60, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301122.

Huang:2018:SPM

[HCC18] Yunying Huang, Zhen Chao, and Guoliang Chen. Spectral properties of the matrix splitting preconditioners for generalized saddle point problems. Journal of Computational and Applied Mathematics, 332(??):1– 12, April 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304922.

Hao:2017:SOD

[HCL17] Zhaopeng Hao, Wanrong Cao, and Guang Lin. A second-order difference scheme for the time fractional substantial diffusion equation. Journal of Computational and Applied Mathematics, 313(??):54–69, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304265.

Hou:2016:AMH

[HCSC16] Jiangyong Hou, Jie Chen, Shuyu Sun, and Zhangxin Chen. Adaptive mixedhybrid and penalty discontinuous Galerkin method for two-phase flow in heterogeneous media. Journal of Computational and Applied Mathematics, 307(??):262–283, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042716300309.

Hervert-Escobar:2018:TDO

[HEA18] Laura Hervert-Escobar and Vassil Alexandrov. Territorial design optimization for business sales plan. Journal of Computational and Applied Mathematics, 340(??):501–507, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300815.

Heintz:2015:NMS

[Hei15] A. Heintz. A numerical method for simulation dynamics of incompressible lipid membranes in viscous fluid. Journal of Computational and Applied Mathematics, 289(??):87–100, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001570.

Hilario:2016:TOA

[HFMM16] L. Hilario, A. Falcó, N. Montés, and M. C. Mora. A tensor optimization algorithm for Bézier shape deformation. Journal of Computational and Applied Mathematics, 291(??):264–280, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001053.

Han:2017:OPV

[HG17] Xuli Han and Xiao Guo. Optimal parameter values for approximating conic sections by the quartic Bézier curves. Journal of Computational and Applied Mathematics, 322(??):86–95, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301528.

Han:2018:CHI

[HG18] Xuli Han and Xiao Guo. Cubic Hermite interpolation with minimal derivative oscillation. Journal of Computational and Applied Mathematics, 331(??):82–87, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304910.

Hu:2019:SIE

[HG19] Xuehua Hu and Wenhao Gui. Statistical inference for the exponentiated half logistic distribution based on censored data. Journal of Computational and Applied Mathematics, 362(??):219–244, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302523.

Han:2015:RPB

[HH15] Xuli Han and Jing Han.

Representation of piecewise biharmonic surfaces using biquadratic B-splines. Journal of Computational and Applied Mathematics, 290(??):403–411, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003192.

He:2016:PEA

[HH16] Yinnian He and Weizhang Huang. A posteriori error analysis for finite element solution of one-dimensional elliptic differential equations using equidistributing meshes. Journal of Computational and Applied Mathematics, 299 (??):101–126, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715005439.

Herceg:2018:EOF

[HH18a] Djordje Herceg and Dragoslav Herceg. Eighth order family of iterative methods for nonlinear equations and their basins of attraction. Journal of Computational and Applied Mathematics, 343(??):458–480, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302358.

Herceg:2018:NFM

[HH18b] Djordje Herceg and Dragoslav Herceg. A new family of methods for single and multiple roots. Journal of Computational and Applied Mathematics, 342(??):1–15, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301808.

Huang:2018:MGS

[HH18c] Zhuo-Hong Huang and Ting-Zhu Huang. A modified generalized shiftsplitting method for nonsymmetric saddle point problems. Journal of Computational and Applied Mathematics, 328 (??):381–399, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303850.

Huang:2018:SCA

[HH18d] Zhuo-Hong Huang and Ting-Zhu Huang. Semiconvergence analysis of the GSS iteration methods for singular saddle point problems. Journal of Computational and Applied Mathematics, 328(??):1–11, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303412.

Han:2019:NAE

[HH19a] Danfu Han and Weimin Han. Numerical analysis of an evolutionary variational-hemivariational inequality with application to a dynamic contact problem. Journal of Computational and Applied Mathematics, 358(??):163–178, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301293.

Huang:2019:NAA

[HH19b] Jie Huang and Ting-Zhu Huang. A nonstationary accelerating alternating direction method for frame-based Poissonian image deblurring. Journal of Computational and Applied Mathematics, 352(??):181–193, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307180.

Hendricks:2017:HOA

[HHEG17] Christian Hendricks, Christof Heuer, Matthias Ehrhardt, and Michael Günther. High-order ADI finite difference schemes for parabolic equations in the combination technique with application in finance. Journal of Computational and Applied Mathematics, 316(??):175–194, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304137.

Heinlein:2015:ACM

[HHKR15] A. Heinlein, U. Hetmaniuk, A. Klawonn, and O. Rheinbach. The approximate component mode synthesis special finite element method in two dimensions: Parallel implementation and numerical results. Journal of Computational and Applied Mathematics, 289(??):116–133, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001405.

Hao:2018:CBI

[HHLS18] Wenrui Hao, Bei Hu, Shuwang Li, and Lingyu Song. Convergence of boundary integral method for a free boundary system. Journal of Computational and Applied Mathematics, 334(??):128–157, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305769.

Hao:2017:DTR

[HHOT17] Dinh Nho Hào, Bui Viet Huong, Nguyen Thi Ngoc Oanh, and Phan Xuan Thanh. Determination of a term in the right-hand side of parabolic equations. Journal of Computational and Applied Mathematics, 309(??):28–43, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302564.

Haghighi:2016:DDI

[HHR16] Amir Haghighi, Seyed Mohammad Hosseini, and Andreas Rößler. Diagonally drift-implicit Runge– Kutta methods of strong order one for stiff stochastic differential systems. Journal of Computational and Applied Mathematics, 293(??):82–93, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001065.

Hansen:2015:NNC

[HHS^{+}15] Jakob K. Hansen, Jarom D. Hogue, Grant K. Sander, Rosemary A. Renaut, and Sudeep C. Popat. Nonnegatively constrained least squares and parameter choice by the residual periodogram for the inversion of electrochemical impedance spectroscopy data. Journal of Computational and Applied Mathematics, 278(??):52–74, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004154.

Hu:2018:SEM

[HHZ18] Xianliang Hu, Danfu Han, and Jiang Zhu. Spline element methods allowing multiple level hanging nodes. Journal of Computational and Applied Mathematics, 337(??):125–134, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300347.

Hussain:2016:SDS

[HIS16] Kasim Hussain, Fudziah Ismail, and Norazak Senu. Solving directly special fourth-order ordinary differential equations using Runge–Kutta type method. Journal of Computational and Applied Mathematics, 306(??):179– 199, November 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716301662.

Hejazi:2017:EVS

[HJ17a] Seyed Amir Hejazi and Kenneth R. Jackson. Efficient valuation of SCR via a neural network approach. Journal of Computational and Applied Mathematics, 313(??):427–439, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304940.

Huang:2017:REM

[HJ17b] Yi Huang and Zhongxiao Jia. On regularizing effects of MINRES and MR-II for large scale symmetric discrete ill-posed problems. Journal of Computational and Applied Mathematics, 320:145–163, August 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Huang:2018:AAE

[HJL18] Yunqing Huang, Hongen Jia, and Jichun Li. Analysis and application of an equivalent Berenger’s PML model. Journal of Computational and Applied Mathematics, 333(??):157–169, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305381.

Hauer:2019:PAS

[HJS19] Michael Hauer, Bert Jüttler, and Josef Schicho. Projective and affine symmetries and equivalences of rational and polynomial surfaces. Journal of Computational and Applied Mathematics, 349(??):424–437, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303789.

Higham:2007:SCR

[HK07] Desmond J. Higham and Peter E. Kloeden. Strong convergence rates for backward Euler on a class of nonlinear jumpdiffusion problems. Journal of Computational and Applied Mathematics, 205(2):949–956, August 15, 2007. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042706004158. See

comments [TGJ19].

Hopkins:2018:AAE

[HK18] Tim Hopkins and Emrah Kiliç. An analytical approach: Explicit inverses of periodic tridiagonal matrices. Journal of Computational and Applied Mathematics, 335(??):207– 226, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305988.

Han:2019:NSA

[HK19] Joo Hyeong Han and Jae Ryong Kweon. A numerical scheme for approximating interior jump discontinuity solution of a compressible Stokes system. Journal of Computational and Applied Mathematics, 345:320–337, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Hazarika:2018:MLS

[HKAR18] Bipan Hazarika, Erdal Karapinar, Reza Arab, and Mohsen Rabbani. Metric-like spaces to prove existence of solution for nonlinear quadratic integral equation and numerical method to solve it. Journal of Computational and Applied Mathematics, 328(??):302–313, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303539.

Hui:2015:DPA

[HKZZ15] Jun-Jun Hui, Xiang-Yu Kong, He-Xin Zhang, and Xin Zhou. Delaypartitioning approach for systems with interval timevarying delay and nonlinear perturbations. Journal of Computational and Applied Mathematics, 281 (??):74–81, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714005408.

Hu:2015:FGS

[HL15] Wenbin Hu and Shenghong Li. Fast Greeks by simulation: the block adjoint method with memory reduction. Journal of Computational and Applied Mathematics, 274(??):70–78, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003124.

Hao:2018:IIQ

[HL18a] Yong-Xia Hao and Chong-Jun Li. The C^{1} and C^{2} quasi-plateau problems. Journal of Computational and Applied Mathematics, 329(??):106–124, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302108.

Hennink:2018:DGM

[HL18b] Aldo Hennink and Danny Lathouwers. A discontinuous Galerkin method for the mono-energetic Fokker–Planck equation based on a spherical interior penalty formulation. Journal of Computational and Applied Mathematics, 330(??):253–267, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304089.

Haelterman:2015:EQL

[HLBP15] R. Haelterman, B. Lauwens, H. Bruyninckx, and J. Petit. Equivalence of QN– LS and BQN–LS for affine problems. Journal of Computational and Applied Mathematics, 278(??):48–51, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004324.

Harizanov:2017:PAP

[HLG^{+}17] Stanislav Harizanov, Ivan Lirkov, Krassimir Georgiev, Marcin Paprzycki, and Maria Ganzha. Performance analysis of a parallel algorithm for restoring large-scale CT images. Journal of Computational and Applied Mathematics, 310(??):104–114, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303132.

Hong:2015:NMP

[HLL15] Yicheng Hong, Sungchul Lee, and Tianguo Li. Numerical method of pricing discretely monitored barrier option. Journal of Computational and Applied Mathematics, 278(??):149–161, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003793.

He:2018:CNL

[HLLS18] Chunlin He, Shuchao Li, Wenjun Luo, and Liqun Sun. Calculating the normalized Laplacian spectrum and the number of spanning trees of linear pentagonal chains. Journal of Computational and Applied Mathematics, 344(??):381–393, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303224.

Huang:2019:IIS

[HLLY19] Wei-Qiang Huang, Wen-Wei Lin, Henry Horng-Shing Lu, and Shing-Tung Yau. iSIRA: Integrated shift-invert residual Arnoldi method for graph Laplacian matrices from big data. Journal of Computational and Applied Mathematics, 346(??):518–531, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304515.

Hu:2018:CRS

[HLM18] Liangjian Hu, Xiaoyue Li, and Xuerong Mao. Convergence rate and stability of the truncated Euler–Maruyama method for stochastic differential equations. Journal of Computational and Applied Mathematics, 337(??):274–289, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300463.

Huska:2019:CNC

[HLMS19] Martin Huska, Alessandro Lanza, Serena Morigi, and Fiorella Sgallari. Convex non-convex segmentation of scalar fields over arbitrary triangulated surfaces. Journal of Computational and Applied Mathematics, 349(??):438–451, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830400X.

Haelterman:2015:SBQ

[HLV^{+}15] R. Haelterman, B. Lauwens, F. Van Utterbeeck, H. Bruyninckx, and J. Vierendeels. On the similarities between the quasi-Newton least squares method and GMRes. Journal of Computational and Applied Mathematics, 273(??):25–28, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002659.

He:2016:PSS

[HLV16] Xiaoming He, Jiangguo Liu, and Erik Van Vleck. Preface — SI: SIAM-2015. Journal of Computational and Applied Mathematics, 307(??):1, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302151.

Huang:2019:CMF

[HLZ19] Yunqing Huang, Jichun Li, and Shangyou Zhang. Continuous mixed finite elements for the second order elliptic equation with a low order term. Journal of Computational and Applied Mathematics, 357(??):273–283, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301153.

Hao:2016:DSE

[HM16] Sijia Hao and Per-Gunnar Martinsson. A direct solver for elliptic PDEs in three dimensions based on hierarchical merging of Poincaré–Steklov operators. Journal of Computational and Applied Mathematics, 308 (??):419–434, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302308.

He:2019:LFA

[HM19a] Yunhui He and Scott P.

MacLachlan. Local Fourier analysis for mixed finiteelement methods for the Stokes equations. Journal of Computational and Applied Mathematics, 357 (??):161–183, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300457.

Huang:2019:ICI

[HM19b] Baohua Huang and Changfeng Ma. Iterative criteria for identifying strong Htensors. Journal of Computational and Applied Mathematics, 352(??):93– 109, May 15, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718307234.

Han:2019:PEI

[HMD19] Minggang Han, Qiang Ma, and Xiaohua Ding.

The projected explicit Itô–Taylor methods for stochastic differential equations under locally Lipschitz conditions and polynomial growth conditions. Journal of Computational and Applied Mathematics, 348(??):161–180, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305442.

Hendy:2019:SHT

[HMDSR19] Ahmed S. Hendy, J. E. Macías-Díaz, and Adán J. Serna-Reyes. On the solution of hyperbolic twodimensional fractional systems via discrete variational schemes of high order of accuracy. Journal of Computational and Applied Mathematics, 354 (??):612–622, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718307003.

Hong:2018:TDC

[HMH18] Phong Luu Hong, Triet Le Minh, and Quan Pham Hoang. On a three dimensional Cauchy problem for inhomogeneous Helmholtz equation associated with perturbed wave number. Journal of Computational and Applied Mathematics, 335(??):86– 98, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306039.

Hoseinzadeh:2019:SRG

[HMKCR19] Akram Hoseinzadeh, Mohsen Maleki, Zahra Khodadadi, and Javier E. Contreras-Reyes. The skew-reflected-Gompertz distribution for analyzing symmetric and asymmetric data. Journal of Computational and Applied Mathematics, 349(??):132–141, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305569.

Harizanov:2017:VCP

[HMMV17] Stanislav Harizanov, Svetozar Margenov, Pencho Marinov, and Yavor Vutov. Volume constrained 2-phase segmentation method utilizing a linear system solver based on the best uniform polynomial approximation of x^{−1/2}. Journal of Computational and Applied Mathematics, 310(??):115–128, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302928.

Hochstenbach:2015:JDM

[HMP15] Michiel E. Hochstenbach, Andrej Muhic, and Bor Plestenjak. Jacobi– Davidson methods for polynomial two-parameter eigenvalue problems. Journal of Computational and Applied Mathematics, 288(??):251–263, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271500237X.

Han:2019:PMU

[HMS19] Weimin Han, Stanisław Migórski, and Mircea Sofonea. On penalty method for unilateral contact problem with nonmonotone contact condition. Journal of Computational and Applied Mathematics, 356(??):293–301, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300676.

Hueso:2015:CED

[HMT15] José L. Hueso, Eulalia Martínez, and Carles Teruel. Convergence, efficiency and dynamics of new fourth and sixth order families of iterative methods for nonlinear systems. Journal of Computational and Applied Mathematics, 275(??):412–420, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002830.

Hanke:2017:LSC

[HMT^{+}17] Michael Hanke, Roswitha März, Caren Tischendorf, Ewa Weinmüller, and Stefan Wurm. Leastsquares collocation for linear higher-index differentialalgebraic equations. Journal of Computational and Applied Mathematics, 317 (??):403–431, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716306276.

Hu:2019:WGF

[HMY19] Xiaozhe Hu, Lin Mu, and Xiu Ye. A weak Galerkin finite element method for the Navier– Stokes equations. Journal of Computational and Applied Mathematics, 362 (??):614–625, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718304990.

Harrat:2018:SOC

[HND18] Aicha Harrat, Juan J. Nieto, and Amar Debbouche. Solvability and optimal controls of impulsive Hilfer fractional delay evolution inclusions with Clarke subdifferential. Journal of Computational and Applied Mathematics, 344 (??):725–737, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303029.

Hosseinpour:2019:MLS

[HNT19]  Soleiman  Hosseinpour,

Alireza Nazemi, and Emran Tohidi. Müntz– Legendre spectral collocation method for solving delay fractional optimal control problems. Journal of Computational and Applied Mathematics, 351 (??):344–363, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718306800.

Huang:2017:EPD

[HOM17] Chun-Sung Huang, John G.

O’Hara, and Sure Mataramvura. Efficient pricing of discrete arithmetic Asian options under mean reversion and jumps based on Fourier-cosine expansions. Journal of Computational and Applied Mathematics, 311(??):230–238, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303387.

Hon:2019:CPF

[Hon19] Sean Hon. Circulant preconditioners for functions of Hermitian Toeplitz matrices. Journal of Computational and Applied Mathematics, 352(??):328–340, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306769.

Hansen:2016:ESD

[HOS16] Eskil Hansen, Alexander Ostermann, and Katharina Schratz. The error structure of the Douglas– Rachford splitting method for stiff linear problems. Journal of Computational and Applied Mathematics, 303(??):140–145, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300930.

Huang:2018:ROI

[HOTZ18] Ya Huang, Yao Ouyang, Lingxiao Tang, and Jieming Zhou. Robust optimal investment and reinsurance problem for the product of the insurer’s and the reinsurer’s utilities. Journal of Computational and Applied Mathematics, 344(??):532–552, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303376.

Huang:2015:TRS

[HP15] Mingxia Huang and Dingguo Pu. A trustregion SQP method without a penalty or a filter for nonlinear programming. Journal of Computational and Applied Mathematics, 281(??):107– 119, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005779.

Haqiri:2017:MVS

[HP17] Tayyebe Haqiri and Federico Poloni. Methods for verified stabilizing solutions to continuous-time algebraic Riccati equations. Journal of Computational and Applied Mathematics, 313(??):515–535, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304411.

Hinrichs:2019:TPW

[HPT19] Aicke Hinrichs, Friedrich Pillichshammer, and Shu Tezuka. Tractability properties of the weighted star discrepancy of the Halton sequence. Journal of Computational and Applied Mathematics, 350(??):46– 54, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305934.

Horvath:2016:SSP

[HPW16] Zoltán Horváth, Helmut Podhaisky, and Rüdiger Weiner. Strong stability preserving explicit peer methods. Journal of Computational and Applied Mathematics, 296(??):776– 788, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005506.

Huang:2018:PDP

[HQ18] Zheng-Hai Huang and Liqun Qi. Positive definiteness of paired symmetric tensors and elasticity tensors. Journal of Computational and Applied Mathematics, 338(??):22–43, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300542.

Hashemizadeh:2015:NSH

[HR15] E. Hashemizadeh and M. Rostami. Numerical solution of Hammerstein integral equations of mixed type using the sinc-collocation method. Journal of Computational and Applied Mathematics, 279(??):31– 39, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004658.

Hintermuller:2016:DBP

[HR16] M. Hintermüller and S. Rösel.

A duality-based pathfollowing semismooth Newton method for elastoplastic contact problems. Journal of Computational and Applied Mathematics, 292(??):150–173, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003325.

Hu:2017:NFE

[HRGZ17] Xiaozhe Hu, Carmen Rodrigo, Francisco J. Gaspar, and Ludmil T. Zikatanov. A nonconforming finite element method for the Biot’s consolidation model in poroelasticity. Journal of Computational and Applied Mathematics, 310(??):143–154, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302734.

Huijskens:2016:ENF

[HRO16] T. P. Huijskens, M. J.

Ruijter, and C. W. Oosterlee. Efficient numerical Fourier methods for coupled forward-backward SDEs. Journal of Computational and Applied Mathematics, 296(??):593–612, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005191.

Hochstenbach:2015:RPD

[HRR15] M. E. Hochstenbach, L. Reichel, and G. Rodriguez. Regularization parameter determination for discrete ill-posed problems. Journal of Computational and Applied Mathematics, 273(??):132–149, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271400274X.

Hernandez:2019:ZFT

[HRRC19] Gregorio Hernández, José Ranilla, and Sandra Ranilla-Cortina. Zero forcing in triangulations. Journal of Computational and Applied Mathematics, 354 (??):123–130, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718303327.

Haslinger:2016:RIM

[HRS16] Jaroslav Haslinger, Sergey Repin, and Stanislav Sysala. A reliable incremental method of computing the limit load in deformation plasticity based on compliance: Continuous and discrete setting. Journal of Computational and Applied Mathematics, 303(??):156–170, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300917.

Hasegawa:2017:UFM

[HS17a] Takemitsu Hasegawa and Hiroshi Sugiura. A userfriendly method for computing indefinite integrals of oscillatory functions. Journal of Computational and Applied Mathematics, 315(??):126–141, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305234.

Hesameddini:2017:SSV

[HS17b] Esmail Hesameddini and Mehdi Shahbazi. Solving system of Volterra– Fredholm integral equations with Bernstein polynomials and hybrid Bernstein block–pulse functions. Journal of Computational and Applied Mathematics, 315(??):182–194, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305283.

Huang:2017:MSS

[HS17c] Zhuo-Hong Huang and Hong Su. A modified shiftsplitting method for nonsymmetric saddle point problems. Journal of Computational and Applied Mathematics, 317(??):535– 546, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630574X.

Hao:2018:PBL

[HS18a] Ruru Hao and Zhixun Su. A patch-based lowrank tensor approximation model for multiframe image denoising. Journal of Computational and Applied Mathematics, 329(??):125–133, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300432.

Hasegawa:2018:UAC

[HS18b] Takemitsu Hasegawa and Hiroshi Sugiura. Uniform approximation to Cauchy principal value integrals with logarithmic singularity. Journal of Computational and Applied Mathematics, 327(??):1–11, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302947.

He:2018:MFE

[HS18c] Mingyan He and Pengtao Sun. Mixed finite element analysis for the Poisson–Nernst–Planck/ Stokes coupling. Journal of Computational and Applied Mathematics, 341(??):61–79, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301705.

Hasegawa:2019:UAF

[HS19] Takemitsu Hasegawa and Hiroshi Sugiura. Uniform approximation to finite Hilbert transform of oscillatory functions and its algorithm. Journal of Computational and Applied Mathematics, 358(??):327–342, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300858.

Heydari:2019:IMK

[HSAA19] Mojgan Heydari, Elyas Shivanian, Babak Azarnavid, and Saeid Abbasbandy. An iterative multistep kernel based method for nonlinear Volterra integral and integro-differential equations of fractional order. Journal of Computational and Applied Mathematics, 361(??):97–112, December 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0377042719302031.

Hazarika:2018:ESI

[HSAR18] Bipan Hazarika, H. M. Srivastava, Reza Arab, and Mohsen Rabbani. Existence of solution for an infinite system of nonlinear integral equations via measure of noncompactness and homotopy perturbation method to solve it. Journal of Computational and Applied Mathematics, 343(??):341–352, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302814.

Sun:2017:OCW

[hShX17] Xian hang Sun and Ming hai Xu. Optimal control of water flooding reservoir using proper orthogonal decomposition. Journal of Computational and Applied Mathematics, 320: 120–137, August 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Heber:2019:ASS

[HSS19] Frederik Heber, Frank Schöpfer, and Thomas Schuster. Acceleration of sequential subspace optimization in Banach spaces by orthogonal search directions. Journal of Computational and Applied Mathematics, 345:1–22, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

He:2015:TGC

[HSWH15] Mingyan He, Pengtao Sun, Cheng Wang, and Ziping Huang. A two-grid combined finite elementupwind finite volume method for a nonlinear convection-dominated diffusion reaction equation. Journal of Computational and Applied Mathematics, 288(??):223–232, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002484.

Hozman:2018:DFP

[HT18] Jirí Hozman and Tomás Tichý. DG framework for pricing European options under one-factor stochastic volatility models. Journal of Computational and Applied Mathematics, 344 (??):585–600, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303479.

Hu:2016:EAC

[Hu16]  Qianqian Hu.  Ex-

plicit G^{1} approximation of conic sections using Bézier curves of arbitrary degree. Journal of Computational and Applied Mathematics, 292(??):505–512, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003787.

Huang:2019:ASP

[Hua19] Rong Huang. Accurate solutions of product linear systems associated with rank-structured matrices. Journal of Computational and Applied Mathematics, 347(??):108–127, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304527.

Huillet:2016:MLD

[Hui16] Thierry E. Huillet. On Mittag-Leffler distributions and related stochastic processes. Journal of Computational and Applied Mathematics, 296 (??):181–211, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715004835.

Hernandez-Veron:2018:IAS

[HVM18] M. A. Hernández-Verón and Eulalia Martínez. Improving the accessibility of Steffensen’s method by decomposition of operators. Journal of Computational and Applied Mathematics, 330(??):536–552, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304569.

Hernandez-Veron:2019:DLC

[HVMR19] M. A. Hernández-Verón, Á.A. Magreñán, and M. J. Rubio. Dynamics and local convergence of a family of derivative-free iterative processes. Journal of Computational and Applied Mathematics, 354 (??):414–430, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718305193.

Haelterman:2015:ASP

[HVV15] R. Haelterman, D. Van Eester, and D. Verleyen. Accelerating the solution of a physics model inside a tokamak using the (inverse) column updating method. Journal of Computational and Applied Mathematics, 279(??):133– 144, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004762.

Huang:2017:NNS

[HW17] Shuai Huang and Zhong Wan. A new nonmonotone spectral residual method for nonsmooth nonlinear equations. Journal of Computational and Applied Mathematics, 313(??):82–101, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304344.

He:2018:PRE

[HW18a] Xubiao He and Jiayue Wang. Pricing real estate index options by compactly supported radialpolynomial basis point interpolation. Journal of Computational and Applied Mathematics, 333 (??):350–361, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717305666.

Huang:2018:MCS

[HW18b] Cong Huang and Rong Wang. Mapped CWENO scheme for hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 344 (??):229–244, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718303030.

Han:2016:WLB

[HWC16a] Li-Wen Han, Ya-Sha Wu, and Ying Chu. Weighted Lupaš q-Bézier curves. Journal of Computational and Applied Mathematics, 308(??):318–329, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302886.

Hu:2016:MFO

[HWC16b] Fuxing Hu, Rong Wang, and Xueyong Chen. A modified fifth-order WENOZ method for hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 303(??):56–68, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300747.

Hao:2015:PPM

[HWCC15] Zijun Hao, Zhongping Wan, Xiaoni Chi, and Jiawei Chen. A power penalty method for secondorder cone nonlinear complementarity problems. Journal of Computational and Applied Mathematics, 290(??):136–149, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002812.

Hong:2019:AMC

[HWH19a] Zhimin Hong, Yanjuan Wang, and Hui Hao. Adaptive Monte Carlo methods for solving hyperbolic telegraph equation. Journal of Computational and Applied Mathematics, 345:405–415, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Hu:2019:IAL

[HWH19b] Shenglong Hu, Jie Wang, and Zheng-Hai Huang. An inexact augmented Lagrangian multiplier method for solving quadratic complementary problems: an adapted algorithmic framework combining specific resolution techniques. Journal of Computational and Applied Mathematics, 361(??):64–78, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302080.

Hoang:2015:PMP

[HWP^{+}15] T.-V. Hoang, L. Wu, S. Paquay, A. C. Obreja, R. C. Voicu, R. Müller, J.- C. Golinval, and L. Noels. A probabilistic model for predicting the uncertainties of the humid stiction phenomenon on hard materials. Journal of Computational and Applied Mathematics, 289(??):173–195, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000928.

He:2015:SPB

[HWT15] Suxiang He, Min Wei, and Hengqing Tong. A smooth penalty-based sample average approximation method for stochastic complementarity problems. Journal of Computational and Applied Mathematics, 287(??):20–31, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001636.

Huang:2019:GDS

[HWXC19a] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. The generalized double steps scale-SOR iteration method for solving complex symmetric linear systems. Journal of Computational and Applied Mathematics, 346(??):284–306, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304424. See

retraction notice [HWXC19b].

Huang:2019:RNN

[HWXC19b] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. Retraction notice to: The generalized double steps scale-SOR iteration method for solving complex symmetric linear systems [CAM 346 (2018) 284–306]. Journal of Computational and Applied Mathematics, 360(??):170, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302110. See

[HWXC19a].

Hao:2018:MLM

[HWZ18a] Yongle Hao, Xiaoshen Wang, and Kai Zhang. Multi-level Monte Carlo weak Galerkin method with nested meshes for stochastic Brinkman problem. Journal of Computational and Applied Mathematics, 330(??):214–227, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304119.

Huang:2018:ENL

[HWZ18b] Shuai Huang, Zhong Wan, and Jing Zhang. An extended nonmonotone line search technique for largescale unconstrained optimization. Journal of Computational and Applied Mathematics, 330(??):586– 604, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717304570.

He:2019:SDS

[HWZ19] Zhuo-Heng He, Qing-Wen Wang, and Yang Zhang. A simultaneous decomposition for seven matrices with applications. Journal of Computational and Applied Mathematics, 349(??):93–113, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0377042718305466.

He:2015:IAE

[HX15] Guo He and Shuhuang Xiang. An improved algorithm for the evaluation of Cauchy principal value integrals of oscillatory functions and its application. Journal of Computational and Applied Mathematics, 280(??):1– 13, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005020.

Hu:2019:RMT

[HXC19] Xiaoping Hu, Ying Xiu, and Jie Cao. Recombined multinomial tree based on saddle-point approximation and its application to Lévy models options pricing. Journal of Computational and Applied Mathematics, 346(??):432–439, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304370.

He:2016:CCM

[HXX16] Guo He, Shuhuang Xiang, and Zhenhua Xu. A Chebyshev collocation method for a class of Fredholm integral equations with highly oscillatory kernels. Journal of Computational and Applied Mathematics, 300(??):354– 368, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006391.

Hu:2019:ASP

[HY19a] Shengrong Hu and Wenjun Yang. Add-sub pivoting triangular factorization for symmetric matrix. Journal of Computational and Applied Mathematics, 355(??):116–127, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300159.

Huang:2019:UAC

[HY19b] Yuxia Huang and Chuancun Yin. A unifying approach to constrained and unconstrained optimal reinsurance. Journal of Computational and Applied Mathematics, 360(??):1–17, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/

science/article/pii/
S0377042719301773.

Hajrajabi:2018:NAM

[HYF18] A. Hajrajabi, A. R. Yazdanian, and R. Farnoosh. Nonlinear autoregressive model with stochastic volatility innovations: Semiparametric and Bayesian approach. Journal of Computational and Applied Mathematics, 344(??):37– 46, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718303133.

Haoran:2018:SLA

[HYQ^{+}18] Zhang Haoran, Liang Yongtu, Liao Qi, Shen Yun, and Yan Xiaohan. A self-learning approach for optimal detailed scheduling of multiproduct pipeline. Journal of Computational and Applied Mathematics, 327(??):41–63, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302935.

Huang:2016:OIP

[HYZ16] Ya Huang, Xiangqun Yang, and Jieming Zhou. Optimal investment and proportional reinsurance for a jump-diffusion risk model with constrained control variables. Journal of Computational and Applied Mathematics, 296 (??):443–461, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715004847.

Hu:2017:FSE

[HZ17a] Jiashun Hu and Chunxiong Zheng. Fast and stable evaluation of the exact absorbing boundary condition for the semidiscrete linear Schrödinger equation in unbounded domains. Journal of Computational and Applied Mathematics, 326 (??):116–125, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302546.

Huang:2017:QDB

[HZ17b] Jiting Huang and Peixin Zhao. QR decomposition based orthogonality estimation for partially linear models with longitudinal data. Journal of Computational and Applied Mathematics, 321(??):406–415, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300894.

Huang:2017:IAS

[HZ17c] Xiaojin Huang and Detong Zhu. An interior affine scaling cubic regularization algorithm for derivativefree optimization subject to bound constraints. Journal of Computational and Applied Mathematics, 321(??):108–127, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300936.

Hadjidimos:2018:CTC

[HZ18a] Apostolos Hadjidimos and Li-Li Zhang. Comparison of three classes of algorithms for the solution of the linear complementarity problem with an H^{+}- matrix. Journal of Computational and Applied Mathematics, 336(??):175– 191, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306568.

He:2018:CFP

[HZ18b] Xin-Jiang He and Song-Ping Zhu. A closedform pricing formula for European options under the Heston model with stochastic interest rate. Journal of Computational and Applied Mathematics, 335(??):323– 333, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717306313. See

note [RZ19b].

Hou:2019:IVA

[HZ19] Guoliang Hou and Shugong Zhang. An improved verification algorithm for nonlinear systems of equations based on Krawczyk operator. Journal of Computational and Applied Mathematics, 359(??):145–152, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301591.

Iyiola:2018:RDP

[IAAW18] O. S. Iyiola, E. O. Asante-Asamani, and B. A. Wade. A real distinct poles rational approximation of generalized Mittag-Leffler functions and their inverses: Applications to fractional calculus. Journal of Computational and Applied Mathematics, 330(??):307–317, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304090.

Ilaya-Ayza:2017:MCO

[IABIPG17] Amilkar E. Ilaya-Ayza, Julio Benítez, Joaquín Izquierdo, and Rafael Pérez-García. Multicriteria optimization of supply schedules in intermittent water supply systems. Journal of Computational and Applied Mathematics, 309(??):695–703, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302266.

Ilaya-Ayza:2018:GTI

[IAMCI18] Amilkar E. Ilaya-Ayza, Carlos Martins, Enrique Campbell, and Joaquín Izquierdo. Gradual transition from intermittent to continuous water supply based on multi-criteria optimization for network sector selection. Journal of Computational and Applied Mathematics, 330(??):1016–1029, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301954.

Ito:2015:OPA

[IANS15] Shinji Ito, Kensuke Aishima, Takaaki Nara, and Masaaki Sugihara. Orthogonal polynomial approach to estimation of poles of rational functions from data on open curves. Journal of Computational and Applied Mathematics, 273(??):326–345, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002799.

Izgi:2018:SIS

[IÇ18] Burhaneddin Izgi and Coskun Çetin. Semiimplicit split-step numerical methods for a class of nonlinear stochastic differential equations with non-Lipschitz drift terms. Journal of Computational and Applied Mathematics, 343(??):62–79, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0377042718301900.

Izquierdo:2016:IPD

[ICMPG16] Joaquín Izquierdo, Enrique Campbell, Idel Montalvo, and Rafael Pérez-García. Injecting problemdependent knowledge to improve evolutionary optimization search ability. Journal of Computational and Applied Mathematics, 291(??):281–292, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500165X.

Izgi:2016:EVA

[ID16] Burhaneddin Izgi and Ahmet Duran. 3D extreme value analysis for stock return, interest rate and speed of mean reversion. Journal of Computational and Applied Mathematics, 297(??):51–64, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005087.

Ilati:2018:EAM

[ID18] Mohammad Ilati and Mehdi Dehghan. Error analysis of a meshless weak form method based on radial point interpolation technique for Sivashinsky equation arising in the alloy solidification problem. Journal of Computational and Applied Mathematics, 327(??):314–324, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303229.

Iampietro:2018:MSI

[IDGH18] D. Iampietro, F. Daude, P. Galon, and J.-M. Hérard. A Mach-sensitive implicit-explicit scheme adapted to compressible multi-scale flows. Journal of Computational and Applied Mathematics, 340(??):122–150, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300980.

Illan-Gonzalez:2017:EFP

[IGRL17] J. Illán-González and J. M. Rebollido-Lorenzo. Evaluation of finite part integrals using a regularization technique that decreases instability. Journal of Computational and Applied Mathematics, 319(??):210–219, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300134.

Iqbal:2015:LMM

[IIA15] Javed Iqbal, Asif Iqbal, and Muhammad Arif. Levenberg–Marquardt method for solving systems of absolute value equations. Journal of Computational and Applied Mathematics, 282(??):134–138, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042714005433.

Iiduka:2015:COF

[Iid15]  Hideaki Iiduka.  Con-

vex optimization over fixed point sets of quasinonexpansive and nonexpansive mappings in utility-based bandwidth allocation problems with operational constraints. Journal of Computational and Applied Mathematics, 282(??):225–236, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042715000175.

Izzo:2018:SSP

[IJ18] Giuseppe Izzo and Zdzislaw Jackiewicz. Strong stability preserving transformed DIMSIMs. Journal of Computational and Applied Mathematics, 343(??):174–188, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830164X.

Itikawa:2015:PPU

[IL15] Jackson Itikawa and Jaume Llibre. Phase portraits of uniform isochronous quartic centers. Journal of Computational and Applied Mathematics, 287(??):98–114, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001235.

Iqbal:2016:WCE

[IN16] Tanveer Iqbal and Arnold Neumaier. Worst case error bounds for the solution of uncertain Poisson equations with mixed boundary conditions. Journal of Computational and Applied Mathematics, 303(??):40–55, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300929.

intHout:2016:CMC

[iW16] Karel J. in ’t Hout and Maarten Wyns. Convergence of the modified Craig–Sneyd scheme for two-dimensional convection– diffusion equations with mixed derivative term. Journal of Computational and Applied Mathematics, 296(??):170–180, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042715004756.

Jovan:2017:DPM

[JA17] Matej Jovan and Ales Ahcan. Default prediction with the Merton-type structural model based on the NIG Lévy process. Journal of Computational and Applied Mathematics, 311(??):414–422, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303673.

Jando:2017:EGO

[Jan17a] Dörte Jando. Efficient goal-oriented global error estimators for BDF methods using discrete adjoints. Journal of Computational and Applied Mathematics, 316(??):195–212, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304526.

Jaradat:2017:ITD

[JAN17b] Ali Jaradat, Fadi Awawdeh, and Mohd Salmi Md Noorani. Identification of time-dependent source terms and control parameters in parabolic equations from overspecified boundary data. Journal of Computational and Applied Mathematics, 313(??):397–409, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304496.

Jenkinson:2018:HOA

[JB18] M. J. Jenkinson and J. W. Banks. High-order accurate FDTD schemes for dispersive Maxwell’s equations in second-order form using recursive convolutions. Journal of Computational and Applied Mathematics, 336(??):192– 218, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306362.

Jenkins:2015:ENM

[JBFS15] S. E. Jenkins, C. J.

Budd, M. A. Freitag, and N. D. Smith. The effect of numerical model error on data assimilation. Journal of Computational and Applied Mathematics, 290(??):567–588, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003143.

Javadi:2016:SGP

[JBT16] Shahnam Javadi, Esmail Babolian, and Zeinab Taheri. Solving generalized pantograph equations by shifted orthonormal Bernstein polynomials. Journal of Computational and Applied Mathematics, 303(??):1–14, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300723.

Jimenez:2015:CRW

[JC15] J. C. Jimenez and F. Carbonell. Convergence rate of weak local linearization schemes for stochastic differential equations with additive noise. Journal of Computational and Applied Mathematics, 279 (??):106–122, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714004671.

Jerez:2018:BMM

[JC18] Silvia Jerez and Ariel Camacho. Bone metastasis modeling based on the interactions between the BMU and tumor cells. Journal of Computational and Applied Mathematics, 330(??):866–876, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306367.

Jerez:2019:SME

[JC19] Silvia Jerez and J. Adrián Cantó. A stochastic model for the evolution of bone metastasis: Persistence and recovery. Journal of Computational and Applied Mathematics, 347(??):12–23, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304734.

Jamil:2015:HAA

[JCC15] Noreen Jamil, Xuemei Chen, and Alexander Cloninger. Hildreth’s algorithm with applications to soft constraints for user interface layout. Journal of Computational and Applied Mathematics, 288(??):193–202, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002320.

Jackson:2017:MPV

[JCC17] Mark Jackson and Benito M. Chen-Charpentier. Modeling plant virus propagation with delays. Journal of Computational and Applied Mathematics, 309(??):611–621, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301972.

Jackson:2018:MBC

[JCC18] Mark Jackson and Benito M. Chen-Charpentier. A model of biological control of plant virus propagation with delays. Journal of Computational and Applied Mathematics, 330(??):855–865, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300092.

Jackson:2019:MPV

[JCC19] Mark Jackson and Benito M. Chen-Charpentier. Modeling plant virus propagation with seasonality. Journal of Computational and Applied Mathematics, 345:310–319, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Jia:2017:CSF

[JCH17] Zehui Jia, Xingju Cai, and Deren Han. Comparison of several fast algorithms for projection onto an ellipsoid. Journal of Computational and Applied Mathematics, 319(??):320–337, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300122.

Jin:2016:SFI

[JD16] Congming Jin and Jiu Ding. Solving Fredholm integral equations via a piecewise linear maximum entropy method. Journal of Computational and Applied Mathematics, 304(??):130–137, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301005.

Jebreen:2019:GRB

[Jeb19] Haifa Bin Jebreen. A Gaussian radial basis function-finite difference technique to simulate the HCIR equation. Journal of Computational and Applied Mathematics, 347(??):181–195, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304965.

Jeon:2015:HSU

[Jeo15]  Youngmok Jeon.  Hy-

bridized SUPG and upwind numerical schemes for convection dominated diffusion problems. Journal of Computational and Applied Mathematics, 275(??):91–99, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042714003628.

Jimenez-Gamero:2019:AND

[JGAFAL19] M. D. Jiménez-Gamero, M. V. Alba-Fernández, and F. J. Ariza-López. Approximating the null distribution of a class of statistics for testing independence. Journal of Computational and Applied Mathematics, 354 (??):131–143, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718301572.

Jones:2016:CED

[JGGS16] Tiffany Jones, Leonel P. Gonzalez, Shekhar Guha, and Qin Sheng. A continuing exploration of a decomposed compact method for highly oscillatory wave problems. Journal of Computational and Applied Mathematics, 299(??):207– 220, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715006044.

Jian:2015:NSC

[JGTB15] Jin-Bao Jian, Chuan-Hao Guo, Chun-Ming Tang, and Yan-Qin Bai. A new superlinearly convergent algorithm of combining QP subproblem with system of linear equations for nonlinear optimization. Journal of Computational and Applied Mathematics, 273(??):88–102, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002829.

Jafari:2017:NAS

[JHJB17] H. Jafari, A. Haghbin, S. J. Johnston, and D. Baleanu. A new algorithm for solving dynamic equations on a time scale. Journal of Computational and Applied Mathematics, 312(??):167–173, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301030.

Jeon:2017:VAF

[JHK17] Junkee Jeon, Heejae Han, and Myungjoo Kang. Valuing American floating strike lookback option and Neumann problem for inhomogeneous Black–Scholes equation. Journal of Computational and Applied Mathematics, 313(??):218–234, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630440X.

Jerez-Hanckes:2015:LMT

[JHPT15] Carlos Jerez-Hanckes, José Pinto, and Simon Tournier. Local multiple traces formulation for highfrequency scattering problems. Journal of Computational and Applied Mathematics, 289(??):306–321, December 1, 2015. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715000102.

Jin:2016:LAH

[Jin16] Yanyi Jin. Local analysis of HDG methods for convection-dominated diffusion problems. Journal of Computational and Applied Mathematics, 304(??):57–72, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301340.

Jiang:2019:IFR

[JJ19] Xianzhen Jiang and Jinbao Jian. Improved Fletcher–Reeves and Dai– Yuan conjugate gradient methods with the strong Wolfe line search. Journal of Computational and Applied Mathematics, 348 (??):525–534, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042718305570.

Jiang:2017:REV

[JJX17] Yunlu Jiang, Qinghua Ji, and Baojian Xie. Robust estimation for the varying coefficient partially nonlinear models. Journal of Computational and Applied Mathematics, 326(??):31–43, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730198X.

Jeong:2018:EHF

[JK18] Darae Jeong and Junseok Kim. An explicit hybrid finite difference scheme for the Allen–Cahn equation. Journal of Computational and Applied Mathematics, 340(??):247–255, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718301055.

Jia:2019:TAM

[JK19] Zhongxiao Jia and Wenjie Kang. A transformation approach that makes SPAI, PSAI and RSAI procedures efficient for large double irregular nonsymmetric sparse linear systems. Journal of Computational and Applied Mathematics, 348(??):200–213, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830520X.

Jiang:2015:CNL

[JKL^{+}15] Nan Jiang, Michaela Kubacki, William Layton, Marina Moraiti, and Hoang Tran. A Crank– Nicolson leapfrog stabilization: Unconditional stability and two applications. Journal of Computational and Applied Mathematics, 281(??):263– 276, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004336.

Jia:2015:HRH

[JL15] Zhongxiao Jia and Cen Li. Harmonic and refined harmonic shift-invert residual Arnoldi and Jacobi– Davidson methods for interior eigenvalue problems. Journal of Computational and Applied Mathematics, 282(??):83– 97, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000084.

Jeong:2016:WNL

[JL16a] SeongHee Jeong and Eunjung Lee. Weighted norm least squares finite element method for Poisson equation in a polyhedral domain. Journal of Computational and Applied Mathematics, 299(??):35– 49, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005117.

Jiang:2016:RMF

[JL16b] Lijian Jiang and Qiuqi Li. Reduced multiscale finite element basis methods for elliptic PDEs with parameterized inputs. Journal of Computational and Applied Mathematics, 301(??):101–120, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300127.

Ji:2017:BCG

[JL17] Hao Ji and Yaohang Li. Block conjugate gradient algorithms for least squares problems. Journal of Computational and Applied Mathematics, 317 (??):203–217, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716305738.

Ji:2019:NMT

[JL19a] Hangjie Ji and Longfei Li. Numerical methods for thermally stressed shallow shell equations. Journal of Computational and Applied Mathematics, 362(??):626–652, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830606X.

Jiang:2019:INS

[JL19b] Yongyue Jiang and Ping Lin. Improvement of numerical simulation for GMAW based on a new model with thermocapillary effect. Journal of Computational and Applied Mathematics, 356(??):37–45, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S037704271930055X.

Jing:2017:NAC

[JLCZ17] Feifei Jing, Jian Li, Zhangxin Chen, and Zhonghua Zhang. Numerical analysis of a characteristic stabilized finite element method for the time-dependent Navier–Stokes equations with nonlinear slip boundary conditions. Journal of Computational and Applied Mathematics, 320: 43–60, August 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Jin:2016:SFE

[JLLZ16] Bangti Jin, Raytcho Lazarov, Xiliang Lu, and Zhi Zhou. A simple finite element method for boundary value problems with a Riemann–Liouville derivative. Journal of Computational and Applied Mathematics, 293(??):94–111, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001454.

Jari:2015:SFE

[JM15] Rabeea Jari and Lin Mu. Superconvergence of H(div) finite element approximations for the Stokes problem by local L^{2}-projection methods. Journal of Computational and Applied Mathematics, 278(??):278–292, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004543.

Jator:2018:BNT

[JM18] S. N. Jator and V. Manathunga. Block Nyström type integrator for Bratu’s equation. Journal of Computational and Applied Mathematics, 327(??):341– 349, January 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717303254.

Jamil:2016:ELR

[JMN^{+}16] Noreen Jamil, Johannes Müller, M. Asif Naeem, Christof Lutteroth, and Gerald Weber. Extending linear relaxation for non-square matrices and soft constraints. Journal of Computational and Applied Mathematics, 308 (??):346–360, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302230.

Jamil:2018:RIO

[JMNB18] Noreen Jamil, Farhaan Mirza, M. Asif Naeem, and Nilufar Baghaei. A refinement of an iterative orthogonal projection method. Journal of Computational and Applied Mathematics, 341(??):31–41, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301043.

Jimenez:2017:WLL

[JMS17] J. C. Jimenez, C. Mora, and M. Selva. A weak local linearization scheme for stochastic differential equations with multiplicative noise. Journal of Computational and Applied Mathematics, 313(??):202–217, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304332.

Jian:2019:LEP

[JMSL19] Ling Jian, Xiaoyu Ma, Yunquan Song, and Shihua Luo. Laplace error penalty-based M-type model detection for a class of high dimensional semiparametric models. Journal of Computational and Applied Mathematics, 347(??):210–221, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304977.

Jung:2018:NAW

[JN18] Chang-Yeol Jung and Thien Binh Nguyen. A new adaptive weighted essentially non-oscillatory WENO-θ scheme for hyperbolic conservation laws. Journal of Computational and Applied Mathematics, 328(??):314–339, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303606.

Jiang:2019:EEM

[JN19] Xiao Jiang and Saralees Nadarajah. Efficient expressions for moments of dependent random sums using copulas. Journal of Computational and Applied Mathematics, 353(??):130– 139, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307684.

Jiang:2017:MMR

[JO17] Lijian Jiang and Na Ou.

Multiscale model reduction method for Bayesian inverse problems of subsurface flow. Journal of Computational and Applied Mathematics, 319(??):188–209, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300110.

Jin:2016:PDF

[JQWW16] Zhuo Jin, Linyi Qian, Wei Wang, and Rongming Wang. Pricing dynamic fund protections with regime switching. Journal of Computational and Applied Mathematics, 297(??):13–25, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005579.

Jagels:2015:CGQ

[JR15] Carl Jagels and Lothar Reichel. On the computation of Gauss quadrature rules for measures with a monomial denominator. Journal of Computational and Applied Mathematics, 286(??):102–113, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001120.

Jagels:2017:GAS

[JRT17] Carl Jagels, Lothar Reichel, and Tunan Tang. Generalized averaged Szegő quadrature rules. Journal of Computational and Applied Mathematics, 311(??):645–654, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304071.

Jax:2017:GRT

[JS17] T. Jax and G. Steinebach. Generalized ROW-type methods for solving semiexplicit DAEs of index-1. Journal of Computational and Applied Mathematics, 316(??):213–228, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303934.

Jeon:2019:FHP

[JS19a] Junkee Jeon and Yong Hyun Shin. Finite horizon portfolio selection with a negative wealth constraint. Journal of Computational and Applied Mathematics, 356(??):329–338, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300792.

Jones:2019:NSS

[JS19b] Tiffany Jones and Qin Sheng. Numerical stabilities study of a decomposed compact method for highly oscillatory Helmholtz equations. Journal of Computational and Applied Mathematics, 354(??):334– 347, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301948.

Jia:2018:STA

[JSC18] Xiaohong Jia, Xiaoran Shi, and Falai Chen. Survey on the theory and applications of μ-bases for rational curves and surfaces. Journal of Computational and Applied Mathematics, 329(??):2–23, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303746.

Jia:2015:GST

[JSL15] Jiteng Jia, Tomohiro Sogabe, and Sumei Li. A generalized symbolic Thomas algorithm for the solution of opposite-bordered tridiagonal linear systems. Journal of Computational and Applied Mathematics, 290(??):423–432, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003246.

Jones:2017:CIT

[JSW17] Aquil D. Jones, Gideon Simpson, and William Wilson. Conservative integrators for a toy model of weak turbulence. Journal of Computational and Applied Mathematics, 325(??):113–124, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042717302091. See

erratum [JSW18b].

Jiang:2018:PEV

[JSW18a] Yiming Jiang, Shiyu Song, and Yongjin Wang. Pricing European vanilla options under a jump-to-default threshold diffusion model. Journal of Computational and Applied Mathematics, 344 (??):438–456, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302346.

Jones:2018:ECI

[JSW18b] Aquil D. Jones, Gideon Simpson, and William Wilson. Erratum to “Conservative integrators for a toy model of weak turbulence” [j. comput. appl. math. 325 (2017) 113– 124]. Journal of Computational and Applied Mathematics, 328(??):214–215, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303552. See

[JSW17].

Jodar:2018:MMC

[JTCC18] Lucas Jodar, Juan R. Torregrosa, Juan C. Cortés, and Regino Criado. Mathematical modeling and computational methods. Journal of Computational and Applied Mathematics, 330(??):661–665, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S037704271730506X.

Jiang:2017:SQO

[JW17] Yingchun Jiang and Suping Wang. Sampling and quasi-optimal approximation for signals in a reproducing kernel space of homogeneous type. Journal of Computational and Applied Mathematics, 319(??):296–307, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300353.

Jiao:2018:MSC

[JWQ18] Xue Jiao, Tieru Wu, and Xuzhou Qin. Mesh segmentation by combining mesh saliency with spectral clustering. Journal of Computational and Applied Mathematics, 329(??):134–146, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302431.

Jia:2018:NRS

[JWZC18] Zhigang Jia, Musheng Wei, Mei-Xiang Zhao, and Yong Chen. A new real structure-preserving quaternion QR algorithm. Journal of Computational and Applied Mathematics, 343(??):26–48, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301870.

Ji:2017:TES

[JY17] Yanting Ji and Chenggui Yuan. Tamed EM scheme of neutral stochastic differential delay equations. Journal of Computational and Applied Mathematics, 326(??):337–357, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303023.

Jeong:2019:ANS

[JY19] Byeongseon Jeong and Jungho Yoon. Analysis of non-stationary Hermite subdivision schemes reproducing exponential polynomials. Journal of Computational and Applied Mathematics, 349(??):452– 469, March 15, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718304771.

Jo:2016:CLI

[JYK16] Sihun Jo, Minsuk Yang, and Geonwoo Kim. On convergence of Laplace inversion for the American put option under the CEV model. Journal of Computational and Applied Mathematics, 305(??):36–43, October 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301571.

Jeon:2017:PVP

[JYK17] Junkee Jeon, Ji-Hun Yoon, and Myungjoo Kang. Pricing vulnerable pathdependent options using integral transforms. Journal of Computational and Applied Mathematics, 313(??):259–272, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304447.

Jeon:2018:PDF

[JYP18] Junkee Jeon, Ji-Hun Yoon, and Chang-Rae Park. The pricing of dynamic fund protection with default risk. Journal of Computational and Applied Mathematics, 333(??):116–130, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305320.

Jia:2018:SCN

[JZ18] Zhi-Gang Jia and Mei-Xiang Zhao. A structured condition number for self-adjoint polynomial matrix equations with applications in linear control. Journal of Computational and Applied Mathematics, 331(??):208–216, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304879.

Jiang:2016:DBI

[JZS16] Xue Jiang, Shugong Zhang, and Baoxin Shang. The discretization for bivariate ideal interpolation. Journal of Computational and Applied Mathematics, 308(??):177–186, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302291.

Kudu:2016:FDM

[KAA16] Mustafa Kudu, Ilhame Amirali, and Gabil M. Amiraliyev. A finitedifference method for a singularly perturbed delay integro-differential equation. Journal of Computational and Applied Mathematics, 308 (??):379–390, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302904.

Kadioglu:2017:ASC

[Kad17] Samet Y. Kadioglu. Analysis of the self-consistent IMEX method for tightly coupled non-linear systems. Journal of Computational and Applied Mathematics, 322(??):148–160, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301413.

Kamocki:2016:NRF

[Kam16a] Rafal Kamocki. A new representation formula for the Hilfer fractional derivative and its application. Journal of Computational and Applied Mathematics, 308(??):39–45, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630231X.

Kim:2016:CSB

[KAM16b] Jeongho Kim, James F.

Antaki, and Mehrdad Massoudi. Computational study of blood flow in microchannels. Journal of Computational and Applied Mathematics, 292(??):174–187, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003398.

Kamalja:2017:RCM

[Kam17] Kirtee K. Kamalja. Reliability computing method for generalized k-out-of-n system. Journal of Computational and Applied Mathematics, 323 (??):111–122, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042717301152.

Kang:2018:NIO

[Kan18] Hongchao Kang. Numerical integration of oscillatory Airy integrals with singularities on an infinite interval. Journal of Computational and Applied Mathematics, 333(??):314– 326, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305691.

Kumar:2018:MNS

[KAS18] Devendra Kumar, Ravi P. Agarwal, and Jagdev Singh. A modified numerical scheme and convergence analysis for fractional model of Lienard’s equation. Journal of Computational and Applied Mathematics, 339(??):405– 413, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301309.

Kaur:2016:SIF

[Kau16] Amandeep Kaur. On solving an isospectral flow. Journal of Computational and Applied Mathematics, 308(??):263–275, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302679.

Kavlak:2016:MWL

[Kav16] Konul Bayramoğlu Kavlak.

The mean wasted life time of a component of system. Journal of Computational and Applied Mathematics, 305(??):44–54, October 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301558.

Kawai:2017:AIS

[Kaw17] Reiichiro Kawai. Adaptive importance sampling Monte Carlo simulation for general multivariate probability laws. Journal of Computational and Applied Mathematics, 319(??):440–459, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730050X.

Kazashi:2018:FDF

[Kaz18] Yoshihito Kazashi. A fully discretised filtered polynomial approximation on spherical shells. Journal of Computational and Applied Mathematics, 333 (??):428–441, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717305654.

Kumar:2016:BSQ

[KB16] Rakesh Kumar and S. Baskar. B-spline quasi-interpolation based numerical methods for some Sobolev type equations. Journal of Computational and Applied Mathematics, 292(??):41–66, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003374.

Kamrani:2017:PCN

[KB17] Minoo Kamrani and Dirk Blömker. Pathwise convergence of a numerical method for stochastic partial differential equations with correlated noise and local Lipschitz condition. Journal of Computational and Applied Mathematics, 323(??):123–135, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S037704271730167X.

Kosinka:2019:GQC

[KB19] Jirí Kosinka and Michael Barton. Gaussian quadrature for C^{1} cubic Clough– Tocher macro-triangles. Journal of Computational and Applied Mathematics, 351(??):6–13, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718306447.

Kooij:2017:BKS

[KBG17] G. L. Kooij, M. A.

Botchev, and B. J. Geurts. A block Krylov subspace implementation of the time-parallel Paraexp method and its extension for nonlinear partial differential equations. Journal of Computational and Applied Mathematics, 316(??):229–246, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304563.

Kui:2019:CFD

[KBL19] Zhiqing Kui, Jean Baccou, and Jacques Liandrat. Construction of finitedimensional multiresolutions based on linear subdivision schemes: Application to the 4-point shifted Lagrange scheme. Journal of Computational and Applied Mathematics, 349(??):494–507, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305387.

Kozlov:2015:DCS

[KBT15] Valery V. Kozlov, Alexander P. Buslaev, and Alexander G. Tatashev. A dynamical communication system on a network. Journal of Computational and Applied Mathematics, 275(??):247–261, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400346X.

Kanaujiya:2017:PEG

[KC17] Ankur Kanaujiya and Siddhartha P. Chakrabarty. Pricing and estimates of Greeks for passport option: a three time level approach. Journal of Computational and Applied Mathematics, 315(??):49– 64, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305167.

Kim:2019:GSP

[KC19] Kyung-Su Kim and Sae-Young Chung. Greedy subspace pursuit for joint sparse recovery. Journal of Computational and Applied Mathematics, 352(??):308–327, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307179.

Ku-Carrillo:2017:EOO

[KCDACC17] Roberto A. Ku-Carrillo, Sandra E. Delgadillo-Aleman, and Benito M. Chen-Charpentier. Effects of the obesity on optimal control schedules of chemotherapy on a cancerous tumor. Journal of Computational and Applied Mathematics, 309(??):603–610, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302278.

Karaaslan:2018:HDG

[KCK18] Mehmet Fatih Karaaslan, Fatih Celiker, and Muhammet Kurulay. A hybridizable discontinuous Galerkin method for a class of fractional boundary value problems. Journal of Computational and Applied Mathematics, 333 (??):20–27, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717304843.

Kundu:2016:SRM

[KCNH16] Amarjit Kundu, Shovan Chowdhury, Asok K. Nanda, and Nil Kamal Hazra. Some results on majorization and their applications. Journal of Computational and Applied Mathematics, 301(??):161–177, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000261.

Kim:2017:BAA

[KCX17] Hyea Hyun Kim, Eric T. Chung, and Chenxiao Xu. A BDDC algorithm with adaptive primal constraints for staggered discontinuous Galerkin approximation of elliptic problems with highly oscillating coefficients. Journal of Computational and Applied Mathematics, 311(??):599–617, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303971.

Khakimzyanov:2017:SPS

[KD17] Gayaz Khakimzyanov and Denys Dutykh. On supraconvergence phenomenon for second order centered finite differences on non-uniform grids. Journal of Computational and Applied Mathematics, 326(??):1–14, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730242X.

Kaciranlar:2019:OEB

[KD19] Selahattin Kaçıranlar and Issam Dawoud. The optimal extended balanced loss function estimators. Journal of Computational and Applied Mathematics, 345:86–98, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Kazemi:2017:AES

[KDB17] Seyed-Mohammad-Mahdi Kazemi, Mehdi Dehghan, and Ali Foroush Bastani. Asymptotic expansion of solutions to the Black–Scholes equation arising from American option pricing near the expiry. Journal of Computational and Applied Mathematics, 311(??):11–37, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303053.

Kazemi:2018:NFR

[KDB18] Seyed-Mohammad-Mahdi Kazemi, Mehdi Dehghan, and Ali Foroush Bastani. On a new family of radial basis functions: Mathematical analysis and applications to option pricing. Journal of Computational and Applied Mathematics, 328(??):75–100, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303126.

Kheybari:2017:SAAb

[KDW17a] S. Kheybari, M. T. Darvishi, and A. M. Wazwaz. A semi-analytical algorithm to solve systems of integro-differential equations under mixed boundary conditions. Journal of Computational and Applied Mathematics, 317 (??):72–89, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716305714.

Kheybari:2017:SAAa

[KDW17b] S. Kheybari, M. T.

Darvishi, and A. M. Wazwaz. A semi-analytical approach to solve integrodifferential equations. Journal of Computational and Applied Mathematics, 317 (??):17–30, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716305441.

Komori:2019:RMS

[KEB19] Yoshio Komori, Alexey Eremin, and Kevin Burrage. S-ROCK methods for stochastic delay differential equations with one fixed delay. Journal of Computational and Applied Mathematics, 353 (??):345–354, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718307738.

Khoveyni:2016:DEA

[KElY16] Mohammad Khoveyni, Robabeh Eslami, and Guo liang Yang. Data envelopment analysis in the absence of convexity: Specifying efficiency status and estimating returns to scale. Journal of Computational and Applied Mathematics, 304(??):172–200, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630139X.

Kautsky:2016:NPP

[KF16] Jaroslav Kautsky and Jan Flusser. Numerical problems with the Pascal triangle in moment computation. Journal of Computational and Applied Mathematics, 306(??):53–68, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301601.

Kassim:2017:NGS

[KFT17] M. D. Kassim, K. M. Furati, and N.-E. Tatar. Nonexistence of global solutions for a fractional differential problem. Journal of Computational and Applied Mathematics, 314(??):61– 68, April 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304952.

Kazemi:2019:NAS

[KGES19] Manochehr Kazemi, Hamid Mottaghi Golshan, Reza Ezzati, and Mohsen Sadatrasoul. New approach to solve two-dimensional Fredholm integral equations. Journal of Computational and Applied Mathematics, 354(??):66– 79, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830760X.

Kayasandik:2019:DMR

[KGL19] Cihan Kayasandik, Kanghui Guo, and Demetrio Labate. Directional multiscale representations and applications in digital neuron reconstruction. Journal of Computational and Applied Mathematics, 349(??):482–493, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830548X.

Kumar:2018:SCS

[KGMS18] Abhimanyu Kumar, D. K.

Gupta, Eulalia Martínez, and Sukhjit Singh. Semilocal convergence of a secant-type method under weak Lipschitz conditions in Banach spaces. Journal of Computational and Applied Mathematics, 330(??):732–741, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042717301073.

Khanh:2018:ABH

[KH18] Tra Quoc Khanh and Ngo Van Hoa. On the axisymmetric backward heat equation with non-zero right hand side: Regularization and error estimates. Journal of Computational and Applied Mathematics, 335(??):156– 167, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305964.

Khastan:2017:NSF

[Kha17] A. Khastan. New solutions for first order linear fuzzy difference equations. Journal of Computational and Applied Mathematics, 312(??):156–166, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301170.

Kohatsu-Higa:2017:WRC

[KHLY17] Arturo Kohatsu-Higa, Antoine Lejay, and Kazuhiro Yasuda. Weak rate of convergence of the Euler–Maruyama scheme for stochastic differential equations with nonregular drift. Journal of Computational and Applied Mathematics, 326 (??):138–158, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302510.

Kugler:2018:NAV

[KHS^{+}18] Michael Kugler, Alexandre Hostettler, Luc Soler, Yves Remond, and Daniel George. A new algorithm for volume mesh refinement on merging geometries: Application to liver and vascularisation. Journal of Computational and Applied Mathematics, 330(??):429–440, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304296.

Kheiri:2019:SAF

[KJ19] Hossein Kheiri and Mohsen Jafari. Stability analysis of a fractional order model for the HIV/AIDS epidemic in a patchy environment. Journal of Computational and Applied Mathematics, 346(??):323–339, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304163.

Kumar:2015:AOD

[KK15] Sunil Kumar and Mukesh Kumar. An analysis of overlapping domain decomposition methods for singularly perturbed reaction–diffusion problems. Journal of Computational and Applied Mathematics, 281(??):250– 262, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005743.

Kulikova:2017:NBA

[KK17] M. V. Kulikova and G. Yu. Kulikov. NIRK-based accurate continuous-discrete extended Kalman filters for estimating continuoustime stochastic target tracking models. Journal of Computational and Applied Mathematics, 316(??):260–270, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304058.

Kamarujjama:2019:CNC

[KK19a] M. Kamarujjama and Owais Khan. Computation of new class of integrals involving generalized Galue type Struve function. Journal of Computational and Applied Mathematics, 351(??):228–236, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306812.

Khan:2019:GCF

[KK19b] Tahir Ullah Khan and Muhammad Adil Khan. Generalized conformable fractional operators. Journal of Computational and Applied Mathematics, 346(??):378–389, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304382.

Kumar:2019:TFF

[KK19c] Pankaj Kumar and Jiten C. Kalita. A transformationfree ψ–v formulation of the Navier–Stokes equations on compact nonuniform grids. Journal of Computational and Applied Mathematics, 353(??):292– 317, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307660.

Krol:2016:ETP

[KKK16] M. Król, A. V. Kunynets, and M. V. Kutniv. Exact three-point difference scheme for singular nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 298(??):175–189, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005981.

Kitkuan:2019:AZT

[KKP^{+}19] Duangkamon Kitkuan, Poom Kumam, Anantachai Padcharoen, Wiyada Kumam, and Phatiphat Thounthong. Algorithms for zeros of two accretive operators for solving convex minimization problems and its application to image restoration problems. Journal of Computational and Applied Mathematics, 354(??):471– 495, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302528.

Khaliq:2016:RAN

[KKS16] Abdul Q. M. Khaliq, JaEun Ku, and Qin Sheng. Recent advances in numerical methods for systems of partial differential equations. Journal of Computational and Applied Mathematics, 299 (??):1, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005968.

Khademi:2019:GSK

[KKV19] Ali Khademi, Sergey Korotov, and Jon Eivind Vatne. On the generalization of the Synge– Krízek maximum angle condition for d-simplices. Journal of Computational and Applied Mathematics, 358(??):29–33, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301220.

Kim:2016:IAF

[KKW16] Jerim Kim, Bara Kim, and In-Suk Wee. Iterative algorithm for the first passage time distribution in a jump-diffusion model with regime-switching, and its applications. Journal of Computational and Applied Mathematics, 294 (??):177–195, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715004227.

Kang:2015:CIO

[KL15] Hongchao Kang and Chen Ling. Computation of integrals with oscillatory singular factors of algebraic and logarithmic type. Journal of Computational and Applied Mathematics, 285(??):72–85, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000679.

Khan:2017:BCB

[KL17] Khalid Khan and D. K. Lobiyal. Bèzier curves based on Lupaš (p, q)- analogue of Bernstein functions in CAGD. Journal of Computational and Applied Mathematics, 317 (??):458–477, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716306264.

Kim:2017:NAM

[KLC17] Sang Dong Kim, Eunjung Lee, and Wonjoon Choi. Newton’s algorithm for magnetohydrodynamic equations with the initial guess from Stokes-like problem. Journal of Computational and Applied Mathematics, 309(??):1–10, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302941.

Kokabifar:2016:NMP

[KLK16a] E. Kokabifar, G. B. Loghmani, and S. M. Karbassi. Nearest matrix with prescribed eigenvalues and its applications. Journal of Computational and Applied Mathematics, 298(??):53– 63, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715005853.

Kong:2016:MPH

[KLK16b] Jae Hoon Kong, Sunhong Lee, and Gwangil Kim. Minkowski Pythagoreanhodograph preserving mappings. Journal of Computational and Applied Mathematics, 308(??):166–176, December 15, 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716302667.

Kuo:2018:CMC

[KLL18] Yueh-Cheng Kuo, Wen-Wei Lin, and Ching-Sung Liu. Continuation methods for computing Z-/H-eigenpairs of nonnegative tensors. Journal of Computational and Applied Mathematics, 340(??):71–88, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301067.

Kanschat:2017:GMD

[KLM17] Guido Kanschat, Raytcho Lazarov, and Youli Mao. Geometric multigrid for Darcy and Brinkman models of flows in highly heterogeneous porous media: a numerical study. Journal of Computational and Applied Mathematics, 310(??):174–185, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302333.

Kelly:2018:ADB

[KLM18] James F. Kelly, Cheng-Gang Li, and Mark M. Meerschaert. Anomalous diffusion with ballistic scaling: a new fractional derivative. Journal of Computational and Applied Mathematics, 339 (??):161–178, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717305721.

Korman:2019:CAS

[KLS19] Philip Korman, Yi Li, and Dieter S. Schmidt. A computer assisted study of uniqueness of nodal ground state solutions. Journal of Computational and Applied Mathematics, 356(??):402–406, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042719300937.

Kapita:2018:PWD

[KM18] Shelvean Kapita and Peter Monk. A plane wave discontinuous Galerkin method with a Dirichletto-Neumann boundary condition for the scattering problem in acoustics. Journal of Computational and Applied Mathematics, 327(??):208–225, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303114.

Ketabchi:2015:SEE

[KMB15] R. Ketabchi, R. Mokhtari, and E. Babolian. Some error estimates for solving Volterra integral equations by using the reproducing kernel method. Journal of Computational and Applied Mathematics, 273(??):245–250, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002891.

Kumar:2018:FHO

[KMB^{+}18] Abhimanyu Kumar, P. Maroju, R. Behl, D. K. Gupta, and S. S. Motsa. A family of higher order iterations free from second derivative for nonlinear equations in R. Journal of Computational and Applied Mathematics, 330(??):676–694, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303461.

Khaledi:2016:SAP

[KMD^{+}16] K. Khaledi, E. Mahmoudi, M. Datcheva, D. König, and T. Schanz. Sensitivity analysis and parameter identification of a time dependent constitutive model for rock salt. Journal of Computational and Applied Mathematics, 293(??):128–138, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002022.

Krakowski:2017:MCA

[KMLB17] Krzysztof A. Krakowski, Luís Machado, Fátima Silva Leite, and Jorge Batista. A modified Casteljau algorithm to solve interpolation problems on Stiefel manifolds. Journal of Computational and Applied Mathematics, 311(??):84–99, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303375.

Kalogiratou:2016:NAC

[KMRS16] Z. Kalogiratou, Th. Monovasilis, Higinio Ramos, and T. E. Simos. A new approach on the construction of trigonometrically fitted two step hybrid methods. Journal of Computational and Applied Mathematics, 303(??):146–155, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300991.

Kleefeld:2016:SSM

[KMV16] B. Kleefeld and J. Martín-Vaquero. SERK2v3: Solving mildly stiff nonlinear partial differential equations. Journal of Computational and Applied Mathematics, 299(??):194– 206, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715006056.

Kuo:2017:IDI

[KNP^{+}17] F. Y. Kuo, D. Nuyens, L. Plaskota, I. H. Sloan, and G. W. Wasilkowski. Infinite-dimensional integration and the multivariate decomposition method. Journal of Computational and Applied Mathematics, 326 (??):217–234, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302765.

Kowalczyk:2017:RFA

[Kow17]  Piotr Kowalczyk.  On

root finding algorithms for complex functions with branch cuts. Journal of Computational and Applied Mathematics, 314 (??):1–9, April 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716305040.

Karaa:2015:OEE

[KP15] Samir Karaa and Amiya K.

Pani. Optimal error estimates of mixed FEMs for second order hyperbolic integro-differential equations with minimal smoothness on initial data. Journal of Computational and Applied Mathematics, 275(??):113–134, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003665.

Karciauskas:2019:FFF

[KP19] Kestutis Karciauskas and Jörg Peters. Fair free-form surfaces that are almost everywhere parametrically C^{2}. Journal of Computational and Applied Mathematics, 349(??):470–481, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304667.

Khan:2019:ZME

[KPP19] Muhammad Adil Khan, Dilda Pecarić, and Josip Pecarić. On Zipf–Mandelbrot entropy. Journal of Computational and Applied Mathematics, 346 (??):192–204, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304217.

Kouibia:2019:ASM

[KPRA19] A. Kouibia, M. Pasadas, L. Reyah, and R. Akhrif. Approximation of surfaces by modified Helmholtz splines. Journal of Computational and Applied Mathematics, 350(??):262–273, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306307.

Korotov:2016:LES

[KPS16] Sergey Korotov, Ángel Plaza, and José P. Suárez. Longest-edge n-section algorithms: properties and open problems. Journal of Computational and Applied Mathematics, 293(??):139–146, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001995.

Kumar:2017:CST

[KPS17] Kamlesh Kumar, Rajesh K. Pandey, and Shiva Sharma. Comparative study of three numerical schemes for fractional integro-differential equations. Journal of Computational and Applied Mathematics, 315(??):287–302, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716305465.

Kolk:2015:MSC

[KPT15] Marek Kolk, Arvet Pedas, and Enn Tamme. Modified spline collocation for linear fractional differential equations. Journal of Computational and Applied Mathematics, 283(??):28–40, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000333.

Krishnapriya:2017:MAI

[KPW17] P. Krishnapriya, M. Pitchaimani, and Tarynn M. Witten. Mathematical analysis of an influenza A epidemic model with discrete delay. Journal of Computational and Applied Mathematics, 324(??):155–172, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302017.

Khan:2016:DIE

[KR16] Subuhi Khan and Mumtaz Riyasat. Differential and integral equations for the 2-iterated Appell polynomials. Journal of Computational and Applied Mathematics, 306(??):116–132, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301820.

Karimi:2017:RCP

[KR17] Milad Karimi and Alireza Rezaee. Regularization of the Cauchy problem for the Helmholtz equation by using Meyer wavelet. Journal of Computational and Applied Mathematics, 320: 76–95, August 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Ku:2018:SES

[KR18] JaEun Ku and Lothar Reichel. Simple efficient solvers for certain ill-conditioned systems of linear equations, including H (div) problems. Journal of Computational and Applied Mathematics, 343(??):240–249, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302723.

Kreuzer:2019:UBS

[Kre19] Wolfgang Kreuzer. Using B-spline frames to represent solutions of acoustics scattering problems. Journal of Computational and Applied Mathematics, 351(??):331–343, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306976.

Kelly:2018:ATP

[KRR18] Cónall Kelly, Alexandra Rodkina, and Eeva Maria Rapoo. Adaptive timestepping for pathwise stability and positivity of strongly discretised nonlinear stochastic differential equations. Journal of Computational and Applied Mathematics, 334(??):39–57, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305873.

Kerber:2015:RRR

[KS15a] Michael Kerber and Michael Sagraloff. Root refinement for real polynomials using quadratic interval refinement. Journal of Computational and Applied Mathematics, 280(??):377–395, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400510X.

Kudryashov:2015:NMO

[KS15b] Nikolay A. Kudryashov and Kirill E. Shilnikov. Numerical modeling and optimization of the cryosurgery operations. Journal of Computational and Applied Mathematics, 290 (??):259–267, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500312X.

Kou:2016:USM

[KS16] Jisheng Kou and Shuyu Sun. Unconditionally stable methods for simulating multi-component twophase interface models with Peng–Robinson equation of state and various boundary conditions. Journal of Computational and Applied Mathematics, 291(??):158–182, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001077.

Kashkaria:2017:ECI

[KS17a] Bothayna S. H. Kashkaria and Muhammed I. Syam.

Evolutionary computational intelligence in solving a class of nonlinear Volterra–Fredholm integro-differential equations. Journal of Computational and Applied Mathematics, 311(??):314–323, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303466.

Kramer:2017:ETI

[KS17b] Patrick Krämer and Katharina Schratz. Efficient time integration of the Maxwell–Klein– Gordon equation in the non-relativistic limit regime. Journal of Computational and Applied Mathematics, 316(??):247–259, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303193.

Kumrod:2017:NCC

[KS17c] Pathaithep Kumrod and Wutiphol Sintunavarat. A new contractive condition approach to φ-fixed point results in metric spaces and its applications. Journal of Computational and Applied Mathematics, 311(??):194–204, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303351.

Kaiser:2018:AEA

[KS18] Klaus Kaiser and Jochen Schütz. Asymptotic error analysis of an IMEX Runge–Kutta method. Journal of Computational and Applied Mathematics, 343(??):139–154, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302395.

Kleinert:2017:NCD

[KSD17] Jan Kleinert, Bernd Simeon, and Klaus Dreßler. Nonsmooth contact dynamics for the large-scale simulation of granular material. Journal of Computational and Applied Mathematics, 316(??):345–357, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304575.

Khakimzyanov:2016:NRA

[KSDM16]  Gayaz  Khakimzyanov,

Nina Yu. Shokina, Denys Dutykh, and Dimitrios Mitsotakis. A new runup algorithm based on local high-order analytic expansions. Journal of Computational and Applied Mathematics, 298(??):82– 96, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005993.

Kokonendji:2015:CMN

[KSR15] Célestin C. Kokonendji, Cyrille C. Moypemna Sembona, and Joachim Sioké Rainaldy. A characterization of multivariate normal stable Tweedie models and their associated polynomials. Journal of Computational and Applied Mathematics, 288(??):159–168, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002356.

Kou:2019:SAP

[KSW19] Jisheng Kou, Shuyu Sun, and Yuanqing Wu. A semi-analytic porosity evolution scheme for simulating wormhole propagation with the Darcy– Brinkman–Forchheimer model. Journal of Computational and Applied Mathematics, 348(??):401–420, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042718305429.

Kravchenko:2015:AAT

[KT15] Vladislav V. Kravchenko and Sergii M. Torba. Analytic approximation of transmutation operators and applications to highly accurate solution of spectral problems. Journal of Computational and Applied Mathematics, 275(??):1–26, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003410.

Kuzmanovic:2016:DOA

[KTT16] Ivana Kuzmanović, Zoran Tomljanović, and Ninoslav Truhar. Damping optimization over the arbitrary time of the excited mechanical system. Journal of Computational and Applied Mathematics, 304(??):120–129, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301182.

Kumar:2018:ACF

[KtTBK18] N. Kumar, J. H. M. ten Thije Boonkkamp, and B. Koren. Approximation of the convective flux in the incompressible Navier– Stokes equations using local boundary-value problems. Journal of Computational and Applied Mathematics, 340(??):523–536, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300591.

Ku:2016:LEE

[Ku16] JaEun Ku. Local error estimates for leastsquares finite element methods for first-order system. Journal of Computational and Applied Mathematics, 299(??):92– 100, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005397.

Koleva:2015:OSK

[KV15] Miglena N. Koleva and Lubin G. Vulkov. Operator splitting kernel based numerical method for a generalized Leland’s model. Journal of Computational and Applied Mathematics, 275(??):294–303, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003380.

Koleva:2016:PFL

[KV16] Miglena N. Koleva and Lubin G. Vulkov. A positive flux limited difference scheme for the uncertain correlation 2D Black–Scholes problem. Journal of Computational and Applied Mathematics, 293(??):112–127, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001417.

Koleva:2017:NSO

[KV17] Miglena N. Koleva and Lubin G. Vulkov. A numerical study for optimal portfolio regime-switching model I. 2D Black– Scholes equation with an exponential non-linear term. Journal of Computational and Applied Mathematics, 318(??):538– 549, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716000236.

Kapl:2018:DBC

[KV18a] Mario Kapl and Vito Vitrih. Dimension and basis construction for C^{2}- smooth isogeometric spline spaces over bilinear-like G^{2} two-patch parameterizations. Journal of Computational and Applied Mathematics, 335(??):289– 311, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717306283.

Koleva:2018:FCA

[KV18b] Miglena N. Koleva and Lubin G. Vulkov. Fast computational approach to the delta Greek of nonlinear black–Scholes equations. Journal of Computational and Applied Mathematics, 340(??):508–522, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305629.

Kapl:2019:STE

[KV19a] Mario Kapl and Vito Vitrih. Solving the triharmonic equation over multipatch planar domains using isogeometric analysis.

Journal of Computational and Applied Mathematics, 358(??):385–404, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301451.

Khieu:2019:RHD

[KV19b] Tran Thi Khieu and Hoang-Hung Vo. Recovering the historical distribution for nonlinear spacefractional diffusion equation with temporally dependent thermal conductivity in higher dimensional space. Journal of Computational and Applied Mathematics, 345: 114–126, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Kostova-Vassilevska:2018:MRD

[KVO18] Tanya Kostova-Vassilevska and Geoffrey M. Oxberry. Model reduction of dynamical systems by proper orthogonal decomposition: Error bounds and comparison of methods using snapshots from the solution and the time derivatives. Journal of Computational and Applied Mathematics, 330(??):553–573, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304181.

Keller:2016:CCP

[KW16] Pawel Keller and Iwona Wróbel. Computing Cauchy principal value integrals using a standard adaptive quadrature. Journal of Computational and Applied Mathematics, 294(??):323–341, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004422.

Kulikov:2018:DQC

[KW18] G. Yu. Kulikov and R. Weiner. Doubly quasiconsistent fixed-stepsize numerical integration of stiff ordinary differential equations with implicit two-step peer methods. Journal of Computational and Applied Mathematics, 340(??):256–275, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830116X.

Kim:2016:SCN

[KYS16] Sihwan Kim, Jaeryun Yim, and Dongwoo Sheen.

Stable cheapest nonconforming finite elements for the Stokes equations. Journal of Computational and Applied Mathematics, 299(??):2– 14, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500343X.

Kato:2015:IPM

[KYY15] Atsushi Kato, Hiroshi Yabe, and Hiroshi Yamashita. An interior point method with a primal-dual quadratic barrier penalty function for nonlinear semidefinite programming. Journal of Computational and Applied Mathematics, 275(??):148–161, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003434.

Lee:2015:GSA

[LA15a] Ryeong Lee and Young Joon Ahn. Geometric shape analysis for convolution curve of two compatible quadratic Bézier curves. Journal of Computational and Applied Mathematics, 288(??):141–150, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002307.

Lee:2015:LCH

[LA15b] Ryeong Lee and Young Joon Ahn. Limit curve of H-Bézier curves and rational Bézier curves in standard form with the same weight. Journal of Computational and Applied Mathematics, 281(??):1– 9, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005494.

Li:2015:TLV

[LA15c] Yuan Li and Rong An. Two-level variational multiscale finite element methods for Navier–Stokes type variational inequality problem. Journal of Computational and Applied Mathematics, 290 (??):656–669, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003404.

Leela-apiradee:2019:NCT

[La19] Worrawate Leela-apiradee. New characterizations of tolerance-control and localized solutions to interval system of linear equations. Journal of Computational and Applied Mathematics, 355(??):11–22, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300123.

Ladics:2015:EAW

[Lad15] Tamás Ladics. Error analysis of waveform relaxation method for semi-linear partial differential equations. Journal of Computational and Applied Mathematics, 285(??):15–31, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500062X.

Li:2015:PHS

[LAHB15] Zhong Li, Rachid Ait-Haddou, and Luc Biard. Pythagorean hodograph spline spirals that match G^{3} Hermite data from circles. Journal of Computational and Applied Mathematics, 278(??):162–180, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042714004440.

Lan:2018:AES

[Lan18] Guangqiang Lan. Asymptotic exponential stability of modified truncated EM method for neutral stochastic differential delay equations. Journal of Computational and Applied Mathematics, 340(??):334–341, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301213.

Leela-apiradee:2015:LSM

[LaT15] Worrawate Leela-apiradee and Phantipa Thipwiwatpotjana. L- and Rlocalized solvabilities of max-separable interval linear equations and its applications. Journal of Computational and Applied Mathematics, 279(??):306– 317, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005032.

Leela-apiradee:2017:CFL

[LaTG17] Worrawate Leela-apiradee, Phantipa Thipwiwatpotjana, and Artur Gorka. Closed form of L-localized solution set of maxplus interval linear system and its application on optimization problem. Journal of Computational and Applied Mathematics, 317(??):113– 127, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305696.

Legaz-Aparicio:2018:AMF

[LAVMA18] Álvar-Ginés Legaz-Aparicio, Rafael Verdú-Monedero, and Jesús Angulo. Adaptive morphological filters based on a multiple orientation vector field dependent on image local features. Journal of Computational and Applied Mathematics, 330(??):965–981, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302303.

Lee:2015:FDM

[LBL15] T. E. Lee, M. J. Baines, and S. Langdon. A finite difference moving mesh method based on conservation for moving boundary problems. Journal of Computational and Applied Mathematics, 288(??):1–17, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001855.

Lee:2015:ALS

[LC15a] Hsueh-Chen Lee and Tsu-Fen Chen. Adaptive least-squares finite element approximations to Stokes equations. Journal of Computational and Applied Mathematics, 280(??):396–412, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005202.

Liu:2015:AMG

[LC15b] Li-Bin Liu and Yanping Chen. An adaptive moving grid method for a system of singularly perturbed initial value problems. Journal of Computational and Applied Mathematics, 274(??):11–22, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003033.

Lu:2016:DCP

[LC16] Kang-Ping Lu and Shao-Tung Chang. Detecting change-points for shifts in mean and variance using fuzzy classification maximum likelihood changepoint algorithms. Journal of Computational and Applied Mathematics, 308 (??):447–463, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630276X.

Li:2017:BGC

[LC17a] Weiwei Li and Wen Chen. Band gap calculations of photonic crystals by singular boundary method. Journal of Computational and Applied Mathematics, 315(??):273– 286, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305489.

Liu:2017:PEE

[LC17b] Li-Bin Liu and Yanping Chen. A-posteriori error estimation in maximum norm for a strongly coupled system of two singularly perturbed convection– diffusion problems. Journal of Computational and Applied Mathematics, 313(??):152–167, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303806.

Lin:2018:ATS

[LC18] Matthew M. Lin and Chun-Yueh Chiang. An accelerated technique for solving one type of discretetime algebraic Riccati equations. Journal of Computational and Applied Mathematics, 338 (??):91–110, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S037704271830075X.

Li:2019:SRU

[LC19] Peng Li and Wengu Chen. Signal recovery under cumulative coherence. Journal of Computational and Applied Mathematics, 346(??):399–417, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304394.

Lee:2017:MCM

[LCA17] Woojoo Lee, Ka Chun Cheung, and Jae Youn Ahn. Multivariate countermonotonicity and the minimal copulas. Journal of Computational and Applied Mathematics, 317 (??):589–602, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716306434.

Li:2015:SSU

[LCC^{+}15] Yusha Li, Wenyu Chen, Yiyu Cai, Ahmad Nasri, and Jianmin Zheng. Surface skinning using periodic T -spline in semi-NURBS form. Journal of Computational and Applied Mathematics, 273(??):116–131, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002751.

Christensen:2018:NMS

[lCCVV18] Max la Cour Christensen, Panayot S. Vassilevski, and Umberto Villa. Nonlinear multigrid solvers exploiting AMGe coarse spaces with approximation properties. Journal of Computational and Applied Mathematics, 340(??):691–708, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305307.

Li:2018:NWG

[LCH18] Guanrong Li, Yanping Chen, and Yunqing Huang. A new weak Galerkin finite element scheme for general second-order elliptic problems. Journal of Computational and Applied Mathematics, 344 (??):701–715, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302929.

Li:2016:HOA

[LCL16] Hefeng Li, Jianxiong Cao, and Changpin Li. High-order approximation to Caputo derivatives and Caputo-type advection–diffusion equations (III). Journal of Computational and Applied Mathematics, 299 (??):159–175, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715005919.

Liu:2018:BCF

[LCX18] Wei Liu, Jintao Cui, and Jie Xin. A block-centered finite difference method for an unsteady asymptotic coupled model in fractured media aquifer system. Journal of Computational and Applied Mathematics, 337(??):319–340, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300037.

Lin:2018:PLS

[LCZ18] Guangliang Lin, Xiaoliang Cheng, and Ye Zhang. A parametric level set based collage method for an inverse problem in elliptic partial differential equations. Journal of Computational and Applied Mathematics, 340(??):101–121, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300797.

Liu:2015:IDM

[LD15] Juan Liu and Lu Dai.

An iterative decomposition method for scatterer reconstruction in R^{3}. Journal of Computational and Applied Mathematics, 287(??):148–161, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001867.

Lakoba:2019:SANa

[LD19a] T. I. Lakoba and Z. Deng. Stability analysis of the numerical method of characteristics applied to a class of energy-preserving hyperbolic systems. Part I: Periodic boundary conditions. Journal of Computational and Applied Mathematics, 356(??):67–80, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300433.

Lakoba:2019:SANb

[LD19b] T. I. Lakoba and Z. Deng. Stability analysis of the numerical method of characteristics applied to a class of energy-preserving hyperbolic systems. Part II: Nonreflecting boundary conditions. Journal of Computational and Applied Mathematics, 356(??):267–292, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300585.

Lai:2018:BNS

[LDDF18] Yisheng Lai, Weiping Du, Dexin Duan, and Xiaoke Fang. Bounds on the number of solutions of polynomial systems and the Betti numbers of real piecewise algebraic hypersurfaces. Journal of Computational and Applied Mathematics, 329(??):147–163, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305568.

Li:2019:APM

[LDG19] Zhibao Li, Yu-Hong Dai, and Huan Gao. Alternating projection method for a class of tensor equations. Journal of Computational and Applied Mathematics, 346(??):490–504, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304333.

Le:2017:DAF

[LDK17] Nhat-Tan Le, Duy-Minh Dang, and Tran-Vu Khanh. A decomposition approach via Fourier sine transform for valuing American knock-out options with rebates. Journal of Computational and Applied Mathematics, 317(??):652– 671, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306409.

Li:2019:IAS

[LDL^{+}19] Chunmei Li, Xuefeng Duan, Linzhang Lu, Qingwen Wang, and Shuqian Shen. Iterative algorithm for solving a class of convex feasibility problem. Journal of Computational and Applied Mathematics, 352(??):352–367, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307192.

Liu:2017:ILC

[LDW17] Shengda Liu, Amar Debbouche, and JinRong Wang. On the iterative learning control for stochastic impulsive differential equations with randomly varying trial lengths. Journal of Computational and Applied Mathematics, 312(??):47– 57, March 1, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0377042715005385.

Liu:2018:IMS

[LDW18] Shengda Liu, Amar Debbouche, and JinRong Wang. ILC method for solving approximate controllability of fractional differential equations with noninstantaneous impulses. Journal of Computational and Applied Mathematics, 339 (??):343–355, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717303928.

Li:2017:CFR

[LDZ17] Xiezhang Li, James Diffenderfer, and Jiehua Zhu. Construction of a full rowrank matrix system for multiple scanning directions in discrete tomography. Journal of Computational and Applied Mathematics, 311(??):529–538, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304083.

Liu:2018:MSS

[LDZ18] Linna Liu, Feiqi Deng, and Quanxin Zhu. Mean square stability of two classes of theta methods for numerical computation and simulation of delayed stochastic Hopfield neural networks. Journal of Computational and Applied Mathematics, 343(??):428–447, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301869.

Leal-Enriquez:2018:MMI

[LE18] E. Leal-Enríquez. Mathematical modeling of intimate partner violence: Simulations of loss of control scenarios. Journal of Computational and Applied Mathematics, 330(??):1052–1062, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303436.

Lee:2018:SOO

[Lee18] Hyun Geun Lee. A secondorder operator splitting Fourier spectral method for fractional-in-space reactiondiffusion equations. Journal of Computational and Applied Mathematics, 333 (??):395–403, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717304247.

Levy:2018:MIG

[Lev18] Edmond Levy. On the moments of the integrated geometric Brownian motion. Journal of Computational and Applied Mathematics, 342(??):263–273, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301730.

Liu:2018:NDD

[LF18] J. K. Liu and Y. M. Feng.

A norm descent derivativefree algorithm for solving large-scale nonlinear symmetric equations. Journal of Computational and Applied Mathematics, 344(??):89–99, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302760.

Liu:2019:PRP

[LF19] Songshu Liu and Lixin Feng. A posteriori regularization parameter choice rule for a modified kernel method for a timefractional inverse diffusion problem. Journal of Computational and Applied Mathematics, 353(??):355– 366, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307696.

Li:2019:ITC

[LFCW19] Yujun Li, Lincong Fang, Juan Cao, and Guozhao Wang. Identification of two classes of planar septic Pythagorean hodograph curves. Journal of Computational and Applied Mathematics, 348(??):383–400, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305478.

Lozovskiy:2016:PBM

[LFKG16] Alexander Lozovskiy, Matthew Farthing, Chris Kees, and Eduardo Gildin. PODbased model reduction for stabilized finite element approximations of shallow water flows. Journal of Computational and Applied Mathematics, 302(??):50–70, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300085.

Liu:2019:PFQ

[LFWC19] Jun Liu, Hongfei Fu, Hong Wang, and Xiaochao Chai. A preconditioned fast quadratic spline collocation method for twosided space-fractional partial differential equations. Journal of Computational and Applied Mathematics, 360(??):138–156, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301797.

Orti:2018:CNB

[LGB18] José A. López Ortí, Vicente Agost Gómez, and Miguel Barreda Rochera. Corrigendum to “A new bi-parametric family of temporal transformations to improve the integration algorithms in the study of the orbital motion” [J. Comput. Appl. Math. 318 (2017) 479– 490]. Journal of Computational and Applied Mathematics, 338(??):324–325, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/

                                        S0377042717306167. See

[OGR17a].

Li:2018:WGF

[LGLC18] Rui Li, Yali Gao, Jian Li, and Zhangxin Chen. A weak Galerkin finite element method for a coupled Stokes–Darcy problem on general meshes. Journal of Computational and Applied Mathematics, 334(??):111–127, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730571X.

Li:2019:CND

[LGYC19] Rui Li, Yali Gao, Wenjing Yan, and Zhangxin Chen. A Crank–Nicolson discontinuous finite volume element method for a coupled non-stationary Stokes– Darcy problem. Journal of Computational and Applied Mathematics, 353 (??):86–112, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718307568.

Li:2016:NIT

[LH16] Li Li and Bo Han. A new iteratively total variational regularization for nonlinear inverse problems. Journal of Computational and Applied Mathematics, 298(??):40–52, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005877.

Li:2018:GFG

[LH18] Sheng-Kun Li and Ting-Zhu Huang. Global FOM and GMRES algorithms for a class of complex matrix equations. Journal of Computational and Applied Mathematics, 335(??):227– 241, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306027.

Li:2019:SCG

[LH19] Sheng-Kun Li and Ting-Zhu Huang. A shifted complex global Lanczos method and the quasiminimal residual variant for the Stein-conjugate matrix equation X + AX̄B = C. Journal of Computational and Applied Mathematics, 357(??):1–11, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300809.

Lesnic:2016:ISD

[LHJ16] D. Lesnic, S. O. Hussein, and B. T. Johansson. Inverse spacedependent force problems for the wave equation. Journal of Computational and Applied Mathematics, 306(??):10–39, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301613.

Li:2015:EPT

[LHJR15] Liang Li, Ting-Zhu Huang, Yan-Fei Jing, and Zhi-Gang Ren. Effective preconditioning through minimum degree ordering interleaved with incomplete factorization. Journal of Computational and Applied Mathematics, 279 (??):225–232, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042714004816.

Lee:2019:LSF

[LHK19] Eunjung Lee, Heonkyu Ha, and Sang Dong Kim. A least squares finite element method using Elsasser variables for magnetohydrodynamic equations. Journal of Computational and Applied Mathematics, 346(??):599–608, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830390X.

Li:2018:ROD

[LHLL18] Kun Li, Ting-Zhu Huang, Liang Li, and Stéphane Lanteri. A reduced-order DG formulation based on POD method for the timedomain Maxwell’s equations in dispersive media. Journal of Computational and Applied Mathematics, 336(??):249– 266, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300190.

Liang:2016:VTP

[LHM16] Hong Liang, Lisheng Hou, and Ju Ming. The velocity tracking problem for Wick-stochastic Navier–Stokes flows using Weiner chaos expansion. Journal of Computational and Applied Mathematics, 307(??):25–36, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302035.

Liu:2018:BIC

[LHPZ18] Hongyan Liu, Jin Huang, Yubin Pan, and Jipei Zhang. Barycentric interpolation collocation methods for solving linear and nonlinear highdimensional Fredholm integral equations. Journal of Computational and Applied Mathematics, 327(??):141–154, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303047.

Li:2016:MAF

[LHY16] Jichun Li, Yunqing Huang, and Wei Yang. Mathematical analysis and finite element simulation of a magnetized ferrite model. Journal of Computational and Applied Mathematics, 292(??):279–291, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003519.

Li:2019:UES

[LHY19] Congying Li, Yunqing Huang, and Nianyu Yi. An unconditionally energy stable second order finite element method for solving the Allen–Cahn equation. Journal of Computational and Applied Mathematics, 353(??):38– 48, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307556.

Liu:2015:CAT

[LHZ15] Qingbing Liu, Jian Huang, and Shouzhen Zeng. Convergence analysis of the two preconditioned iterative methods for M-matrix linear systems. Journal of Computational and Applied Mathematics, 281 (??):49–57, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714005135.

Llibre:2015:LCC

[LI15] Jaume Llibre and Jackson Itikawa. Limit cycles for continuous and discontinuous perturbations of uniform isochronous cubic centers. Journal of Computational and Applied Mathematics, 277(??):171–191, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004038.

Li:2018:SCN

[Li18]  Hanyu Li.  Structured

condition numbers for some matrix factorizations of structured matrices. Journal of Computational and Applied Mathematics, 336(??):219– 234, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306301.

Lin:2016:DGL

[Lin16] Runchang Lin. A discontinuous Galerkin leastsquares finite element method for solving coupled singularly perturbed reaction–diffusion equations. Journal of Computational and Applied Mathematics, 307(??):134–142, December 1, 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S037704271630108X.

Liu:2016:RFD

[Liu16] Wei Liu. Recursive filtering for discrete-time linear systems with fading measurement and timecorrelated channel noise. Journal of Computational and Applied Mathematics, 298(??):123–137, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005361.

Liu:2017:PTM

[Liu17] Chein-Shan Liu. The polynomial Trefftz method for solving backward and inverse source wave problems. Journal of Computational and Applied Mathematics, 321(??):211–225, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300572.

Liu:2018:WMM

[Liu18] Tao Liu. A wavelet multiscale method for the inverse problem of a nonlinear convection-diffusion equation. Journal of Computational and Applied Mathematics, 330(??):165– 176, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717304053.

Li:2015:ECL

[LJ15] Liangliang Li and Jigui Jian. Exponential convergence and Lagrange stability for impulsive Cohen– Grossberg neural networks with time-varying delays. Journal of Computational and Applied Mathematics, 277(??):23–35, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003860.

Lu:2016:IAM

[LJ16] Lizheng Lu and Chengkai Jiang. An iterative algorithm for G^{2} multiwise merging of Bézier curves. Journal of Computational and Applied Mathematics, 296(??):352–361, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005063.

Lu:2017:SRW

[LJX17] Lizheng Lu, Chengkai Jiang, and Xueyan Xiang. Some remarks on weighted Lupaš q-Bézier curves. Journal of Computational and Applied Mathematics, 313(??):393–396, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304733.

Lukassen:2018:OSC

[LK18] Axel Ariaan Lukassen and Martin Kiehl. Operator splitting for chemical reaction systems with fast chemistry. Journal of Computational and Applied Mathematics, 344 (??):495–511, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303492.

Lee:2017:OFE

[LKN17] Sang Deok Lee, Young Ik Kim, and Beny Neta. An optimal family of eighthorder simple-root finders with weight functions dependent on functionto-function ratios and their dynamics underlying extraneous fixed points. Journal of Computational and Applied Mathematics, 317(??):31– 54, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305787.

Lee:2017:IFE

[LKS17] Seungwoo Lee, Do Y. Kwak, and Imbo Sim. Immersed finite element method for eigenvalue problem. Journal of Computational and Applied Mathematics, 313(??):410– 426, March 15, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716304551.

Liu:2015:ASR

[LL15] X. Liu and M. Z. Liu. Asymptotic stability of Runge–Kutta methods for nonlinear differential equations with piecewise continuous arguments. Journal of Computational and Applied Mathematics, 280(??):265–274, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271400538X.

Li:2016:MFA

[LL16a] Shengtai Li and Hui Li. Modified FARGO algorithm and its combination with adaptive mesh refinement. Journal of Computational and Applied Mathematics, 307(??):170–182, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300772.

Li:2016:CSE

[LL16b] Yibiao Li and Xin Kai Li. The Chebyshev spectral element approximation with exact quadratures. Journal of Computational and Applied Mathematics, 296(??):320–333, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004732.

Long:2017:FTC

[LL17] Guangqing Long and Pailing Li. A fast trigonometric collocation method for some elliptic pseudodifferential equations. Journal of Computational and Applied Mathematics, 313(??):294–306, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716304460.

Liu:2018:SEG

[LL18] Zexian Liu and Hongwei Liu. Several efficient gradient methods with approximate optimal stepsizes for large scale unconstrained optimization. Journal of Computational and Applied Mathematics, 328(??):400–413, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303862.

Liu:2018:WGF

[LLC18] Xin Liu, Jian Li, and Zhangxin Chen. A weak Galerkin finite element method for the Navier– Stokes equations. Journal of Computational and Applied Mathematics, 333 (??):442–457, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717305708.

Li:2016:SFE

[LLCG16] Rui Li, Jian Li, Zhangxin Chen, and Yali Gao. A stabilized finite element method based on two local Gauss integrations for a coupled Stokes–Darcy problem. Journal of Computational and Applied Mathematics, 292(??):92–104, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003362.

Li:2018:RIW

[LLD18] Fang Li, Xiao-Guang Lv, and Ziwei Deng. Regularized iterative Weiner filter method for blind image deconvolution. Journal of Computational and Applied Mathematics, 336 (??):425–438, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717306544.

Liu:2018:PIA

[LLG^{+}18] Mingzeng Liu, Baojun Li, Qingjie Guo, Chungang Zhu, Ping Hu, and Yuanhai Shao. Progressive iterative approximation for regularized least square bivariate B-spline surface fitting. Journal of Computational and Applied Mathematics, 327(??):175–187, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042717303138.

Li:2019:EMF

[LLG19] Na Li, Ping Lin, and Fuzheng Gao. An expanded mixed finite element method for twodimensional Sobolev equations. Journal of Computational and Applied Mathematics, 348(??):342–355, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305284.

Liu:2018:RSC

[LLHO18] Li-Bin Liu, Guangqing Long, Zaitang Huang, and Aijia Ouyang. Rational spectral collocation and differential evolution algorithms for singularly perturbed problems with an interior layer. Journal of Computational and Applied Mathematics, 335(??):312– 322, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717306325.

Li:2019:CHO

[LLL19a] Keran Li, Wenyuan Liao, and Yaoting Lin. A compact high order Alternating Direction Implicit method for threedimensional acoustic wave equation with variable coefficient. Journal of Computational and Applied Mathematics, 361(??):113– 129, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0377042719301992.

Li:2019:EII

[LLL19b] Suhua Li, Chaoqian Li, and Yaotang Li. Meigenvalue inclusion intervals for a fourth-order partially symmetric tensor. Journal of Computational and Applied Mathematics, 356(??):391–401, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300226.

Lin:2015:CSW

[LLSM15] Guang Lin, Jiangguo Liu, and Farrah Sadre-Marandi. A comparative study on the weak Galerkin, discontinuous Galerkin, and mixed finite element methods. Journal of Computational and Applied Mathematics, 273(??):346–362, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003057.

Li:2017:NSM

[LLSZ17] Lin Li, Ping Lin, Xinhui Si, and Liancun Zheng. A numerical study for multiple solutions of a singular boundary value problem arising from laminar flow in a porous pipe with moving wall. Journal of Computational and Applied Mathematics, 313(??):536–549, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304812.

Layton:2016:RDI

[LLT16] William Layton, Yong Li, and Catalin Trenchea. Recent developments in IMEX methods with time filters for systems of evolution equations. Journal of Computational and Applied Mathematics, 299 (??):50–67, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715005105.

Li:2018:SBE

[LLV18a] Wen Li, Weihui Liu, and Seak-Weng Vong. Some bounds for H-eigenpairs and Z-eigenpairs of a tensor. Journal of Computational and Applied Mathematics, 342(??):37–57, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301717.

Liu:2018:TSA

[LLV18b] Dongdong Liu, Wen Li, and Seak-Weng Vong. The tensor splitting with application to solve multilinear systems. Journal of Computational and Applied Mathematics, 330 (??):75–94, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717303989.

Liu:2017:RCM

[LLW17] Xifu Liu, Wen Li, and Hongxing Wang. Rank constrained matrix best approximation problem with respect to (skew) Hermitian matrices. Journal of Computational and Applied Mathematics, 319(??):77–86, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306392.

Lian:2018:WCR

[LLW18] Ying-Ying Lian, Cui-Xia Li, and Shi-Liang Wu. Weaker convergent results of the generalized Newton method for the generalized absolute value equations. Journal of Computational and Applied Mathematics, 338(??):221–226, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717306556.

Li:2017:NMM

[LLX17] Buyang Li, Jun Liu, and Mingqing Xiao. A new multigrid method for unconstrained parabolic optimal control problems. Journal of Computational and Applied Mathematics, 326(??):358–373, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303084.

Liu:2019:OCR

[LLY19] Fang Liu, Zongqi Liang, and Yubin Yan. Optimal convergence rates for semidiscrete finite element approximations of linear space-fractional partial differential equations under minimal regularity assumptions. Journal of Computational and Applied Mathematics, 352(??):409–425, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307295.

Lee:2016:SSH

[LLZ16] Young-Ju Lee, Wei Leng, and Chen-Song Zhang. A stable and scalable hybrid solver for rate-type non-Newtonian fluid models. Journal of Computational and Applied Mathematics, 300(??):103– 118, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500638X.

Li:2018:MSI

[LLZ18] Peng Li, Keying Liu, and Weizhou Zhong. Marching schemes for inverse scattering problems in waveguides with curved boundaries. Journal of Computational and Applied Mathematics, 328(??):287–301, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303813.

Liu:2015:NSD

[LM15] Wei Liu and Xuerong Mao. Numerical stationary distribution and its convergence for nonlinear stochastic differential equations. Journal of Computational and Applied Mathematics, 276 (??):16–29, March 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042714003768.

Liu:2017:POS

[LM17a] Tianmiao Liu and Yoshifumi Muroi. Pricing of options in the singular perturbed stochastic volatility model. Journal of Computational and Applied Mathematics, 320: 138–144, August 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Lodhi:2017:QBS

[LM17b] Ram Kishun Lodhi and Hradyesh Kumar Mishra. Quintic B-spline method for solving second order linear and nonlinear singularly perturbed two-point boundary value problems. Journal of Computational and Applied Mathematics, 319(??):170–187, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300158.

Lv:2018:LMM

[LM18] Chang-Qing Lv and Chang-Feng Ma. A Levenberg– Marquardt method for solving semi-symmetric tensor equations. Journal of Computational and Applied Mathematics, 332 (??):13–25, April 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717304958.

Lizama:2018:WPS

[LMA18] Carlos Lizama and Marina Murillo-Arcila. Well posedness for semidiscrete fractional Cauchy problems with finite delay. Journal of Computational and Applied Mathematics, 339(??):356–366, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303783.

Lackner:2019:DEB

[LMM19] C. Lackner, S. Meng, and P. Monk. Determination of electromagnetic Bloch variety in a medium with frequencydependent coefficients. Journal of Computational and Applied Mathematics, 358(??):359–373, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301529.

Li:2017:NSS

[LMS17] Jichun Li, Eric A. Machorro, and Sidney Shields. Numerical study of signal propagation in corrugated coaxial cables. Journal of Computational and Applied Mathematics, 309(??):230–243, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303156.

Li:2018:TND

[LMUZ18] Jian Li, Mingzhi Mao, Frank Uhlig, and Yunong Zhang. Z-type neuraldynamics for time-varying nonlinear optimization under a linear equality constraint with robot application. Journal of Computational and Applied Mathematics, 327(??):155–166, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303175.

Li:2018:EPC

[LMW18] Haochen Li, Zhenguo Mu, and Yushun Wang. An energy-preserving Crank– Nicolson Galerkin spectral element method for the two dimensional nonlinear Schrödinger equation. Journal of Computational and Applied Mathematics, 344(??):245–258, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302966.

Leobacher:2016:UIP

[LN16] Gunther Leobacher and Philip Ngare. Utility indifference pricing of derivatives written on industrial loss indices. Journal of Computational and Applied Mathematics, 300(??):68– 82, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005828.

Loisel:2017:OSP

[LN17] Sébastien Loisel and Hieu Nguyen. An optimal Schwarz preconditioner for a class of parallel adaptive finite elements. Journal of Computational and Applied Mathematics, 321(??):90–107, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300833.

Loisel:2019:COA

[LN19] Sébastien Loisel and Hieu Nguyen. On the convergence of an optimal additive Schwarz method for parallel adaptive finite elements. Journal of Computational and Applied Mathematics, 355(??):193–200, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300378.

Losada:2017:ASC

[LNP17] J. Losada, J. J. Nieto, and E. Pourhadi. On the attractivity of solutions for a class of multi-term fractional functional differential equations. Journal of Computational and Applied Mathematics, 312(??):2–12, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003763.

Lemire:2019:XXX

[LO19] Daniel Lemire and Melissa E.

O’Neill. Xorshift1024*, xorshift1024+, xorshift128+ and xoroshiro128+ fail statistical tests for linearity. Journal of Computational and Applied Mathematics, 350(??):139–142, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306265.

Lopez-Orti:2015:IAD

[LOAGBR15] José A. López-Ortí, Vicente Agost-Gómez, and Miguel Barreda-Rochera. An improved algorithm to develop semi-analytical planetary theories using Sundman generalized variables. Journal of Computational and Applied Mathematics, 275(??):403–411, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002842.

Lagomasino:2015:CTH

[LP15a] G. López Lagomasino and S. Medina Peralta. On the convergence of type I Hermite–Padé approximants for rational perturbations of a Nikishin system. Journal of Computational and Applied Mathematics, 284(??):216–227, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000229.

Leonenko:2015:NAH

[LP15b] G. M. Leonenko and T. N. Phillips. Numerical approximation of high-dimensional Fokker– Planck equations with polynomial coefficients. Journal of Computational and Applied Mathematics, 273(??):296–312, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002696.

Liu:2018:PCK

[LP18] Gang Liu and Saminathan Ponnusamy. Prescribed cycles of König’s method for polynomials. Journal of Computational and Applied Mathematics, 336(??):468– 476, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730643X.

Lathouwers:2019:AMP

[LP19] Danny Lathouwers and Zoltán Perkó. An angular multigrid preconditioner for the radiation transport equation with Fokker–Planck scattering. Journal of Computational and Applied Mathematics, 350(??):165–177, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306174.

Lemos-Paiao:2017:EMC

[LPST17] Ana P. Lemos-Paião, Cristiana J. Silva, and Delfim F. M. Torres. An epidemic model for cholera with optimal control treatment. Journal of Computational and Applied Mathematics, 318(??):168– 180, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630526X.

Lizama:2017:ESS

[LPT17] Carlos Lizama, Jorge Pereira, and Eduard Toon. On the exponential stability of Samuelson model on some classes of times scales. Journal of Computational and Applied Mathematics, 325(??):1–17, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302054.

Liu:2019:EAM

[LPW^{+}19] Jiangguo Liu, Stephen Pankavich, Ying Wang, Anna Zemlyanova, and Xu Zhang. Editorial: the 3rd Annual Meeting of SIAM Central States Section. Journal of Computational and Applied Mathematics, 362(??):245, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719302961.

Liu:2019:AEM

[LQ19] Hailiang Liu and Linrui Qian. Alternating evolution methods for static Hamilton–Jacobi equations. Journal of Computational and Applied Mathematics, 351(??):270–289, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306204.

Li:2016:EAB

[LQZ16] Lingfei Li, Xianjun Qu, and Gongqiu Zhang. An efficient algorithm based on eigenfunction expansions for some optimal timing problems in finance. Journal of Computational and Applied Mathematics, 294(??):225–250, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004173.

Li:2016:SIA

[LRKV16] X. Li, J. K. Ryan, R. M.

Kirby, and C. Vuik. Smoothness-Increasing Accuracy-Conserving (SIAC) filters for derivative approximations of discontinuous Galerkin (DG) solutions over nonuniform meshes and near boundaries. Journal of Computational and Applied Mathematics, 294(??):275–296, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004185.

Li:2018:EHT

[LRL^{+}18] Shengguo Li, François-Henry Rouet, Jie Liu, Chun Huang, Xingyu Gao, and Xuebin Chi. An efficient hybrid tridiagonal divide-and-conquer algorithm on distributed memory architectures. Journal of Computational and Applied Mathematics, 344 (??):512–520, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303285.

Latha:2018:FOM

[LRRV18] V. Preethi Latha, Fathalla A. Rihan, R. Rakkiyappan, and G. Velmurugan. A fractional-order model for Ebola virus infection with delayed immune response on heterogeneous complex networks. Journal of Computational and Applied Mathematics, 339(??):134– 146, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0377042717305927.

Lebtahi:2017:ASI

[LRT17] Leila Lebtahi, Óscar Romero, and Néstor Thome. Algorithms for solving the inverse problem associated with KAK = A^{s+1}. Journal of Computational and Applied Mathematics, 309(??):333–341, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301145.

Li:2016:NMM

[LRX16] Xindong Li, Hongxing Rui, and Wenwen Xu. A new MCC–MFE method for compressible miscible displacement in porous media. Journal of Computational and Applied Mathematics, 302(??):139–156, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300322.

Li:2015:TCR

[LRZ15] Danping Li, Ximin Rong, and Hui Zhao. Timeconsistent reinsuranceinvestment strategy for an insurer and a reinsurer with mean-variance criterion under the CEV model. Journal of Computational and Applied Mathematics, 283(??):142–162, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000503.

Laadhari:2017:EFE

[LS17a] Aymen Laadhari and Gábor Székely. Eulerian finite element method for the numerical modeling of fluid dynamics of natural and pathological aortic valves. Journal of Computational and Applied Mathematics, 319(??):236–261, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305842.

Lee:2017:NGM

[LS17b] Chak Shing Lee and Dongwoo Sheen. Nonconforming Generalized Multiscale Finite Element Methods. Journal of Computational and Applied Mathematics, 311(??):215–229, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042716303478.

Lin:2018:PCG

[LS18a] Fu-Rong Lin and Yong-Jie Shi. Preconditioned conjugate gradient methods for the solution of Love’s integral equation with very small parameter. Journal of Computational and Applied Mathematics, 327(??):295–305, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303205.

Liu:2018:HOS

[LS18b] Kai Liu and Wei Shi. Highorder skew-symmetric differentiation matrix on symmetric grid. Journal of Computational and Applied Mathematics, 343(??):206–216, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302334.

Liu:2018:SSRa

[LSBL18a] Yiming Liu, Yimin Shi, Xuchao Bai, and Bin Liu. Stress-strength reliability analysis of multistate system based on generalized survival signature. Journal of Computational and Applied Mathematics, 342(??):274–291, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830222X.

Liu:2018:SSRb

[LSBL18b] Yiming Liu, Yimin Shi, Xuchao Bai, and Bin Liu. Stress-strength reliability analysis of system with multiple types of components using survival signature. Journal of Computational and Applied Mathematics, 342(??):375–398, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302103.

Li:2015:BPS

[LSD15] Minghao Li, Dongyang Shi, and Ying Dai. The Brezzi–Pitkäranta stabilization scheme for the elasticity problem. Journal of Computational and Applied Mathematics, 286(??):7–16, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042715000941.

Li:2019:NCG

[LSDY19] Xiangli Li, Juanjuan Shi, Xiaoliang Dong, and Jianglan Yu. A new conjugate gradient method based on quasi-Newton equation for unconstrained optimization. Journal of Computational and Applied Mathematics, 350(??):372–379, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306435.

Li:2018:HOW

[LSG18] Gang Li, Lina Song, and Jinmei Gao. High order well-balanced discontinuous Galerkin methods based on hydrostatic reconstruction for shallow water equations. Journal of Computational and Applied Mathematics, 340(??):546–560, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305289.

Li:2018:LTR

[LSH18a] Yu-Fan Li, Kun Shang, and Zheng-Hai Huang.

Low Tucker rank tensor recovery via ADMM based on exact and inexact iteratively reweighted algorithms. Journal of Computational and Applied Mathematics, 331(??):64–81, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304600.

Lin:2018:PFC

[LSH^{+}18b] P. T. Lin, J. N. Shadid, J. J. Hu, R. P. Pawlowski, and E. C. Cyr. Performance of fully-coupled algebraic multigrid preconditioners for largescale VMS resistive MHD. Journal of Computational and Applied Mathematics, 344(??):782–793, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304594.

Li:2019:MVR

[LSH^{+}19] Zibo Li, Guangmin Sun, Cunfu He, Xiucheng Liu, Ruihuan Zhang, Yu Li, Dequn Zhao, Hao Liu, and Fan Zhang. Multivariable regression methods using modified Chebyshev polynomials of class 2. Journal of Computational and Applied Mathematics, 346(??):609–619, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302024.

Lv:2015:ENR

[LSL15] Xiao-Guang Lv, Yong-Zhong Song, and Fang Li. An efficient nonconvex regularization for wavelet frame and total variation based image restoration. Journal of Computational and Applied Mathematics, 290(??):553–566, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003283.

Lu:2017:CSS

[LSL17] Y. L. Lu, M. H. Song, and M. Z. Liu. Convergence and stability of the split-step theta method for stochastic differential equations with piecewise continuous arguments. Journal of Computational and Applied Mathematics, 317(??):55– 71, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716305751.

Lu:2018:CSC

[LSL18] Yulan Lu, Minghui Song, and Mingzhu Liu. Convergence and stability of the compensated split-step theta method for stochastic differential equations with piecewise continuous arguments driven by Poisson random measure. Journal of Computational and Applied Mathematics, 340(??):296–317, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301183.

Li:2017:LBL

[LSLZ17] Chunguang Li, Cong Sun, Cuihua Li, and Hong Zheng. Lower bound limit analysis by quadrilateral elements. Journal of Computational and Applied Mathematics, 315(??):319– 326, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630557X.

Los:2018:PST

[LSP18] M. Lo’s, R. Schaefer, and M. Paszyński. Parallel space-time hp adaptive discretization scheme for parabolic problems. Journal of Computational and Applied Mathematics, 344(??):819–835, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306192.

Li:2019:FRL

[LSS19] Zijia Li, Daniel F. Scharler, and Hans-Peter Schröcker. Factorization results for left polynomials in some associative real algebras: State of the art, applications, and open questions. Journal of Computational and Applied Mathematics, 349(??):508–522, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305971.

Li:2019:EES

[LSW19] Jiyong Li, Wei Shi, and Xinyuan Wu. The existence of explicit symplectic ARKN methods with several stages and algebraic order greater than two. Journal of Computational and Applied Mathematics, 353(??):204– 209, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830757X.

Liu:2019:CTS

[LSXL19] Hanze Liu, Bo Sang, Xiangpeng Xin, and Xiqiang Liu. CK transformations, symmetries, exact solutions and conservation laws of the generalized variable-coefficient KdV types of equations. Journal of Computational and Applied Mathematics, 345: 127–134, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Li:2019:FEM

[LSYY19] Ruo Li, Zhiyuan Sun, Fanyi Yang, and Zhijian Yang. A finite element method by patch reconstruction for the Stokes problem using mixed formulations. Journal of Computational and Applied Mathematics, 353 (??):1–20, June 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718307489.

Li:2016:SAI

[LSZ16] Tingting Li, Chi-Wang Shu, and Mengping Zhang.

Stability analysis of the inverse Lax–Wendroff boundary treatment for high order upwind-biased finite difference schemes. Journal of Computational and Applied Mathematics, 299 (??):140–158, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715005920.

Liu:2017:RNB

[LSZB17] Bin Liu, Yimin Shi, Fode Zhang, and Xuchao Bai. Reliability nonparametric Bayesian estimation for the masked data of parallel systems in step-stress accelerated life tests. Journal of Computational and Applied Mathematics, 311(??):375–386, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630348X.

Lazarov:2018:NUI

[LT18] Raytcho Lazarov and Aziz Takhirov. A note on the uniform inf-sup condition for the Brinkman problem in highly heterogeneous media. Journal of Computational and Applied Mathematics, 340(??):537–545, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305976.

Liu:2017:AFT

[LTY17] Zhongli Liu, Hongjiong Tian, and Xiong You. Adiabatic Filon-type methods for highly oscillatory second-order ordinary differential equations. Journal of Computational and Applied Mathematics, 320: 1–14, August 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Lu:2015:EAM

[Lu15a]  Lizheng Lu.  Explicit

algorithms for multiwise merging of Bézier curves. Journal of Computational and Applied Mathematics, 278(??):138–148, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004397.

Lu:2015:GMB

[Lu15b] Lizheng Lu. Gram matrix of Bernstein basis: Properties and applications. Journal of Computational and Applied Mathematics, 280(??):37– 41, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005160.

Lu:2015:PQH

[Lu15c] Lizheng Lu. Planar quintic G^{2} Hermite interpolation with minimum strain energy. Journal of Computational and Applied Mathematics, 274(??):109–117, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003343.

Lu:2016:HOM

[Lu16]  De-Yong Lu.  A hy-

brid optimization method for multiplicative noise and blur removal. Journal of Computational and Applied Mathematics, 302(??):224–233, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300607.

Lu:2018:SEP

[Lu18] Peng Lu. Structural effects of participation propensity in online collective actions: Based on big data and Delphi methods.

Journal of Computational and Applied Mathematics, 344(??):288–300, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302437.

Luo:2018:EAW

[Luo18] Xue Luo. Error analysis of the Wiener–Askey polynomial chaos with hyperbolic cross approximation and its application to differential equations with random input. Journal of Computational and Applied Mathematics, 335(??):242– 269, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306271.

Le:2016:COM

[LV16a] Thanh Hieu Le and Marc Van Barel. A convex optimization model for finding non-negative polynomials. Journal of Computational and Applied Mathematics, 301(??):121–134, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000297.

Lopez:2016:SPC

[LV16b] Luciano Lopez and Giuseppe Vacca. Spectral properties and conservation laws in mimetic finite difference methods for PDEs. Journal of Computational and Applied Mathematics, 292(??):760–784, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000369.

Li:2015:AAA

[LW15] Zhenquan Li and Robert Wood. Accuracy analysis of an adaptive mesh refinement method using benchmarks of 2-D steady incompressible liddriven cavity flows and coarser meshes. Journal of Computational and Applied Mathematics, 275(??):262–271, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003458.

Li:2017:NRK

[LW17a] Xiuying Li and Boying Wu. A new reproducing kernel method for variable order fractional boundary value problems for functional differential equations. Journal of Computational and Applied Mathematics, 311(??):387–393, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303703.

Li:2017:AVA

[LW17b] Zhenquan Li and Robert Wood. Accuracy verification of a 2D adaptive mesh refinement method for incompressible or steady flow. Journal of Computational and Applied Mathematics, 318(??):259– 265, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304423.

Lan:2019:SCR

[LW19a] Guangqiang Lan and Qiushi Wang. Strong convergence rates of modified truncated EM methods for neutral stochastic differential delay equations. Journal of Computational and Applied Mathematics, 362(??):83–98, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/

science/article/pii/
S0377042719302560.

Liu:2019:NCD

[LW19b] Jun Liu and Zhu Wang.

Non-commutative discretizethen-optimize algorithms for elliptic PDE-constrained optimal control problems. Journal of Computational and Applied Mathematics, 362(??):596–613, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718304485.

Liu:2019:SWG

[LW19c] Yujie Liu and Junping Wang. Simplified weak Galerkin and new finite difference schemes for the Stokes equation. Journal of Computational and Applied Mathematics, 361(??):176–206, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302195.

Liu:2017:IMB

[LWC17] Chein-Shan Liu, Peifang Wang, and Jiang-Ren Chang. An iterative method based on coupled closed-form coefficients expansions for recovering the pollutant source and initial pollution profile. Journal of Computational and Applied Mathematics, 326(??):87–98, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302728.

Li:2018:STF

[LWDW18] Jiyong Li, Xianfen Wang, Shuo Deng, and Bin Wang. Symmetric trigonometricallyfitted two-step hybrid methods for oscillatory problems. Journal of Computational and Applied Mathematics, 344 (??):115–131, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303157.

Ling:2015:PAM

[LWJ15] Sitao Ling, Musheng Wei, and Zhigang Jia. Perturbation analysis for the matrix least squares problem AXB = C. Journal of Computational and Applied Mathematics, 273(??):150–159, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002805.

Li:2017:SVI

[LWL17] Yiqun Li, Boying Wu, and Melvin Leok. Spectral variational integrators for semi-discrete Hamiltonian wave equations. Journal of Computational and Applied Mathematics, 325(??):56–73, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302145.

Li:2018:CDC

[LWLL18] Hanhui Li, Hefeng Wu, Shujin Lin, and Xiaonan Luo. Coupling deep correlation filter and online discriminative learning for visual object tracking. Journal of Computational and Applied Mathematics, 329(??):191–201, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302443.

Liu:2016:NSI

[LWS16] Xin-Ge Liu, Feng-Xian Wang, and Yan-Jun Shu. A novel summation inequality for stability analysis of discrete-time neural networks. Journal of Computational and Applied Mathematics, 304(??):160–171, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301194.

Liu:2017:SSA

[LWTQ17] Xin-Ge Liu, Feng-Xian Wang, Mei-Lan Tang, and Sai-Bing Qiu. Stability and synchronization analysis of neural networks via Halanay-type inequality. Journal of Computational and Applied Mathematics, 319(??):14–23, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630646X.

Li:2018:MMR

[LWV18] Qiuqi Li, Yuhe Wang, and Maria Vasilyeva. Multiscale model reduction for fluid infiltration simulation through dualcontinuum porous media with localized uncertainties. Journal of Computational and Applied Mathematics, 336(??):127– 146, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300086.

Lee:2017:ICF

[LWW17] Sanghyun Lee, Mary F.

Wheeler, and Thomas Wick. Iterative coupling of flow, geomechanics and adaptive phasefield fracture including level-set crack width approaches. Journal of Computational and Applied Mathematics, 314(??):40– 60, April 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305118.

Li:2018:PLA

[LWXF18] Junbin Li, Renhong Wang, Min Xu, and Qin Fang. Piecewise linear approximation methods with stochastic sampling sites. Journal of Computational and Applied Mathematics, 329(??):173–178, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301061.

Li:2015:VAH

[LWZ^{+}15] Wen-Ping Li, Zheng-Ming Wang, Tao Zhang, Bao-Cun Bai, Yong-He Ding, and Ya Deng. Vectorial additive halfquadratic minimization for isotropic regularization. Journal of Computational and Applied Mathematics, 281(??):152– 168, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005676.

Liu:2017:ACT

[LWZ17] Dongjie Liu, Jiming Wu, and Xiaoping Zhang. The adaptive composite trapezoidal rule for Hadamard finite-part integrals on an interval. Journal of Computational and Applied Mathematics, 325(??):165–174, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302121.

Li:2019:CBO

[LWZ19a] Anshui Li, Yuanyuan Wang, and Minzhi Zhao. On the convergence of bivariate order statistics: Almost sure convergence and convergence rate. Journal of Computational and Applied Mathematics, 348(??):445–452, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305508.

Luo:2019:MRG

[LWZ19b] Shangzhen Luo, Mingming Wang, and Wei Zhu. Maximizing a robust goalreaching probability with penalization on ambiguity. Journal of Computational and Applied Mathematics, 348(??):261–281, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305363.

Li:2016:RSP

[LWZZ16] Ying Li, Musheng Wei, Fengxia Zhang, and Jianli Zhao. Real structurepreserving algorithms of Householder based transformations for quaternion matrices. Journal of Computational and Applied Mathematics, 305(??):82–91, October 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301583.

Li:2019:PMQ

[LWZZ19] Ying Li, Musheng Wei, Fengxia Zhang, and Jianli Zhao. On the power method for quaternion right eigenvalue problem. Journal of Computational and Applied Mathematics, 345:59–69, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Li:2015:TLA

[LX15a] Binjie Li and Xiaoping Xie. A two-level algorithm for the weak Galerkin discretization of diffusion problems. Journal of Computational and Applied Mathematics, 287(??):179–195, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271500196X.

Lu:2015:REP

[LX15b] Xin Lu and Shufang Xu. Rounding errors of partial derivatives of simple eigenvalues of the quadratic eigenvalue problem. Journal of Computational and Applied Mathematics, 282(??):268– 277, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000151.

Li:2016:AFH

[LX16a] Binjie Li and Xiaoping Xie. Analysis of a family of HDG methods for second order elliptic problems. Journal of Computational and Applied Mathematics, 307(??):37–51, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630200X.

Liu:2016:LMA

[LX16b] Jun Liu and Mingqing Xiao. A leapfrog multigrid algorithm for the optimal control of parabolic PDEs with Robin boundary conditions. Journal of Computational and Applied Mathematics, 307(??):216–234, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300577.

Lu:2016:QPA

[LX16c] Lizheng Lu and Xueyan Xiang. Quintic polynomial approximation of logaesthetic curves by curvature deviation. Journal of Computational and Applied Mathematics, 296 (??):389–396, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715005002.

Lu:2017:NMD

[LX17] Lizheng Lu and Xueyan Xiang. Note on multidegree reduction of Bézier curves via modified Jacobi– Bernstein basis transformation. Journal of Computational and Applied Mathematics, 315(??):65– 69, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305246.

Lan:2018:SCR

[LX18] Guangqiang Lan and Fang Xia. Strong convergence rates of modified truncated EM method for stochastic differential equations. Journal of Computational and Applied Mathematics, 334(??):1–17, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305848.

Li:2018:CSF

[LXD^{+}18] Chong-Jun Li, Lin-Lin Xie, Wen-Bin Du, Hai-Dong Li, and Huan Bao. Curve and surface fitting models based on the diagonalizable differential systems. Journal of Computational and Applied Mathematics, 329(??):179–190, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730208X.

Li:2015:NMC

[LXL15] Wei Li, Mengxue Xia, and Haohao Li. New method for computing the upper bound of optimal value in interval quadratic program. Journal of Computational and Applied Mathematics, 288(??):70–80, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002083.

Li:2016:SRU

[LXL16] Wei Li, Mengxue Xia, and Haohao Li. Some results on the upper bound of optimal values in interval convex quadratic programming. Journal of Computational and Applied Mathematics, 302(??):38–49, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300243.

Lan:2019:PSE

[LXW19] Guangqiang Lan, Fang Xia, and Qiushi Wang. Polynomial stability of exact solution and a numerical method for stochastic differential equations with time-dependent delay. Journal of Computational and Applied Mathematics, 346(??):340–356, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304448.

Liu:2017:MSM

[LXZ17] Keji Liu, Yongzhi Xu, and Jun Zou. A multilevel sampling method for detecting sources in a stratified ocean waveguide. Journal of Computational and Applied Mathematics, 309(??):95–110, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303119.

Liu:2018:DRW

[LXZ18] Keji Liu, Yongzhi Xu, and Jun Zou. Direct recovery of wavepenetrable scatterers in a stratified ocean waveguide. Journal of Computational and Applied Mathematics, 338(??):239–257, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300682.

Lan:2015:ESE

[LY15a] Guangqiang Lan and Chenggui Yuan. Exponential stability of the exact solutions and θ-EM approximations to neutral SDDEs with Markov switching. Journal of Computational and Applied Mathematics, 285 (??):230–242, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000989.

Langer:2015:PSA

[LY15b] Ulrich Langer and Huidong Yang. Partitioned solution algorithms for fluidstructure interaction problems with hyperelastic models. Journal of Computational and Applied Mathematics, 276(??):47– 61, March 1, 2015. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714003641.

Li:2018:CMB

[LY18] Rui Li and Jun-Feng Yin. On the convergence of modulus-based matrix splitting iteration methods for a class of nonlinear complementarity problems with H_{+}-matrices. Journal of Computational and Applied Mathematics, 342(??):202–209, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730657X.

Lv:2015:RST

[LYG15] Jing Lv, Hu Yang, and Chaohui Guo. Robust smooth-threshold estimating equations for generalized varying-coefficient partially linear models based on exponential score function. Journal of Computational and Applied Mathematics, 280(??):125–140, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004749.

Liang:2017:SSE

[LYG17] Hui Liang, Zhanwen Yang, and Jianfang Gao. Strong superconvergence of the Euler–Maruyama method for linear stochastic Volterra integral equations. Journal of Computational and Applied Mathematics, 317 (??):447–457, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716305295.

Li:2019:MVE

[LYS19] Changfeng Li, Yirang Yuan, and Huailing Song. A mixed volume element with upwind multistep mixed volume element and convergence analysis for numerical simulation of nuclear waste contaminant disposal. Journal of Computational and Applied Mathematics, 356(??):164–181, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300512.

Liu:2015:LPA

[LZ15] Chunxiao Liu and Shengfeng Zhu. Laguerre pseudospectral approximation to the Thomas–Fermi equation. Journal of Computational and Applied Mathematics, 282(??):251– 261, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000138.

Liang:2016:SLP

[LZ16a] Zhao-Zheng Liang and Guo-Feng Zhang. SIMPLElike preconditioners for saddle point problems from the steady Navier– Stokes equations. Journal of Computational and Applied Mathematics, 302(??):211–223, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300590.

Liu:2016:MSS

[LZ16b] Linna Liu and Quanxin Zhu. Mean square stability of two classes of theta method for neutral stochastic differential delay equations. Journal of Computational and Applied Mathematics, 305(??):55–67, October 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301418.

Liu:2016:EAG

[LZ16c] Sanyang Liu and Yingfeng Zhao. An efficient algorithm for globally solving generalized linear multiplicative programming. Journal of Computational and Applied Mathematics, 296(??):840–847, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005543.

Leduc:2017:CRR

[LZ17a] Guillaume Leduc and Xiangchen Zeng. Convergence rate of regimeswitching trees. Journal of Computational and Applied Mathematics, 319(??):56–76, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306446.

Liang:2017:CBG

[LZ17b] Zhao-Zheng Liang and Guo-Feng Zhang. Convergence behavior of generalized parameterized Uzawa method for singular saddle-point problems. Journal of Computational and Applied Mathematics, 311(??):293–305, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303417.

Li:2018:CFP

[LZ18a] Cai-Yun Li and Chun-Gang Zhu. G^{1} continuity of four pieces of developable surfaces with Bézier boundaries. Journal of Computational and Applied Mathematics, 329(??):164–172, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301097.

Li:2018:PHP

[LZ18b] Chong-Jun Li and Yi-Jun Zhong. A pseudo-heuristic parameter selection rule for ℓ^{1}-regularized minimization problems. Journal of Computational and Applied Mathematics, 333 (??):1–19, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S037704271730496X.

Liang:2018:PAB

[LZ18c] Zhao-Zheng Liang and Guo-Feng Zhang. Parameterized approximate block LU preconditioners for generalized saddle point problems. Journal of Computational and Applied Mathematics, 336(??):281– 296, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306593.

Lu:2019:HQP

[LZ19] Lizheng Lu and Shiqing Zhao. High-quality point sampling for B-spline fitting of parametric curves with feature recognition. Journal of Computational and Applied Mathematics, 345:286–294, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Li:2019:AFT

[LZA19] Haitao Li, Yating Zheng, and Fuad E. Alsaadi. Algebraic formulation and topological structure of Boolean networks with state-dependent delay. Journal of Computational and Applied Mathematics, 350 (??):87–97, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718306046.

Li:2015:NAF

[LZC15] Jian Li, Xin Zhao, and Zhangxin Chen. A novel L^{∞} analysis for finite volume approximations of the Stokes problem. Journal of Computational and Applied Mathematics, 279(??):97–105, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004634.

Li:2016:MGT

[LZL16] Chaoqian Li, Chengyi Zhang, and Yaotang Li. Minimal Gersgorin tensor eigenvalue inclusion set and its approximation. Journal of Computational and Applied Mathematics, 302(??):200–210, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300553.

Lin:2018:ROF

[LZL^{+}18a] Xiaoran Lin, Shangbo Zhou, Hua Li, Hongbin Tang, and Ying Qi. Rhythm oscillation in fractional-order relaxation oscillator and its application in image enhancement. Journal of Computational and Applied Mathematics, 339(??):69– 84, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300566.

Liu:2018:CDO

[LZL^{+}18b] Yansheng Liu, Yating Zheng, Haitao Li, Fuad E. Alsaadi, and Bashir Ahmad. Control design for output tracking of delayed Boolean control networks. Journal of Computational and Applied Mathematics, 327(??):188–195, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303163.

Lu:2019:NSP

[LZRH19] Jing Lu, Botao Zhou, Md. Motiur Rahman, and Xiaoming He. New solution to the pressure transient equation in a two-layer reservoir with crossflow. Journal of Computational and Applied Mathematics, 362 (??):680–693, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718303480.

Liu:2015:IIR

[LZS15] Xiaomin Liu, Bin Zhang, and Jinju Sun. An improved implicit reinitialization method for the level set function applied to shape and topology optimization of fluid. Journal of Computational and Applied Mathematics, 281(??):207– 229, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005731.

Li:2018:GCG

[LZW18] Ji Li, Tie Zhou, and Chao Wang. On global convergence of gradient descent algorithms for generalized phase retrieval problem. Journal of Computational and Applied Mathematics, 329(??):202–222, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303497.

Liu:2019:DSH

[LZX19] Shitao Liu, Liang Zhang, and Yifan Xing. Dynamics of a stochastic heroin epidemic model. Journal of Computational and Applied Mathematics, 351 (??):260–269, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718306708.

Liu:2015:EOS

[LZY15] Youjun Liu, Jianwen Zhang, and Jurang Yan. Existence of oscillatory solutions of second order delay differential equations. Journal of Computational and Applied Mathematics, 277(??):17–22, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003823.

Liu:2017:NDT

[LZY17] Chaolin Liu, Zhimin Zhang, and Hu Yang. A note on a discrete time MAP risk model. Journal of Computational and Applied Mathematics, 309(??):111–121, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303065.

Liao:2019:BPP

[LZZ19] Li-Dan Liao, Guo-Feng Zhang, and Mu-Zheng Zhu. A block product preconditioner for saddle point problems. Journal of Computational and Applied Mathematics, 352(??):426–436, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307167.

Liu:2017:FBM

[LZZL17] Minghui Liu, Deyue Zhang, Xu Zhou, and Feng Liu. The Fourier– Bessel method for solving the Cauchy problem connected with the Helmholtz equation. Journal of Computational and Applied Mathematics, 311(??):183–193, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303429.

Magrenan:2016:LCD

[MA16] Á. Alberto Magreñán and Ioannis K. Argyros. On the local convergence and the dynamics of Chebyshev–Halley methods with six and eight order of convergence. Journal of Computational and Applied Mathematics, 298(??):236–251, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005907.

Mirzaee:2018:ASN

[MA18] Farshid Mirzaee and Sahar Alipour. Approximate solution of nonlinear quadratic integral equations of fractional order via piecewise linear functions. Journal of Computational and Applied Mathematics, 331(??):217–227, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304697.

Marquez:2019:CAO

[MA19a] Miguel Marquez and Henry Arguello. Coded aperture optimization for single pixel compressive computed tomography. Journal of Computational and Applied Mathematics, 348 (??):58–69, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718305211.

Mehrez:2019:NHH

[MA19b] Khaled Mehrez and Praveen Agarwal. New Hermite– Hadamard type integral inequalities for convex functions and their applications. Journal of Computational and Applied Mathematics, 350(??):274– 285, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306290.

Mohammadi:2017:RKR

[MAF17]  Fatemeh  Mohammadi,

Carmen Arévalo, and Claus Führer. Runge– Kutta restarters for multistep methods in presence of frequent discontinuities. Journal of Computational and Applied Mathematics, 316(??):287–297, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304691.

Ma:2019:EAM

[MAK19] Xiaofeng Ma, Manuchehr Aminian, and Michael Kirby. Error-adaptive modeling of streaming time-series data using radial basis functions. Journal of Computational and Applied Mathematics, 362 (??):295–308, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718306642.

Mantica:2015:OPE

[Man15]  Giorgio Mantica.  Or-

thogonal polynomials of equilibrium measures supported on Cantor sets. Journal of Computational and Applied Mathematics, 290(??):239–258, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003088.

Mao:2015:TEM

[Mao15] Xuerong Mao. The truncated Euler–Maruyama method for stochastic differential equations. Journal of Computational and Applied Mathematics, 290 (??):370–384, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003210.

Mao:2016:CRT

[Mao16] Xuerong Mao. Convergence rates of the truncated Euler–Maruyama method for stochastic differential equations. Journal of Computational and Applied Mathematics, 296 (??):362–375, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715004884.

Marz:2017:NAO

[Mär17]  Roswitha März.  New

answers to an old question in the theory of differential-algebraic equations: Essential underlying ODE versus inherent ODE. Journal of Computational and Applied Mathematics, 316(??):271–286, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304599.

Marcuzzi:2018:LEP

[Mar18]  Fabio Marcuzzi.  Lin-

ear estimation of physical parameters with subsampled and delayed data. Journal of Computational and Applied Mathematics, 331(??):11–22, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304673.

Marques:2019:PRG

[Mar19a] Filipe J. Marques. Products of ratios of gamma functions — an application to the distribution of the test statistic for testing the equality of covariance matrices. Journal of Computational and Applied Mathematics, 354(??):86– 95, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300499.

Martin:2019:SAF

[Mar19b]  Olga Martin.  Stabil-

ity approach to the fractional variational iteration method used for the dynamic analysis of viscoelastic beams. Journal of Computational and Applied Mathematics, 346(??):261–276, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303765.

Mirzaee:2019:NSB

[MAS19] Farshid Mirzaee, Sahar Alipour, and Nasrin Samadyar. Numerical solution based on hybrid of block-pulse and parabolic functions for solving a system of nonlinear stochastic Itô–Volterra integral equations of fractional order. Journal of Computational and Applied Mathematics, 349(??):157–171, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305910.

Matculevich:2018:FAE

[Mat18] S. Matculevich. Functional approach to the error control in adaptive IgA schemes for elliptic boundary value problems. Journal of Computational and Applied Mathematics, 344(??):394–423, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303005.

Mazure:2018:CTP

[Maz18] Marie-Laurence Mazure.

Constructing totally positive piecewise Chebyshevian B-spline bases. Journal of Computational and Applied Mathematics, 342(??):550–586, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830195X.

Michalkova:2015:TBS

[MB15a] Kristýna Michálková and Bohumír Bastl. A tetrahedron-based subdivision scheme for spatial G^{1} curves. Journal of Computational and Applied Mathematics, 281 (??):196–206, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714005809.

Moreno-Balcazar:2015:MSO

[MB15b] Juan J. Moreno-Balcázar.

Δ-Meixner–Sobolev orthogonal polynomials: Mehler– Heine type formula and zeros. Journal of Computational and Applied Mathematics, 284(??):228–234, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271400497X.

Menniche:2017:LBA

[MB17] Linda Menniche and Djamel Benterki. A logarithmic barrier approach for linear programming. Journal of Computational and Applied Mathematics, 312 (??):267–275, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S037704271630259X.

Maroju:2017:SNO

[MBM17] P. Maroju, Ramandeep Behl, and S. S. Motsa. Some novel and optimal families of King’s method with eighth and sixteenth-order of convergence. Journal of Computational and Applied Mathematics, 318(??):136– 148, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305519.

Mohammadi:2018:GFO

[MC18] Fakhrodin Mohammadi and Carlo Cattani. A generalized fractional-order Legendre wavelet Tau method for solving fractional differential equations. Journal of Computational and Applied Mathematics, 339(??):306–316, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304624.

March:2019:FVS

[MC19] Nathan G. March and Elliot J. Carr. Finite volume schemes for multilayer diffusion. Journal of Computational and Applied Mathematics, 345: 206–223, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Martinez-Casas:2017:IRW

[MCGNBD17] José Martínez-Casas, Juan Giner-Navarro, L. Baeza, and F. D. Denia. Improved railway wheelsettrack interaction model in the high-frequency domain. Journal of Computational and Applied Mathematics, 309(??):642–653, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630214X.

Ma:2016:SOA

[MCLW16] Qiang Ma, Junzhi Cui, Zhihui Li, and Ziqiang Wang. Second-order asymptotic algorithm for heat conduction problems of periodic composite materials in curvilinear coordinates. Journal of Computational and Applied Mathematics, 306(??):87–115, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301728.

Maleknejad:2018:ANA

[MD18] K. Maleknejad and R. Dehbozorgi. Adaptive numerical approach based upon Chebyshev operational vector for nonlinear Volterra integral equations and its convergence analysis. Journal of Computational and Applied Mathematics, 344 (??):356–366, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303170.

Makarov:2019:EAS

[MD19] Volodymyr L. Makarov and Denys V. Dragunov.

An efficient approach for solving stiff nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 345:452–470, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Macias-Diaz:2019:SPB

[MDAP19] J. E. Macías-Díaz and J. Alba-Pérez. A structurepreserving Bhattacharya method for nonlinear parabolic equations with fractional diffusion and advection. Journal of Computational and Applied Mathematics, 354(??):623– 640, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042719300093.

Macias-Diaz:2017:CDC

[MDG17a] J. E. Macías-Díaz and A. E. González. A convergent and dynamically consistent finite-difference method to approximate the positive and bounded solutions of the classical Burgers–Fisher equation. Journal of Computational and Applied Mathematics, 318(??):604– 615, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005634.

Macias-Diaz:2017:BLI

[MDG17b] J. E. Macías-Díaz and J. A. Guerrero. A bounded linear integrator for some diffusive nonlinear time-dependent partial differential equations. Journal of Computational and Applied Mathematics, 318(??):515– 528, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300115.

Macias-Diaz:2019:PPN

[MDG19] J. E. Macías-Díaz and Armando Gallegos. On a positivity-preserving numerical model for a linearized hyperbolic Fisher– Kolmogorov–Petrovski–Piscounov equation. Journal of Computational and Applied Mathematics, 354(??):603– 611, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304795.

Macias-Diaz:2017:MBE

[MDGVR17] J. E. Macías-Díaz, Armando Gallegos, and H. Vargas-Rodríguez. A modified Bhattacharya exponential method to approximate positive and bounded solutions of the Burgers–Fisher equation. Journal of Computational and Applied Mathematics, 318(??):366–377, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042716303454.

Mei:2019:SIF

[MDH19] Jin-Jin Mei, Yiqiu Dong, and Ting-Zhu Huang. Simultaneous image fusion and denoising by using fractional-order gradient information. Journal of Computational and Applied Mathematics, 351 (??):212–227, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718306770.

Macias-Diaz:2016:DQB

[MDT16] J. E. Macías-Díaz and Stefania Tomasiello. A differential quadraturebased approach à la Picard for systems of partial differential equations associated with fuzzy differential equations. Journal of Computational and Applied Mathematics, 299(??):15– 23, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004161.

Macias-Diaz:2017:DMD

[MDVM17] J. E. Macías-Díaz and J. Villa-Morales. A deterministic model for the distribution of the stopping time in a stochastic equation and its numerical solution. Journal of Computational and Applied Mathematics, 318(??):93– 106, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305581.

Macias-Diaz:2019:SES

[MDVR19] J. E. Macías-Díaz and Héctor Vargas-Rodríguez. Some exact solutions of a hyperbolic model of energy transmission in non-homogeneous media. Journal of Computational and Applied Mathematics, 354(??):597– 602, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304254.

Mansoori:2018:DMS

[ME18] Amin Mansoori and Majid Erfanian. A dynamic model to solve the absolute value equations. Journal of Computational and Applied Mathematics, 333(??):28– 35, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304636.

Mehrez:2019:NPS

[Meh19] Khaled Mehrez. New properties for several classes of functions related to the Fox–Wright functions. Journal of Computational and Applied Mathematics, 362(??):161–171, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302602.

Mejjaoli:2017:HAA

[Mej17] Hatem Mejjaoli. Harmonic analysis associated with the modified Cherednik type operator and quantitative uncertainty principles for its Hartley transform. Journal of Computational and Applied Mathematics, 321(??):508–531, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301000.

Melas:2019:CEF

[Mel19] Evangelos Melas. Classes of elementary function solutions to the CEV model I. Journal of Computational and Applied Mathematics, 360(??):62–77, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301645.

Mennouni:2017:PCG

[Men17] Abdelaziz Mennouni. Piecewise constant Galerkin method for a class of Cauchy singular integral equations of the second kind in L^{2}. Journal of Computational and Applied Mathematics, 326 (??):268–272, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730273X.

Meral:2019:DFS

[Mer19] Gülnihal Meral. DRBEM– FDM solution of a chemotaxishaptotaxis model for cancer invasion. Journal of Computational and Applied Mathematics, 354 (??):299–309, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718302425.

Myerscough:2016:EPP

[MF16] Keith W. Myerscough and Jason Frank. Explicit, parallel Poisson integration of point vortices on the sphere. Journal of Computational and Applied Mathematics, 304(??):100–119, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301108.

Mangold:2015:NMR

[MFK^{+}15] Michael Mangold, Lihong Feng, Dmytro Khlopov, Stefan Palis, Peter Benner, Daniel Binev, and Andreas Seidel-Morgenstern. Nonlinear model reduction of a continuous fluidized bed crystallizer. Journal of Computational and Applied Mathematics, 289(??):253–266, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715000400.

Mollapourasl:2018:RPM

[MFLL18] Reza Mollapourasl, Ali Fereshtian, Hengguang Li, and Xun Lu. RBF-PU method for pricing options under the jump-diffusion model with local volatility. Journal of Computational and Applied Mathematics, 337(??):98–118, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300232.

Mouloud:2018:TDS

[MFWK18] A. Mouloud, H. Fellouah, B. A. Wade, and M. Kessal. Time discretization and stability regions for dissipativedispersive Kuramoto–Sivashinsky equation arising in turbulent gas flow over laminar liquid. Journal of Computational and Applied Mathematics, 330(??):605–617, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304314.

Ma:2015:MAO

[MFX15] Junjie Ma, Chunhua Fang, and Shuhuang Xiang.

Modified asymptotic orders of the direct Filon method for a class of Volterra integral equations. Journal of Computational and Applied Mathematics, 281(??):120– 125, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005664.

Maclean:2015:CHO

[MG15a] John Maclean and Georg A.

Gottwald. On convergence of higher order schemes for the projective integration method for stiff ordinary differential equations. Journal of Computational and Applied Mathematics, 288(??):44–69, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002149.

Magrenan:2015:RDD

[MG15b] Ángel Alberto Magreñán and José Manuel Gutiérrez. Real dynamics for damped Newton’s method applied to cubic polynomials. Journal of Computational and Applied Mathematics, 275(??):527–538, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042713006511.

Martin:2015:CAC

[MG15c] Raúl Martín Martín and Irene García Camacha Gutiérrez. Combined algorithm to compute Doptimal designs. Journal of Computational and Applied Mathematics, 278(??):248–257, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004567.

Mnatsakanov:2017:MRA

[MG17] Robert M. Mnatsakanov and Broti Garai. On the moment-recovered approximations of regression and derivative functions with applications. Journal of Computational and Applied Mathematics, 315 (??):17–31, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S037704271630512X.

Medina-Guevara:2017:ACO

[MGMDGVR17] M. G. Medina-Guevara, J. E. Macías-Díaz, A. Gallegos, and H. Vargas-Rodríguez. On S^{1} as an alternative continuous opinion space in a three-party regime. Journal of Computational and Applied Mathematics, 318(??):230– 241, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304782.

Moreno-Garcia:2017:TSR

[MGMGJLRB17] Antonio Moreno-Garcia, Juan Moreno-Garcia, Luis Jimenez-Linares, and Luis Rodriguez-Benitez. Time series represented by means of fuzzy piecewise lineal segments. Journal of Computational and Applied Mathematics, 318(??):156– 167, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305271.

Mirzaee:2016:ASN

[MH16a] Farshid Mirzaee and Elham Hadadiyan. Approximation solution of nonlinear Stratonovich Volterra integral equations by applying modification of hat functions. Journal of Computational and Applied Mathematics, 302(??):272–284, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300620.

Mirzaee:2016:CMS

[MH16b] Farshid Mirzaee and Afsun Hamzeh. A computational method for solving nonlinear stochastic Volterra integral equations. Journal of Computational and Applied Mathematics, 306(??):166–178, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301856.

Maleki:2015:DST

[MHAA15] Mohammad Maleki, Ishak Hashim, Saeid Abbasbandy, and A. Alsaedi. Direct solution of a type of constrained fractional variational problems via an adaptive pseudospectral method. Journal of Computational and Applied Mathematics, 283(??):41–57, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271500031X.

Mostafaee:2016:ILP

[MHC16] Amin Mostafaee, Milan Hladík, and Michal Cerný. Inverse linear programming with interval coefficients. Journal of Computational and Applied Mathematics, 292(??):591–608, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004069.

Ming:2016:SCM

[MHL16] Wanyuan Ming, Chengming Huang, and Meng Li. Superconvergence in collocation methods for Volterra integral equations with vanishing delays. Journal of Computational and Applied Mathematics, 308(??):361–378, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302801.

Muoi:2016:DGM

[MHMP16] Pham Quy Muoi, Dinh Nho Hào, Peter Maass, and Michael Pidcock. Descent gradient methods for nonsmooth minimization problems in ill-posed problems. Journal of Computational and Applied Mathematics, 298(??):105–122, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005932.

Matus:2017:ASO

[MHV17] Piotr Matus, Le Minh Hieu, and Lubin G. Vulkov. Analysis of second order difference schemes on non-uniform grids for quasilinear parabolic equations. Journal of Computational and Applied Mathematics, 310(??):186–199, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301716.

Matsue:2017:CLF

[MHY17] Kaname Matsue, Tomohiro Hiwaki, and Nobito Yamamoto. On the construction of Lyapunov functions with computer assistance. Journal of Computational and Applied Mathematics, 319(??):385–412, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/

S0377042717300067. See

errata [MHY21].

Matsue:2021:ECL

[MHY21] Kaname Matsue, Tomohiro Hiwaki, and Nobito Yamamoto. Errata to “On the construction of Lyapunov functions with computer assistance” [j. comp. appl. math. 319 (2017) 385–412]. Journal of Computational and Applied Mathematics, 384(??):Article 113175, March 1, 2021. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042720304660. See

[MHY17].

Miao:2018:CEA

[Mia18] Cun-Qiang Miao. Computing eigenpairs in augmented Krylov subspace produced by Jacobi–Davidson correction equation. Journal of Computational and Applied Mathematics, 343(??):363–372, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830270X.

Micula:2018:INM

[Mic18] Sanda Micula. An iterative numerical method for fractional integral equations of the second kind. Journal of Computational and Applied Mathematics, 339(??):124–133, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306209.

Mehdiyeva:2016:SRN

[MII16] G. Yu. Mehdiyeva, V. R. Ibrahimov, and M. N. Imanova. Some refinement of the notion of symmetry for the Volterra integral equations and the construction of symmetrical methods to solve them. Journal of Computational and Applied Mathematics, 306(??):1–9, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301534.

Milosevic:2015:CAS

[Mil15] Marija Milosević. Convergence and almost sure exponential stability of implicit numerical methods for a class of highly nonlinear neutral stochastic differential equations with constant delay. Journal of Computational and Applied Mathematics, 280(??):248–264, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005421.

Milosevic:2016:EMA

[Mil16] Marija Milosević. The Euler–Maruyama approximation of solutions to stochastic differential equations with piecewise constant arguments. Journal of Computational and Applied Mathematics, 298(??):1–12, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005646.

Mirzaei:2015:AML

[Mir15] Davoud Mirzaei. Analysis of moving least squares approximation revisited. Journal of Computational and Applied Mathematics, 282(??):237– 250, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000199.

Mirzaei:2016:EBG

[Mir16]  Davoud Mirzaei.  Er-

ror bounds for GMLS derivatives approximations of Sobolev functions. Journal of Computational and Applied Mathematics, 294(??):93–101, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004100.

Milovanovic:2015:GQR

[MIT15a] Gradimir V. Milovanović, Tomislav S. Igić, and Dragana Turnić. Generalized quadrature rules of Gaussian type for numerical evaluation of singular integrals. Journal of Computational and Applied Mathematics, 278(??):306–325, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004555.

Mitchell:2015:ACI

[Mit15b] S. L. Mitchell. Applying the combined integral method to two-phase Stefan problems with delayed onset of phase change. Journal of Computational and Applied Mathematics, 281(??):58– 73, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005305.

Miyajima:2015:SEB

[Miy15] Shinya Miyajima. A sharp error bound of the approximate solutions for saddle point linear systems. Journal of Computational and Applied Mathematics, 277(??):36–46, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003926.

Miyajima:2017:FVC

[Miy17] Shinya Miyajima. Fast verified computation for stabilizing solutions of discrete-time algebraic Riccati equations. Journal of Computational and Applied Mathematics, 319(??):352–364, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300468.

Miyajima:2018:FVC

[Miy18] Shinya Miyajima. Fast verified computation for the matrix principal p-th root. Journal of Computational and Applied Mathematics, 330(??):276–288, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304077.

Miyajima:2019:VCH

[Miy19] Shinya Miyajima. Verified computation for the Hermitian positive definite solution of the conjugate discrete-time algebraic Riccati equation. Journal of Computational and Applied Mathematics, 350(??):80– 86, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306150.

Ma:2018:CAG

[MJ18a] Lingling Ma and Lijian Jiang. Convergence analysis for GMsFEM approximation of elliptic eigenvalue problems. Journal of Computational and Applied Mathematics, 327(??):109–126, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303059.

Majumdar:2018:GAT

[MJ18b] Kaushik Majumdar and Srinath Jayachandran. A geometric analysis of time series leading to information encoding and a new entropy measure. Journal of Computational and Applied Mathematics, 328(??):469–484, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303473.

Mortari:2019:HAL

[MJS19] Daniele Mortari, Hunter Johnston, and Lidia Smith. High accuracy least-squares solutions of nonlinear differential equations. Journal of Computational and Applied Mathematics, 352(??):293–307, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307325.

Meng:2017:PDS

[MK17] Jie Meng and Hyun-Min Kim. The positive definite solution of the nonlinear matrix equation X^{p} = A + M(B + X^{−1})^{−1}M^{∗}. Journal of Computational and Applied Mathematics, 322(??):139–147, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301371.

Minh:2019:SFB

[MKKV19] Triet Le Minh, Tran Thi Khieu, Tra Quoc Khanh, and Hoang-Hung Vo. On a space fractional backward diffusion problem and its approximation of local solution. Journal of Computational and Applied Mathematics, 346(??):440–455, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304369.

Meinecke:2016:SDP

[ML16] Lina Meinecke and Per Lötstedt. Stochastic diffusion processes on Cartesian meshes. Journal of Computational and Applied Mathematics, 294 (??):1–11, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715004070.

Ma:2017:GJE

[ML17a] Jie Ma and Yongshu Li. Gauss–Jordan elimination method for computing all types of generalized inverses related to the 1-inverse. Journal of Computational and Applied Mathematics, 321(??):26–43, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300687.

Ma:2017:ANT

[ML17b] Xin Ma and Zhibin Liu.

Application of a novel time-delayed polynomial grey model to predict the natural gas consumption in China. Journal of Computational and Applied Mathematics, 324(??):17–24, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301905.

Ma:2018:WCL

[ML18] Junjie Ma and Huilan Liu. A well-conditioned Levin method for calculation of highly oscillatory integrals and its application. Journal of Computational and Applied Mathematics, 342(??):451–462, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302255.

Ma:2019:LFF

[ML19a] Weijun Ma and Zhiping Li. A locking-free FEM for cavitation computation in nearly incompressible nonlinear elasticity. Journal of Computational and Applied Mathematics, 353(??):210– 218, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307507.

Mei:2019:SRK

[ML19b] Liangcai Mei and Yingzhen Lin. Simplified reproducing kernel method and convergence order for linear Volterra integral equations with variable coefficients. Journal of Computational and Applied Mathematics, 346(??):390–398, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304473.

Meng:2016:NDB

[MLJ16] Jing Meng, Hou-Biao Li, and Yan-Fei Jing. A new deflated block GCROT(m, k) method for the solution of linear systems with multiple righthand sides. Journal of Computational and Applied Mathematics, 300 (??):155–171, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S037704271500641X.

Mishra:2015:ODM

[MLS15] S. K. Mishra, K. K. Lai, and Vinay Singh. Optimality and duality for minimax fractional programming with support function under (C, α, ρ, d)- convexity. Journal of Computational and Applied Mathematics, 274 (??):1–10, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042714003069.

Meng:2016:CAF

[MLSK16] Zhaoliang Meng, Zhongxuan Luo, Dongwoo Sheen, and Sihwan Kim. Convergence analysis of a family of 14-node brick elements. Journal of Computational and Applied Mathematics, 301(??):53–63, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300024.

Mallela:2016:HTC

[MLV16] Abhishek Mallela, Suzanne Lenhart, and Naveen K. Vaidya. HIV–TB coinfection treatment: Modeling and optimal control theory perspectives. Journal of Computational and Applied Mathematics, 307(??):143–161, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301078.

Ma:2019:ANN

[MLW19] Xin Ma, Zhibin Liu, and Yong Wang. Application of a novel nonlinear multivariate grey Bernoulli model to predict the tourist income of China. Journal of Computational and Applied Mathematics, 347(??):84–94, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304709.

Maruster:2015:EES

[MM15a] L. Maruster and St.

Maruster. On the error estimation and T - stability of the Mann iteration. Journal of Computational and Applied Mathematics, 276(??):110–116, March 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003884.

Millar:2015:FEM

[MM15b] Felipe Millar and David Mora. A finite element method for the buckling problem of simply supported Kirchhoff plates. Journal of Computational and Applied Mathematics, 286(??):68–78, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000886.

Miyatake:2015:NAC

[MM15c] Yuto Miyatake and Takayasu Matsuo. A note on the adaptive conservative/dissipative discretization for evolutionary partial differential equations. Journal of Computational and Applied Mathematics, 274(??):79–87, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003148.

Meshkat:2017:JRW

[MM17] R. S. Meshkat and E. Mahmoudi. Joint reliability and weighted importance measures of a k-out-of-n system with random weights for components. Journal of Computational and Applied Mathematics, 326(??):273–283, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302959.

Marinov:2018:SWS

[MM18] Tchavdar T. Marinov and Rossitza S. Marinova. Solitary wave solutions with non-monotone shapes for the modified Kawahara equation. Journal of Computational and Applied Mathematics, 340(??):561–570, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304168.

Marco:2019:ACM

[MM19] Ana Marco and José-Javier Martínez. Accurate computation of the Moore–Penrose inverse of strictly totally positive matrices. Journal of Computational and Applied Mathematics, 350(??):299– 308, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306162.

Milosevic:2015:EIM

[MMD15] Dusan M. Milosević, Mimica R. Milosević, and Jovana Dzunić. On an efficient inclusion method for finding polynomial zeros. Journal of Computational and Applied Mathematics, 290(??):298–309, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500309X.

Manas-Manas:2016:ABV

[MMMMB16] Juan F. Mañas-Mañas, Francisco Marcellán, and Juan J. Moreno-Balcázar. Asymptotic behavior of varying discrete Jacobi– Sobolev orthogonal polynomials. Journal of Computational and Applied Mathematics, 300(??):341– 353, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000212.

Maleknejad:2015:CAS

[MMO15] K. Maleknejad, R. Mollapourasl, and A. Ostadi. Convergence analysis of sinc-collocation methods for nonlinear Fredholm integral equations with a weakly singular kernel. Journal of Computational and Applied Mathematics, 278(??):1–11, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004166.

Mora:2017:MMI

[MMPGCSP17] H. Mora, J. Mora-Pascual, J. M. García-Chamizo, and M. T. Signes-Pont. Mathematical model and implementation of rational processing. Journal of Computational and Applied Mathematics, 309(??):575–586, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042716302187.

Moerkotte:2015:POQ

[MMRS15] Guido Moerkotte, Martin Montag, Audrey Repetti, and Gabriele Steidl. Proximal operator of quotient functions with application to a feasibility problem in query optimization. Journal of Computational and Applied Mathematics, 285(??):243–255, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001004.

Malesza:2019:ASF

[MMS19] W. Malesza, M. Macias, and D. Sierociuk. Analytical solution of fractional variable order differential equations. Journal of Computational and Applied Mathematics, 348 (??):214–236, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042718305223.

Migda:2019:APS

[MMZ19] Janusz Migda, Małgorzata Migda, and Małgorzata Zdanowicz. Asymptotic properties of solutions to fourth order difference equations. Journal of Computational and Applied Mathematics, 362(??):68–82, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302535.

Mandal:2017:SLS

[MN17] Moumita Mandal and Gnaneshwar Nelakanti. Superconvergence of Legendre spectral projection methods for Fredholm– Hammerstein integral equations. Journal of Computational and Applied Mathematics, 319(??):423–439, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300481.

Mandal:2019:SRL

[MN19a] Moumita Mandal and Gnaneshwar Nelakanti. Superconvergence results of Legendre spectral projection methods for weakly singular Fredholm–Hammerstein integral equations. Journal of Computational and Applied Mathematics, 349(??):114–131, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305831.

McGregor:2019:APG

[MN19b] Geoffrey McGregor and Jean-Christophe Nave. Area-preserving geometric Hermite interpolation. Journal of Computational and Applied Mathematics, 361(??):236–248, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301244.

Mehrdoust:2018:MFH

[MNFS18] F. Mehrdoust, A. R.

Najafi, S. Fallah, and O. Samimi. Mixed fractional Heston model and the pricing of American options. Journal of Computational and Applied Mathematics, 330(??):141– 154, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717303904.

Malinowska:2018:OCD

[MO18a] Agnieszka B. Malinowska and Tatiana Odzijewicz. Optimal control of discrete-time fractional multi-agent systems. Journal of Computational and Applied Mathematics, 339 (??):258–274, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717305058.

Matychyn:2018:TOC

[MO18b] Ivan Matychyn and Viktoriia Onyshchenko. On time-optimal control of fractional-order systems. Journal of Computational and Applied Mathematics, 339(??):245–257, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717305174.

Mohsen:2017:AFS

[Moh17] Adel A. K. Mohsen. Accurate function Sinc interpolation and derivative estimations over finite intervals. Journal of Computational and Applied Mathematics, 324(??):216–224, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301425.

Mohebbi:2019:CNL

[Moh19] Akbar Mohebbi. Crank– Nicolson and Legendre spectral collocation methods for a partial integrodifferential equation with a singular kernel. Journal of Computational and Applied Mathematics, 349(??):197–206, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305855.

Mokhtary:2015:RER

[Mok15] P. Mokhtary. Reconstruction of exponentially rate of convergence to Legendre collocation solution of a class of fractional integro-differential equations. Journal of Computational and Applied Mathematics, 279(??):145–158, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004713.

Molabahrami:2015:DCM

[Mol15] Ahmad Molabahrami. Direct computation method for solving a general nonlinear Fredholm integrodifferential equation under the mixed conditions: Degenerate and non-degenerate kernels. Journal of Computational and Applied Mathematics, 282(??):34–43, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042714005810.

Molabahrami:2017:GLM

[Mol17] Ahmad Molabahrami. Galerkin– Levin method for highly oscillatory integrals. Journal of Computational and Applied Mathematics, 321 (??):499–507, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301140.

Moore:2017:MCS

[Moo17] Brian E. Moore. Multiconformal-symplectic PDEs and discretizations. Journal of Computational and Applied Mathematics, 323(??):1–15, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301632.

Mena:2018:NLR

[MOPP18] Hermann Mena, Alexander Ostermann, Lena-Maria Pfurtscheller, and Chiara Piazzola. Numerical low-rank approximation of matrix differential equations. Journal of Computational and Applied Mathematics, 340(??):602–614, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301201.

Matsumoto:2018:AQM

[MOY18] Makoto Matsumoto, Ryuichi Ohori, and Takehito Yoshiki. Approximation of quasi-Monte Carlo worst case error in weighted spaces of infinitely times smooth functions. Journal of Computational and Applied Mathematics, 330(??):155– 164, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717303990.

Matonoha:2015:CET

[MP15] C. Matonoha and S. Papácek.

On the connection and equivalence of two methods for solving an illposed inverse problem based on FRAP data. Journal of Computational and Applied Mathematics, 290(??):598–608, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003271.

Megan:2017:ESN

[MP17a] Mihail Megan and Ioan-Lucian Popa. Exponential splitting for nonautonomous linear discretetime systems in Banach spaces. Journal of Computational and Applied Mathematics, 312(??):181–191, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301637.

Morkisz:2017:OPA

[MP17b] Paweł M. Morkisz and Paweł Przybylowicz. Optimal pointwise approximation of SDE’s from inexact information. Journal of Computational and Applied Mathematics, 324(??):85–100, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301930.

Medina:2018:SIE

[MP18] F. Patricia Medina and Małgorzata Peszynska.

Stability for implicit– explicit schemes for nonequilibrium kinetic systems in weighted spaces with symmetrization. Journal of Computational and Applied Mathematics, 328(??):216–231, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303618.

Matus:2018:CTS

[MPH18] Piotr Matus, Dmitriy Poliakov, and Le Minh Hieu. On the consistent twoside estimates for the solutions of quasilinear convection-diffusion equations and their approximations on non-uniform grids. Journal of Computational and Applied Mathematics, 340(??):571–581, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730451X.

Martinez:2017:IMI

[MPM^{+}17] J. Martínez, F. Piscaglia, A. Montorfano, A. Onorati, and S. M. Aithal. Influence of momentum interpolation methods on the accuracy and convergence of pressure-velocity coupling algorithms in OpenFOAM ^{(}R). Journal of Computational and Applied Mathematics, 309(??):654–673, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301649.

Marquez:2019:HTR

[MPS19] Alberto Márquez, Ángel Plaza, and José P. Suárez. Hamiltonian triangular refinements and space-filling curves. Journal of Computational and Applied Mathematics, 346(??):18–25, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303819.

Morgado:2015:NAD

[MR15] M. L. Morgado and M. Rebelo. Numerical approximation of distributed order reaction– diffusion equations. Journal of Computational and Applied Mathematics, 275(??):216–227, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0377042714003495.

Makarov:2019:ECS

[MR19] Volodymyr L. Makarov and Nataliia M. Romaniuk. Exponentially convergent symbolic algorithm of the functionaldiscrete method for the fourth order Sturm–Liouville problems with polynomial coefficients. Journal of Computational and Applied Mathematics, 358(??):405–423, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301499.

Martinez:2015:DOC

[MRMA15] S. Martínez, M. Rueda, H. Martínez, and A. Arcos. Determining P optimum calibration points to construct calibration estimators of the distribution function. Journal of Computational and Applied Mathematics, 275(??):281–293, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003392.

Martinez:2017:ODO

[MRMA17] S. Martínez, M. Rueda, H. Martínez, and A. Arcos. Optimal dimension and optimal auxiliary vector to construct calibration estimators of the distribution function. Journal of Computational and Applied Mathematics, 318(??):444– 459, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300413.

Morigi:2017:FTR

[MRS17] Serena Morigi, Lothar Reichel, and Fiorella Sgallari. Fractional Tikhonov regularization with a nonlinear penalty term. Journal of Computational and Applied Mathematics, 324(??):142–154, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301723.

Mach:2016:ACA

[MRVV16] T. Mach, L. Reichel, M. Van Barel, and R. Vandebril. Adaptive cross approximation for ill-posed problems. Journal of Computational and Applied Mathematics, 303 (??):206–217, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630067X.

Meng:2015:OCM

[MS15a] Qingxin Meng and Yang Shen. Optimal control of mean-field jump-diffusion systems with delay: a stochastic maximum principle approach. Journal of Computational and Applied Mathematics, 279 (??):13–30, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714004579.

Muller:2015:FEM

[MS15b] Fabian L. Müller and Christoph Schwab. Finite elements with mesh refinement for wave equations in polygons. Journal of Computational and Applied Mathematics, 283(??):163–181, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000114.

Maleknejad:2017:HFM

[MS17a] K. Maleknejad and E. Saeedipoor. Hybrid function method and convergence analysis for twodimensional nonlinear integral equations. Journal of Computational and Applied Mathematics, 322(??):96–108, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301310.

Merrien:2017:EHS

[MS17b] Jean-Louis Merrien and Tomas Sauer. Extended Hermite subdivision schemes. Journal of Computational and Applied Mathematics, 317 (??):343–361, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716305957.

Mirzaee:2017:AOM

[MS17c] Farshid Mirzaee and Nasrin Samadyar. Application of operational matrices for solving system of linear Stratonovich Volterra integral equation. Journal of Computational and Applied Mathematics, 320: 164–175, August 2017.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Mnatsakanov:2017:RQF

[MS17d] Robert M. Mnatsakanov and Aleksandre Sborshchikovi. Recovery of a quantile function from moments. Journal of Computational and Applied Mathematics, 315(??):354–364, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305702.

Mohammadi:2017:CAG

[MS17e] M. Mohammadi and R. Schaback. Convergence analysis of general spectral methods. Journal of Computational and Applied Mathematics, 313(??):284–293, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304514.

Maulik:2018:EIC

[MS18] Romit Maulik and Omer San. Explicit and implicit LES closures for Burgers turbulence. Journal of Computational and Applied Mathematics, 327(??):12–40, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303035.

Medina:2019:SHO

[MS19a] Arcesio Castañeda Medina and Rochus Schmid. Solution of high order compact discretized 3D elliptic partial differential equations by an accelerated multigrid method. Journal of Computational and Applied Mathematics, 350(??):343– 352, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830640X.

Meurant:2019:CSP

[MS19b] Gérard Meurant and Alvise Sommariva. On the computation of sets of points with low Lebesgue constant on the unit disk. Journal of Computational and Applied Mathematics, 345:388–404, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Mozartova:2015:CBM

[MSH15] A. Mozartova, I. Savostianov, and W. Hundsdorfer. Comparison of boundedness and monotonicity properties of one-leg and linear multistep methods. Journal of Computational and Applied Mathematics, 279(??):159–172, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004737.

Mirzaee:2018:EPS

[MSH18] Farshid Mirzaee, Nasrin Samadyar, and Seyede Fatemeh Hoseini. Euler polynomial solutions of nonlinear stochastic Itô-Volterra integral equations. Journal of Computational and Applied Mathematics, 330(??):574–585, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304223.

Minev:2018:FFP

[MSV18] Peter Minev, Shriram Srinivasan, and Petr N. Vabishchevich. Flux formulation of parabolic equations with highly heterogeneous coefficients. Journal of Computational and Applied Mathematics, 340(??):582–601, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717306179.

Meng:2016:OIR

[MSY16] Hui Meng, Tak Kuen Siu, and Hailiang Yang. Optimal insurance risk control with multiple reinsurers. Journal of Computational and Applied Mathematics, 306(??):40–52, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301698.

Meng:2017:NOI

[MSY17] Hui Meng, Tak Kuen Siu, and Hailiang Yang. A note on optimal insurance risk control with multiple reinsurers. Journal of Computational and Applied Mathematics, 319(??):38–42, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306458.

Miyatake:2018:EBS

[MSZ18] Yuto Miyatake, Tomohiro Sogabe, and Shao-Liang Zhang. On the equivalence between SOR-type methods for linear systems and the discrete gradient methods for gradient systems. Journal of Computational and Applied Mathematics, 342(??):58–69, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830181X.

Marrero:2015:SRD

[MT15] J. Abderramán Marrero and V. Tomeo. Some results on determinants and inverses of nonsingular pentadiagonal matrices. Journal of Computational and Applied Mathematics, 275(??):447–455, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714001563.

Maeda:2016:GEA

[MT16a] Kazuki Maeda and Satoshi Tsujimoto. A generalized eigenvalue algorithm for tridiagonal matrix pencils based on a nonautonomous discrete integrable system. Journal of Computational and Applied Mathematics, 300(??):134– 154, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715006445.

Milstein:2016:LMS

[MT16b] G. N. Milstein and M. V. Tretyakov. Layer methods for stochastic Navier– Stokes equations using simplest characteristics. Journal of Computational and Applied Mathematics, 302(??):1–23, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300310.

Marrero:2017:FRN

[MT17a] J. Abderramán Marrero and V. Tomeo. A fast and reliable numerical solver for general bordered ktridiagonal matrix linear equations. Journal of Computational and Applied Mathematics, 318 (??):211–219, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716304939.

Marrero:2017:CRS

[MT17b] J. Abderramán Marrero and V. Tomeo. On compact representations for the solutions of linear difference equations with variable coefficients. Journal of Computational and Applied Mathematics, 318(??):411– 421, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301054.

Mojarrab:2017:GLL

[MT17c] M. Mojarrab and F. Toutounian. Global LSMR(Gl-LSMR) method for solving general linear systems with several right-hand sides. Journal of Computational and Applied Mathematics, 321(??):78–89, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300699.

Melnikov:2018:NCS

[MT18] Yuri B. Melnikov and Irina V. Turuntaeva. On a new class of special functions generated by integral-difference operators. Journal of Computational and Applied Mathematics, 327(??):167–174, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042717302893.

Mukaihira:2019:DCB

[MT19] Atsushi Mukaihira and Yoshimasa Tsuboi. DSF chain for biorthogonal polynomials and its application to matrix eigenvalue problems. Journal of Computational and Applied Mathematics, 347(??):273–295, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301699.

Mac:2015:RBP

[MTCC15] D. H. Mac, Z. Tang, S. Clénet, and E. Creusé. Residual-based a posteriori error estimation for stochastic magnetostatic problems. Journal of Computational and Applied Mathematics, 289(??):51–67, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001806.

Mizuguchi:2017:NVE

[MTKO17] Makoto Mizuguchi, Akitoshi Takayasu, Takayuki Kubo, and Shin’ichi Oishi. Numerical verification for existence of a global-in-time solution to semilinear parabolic equations. Journal of Computational and Applied Mathematics, 315(??):1– 16, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305131.

Meng:2019:PCV

[MTMZ19] Fanyong Meng, Jie Tang, Beiling Ma, and Qiang Zhang. Proportional coalition values for monotonic games on convex geometries with a coalition structure. Journal of Computational and Applied Mathematics, 348(??):34–47, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302450.

Muftahov:2017:NSV

[MTS17] Ildar Muftahov, Aleksandr Tynda, and Denis Sidorov. Numeric solution of Volterra integral equations of the first kind with discontinuous kernels. Journal of Computational and Applied Mathematics, 313(??):119–128, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630423X.

Marrero:2015:IIH

[MTT15] J. Abderramán Marrero, V. Tomeo, and E. Torrano. On inverses of infinite Hessenberg matrices. Journal of Computational and Applied Mathematics, 275(??):356–365, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271400329X.

M:2017:PRH

[MTV17] Chandru M., Prabha T., and Shanthi V. A parameter robust higher order numerical method for singularly perturbed two parameter problems with non-smooth data. Journal of Computational and Applied Mathematics, 309(??):11–27, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302795.

Mu:2019:PPE

[Mu19a] Lin Mu. A priori and a posterior error estimate of new weak Galerkin finite element methods for second order elliptic interface problems on polygonal meshes. Journal of Computational and Applied Mathematics, 362 (??):423–442, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718305521.

Mu:2019:WGB

[Mu19b] Lin Mu. Weak Galerkin based a posteriori error estimates for second order elliptic interface problems on polygonal meshes. Journal of Computational and Applied Mathematics, 361(??):413–425, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302213.

Mudzimbabwe:2019:SNS

[Mud19] Walter Mudzimbabwe. A simple numerical solution for an optimal investment strategy for a DC pension plan in a jump diffusion model. Journal of Computational and Applied Mathematics, 360(??):55–61, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301748.

Muslu:2017:NSB

[Mus17] Gulcin M. Muslu. Numerical study of blow-up to the purely elliptic generalized Davey–Stewartson system. Journal of Computational and Applied Mathematics, 317(??):331– 342, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305969.

Matysik:2016:SIM

[MV16a] Oleg Matysik and Marc M. Van Hulle. Simpleiteration method with alternating step size for solving operator equations in Hilbert space. Journal of Computational and Applied Mathematics, 300 (??):290–299, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716000030.

Minev:2016:OSS

[MV16b] Peter Minev and Petr N. Vabishchevich. An operatorsplitting scheme for the stream function-vorticity formulation of the unsteady Navier–Stokes equations. Journal of Computational and Applied Mathematics, 293(??):147–163, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000278.

Mitchell:2016:ANS

[MV16c] S. L. Mitchell and M. Vynnycky. On the accurate numerical solution of a twophase Stefan problem with phase formation and depletion. Journal of Computational and Applied Mathematics, 300(??):259– 274, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006330.

Mudzimbabwe:2016:ISP

[MV16d] Walter Mudzimbabwe and Lubin Vulkov. IMEX schemes for a parabolic-ODE system of European options with liquidity shocks. Journal of Computational and Applied Mathematics, 299(??):245– 256, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715006202.

Minev:2018:SSS

[MV18] Peter Minev and Petr N. Vabishchevich. Splitting schemes for the stress formulation of the incompressible Navier–Stokes equations. Journal of Computational and Applied Mathematics, 344 (??):807–818, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301991.

McKee:2019:EDP

[MVC19] S. McKee, M. Vynnycky, and J. A. Cuminato. An elementary diffusion problem, Laplace transforms and novel mathematical identities. Journal of Computational and Applied Mathematics, 353(??):113– 119, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307477.

Martin-Vaquero:2017:HON

[MVdRE^{+}17] J. Martín-Vaquero, A. Martín del Rey, A. H. Encinas, J. D. Hernández Guillén, A. Queiruga-Dios, and G. Rodríguez Sánchez. Higher-order nonstandard finite difference schemes for a MSEIR model for a malware propagation. Journal of Computational and Applied Mathematics, 317(??):146– 156, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305866.

Martin-Vaquero:2019:EFO

[MVK19] J. Martín-Vaquero and A. Kleefeld. ESERK5: a fifth-order extrapolated stabilized explicit Runge– Kutta method. Journal of Computational and Applied Mathematics, 356(??):22–36, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300561.

Macias:2019:TMS

[MVM19] Siegfried Macías and José Villa-Morales. Temporal monotonicity of the solutions of some semilinear parabolic equations with fractional diffusion. Journal of Computational and Applied Mathematics, 354(??):163– 173, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830596X.

Mendoza:2015:RLT

[MVPN15] C. Díaz Mendoza, P. González Vera, M. Jiménez Paiz, and O. Njåstad. Radau and Lobatto-type quadratures associated with strong Stieltjes distributions. Journal of Computational and Applied Mathematics, 284(??):184–201, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004853.

Martin-Vaquero:2018:VSL

[MVQDdR^{+}18] J. Martín-Vaquero, A. Queiruga-Dios, A. Martín del Rey, A. H. Encinas, J. D. Hernández Guillén, and G. Rodríguez Sánchez. Variable step length algorithms with high-order extrapolated non-standard finite difference schemes for a SEIR model. Journal of Computational and Applied Mathematics, 330 (??):848–854, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717301541.

Mu:2015:WGF

[MWY15a] Lin Mu, Junping Wang, and Xiu Ye. A weak Galerkin finite element method with polynomial reduction. Journal of Computational and Applied Mathematics, 285(??):45–58, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500059X.

Mu:2015:MWG

[MWY15b] Lin Mu, Xiaoshen Wang, and Xiu Ye. A modified weak Galerkin finite element method for the Stokes equations. Journal of Computational and Applied Mathematics, 275(??):79–90, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400363X.

Mu:2016:HFW

[MWY16a] Lin Mu, Junping Wang, and Xiu Ye. A hybridized formulation for the weak Galerkin mixed finite element method. Journal of Computational and Applied Mathematics, 307(??):335–345, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271600008X.

Mu:2016:WGG

[MWY16b] Lin Mu, Junping Wang, and Xiu Ye. A weak Galerkin Generalized Multiscale Finite Element Method. Journal of Computational and Applied Mathematics, 305(??):68– 81, October 15, 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716301376.

Ma:2016:SCR

[MX16] Yong Ma and Weidong Xu. Structural credit risk modelling with Hawkes jump diffusion processes. Journal of Computational and Applied Mathematics, 303(??):69–80, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300796.

Mu:2017:GEM

[MXW17] Lin Mu, Kelin Xia, and Guowei Wei. Geometric and electrostatic modeling using molecular rigidity functions. Journal of Computational and Applied Mathematics, 313(??):18–37, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630379X.

Mu:2018:SFE

[MY18] Lin Mu and Xiu Ye. A simple finite element method for linear hyperbolic problems. Journal of Computational and Applied Mathematics, 330(??):330–339, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304144.

Ma:2018:CIP

[MYG18] Qinghua Ma, Fahuai Yi, and Chonghu Guan. A consumption-investment problem with constraints on minimum and maximum consumption rates. Journal of Computational and Applied Mathematics, 338(??):185–198, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718300773.

Mao:2016:ASN

[MYM16] Wei Mao, Surong You, and Xuerong Mao. On the asymptotic stability and numerical analysis of solutions to nonlinear stochastic differential equations with jumps. Journal of Computational and Applied Mathematics, 301(??):1–15, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271600039X.

Ma:2016:MMM

[MZ16] Jingtang Ma and Zhiqiang Zhou. Moving mesh methods for pricing Asian options with regime switching. Journal of Computational and Applied Mathematics, 298(??):211–221, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005816.

Mu:2019:IWG

[MZ19] Lin Mu and Xu Zhang. An immersed weak Galerkin method for elliptic interface problems. Journal of Computational and Applied Mathematics, 362 (??):471–483, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718305004.

Nassar:2018:NEW

[NADK18] Mazen Nassar, Ahmed Z.

Afify, Sanku Dey, and Devendra Kumar. A new extension of Weibull distribution: Properties and different methods of estimation. Journal of Computational and Applied Mathematics, 336(??):439– 457, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306076.

Nasab:2018:SSD

[NAH18] M. Hosseini Nasab, A. Abdi, and G. Hojjati. Symmetric second derivative integration methods. Journal of Computational and Applied Mathematics, 330(??):618–629, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304478.

Nakano:2015:QMC

[Nak15] Yumiharu Nakano. Quasi-Monte Carlo methods for Choquet integrals. Journal of Computational and Applied Mathematics, 287(??):63–66, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271500179X.

Nassar:2017:EIW

[NAK17] M. Nassar and O. E. Abo-Kasem. Estimation of the inverse Weibull parameters under adaptive type-II progressive hybrid censoring scheme. Journal of Computational and Applied Mathematics, 315 (??):228–239, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716305453.

Novin:2018:NFM

[NAM18] Reza Novin, Mohammad Ali Fariborzi Araghi, and Yaghoub Mahmoudi. A novel fast modification of the Adomian decomposition method to solve integral equations of the first kind with hypersingular kernels. Journal of Computational and Applied Mathematics, 343(??):619–634, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302504.

Nategh:2019:RDC

[Nat19] Mehdi Nategh. Recursive dynamics in a class of systems on q-timescales with unbounded delays. Journal of Computational and Applied Mathematics, 352(??):401–408, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307465.

Narang:2019:GEC

[NBAK19] Mona Narang, Saurabh Bhatia, Ali Saleh Alshomrani, and V. Kanwar. General efficient class of Steffensen type methods with memory for solving systems of nonlinear equations. Journal of Computational and Applied Mathematics, 352(??):23– 39, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306563.

Nogueras:2016:SSB

[NC16] Rafael Nogueras and Carlos Cotta. Studying self-balancing strategies in island-based multimemetic algorithms. Journal of Computational and Applied Mathematics, 293(??):180–191, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002009.

Nadarajah:2019:EDS

[NC19] Saralees Nadarajah and Stephen Chan. The exact distribution of the sum of stable random variables. Journal of Computational and Applied Mathematics, 349(??):187–196, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718305958.

Nunes:2019:ECV

[NCF^{+}19] Célia Nunes, Gilberto Capistrano, Dário Ferreira, Sandra S. Ferreira, and João Tiago Mexia. Exact critical values for one-way fixed effects models with random sample sizes. Journal of Computational and Applied Mathematics, 354(??):112– 122, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303340.

Nadarajah:2016:MBD

[NCJ16] Saralees Nadarajah, Jeffrey Chu, and Xiao Jiang. On moment based density approximations for aggregate losses. Journal of Computational and Applied Mathematics, 298(??):152–166, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006196.

Nicolopoulos:2019:SFR

[NCPD19] A. Nicolopoulos, M. Campos-Pinto, and B. Després. A stable formulation of resonant Maxwell’s equations in cold plasma. Journal of Computational and Applied Mathematics, 362(??):185–204, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302419.

Naitsat:2018:GAD

[NCQ^{+}18] Alexander Naitsat, Shichao Cheng, Xiaofeng Qu, Xin Fan, Emil Saucan, and Yehoshua Y. Zeevi. Geometric approach to detecting volumetric changes in medical images. Journal of Computational and Applied Mathematics, 329(??):37–50, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303758.

Nikooeinejad:2016:NSO

[NDH16] Z. Nikooeinejad, A. Delavarkhalafi, and M. Heydari. A numerical solution of open-loop Nash equilibrium in nonlinear differential games based on Chebyshev pseudospectral method. Journal of Computational and Applied Mathematics, 300(??):369– 384, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716000303.

Nahid:2019:PMP

[NDN19] Nilofar Nahid, Payel Das, and Gnaneshwar Nelakanti. Projection and multi projection methods for nonlinear integral equations on the half-line. Journal of Computational and Applied Mathematics, 359(??):119–144, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301736.

Najafalizadeh:2017:BPO

[NE17] S. Najafalizadeh and R. Ezzati. A block pulse operational matrix method for solving two-dimensional nonlinear integro-differential equations of fractional order. Journal of Computational and Applied Mathematics, 326(??):159–170, December 15, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717302923.

Nemati:2015:NSV

[Nem15]  S. Nemati.  Numeri-

cal solution of Volterra– Fredholm integral equations using Legendre collocation method. Journal of Computational and Applied Mathematics, 278(??):29–36, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004373.

Noack:2017:HGD

[NF17] Marcus M. Noack and Simon W. Funke. Hybrid genetic deflated Newton method for global optimisation. Journal of Computational and Applied Mathematics, 325(??):97–112, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730225X.

Nieto:2018:CSM

[NGGA^{+}18] P. J. García Nieto, E. García-Gonzalo, J. C. Álvarez Antón, V. M. González Suárez, R. Mayo Bayón, and F. Mateos Martín. A comparison of several machine learning techniques for the centerline segregation prediction in continuous cast steel slabs and evaluation of its performance. Journal of Computational and Applied Mathematics, 330(??):877–895, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300961.

Nieto:2016:HPO

[NGGFM16] P. J. García Nieto, E. García-Gonzalo, J. R. Alonso Fernández, and C. Díaz Muñiz. A hybrid PSO optimized SVM-based model for predicting a successful growth cycle of the spirulina platensis from raceway experiments data. Journal of Computational and Applied Mathematics, 291(??):293–303, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000217.

Nieto:2017:HWK

[NGGFM17] P. J. García Nieto, E. García-Gonzalo, J. R. Alonso Fernández, and C. Díaz Muñiz. A hybrid wavelet kernel SVM-based method using artificial bee colony algorithm for predicting the cyanotoxin content from experimental cyanobacteria concentrations in the Trasona reservoir (Northern Spain). Journal of Computational and Applied Mathematics, 309(??):587–602, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716300255.

Nieto:2019:FHH

[NGGL^{+}19] P. J. García Nieto, E. García-Gonzalo, F. Sánchez Lasheras, J. P. Paredes-Sánchez, and P. Riesgo Fernández. Forecast of the higher heating value in biomass torrefaction by means of machine learning techniques. Journal of Computational and Applied Mathematics, 357(??):284–301, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301281.

Ngondiep:2019:LTU

[Ngo19a]  Eric Ngondiep.  Long

time unconditional stability of a two-level hybrid method for nonstationary incompressible Navier– Stokes equations. Journal of Computational and Applied Mathematics, 345: 501–514, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Ngondiep:2019:SAM

[Ngo19b] Eric Ngondiep. Stability analysis of MacCormack rapid solver method for evolutionary Stokes– Darcy problem. Journal of Computational and Applied Mathematics, 345: 269–285, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Nikooeinejad:2019:NEA

[NH19] Z. Nikooeinejad and M. Heydari. Nash equilibrium approximation of some class of stochastic differential games: a combined Chebyshev spectral collocation method with policy iteration. Journal of Computational and Applied Mathematics, 362(??):41–54, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302493.

Nasab:2017:GSS

[NHA17] M. Hosseini Nasab, G. Hojjati, and A. Abdi. Gsymplectic second derivative general linear methods for Hamiltonian problems. Journal of Computational and Applied Mathematics, 313(??):486–498, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305003.

Nievergelt:2015:UIC

[Nie15] Yves Nievergelt. Uncorrelated, or increasing convex, positive data with multiple least-squares exponential curves. Journal of Computational and Applied Mathematics, 287(??):44–47, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001594.

Navascues:2019:GTI

[NJCS19] M. A. Navascués, Sangita Jha, A. K. B. Chand, and M. V. Sebastián. Generalized trigonometric interpolation. Journal of Computational and Applied Mathematics, 354(??):152– 162, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304801.

Nassar:2019:MOA

[NKD^{+}19] Mazen Nassar, Devendra Kumar, Sanku Dey, Gauss M. Cordeiro, and Ahmed Z. Afify. The Marshall–Olkin alpha power family of distributions with applications. Journal of Computational and Applied Mathematics, 351 (??):41–53, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718306605.

Nkeki:2018:OIR

[Nke18a] Charles I. Nkeki. Optimal investment risks and debt management with backup security in a financial crisis. Journal of Computational and Applied Mathematics, 338(??):129–152, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300694.

Nkeki:2018:OPF

[Nke18b] Charles I. Nkeki. Optimal pension fund management in a jump-diffusion environment: Theoretical and empirical studies. Journal of Computational and Applied Mathematics, 330(??):228–252, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730359X.

Nadarajah:2017:EDS

[NL17] Saralees Nadarajah and Rui Li. The exact density of the sum of independent skew normal random variables. Journal of Computational and Applied Mathematics, 311(??):1–10, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303041.

Nivoliers:2015:AFS

[NLG15] Vincent Nivoliers, Bruno Lévy, and Christophe Geuzaine. Anisotropic and feature sensitive triangular remeshing using normal lifting. Journal of Computational and Applied Mathematics, 289(??):225–240, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000631.

Najafi:2017:BPU

[NM17] Ali Reza Najafi and Farshid Mehrdoust. Bond pricing under mixed generalized CIR model with mixed Wishart volatility process. Journal of Computational and Applied Mathematics, 319(??):108– 116, August 1, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717300031.

Nanda:2019:PEG

[NMKK19] Asok K. Nanda, Sudhansu S. Maiti, Chanchal Kundu, and Amarjit Kundu. Parameter estimates of general failure rate model: a Bayesian approach. Journal of Computational and Applied Mathematics, 351(??):317–330, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306678.

Niu:2018:NSS

[NML18] Datian Niu, Jiana Meng, and Hongying Li. A new shift strategy for the implicitly restarted refined harmonic Lanczos method. Journal of Computational and Applied Mathematics, 336(??):458– 467, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306581.

Nguyen:2017:MIE

[NN17] Dong T. P. Nguyen and Dirk Nuyens. Multivariate integration over R^{s} with exponential rate of convergence. Journal of Computational and Applied Mathematics, 315(??):327–342, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305490.

Negarchi:2018:NSV

[NN18a] Neda Negarchi and Kazem Nouri. Numerical solution of Volterra–Fredholm integral equations using the collocation method based on a special form of the Müntz–Legendre polynomials. Journal of Computational and Applied Mathematics, 344(??):15– 24, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718303066.

Nikseresht:2018:NNN

[NN18b] Asiye Nikseresht and Alireza Nazemi. A novel neural network model for solving a class of nonlinear semidefinite programming problems. Journal of Computational and Applied Mathematics, 338(??):69–79, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300529.

Nikseresht:2019:NNN

[NN19] Asiye Nikseresht and Alireza Nazemi. A novel neural network for solving semidefinite programming problems with some applications. Journal of Computational and Applied Mathematics, 350(??):309– 323, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306320.

Nguyen:2015:EGT

[NP15] Hao Nguyen and Guergana Petrova. Extended Gaussian type cubatures for the ball. Journal of Computational and Applied Mathematics, 290 (??):209–223, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003040.

Nowak:2019:PEO

[NP19] Piotr Nowak and Michał Pawłowski. Pricing European options under uncertainty with application of Levy processes and the minimal L^{q} equivalent martingale measure. Journal of Computational and Applied Mathematics, 345:

416–433, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Nadeem:2018:SMI

[NPR18] Aamir Nadeem, Roland Potthast, and Andreas Rhodin. On sequential multiscale inversion and data assimilation. Journal of Computational and Applied Mathematics, 336 (??):338–352, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717304028.

Novara:2015:BBD

[NR15] Paola Novara and Lucia Romani. Building blocks for designing arbitrarily smooth subdivision schemes with conic precision. Journal of Computational and Applied Mathematics, 279(??):67– 79, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004701.

Noschese:2019:CUS

[NR19] Silvia Noschese and Lothar Reichel. Computing unstructured and structured polynomial pseudospectrum approximations. Journal of Computational and Applied Mathematics, 350(??):57– 68, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305843.

Nunez:2018:ARS

[NRV18] Luigi Nuñez, Rommel G.

Regis, and Kayla Varela. Accelerated random search for constrained global optimization assisted by radial basis function surrogates. Journal of Computational and Applied Mathematics, 340(??):276–295, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300888.

Neitzel:2016:GHM

[NS16] Frank Neitzel and Burkhard Schaffrin. On the Gauss– Helmert model with a singular dispersion matrix where BQ is of smaller rank than B. Journal of Computational and Applied Mathematics, 291(??):458–467, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001521.

Nechepurenko:2017:CHM

[NS17] Yu. M. Nechepurenko and M. Sadkane. Computing humps of the matrix exponential. Journal of Computational and Applied Mathematics, 319(??):87–96, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306422.

Nikolovski:2018:CSD

[NS18] Filip Nikolovski and Irena Stojkovska. Complexstep derivative approximation in noisy environment. Journal of Computational and Applied Mathematics, 327(??):64–78, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302996.

Nazemi:2019:NGB

[NS19] Alireza Nazemi and Atiye Sabeghi. A novel gradientbased neural network for solving convex secondorder cone constrained variational inequality problems. Journal of Computational and Applied Mathematics, 347(??):343–356, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830517X.

Nemati:2016:FNA

[NSM16] S. Nemati, S. Sedaghat, and I. Mohammadi. A fast numerical algorithm based on the second kind Chebyshev polynomials for fractional integrodifferential equations with weakly singular kernels. Journal of Computational and Applied Mathematics, 308(??):231–242, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302825.

Nhan:2018:OUC

[NSV18] Thái Anh Nhan, Martin Stynes, and Relja Vulanović. Optimal uniformconvergence results for convection-diffusion problems in one dimension using preconditioning. Journal of Computational and Applied Mathematics, 338(??):227–238, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300839.

Ngo:2015:BTP

[NSW15] L. X. C. Ngô, J. R. Sendra, and F. Winkler. Birational transformations preserving rational solutions of algebraic ordinary differential equations. Journal of Computational and Applied Mathematics, 286(??):114–127, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001533.

Nie:2017:SAE

[NSYX17] Cunyun Nie, Shi Shu, Haiyuan Yu, and Wenhua Xia. Superconvergence and asymptotic expansions for bilinear finite volume element approximation on non-uniform grids. Journal of Computational and Applied Mathematics, 321(??):323–335, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306343.

Ng:2015:CBP

[NT15] Tuen Wai Ng and Chiu Yin Tsang. Chebyshev– Blaschke products: Solutions to certain approximation problems and differential equations. Journal of Computational and Applied Mathematics, 277(??):106–114, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003859.

Nguyen-Thoi:2015:UBL

[NTPVNTDT15] T. Nguyen-Thoi, P. Phung-Van, M. H. Nguyen-Thoi, and H. Dang-Trung. An upper-bound limit analysis of Mindlin plates using CS–DSG3 method and second-order cone programming. Journal of Computational and Applied Mathematics, 281 (??):32–48, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714005470.

Navarro:2019:LTN

[NV19] Juan F. Navarro and Juan Vargas. On the long term numerical integration of planetary rings. Journal of Computational and Applied Mathematics, 354(??):390– 401, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304229.

Neilan:2019:DMT

[NW19] Michael Neilan and Mohan Wu. Discrete Miranda– Talenti estimates and applications to linear and nonlinear PDEs. Journal of Computational and Applied Mathematics, 356(??):358–376, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300482.

Niu:2018:NSN

[NXLX18] Jing Niu, Minqiang Xu, Yingzhen Lin, and Qing Xue. Numerical solution of nonlinear singular boundary value problems. Journal of Computational and Applied Mathematics, 331(??):42–51, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304715.

Nevruz:2019:SIT

[NY19] Ezgi Nevruz and S. Kasirga Yildirak. Spatiotemporal interpolation through an extension of differential evolution algorithm for agricultural insurance claims. Journal of Computational and Applied Mathematics, 352(??):278–292, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830712X.

Nguyen:2015:NHO

[NZ15] Duc Duy Nguyen and Shan Zhao. A new high order dispersive FDTD method for Drude material with complex interfaces. Journal of Computational and Applied Mathematics, 285(??):1–14, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000606.

Nedjar:2016:GLD

[NZ16] Sihem Nedjar and Halim Zeghdoudi. On gamma Lindley distribution: Properties and simulations. Journal of Computational and Applied Mathematics, 298(??):167–174, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715006184.

Ning:2017:ISC

[NZL17] Boming Ning, Di Zhao, and Hongyi Li. Improved Schur complement preconditioners for block-Toeplitz systems with small size blocks. Journal of Computational and Applied Mathematics, 311(??):655–663, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303995.

Omidi:2017:EDM

[OAN17] Farahnaz Omidi, Behzad Abbasi, and Alireza Nazemi. An efficient dynamic model for solving a portfolio selection with uncertain chance constraint models. Journal of Computational and Applied Mathematics, 319(??):43–55, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306306.

Okasha:2019:GLE

[OAS19] Hassan M. Okasha and Ali A. Al-Shomrani. Generalized linear exponential geometric distributions and its applications. Journal of Computational and Applied Mathematics, 351(??):198–211, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306496.

Orti:2017:GDC

[OCU17] José A. López Ortí, Francisco José Marco Castillo, and María José Martínez Usó. Geometrical definition of a continuous family of time transformations generalizing and including the classic anomalies of the elliptic two-body problem. Journal of Computational and Applied Mathematics, 309(??):482–492, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300978.

Orti:2018:GDC

[OCU18] José A. López Ortí, Francisco José Marco Castillo, and María José Martínez Usó. Geometrical definition of a continuous family of time transformations on the hyperbolic two-body problem.

Journal of Computational and Applied Mathematics, 330(??):1081–1092, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302261.

Okan:2017:SCH

[OD17] Ayla Okan and Metin Demiralp. A self-consistent high dimensional modelling based decomposition approach for univariate linear integral operators: Tridiagonal Kernel Enhanced Multivariance Products Representation (TKEMPR). Journal of Computational and Applied Mathematics, 326(??):99–115, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302698.

Ozkut:2019:RAU

[OE19] Murat Ozkut and Serkan Eryilmaz. Reliability analysis under Marshall– Olkin run shock model. Journal of Computational and Applied Mathematics, 349(??):52–59, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305673.

Ortiz-Gracia:2016:EWB

[OG16] Luis Ortiz-Gracia. Efficient wavelets-based valuation of synthetic CDO tranches. Journal of Computational and Applied Mathematics, 292 (??):562–575, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003878.

Orti:2017:NBP

[OGR17a] José A. López Ortí, Vicente Agost Gómez, and Miguel Barreda Rochera. A new bi-parametric family of temporal transformations to improve the integration algorithms in the study of the orbital motion. Journal of Computational and Applied Mathematics, 318(??):479– 490, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300371. See

corrigendum [?].

Orti:2017:TAC

[OGR17b] José A. López Ortí, Manuel Forner Gumbau, and Miguel Barreda Rochera. Two algorithms to construct a consistent first order theory of equilibrium figures of close binary systems. Journal of Computational and Applied Mathematics, 318(??):14– 25, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306252.

Orti:2019:IAS

[OGR19] José A. López Ortí, Manuel Forner Gumbau, and Miguel Barreda Rochera. An improved algorithm of second order to construct consistent theories of equilibrium figures of rotating celestial bodies. Journal of Computational and Applied Mathematics, 354(??):402– 413, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305156.

Ogulenko:2017:APS

[Ogu17] Aleksey Ogulenko. Asymptotical properties of social network dynamics on time scales. Journal of Computational and Applied Mathematics, 319(??):413–422, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300523.

Ogata:2018:NIB

[OH18] Hidenori Ogata and Hiroshi Hirayama. Numerical integration based on hyperfunction theory. Journal of Computational and Applied Mathematics, 327(??):243–259, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303187.

Okasha:2016:NFM

[OK16] Hassan M. Okasha and M. Kayid. A new family of Marshall–Olkin extended generalized linear exponential distribution. Journal of Computational and Applied Mathematics, 296(??):576–592, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005178.

Ouannas:2019:CFO

[OKGB19] Adel Ouannas, Amina-Aicha Khennaoui, Giuseppe Grassi, and Samir Bendoukha. On chaos in the fractional-order Grassi– Miller map and its control. Journal of Computational and Applied Mathematics, 358(??):293–305, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301566.

Ortobelli:2016:ASD

[OLPT16] Sergio Ortobelli, Tommaso Lando, Filomena Petronio, and Tomas Tichý. Asymptotic stochastic dominance rules for sums of i.i.d. random variables. Journal of Computational and Applied Mathematics, 300(??):432– 448, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006299.

Ordonez:2019:HAS

[OLRPdCJ19] Celestino Ordóñez, Fernando Sánchez Lasheras, Javier Roca-Pardiñas, and Francisco Javier de Cos Juez. A hybrid ARIMA–SVM model for the study of the remaining useful life of aircraft engines. Journal of Computational and Applied Mathematics, 346(??):184–191, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830428X.

Obradovic:2017:SCN

[OM17a] Maja Obradović and Marija Milosević. Stability of a class of neutral stochastic differential equations with unbounded delay and Markovian switching and the Euler–Maruyama method. Journal of Computational and Applied Mathematics, 309(??):244–266, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303107.

Ozalp:2017:FLT

[OM17b] Nuri Ozalp and Ozlem Ozturk Mizrak. Fractional Laplace transform method in the framework of the CTIT transformation. Journal of Computational and Applied Mathematics, 317(??):90– 99, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716305817.

Ozaki:2015:IEF

[OOO15] Katsuhisa Ozaki, Takeshi Ogita, and Shin’ichi Oishi. Improvement of errorfree splitting for accurate matrix multiplication. Journal of Computational and Applied Mathematics, 288(??):127–140, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002289.

ORiordan:2019:NAS

[OP19] E. O’Riordan and M. L. Pickett. Numerical approximations to the scaled first derivatives of the solution to a two parameter singularly perturbed problem. Journal of Computational and Applied Mathematics, 347(??):128–149, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304813.

Orr:2019:GLS

[Orr19]  Derek Orr.  General-

ized log-sine integrals and Bell polynomials. Journal of Computational and Applied Mathematics, 347(??):330–342, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305132.

Ottino:2015:BPB

[OS15] D. Ottino and S. Scacchi. BPX preconditioners for the bidomain model of electrocardiology. Journal of Computational and Applied Mathematics, 285(??):151–168, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000722.

Orava:2015:MPM

[OSC15] Vít Orava, Ondrej Soucek, and Peter Cendula. Multiphase modeling of nonisothermal reactive flow in fluidized bed reactors. Journal of Computational and Applied Mathematics, 289(??):282–295, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000242.

Oozawa:2018:RBC

[OSMZ18] Masaya Oozawa, Tomohiro Sogabe, Yuto Miyatake, and Shao-Liang Zhang. On a relationship between the t-congruence Sylvester equation and the Lyapunov equation. Journal of Computational and Applied Mathematics, 329(??):51–56, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302972.

Orlando:2017:RIV

[OT17] Giuseppe Orlando and Giovanni Taglialatela. A review on implied volatility calculation. Journal of Computational and Applied Mathematics, 320: 202–220, August 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Ou:2018:NGC

[Ou18]  Yigui Ou.  A note

on the global convergence theorem of accelerated adaptive Perry conjugate gradient methods. Journal of Computational and Applied Mathematics, 332(??):101–106, April 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042717305253.

Oste:2017:TTM

[OV17] Roy Oste and Joris Van der Jeugt. Tridiagonal test matrices for eigenvalue computations: Two-parameter extensions of the Clement matrix. Journal of Computational and Applied Mathematics, 314(??):30– 39, April 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305088.

Oyarzua:2017:ACF

[OZ17] Ricardo Oyarzúa and Paulo Zúñiga. Analysis of a conforming finite element method for the Boussinesq problem with temperaturedependent parameters. Journal of Computational and Applied Mathematics, 323(??):71–94, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301644.

Onchis:2018:RAA

[OZ18] Darian M. Onchis and Simone Zappalà. Realizable algorithm for approximating Hilbert–Schmidt operators via Gabor multipliers. Journal of Computational and Applied Mathematics, 337(??):119–124, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830027X.

Pekalp:2019:ASI

[PA19] Mustafa Hilmi Pekalp and Halil Aydogdu. An asymptotic solution of the integral equation for the second moment function in geometric processes. Journal of Computational and Applied Mathematics, 353(??):179– 190, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307453.

Pathak:2015:CFP

[PAC15] Hemant Kumar Pathak, Ravi P. Agarwal, and Yeol Je Cho. Coincidence and fixed points for multi-valued mappings and its application to nonconvex integral inclusions. Journal of Computational and Applied Mathematics, 283(??):201–217, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005755. See

corrigendum [BP17a].

Papanicolopulos:2016:ECC

[Pap16a]  Stefanos-Aldo  Papani-

colopulos. Efficient computation of cubature rules with application to new asymmetric rules on the triangle. Journal of Computational and Applied Mathematics, 304(??):73– 83, October 1, 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716301339.

Papanicolopulos:2016:NFS

[Pap16b]  Stefanos-Aldo  Papani-

colopulos. New fully symmetric and rotationally symmetric cubature rules on the triangle using minimal orthonormal bases. Journal of Computational and Applied Mathematics, 294(??):39–48, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004082.

Park:2015:SAE

[Par15]  Moongyu Park.  Sta-

bility analysis and error estimate of flowfielddependent variation (FDV) method for first order linear hyperbolic equations. Journal of Computational and Applied Mathematics, 285(??):125–131, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000837.

Pazos:2015:ACT

[PB15] Fernando Pazos and Amit Bhaya. Adaptive choice of the Tikhonov regularization parameter to solve ill-posed linear algebraic equations via Liapunov optimizing control. Journal of Computational and Applied Mathematics, 279(??):123–132, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004683.

Perez:2017:SSL

[PBGG17] C. Pérez, F. Benítez, and J. B. García-Gutiérrez. Stabilization of switched linear systems by using projections. Journal of Computational and Applied Mathematics, 318 (??):117–123, July 2017.

CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716305532.

Perez-Benito:2019:GBM

[PBJCM19] Cristina Pérez-Benito, Cristina Jordán, J. Alberto Conejero, and Samuel Morillas. Graph-based methods for simultaneous smoothing and sharpening of color images. Journal of Computational and Applied Mathematics, 350(??):380– 395, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306393.

Perez-Benito:2018:MBL

[PBMJC18] Cristina Pérez-Benito, Samuel Morillas, Cristina Jordán, and J. Alberto Conejero. A model based on local graphs for colour images and its application for Gaussian noise smoothing. Journal of Computational and Applied Mathematics, 330(??):955–964, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302492.

Pany:2016:OEE

[PBP16] Ambit K. Pany, Saumya Bajpai, and Amiya K. Pani. Optimal error estimates for semidiscrete Galerkin approximations to equations of motion described by Kelvin–Voigt viscoelastic fluid flow model. Journal of Computational and Applied Mathematics, 302(??):234–257, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300164.

Patvardhan:2015:SQK

[PBS15] C. Patvardhan, Sulabh Bansal, and A. Srivastav. Solving the 0–1 quadratic knapsack problem with a competitive quantum inspired evolutionary algorithm. Journal of Computational and Applied Mathematics, 285(??):86–99, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000862.

Pinto:2017:LFO

[PC17] Carla M. A. Pinto and Ana R. M. Carvalho. A latency fractional order model for HIV dynamics. Journal of Computational and Applied Mathematics, 312(??):240–256, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302539.

Pinto:2018:IPE

[PC18a] Carla M. A. Pinto and Ana R. M. Carvalho. The impact of pre-exposure prophylaxis (PrEP) and screening on the dynamics of HIV. Journal of Computational and Applied Mathematics, 339(??):231–244, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305204.

Portelenelle:2018:EQR

[PC18b] Brice Portelenelle and Jean-Pierre Croisille. An efficient quadrature rule on the Cubed Sphere. Journal of Computational and Applied Mathematics, 328(??):59–74, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303278.

Piqueras:2017:FFN

[PCJ17] M.-A. Piqueras, R. Company, and L. Jódar. A front-fixing numerical method for a free boundary nonlinear diffusion logistic population model. Journal of Computational and Applied Mathematics, 309(??):473–481, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300760.

Piqueras:2018:CPS

[PCJ18a] M. A. Piqueras, R. Company, and L. Jódar. Computing positive stable numerical solutions of moving boundary problems for concrete carbonation. Journal of Computational and Applied Mathematics, 330(??):794–805, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301176.

Piqueras:2018:NAC

[PCJ18b] M.-A. Piqueras, R. Company, and L. Jódar. Numerical analysis and computing of free boundary problems for concrete carbonation chemical corrosion. Journal of Computational and Applied Mathematics, 336(??):297– 316, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300049.

Prabha:2019:DAT

[PCSR19] T. Prabha, M. Chandru, V. Shanthi, and H. Ramos. Discrete approximation for a two-parameter singularly perturbed boundary value problem having discontinuity in convection coefficient and source term. Journal of Computational and Applied Mathematics, 359(??):102–118, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301712.

Parand:2017:AST

[PD17] Kourosh Parand and Mehdi Delkhosh. Accurate solution of the Thomas– Fermi equation using the fractional order of rational Chebyshev functions. Journal of Computational and Applied Mathematics, 317(??):624– 642, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305775.

Pak:2019:SIP

[PD19] Abbas Pak and Sanku Dey. Statistical inference for the power Lindley model based on record values and inter-record times. Journal of Computational and Applied Mathematics, 347(??):156–172, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304898.

Perez-Diaz:2015:CBB

[PDB15] Sonia Pérez-Díaz and Angel Blasco. A connection between birational automorphisms of the plane and linear systems of curves. Journal of Computational and Applied Mathematics, 283(??):91–105, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000382.

Phillips:2019:TGS

[PDO19] T. N. Phillips, J. H. Davies, and D. Oldham. Towards global SEM mantle convection simulations on polyhedral-based grids. Journal of Computational and Applied Mathematics, 348(??):48–57, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305296.

Pereyra:2016:MOR

[Per16] Victor Pereyra. Model order reduction with oblique projections for large scale wave propagation. Journal of Computational and Applied Mathematics, 295(??):103–114, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000412.

Presho:2016:MCG

[PG16] Michael Presho and Juan Galvis. A mass conservative Generalized Multiscale Finite Element Method applied to twophase flow in heterogeneous porous media. Journal of Computational and Applied Mathematics, 296 (??):376–388, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0377042715005014.

Pasricha:2019:PVP

[PG19] Puneet Pasricha and Anubha Goel. Pricing vulnerable power exchange options in an intensity based framework. Journal of Computational and Applied Mathematics, 355(??):106–115, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300287.

Pinilla:2018:CAD

[PGD^{+}18] Samuel Pinilla, Hans García, Luis Díaz, Juan Poveda, and Henry Arguello. Coded aperture design for solving the phase retrieval problem in X-ray crystallography. Journal of Computational and Applied Mathematics, 338(??):111–128, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300736.

Pedroche:2018:STL

[PGRC18a] Francisco Pedroche, Esther García, Miguel Romance, and Regino Criado. On the spectrum of two-layer approach and Multiplex PageRank. Journal of Computational and Applied Mathematics, 344(??):161–172, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303042.

Pedroche:2018:SEP

[PGRC18b] Francisco Pedroche, Esther García, Miguel Romance, and Regino Criado. Sharp estimates for the personalized multiplex PageRank. Journal of Computational and Applied Mathematics, 330(??):1030–1040, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300717.

Poveda:2016:AEH

[PHCG16] Leonardo A. Poveda, Sebastian Huepo, Victor M. Calo, and Juan Galvis. Asymptotic expansions for high-contrast linear elasticity. Journal of Computational and Applied Mathematics, 295(??):25–34, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715001740.

Pimenov:2017:CNL

[PHD17] V. G. Pimenov, A. S. Hendy, and R. H. De Staelen. On a class of nonlinear delay distributed order fractional diffusion equations. Journal of Computational and Applied Mathematics, 318 (??):433–443, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716300954.

Proinov:2019:CAS

[PI19] Petko D. Proinov and Stoil I. Ivanov. Convergence analysis of Sakurai– Torii–Sugiura iterative method for simultaneous approximation of polynomial zeros. Journal of Computational and Applied Mathematics, 357(??):56–70, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300949.

Pishbin:2015:OCR

[Pis15] S. Pishbin. Optimal convergence results of piecewise polynomial collocation solutions for integralalgebraic equations of index-3. Journal of Computational and Applied Mathematics, 279(??):209– 224, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400483X.

Pishbin:2017:OTM

[Pis17] S. Pishbin. Operational Tau method for singular system of Volterra integro-differential equations. Journal of Computational and Applied Mathematics, 311(??):205–214, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303363.

Padcharoen:2019:ALM

[PKMM19] Anantachai Padcharoen, Poom Kumam, and Juan Martínez-Moreno. Augmented Lagrangian method for TV–l_{1}–l_{2} based colour image restoration. Journal of Computational and Applied Mathematics, 354 (??):507–519, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718307039.

Petkovic:2018:RGS

[PKR18] Marko D. Petković, Mihailo A. Krstić, and Kostadin P. Rajković. Rapid generalized schultz iterative methods for the computation of outer inverses. Journal of Computational and Applied Mathematics, 344 (??):572–584, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830325X.

Plekhanova:2017:DCP

[Ple17] M. V. Plekhanova. Distributed control problems for a class of degenerate semilinear evolution equations. Journal of Computational and Applied Mathematics, 312(??):39–46, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004860.

Peng:2017:IMS

[PLP17] Jingjing Peng, Anping Liao, and Zhenyun Peng. An iteration method to solve multiple constrained least squares problems. Journal of Computational and Applied Mathematics, 322(??):129–138, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301346.

Panja:2015:GTQ

[PM15] M. M. Panja and B. N.

Mandal. Gauss-type quadrature rule with complex nodes and weights for integrals involving Daubechies scale functions and wavelets. Journal of Computational and Applied Mathematics, 290 (??):609–632, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003180.

Panigrahi:2019:LMG

[PMN19] Bijaya Laxmi Panigrahi, Moumita Mandal, and Gnaneshwar Nelakanti. Legendre multi-Galerkin methods for Fredholm integral equations with weakly singular kernel and the corresponding eigenvalue problem. Journal of Computational and Applied Mathematics, 346(??):224–236, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304308.

Petkovic:2016:ASC

[PN16] Ivan Petković and Beny Neta. On an application of symbolic computation and computer graphics to root-finders: the case of multiple roots of unknown multiplicity. Journal of Computational and Applied Mathematics, 308(??):215–230, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302783.

Pollock:2016:SIA

[Pol16]  Sara Pollock.  Stabi-

lized and inexact adaptive methods for capturing internal layers in quasilinear PDE. Journal of Computational and Applied Mathematics, 308 (??):243–262, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302813.

Petkovic:2015:HPM

[PP15] Marko D. Petković and Miodrag S. Petković.

Hyper-power methods for the computation of outer inverses. Journal of Computational and Applied Mathematics, 278(??):110– 118, April 15, 2015. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714004312.

Pitolli:2016:NSL

[PP16] F. Pitolli and C. Pocci.

Neuroelectric source localization by random spatial sampling. Journal of Computational and Applied Mathematics, 296(??):237– 246, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500480X.

Petkovic:2018:NDR

[PP18a] M. S. Petković and L. D.

Petković. A note on determinantal representation of a Schröder–König-like simultaneous method for finding polynomial zeros. Journal of Computational and Applied Mathematics, 343(??):481–487, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718302784.

Petkovic:2018:TGF

[PP18b] M. S. Petković and L. D. Petković. Traub–Gander’s family for the simultaneous determination of multiple zeros of polynomials. Journal of Computational and Applied Mathematics, 327(??):362–371, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303424.

Petkovic:2019:CEM

[PP19] Miodrag S. Petković and Ljiljana D. Petković. Construction and efficiency of multipoint root-ratio methods for finding multiple zeros. Journal of Computational and Applied Mathematics, 351(??):54– 65, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306502.

Paul:2016:MAS

[PPM16] Swaraj Paul, M. M. Panja, and B. N. Mandal. Multiscale approximation of the solution of weakly singular second kind Fredholm integral equation in Legendre multiwavelet basis. Journal of Computational and Applied Mathematics, 300(??):275– 289, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006342.

Pranic:2015:GAG

[PR15] Miroslav S. Pranić and Lothar Reichel. Generalized anti-Gauss quadrature rules. Journal of Computational and Applied Mathematics, 284 (??):235–243, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042714004956.

Paluszny:2016:OGM

[PR16] Marco Paluszny and Gustavo Restrepo. Orthogonal grids on meander-like regions: a database approach. Journal of Computational and Applied Mathematics, 295(??):62–69, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001041.

Patricio:2019:SIB

[PRP19] M. F. Simões Patrício, Higinio Ramos, and Miguel Patrício. Solving initial and boundary value problems of fractional ordinary differential equations by using collocation and fractional powers. Journal of Computational and Applied Mathematics, 354(??):348– 359, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304606.

Park:2018:PDT

[PRRY18] Y. Park, L. Reichel, G. Rodriguez, and X. Yu. Parameter determination for Tikhonov regularization problems in general form. Journal of Computational and Applied Mathematics, 343(??):12–25, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302449.

Przybylowicz:2015:MAE

[Prz15] Pawel Przybylowicz. Minimal asymptotic error for one-point approximation of SDEs with timeirregular coefficients. Journal of Computational and Applied Mathematics, 282 (??):98–110, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715000126.

Przybylowicz:2019:EAS

[Prz19] Pawel Przybylowicz. Efficient approximate solution of jump-diffusion SDEs via path-dependent adaptive step-size control. Journal of Computational and Applied Mathematics, 350(??):396–411, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306587.

Panopoulos:2015:ESS

[PS15] G. A. Panopoulos and T. E. Simos. An eight-step semi-embedded predictor– corrector method for orbital problems and related IVPs with oscillatory solutions for which the frequency is unknown. Journal of Computational and Applied Mathematics, 290(??):1–15, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715002642.

Palitta:2018:CEP

[PS18] Davide Palitta and Valeria Simoncini. Computationally enhanced projection methods for symmetric Sylvester and Lyapunov matrix equations. Journal of Computational and Applied Mathematics, 330(??):648–659, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717304004.

Pandey:2016:CMG

[PSK16] Rajesh K. Pandey, Shiva Sharma, and Kamlesh Kumar. Collocation method for generalized Abel’s integral equations. Journal of Computational and Applied Mathematics, 302(??):118–128, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300152.

Patel:2017:TDW

[PSS17] Vijay Kumar Patel, Somveer Singh, and Vineet Kumar Singh. Two-dimensional wavelets collocation method for electromagnetic waves in dielectric media. Journal of Computational and Applied Mathematics, 317 (??):307–330, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716305593.

Popuzin:2018:FNA

[PST18] V. V. Popuzin, M. A.

Sumbatyan, and A. E. Tarasov. A fast numerical algorithm for a basic dual integral equation of the flapping wing in a flow of non-viscous incompressible fluid. Journal of Computational and Applied Mathematics, 344 (??):457–472, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303315.

Peng:2018:ABA

[PSW18] Jie Peng, Shi Shu, and Junxian Wang. An adaptive BDDC algorithm in variational form for mortar discretizations. Journal of Computational and Applied Mathematics, 335(??):185– 206, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305915.

Pektas:2017:HFI

[PT17a] Burhan Pektas and Ege Tamci. The heat flux identification problem for a nonlinear parabolic equation in 2D. Journal of Computational and Applied Mathematics, 312 (??):134–142, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716300218.

Psarrakos:2017:GCR

[PT17b] Georgios Psarrakos and Abdolsaeed Toomaj. On the generalized cumulative residual entropy with applications in actuarial science. Journal of Computational and Applied Mathematics, 309(??):186–199, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303090.

Pedas:2017:STS

[PTV17] Arvet Pedas, Enn Tamme, and Mikk Vikerpuur. Smoothing transformation and spline collocation for nonlinear fractional initial and boundary value problems. Journal of Computational and Applied Mathematics, 317(??):1– 16, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305556.

Pulch:2017:HNQ

[Pul17] Roland Pulch. A Hankel norm for quadrature rules solving random linear dynamical systems. Journal of Computational and Applied Mathematics, 316(??):307–318, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305258.

Puccetti:2015:DCJ

[PW15a] Giovanni Puccetti and Ruodu Wang. Detecting complete and joint mixability. Journal of Computational and Applied Mathematics, 280(??):174–187, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005299.

Purnomo:2015:SGO

[PW15b] Hindriyanto Dwi Purnomo and Hui-Ming Wee. Soccer game optimization with substitute players. Journal of Computational and Applied Mathematics, 283(??):79–90, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000205.

Pao:2016:NMF

[PW16] C. V. Pao and Yuan-Ming Wang. Numerical methods for fourth-order elliptic equations with nonlocal boundary conditions. Journal of Computational and Applied Mathematics, 292(??):447–468, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003805.

Plonka:2017:SFF

[PW17] Gerlind Plonka and Katrin Wannenwetsch. A sparse fast Fourier algorithm for real non-negative vectors. Journal of Computational and Applied Mathematics, 321(??):532–539, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301383.

Wu:2019:NAA

[pWmGkY^{+}19] Jian ping Wu, Pei ming Guo, Fu kang Yin, Jun Peng, and Jin hui Yang. A new aggregation algorithm based on coordinates partitioning recursively for algebraic multigrid method. Journal of Computational and Applied Mathematics, 345: 184–195, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Pan:2017:OAL

[PX17] Jian Pan and Qingxian Xiao. Optimal assetliability management with liquidity constraints and stochastic interest rates in the expected utility framework. Journal of Computational and Applied Mathematics, 317(??):371– 387, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305799.

Perman:2018:OHS

[PZ18] Mihael Perman and Ana Zalokar. Optimal hedging strategies in equitylinked products. Journal of Computational and Applied Mathematics, 344 (??):601–607, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303212.

Qiu:2018:IPM

[QC18] Songqiang Qiu and Zhongwen Chen. An interior point method for nonlinear optimization with a quasi-tangential subproblem. Journal of Computational and Applied Mathematics, 334(??):77–96, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305617.

Qi:2018:SPC

[QCSG18] Feng Qi, Viera Cernanová, Xiao-Ting Shi, and Bai-Ni Guo. Some properties of central Delannoy numbers. Journal of Computational and Applied Mathematics, 328(??):101–115, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303540.

Qu:2018:IBE

[QDBG18] X. Y. Qu, C. Y. Dong, Y. Bai, and Y. P. Gong. Isogeometric boundary element method for calculating effective property of steady state thermal conduction in 2D heterogeneities with a homogeneous interphase. Journal of Computational and Applied Mathematics, 343(??):124–138, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302486.

Quarati:2015:SSE

[QDG^{+}15] Alfonso Quarati, Emanuele Danovaro, Antonella Galizia, Andrea Clematis, Daniele D’Agostino, and Antonio Parodi. Scheduling strategies for enabling meteorological simulation on hybrid clouds. Journal of Computational and Applied Mathematics, 273(??):438–451, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002349.

Qarariyah:2019:IAI

[QDY^{+}19] Ammar Qarariyah, Fang Deng, Tianhui Yang, Yuan Liu, and Jiansong Deng. Isogeometric analysis on implicit domains using weighted extended PHTsplines. Journal of Computational and Applied Mathematics, 350(??):353–371, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306198.

Qi:2017:RGM

[QG17] Feng Qi and Bai-Ni Guo. The reciprocal of the geometric mean of many positive numbers is a Stieltjes transform. Journal of Computational and Applied Mathematics, 311(??):165–170, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303181.

Qi:2018:IRM

[Qi18] Feng Qi. Integral representations for multivariate logarithmic polynomials. Journal of Computational and Applied Mathematics, 336(??):54– 62, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306465.

Qi:2019:DIR

[Qi19]  Feng Qi.  A double

inequality for the ratio of two non-zero neighbouring Bernoulli numbers. Journal of Computational and Applied Mathematics, 351(??):1–5, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306575.

Qi:2018:IRB

[QL18] Feng Qi and Dongkyu Lim.

Integral representations of bivariate complex geometric mean and their applications. Journal of Computational and Applied Mathematics, 330(??):41–58, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303941.

Qi:2019:CMD

[QL19] Feng Qi and Ai-Qi Liu.

Completely monotonic degrees for a difference between the logarithmic and psi functions. Journal of Computational and Applied Mathematics, 361(??):366–371, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302298.

Qin:2017:MTT

[QLT^{+}17] Shanlin Qin, Fawang Liu, Ian Turner, Viktor Vegh, Qiang Yu, and Qianqian Yang. Multi-term timefractional Bloch equations and application in magnetic resonance imaging. Journal of Computational and Applied Mathematics, 319(??):308–319, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300390.

Qi:2019:MTS

[QLZ19] Xiaohong Qi, Qi Li, and Fukang Zhu. Modeling time series of count with excess zeros and ones based on INAR(1) model with zero-and-one inflated Poisson innovations. Journal of Computational and Applied Mathematics, 346(??):572–590, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042718304692.

Qiu:2015:TLD

[QM15] Hailong Qiu and Liquan Mei. Two-level defect-correction stabilized finite element method for Navier–Stokes equations with friction boundary conditions. Journal of Computational and Applied Mathematics, 280(??):80–93, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400524X.

Quinn:2015:NMN

[Qui15] J. Quinn. A numerical method for a nonlinear singularly perturbed interior layer problem using an approximate layer location. Journal of Computational and Applied Mathematics, 290(??):500–515, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003313.

Qian:2015:SQA

[QWS15] Xiaoyan Qian, Li Wang, and Yongzhong Song. A successive quadratic approximations method for nonlinear eigenvalue problems. Journal of Computational and Applied Mathematics, 290 (??):268–277, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003052.

Qiu:2016:DCS

[QWZ16] Weifeng Qiu, Minglei Wang, and Jiahao Zhang. Direct computation of stresses in linear elasticity. Journal of Computational and Applied Mathematics, 292(??):363–368, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003544.

Qian:2018:SDI

[QWZ18] Jiang Qian, Fan Wang, and Chungang Zhu. Scattered data interpolation based upon bivariate recursive polynomials. Journal of Computational and Applied Mathematics, 329(??):223–243, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042717302868.

Raqab:2019:MLP

[RABB19] Mohammad Z. Raqab, Laila A. Alkhalfan, Omar M. Bdair, and Narayanaswamy Balakrishnan. Maximum likelihood prediction of records from 3parameter Weibull distribution and some approximations. Journal of Computational and Applied Mathematics, 356(??):118–132, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300706.

Rad:2017:OIR

[Rad17] Helia Niroomand Rad.

Operator index reduction in electromagnetism. Journal of Computational and Applied Mathematics, 316(??):298–306, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305222.

Rang:2015:ITR

[Ran15] Joachim Rang. Improved traditional Rosenbrockwanner methods for stiff ODEs and DAEs. Journal of Computational and Applied Mathematics, 286(??):128–144, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001569.

Re:2015:EEV

[RANG15] B. Re, R. Armellin, N. R. Nannan, and A. Guardone. Efficient evaluation of vapourliquid equilibria from multi-parameter thermodynamic models using differential algebra. Journal of Computational and Applied Mathematics, 273(??):404–413, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002386.

Ratz:2015:TTI

[Rät15] Andreas Rätz. Turingtype instabilities in bulksurface reaction–diffusion systems. Journal of Computational and Applied Mathematics, 289(??):142– 152, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0377042715001272.

Rao:2019:PUC

[RC19] S. Chandra Sekhara Rao and Sheetal Chawla. Parameteruniform convergence of a numerical method for a coupled system of singularly perturbed semilinear reaction–diffusion equations with boundary and interior layers. Journal of Computational and Applied Mathematics, 352(??):223–239, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307118.

Rosa:2017:CPS

[RCBG17] M. Rosa, J. C. Camacho, M. S. Bruzón, and M. L. Gandarias. Classical and potential symmetries for a generalized Fisher equation. Journal of Computational and Applied Mathematics, 318(??):181– 188, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305179.

Rosa:2019:CLS

[RCBG19] M. Rosa, J. C. Camacho, M. S. Bruzón, and M. L. Gandarias. Conservation laws, symmetries, and exact solutions of the classical Burgers– Fisher equation in two dimensions. Journal of Computational and Applied Mathematics, 354(??):545– 550, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306733.

Ren:2017:SAG

[RCN17] Zhi-Ru Ren, Yang Cao, and Qiang Niu. Spectral analysis of the generalized shift-splitting preconditioned saddle point problem. Journal of Computational and Applied Mathematics, 311(??):539–550, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304009.

Rivero-Castillo:2015:EDB

[RCPA15] Daniel Rivero-Castillo, Héctor Pijeira, and Pedro Assunçao. Edge detection based on Krawtchouk polynomials. Journal of Computational and Applied Mathematics, 284 (??):244–250, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042714004968.

Ralescu:2015:CAB

[RDRM15] Anca Ralescu, Irene Díaz, and Luis J. Rodríguez-Muñiz. A classification algorithm based on geometric and statistical information. Journal of Computational and Applied Mathematics, 275(??):335–344, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003318.

Reis:2016:NSG

[RdSLL16] Ruy Freitas Reis, Felipe dos Santos Loureiro, and Marcelo Lobosco. 3D numerical simulations on GPUs of hyperthermia with nanoparticles by a nonlinear bioheat model. Journal of Computational and Applied Mathematics, 295(??):35–47, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001247.

Rathie:2017:EAE

[RdSMO17] P. N. Rathie and L. C.

de S. M. Ozelim. Exact and approximate expressions for the reliability of stable Lévy random variables with applications to stock market modelling. Journal of Computational and Applied Mathematics, 321(??):314–322, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301085. See

corrigendum [RdSMO18].

Rathie:2018:CEA

[RdSMO18] P. N. Rathie and L. C. de S. M. Ozelim. Corrigendum to “Exact and approximate expressions for the reliability of stable Levy random variables with applications to stock market modelling” [J. Comput. Appl. Math. 321 (2017) 314–322]. Journal of Computational and Applied Mathematics, 343(??):771–773, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302371. See

[RdSMO17].

Ranirina:2019:HVS

[RdV19] Dinna Ranirina and Johan de Villiers. On Hermite vector splines and multi-wavelets. Journal of Computational and Applied Mathematics, 349(??):366–378, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304849.

Rashedi:2016:SRK

[REBF16] Somaiyeh Rashedi, Ghodrat Ebadi, Sebastian Birk, and Andreas Frommer. On short recurrence Krylov type methods for linear systems with many right-hand sides. Journal of Computational and Applied Mathematics, 300(??):18– 29, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005944.

Reutskiy:2016:BSM

[Reu16a] S. Yu. Reutskiy. The backward substitution method for multipoint problems with linear Volterra– Fredholm integro-differential equations of the neutral type. Journal of Computational and Applied Mathematics, 296(??):724–738, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005130.

Reutskiy:2016:NMS

[Reu16b]  S. Yu. Reutskiy.  A

novel method for solving second order fractional eigenvalue problems. Journal of Computational and Applied Mathematics, 306(??):133–153, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301674.

Reutskiy:2017:NNM

[Reu17]  S. Yu. Reutskiy.  A

new numerical method for solving high-order fractional eigenvalue problems. Journal of Computational and Applied Mathematics, 317(??):603– 623, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306379.

Roeva:2018:CDM

[RF18] Olympia Roeva and Stefka Fidanova. Comparison of different metaheuristic algorithms based on InterCriteria analysis.

Journal of Computational and Applied Mathematics, 340(??):615–628, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303795.

Royo:2016:RPM

[RFLM16] Beatriz Royo, Alberto Fraile, Emilio Larrodé, and Victoria Muerza. Route planning for a mixed delivery system in long distance transportation and comparison with pure delivery systems. Journal of Computational and Applied Mathematics, 291(??):488–496, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001958.

Romanovski:2017:BCP

[RFO17] Valery G. Romanovski, Wilker Fernandes, and Regilene Oliveira. Bicenter problem for some classes of Z_{2}-equivariant systems. Journal of Computational and Applied Mathematics, 320:61–75, August 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Rahmani:2016:ICS

[RIK16] Rabi-Allah Rahmani, Muhyiddin Izadi, and Baha-Eldin Khaledi. Importance of components in kout-of-n system with components having random weights. Journal of Computational and Applied Mathematics, 296(??):1– 9, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004616.

Rahpeymaii:2016:LMQ

[RKB16] Farzad Rahpeymaii, Morteza Kimiaei, and Alireza Bagheri. A limited memory quasi-Newton trustregion method for box constrained optimization. Journal of Computational and Applied Mathematics, 303(??):105–118, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300735.

Ritter:2016:PMM

[RKS16] Juliane Ritter, Axel Klar, and Florian Schneider. Partial-moment minimumentropy models for kinetic chemotaxis equations in one and two dimensions. Journal of Computational and Applied Mathematics, 306(??):300–315, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301923.

Reshniak:2019:SSS

[RKV19] Viktor Reshniak, Abdul Khaliq, and David Voss. Slow-scale split-step tau-leap method for stiff stochastic chemical systems. Journal of Computational and Applied Mathematics, 361(??):79–96, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S037704271930175X.

Rudoy:2018:DDT

[RL18] E. M. Rudoy and N. P.

Lazarev. Domain decomposition technique for a model of an elastic body reinforced by a Timoshenko’s beam. Journal of Computational and Applied Mathematics, 334(??):18–26, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042717305794.

Ramos:2017:NAB

[RM17] Higinio Ramos and M. T. T. Monteiro. A new approach based on the Newton’s method to solve systems of nonlinear equations. Journal of Computational and Applied Mathematics, 318(??):3– 13, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271630629X.

Ruiz:2019:IPS

[RM19] A. Ruiz and C. Muriel. Involutive pairs of λsymmetries for n th-order ordinary differential equations. Journal of Computational and Applied Mathematics, 354(??):562– 568, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306630.

Roul:2019:FCR

[RMA19] Pradip Roul, Harshita Madduri, and Ravi Agarwal. A fast-converging recursive approach for Lane–Emden type initial value problems arising in astrophysics. Journal of Computational and Applied Mathematics, 359(??):182–195, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301621.

Roul:2019:NIA

[RMK19] Pradip Roul, Harshita Madduri, and Klaus Kassner. A new iterative algorithm for a strongly nonlinear singular boundary value problem. Journal of Computational and Applied Mathematics, 351(??):167–178, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306514.

Roldan:2015:MFP

[RMMRC15] A. Roldán, J. Martínez-Moreno, C. Roldán, and Y. J. Cho. Multidimensional fixed point theorems under (ψ, ϕ)-contractive conditions in partially ordered complete metric spaces. Journal of Computational and Applied Mathematics, 273(??):76–87, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002672.

Ramos:2017:UAD

[RMVA17] Higinio Ramos, Shubham Mehta, and J. Vigo-Aguiar. A unified approach for the development of k-step block Falkner-type methods for solving general secondorder initial-value problems in ODEs. Journal of Computational and Applied Mathematics, 318 (??):550–564, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715006305.

Roozegar:2017:NCP

[RN17] Rasool Roozegar and Saralees Nadarajah. New classes of power series bivariate copulas. Journal of Computational and Applied Mathematics, 326 (??):235–246, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302650.

Ramirez-Nafarrate:2016:PEO

[RNM16] Adrian Ramirez-Nafarrate and David F. Muñoz. Performance evaluation of output analysis methods in steady-state simulations. Journal of Computational and Applied Mathematics, 301(??):64–73, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630019X.

Radu:2015:RLS

[RNPK15] Florin Adrian Radu, Jan Martin Nordbotten, Iuliu Sorin Pop, and Kundan Kumar. A robust linearization scheme for finite volume based discretizations for simulation of two-phase flow in porous media. Journal of Computational and Applied Mathematics, 289(??):134–141, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001387.

Rahimkhani:2017:NSF

[ROB17] P. Rahimkhani, Y. Ordokhani, and E. Babolian. Numerical solution of fractional pantograph differential equations by using generalized fractionalorder Bernoulli wavelet. Journal of Computational and Applied Mathematics, 309(??):493–510, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302758.

Romagnoli:2018:MIC

[Rom18] Silvia Romagnoli. Measureinvariance of copula functions as tool for testing no-arbitrage assumption. Journal of Computational and Applied Mathematics, 338(??):80–90, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300724.

Roshan:2017:ESC

[Ros17] Jamal Rezaei Roshan. Existence of solutions for a class of system of functional integral equation via measure of noncompactness. Journal of Computational and Applied Mathematics, 313(??):129–141, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304319.

Ross:2019:OCT

[Ros19] I. M. Ross. An optimal control theory for nonlinear optimization. Journal of Computational and Applied Mathematics, 354 (??):39–51, July 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042719300019.

Roth:2015:CPB

[Rót15]  Ágoston Róth.  Con-

trol point based exact description of trigonometric/hyperbolic curves, surfaces and volumes. Journal of Computational and Applied Mathematics, 290(??):74–91, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002769.

Ren:2015:GMF

[RP15] Jun Ren and Michael Presho. A Generalized Multiscale Finite Element Method for high-contrast single-phase flow problems in anisotropic media. Journal of Computational and Applied Mathematics, 277(??):202–214, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042714004026.

Rambeerich:2016:HOF

[RP16a] Nisha Rambeerich and Athanasios A. Pantelous. A high order finite element scheme for pricing options under regime switching jump diffusion processes. Journal of Computational and Applied Mathematics, 300(??):83– 96, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715006317.

Rancic:2016:NSR

[RP16b] Lidija Z. Rancić and Miodrag S. Petković. New simultaneous root-finding methods with accelerated convergence for analytic functions. Journal of Computational and Applied Mathematics, 296(??):228– 236, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004823.

Ramsden:2017:ARM

[RP17] Lewis Ramsden and Apostolos D. Papaioannou. Asymptotic results for a Markov-modulated risk process with stochastic investment. Journal of Computational and Applied Mathematics, 313(??):38–53, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304307.

Rincon:2016:NAS

[RQ16] M. A. Rincon and N. P.

Quintino. Numerical analysis and simulation for a nonlinear wave equation. Journal of Computational and Applied Mathematics, 296(??):247–264, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004768.

Ruiz-Ramirez:2016:PNA

[RR16] Javier Ruiz-Ramírez. Preserving nonnegativity of an affine finite element approximation for a convection– diffusion–reaction problem. Journal of Computational and Applied Mathematics, 306(??):286–299, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301911.

Rukavishnikov:2018:WFE

[RR18] Viktor A. Rukavishnikov and Alexey V. Rukavishnikov. Weighted finite element method for the Stokes problem with corner singularity. Journal of Computational and Applied Mathematics, 341(??):144–156, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301821.

Ramirez:2018:TNR

[RRM18] J. Ramírez, J. L. Romero, and C. Muriel. Two new reductions methods for polynomial differential equations and applications to nonlinear PDEs. Journal of Computational and Applied Mathematics, 333(??):36– 50, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305265.

Ramirez:2019:NMO

[RRM19] J. Ramírez, J. L. Romero, and C. Muriel. A new method to obtain either first- or secondorder reductions for parametric polynomial ODEs.

Journal of Computational and Applied Mathematics, 358(??):146–162, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301256.

Romani:2019:EDM

[RRS19] Lucia Romani, Milvia Rossini, and Daniela Schenone. Edge detection methods based on RBF interpolation. Journal of Computational and Applied Mathematics, 349(??):532–547, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304837.

Rasmussen:2017:ARS

[RRW17] Martin Rasmussen, Janosch Rieger, and Kevin N. Webster. Approximation of reachable sets using optimal control and support vector machines. Journal of Computational and Applied Mathematics, 311(??):68–83, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716302850.

Ruiz:2016:DFT

[RS16] Elias D. Nino Ruiz and Adrian Sandu. A derivativefree trust region framework for variational data assimilation. Journal of Computational and Applied Mathematics, 293 (??):164–179, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S037704271500103X.

Ravibabu:2018:NVA

[RS18] Mashetti Ravibabu and Arindama Singh. A new variant of Arnoldi method for approximation of eigenpairs. Journal of Computational and Applied Mathematics, 344 (??):424–437, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303248.

Reif:2019:OPN

[RS19] Ulrich Reif and Malcolm A. Sabin. Old problems and new challenges in subdivision. Journal of Computational and Applied Mathematics, 349(??):523–531, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303649.

Ruiz:2016:HPC

[RSID16] P. Ruiz, J. Sastre, J. Ibáñez, and E. Defez. High performance computing of the matrix exponential. Journal of Computational and Applied Mathematics, 291(??):370–379, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002101.

Roul:2018:NHO

[RT18a] Pradip Roul and Kiran Thula. A new high-order numerical method for solving singular two-point boundary value problems. Journal of Computational and Applied Mathematics, 343(??):556–574, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302516.

Rzadkowski:2018:FOP

[RT18b] Grzegorz Rzadkowski and Emran Tohidi. A fourth order product integration rule by using the generalized Euler–Maclaurin summation formula. Journal of Computational and Applied Mathematics, 335 (??):334–348, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717306374.

Rauwoens:2015:GMS

[RTV15] Pieter Rauwoens, Peter Troch, and Jan Vierendeels. A geometric multigrid solver for the freesurface equation in environmental models featuring irregular coastlines. Journal of Computational and Applied Mathematics, 289(??):22–36, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500182X.

Rujivan:2016:CFF

[Ruj16] Sanae Rujivan. A closedform formula for the conditional moments of the extended CIR process. Journal of Computational and Applied Mathematics, 297(??):75–84, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005452.

Ramos:2015:ANM

[RVA15] Higinio Ramos and J. Vigo-Aguiar. The application of Newton’s method in vector form for solving nonlinear scalar equations where the classical Newton method fails. Journal of Computational and Applied Mathematics, 275(??):228–237, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003483.

Reddy:2019:EAB

[RVC19] G. M. M. Reddy, M. Vynnycky, and J. A. Cuminato. An efficient adaptive boundary algorithm to reconstruct Neumann boundary data in the MFS for the inverse Stefan problem. Journal of Computational and Applied Mathematics, 349(??):21–40, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305491.

Raducan:2015:RPN

[RVZ15] Anisoara Maria Raducan, Raluca Vernic, and Gheorghita Zbaganu. On the ruin probability for nonhomogeneous claims and arbitrary inter-claim revenues. Journal of Computational and Applied Mathematics, 290 (??):319–333, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003155.

Richter:2015:VLD

[RW15] Thomas Richter and Thomas Wick. Variational localizations of the dual weighted residual estimator. Journal of Computational and Applied Mathematics, 279(??):192–208, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004798.

Roul:2016:NNA

[RW16] Pradip Roul and Ujwal Warbhe. A novel numerical approach and its convergence for numerical solution of nonlinear doubly singular boundary value problems. Journal of Computational and Applied Mathematics, 296(??):661– 676, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005208.

Ran:2017:PIH

[RWW17] Yu-Hong Ran, Jun-Gang Wang, and Dong-Ling Wang. On partially inexact HSS iteration methods for the complex symmetric linear systems in space fractional CNLS equations. Journal of Computational and Applied Mathematics, 317(??):128– 136, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305726.

Ran:2019:LCN

[RZ19a] Maohua Ran and Chengjian Zhang. Linearized Crank– Nicolson scheme for the nonlinear time-space fractional Schrödinger equations. Journal of Computational and Applied Mathematics, 355(??):218–231, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300615.

Ruan:2019:NCF

[RZ19b] Xinfeng Ruan and Wenjun Zhang. A note on “A closed-form pricing formula for European options under the Heston model with stochastic interest rate”. Journal of Computational and Applied Mathematics, 350(??):55– 56, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306034. See

[HZ18b].

Schutz:2017:HSS

[SA17] Jochen Schütz and Vadym Aizinger. A hierarchical scale separation approach for the hybridized discontinuous Galerkin method. Journal of Computational and Applied Mathematics, 317(??):500– 509, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306288.

Sahihi:2018:RKM

[SAA18] Hussein Sahihi, Saeid Abbasbandy, and Tofigh Allahviranloo. Reproducing kernel method for solving singularly perturbed differential-difference equations with boundary layer behavior in Hilbert space. Journal of Computational and Applied Mathematics, 328(??):30–43, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303308.

Sadri:2017:LCN

[SAC17] K. Sadri, A. Amini, and C. Cheng. Low cost numerical solution for three-dimensional linear and nonlinear integral equations via threedimensional Jacobi polynomials. Journal of Computational and Applied Mathematics, 319(??):493– 513, August 1, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717300511.

Sadkane:2019:AMA

[Sad19] Miloud Sadkane. On alternating maximization algorithm for computing the hump of matrix powers. Journal of Computational and Applied Mathematics, 353(??):154– 163, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307635.

Sharifi:2015:FGI

[SAH15] M. Sharifi, M. Arab, and F. Khaksar Haghani. Finding generalized inverses by a fast and efficient numerical method. Journal of Computational and Applied Mathematics, 279 (??):187–191, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042714004804.

Soltani:2017:TIR

[SAH17] Sara Soltani, Martin S. Andersen, and Per Christian Hansen. Tomographic image reconstruction using training images. Journal of Computational and Applied Mathematics, 313(??):243–258, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716304393.

Said:2016:IGG

[Sai16] Samia M. Said. Influence of gravity on generalized magneto-thermoelastic medium for three-phase-lag model. Journal of Computational and Applied Mathematics, 291(??):142–157, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400572X.

Saiedinezhad:2019:MNH

[Sai19] S. Saiedinezhad. On a measure of noncompactness in the Holder space C^{k,γ}(Ω) and its application. Journal of Computational and Applied Mathematics, 346(??):566–571, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304503.

Sakar:2017:IRK

[Sak17] Mehmet Giyas Sakar. Iterative reproducing kernel Hilbert spaces method for Riccati differential equations. Journal of Computational and Applied Mathematics, 309(??):163–174, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303016.

Saray:2019:EAS

[Sar19] Behzad Nemati Saray.

An efficient algorithm for solving Volterra integrodifferential equations based on Alpert’s multi-wavelets Galerkin method. Journal of Computational and Applied Mathematics, 348 (??):453–465, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042718305612.

Sokolov:2015:ASI

[SAT15] Andriy Sokolov, Ramzan Ali, and Stefan Turek. An AFC-stabilized implicit finite element method for partial differential equations on evolvingin-time surfaces. Journal of Computational and Applied Mathematics, 289(??):101–115, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001478.

Solis:2016:DMD

[SB16] Francisco J. Solis and Ignacio Barradas. Discrete multiple delay advectionreaction operators. Journal of Computational and Applied Mathematics, 291(??):441–448, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500206X.

Sazmand:2018:AJS

[SB18] Afsane Sazmand and Mahmoud Behroozifar. Application Jacobi spectral method for solving the time-fractional differential equation. Journal of Computational and Applied Mathematics, 339(??):49– 68, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830089X.

Salam:2019:TBN

[SB19] Ahmed Salam and Haithem Ben Kahla. A treatment of breakdowns and near breakdowns in a reduction of a matrix to upper J-Hessenberg form and related topics. Journal of Computational and Applied Mathematics, 359(??):16–29, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S037704271930130X.

Silva:2016:NCF

[SBC16] Rodrigo B. Silva, Marcelo Bourguignon, and Gauss M. Cordeiro. A new compounding family of distributions: the generalized gamma power series distributions. Journal of Computational and Applied Mathematics, 303 (??):119–139, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300802.

Sun:2018:PCN

[SBCJ18] Tao Sun, Roberto Barrio, Lizhi Cheng, and Hao Jiang. Precompact convergence of the nonconvex primal-dual hybrid gradient algorithm. Journal of Computational and Applied Mathematics, 330(??):15–27, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303886.

Singh:2017:NFC

[SBP17] Heisnam Rohen Singh, Saroj Kr. Biswas, and Biswajit Purkayastha. A neuro-fuzzy classification technique using dynamic clustering and GSS rule generation. Journal of Computational and Applied Mathematics, 309 (??):683–694, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042716301960.

Salete:2017:SPM

[SBU^{+}17] E. Salete, J. J. Benito, F. Ureña, L. Gavete, M. Ureña, and A. García. Stability of perfectly matched layer regions in generalized finite difference method for wave problems. Journal of Computational and Applied Mathematics, 312(??):231–239, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302618.

Song:2018:SME

[SC18] Caiqin Song and Guoliang Chen. Solutions to matrix equations X − AXB = CY + R and X − AX̂B = CY + R. Journal of Computational and Applied Mathematics, 343(??):488–500, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302735.

Sousa:2018:BOP

[SCG18] Rúben Sousa, Ana Bela Cruzeiro, and Manuel Guerra. Barrier option pricing under the 2hypergeometric stochastic volatility model. Journal of Computational and Applied Mathematics, 328(??):197–213, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303643.

Scholz:2018:SMD

[Sch18] Lena Scholz. The signature method for DAEs arising in the modeling of electrical circuits. Journal of Computational and Applied Mathematics, 332(??):107–139, April 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305022.

Sanz:2018:SDS

[SCMS18] María T. Sanz, Antonio Caselles, Joan C. Micó, and David Soler. A stochastic dynamical social model involving a human happiness index. Journal of Computational and Applied Mathematics, 340(??):231–246, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301158.

Solis:2019:CTG

[SCP19] Cesar U. Solis, Julio B.

Clempner, and Alexander S. Poznyak. Continuoustime gradient-like descent algorithm for constrained convex unknown functions: Penalty method application. Journal of Computational and Applied Mathematics, 355(??):268– 282, August 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

https://www.sciencedirect.
com/science/article/
pii/S0377042719300391.

Sellier:2015:SSW

[SD15] J. M. Sellier and I. Dimov. A sensitivity study of the Wigner Monte Carlo method. Journal of Computational and Applied Mathematics, 277(??):87–93, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004063.

Slodicka:2015:RTA

[SDG15] Marián Slodicka, Rob H. De Staelen, and Christophe Geuzaine. Recent trends in Advanced Computational Methods in Engineering (ACOMEN). Journal of Computational and Applied Mathematics, 289(??):1–2, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002344.

Sedaghatjoo:2017:UNS

[SDH17] Zeynab Sedaghatjoo, Mehdi Dehghan, and Hossein Hosseinzadeh. On uniqueness of numerical solution of boundary integral equations with 3times monotone radial kernels. Journal of Computational and Applied Mathematics, 311(??):664–681, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303910.

Sadatrasoul:2016:NST

[SE16] S. M. Sadatrasoul and R. Ezzati. Numerical solution of two-dimensional nonlinear Hammerstein fuzzy integral equations based on optimal fuzzy quadrature formula. Journal of Computational and Applied Mathematics, 292(??):430–446, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003854.

Salman:2018:BMM

[SE18] S. M. Salman and A. A.

Elsadany. On the bifurcation of Marotto’s map and its application in image encryption. Journal of Computational and Applied Mathematics, 328(??):177–196, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303515.

Seydaoglu:2016:NSB

[SEÖ16] M. Seydaoğlu, U. Erdogan, and T. Özis. Numerical solution of Burgers’ equation with high order splitting methods. Journal of Computational and Applied Mathematics, 291(??):410–421, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715002393.

Sarmiento:2018:ESG

[SEV^{+}18] A. F. Sarmiento, L. F. R. Espath, P. Vignal, L. Dalcin, M. Parsani, and V. M. Calo. An energy-stable generalized-α method for the Swift–Hohenberg equation. Journal of Computational and Applied Mathematics, 344(??):836–851, December 15, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717305642.

Seydaoglu:2018:AAA

[Sey18]  Muaz Seydaoglu.  An

accurate approximation algorithm for Burgers’ equation in the presence of small viscosity. Journal of Computational and Applied Mathematics, 344(??):473–481, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718303467.

Sun:2016:DGF

[SF16] Tong Sun and Adamou Fode. A discontinuous Galerkin-front tracking scheme and its optimaloptimal error estimation.

Journal of Computational and Applied Mathematics, 292(??):369–386, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003866.

Sabatier:2018:CDI

[SF18] Jocelyn Sabatier and Christophe Farges. Comments on the description and initialization of fractional partial differential equations using Riemann– Liouville’s and Caputo’s definitions. Journal of Computational and Applied Mathematics, 339 (??):30–39, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718301092.

Solano-Feo:2016:NMS

[SFGJR^{+}16] F. Solano-Feo, J. M.

Guevara-Jordan, O. Rojas, B. Otero, and R. Rodriguez. A new mimetic scheme for the acoustic wave equation. Journal of Computational and Applied Mathematics, 295(??):2–12, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005051.

Sebestyen:2016:SPC

[SFH^{+}16] G. Svantnerné Sebestyén, István Faragó, Róbert Horváth, R. Kersner, and M. Klincsik. Stability of patterns and of constant steady states for a cross-diffusion system. Journal of Computational and Applied Mathematics, 293(??):208–216, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001946.

Singh:2019:HWC

[SGG19] Randhir Singh, Himanshu Garg, and Vandana Guleria. Haar wavelet collocation method for Lane– Emden equations with Dirichlet, Neumann and Neumann–Robin boundary conditions. Journal of Computational and Applied Mathematics, 346(??):150–161, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304242.

Sarakhsi:2016:NHA

[SGK16] M. Khojaste Sarakhsi, S. M. T. Fatemi Ghomi, and B. Karimi. A new hybrid algorithm of scatter search and Nelder– Mead algorithms to optimize joint economic lot sizing problem. Journal of Computational and Applied Mathematics, 292(??):387–401, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003891.

Schroder:2017:STA

[SGL17] Dirk Schröder, Alf Gerisch, and Jens Lang. Spacetime adaptive linearly implicit peer methods for parabolic problems. Journal of Computational and Applied Mathematics, 316(??):330–344, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303922.

Svacek:2015:MMF

[SH15] Petr Svácek and Jaromír Horácek. On mathematical modeling of fluidstructure interactions with nonlinear effects: Finite element approximations of gust response. Journal of Computational and Applied Mathematics, 273(??):394–403, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002404.

Sugiura:2018:GCS

[SH18] Hiroshi Sugiura and Takemitsu Hasegawa. On the global convergence of Schröder’s iterative formulae for real roots of algebraic equations. Journal of Computational and Applied Mathematics, 344 (??):313–322, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303182.

Sugiura:2019:GCS

[SH19] Hiroshi Sugiura and Takemitsu Hasegawa. On the global convergence of Schröder’s iteration formula for real zeros of entire functions. Journal of Computational and Applied Mathematics, 358(??):136–145, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301177.

Shaikhet:2015:SES

[Sha15] Leonid Shaikhet. Stability of equilibrium states for a stochastically perturbed exponential type system of difference equations. Journal of Computational and Applied Mathematics, 290(??):92–103, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002757.

Shen:2019:DLR

[SHC^{+}19] Zhao-Li Shen, Ting-Zhu Huang, Bruno Carpentieri, Chun Wen, Xian-Ming Gu, and Xue-Yuan Tan. Off-diagonal lowrank preconditioner for difficult PageRank problems. Journal of Computational and Applied Mathematics, 346(??):456–470, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304357.

Sun:2019:FDV

[SHCJ19] Dong-Lin Sun, Ting-Zhu Huang, Bruno Carpentieri, and Yan-Fei Jing. Flexible and deflated variants of the block shifted GM-RES method. Journal of Computational and Applied Mathematics, 345: 168–183, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Shokrollahi:2018:EGA

[Sho18] Foad Shokrollahi. The evaluation of geometric Asian power options under time changed mixed fractional Brownian motion. Journal of Computational and Applied Mathematics, 344(??):716–724, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303194.

Salimi:2015:FPR

[SHS^{+}15] P. Salimi, N. Hussain, S. Shukla, Sh. Fathollahi, and S. Radenović. Fixed point results for cyclic α − ψφ-contractions with application to integral equations. Journal of Computational and Applied Mathematics, 290 (??):445–458, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042715003118.

Shen:2019:RDL

[SHW19] Peiping Shen, Bingdi Huang, and Lufan Wang. Range division and linearization algorithm for a class of linear ratios optimization problems. Journal of Computational and Applied Mathematics, 350(??):324–342, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306460.

Szekeres:2016:FEA

[SI16] Béla J. Szekeres and Ferenc Izsák. Finite element approximation of fractional order elliptic boundary value problems. Journal of Computational and Applied Mathematics, 292(??):553–561, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500388X.

Sastre:2019:FTP

[SIAJ^{+}19] Jorge Sastre, Javier Ibáñez, Pedro Alonso-Jordá, Jesús Peinado, and Emilio Defez. Fast Taylor polynomial evaluation for the computation of the matrix cosine. Journal of Computational and Applied Mathematics, 354(??):641– 650, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307726.

Sir:2015:CTA

[SJ15] Zbynek Sír and Bert Jüttler. On de casteljautype algorithms for rational Bézier curves. Journal of Computational and Applied Mathematics, 288(??):244–250, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000497.

Shivanian:2017:ISM

[SJ17] Elyas Shivanian and Ahmad Jafarabadi. An improved spectral meshless radial point interpolation for a class of timedependent fractional integral equations: 2D fractional evolution equation. Journal of Computational and Applied Mathematics, 325(??):18–33, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302030.

Shivanian:2018:ASM

[SJ18] Elyas Shivanian and Ahmad Jafarabadi. Analysis of the spectral meshless radial point interpolation for solving fractional reactionsubdiffusion equation. Journal of Computational and Applied Mathematics, 336 (??):98–113, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717306453.

Song:2016:ROM

[SJL16a] Huailing Song, Lijian Jiang, and Qiuqi Li. A reduced order method for Allen–Cahn equations. Journal of Computational and Applied Mathematics, 292(??):213–229, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003714.

Song:2016:RES

[SJL16b] Yunquan Song, Ling Jian, and Lu Lin. Robust exponential squared loss-based variable selection for highdimensional single-index varying-coefficient model.

Journal of Computational and Applied Mathematics, 308(??):330–345, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302643.

Salvador:2017:UHE

[SJRR17] F. J. Salvador, D. Jaramillo, J.-V. Romero, and M.- D. Roselló. Using a homogeneous equilibrium model for the study of the inner nozzle flow and cavitation pattern in convergent-divergent nozzles of diesel injectors. Journal of Computational and Applied Mathematics, 309(??):630–641, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301832.

Smarda:2015:ECA

[SK15] Zdenek Smarda and Yasir Khan. An efficient computational approach to solving singular initial value problems for Lane– Emden type equations. Journal of Computational and Applied Mathematics, 290(??):65–73, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500271X.

Singh:2016:HWM

[SK16] Inderdeep Singh and Sheo Kumar. Haar wavelet method for some nonlinear Volterra integral equations of the first kind. Journal of Computational and Applied Mathematics, 292(??):541–552, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003842.

Suchde:2018:PCM

[SK18] Pratik Suchde and Jörg Kuhnert. Point cloud movement for fully Lagrangian meshfree methods. Journal of Computational and Applied Mathematics, 340(??):89–100, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300992.

Schroder:2015:MFE

[SKB15] Andreas Schröder, Heiko Kleemann, and Heribert Blum. Mixed finite element methods for twobody contact problems. Journal of Computational and Applied Mathematics, 283(??):58–70, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005196.

Singh:2016:IEF

[SKBT16] Mehakpreet Singh, Jitendra Kumar, Andreas Bück, and Evangelos Tsotsas. An improved and efficient finite volume scheme for bivariate aggregation population balance equation. Journal of Computational and Applied Mathematics, 308(??):83–97, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302175.

Saboor:2019:SNR

[SKC^{+}19] Abdus Saboor, Muhammad Nauman Khan, Gauss M. Cordeiro, Marcelino A. R. Pascoa, Pedro L. Ramos, and Mustafa Kamal. Some new results for the transmuted generalized gamma distribution. Journal of Computational and Applied Mathematics, 352 (??):165–180, May 15, 2019. CODEN JCAMDI.

ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042718307271.

Skrzipek:2017:SRD

[Skr17] M. R. Skrzipek. Signal recovery by discrete approximation and a Pronylike method. Journal of Computational and Applied Mathematics, 326 (??):193–203, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302741.

Su:2018:MSQ

[SL18] Chunmei Su and Zhiping Li. A meshing strategy for a quadratic iso-parametric FEM in cavitation computation in nonlinear elasticity. Journal of Computational and Applied Mathematics, 330(??):630–647, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304235.

Shen:2019:ASN

[SLA19] Shujun Shen, Fawang Liu, and Vo V. Anh. The analytical solution and numerical solutions for a two-dimensional multiterm time fractional diffusion and diffusion-wave equation. Journal of Computational and Applied Mathematics, 345: 515–534, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Smith:2015:INB

[SLH15] B. Smith, R. Laoulache, and A. Heryudono. Implementation of Neumann boundary condition with influence matrix method for viscous annular flow using pseudospectral collocation. Journal of Computational and Applied Mathematics, 285(??):100–115, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000734.

Saffarzadeh:2018:ITN

[SLH18] M. Saffarzadeh, G. B. Loghmani, and M. Heydari. An iterative technique for the numerical solution of nonlinear stochastic Itô– Volterra integral equations. Journal of Computational and Applied Mathematics, 333(??):74– 86, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304661.

Slodicka:2015:DST

[Slo15] M. Slodicka. Determination of a solely timedependent source in a semilinear parabolic problem by means of boundary measurements. Journal of Computational and Applied Mathematics, 289(??):433–440, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004439.

Sun:2016:FRG

[SLQ16] Zhongfeng Sun, Aijuan Li, and Huizeng Qin. Further results on generalized multiple fractional part integrals for complex values. Journal of Computational and Applied Mathematics, 302(??):186–199, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300619.

Shi:2018:DGM

[SLS18] Cengke Shi, Jichun Li, and Chi-Wang Shu. Discontinuous Galerkin methods for Maxwell’s equations in Drude metamaterials on unstructured meshes. Journal of Computational and Applied Mathematics, 342(??):147–163, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301791.

Sun:2017:EUP

[SLW17] Yan Sun, Lishan Liu, and Yonghong Wu. The existence and uniqueness of positive monotone solutions for a class of nonlinear Schrödinger equations on infinite domains. Journal of Computational and Applied Mathematics, 321(??):478–486, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301012.

Speleers:2015:ODP

[SM15a] Hendrik Speleers and Carla Manni. Optimizing domain parameterization in isogeometric analysis based on Powell– Sabin splines. Journal of Computational and Applied Mathematics, 289(??):68–86, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001776.

Suarez:2015:LPI

[SM15b] José P. Suárez and Tania Moreno. The limit property for the interior solid angles of some refinement schemes for simplicial meshes. Journal of Computational and Applied Mathematics, 275(??):135–138, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003616.

Stoykov:2016:SPI

[SM16] S. Stoykov and S. Margenov.

Scalable parallel implementation of shooting method for large-scale dynamical systems. Application to bridge components. Journal of Computational and Applied Mathematics, 293(??):223–231, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002332.

Stoykov:2017:NMP

[SM17] S. Stoykov and S. Margenov. Numerical methods and parallel algorithms for computation of periodic responses of plates. Journal of Computational and Applied Mathematics, 310(??):200–212, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303120.

Sobhani:2018:NMP

[SM18] Amirhossein Sobhani and Mariyan Milev. A numerical method for pricing discrete double barrier option by Legendre multiwavelet. Journal of Computational and Applied Mathematics, 328(??):355–364, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303849.

Sato:2019:SDE

[SM19a] S. Sato and T. Matsuo. On spatial discretization of evolutionary differential equations on the periodic domain with a mixed derivative. Journal of Computational and Applied Mathematics, 358(??):221–240, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301463.

Silva:2019:BMC

[SM19b] Ivair R. Silva and Reinaldo Marques. Bayesian Monte Carlo testing with onedimensional measures of evidence. Journal of Computational and Applied Mathematics, 351(??):250– 259, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306964.

Stefanescu:2018:PDD

[SMS18] Razvan Stefanescu, Azam Moosavi, and Adrian Sandu. Parametric domain decomposition for accurate reduced order models: Applications of MP-LROM methodology. Journal of Computational and Applied Mathematics, 340(??):629–644, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305782.

Silva:2018:FBM

[SMTdF18] Ivair R. Silva, Edgard M. Maboudou-Tchao, and Weslei L. de Figueiredo. Frequentist-Bayesian Monte Carlo test for mean vectors in high dimension. Journal of Computational and Applied Mathematics, 333(??):51–64, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730523X.

Sulman:2019:PPA

[SN19] M. Sulman and T. Nguyen. A positivity preserving adaptive moving mesh method for cancer cell invasion models. Journal of Computational and Applied Mathematics, 361(??):487–501, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719302456.

Sheng:2019:LCA

[SNY19] Zhou Sheng, Qin Ni, and Gonglin Yuan. Local convergence analysis of inverse iteration algorithm for computing the h-spectral radius of a nonnegative weakly irreducible tensor. Journal of Computational and Applied Mathematics, 357(??):26–37, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300871.

Suzuki:2016:GFD

[SO16] Yukihito Suzuki and Masashi Ohnawa. GENERIC formalism and discrete variational derivative method for the two-dimensional vorticity equation. Journal of Computational and Applied Mathematics, 296 (??):690–708, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S037704271500518X.

Sanchez-Orgaz:2019:NMM

[SODBK19] E. M. Sánchez-Orgaz, F. D. Denia, L. Baeza, and R. Kirby. Numerical mode matching for sound propagation in silencers with granular material. Journal of Computational and Applied Mathematics, 350(??):233–246, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718306381.

Sato:2019:RNI

[SOMF19] S. Sato, K. Oguma, T. Matsuo, and B.-F. Feng. A robust numerical integrator for the short pulse equation near criticality. Journal of Computational and Applied Mathematics, 361(??):343–365, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301827.

Song:2016:ESI

[Son16]  Huailing Song.  En-

ergy SSP–IMEX Runge– Kutta methods for the Cahn–Hilliard equation. Journal of Computational and Applied Mathematics, 292(??):576–590, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003921.

Soll:2017:SSB

[SP17] Tino Soll and Roland Pulch. Sample selection based on sensitivity analysis in parameterized model order reduction. Journal of Computational and Applied Mathematics, 316(??):369–379, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304757.

Shakeri:2016:GNM

[SPCSA16] Heman Shakeri, Pietro Poggi-Corradini, Caterina Scoglio, and Nathan Albin. Generalized network measures based on modulus of families of walks. Journal of Computational and Applied Mathematics, 307(??):307–318, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300061.

Shen:2015:NPR

[SPD15] Li-Yong Shen and Sonia Pérez-Díaz. Numerical proper reparametrization of parametric plane curves. Journal of Computational and Applied Mathematics, 277(??):138–161, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004087.

SanJose:2016:IRG

[SPG^{+}16a] Roberto San José, Juan L. Pérez, Rosa M. González, Julia Pecci, Antonio Garzón, and Marino Palacios. Impacts of the 4.5 and 8.5 RCP global climate scenarios on urban meteorology and air quality: Application to Madrid, Antwerp, Milan, Helsinki and London. Journal of Computational and Applied Mathematics, 293(??):192–207, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002423.

Syrakov:2016:SEA

[SPG^{+}16b] Dimiter Syrakov, Maria Prodanova, Emilia Georgieva, Iglika Etropolska, and Kiril Slavov. Simulation of European air quality by WRF–CMAQ models using AQMEII-2 infrastructure. Journal of Computational and Applied Mathematics, 293(??):232–245, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000448.

Sharma:2018:CMC

[SPK18] Shiva Sharma, Rajesh K.

Pandey, and Kamlesh Kumar. Collocation method with convergence for generalized fractional integro-differential equations. Journal of Computational and Applied Mathematics, 342(??):419–430, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830219X.

Sviercoski:2015:ACG

[SPM15] R. F. Sviercoski, P. Popov, and S. Margenov. An analytical coarse grid operator applied to a multiscale multigrid method. Journal of Computational and Applied Mathematics, 287(??):207–219, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001466.

Sicilia:2016:OAS

[SQRE16] Juan Antonio Sicilia, Carlos Quemada, Beatriz Royo, and David Escuín. An optimization algorithm for solving the rich vehicle routing problem based on variable neighborhood search and tabu search metaheuristics. Journal of Computational and Applied Mathematics, 291(??):468–477, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002034.

Sanchez-Reyes:2015:DSP

[SR15] J. Sánchez-Reyes. Detecting symmetries in polynomial Bézier curves. Journal of Computational and Applied Mathematics, 288(??):274–283, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002435.

Sahu:2016:CEN

[SR16a] P. K. Sahu and S. Saha Ray. Comparative experiment on the numerical solutions of Hammerstein integral equation arising from chemical phenomenon. Journal of Computational and Applied Mathematics, 291(??):402–409, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042715001508.

Sahu:2016:LSC

[SR16b] P. K. Sahu and S. Saha Ray. Legendre spectral collocation method for the solution of the model describing biological species living together. Journal of Computational and Applied Mathematics, 296(??):47– 55, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500463X.

Selvi:2016:INM

[SR16c] P. Avudai Selvi and N. Ramanujam. An iterative numerical method for singularly perturbed reaction– diffusion equations with negative shift. Journal of Computational and Applied Mathematics, 296 (??):10–23, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715004550.

Suciu:2015:FPA

[SRA^{+}15] N. Suciu, F. A. Radu, S. Attinger, L. Schüler, and P. Knabner. A Fokker–Planck approach for probability distributions of species concentrations transported in heterogeneous media. Journal of Computational and Applied Mathematics, 289(??):241–252, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000424.

Singh:2018:ASG

[SRBD18] Abhinoy Kumar Singh, Rahul Radhakrishnan, Shovan Bhaumik, and Paresh Date. Adaptive sparse-grid Gauss– Hermite filter. Journal of Computational and Applied Mathematics, 342(??):305–316, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301742.

Soni:2017:AEP

[SRH17] V. Soni, O. Roussel, and A. Hadjadj. On the accuracy and efficiency of point-value multiresolution algorithms for solving scalar wave and Euler equations. Journal of Computational and Applied Mathematics, 323(??):159–175, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301656.

Salvador:2016:NSP

[SRRJ16] F. J. Salvador, J.-V.

Romero, M.-D. Roselló, and D. Jaramillo. Numerical simulation of primary atomization in diesel spray at low injection pressure. Journal of Computational and Applied Mathematics, 291(??):94–102, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001971.

Sadkane:2016:ECS

[SS16a] Miloud Sadkane and Roger B.

Sidje. Efficient computation of the spectral projections of regular matrix pairs. Journal of Computational and Applied Mathematics, 298(??):72– 81, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005890.

Sendra:2016:SCD

[SS16b] J. Rafael Sendra and Juana Sendra. Symbolic computation of Drazin inverses by specializations. Journal of Computational and Applied Mathematics, 301(??):201–212, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300395.

Settanni:2016:DEN

[SS16c] Giuseppina Settanni and Ivonne Sgura. Devising efficient numerical methods for oscillating patterns in reaction–diffusion systems. Journal of Computational and Applied Mathematics, 292(??):674–693, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002708.

Slodicka:2017:IMK

[SS17] M. Slodicka and L. Seliga. Identification of memory kernels in hyperbolic problems. Journal of Computational and Applied Mathematics, 311(??):618–629, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042716303752.

Safarzadeh:2018:DIN

[SS18a] Malihe Safarzadeh and Abbas Salemi. DGM-RES and index numerical range of matrices. Journal of Computational and Applied Mathematics, 335 (??):349–360, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717306295.

Salandin:2018:CMC

[SS18b] Andrea Salandin and David Soler. Computing the minimum construction cost of a building’s external wall taking into account its energy efficiency. Journal of Computational and Applied Mathematics, 338(??):199–211, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300748.

Sharma:2018:NSA

[SS18c] Vaishali Sharma and Amit Setia. Numerical solution and its analysis for a system of hypersingular integral equations. Journal of Computational and Applied Mathematics, 343(??):520–538, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302474.

Solis:2018:BMC

[SS18d] Francisco J. Solis and Fernando Saldaña. Biological mechanisms of coexistence for a family of age structured population models. Journal of Computational and Applied Mathematics, 343(??):708–718, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302681.

Sarmavuori:2019:NIF

[SS19a] Juha Sarmavuori and Simo Särkkä. Numerical integration as a finite matrix approximation to multiplication operator. Journal of Computational and Applied Mathematics, 353(??):283– 291, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307623.

Siskova:2019:ISF

[SS19b] K. Sisková and M. Slodicka. Identification of a source in a fractional wave equation from a boundary measurement. Journal of Computational and Applied Mathematics, 349(??):172–186, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830565X.

Soler:2018:SDM

[SSCM18] David Soler, María T.

Sanz, Antonio Caselles, and Joan C. Micó. A stochastic dynamic model to evaluate the influence of economy and well-being on unemployment control. Journal of Computational and Applied Mathematics, 330(??):1063–1080, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302042.

Sellier:2016:WMC

[SSD16] J. M. Sellier, R. F. Sviercoski, and I. Dimov. On the Wigner Monte Carlo method coupled to pseudopotential models. Journal of Computational and Applied Mathematics, 293(??):217–222, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500045X.

Sonntag:2016:SDC

[SSG16] Matthias Sonntag, Stephan Schmidt, and Nicolas R. Gauger. Shape derivatives for the compressible Navier–Stokes equations in variational form. Journal of Computational and Applied Mathematics, 296 (??):334–351, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715004628.

Stanimirovic:2018:ZNN

[SSG18] Predrag S. Stanimirović, Shwetabh Srivastava, and D. K. Gupta. From Zhang neural network to scaled hyperpower iterations. Journal of Computational and Applied Mathematics, 331(??):133–155, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304892.

Stanimirovic:2016:COI

[SSH16] Predrag S. Stanimirović, F. Soleymani, and F. Khaksar Haghani. Computing outer inverses by scaled matrix iterations. Journal of Computational and Applied Mathematics, 296(??):89–101, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004653.

Sun:2017:MRD

[SSKV17] Xianming Sun, Thorsten Schulz, Asma Khedher, and Michèle Vanmaele. Model risk and discretisation of locally riskminimising strategies. Journal of Computational and Applied Mathematics, 311(??):38–53, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303211.

Sbibih:2015:STQ

[SST15] D. Sbibih, A. Serghini, and A. Tijini. Superconvergent trivariate quadratic spline quasi-interpolants on Worsey–Piper split. Journal of Computational and Applied Mathematics, 276(??):117–128, March 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042714003811.

Shao:2019:ECM

[SSW19] Wenting Shao, Shi Sun, and Yingwei Wang. An economical cascadic multigrid method for the weak Galerkin finite element approximation of second order elliptic problems. Journal of Computational and Applied Mathematics, 362(??):341–353, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718306253.

Shen:2019:HPR

[SSZ19] Jiguang Shen, John R. Singler, and Yangwen Zhang. HDG–POD reduced order model of the heat equation. Journal of Computational and Applied Mathematics, 362 (??):663–679, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S037704271830582X.

Sun:2016:EAF

[SSZL16] Yuzhou Sun, Pengtao Sun, Bin Zheng, and Guang Lin. Error analysis of finite element method for Poisson–Nernst–Planck equations. Journal of Computational and Applied Mathematics, 301 (??):28–43, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042716300073.

Saad:2016:GLP

[ST16a] Nasser Saad and K. Thirulogasanthar. Generalized 2D Laguerre polynomials and their quaternionic extensions. Journal of Computational and Applied Mathematics, 308 (??):301–317, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302862.

Shiraya:2016:AFB

[ST16b] Kenichiro Shiraya and Akihiko Takahashi. An approximation formula for basket option prices under local stochastic volatility with jumps: an application to commodity markets. Journal of Computational and Applied Mathematics, 292(??):230–256, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003507.

Stamatiou:2018:BPN

[Sta18] I. S. Stamatiou. A boundary preserving numerical scheme for the Wright– Fisher model. Journal of Computational and Applied Mathematics, 328(??):132–150, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303527.

Stamatiou:2019:EPP

[Sta19]  I. S. Stamatiou.  An

explicit positivity preserving numerical scheme for CIR/CEV type delay models with jump. Journal of Computational and Applied Mathematics, 360(??):78–98, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301852.

Stoykov:2018:BAG

[Sto18]  S. Stoykov.  Buckling

analysis of geometrically nonlinear curved beams. Journal of Computational and Applied Mathematics, 340(??):653–663, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730417X.

Soheili:2015:FCN

[STS15] Ali R. Soheili, F. Toutounian, and F. Soleymani. A fast convergent numerical method for matrix sign function with application in SDEs. Journal of Computational and Applied Mathematics, 282(??):167– 178, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000060.

Sacerdote:2016:FPT

[STZ16] Laura Sacerdote, Massimiliano Tamborrino, and Cristina Zucca. First passage times of twodimensional correlated processes: Analytical results for the Wiener process and a numerical method for diffusion processes. Journal of Computational and Applied Mathematics, 296(??):275–292, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715004859.

Su:2016:ISP

[Su16] Qifang Su. Inverse spectral problem for pseudo-Jacobi matrices with partial spectral data. Journal of Computational and Applied Mathematics, 297(??):1– 12, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005567.

Su:2018:CNM

[Su18] Chunmei Su. Comparison of numerical methods for the Zakharov system in the subsonic limit regime. Journal of Computational and Applied Mathematics, 330(??):441–455, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730420X.

Sun:2016:MPF

[Sun16] Zhongyang Sun. Maximum principle for forwardbackward stochastic control system under Gexpectation and relation to dynamic programming. Journal of Computational and Applied Mathematics, 296(??):753–775, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005440.

Sun:2019:FDF

[Sun19]  Pengtao Sun.  Ficti-

tious domain finite element method for Stokes/ elliptic interface problems with jump coefficients. Journal of Computational and Applied Mathematics, 356(??):81–97, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300469.

Siddiqi:2016:HFT

[SuSR16] Shahid S. Siddiqi, Wardat us Salam, and Kashif Rehan. Hyperbolic forms of ternary non-stationary subdivision schemes originated from hyperbolic Bsplines. Journal of Computational and Applied Mathematics, 301(??):16–27, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716000054.

Steyer:2016:SSS

[SV16a] Andrew J. Steyer and Erik S. Van Vleck. A step-size selection strategy for explicit Runge– Kutta methods based on Lyapunov exponent theory. Journal of Computational and Applied Mathematics, 292(??):703–719, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002095.

Stykel:2016:TSM

[SV16b] Tatjana Stykel and Alexander Vasilyev. A twostep model reduction approach for mechanical systems with moving loads. Journal of Computational and Applied Mathematics, 297(??):85– 97, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005592.

Schenkels:2019:RFB

[SV19] Nick Schenkels and Wim Vanroose. Regula falsi based automatic regularization method for PDE constrained optimization. Journal of Computational and Applied Mathematics, 348(??):14– 25, March 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718305375.

Shokri:2019:NCT

[SVAKGR19a] Ali Shokri, Jesús Vigo-Aguiar, Mohammad Mehdizadeh Khalsaraei, and Raquel Garcia-Rubio. A new class of two-step p-stable TFPL methods for the numerical solution of second-order IVPs with oscillating solutions. Journal of Computational and Applied Mathematics, 354(??):551– 561, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301936.

Shokri:2019:NFS

[SVAKGR19b] Ali Shokri, Jesús Vigo-Aguiar, Mohammad Mehdizadeh Khalsaraei, and Raquel Garcia-Rubio. A new fourstep P -stable Obrechkoff method with vanished phase-lag and some of its derivatives for the numerical solution of radial Schrödinger equation. Journal of Computational and Applied Mathematics, 354(??):569– 586, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302048.

Sangers:2015:ECS

[SvG15] Alex Sangers and Martin B. van Gijzen. The eigenvectors corresponding to the second eigenvalue of the Google matrix and their relation to link spamming. Journal of Computational and Applied Mathematics, 277(??):192–201, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004105.

SanJuan:2017:UDN

[SVGM17] P. San Juan, A. M. Vidal, and V. M. Garcia-Molla. Updating/downdating the NonNegative matrix factorization. Journal of Computational and Applied Mathematics, 318 (??):59–68, July 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/
pii/S0377042716305908.

Spiridonov:2019:GMF

[SVL19] Denis Spiridonov, Maria Vasilyeva, and Wing Tat Leung. A Generalized Multiscale Finite Element Method (GMsFEM) for perforated domain flows with Robin boundary conditions. Journal of Computational and Applied Mathematics, 357 (??):319–328, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301268.

Stepanov:2018:GMD

[SVV18] Sergei Stepanov, Maria Vasilyeva, and Vasiliy I. Vasil’ev. Generalized multiscale discontinuous Galerkin method for solving the heat problem with phase change. Journal of Computational and Applied Mathematics, 340(??):645–652, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306180.

Sereeter:2019:CNR

[SVW19] Baljinnyam Sereeter, Cornelis Vuik, and Cees Witteveen. On a comparison of Newton–Raphson solvers for power flow problems. Journal of Computational and Applied Mathematics, 360(??):157–169, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301876.

Sendra:2016:CHE

[SW16a] J. Rafael Sendra and Stephan M. Winkler. Corrigendum to: A heuristic and evolutionary algorithm to optimize the coefficients of curve parametrizations [J. Comput. Appl. Math. 305 (2016) 18–35]. Journal of Computational and Applied Mathematics, 308(??):499–500, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302011. See

[SW16b].

Sendra:2016:HEA

[SW16b] J. Rafael Sendra and Stephan M. Winkler. A heuristic and evolutionary algorithm to optimize the coefficients of curve parametrizations. Journal of Computational and Applied Mathematics, 305(??):18–35, October 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301406. See

[SW16a].

Schmitt:2017:ESP

[SW17a] Bernhard A. Schmitt and Rüdiger Weiner. Efficient A-stable peer twostep methods. Journal of Computational and Applied Mathematics, 316(??):319–329, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304149.

Soleimani:2017:CIP

[SW17b] Behnam Soleimani and Rüdiger Weiner. A class of implicit peer methods for stiff systems. Journal of Computational and Applied Mathematics, 316(??):358–368, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302849.

Sun:2017:SSM

[SW17c] Liying Sun and Lijin Wang. Stochastic symplectic methods based on the Padé approximations for linear stochastic Hamiltonian systems. Journal of Computational and Applied Mathematics, 311(??):439–456, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303715.

Sun:2017:OEA

[SW17d] Weiwei Sun and Jilu Wang. Optimal error analysis of Crank–Nicolson schemes for a coupled nonlinear Schrödinger system in 3D. Journal of Computational and Applied Mathematics, 317(??):685– 699, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305970.

Sheng:2016:OCM

[SWH16] Qiwei Sheng, Cheng Wang, and Weimin Han. An optimal cascadic multigrid method for the radiative transfer equation. Journal of Computational and Applied Mathematics, 303(??):189–205, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301029.

Shen:2015:MRS

[SWL15] Dongmei Shen, Musheng Wei, and Yonghui Liu. Minimum rank (skew) Hermitian solutions to the matrix approximation problem in the spectral norm. Journal of Computational and Applied Mathematics, 288(??):351–365, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002526.

Shi:2015:NPP

[SWZW15] Cong Shi, Chen Wang, Guanghui Zheng, and Ting Wei. A new a posteriori parameter choice strategy for the convolution regularization of the space-fractional backward diffusion problem. Journal of Computational and Applied Mathematics, 279(??):233–248, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042714004841.

Sun:2018:HMC

[SX18] Yongchao Sun and Chenglong Xu. A hybrid Monte Carlo acceleration method of pricing basket options based on splitting. Journal of Computational and Applied Mathematics, 342(??):292–304, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302267.

Sahin:2018:EPT

[SY18] Murat Sahin and Semih Yilmaz. The eigenvalues of r-periodic tridiagonal matrices by factorization of some recursive sequences. Journal of Computational and Applied Mathematics, 335(??):99– 113, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305940.

Shen:2019:SSN

[SY19] Yunqiu Shen and Tjalling J. Ypma. Solving separable nonlinear least squares problems using the QR factorization. Journal of Computational and Applied Mathematics, 345: 48–58, January 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Solodushkin:2015:FOP

[SYD15] Svyatoslav I. Solodushkin, Irina F. Yumanova, and Rob H. De Staelen. First order partial differential equations with time delay and retardation of a state variable. Journal of Computational and Applied Mathematics, 289(??):322–330, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005883.

Solodushkin:2017:DSM

[SYD17] Svyatoslav I. Solodushkin, Irina F. Yumanova, and Rob H. De Staelen. A difference scheme for multidimensional transfer equations with time delay. Journal of Computational and Applied Mathematics, 318(??):580– 590, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006238.

Shen:2016:DEC

[SYT16] Wanqiang Shen, Ping Yin, and Chengjie Tan. Degree elevation of changeable degree spline. Journal of Computational and Applied Mathematics, 300 (??):56–67, July 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715005841.

Sun:2019:ITD

[SYW19] Liangliang Sun, Xiongbin Yan, and Ting Wei. Identification of time-dependent convection coefficient in a time-fractional diffusion equation. Journal of Computational and Applied Mathematics, 346(??):505–517, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304497.

Shen:2019:CQO

[SYZ19] Yue Shen, Ningning Yan, and Zhaojie Zhou. Convergence and quasi-optimality of an adaptive finite element method for elliptic Robin boundary control problem. Journal of Computational and Applied Mathematics, 356(??):1–21, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300548.

Spiteri:2016:OSB

[SZ16] Raymond J. Spiteri and Saeed Torabi Ziaratgahi. Operator splitting for the bidomain model revisited. Journal of Computational and Applied Mathematics, 296(??):550–563, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004677.

Song:2018:SIP

[SZ18] Lunji Song and Shan Zhao. Symmetric interior penalty Galerkin approaches for two-dimensional parabolic interface problems with low regularity solutions. Journal of Computational and Applied Mathematics, 330(??):356– 379, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717304491.

Song:2019:IRC

[SZJ19] Xiaoyan Song, Guang-Hui Zheng, and Lijian Jiang. Identification of the reaction coefficient in time fractional diffusion equations. Journal of Computational and Applied Mathematics, 345:295–309, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Su:2018:RRS

[SZM^{+}18] Zhuo Su, Biyi Zeng, Jiaxin Miao, Xiaonan Luo, Baocai Yin, and Qiang Chen. Relative reductive structure-aware regression filter. Journal of Computational and Applied Mathematics, 329(??):244–255, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303011.

Shi:2016:STV

[SZZ^{+}16] Lin Shi, Hong Zhu, Shouming Zhong, Yong Zeng, and Jun Cheng. Synchronization for timevarying complex networks based on control. Journal of Computational and Applied Mathematics, 301(??):178–187, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000285.

Takhedmit:2017:PUA

[TA17] Baya Takhedmit and Karim Abbas. A parametric uncertainty analysis method for queues with vacations. Journal of Computational and Applied Mathematics, 312(??):143– 155, March 1, 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716300784.

Tan:2017:NEM

[Tan17] Xueyuan Tan. A new extrapolation method for PageRank computations. Journal of Computational and Applied Mathematics, 313(??):383–392, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304034.

Tan:2018:ASC

[Tan18]  Li Tan.  Almost sure

convergence rate of θ-EM scheme for neutral SD-DEs. Journal of Computational and Applied Mathematics, 342(??):25–36, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301754.

Tavares:2018:CFV

[TAT18] Dina Tavares, Ricardo Almeida, and Delfim F. M. Torres. Combined fractional variational problems of variable order and some computational aspects. Journal of Computational and Applied Mathematics, 339(??):374–388, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302133.

Thakoor:2018:FQM

[TB18] Deeveya Thakoor and Muddun Bhuruth. Fast quadrature methods for options with discrete dividends. Journal of Computational and Applied Mathematics, 330(??):1–14, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303953.

Torregrosa:2016:NLQ

[TBAH16] A. J. Torregrosa, A. Broatch, F. J. Arnau, and M. Hernández. A non-linear quasi-3D model with Flux-Corrected-Transport for engine gasexchange modelling. Journal of Computational and Applied Mathematics, 291(??):103–111, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001879.

Temimi:2016:IFD

[TBR16] H. Temimi and M. Ben-Romdhane. An iterative finite difference method for solving Bratu’s problem. Journal of Computational and Applied Mathematics, 292(??):76–82, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003453.

Torrente:2018:RNB

[TBS18] M. Torrente, M. C. Beltrametti, and J. R. Sendra. r-norm bounds and metric properties for zero loci of real analytic functions. Journal of Computational and Applied Mathematics, 336(??):375– 393, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300220.

Touil:2017:FPD

[TBY17] Imene Touil, Djamel Benterki, and Adnan Yassine. A feasible primal-dual interior point method for linear semidefinite programming. Journal of Computational and Applied Mathematics, 312(??):216–230, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716302254.

Thompson:2019:MRM

[TC19] Tony Thompson and Ke Chen. A more robust multigrid algorithm for diffusion type registration models. Journal of Computational and Applied Mathematics, 361(??):502–527, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301864.

Tang:2018:PAO

[TDA18] H. S. Tang, W. B. Dong, and A. Agrawal. A phenomenon of artificial oddeven grid oscillation and its presence in domain decomposition computation: Algebraic analysis and numerical illustration. Journal of Computational and Applied Mathematics, 333(??):404–427, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305186.

Teng:2015:PQO

[TEG15] Long Teng, Matthias Ehrhardt, and Michael Günther. The pricing of quanto options under dynamic correlation. Journal of Computational and Applied Mathematics, 275(??):304–310, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003367.

Tsai:2018:NSS

[TF18] Wan-Yu Tsai and Arash Fahim. A numerical scheme for a singular control problem: Investment-consumption under proportional transaction costs. Journal of Computational and Applied Mathematics, 333 (??):170–184, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S037704271730537X.

Tsitouras:2017:PFR

[TFS17] Ch. Tsitouras, I. Th. Famelis, and T. E. Simos. Phase-fitted Runge–Kutta pairs of orders 8(7). Journal of Computational and Applied Mathematics, 321(??):226–231, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730095X.

Tahmasebi:2019:CSC

[TGJ19] Mahdieh Tahmasebi, Azadeh Ghasemifard, and Mohammad Taghi Jahandideh. Comments on “Strong convergence rates for backward Euler on a class of nonlinear jumpdiffusion problems” [j. comput. appl. math. 205 (2007) 949–956]. Journal of Computational and Applied Mathematics, 359(??):69–72, October 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301633. See

[HK07].

Tranquilli:2017:AJV

[TGSS17] Paul Tranquilli, S. Ross Glandon, Arash Sarshar, and Adrian Sandu. Analytical Jacobian-vector products for the matrixfree time integration of partial differential equations. Journal of Computational and Applied Mathematics, 310(??):213–223, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302199.

Thong:2018:IEA

[TH18] Duong Viet Thong and Dang Van Hieu. Inertial extragradient algorithms for strongly pseudomonotone variational inequalities. Journal of Computational and Applied Mathematics, 341(??):80–98, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301651.

Tan:2019:HDF

[TH19] Wee Chin Tan and Viet Ha Hoang. High dimensional finite element method for multiscale nonlinear monotone parabolic equations. Journal of Computational and Applied Mathematics, 345:471–500, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Tong:2018:RBP

[THC^{+}18] Shanshan Tong, Bo Han, Yong Chen, Jinping Tang, Bo Bi, and Ruixue Gu. RTE-based parameter reconstruction with TV+L1 regularization. Journal of Computational and Applied Mathematics, 337(??):256–273, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300323.

Tuan:2017:RFS

[THL^{+}17] Nguyen Huy Tuan, Dinh Nguyen Duy Hai, Le Dinh Long, Van Thinh Nguyen, and Mokhtar Kirane. On a Riesz–Feller space fractional backward diffusion problem with a nonlinear source. Journal of Computational and Applied Mathematics, 312(??):103–126, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000078.

Thompson:2016:DCT

[Tho16] Travis Thompson. A discrete commutator theory for the consistency and phase error analysis of semi-discrete C^{0} finite element approximations to the linear transport equation. Journal of Computational and Applied Mathematics, 303(??):229–248, September 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630098X.

Taheri:2018:PEB

[TJA18] B. M. Taheri, H. Jabbari, and M. Amini. Parameter estimation of bivariate distributions in presence of outliers: an application to FGM copula. Journal of Computational and Applied Mathematics, 343(??):155–173, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302383.

Taheri:2017:NSS

[TJB17] Z. Taheri, S. Javadi, and E. Babolian. Numerical solution of stochastic fractional integro-differential equation by the spectral collocation method. Journal of Computational and Applied Mathematics, 321(??):336–347, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300924.

Torregrosa:2016:MMC

[TJCC16] Juan R. Torregrosa, Lucas Jódar, Juan C. Cortés, and Regino Criado. Mathematical modeling and computational methods. Journal of Computational and Applied Mathematics, 291(??):1–4, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002411.

Torregrosa:2017:MMC

[TJCC17] Juan R. Torregrosa, Lucas Jodar, Juan C. Cortés, and Regino Criado. Mathematical modeling and computational methods. Journal of Computational and Applied Mathematics, 309(??):320–324, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042716303533.

Toopchi:2019:FPP

[TJSW19] Yashar Toopchi, Mahdi Jalili, Jalil Sadati, and Jidong Wang. Fractional PI pinning synchronization of fractional complex dynamical networks. Journal of Computational and Applied Mathematics, 347(??):357–368, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830493X.

Tutuncu:2015:AFN

[TK15] G. Yazgi Tütüncü and Necla Kayaalp. An aggregated fuzzy naive Bayes data classifier. Journal of Computational and Applied Mathematics, 286(??):17–27, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000655.

Toulias:2016:FMM

[TK16] Thomas L. Toulias and Christos P. Kitsos. Fitting the Michaelis–Menten model. Journal of Computational and Applied Mathematics, 296(??):303–319, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005026.

Tiwari:2016:NSW

[TKH16] Sudarshan Tiwari, Axel Klar, and Steffen Hardt. Numerical simulation of wetting phenomena by a meshfree particle method. Journal of Computational and Applied Mathematics, 292(??):469–485, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003830.

Tuan:2017:REF

[TKMT17] Nguyen Huy Tuan, Vo Anh Khoa, Mach Nguyet Minh, and Thanh Tran. Reconstruction of the electric field of the Helmholtz equation in three dimensions. Journal of Computational and Applied Mathematics, 309(??):56–78, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302552.

Trofimov:2019:MMO

[TLE19] Vyacheslav A. Trofimov, Maria M. Loginova, and Vladimir A. Egorenkov. A mathematical model of optical bistability and the multiplicity of its solutions. Journal of Computational and Applied Mathematics, 354(??):663– 681, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830726X.

Tian:2018:MPS

[TLFC18] Zhaolu Tian, Xinxiang Li, C. M. Fan, and C. S. Chen. The method of particular solutions using trigonometric basis functions. Journal of Computational and Applied Mathematics, 335(??):20– 32, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305885.

Torres:2016:EWD

[TLL16] Carlos R. Torres, Edgar E.

Lanz, and Sergio I. Larios. Effect of wind direction and orography on flow structures at baja California Coast: a numerical approach. Journal of Computational and Applied Mathematics, 295(??):48–61, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001090.

Turian:2019:MSE

[TLLL19] Emma Turian, Kai Liu, John Lowengrub, and Shuwang Li. Morphological stability of an elastic tumor-host interface. Journal of Computational and Applied Mathematics, 362(??):410–422, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718303960.

Tian:2018:TMZ

[TLTJ18] Guo-Liang Tian, Yin Liu, Man-Lai Tang, and Xuejun Jiang. Type I multivariate zero-truncated/ adjusted Poisson distributions with applications. Journal of Computational and Applied Mathematics, 344(??):132–153, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302851.

Takayasu:2017:NVB

[TMS^{+}17] Akitoshi Takayasu, Kaname Matsue, Takiko Sasaki, Kazuaki Tanaka, Makoto Mizuguchi, and Shin’ichi Oishi. Numerical validation of blow-up solutions of ordinary differential equations. Journal of Computational and Applied Mathematics, 314(??):10– 29, April 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305027.

Tsuchiya:2019:NEC

[TN19] Takuya Tsuchiya and Makoto Nakamura. On the numerical experiments of the Cauchy problem for semi-linear Klein–Gordon equations in the de Sitter spacetime. Journal of Computational and Applied Mathematics, 361(??):396–412, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302390.

Tuan:2018:RSC

[TNTL18] Nguyen Huy Tuan, Tran Bao Ngoc, Salih Tatar, and Le Dinh Long. Recovery of the solute concentration and dispersion flux in an inhomogeneous time fractional diffusion equation. Journal of Computational and Applied Mathematics, 342(??):96–118, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301687.

Tomar:2019:EIM

[TP19] Saurabh Tomar and R. K. Pandey. An efficient iterative method for solving Bratu-type equations. Journal of Computational and Applied Mathematics, 357(??):71–84, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300986.

Tichavsky:2017:NCD

[TPC17] Petr Tichavský, Anh-Huy Phan, and Andrzej Cichocki. Numerical CP decomposition of some difficult tensors. Journal of Computational and Applied Mathematics, 317(??):362– 370, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042716306008.

Tran:2015:PAF

[Tra15] Dong Xuan Tran. Padé approximants for finite time ruin probabilities. Journal of Computational and Applied Mathematics, 278(??):130–137, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004051.

Truhar:2019:NUL

[Tru19] Ninoslav Truhar. A note on an upper and a lower bound on sines between eigenspaces for regular Hermitian matrix pairs. Journal of Computational and Applied Mathematics, 358(??):374–384, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301311.

Teodoro:2017:PAP

[TS17] M. Filomena Teodoro and Carla Simão. Perception about pediatric hypertension. Journal of Computational and Applied Mathematics, 312(??):209–215, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301364.

Tanaka:2017:SNI

[TSMO17] Kazuaki Tanaka, Kouta Sekine, Makoto Mizuguchi, and Shin’ichi Oishi. Sharp numerical inclusion of the best constant for embedding H_{0}^{1}(Ω) ↪ L^{p}(Ω) on bounded convex domain. Journal of Computational and Applied Mathematics, 311(??):306–313, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303405.

Tung:2018:MAP

[Tun18] Michael M. Tung. Modelling acoustics on the Poincaré half-plane. Journal of Computational and Applied Mathematics, 337(??):366–372, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305460.

Talonov:2016:NHS

[TV16a] Alexey Talonov and Maria Vasilyeva. On numerical homogenization of shale gas transport. Journal of Computational and Applied Mathematics, 301(??):44–52, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000406.

Themistoclakis:2016:NSN

[TV16b] W. Themistoclakis and A. Vecchio. On the numerical solution of a nonlocal boundary value problem. Journal of Computational and Applied Mathematics, 292(??):720–731, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001259.

Telen:2018:SNF

[TV18] Simon Telen and Marc Van Barel. A stabilized normal form algorithm for generic systems of polynomial equations. Journal of Computational and Applied Mathematics, 342(??):119–132, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302012.

Tong:2017:LPR

[TW17] Hongzhi Tong and Qiang Wu. Learning performance of regularized moving least square regression. Journal of Computational and Applied Mathematics, 325(??):42–55, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302248.

Tung:2019:AMM

[TW19] Michael M. Tung and Ewa B. Weinmüller. Acoustic metamaterial models on the (2 + 1)D Schwarzschild plane. Journal of Computational and Applied Mathematics, 346(??):162–170, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304291.

Tao:2019:EEP

[TX19] Qi Tao and Yinhua Xia. Error estimates and postprocessing of local discontinuous Galerkin method for Schrödinger equations. Journal of Computational and Applied Mathematics, 356(??):198–218, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300494.

Tan:2018:SOI

[TXFH18] Jieqing Tan, Yan Xing, Wen Fan, and Peilin Hong. Smooth orientation interpolation using parametric quintic-polynomial-based quaternion spline curve. Journal of Computational and Applied Mathematics, 329(??):256–267, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303485.

Tan:2017:CPP

[TYMG17] Jianguo Tan, Hua Yang, Weiwei Men, and Yongfeng Guo. Construction of positivity preserving numerical method for jump–diffusion option pricing models. Journal of Computational and Applied Mathematics, 320:96–100, August 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Teng:2019:NFN

[TYQ^{+}19] Yueyang Teng, Yudong Yao, Shouliang Qi, Chen Li, Lisheng Xu, Wei Qian, and Ge Wang. A novel framework for the NMF methods with experiments to unmixing signals and feature representation. Journal of Computational and Applied Mathematics, 362(??):205–218, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302444.

Takeuchi:2017:SOA

[TYS17] Yuki Takeuchi, Yoshihide Yoshimoto, and Reiji Suda. Second order accuracy finite difference methods for space-fractional partial differential equations. Journal of Computational and Applied Mathematics, 320:101–119, August 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Tian:2018:NMW

[TZZ18] Tian Tian, Qilong Zhai, and Ran Zhang. A new modified weak Galerkin finite element scheme for solving the stationary Stokes equations. Journal of Computational and Applied Mathematics, 329(??):268–279, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300420.

Uda:2018:SCV

[Uda18] Tomoki Uda. Shape calculus for vortex patch equilibria and its application to lattice configurations. Journal of Computational and Applied Mathematics, 332(??):86–100, April 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305241.

Uteshev:2018:PEP

[UG18] Alexei Yu. Uteshev and Marina V. Goncharova. Point-to-ellipse and pointto-ellipsoid distance equation analysis. Journal of Computational and Applied Mathematics, 328(??):232–251, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730362X.

Urena:2019:SSO

[UGG^{+}19] F. Ureña, L. Gavete, A. García, J. J. Benito, and A. M. Vargas. Solving second order non-linear parabolic PDEs using generalized finite difference method (GFDM). Journal of Computational and Applied Mathematics, 354(??):221– 241, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300876.

Ugurlu:2017:CMD

[Ugu17]  Kerem Ugurlu.  Con-

trolled Markov Decision Processes with AVaR criteria for unbounded costs. Journal of Computational and Applied Mathematics, 319(??):24–37, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306410.

Ugurlu:2018:ROC

[Ugu18]  Kerem Ugurlu.  Ro-

bust optimal control using conditional risk mappings in infinite horizon. Journal of Computational and Applied Mathematics, 344(??):275–287, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303017.

Siraj-ul-Islam:2018:MMT

[uIuD18] Siraj ul Islam and Zaheer ud Din. Meshless methods for twodimensional oscillatory Fredholm integral equations. Journal of Computational and Applied Mathematics, 335(??):33– 50, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305812.

Siraj-ul-Islam:2015:NQR

[uIZ15] Siraj ul Islam and Sakhi Zaman. New quadrature rules for highly oscillatory integrals with stationary points. Journal of Computational and Applied Mathematics, 278(??):75–89, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004178.

Uzinski:2015:SSD

[UPD^{+}15] Julio C. Uzinski, Henrique M. Paiva, Marco A. Q. Duarte, Roberto K. H. Galvão, and Francisco Villarreal. A state-space description for perfect-reconstruction wavelet FIR filter banks with special orthonormal basis functions. Journal of Computational and Applied Mathematics, 290 (??):290–297, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002721.

Ueda:2019:SEE

[US19] Yuki Ueda and Norikazu Saito. Stability and error estimates for the successive-projection technique with B-splines in time. Journal of Computational and Applied Mathematics, 358(??):266–278, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301517.

Vabishchevich:2018:TCD

[Vab18] Petr N. Vabishchevich. Two-component domain decomposition scheme with overlapping subdomains for parabolic equations. Journal of Computational and Applied Mathematics, 340(??):664–675, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042717304326.

Vigo-Aguiar:2015:RTC

[VACR15] Jesús Vigo-Aguiar, Alicia Cordero, and Higinio Ramos. Recent trends on Computational and Mathematical Methods in Science and Engineering (CMMSE). Journal of Computational and Applied Mathematics, 275(??):213–215, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003446.

Vigo-Aguiar:2019:SIS

[VAMDGR19] Jesús Vigo-Aguiar, Jorge E.

Macías-Díaz, and Raquel Garcia-Rubio. Special issue: Selected papers of CMMSE. Journal of Computational and Applied Mathematics, 354(??):81– 85, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271930072X.

Vigo-Aguiar:2017:CCT

[VAMGR17] Jesús Vigo-Aguiar, Jesús Medina, and Raquel Garcia-Rubio. Current computational tools for science engineering and economics at CMMSE. Journal of Computational and Applied Mathematics, 318(??):1– 2, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306100.

Vigo-Aguiar:2015:CFT

[VAR15] Jesús Vigo-Aguiar and Higinio Ramos. On the choice of the frequency in trigonometrically-fitted methods for periodic problems. Journal of Computational and Applied Mathematics, 277(??):94–105, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400404X.

Vigo-Aguiar:2017:FAS

[VARC17] J. Vigo-Aguiar, Higinio Ramos, and C. Clavero. A first approach in solving initial-value problems in ODEs by elliptic fitting methods. Journal of Computational and Applied Mathematics, 318(??):599– 603, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006007.

Vasilyeva:2019:NMU

[VCC^{+}19] Maria Vasilyeva, Eric T.

Chung, Siu Wun Cheung, Yating Wang, and Georgy Prokopev. Nonlocal multicontinua upscaling for multicontinua flow problems in fractured porous media. Journal of Computational and Applied Mathematics, 355(??):258–267, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300408.

Vasilyeva:2019:UMP

[VCL^{+}19] Maria Vasilyeva, Eric T.

Chung, Wing Tat Leung, Yating Wang, and Denis Spiridonov. Upscaling method for problems in perforated domains with non-homogeneous boundary conditions on perforations using Non-Local Multi-Continuum method (NLMC). Journal of Computational and Applied Mathematics, 357 (??):215–227, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301037.

vanDoorn:2015:SPB

[vD15] Erik A. van Doorn. Spectral properties of birthdeath polynomials. Journal of Computational and Applied Mathematics, 284(??):251–258, August 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003719.

VanBockstal:2015:IMK

[VDS15] K. Van Bockstal, R. H. De Staelen, and M. Slodicka. Identification of a memory kernel in a semilinear integrodifferential parabolic problem with applications in heat conduction with memory. Journal of Computational and Applied Mathematics, 289(??):196–207, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000898.

Viet:2019:SSC

[VDT19] Tran Quoc Viet, Nguyen Minh Dien, and Dang Duc Trong. Stability of solutions of a class of nonlinear fractional Laplacian parabolic problems. Journal of Computational and Applied Mathematics, 355(??):51–76, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300160.

Vejchodsky:2018:FRL

[Vej18] Tomás Vejchodský. Flux reconstructions in the Lehmann–Goerisch method for lower bounds on eigenvalues. Journal of Computational and Applied Mathematics, 340(??):676–690, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301134.

Verdugo:2016:MAB

[Ver16] Anael Verdugo. Mathematical analysis of a biochemical oscillator with delay. Journal of Computational and Applied Mathematics, 291(??):66–75, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002472.

Vishnevsky:2017:SAS

[VF17] Alexander M. Vishnevsky and Anna A. Firsova. Statistical analysis of solution accuracy for inverse problems in electrodynamics. Journal of Computational and Applied Mathematics, 312(??):174–180, March 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301352.

Vidal-Ferrandiz:2016:MMS

[VFFGV16] A. Vidal-Ferràndiz, R. Fayez, D. Ginestar, and G. Verdú. Moving meshes to solve the time-dependent neutron diffusion equation in hexagonal geometry. Journal of Computational and Applied Mathematics, 291(??):197–208, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001934.

Vahdati:2018:OPU

[VFG18] S. Vahdati, M. Fardi, and M. Ghasemi. Option pricing using a computational method based on reproducing kernel. Journal of Computational and Applied Mathematics, 328(??):252–266, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302777.

Vidal-Ferrandiz:2017:STP

[VFGPG^{+}17] A. Vidal-Ferràndiz, S. González-Pintor, D. Ginestar, G. Verdú, and C. Demazière. Schwarz type preconditioners for the neutron diffusion equation. Journal of Computational and Applied Mathematics, 309(??):563–574, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301157.

Vidal-Ferrandiz:2018:PWH

[VFGPG^{+}18] A. Vidal-Ferràndiz, S. González-Pintor, D. Ginestar, C. Demazière, and G. Verdú. Pin-wise homogenization for SP_{N} neutron transport approximation using the finite element method. Journal of Computational and Applied Mathematics, 330(??):806– 821, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717303230.

Vignati:2018:NSI

[VFN18] Federica Vignati, Damiano Fustinoni, and Alfonso Niro. A novel scale-invariant, dynamic method for hierarchical clustering of data affected by measurement uncertainty. Journal of Computational and Applied Mathematics, 344 (??):521–531, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830339X.

Vignati:2015:MDS

[VG15] F. Vignati and A. Guardone. Multi-domain simulations of shock wave interaction with aerodynamic obstacles in cylindrical implosions. Journal of Computational and Applied Mathematics, 283(??):218–227, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000035.

Vargas:2017:AGA

[VG17] Luisa Fernanda Vargas and Juan Carlos Muñoz Grajales. Analysis of a Galerkin approach applied to a system of coupled Schrödinger equations. Journal of Computational and Applied Mathematics, 313(??):318–342, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304502.

Vigna:2017:FSM

[Vig17] Sebastiano Vigna. Further scramblings of Marsaglia’s xorshift generators. Journal of Computational and Applied Mathematics, 315 (??):175–181, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716305301.

Viswanath:2015:SIL

[Vis15] Divakar Viswanath. Spectral integration of linear boundary value problems. Journal of Computational and Applied Mathematics, 290(??):159–173, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002691.

Viscardi:2019:SRD

[Vis19] Alberto Viscardi. Semi– regular Dubuc-Deslauriers wavelet tight frames. Journal of Computational and Applied Mathematics, 349(??):548–562, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830476X.

VanBarel:2016:NEP

[VK16] Marc Van Barel and Peter Kravanja. Nonlinear eigenvalue problems and contour integrals. Journal of Computational and Applied Mathematics, 292(??):526–540, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271500374X.

Vrsek:2015:DSR

[VL15] Jan Vrsek and Miroslav Lávicka. Determining surfaces of revolution from their implicit equations. Journal of Computational and Applied Mathematics, 290(??):125–135, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002800.

Vlcek:2019:LMO

[VL19] Jan Vlcek and Ladislav Luksan. A limitedmemory optimization method using the infinitely many times repeated BNS update and conjugate directions. Journal of Computational and Applied Mathematics, 351(??):14– 28, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306629.

Vynnycky:2015:NSS

[VM15] M. Vynnycky and S. L.

Mitchell. On the numerical solution of a Stefan problem with finite extinction time. Journal of Computational and Applied Mathematics, 276(??):98–109, March 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400380X.

Vasilyeva:2018:MMR

[VMM18] Maria Vasilyeva, Aashutosh Mistry, and Partha P. Mukherjee. Multiscale model reduction for porescale simulation of Li-ion batteries using GMsFEM. Journal of Computational and Applied Mathematics, 344(??):73–88, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302875.

Vodicka:2017:QNC

[VMR17] Roman Vodicka, Vladislav Mantic, and Tomás Roubícek. Quasistatic normal-compliance contact problem of viscoelastic bodies with Coulomb friction implemented by QP and SGBEM. Journal of Computational and Applied Mathematics, 315 (??):249–272, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S037704271630499X.

Villalba:2016:MBC

[VOB16] Luis Javier García Villalba, Ana Lucila Sandoval Orozco, and Mario Blaum. On multiple burstcorrecting MDS codes. Journal of Computational and Applied Mathematics, 295(??):170–174, March 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005780.

Vela:2019:PAS

[VRBS19] Luis Vela Vela, J. M. Reynolds-Barredo, and Raul Sánchez. A positioning algorithm for SPH ghost particles in smoothly curved geometries. Journal of Computational and Applied Mathematics, 353(??):140– 153, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307520.

VanBockstal:2015:EEF

[VS15a] K. Van Bockstal and M. Slodicka. Error estimates for the full discretization of a nonlocal parabolic model for Type-I superconductors. Journal of Computational and Applied Mathematics, 275(??):516–526, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714000508.

Viquerat:2015:CDG

[VS15b] J. Viquerat and C. Scheid.

A 3D curvilinear discontinuous Galerkin timedomain solver for nanoscale light-matter interactions. Journal of Computational and Applied Mathematics, 289(??):37–50, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001818.

Vermolen:2018:IRP

[VS18] F. J. Vermolen and A. Segal. On an integration rule for products of barycentric coordinates over simplexes in R^{n}. Journal of Computational and Applied Mathematics, 330(??):289–294, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304302.

vanSlingerland:2015:STL

[vSV15] P. van Slingerland and C. Vuik. Scalable twolevel preconditioning and deflation based on a piecewise constant subspace for (SIP)DG systems for diffusion problems. Journal of Computational and Applied Mathematics, 275(??):61–78, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271400315X.

Wout:2016:SAM

[vtWvdHvdVV16] Elwin van ’t Wout, Duncan R. van der Heul, Harmen van der Ven, and Cornelis Vuik. Stability analysis of the marching-on-in-time boundary element method for electromagnetics. Journal of Computational and Applied Mathematics, 294(??):358–371, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004549.

vanVeen:2016:GFB

[vV16] L. van Veen. On Green’s function-based time stepping for semilinear initialboundary value problems. Journal of Computational and Applied Mathematics, 308(??):46–58, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302163.

Volkmer:2019:SLN

[VX19] Hans W. Volkmer and Dexuan Xie. Study of local and nonlocal Poisson–Boltzmann equations for linear and nonlinear models with spherical symmetry. Journal of Computational and Applied Mathematics, 362 (??):653–662, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304023.

Vavpetic:2019:GFO

[VŽ19] Aleš Vavpetič and Emil Žagar. A general framework for the optimal approximation of circular arcs by parametric polynomial curves. Journal of Computational and Applied Mathematics, 345: 146–158, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Verdiere:2018:DIO

[VZDV18] N. Verdière, S. Zhu, and L. Denis-Vidal. A distribution input-output polynomial approach for estimating parameters in nonlinear models. Application to a chikungunya model. Journal of Computational and Applied Mathematics, 331(??):104–118, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304855.

Wheaton:2017:NIM

[WA17] Ira Wheaton and Samuel Awoniyi. A new iterative method for solving non-square systems of linear equations. Journal of Computational and Applied Mathematics, 322(??):1–6, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301401.

Woodfield:2019:SFE

[WAGVRB19] James Woodfield, Mario Alvarez, Bryan Gómez-Vargas, and Ricardo Ruiz-Baier. Stability and finite element approximation of phase change models for natural convection in porous media. Journal of Computational and Applied Mathematics, 360(??):117–137, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301839.

Wang:2015:NEE

[Wan15a]  Chao Wang.  A note

on the error estimation of the Mann iteration. Journal of Computational and Applied Mathematics, 285(??):226–229, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042715000977.

Wang:2015:OIE

[Wan15b] Liang Wang. Optimal interval estimation for a family of lower truncated distribution under progressive censoring. Journal of Computational and Applied Mathematics, 287(??):67–77, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001764.

Wang:2017:IKD

[Wan17a] Liang Wang. Inference for the Kumaraswamy distribution under k-record values. Journal of Computational and Applied Mathematics, 321(??):246–260, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301024.

Wang:2017:UUB

[Wan17b] Wansheng Wang. Uniform ultimate boundedness of numerical solutions to nonlinear neutral delay differential equations. Journal of Computational and Applied Mathematics, 309(??):132–144, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302990.

Wang:2018:NDS

[Wan18a]  Chunmei Wang.  New

discretization schemes for time-harmonic Maxwell equations by weak Galerkin finite element methods. Journal of Computational and Applied Mathematics, 341(??):127–143, October 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301833.

Wang:2018:CRC

[Wan18b]  Haiyong Wang.  On

the convergence rate of Clenshaw–Curtis quadrature for integrals with algebraic endpoint singularities. Journal of Computational and Applied Mathematics, 333(??):87–98, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305368.

Wang:2018:ECS

[Wan18c] Liang Wang. Estimation of constant-stress accelerated life test for Weibull distribution with nonconstant shape parameter. Journal of Computational and Applied Mathematics, 343(??):539–555, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302838.

Wang:2018:EEP

[Wan18d] Liang Wang. Estimation of exponential population with nonconstant parameters under constant-stress model. Journal of Computational and Applied Mathematics, 342(??):478–494, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302206.

Wang:2018:IPC

[Wan18e]  Liang Wang.  Infer-

ence of progressively censored competing risks data from Kumaraswamy distributions. Journal of Computational and Applied Mathematics, 343(??):719–736, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830284X.

Wang:2018:NAT

[Wan18f]  Xiaofeng Wang.  A

new accelerating technique applied to a variant of cordero-torregrosa method. Journal of Computational and Applied Mathematics, 330(??):695– 709, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717303448.

Wang:2018:OTM

[Wan18g]  Xiaofeng Wang.  An

Ostrowski-type method with memory using a novel self-accelerating parameter. Journal of Computational and Applied Mathematics, 330(??):710–720, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301917.

Wang:2018:OHA

[Wan18h] Zhanping Wang. Optimal harvesting for age distribution and weighted size competitive species with diffusion. Journal of Computational and Applied Mathematics, 328(??):485–496, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303771.

Watanabe:2016:ENV

[Wat16] Yoshitaka Watanabe. An efficient numerical verification method for the Kolmogorov problem of incompressible viscous fluid. Journal of Computational and Applied Mathematics, 302(??):157–170, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300358.

Wang:2017:FNV

[WC17] Wansheng Wang and Yingzi Chen. Fast numerical valuation of options with jump under Merton’s model. Journal of Computational and Applied Mathematics, 318(??):79– 92, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305805.

Wang:2018:FNP

[WCLL18] Yinkun Wang, Xiangling Chen, Ying Li, and Jianshu Luo. Fast non-polynomial interpolation and integration for functions with logarithmic singularities. Journal of Computational and Applied Mathematics, 343(??):448–457, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302292.

Wan:2017:ACD

[WCS17] Yuan Wan, Xiao-Li Chen, and Ying Shi. Adaptive cost dynamic time warping distance in time series analysis for classification. Journal of Computational and Applied Mathematics, 319(??):514–520, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300080.

Wei:2018:SCM

[WCS18] Yunxia Wei, Yanping Chen, and Xiulian Shi. A spectral collocation method for multidimensional nonlinear weakly singular Volterra integral equation. Journal of Computational and Applied Mathematics, 331(??):52–63, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304685.

Wang:2015:APB

[WCSL15] Bo Wang, Zemin Cai, Lixin Shen, and Tianshu Liu. An analysis of physics-based optical flow. Journal of Computational and Applied Mathematics, 276(??):62–80, March 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400377X.

Chu:2015:LSD

[wCW15a] Eric King wah Chu and Peter Chang-Yi Weng. Large-scale discrete-time algebraic Riccati equationsdoubling algorithm and error analysis. Journal of Computational and Applied Mathematics, 277(??):115–126, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042714004014.

Wang:2015:ECS

[WCW15b] Lingdi Wang, Wenbin Chen, and Cheng Wang. An energy-conserving second order numerical scheme for nonlinear hyperbolic equation with an exponential nonlinear term. Journal of Computational and Applied Mathematics, 280(??):347–366, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005226.

Wang:2018:POR

[WCW18] Xuezhong Wang, Maolin Che, and Yimin Wei. Partial orthogonal rank-one decomposition of complex symmetric tensors based on the Takagi factorization. Journal of Computational and Applied Mathematics, 332(??):56– 71, April 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305046.

Wang:2019:FES

[WCW19] Feng Wang, Huanzhen Chen, and Hong Wang. Finite element simulation and efficient algorithm for fractional Cahn– Hilliard equation. Journal of Computational and Applied Mathematics, 356(??):248–266, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300536.

Wei:2016:IEP

[WD16] Ying Wei and Hua Dai. An inverse eigenvalue problem for the finite element model of a vibrating rod. Journal of Computational and Applied Mathematics, 300(??):172– 182, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000042.

Wang:2018:TLL

[WD18] Xiaofeng Wang and Weizhong Dai. A three-level linear implicit conservative scheme for the Rosenau-KdV-RLW equation. Journal of Computational and Applied Mathematics, 330(??):295–306, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042717304260.

Wang:2019:CFOa

[WD19] Xiaofeng Wang and Weizhong Dai. A conservative fourth-order stable finite difference scheme for the generalized Rosenau-KdV equation in both 1D and 2D. Journal of Computational and Applied Mathematics, 355(??):310–331, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300573.

Wang:2018:OAN

[WG18] Yunzhu Wang and Jianfang Gao. Oscillation analysis of numerical solutions for delay differential equations with real coefficients. Journal of Computational and Applied Mathematics, 337(??):73–86, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300244.

Wang:2017:IIU

[WGHR17] Si Wang, Weihong Guo, Ting-Zhu Huang, and Garvesh Raskutti. Image inpainting using reproducing kernel Hilbert space and Heaviside functions. Journal of Computational and Applied Mathematics, 311(??):551–564, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304010.

Wang:2016:ELB

[WGLV16] Xuansheng Wang, François Glineur, Linzhang Lu, and Paul Van Dooren. Extended Lanczos bidiagonalization algorithm for low rank approximation and its applications. Journal of Computational and Applied Mathematics, 301(??):213–229, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300176.

Wu:2015:UBB

[WGZ15] Jinming Wu, Dianxuan Gong, and Xiaolei Zhang. An upper bound of the Bezout number for piecewise algebraic curves. Journal of Computational and Applied Mathematics, 273(??):214–224, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271400288X.

Wang:2016:GLM

[WH16] Xiandong Wang and Jianmin He. A geometric Levy model for n-fold compound option pricing in a fuzzy framework. Journal of Computational and Applied Mathematics, 306(??):248–264, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301947.

Wu:2019:ALP

[WH19] Shu-Fei Wu and Ya-Ting Hsieh. The assessment on the lifetime performance index of products with Gompertz distribution based on the progressive type I interval censored sample. Journal of Computational and Applied Mathematics, 351(??):66– 76, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306526.

Wen:2017:NTS

[WHS17] Chun Wen, Ting-Zhu Huang, and Zhao-Li Shen. A note on the two-step matrix splitting iteration for computing PageRank. Journal of Computational and Applied Mathematics, 315(??):87–97, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630509X.

Wu:2018:DML

[WHWY18] Zhizhang Wu, Zhongyi Huang, Wei-Cheng Wang, and Yi Yang. The direct method of lines for elliptic problems in starshaped domains. Journal of Computational and Applied Mathematics, 327(??):350–361, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730328X.

Wang:2016:PWE

[WHZ16] Pengde Wang, Chengming Huang, and Longbin Zhao. Point-wise error estimate of a conservative difference scheme for the fractional Schrödinger equation. Journal of Computational and Applied Mathematics, 306(??):231–247, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630190X.

Weiner:2017:NTF

[WKBB17] R. Weiner, G. Yu. Kulikov, S. Beck, and J. Bruder. New third- and fourthorder singly diagonally implicit two-step peer triples with local and global error controls for solving stiff ordinary differential equations. Journal of Computational and Applied Mathematics, 316(??):380–391, May 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302837.

Wang:2015:FCV

[WKXT15] B. Wang, B. C. Khoo, Z. Q. Xie, and Z. J. Tan. Fast centroidal Voronoi Delaunay triangulation for unstructured mesh generation. Journal of Computational and Applied Mathematics, 280(??):158–173, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005147.

Wu:2016:TSM

[WL16] Shi-Liang Wu and Cui-Xia Li. Two-sweep modulusbased matrix splitting iteration methods for linear complementarity problems. Journal of Computational and Applied Mathematics, 302(??):327–339, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300589.

Wen:2017:DOL

[WL17a] Liping Wen and Qing Liao. Dissipativity of oneleg methods for a class of nonlinear functionalintegro-differential equations. Journal of Computational and Applied Mathematics, 318(??):26– 37, July 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306021.

Wu:2017:SMC

[WL17b] Shi-Liang Wu and Cui-Xia Li. A splitting method for complex symmetric indefinite linear system. Journal of Computational and Applied Mathematics, 313(??):343–354, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304484.

Wu:2017:CTA

[WL17c] Shu-Fei Wu and Meng-Ju Lin. Computational testing algorithmic procedure of assessment for lifetime performance index of products with Weibull distribution under progressive type I interval censoring. Journal of Computational and Applied Mathematics, 311(??):364–374, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630365X.

Wang:2019:CNE

[WL19a] Na-Na Wang and Ji-Cheng Li. A class of new extended shiftsplitting preconditioners for saddle point problems. Journal of Computational and Applied Mathematics, 357(??):123–145, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300883.

Wang:2019:SQF

[WL19b] Xiang Wang and Yonghai Li. Superconvergence of quadratic finite volume method on triangular meshes. Journal of Computational and Applied Mathematics, 348(??):181– 199, March 1, 2019. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718305120.

Wu:2018:CAE

[WLC^{+}18] Shu-Fei Wu, Ying-Tzu Lin, Wen-Jui Chang, Chia-Wei Chang, and Chen Lin. A computational algorithm for the evaluation on the lifetime performance index of products with Rayleigh distribution under progressive type I interval censoring. Journal of Computational and Applied Mathematics, 328(??):508–519, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730345X.

Wang:2018:ECT

[WLLC18] Kun Wang, Hui Liu, Jia Luo, and Zhangxin Chen. Efficient CPRtype preconditioner and its adaptive strategies for large-scale parallel reservoir simulations. Journal of Computational and Applied Mathematics, 328(??):443–468, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303734.

Wu:2016:NFS

[WLM16] Xinyuan Wu, Changying Liu, and Lijie Mei. A new framework for solving partial differential equations using semi-analytical explicit RK(N)-type integrators. Journal of Computational and Applied Mathematics, 301(??):74–90, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300012.

Wang:2019:ACH

[WLW19] Jialing Wang, Dong Liang, and Yushun Wang. Analysis of a conservative highorder compact finite difference scheme for the Klein– Gordon–Schrödinger equation. Journal of Computational and Applied Mathematics, 358(??):84–96, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042719300913.

Wang:2015:EFI

[WLZ15] Chaojie Wang, Hongyi Li, and Di Zhao. An explicit formula for the inverse of a pentadiagonal Toeplitz matrix. Journal of Computational and Applied Mathematics, 278(??):12–18, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003677.

Wang:2017:PTP

[WLZ17] Chaojie Wang, Hongyi Li, and Di Zhao. Preconditioning Toeplitz-plus-diagonal linear systems using the Sherman–Morrison– Woodbury formula. Journal of Computational and Applied Mathematics, 309(??):312–319, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303028.

Wang:2018:NLG

[WLZS18] Zhao Wang, Zhongxuan Luo, Jielin Zhang, and Emil Saucan. A novel local/global approach to spherical parameterization. Journal of Computational and Applied Mathematics, 329(??):294–306, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302479.

Wang:2016:APM

[WMI16] Zhu Wang, Brian McBee, and Traian Iliescu. Approximate partitioned method of snapshots for POD. Journal of Computational and Applied Mathematics, 307(??):374–384, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005774.

Wang:2016:SPC

[WNML16] Rui-Rui Wang, Qiang Niu, Fei Ma, and Lin-Zhang Lu. Spectral properties of a class of matrix splitting preconditioners for saddle point problems. Journal of Computational and Applied Mathematics, 298(??):138–151, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500607X.

Wang:2019:DGM

[WO19] Dong Wang and Braxton Osting. A diffusion generated method for computing Dirichlet partitions. Journal of Computational and Applied Mathematics, 351(??):302–316, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306824.

Wollman:2017:NAS

[Wol17] Stephen Wollman. Numerical approximation of the spatially homogeneous Fokker–Planck– Landau equation. Journal of Computational and Applied Mathematics, 324(??):173–203, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301711.

Weng:2016:SSS

[WP16a] Peter Chang-Yi Weng and Frederick Kin Hing Phoa. Small-sample statistical condition estimation of large-scale generalized eigenvalue problems. Journal of Computational and Applied Mathematics, 298(??):24–39, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005671.

Wischerhoff:2016:RPS

[WP16b] Marius Wischerhoff and Gerlind Plonka. Reconstruction of polygonal shapes from sparse Fourier samples. Journal of Computational and Applied Mathematics, 297(??):117– 131, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005580.

Wang:2017:ROR

[WP17] Wenyuan Wang and Xingchun Peng. Reinsurer’s optimal reinsurance strategy with upper and lower premium constraints under distortion risk measures. Journal of Computational and Applied Mathematics, 315(??):142–160, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305064.

Wang:2018:NHF

[WPA^{+}18] Guotao Wang, Ke Pei, Ravi P. Agarwal, Lihong Zhang, and Bashir Ahmad. Nonlocal Hadamard fractional boundary value problem with Hadamard integral and discrete boundary conditions on a halfline. Journal of Computational and Applied Mathematics, 343(??):230–239, December 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718302656.

Wazwaz:2016:TRM

[WR16] Abdul-Majid Wazwaz and Randolph Rach. Two reliable methods for solving the Volterra integral equation with a weakly singular kernel. Journal of Computational and Applied Mathematics, 302(??):71–80, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300449.

Wang:2018:OIS

[WRZ18] Yajie Wang, Ximin Rong, and Hui Zhao. Optimal investment strategies for an insurer and a reinsurer with a jump diffusion risk process under the CEV model. Journal of Computational and Applied Mathematics, 328(??):414–431, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303898.

Wu:2016:BEE

[WS16] Min Wu and Yimin Shi.

Bayes estimation and expected termination time for the competing risks model from Gompertz distribution under progressively hybrid censoring with binomial removals. Journal of Computational and Applied Mathematics, 300(??):420– 431, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716000273.

Wang:2017:ATG

[WS17] Jin Jiang Wang and Wen Song. An algorithm twisted from generalized ADMM for multi-block separable convex minimization models. Journal of Computational and Applied Mathematics, 309(??):342–358, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042716300401.

Wu:2018:GRL

[WS18] Yong-Hong Wu and Hui Shen. Grey-related least squares support vector machine optimization model and its application in predicting natural gas consumption demand. Journal of Computational and Applied Mathematics, 338(??):212–220, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300700.

Wang:2019:ORE

[WS19] Xin Wang and Shuli Sun. Optimal recursive estimation for networked stochastic uncertain systems with fading measurements and time-correlated channel noises. Journal of Computational and Applied Mathematics, 346(??):549–565, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304461.

Wang:2019:FMI

[WSG^{+}19] Xianchao Wang, Minghui Song, Yukun Guo, Hongjie Li, and Hongyu Liu.

Fourier method for identifying electromagnetic sources with multi-frequency far-field data. Journal of Computational and Applied Mathematics, 358(??):279–292, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301323.

Wang:2017:HEE

[WSL17] Xiuli Wang, Yunquan Song, and Lu Lin. Handling estimating equation with nonignorably missing data based on SIR algorithm. Journal of Computational and Applied Mathematics, 326(??):62–70, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302522.

Wang:2019:CFOb

[WSL19] Bo Wang, Tongjun Sun, and Dong Liang. The conservative and fourthorder compact finite difference schemes for regularized long wave equation. Journal of Computational and Applied Mathematics, 356(??):98–117, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300524.

Wang:2015:SOU

[WSW15] Suxin Wang, Shiyu Song, and Yongjin Wang. Skew Ornstein–Uhlenbeck processes and their financial applications. Journal of Computational and Applied Mathematics, 273(??):363–382, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003045.

Wu:2016:RDB

[Wu16a]  Jibo Wu.  Restricted

difference-based Liu estimator in partially linear model. Journal of Computational and Applied Mathematics, 300(??):97– 102, July 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715006354.

Wu:2016:TEI

[Wu16b] Shu-Lin Wu. Towards essential improvement for the Parareal–TR and Parareal–Gauss4 algorithms.

Journal of Computational and Applied Mathematics, 308(??):391–407, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302709.

Wu:2017:DMC

[Wu17] Tingting Wu. A dispersion minimizing compact finite difference scheme for the 2D Helmholtz equation. Journal of Computational and Applied Mathematics, 311(??):497–512, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303788.

Wu:2018:PAL

[Wu18] Shu-Fei Wu. The performance assessment on the lifetime performance index of products following Chen lifetime distribution based on the progressive Type I interval censored sample. Journal of Computational and Applied Mathematics, 334(??):27–38, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042717305824.

Wu:2019:PSB

[Wu19] Cheng-Hsun Wu. Perturbation solutions for bondpricing equations under a multivariate CIR model with weak dependences. Journal of Computational and Applied Mathematics, 361(??):207–226, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271930202X.

Wang:2016:MMC

[WW16a] Qisheng Wang and Huasheng Wang. Meshless method and convergence analysis for 2-dimensional Fredholm integral equation with complex factors. Journal of Computational and Applied Mathematics, 304(??):18–25, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716301017.

Wei:2016:ISP

[WW16b] Zhaoying Wei and Guangsheng Wei. Inverse spectral problem for nonselfadjoint Dirac operator with boundary and jump conditions dependent on the spectral parameter. Journal of Computational and Applied Mathematics, 308(??):199–214, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302527.

Wang:2019:ECM

[WW19a] Bin Wang and Xinyuan Wu. Exponential collocation methods for conservative or dissipative systems. Journal of Computational and Applied Mathematics, 360(??):99–116, November 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302018.

Wang:2019:MVP

[WW19b] Tianxiao Wang and Jiaqin Wei. Mean-variance portfolio selection under a non-Markovian regimeswitching model. Journal of Computational and Applied Mathematics, 350 (??):442–455, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718306484.

Wang:2019:IRM

[WWH19] Jing Wang, Wei Wang, and Bo Han. An iteration regularization method with general convex penalty for nonlinear inverse problems in Banach spaces. Journal of Computational and Applied Mathematics, 361(??):472–486, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719302420.

Wu:2015:FMM

[WWL15] Ruifeng Wu, Tieru Wu, and Huilai Li. A family of multivariate multiquadric quasi-interpolation operators with higher degree polynomial reproduction. Journal of Computational and Applied Mathematics, 274(??):88–108, January 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003173.

Wang:2017:TCM

[WWM17] Bin Wang, Xinyuan Wu, and Fanwei Meng. Trigonometric collocation methods based on Lagrange basis polynomials for multi-frequency oscillatory second-order differential equations. Journal of Computational and Applied Mathematics, 313(??):185–201, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630437X.

Wang:2018:PIC

[WWST18] Renhong Wang, Zongmin Wu, Xiquan Shi, and Jieqing Tan. Preface: the International Conference on Information and Computational Science, 2– 6 August 2016, Dalian, China. Journal of Computational and Applied Mathematics, 329(??):1, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304727.

Wang:2016:LFW

[WWWZ16] Chunmei Wang, Junping Wang, Ruishu Wang, and Ran Zhang. A lockingfree weak Galerkin finite element method for elasticity problems in the primal formulation. Journal of Computational and Applied Mathematics, 307(??):346–366, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715006275.

Wang:2015:OEB

[WWZ15] Jun-Gang Wang, Ting Wei, and Yu-Bin Zhou. Optimal error bound and simplified Tikhonov regularization method for a backward problem for the time-fractional diffusion equation. Journal of Computational and Applied Mathematics, 279(??):277– 292, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005056.

Wang:2016:WGF

[WWZZ16] Ruishu Wang, Xiaoshen Wang, Qilong Zhai, and Ran Zhang. A weak Galerkin finite element scheme for solving the stationary Stokes equations. Journal of Computational and Applied Mathematics, 302(??):171–185, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716300048.

Wang:2019:EAB

[WXJ19] Shuangyue Wang, Yunhai Xiao, and Zhengfen Jin. An efficient algorithm for batch images alignment with adaptive rank-correction term. Journal of Computational and Applied Mathematics, 346(??):171–183, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304278.

Wang:2015:MIB

[WXW15] Bao Wang, Kelin Xia, and Guo-Wei Wei. Matched interface and boundary method for elasticity interface problems. Journal of Computational and Applied Mathematics, 285 (??):203–225, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000667.

Woo:2017:GSA

[WXY17] Jae-Kyung Woo, Ran Xu, and Hailiang Yang. Gerber–Shiu analysis with two-sided acceptable levels. Journal of Computational and Applied Mathematics, 321(??):185–210, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300729.

Wang:2019:SSI

[WXZ19] Xiang Wang, Xiao-Yong Xiao, and Qing-Qing Zheng. A single-step iteration method for non-Hermitian positive definite linear systems. Journal of Computational and Applied Mathematics, 346(??):471–482, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304412.

Wang:2019:MRH

[WYLD19] Dongqing Wang, Yaru Yan, Yanjun Liu, and Junhang Ding. Model recovery for Hammerstein systems using the hierarchical orthogonal matching pursuit method. Journal of Computational and Applied Mathematics, 345: 135–145, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Wan:2015:PSJ

[WYW15] Zhong Wan, Min Yuan, and Chang Wang. A partially smoothing Jacobian method for nonlinear complementarity problems with P_{0} function. Journal of Computational and Applied Mathematics, 286(??):158–171, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001612.

Wu:2016:OTH

[WZ16] Lan Wu and Jiang Zhou.

Occupation times of hyperexponential jump diffusion processes with application to price step options. Journal of Computational and Applied Mathematics, 294(??):251–274, March 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004537.

Wen:2017:COL

[WZ17] Liping Wen and Yan Zhou.

Convergence of one-leg methods for neutral delay integro-differential equations. Journal of Computational and Applied Mathematics, 317(??):432– 446, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716306215.

Wu:2018:MND

[WZ18] Jinming Wu and Chungang Zhu. The maximum number and its distribution of singular points for parametric piecewise algebraic curves. Journal of Computational and Applied Mathematics, 329(??):322–330, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300444.

Wang:2018:HSF

[WZBM18] Di Wang, Zuoquan Zhang, Rongquan Bai, and Yanan Mao. A hybrid system with filter approach and multiple population genetic algorithm for feature selection in credit scoring. Journal of Computational and Applied Mathematics, 329(??):307–321, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302078.

Wang:2019:CBC

[WZH19] Aizeng Wang, Gang Zhao, and Fei Hou. Constructing Bézier curves with monotone curvature. Journal of Computational and Applied Mathematics, 355(??):1–10, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300111.

Wang:2017:NIM

[WZL17] Teng Wang, Qingqing Zheng, and Linzhang Lu. A new iteration method for a class of complex symmetric linear systems. Journal of Computational and Applied Mathematics, 325(??):188–197, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302388.

Wang:2018:TOI

[WZS18a] Haijin Wang, Qiang Zhang, and Chi-Wang Shu. Third order implicit-explicit Runge– Kutta local discontinuous Galerkin methods with suitable boundary treatment for convection– diffusion problems with Dirichlet boundary conditions. Journal of Computational and Applied Mathematics, 342(??):164–179, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718301729.

Wang:2018:FSL

[WZS^{+}18b] Shuliang Wang, Yiping Zhao, Yue Shu, Hanning Yuan, Jing Geng, and Shaopeng Wang. Fast search local extremum for maximal information coefficient (MIC). Journal of Computational and Applied Mathematics, 327(??):372–387, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302911.

Wang:2019:CNB

[WZS19] Xiaoyue Wang, Xian Zhao, and Jinglei Sun. A compound negative binomial distribution with mutative termination conditions based on a change point. Journal of Computational and Applied Mathematics, 351(??):237–249, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718306745.

Wang:2019:PPR

[WZWJ19] Ruishu Wang, Ran Zhang, Xiuli Wang, and Jiwei Jia. Polynomial preserving recovery for a class of weak Galerkin finite element methods. Journal of Computational and Applied Mathematics, 362(??):528–539, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718305922.

Wang:2016:WGM

[WZZ16] Xiuli Wang, Qilong Zhai, and Ran Zhang. The weak Galerkin method for solving the incompressible Brinkman flow. Journal of Computational and Applied Mathematics, 307(??):13–24, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302047.

Wang:2018:NRC

[WZZ18] Han Wang, Chun-Gang Zhu, and Xuan-Yi Zhao. The number of regular control surfaces of toric patch. Journal of Computational and Applied Mathematics, 329(??):280–293, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301450.

Xiong:2018:MLI

[XC18] Xiangtuan Xiong and Qiang Cheng. A modified Lavrentiev iterative regularization method for analytic continuation. Journal of Computational and Applied Mathematics, 327(??):127–140, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730314X.

Xu:2019:SCL

[XCG19] Wei-Ru Xu, Guo-Liang Chen, and Yi Gong. Submatrix constrained leastsquares inverse problem for symmetric matrices from the design of vibrating structures. Journal of Computational and Applied Mathematics, 355(??):128–142, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300196.

Xue:2017:UCG

[XD17] Changfeng Xue and Shaozhong Deng. Unified construction of Green’s functions for Poisson’s equation in inhomogeneous media with diffuse interfaces. Journal of Computational and Applied Mathematics, 326(??):296–319, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303072.

Xu:2018:CFE

[XDW18] Shipeng Xu, Weibing Deng, and Haijun Wu. A combined finite element method for elliptic problems posted in domains with rough boundaries. Journal of Computational and Applied Mathematics, 336(??):235– 248, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300177.

Xu:2016:EEE

[XDY16] Zhiguo Xu, Xuanchun Dong, and Yongjun Yuan. Error estimates in the energy space for a Gautschitype integrator spectral discretization for the coupled nonlinear Klein– Gordon equations. Journal of Computational and Applied Mathematics, 292(??):402–416, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003799.

Xue:2019:UPS

[XDZZ19] Li Xue, Hao Di, Xuejun Zhao, and Zili Zhang. Uncertain portfolio selection with mental accounts and realistic constraints. Journal of Computational and Applied Mathematics, 346(??):42–52, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304011.

Xu:2016:NSV

[XF16] Haitao Xu and Liya Fan. Numerical solution of Volterra–Fredholm integral equations based on ε-SVR method. Journal of Computational and Applied Mathematics, 298(??):201–210, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S037704271500597X.

Xiang:2018:DAP

[XGH18] Ling Xiang, Nan Gao, and Aijun Hu. Dynamic analysis of a planetary gear system with multiple nonlinear parameters. Journal of Computational and Applied Mathematics, 327(??):325–340, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303217.

Xia:2015:STF

[XH15] Bingxing Xia and Viet Ha Hoang. Sparse tensor finite elements for elastic wave equation with multiple scales. Journal of Computational and Applied Mathematics, 282(??):179– 214, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000047.

Xu:2019:SDG

[XH19] Xiuxiu Xu and Qiumei Huang. Superconvergence of discontinuous Galerkin methods for nonlinear delay differential equations with vanishing delay. Journal of Computational and Applied Mathematics, 348(??):314–327, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305168.

Xu:2019:NAH

[XHH^{+}19] Wei Xu, Ziping Huang, Weimin Han, Wenbin Chen, and Cheng Wang. Numerical analysis of history-dependent variationalhemivariational inequalities with applications in contact mechanics. Journal of Computational and Applied Mathematics, 351 (??):364–377, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718305338.

Xie:2017:NSN

[XHZ17] Jiaquan Xie, Qingxue Huang, and Fuqiang Zhao. Numerical solution of nonlinear Volterra–Fredholm– Hammerstein integral equations in two-dimensional spaces based on block pulse functions. Journal of Computational and Applied Mathematics, 317 (??):565–572, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716306380.

Xia:2017:HUS

[Xia17] Chao Xia. Hyers–Ulam stability of the iterative equation with a general boundary restriction. Journal of Computational and Applied Mathematics, 322(??):7–17, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301437.

Xiang:2019:VCT

[Xia19]  Shuhuang Xiang.  On

van der Corput-type lemmas for Bessel and Airy transforms and applications. Journal of Computational and Applied Mathematics, 351(??):179–185, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306721.

Xie:2015:CAC

[Xie15] Qichang Xie. Computation and application of copula-based weighted average quantile regression. Journal of Computational and Applied Mathematics, 281(??):182– 195, June 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005718.

Xie:2019:IMS

[Xie19] Huiqing Xie. An iterative method for simultaneously computing the derivatives of simple and multiple eigenpairs of a quadratic eigenvalue problem. Journal of Computational and Applied Mathematics, 356(??):133–151, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300263.

Xing:2016:HOF

[Xin16] Yulong Xing. High order finite volume WENO schemes for the shallow water flows through channels with irregular geometry. Journal of Computational and Applied Mathematics, 299(??):229– 244, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006020.

Xu:2019:EAI

[XLS^{+}19] Gang Xu, Bojian Li, Laixin Shu, Long Chen, Jinlan Xu, and Tahsin Khajah. Efficient radaptive isogeometric analysis with Winslow’s mapping and monitor function approach. Journal of Computational and Applied Mathematics, 351(??):186– 197, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830668X.

Xu:2016:EMC

[XM16] Zhenhua Xu and Gradimir V.

Milovanović. Efficient method for the computation of oscillatory Bessel transform and Bessel Hilbert transform. Journal of Computational and Applied Mathematics, 308(??):117–137, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302655.

Xu:2015:ECM

[XMW^{+}15] Gang Xu, Bernard Mourrain, Xiangyang Wu, Long Chen, and Kin chuen Hui. Efficient construction of multi-block volumetric spline parameterization by discrete mask method. Journal of Computational and Applied Mathematics, 290(??):589–597, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003465.

Xu:2015:EHM

[XqYL^{+}15] Gang Xu, Yu qun Yang, Bin-Bin Liu, Yi hong Xu, and Ai jun Wu. An efficient hybrid multiobjective particle swarm optimization with a multiobjective dichotomy line search. Journal of Computational and Applied Mathematics, 280(??):310–326, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005354.

Xu:2015:ANI

[Xu15] Ling Xu. Application of the Newton iteration algorithm to the parameter estimation for dynamical systems. Journal of Computational and Applied Mathematics, 288(??):33–43, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002125.

Xu:2017:EAC

[Xu17] Bo Xu. An efficient ant colony algorithm based on wake-vortex modeling method for aircraft scheduling problem. Journal of Computational and Applied Mathematics, 317(??):157– 170, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305854.

Xu:2018:OSM

[Xu18]  Yingxiang Xu.  Op-

timized Schwarz methods with Ventcell transmission conditions for model problems with continuously variable coefficients. Journal of Computational and Applied Mathematics, 334(??):97–110, May 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305757.

Xiao:2018:CQM

[XW18] Ye Xiao and Xiaoqun Wang. Conditional quasi-Monte Carlo methods and dimension reduction for option pricing and hedging with discontinuous functions. Journal of Computational and Applied Mathematics, 343(??):289–308, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302747.

Xie:2018:NAL

[XWWI18] Xuping Xie, David Wells, Zhu Wang, and Traian Iliescu. Numerical analysis of the Leray reduced order model. Journal of Computational and Applied Mathematics, 328(??):12–29, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303266.

Xing:2015:EVC

[XY15] Yu Xing and Xiaoping Yang. Equilibrium valuation of currency options under a jump-diffusion model with stochastic volatility. Journal of Computational and Applied Mathematics, 280(??):231–247, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005445.

Xie:2016:NBI

[XY16] Dexuan Xie and Jinyong Ying. A new box iterative method for a class of nonlinear interface problems with application in solving Poisson–Boltzmann equation. Journal of Computational and Applied Mathematics, 307(??):319–334, December 1, 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716000091.

Xu:2018:CAP

[XY18a] Gang Xu and Guosong Yu. On convergence analysis of particle swarm optimization algorithm. Journal of Computational and Applied Mathematics, 333 (??):65–73, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717305277.

See reprint [XY18b].

Xu:2018:RCA

[XY18b] Gang Xu and Guosong Yu.

Reprint of: On convergence analysis of particle swarm optimization algorithm. Journal of Computational and Applied Mathematics, 340(??):709–717, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302310. See

[XY18a].

Xie:2019:NRN

[XY19] Jiaquan Xie and Mingxu Yi. Numerical research of nonlinear system of fractional Volterra–Fredholm integral–differential equations via Block-Pulse functions and error analysis. Journal of Computational and Applied Mathematics, 345:159–167, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Xie:2015:AST

[XYZ15] Ziqing Xie, Wenfan Yi, and Jianxin Zhou. An augmented singular transform and its partial Newton method for finding new solutions. Journal of Computational and Applied Mathematics, 286(??):145–157, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042715001260.

Xu:2018:AAG

[XZ18] Wenfang Xu and Ludmil T. Zikatanov. Adaptive aggregation on graphs. Journal of Computational and Applied Mathematics, 340(??):718–730, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305344.

Xu:2019:SSI

[XZ19] Zhen Xu and Hui Zhang.

Stabilized semi-implicit numerical schemes for the Cahn–Hilliard-like equation with variable interfacial parameter. Journal of Computational and Applied Mathematics, 346(??):307–322, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303832.

Xu:2019:SRK

[XZN^{+}19] Minqiang Xu, Zhihong Zhao, Jing Niu, Li Guo, and Yingzhen Lin. A simplified reproducing kernel method for 1-D elliptic type interface problems. Journal of Computational and Applied Mathematics, 351(??):29– 40, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306344.

Zhong:2015:FAS

[xZWlC15] Hong xiu Zhong, Gang Wu, and Guo liang Chen. A flexible and adaptive simpler block GM-RES with deflated restarting for linear systems with multiple right-hand sides. Journal of Computational and Applied Mathematics, 282(??):139– 156, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000059.

Xiong:2018:RFM

[XZXQ18] Xiangtuan Xiong, E. Zhuang, Xuemin Xue, and Zhi Qian. A regularization framework for mildly illposed problems connected with pseudo-differential operator. Journal of Computational and Applied Mathematics, 341(??):1– 11, October 15, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0377042718301444.

Yapman:2019:CAF

[YAA19] Ömer Yapman, Gabil M.

Amiraliyev, and Ilhame Amirali. Convergence analysis of fitted numerical method for a singularly perturbed nonlinear Volterra integrodifferential equation with delay. Journal of Computational and Applied Mathematics, 355(??):301–309, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300421.

Yakubovich:2019:WTI

[Yak19]  S. Yakubovich.  On

the Weber type integral equation related to the Weber–Titchmarsh problem. Journal of Computational and Applied Mathematics, 347(??):150–155, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304916.

Yamada:2017:HOW

[Yam17]  Toshihiro Yamada.  A

higher order weak approximation scheme of multidimensional stochastic differential equations using Malliavin weights. Journal of Computational and Applied Mathematics, 321(??):427–447, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301103.

Yamazaki:2019:TEP

[Yam19] Kazuo Yamazaki. Two examples on the property of the noise in the systems of equations of fluid mechanics. Journal of Computational and Applied Mathematics, 362 (??):460–470, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718305764.

Yang:2017:WHI

[Yan17] Xi Yang. The WR–HSS iteration method for a system of linear differential equations and its applications to the unsteady discrete elliptic problem. Journal of Computational and Applied Mathematics, 315(??):208– 227, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305416.

Yao:2016:MMF

[Yao16] Xudong Yao. A minimax method for finding saddle points of upper semi-differentiable locally Lipschitz continuous functional in Banach space and its convergence. Journal of Computational and Applied Mathematics, 296(??):528–549, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500477X.

Yao:2018:TCL

[Yao18] Xudong Yao. Two classes of Ljusternik–Schnirelman minimax algorithms and an application for finding multiple negative energy solutions of a class of p-Laplacian equations. Journal of Computational and Applied Mathematics, 342(??):495–520, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301961.

Yao:2015:SIE

[YB15] Teng-Teng Yao and Zheng-Jian Bai. Semidefinite inverse eigenvalue problems with prescribed entries and partial eigendata. Journal of Computational and Applied Mathematics, 287(??):115–124, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001909.

Yi:2017:MCS

[YB17] Dokkyun Yi and Sunyoung Bu. A mass conservative scheme for solving the Vlasov–Poisson equation using characteristic curve. Journal of Computational and Applied Mathematics, 324(??):1–16, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301899.

Yao:2016:RBF

[YBC^{+}16] G. Yao, K. M. Bliss, M. Crimi, K. R. Fowler, J. Clark-Stone, W. Li, and P. J. Evans. Radial basis function simulation of slow-release permanganate for groundwater remediation via oxidation.

Journal of Computational and Applied Mathematics, 307(??):235–247, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630053X.

Yilmaz:2016:RSS

[YBD16]  Şerife Yilmaz, Taner

Büyükköroğlu, and Vakif Dzhafarov. Random search of stable member in a matrix polytope. Journal of Computational and Applied Mathematics, 308(??):59–68, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302540.

Yang:2017:IME

[YBLH17] Yidu Yang, Hai Bi, Hao Li, and Jiayu Han. A C^{0} IPG method and its error estimates for the Helmholtz transmission eigenvalue problem. Journal of Computational and Applied Mathematics, 326(??):71–86, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042717301942.

You:2019:NES

[YC19] Honglong You and Chunhao Cai. Nonparametric estimation for a spectrally negative Lévy process based on high frequency data. Journal of Computational and Applied Mathematics, 345: 196–205, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Yang:2018:RDO

[YCF18] Yanfang Yang, Eric T. Chung, and Shubin Fu. Residual driven online mortar mixed finite element methods and applications. Journal of Computational and Applied Mathematics, 340(??):318– 333, October 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718301110.

Yuen:2017:EPF

[YCW17] Kam Chuen Yuen, Mi Chen, and Kam Pui Wat. On the expected penalty functions in a discrete semi-Markov risk model with randomized dividends. Journal of Computational and Applied Mathematics, 311(??):239–251, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303430.

Yang:2019:EMA

[YCX19] Lihong Yang, Zhong Chen, and Kechao Xie. An efficient method for approximate solution of a singular integral equation with Cauchy kernel. Journal of Computational and Applied Mathematics, 352(??):50–61, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307040.

Yu:2018:NST

[YFS18] Ching-Hao Yu, Bao-Feng Feng, and Tony W. H. Sheu. Numerical solutions to a two-component Camassa–Holm equation. Journal of Computational and Applied Mathematics, 336(??):317–337, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.
com/science/article/
pii/S0377042718300116.

Yu:2019:LSA

[YFwC19] Bo Yu, Hung-Yuan Fan, and Eric King wah Chu. Large-scale algebraic Riccati equations with highrank constant terms. Journal of Computational and Applied Mathematics, 361(??):130–143, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302006.

Yin:2019:EPE

[YGC19] Li Yin, Zhiming Gao, and Yanzhao Cao. An explicit a posteriori error estimations for the cell functional minimization scheme of elliptic problems. Journal of Computational and Applied Mathematics, 362 (??):309–323, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718306423.

Yang:2018:NCA

[YGS18] Xiao-Jun Yang, Feng Gao, and H. M. Srivastava. A new computational approach for solving nonlinear local fractional PDEs. Journal of Computational and Applied Mathematics, 339(??):285–296, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304971.

You:2018:FEA

[YH18a] Cuilian You and Yangyang Hao. Fuzzy Euler approximation and its local convergence. Journal of Computational and Applied Mathematics, 343(??):55–61, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302139.

Yuan:2018:CTK

[YH18b] Long Yuan and Qiya Hu. Comparisons of three kinds of plane wave methods for the Helmholtz equation and time-harmonic Maxwell equations with complex wave numbers. Journal of Computational and Applied Mathematics, 344 (??):323–345, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302954.

Yang:2018:NSS

[YHZ18] Yin Yang, Yunqing Huang, and Yong Zhou. Numerical solutions for solving time fractional Fokker– Planck equations based on spectral collocation methods. Journal of Computational and Applied Mathematics, 339(??):389–404, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301589.

Yi:2019:LOW

[Yi19] Son-Young Yi. A lowestorder weak Galerkin method for linear elasticity. Journal of Computational and Applied Mathematics, 350 (??):286–298, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S037704271830623X.

Yin:2019:NHF

[Yin19] Guojian Yin. On the non-Hermitian FEAST algorithms with oblique projection for eigenvalue problems. Journal of Computational and Applied Mathematics, 355(??):23–35, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/

science/article/pii/
S0377042719300251.

Yousefi:2019:CAC

[YJBM19] A. Yousefi, S. Javadi, E. Babolian, and E. Moradi. Convergence analysis of the Chebyshev–Legendre spectral method for a class of Fredholm fractional integro-differential equations. Journal of Computational and Applied Mathematics, 358(??):97–110, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300950.

Yang:2018:GBP

[YL18a] Guang Yang and Xinwang Liu. Goal-based portfolio choice model with discounted preference. Journal of Computational and Applied Mathematics, 328(??):365–380, January 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303801.

Yang:2018:DBW

[YL18b] Jiaqing Yang and Keji Liu. Detecting buried wave-penetrable scatterers in a two-layered medium. Journal of Computational and Applied Mathematics, 330(??):318– 329, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717304107.

Yang:2018:GWG

[YL18c] Zhaochen Yang and Shijun Liao. On the generalized wavelet-Galerkin method. Journal of Computational and Applied Mathematics, 331(??):178–195, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304831.

Yuan:2018:CMA

[YL18d] Weipeng Yuan and Shaoyong Lai. The CEV model and its application to financial markets with volatility uncertainty. Journal of Computational and Applied Mathematics, 344(??):25–36, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302863.

Yang:2019:ESM

[YL19a] Qigui Yang and Guangjie Li. Exponential stability of θ-method for stochastic differential equations in the G-framework. Journal of Computational and Applied Mathematics, 350 (??):195–211, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042718306277.

Yuan:2019:FOI

[YL19b] Weipeng Yuan and Shaoyong Lai. Family optimal investment strategy for a random household expenditure under the CEV model. Journal of Computational and Applied Mathematics, 354(??):1– 14, July 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042719300081.

Yang:2015:MTP

[YLL15] Ximei Yang, Hongwei Liu, and Changhe Liu. A Mehrotra-type predictor– corrector infeasible-interior-point method with a new one-norm neighborhood for symmetric optimization. Journal of Computational and Applied Mathematics, 283(??):106–121, August 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000394.

Yu:2018:SPB

[YLZG18] Xiaokang Yu, Na Lei, Xiaopeng Zheng, and Xianfeng Gu. Surface parameterization based on polar factorization. Journal of Computational and Applied Mathematics, 329(??):24–36, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717300079.

Yepez-Martinez:2019:NMD

[YMGA19] H. Yépez-Martínez and J. F. Gómez-Aguilar. A new modified definition of Caputo–Fabrizio fractional-order derivative and their applications to the Multi Step Homotopy Analysis Method (MHAM). Journal of Computational and Applied Mathematics, 346(??):247–260, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304436.

Yao:2018:ATT

[YN18] Shengwei Yao and Liangshuo Ning. An adaptive three-term conjugate gradient method based on self-scaling memoryless BFGS matrix. Journal of Computational and Applied Mathematics, 332 (??):72–85, April 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717305034.

Yla-Oijala:2015:CDE

[YOK15] Pasi Ylä-Oijala and Sami P.

Kiminki. Challenges in developing efficient Calderon preconditioners for resonating or high material contrast penetrable objects. Journal of Computational and Applied Mathematics, 289(??):296–305, December 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500014X.

Yoshikawa:2017:EMS

[Yos17]  Shuji Yoshikawa.  En-

ergy method for structurepreserving finite difference schemes and some properties of difference quotient. Journal of Computational and Applied Mathematics, 311(??):394–413, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303685.

Yousef:2019:SFA

[You19] A. M. Yousef. Stability and further analytical bifurcation behaviors of Moran–Ricker model with delayed density dependent birth rate regulation. Journal of Computational and Applied Mathematics, 355(??):143–161, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300214.

Yu:2017:MEB

[YP17] Yongchao Yu and Jigen Peng. The Moreau envelope based efficient first-order methods for sparse recovery. Journal of Computational and Applied Mathematics, 322(??):109–128, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301334.

Yadav:2018:SCE

[YP18] Sangita Yadav and Amiya K.

Pani. Superconvergence of a class of expanded discontinuous Galerkin methods for fully nonlinear elliptic problems in divergence form. Journal of Computational and Applied Mathematics, 333(??):215–234, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305496.

Yun:2016:GTM

[YPD16] Beong In Yun, Miodrag Petković, and Jovana Dzunić. On a general transformation of multipoint root-solvers. Journal of Computational and Applied Mathematics, 292(??):136–149, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003490.

Yang:2019:DDC

[YS19] Qi Yang and Yuping Song. Data driven confidence intervals for diffusion process using double smoothing empirical likelihood. Journal of Computational and Applied Mathematics, 348(??):282–297, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305144.

Yang:2017:HFM

[YSCY17] Zhiqiang Yang, Yi Sun, Junzhi Cui, and Zihao Yang. The hole-filling method and multiscale algorithm for the heat transfer performance of periodic porous materials. Journal of Computational and Applied Mathematics, 311(??):583–598, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303776.

Yang:2018:IRI

[YSG18] Liu Yang, Yuquan Sun, and Fanghui Gong. The inexact residual iteration method for quadratic eigenvalue problem and the analysis of convergence. Journal of Computational and Applied Mathematics, 332(??):45– 55, April 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042717304934.

Yang:2017:PAI

[YSL17] Fan Yang, Yi Shen, and Zhi Song Liu. The proximal alternating iterative hard thresholding method for l_{0} minimization, with complexity O(1/\sqrt{k}). Journal of Computational and Applied Mathematics, 311(??):115–129, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303326.

Yuan:2018:MVE

[YSL^{+}18] Yirang Yuan, Tongjun Sun, Changfeng Li, Yunxin Liu, and Qing Yang. Mixed volume element combined with characteristic mixed finite volume element method for oilwater two phase displacement problem. Journal of Computational and Applied Mathematics, 340(??):404–419, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301171.

Yuan:2018:GCM

[YSW^{+}18] Gonglin Yuan, Zhou Sheng, Bopeng Wang, Wujie Hu, and Chunnian Li. The global convergence of a modified BFGS method for nonconvex functions. Journal of Computational and Applied Mathematics, 327(??):274–294, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302753.

Yang:2016:HTI

[YT16] Jin Yang and Sanyi Tang. Holling type II predator– prey model with nonlinear pulse as statedependent feedback control. Journal of Computational and Applied Mathematics, 291(??):225–241, January 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000291.

Yin:2015:SMU

[YTSZ15] Fukang Yin, Tian Tian, Junqiang Song, and Min Zhu. Spectral methods using Legendre wavelets for nonlinear Klein/Sine-Gordon equations. Journal of Computational and Applied Mathematics, 275(??):321–334, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003331.

Yuan:2015:CML

[Yua15]  Quan Yuan.  A con-

strained matrix leastsquares problem in structural dynamics model updating. Journal of Computational and Applied Mathematics, 280(??):367–376, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005068.

Yang:2015:CBS

[YW15a] Huogen Yang and Guozhao Wang. Construction of B-spline surface with Bspline curves as boundary geodesic quadrilateral. Journal of Computational and Applied Mathematics, 290(??):104–113, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002770.

Yang:2015:ODB

[YW15b] Huogen Yang and Guozhao Wang. Optimized design of Bézier surface through Bézier geodesic quadrilateral. Journal of Computational and Applied Mathematics, 273(??):264–273, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002908.

Yang:2015:MTR

[YW15c] Xiao-Juan Yang and Li Wang. A modified Tikhonov regularization method. Journal of Computational and Applied Mathematics, 288(??):180–192, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002290.

Yang:2016:FCN

[YW16] Hu Yang and Shaoxin Wang. A flexible condition number for weighted linear least squares problem and its statistical estimation. Journal of Computational and Applied Mathematics, 292(??):320–328, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003751.

Yuan:2019:PRS

[YWW19] Ziyang Yuan, Hongxia Wang, and Qi Wang. Phase retrieval via sparse Wirtinger flow. Journal of Computational and Applied Mathematics, 355(??):162–173, August 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719300184.

Yuan:2019:GCP

[YWY19] Gonglin Yuan, Zengxin Wei, and Yuning Yang. The global convergence of the Polak–Ribière– Polyak conjugate gradient algorithm under inexact line search for nonconvex functions. Journal of Computational and Applied Mathematics, 362 (??):262–275, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718306666.

Yang:2019:LSA

[YX19] Shaojie Yang and Tianzhou Xu. Lie symmetry analysis for a parabolic Monge– Ampère equation in the optimal investment theory. Journal of Computational and Applied Mathematics, 346(??):483–489, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304618.

Yang:2016:STE

[YY16] Jing Yang and Hu Yang. Smooth-threshold estimating equations for varying coefficient partially nonlinear models based on orthogonality-projection method. Journal of Computational and Applied Mathematics, 302(??):24– 37, August 15, 2016. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716300188.

Yang:2015:SNM

[YYL15] Zhixin Yang, G. Yin, and Haibo Li. Stability of numerical methods for jump diffusions and Markovian switching jump diffusions. Journal of Computational and Applied Mathematics, 275(??):197–212, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003690.

Yang:2015:OEN

[YZ15a] Yi-Jun Yang and Wei Zeng. Optimizing equiareality of NURBS surfaces using composite Möbius transformations. Journal of Computational and Applied Mathematics, 279 (??):1–12, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042714004646.

Ye:2015:DHE

[YZ15b] Qiang Ye and Weifeng Zhi. Discrete Hessian eigenmaps method for dimensionality reduction. Journal of Computational and Applied Mathematics, 278(??):197–212, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004075.

Yuan:2015:TTP

[YZ15c] Gonglin Yuan and Maojun Zhang. A three-terms Polak–Ribière–Polyak conjugate gradient algorithm for large-scale nonlinear equations. Journal of Computational and Applied Mathematics, 286(??):186–195, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001600.

Yang:2016:DGI

[YZ16] Qing Yang and Xu Zhang. Discontinuous Galerkin immersed finite element methods for parabolic interface problems. Journal of Computational and Applied Mathematics, 299 (??):127–139, June 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715005658.

Zhang:2017:CAM

[yZ17] Xiao yong Zhang. Convergence analysis of the multistep Legendre pseudospectral method for Volterra integral equations with vanishing delays. Journal of Computational and Applied Mathematics, 321 (??):284–301, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730105X.

Yu:2018:FQP

[YZ18] Zhiheng Yu and Weinian Zhang. Forts of quadratic polynomials under iteration. Journal of Computational and Applied Mathematics, 331(??):1–10, March 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304259.

Yang:2018:LSO

[YZH18] Xiaofeng Yang, Jia Zhao, and Xiaoming He. Linear, second order and unconditionally energy stable schemes for the viscous Cahn–Hilliard equation with hyperbolic relaxation using the invariant energy quadratization method. Journal of Computational and Applied Mathematics, 343(??):80–97, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302085.

Yang:2015:UAB

[YZJC15] Yang Yang, Zhimin Zhang, Tao Jiang, and Dongya Cheng. Uniformly asymptotic behavior of ruin probabilities in a timedependent renewal risk model with stochastic return. Journal of Computational and Applied Mathematics, 287(??):32–43, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001661.

Zhang:2015:PMB

[yZLX15] Cheng yi Zhang, Shuanghua Luo, and Zongben Xu. On parallel multisplitting block iterative methods for linear systems arising in the numerical solution of Euler equations. Journal of Computational and Applied Mathematics, 279(??):249–260, May 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004828.

Yan:2018:RAM

[YZPL18] Zaizai Yan, Tiefeng Zhu, Xiuyun Peng, and Xiaoyan Li. Reliability analysis for multilevel stress testing with Weibull regression model under the general progressively Type-II censored data. Journal of Computational and Applied Mathematics, 330(??):28–40, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303631.

Yang:2015:GCP

[YZW15] Ai-Li Yang, Guo-Feng Zhang, and Yu-Jiang Wu. General constraint preconditioning iteration method for singular saddle-point problems. Journal of Computational and Applied Mathematics, 282(??):157– 166, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000072.

Yang:2017:AFT

[YZY17] Yang Yang, Ting Zhang, and Kam C. Yuen. Approximations for finitetime ruin probability in a dependent discrete-time risk model with CMC simulations. Journal of Computational and Applied Mathematics, 321 (??):143–159, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300626.

Zuo:2016:MCP

[yZyMxG^{+}16] Xian yu Zuo, Ze yao Mo, Tong xiang Gu, Xiao wen Xu, and Ai qing Zhang. Multi-core parallel robust structured multifrontal factorization method for large discretized PDEs. Journal of Computational and Applied Mathematics, 296(??):36–46, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042715004641.

Yang:2019:NBA

[YZZ19] Zhanwen Yang, Jiwei Zhang, and Chengchao Zhao. Numerical blowup analysis of linearly implicit Euler method for nonlinear parabolic integro-differential equations. Journal of Computational and Applied Mathematics, 358(??):343–358, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719301402.

Yang:2019:IIP

[YZZL19] Zhen-Ping Yang, Jin Zhang, Xide Zhu, and Gui-Hua Lin. Infeasible interior-point algorithms based on sampling average approximations for a class of stochastic complementarity problems and their applications. Journal of Computational and Applied Mathematics, 352(??):382–400, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307386.

Zagar:2018:CSA

[Zag18]  Emil Zagar.  Circular

sector area preserving approximation of circular arcs by geometrically smooth parametric polynomials. Journal of Computational and Applied Mathematics, 336(??):63– 71, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717306441.

Zaky:2019:RHO

[Zak19] Mahmoud A. Zaky. Recovery of high order accuracy in Jacobi spectral collocation methods for fractional terminal value problems with non-smooth solutions. Journal of Computational and Applied Mathematics, 357(??):103–122, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300834.

Zhai:2018:MRM

[ZB18] Jia Zhai and Manying Bai. Mean-risk model for uncertain portfolio selection with background risk. Journal of Computational and Applied Mathematics, 330(??):59–69, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303916.

Zhao:2016:NSE

[ZBL^{+}16] Meng Zhao, Andrew Belmonte, Shuwang Li, Xiaofan Li, and John Lowengrub. Nonlinear simulations of elastic fingering in a Hele–Shaw cell. Journal of Computational and Applied Mathematics, 307(??):394–407, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715005610.

Zhao:2016:SLC

[ZC16a] Qi Zhao and Zhongwen Chen. On the superlinear local convergence of a penalty-free method for nonlinear semidefinite programming. Journal of Computational and Applied Mathematics, 308(??):1–19, December 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302242.

Zhu:2016:EAH

[ZC16b] Huiqing Zhu and Fatih Celiker. Error analysis of an HDG method for a distributed optimal control problem. Journal of Computational and Applied Mathematics, 307(??):2–12, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716302059.

Zhu:2018:NSL

[ZC18] Huiqing Zhu and Fatih Celiker. Nodal superconvergence of the local discontinuous Galerkin method for singularly perturbed problems. Journal of Computational and Applied Mathematics, 330(??):95– 116, March 1, 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717303825.

Zhang:2019:ROP

[ZC19] Qiang Zhang and Ping Chen. Robust optimal proportional reinsurance and investment strategy for an insurer with defaultable risks and jumps. Journal of Computational and Applied Mathematics, 356(??):46–66, August 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719300500.

Zhao:2015:ICR

[ZCC15] Chen Zhao, Bencheng Chai, and Xuelin Chen. An improved convergence rate of A-P reconstruction algorithm and application in signal processing. Journal of Computational and Applied Mathematics, 287(??):12–19, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715000709.

Zhou:2017:SYB

[ZCD17] Duanmei Zhou, Guoliang Chen, and Jiu Ding. Solving the Yang– Baxter-like matrix equation for rank-two matrices. Journal of Computational and Applied Mathematics, 313(??):142–151, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304277.

Zhang:2017:MAH

[ZCFW17] Yongwei Zhang, Liqun Cao, Yangde Feng, and Wu Wang. A multiscale approach and a hybrid FE–BE algorithm for heterogeneous scattering of Maxwell’s equations. Journal of Computational and Applied Mathematics, 319(??):460–479, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300389.

Zeng:2018:TGM

[ZCH18] Jiaoyan Zeng, Yanping Chen, and Hanzhang Hu. Two-grid method for compressible miscible displacement problem by CFEM–MFEM. Journal of Computational and Applied Mathematics, 337(??):175–189, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300098.

Zhao:2016:MNC

[ZCL16] Xiukun Zhao, Yanli Chen, and Junliang Lv. A mixed nonoverlapping covolume method on quadrilateral grids for elliptic problems. Journal of Computational and Applied Mathematics, 292(??):23–40, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715003477.

Zhang:2019:BTE

[ZCS^{+}19] Jianming Zhang, Baotao Chi, Krishna M. Singh, YuDong Zhong, and Chuanming Ju. A binary-tree element subdivision method for evaluation of nearly singular domain integrals with continuous or discontinuous kernel. Journal of Computational and Applied Mathematics, 362(??):22–40, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042719302249.

Zafar:2018:OIM

[ZCST18] Fiza Zafar, Alicia Cordero, Sana Sultana, and Juan R. Torregrosa. Optimal iterative methods for finding multiple roots of nonlinear equations using weight functions and dynamics. Journal of Computational and Applied Mathematics, 342(??):352–374, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830205X.

Zhang:2017:PEA

[ZCZM17] Bei Zhang, Shaochun Chen, Jikun Zhao, and Shipeng Mao. A posteriori error analysis of nonconforming finite element methods for convectiondiffusion problems. Journal of Computational and Applied Mathematics, 321 (??):416–426, September 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301115.

Zurnaci:2019:NPD

[ZD19] Fatma Zürnaci and Çetin Di̇si̇büyük. Non-polynomial divided differences and Bspline functions. Journal of Computational and Applied Mathematics, 349(??):579–592, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305776.

Zlatev:2017:SRE

[ZDF^{+}17] Zahari Zlatev, Ivan Dimov, István Faragó, Krassimir Georgiev, and Ágnes Havasi. Stability of the Richardson extrapolation combined with some implicit Runge–Kutta methods. Journal of Computational and Applied Mathematics, 310(??):224–240, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301388.

Zlatev:2018:NMA

[ZDFF18] Zahari Zlatev, Pasqua D’Ambra, István Faragó, and Imre Fekete. Numerical methods for advanced scientific models: Editorial introduction. Journal of Computational and Applied Mathematics, 340(??):421–423, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830147X.

Zlatev:2017:NAS

[ZDGM17] Zahari Zlatev, Ivan Dimov, Krassimir Georgiev, and Svetozar Margenov. Numerical algorithms for scientific and engineering applications. Journal of Computational and Applied Mathematics, 310(??):1–4, January 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716303338.

Zlatev:2016:ENM

[ZDL16] Zahari Zlatev, Ivan Dimov, and Ivan Lirkov. Efficient numerical methods for large-scale scientific computations: Introduction. Journal of Computational and Applied Mathematics, 293(??):1–6, February 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002745.

Zeng:2015:DBB

[ZDLD15] Chao Zeng, Fang Deng, Xin Li, and Jiansong Deng. Dimensions of biquadratic and bicubic spline spaces over hierarchical T -meshes. Journal of Computational and Applied Mathematics, 287(??):162–178, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715001442.

Ziadi:2017:GOT

[ZEBM17] Raouf Ziadi, Rachid Ellaia, and Abdelatif Bencherif-Madani. Global optimization through a stochastic perturbation of the Polak– Ribière conjugate gradient method. Journal of Computational and Applied Mathematics, 317(??):672– 684, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306318.

Zerbetto:2015:STS

[ZF15] Mirco Zerbetto and Diego Frezzato. Summation through stochastic drawing of addends under steered morphing. Journal of Computational and Applied Mathematics, 278(??):101–109, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714004361.

Ziebell:2018:RTF

[ZFK18] Juliana S. Ziebell, Leandro Farina, and Sergey Korotov. Resuspension and transport of fine sediments by waves over a thin layer of viscoelastic mud with erosion. Journal of Computational and Applied Mathematics, 340(??):745–752, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304909.

Zhang:2017:PSB

[ZG17] Qingye Zhang and Yan Gao. Portfolio selection based on a benchmark process with dynamic value-at-risk constraints. Journal of Computational and Applied Mathematics, 313(??):440–447, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304800.

Zhang:2019:SIR

[ZG19a] Zheng Zhang and Wenhao Gui. Statistical inference of reliability of generalized Rayleigh distribution under progressively type-II censoring. Journal of Computational and Applied Mathematics, 361(??):295–312, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302286.

Zhou:2019:ISP

[ZG19b] Y. C. Zhou and Varun Gupta. Interface solutions of partial differential equations with point singularity. Journal of Computational and Applied Mathematics, 362 (??):400–409, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042718306071.

Zhuang:2019:HDD

[ZG19c] Qiao Zhuang and Ruchi Guo. High degree discontinuous Petrov–Galerkin immersed finite element methods using fictitious elements for elliptic interface problems. Journal of Computational and Applied Mathematics, 362 (??):560–573, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S037704271830579X.

Zuo:2015:TGD

[ZH15] Liyun Zuo and Yanren Hou. A two-grid decoupling method for the mixed Stokes–Darcy model. Journal of Computational and Applied Mathematics, 275(??):139–147, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003653.

Zhang:2016:PRC

[ZH16] Yiteng Zhang and Randy Hazlett. Parametric representation of cell boundary flux distribution in well equations. Journal of Computational and Applied Mathematics, 307(??):65–71, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301984.

Zhang:2018:VSM

[ZH18] Sheng Zhang and Siyu Hong. Variable separation method for a nonlinear time fractional partial differential equation with forcing term. Journal of Computational and Applied Mathematics, 339 (??):297–305, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042717304867.

Zhang:2015:PAM

[Zha15a] Naimin Zhang. On parameter acceleration methods for saddle point problems. Journal of Computational and Applied Mathematics, 288(??):169–179, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002460.

Zhang:2015:PGF

[Zha15b]  Ting Zhang.  Petrov–

Galerkin finite element approximation for the three dimensional well model. Journal of Computational and Applied Mathematics, 277(??):47–56, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003896.

Zhang:2016:SRK

[Zha16] G. L. Zhang. Stability of Runge–Kutta methods for linear impulsive delay differential equations with piecewise constant arguments. Journal of Computational and Applied Mathematics, 297(??):41– 50, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005488.

Zhang:2017:RBF

[Zha17a] Shengliang Zhang. Radial basis functions method for valuing options: a multinomial tree approach. Journal of Computational and Applied Mathematics, 319(??):97–107, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716306471.

Zhao:2017:CMT

[Zha17b] Xiaofei Zhao. A combination of multiscale time integrator and two-scale formulation for the nonlinear Schrödinger equation with wave operator. Journal of Computational and Applied Mathematics, 326(??):320–336, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717303060.

Zhang:2018:EIP

[Zha18]  Xuemei Zhang.  Ex-

act interval of parameter and two infinite families of positive solutions for a n-th order impulsive singular equation. Journal of Computational and Applied Mathematics, 330(??):896–908, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717302662.

Zhang:2019:NFF

[Zha19a]  Wensheng Zhang.  A

new family of fourth-order locally one-dimensional schemes for the 3D elastic wave equation. Journal of Computational and Applied Mathematics, 348 (??):246–260, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042718305430.

Zhao:2019:SIM

[Zha19b] Quan Zhao. A sharpinterface model and its numerical approximation for solid-state dewetting with axisymmetric geometry. Journal of Computational and Applied Mathematics, 361(??):144–156, December 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301888.

Zhang:2019:BEM

[ZHC19] Yarong Zhang, Yinnian He, and Hongbin Chen. Boundary element method for a free third boundary problem modeling tumor growth with spectral accuracy. Journal of Computational and Applied Mathematics, 345: 434–451, January 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Zhang:2019:GFS

[ZHH^{+}19] Yunong Zhang, Liu He, Chaowei Hu, Jinjin Guo, Jian Li, and Yang Shi. General four-step discretetime zeroing and derivative dynamics applied to time-varying nonlinear optimization. Journal of Computational and Applied Mathematics, 347(??):314–329, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304941.

Zhu:2015:CCB

[ZHL15] Yuanpeng Zhu, Xuli Han, and Shengjun Liu. Curve construction based on four αβ–Bernstein-like basis functions. Journal of Computational and Applied Mathematics, 273(??):160–181, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714002878.

Zhou:2016:IOC

[Zho16]  Jianwei Zhou.  Im-

proved optimal conditions and iterative parameters for the optimal control problems with an integral constraint in square. Journal of Computational and Applied Mathematics, 307(??):367–373, December 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715006263.

Zhang:2018:EED

[ZHS18] Yuhong Zhang, Yanren Hou, and Li Shan. Error estimates of a decoupled algorithm for a fluidfluid interaction problem. Journal of Computational and Applied Mathematics, 333(??):266–291, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717305484.

Zhang:2016:FAE

[ZHWS16] Hong-Fan Zhang, Ting-Zhu Huang, Chun Wen, and Zhao-Li Shen. FOM accelerated by an extrapolation method for solving PageRank problems. Journal of Computational and Applied Mathematics, 296(??):397–409, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715004793.

Zhang:2017:MMF

[ZHY^{+}17] Qingfu Zhang, Zhaoqin Huang, Jun Yao, Yueying Wang, and Yang Li. A multiscale mixed finite element method with oversampling for modeling flow in fractured reservoirs using discrete fracture model. Journal of Computational and Applied Mathematics, 323(??):95–110, October 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717301607.

Zhang:2017:NFF

[ZJ17a] Wensheng Zhang and Jiangjun Jiang. A new family of fourth-order locally one-dimensional schemes for the threedimensional wave equation. Journal of Computational and Applied Mathematics, 311(??):130–147, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716303399.

Zhou:2017:SCI

[ZJ17b] Shaobo Zhou and Hai Jin. Strong convergence of implicit numerical methods for nonlinear stochastic functional differential equations. Journal of Computational and Applied Mathematics, 324(??):241–257, November 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730170X.

Zhou:2019:INS

[ZJ19] Shaobo Zhou and Hai Jin. Implicit numerical solutions to neutral-type stochastic systems with superlinearly growing coefficients. Journal of Computational and Applied Mathematics, 350(??):423–441, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830637X.

Zhang:2015:TTS

[ZJG^{+}15] Yunong Zhang, Long Jin, Dongsheng Guo, Yonghua Yin, and Yao Chou. Taylor-type 1-step-ahead numerical differentiation rule for first-order derivative approximation and ZNN discretization. Journal of Computational and Applied Mathematics, 273(??):29–40, January 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714002763.

Zore:2016:LIT

[ZJK16] Urska Zore, Bert Jüttler, and Jirí Kosinka. On the linear independence of truncated hierarchical generating systems.

Journal of Computational and Applied Mathematics, 306(??):200–216, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630187X.

Zhang:2018:OQS

[ZJQW18] Nan Zhang, Zhuo Jin, Linyi Qian, and Rongming Wang. Optimal quota-share reinsurance based on the mutual benefit of insurer and reinsurer. Journal of Computational and Applied Mathematics, 342(??):337–351, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302115.

Zheng:2015:CVP

[ZKRL15] Xiaoming Zheng, Yeonhyang Kim, Leela Rakesh, and En-Bing Lin. A conservative and variation preserving finite volume method for nonoverlapping meshes of reaction and diffusion in angiogenesis. Journal of Computational and Applied Mathematics, 275(??):183–196, February 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714003598.

Zhang:2015:NMNb

[ZL15a] Ruimin Zhang and Yingzhen Lin. A novel method for nonlinear boundary value problems. Journal of Computational and Applied Mathematics, 282(??):77– 82, July 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005913.

Zheng:2015:MBN

[ZL15b] Hua Zheng and Wen Li.

The modulus-based nonsmooth Newton’s method for solving linear complementarity problems. Journal of Computational and Applied Mathematics, 288(??):116–126, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002162.

Zhu:2015:APV

[ZL15c] Song-Ping Zhu and Guang-Hua Lian. Analytically pricing volatility swaps under stochastic volatility. Journal of Computational and Applied Mathematics, 288(??):332–340, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002551.

Zhao:2017:GOA

[ZL17a] Yingfeng Zhao and Sanyang Liu. Global optimization algorithm for mixed integer quadratically constrained quadratic program. Journal of Computational and Applied Mathematics, 319(??):159–169, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300018.

Zheng:2017:ESS

[ZL17b] Qingqing Zheng and Linzhang Lu. Extended shiftsplitting preconditioners for saddle point problems. Journal of Computational and Applied Mathematics, 313(??):70–81, March 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304289.

Zhou:2017:MPM

[ZL17c] Zhongguo Zhou and Dong Liang. The mass-preserving and modified-upwind splitting DDM scheme for time-dependent convection– diffusion equations. Journal of Computational and Applied Mathematics, 317 (??):247–273, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042716305209.

Zainali:2018:DSQ

[ZL18a] Nahid Zainali and Taher Lotfi. On developing a stable and quadratic convergent method for solving absolute value equation. Journal of Computational and Applied Mathematics, 330(??):742–747, March 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303503.

Zhang:2018:SWG

[ZL18b] Tie Zhang and Tao Lin. A stable weak Galerkin finite element method for Stokes problem. Journal of Computational and Applied Mathematics, 333(??):235– 246, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305514.

Zhang:2019:AWG

[ZL19] Tie Zhang and Tao Lin. An analysis of a weak Galerkin finite element method for stationary Navier–Stokes problems. Journal of Computational and Applied Mathematics, 362(??):484–497, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718304631.

Zhu:2015:EAR

[ZLAA15] Jiehua Zhu, Xiezhang Li, Fangjun Arroyo, and Edward Arroyo. Error analysis of reweighted l_{1} greedy algorithm for noisy reconstruction. Journal of Computational and Applied Mathematics, 286(??):93–101, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001089.

Zhang:2018:EST

[ZLF18] Guo-Bao Zhang, Yan Li, and Zhaosheng Feng. Exponential stability of traveling waves in a nonlocal dispersal epidemic model with delay. Journal of Computational and Applied Mathematics, 344(??):47–72, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302899.

Zhao:2019:ELB

[ZLPF19] Zhi-Wen Zhao, Yang Li, Cui-Xin Peng, and Zhi-Hui Fu. Empirical likelihood-based inference in generalized random coefficient autoregressive model with conditional moment restrictions. Journal of Computational and Applied Mathematics, 348(??):146–160, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305351.

Zhang:2015:NMNa

[ZLS^{+}15] Yaoming Zhang, Xiaochao Li, Vladimir Sladek, Jan Sladek, and Xiaowei Gao. A new method for numerical evaluation of nearly singular integrals over high-order geometry elements in 3D BEM.

Journal of Computational and Applied Mathematics, 277(??):57–72, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714003847.

Zhang:2017:UEE

[ZLZ17] Xiaolong Zhang, Songxin Liang, and Li Zou. Uniqueness and error estimates for solutions to higherorder boundary value problems. Journal of Computational and Applied Mathematics, 321(??):44– 59, September 2017. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717300882.

Zhang:2015:BIE

[ZM15a] Lei Zhang and FuMing Ma. Boundary integral equation methods for the scattering problem by an unbounded sound soft rough surface with tapered wave incidence. Journal of Computational and Applied Mathematics, 277(??):1–16, March 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/

science/article/pii/
S0377042714003835.

Zhang:2015:PRK

[ZM15b] Limei Zhang and Fuming Ma. Pouzet–Runge– Kutta–Chebyshev method for Volterra integral equations of the second kind. Journal of Computational and Applied Mathematics, 288(??):323–331, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715002654.

Zhang:2016:SMF

[ZM16a] Shangyou Zhang and Mo Mu. Stable Q_{k}–Q_{k−1} mixed finite elements with discontinuous pressure. Journal of Computational and Applied Mathematics, 301(??):188–200, August 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300097.

Zheng:2016:CTS

[ZM16b] Qing-Qing Zheng and Chang-Feng Ma. A class of triangular splitting methods for saddle point problems. Journal of Computational and Applied Mathematics, 298(??):13–23, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005804.

Zhou:2016:NNF

[ZML16] Xinchen Zhou, Zhaoliang Meng, and Zhongxuan Luo. New nonconforming finite elements on arbitrary convex quadrilateral meshes. Journal of Computational and Applied Mathematics, 296(??):798– 814, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500549X.

Zhang:2019:CAI

[ZMTT19] Li Zhang, Huanhuan Ma, Shuo Tang, and Jieqing Tan. A combined approximating and interpolating ternary 4-point subdivision scheme. Journal of Computational and Applied Mathematics, 349(??):563–578, March 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305594.

Zolfaghari:2019:SAL

[ZN19] Reza Zolfaghari and Nedialko S. Nedialkov. Structural analysis of linear integral-algebraic equations. Journal of Computational and Applied Mathematics, 353(??):243– 252, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271830774X.

Zhang:2015:AFE

[ZNG15] Weiwei Zhang, Yufeng Nie, and Yuantong Gu. Adaptive finite element analysis of elliptic problems based on bubbletype local mesh generation. Journal of Computational and Applied Mathematics, 280(??):42–58, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042714005287.

Zhang:2018:BDR

[ZNS18] Fode Zhang, Hon Keung Tony Ng, and Yimin Shi. Bayesian duality and risk analysis on the statistical manifold of exponential family with censored data. Journal of Computational and Applied Mathematics, 342(??):534–549, November 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302243.

Zia:2016:RTN

[ZP16] Qazi Muhammad Zaigham Zia and Roland W. E. Potthast. The range test and the no response test for Oseen problems: Theoretical foundation. Journal of Computational and Applied Mathematics, 304(??):201–211, October 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271500583X.

Zhao:2019:PPE

[ZP19] Lina Zhao and Eun-Jae Park. A priori and a posteriori error analysis of a staggered discontinuous Galerkin method for convection dominant diffusion equations. Journal of Computational and Applied Mathematics, 346(??):63–83, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

science/article/pii/
S0377042718303923.

Zeng:2017:GGS

[ZPL17] Jinshan Zeng, Zhiming Peng, and Shaobo Lin. GAITA: a Gauss–Seidel iterative thresholding algorithm for ℓ_{1} regularized least squares regression. Journal of Computational and Applied Mathematics, 319(??):220–235, August 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717300146. See

corrigendum [ZPL18].

Zeng:2018:CGG

[ZPL18] Jinshan Zeng, Zhimin Peng, and Shaobo Lin. Corrigendum to “GAITA: a Gauss–Seidel iterative thresholding algorithm for ℓ_{1} regularized least squares regression” [J. Comput. Appl. Math. 319 (2017) 220–235]. Journal of Computational and Applied Mathematics, 333(??):458, May 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271730585X. See

[ZPL17].

Zhang:2018:SCH

[ZQJY18] Tong Zhang, Yanxia Qian, Tao Jiang, and JinYun Yuan. Stability and convergence of the higher projection method for the time-dependent viscoelastic flow problem. Journal of Computational and Applied Mathematics, 338(??):1–21, August 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271830013X.

Zhang:2015:SAA

[ZqZwZ15] Jie Zhang, Ya qi Zhang, and Li wei Zhang. A sample average approximation regularization method for a stochastic mathematical program with general vertical complementarity constraints. Journal of Computational and Applied Mathematics, 280(??):202–216, May 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714005366.

Zhang:2017:FSO

[ZS17a] Sumei Zhang and Yudong Sun. Forward starting options pricing with double stochastic volatility, stochastic interest rates and double jumps. Journal of Computational and Applied Mathematics, 325(??):34–41, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301681.

Zhu:2017:PFP

[ZS17b] Yao Zhu and Ahmed H.

Sameh. PSPIKE+: a family of parallel hybrid sparse linear system solvers. Journal of Computational and Applied Mathematics, 311(??):682–703, February 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716304022.

Zolfaghari:2017:IFE

[ZS17c] Mehdi Zolfaghari and Bahram Sahabi. Impact of foreign exchange rate on oil companies risk in stock market: a Markov-switching approach. Journal of Computational and Applied Mathematics, 317(??):274– 289, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305015.

Zeinali:2018:ELC

[ZS18a] Masoumeh Zeinali and Sedaghat Shahmorad. An equivalence lemma for a class of fuzzy implicit integro-differential equations. Journal of Computational and Applied Mathematics, 327(??):388–399, January 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730300X.

Zhang:2018:BRP

[ZS18b] Zhengyi Zhang and Ahmed H. Sameh. Block row projection method based on M-matrix splitting. Journal of Computational and Applied Mathematics, 340(??):731–744, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717304041.

Zheng:2018:UDP

[ZS18c] Peng Zheng and Xiquan Shi. The use of the dyadic partition in elementary real analysis. Journal of Computational and Applied Mathematics, 329(??):344–352, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302467.

Zhang:2019:EGS

[ZS19] Zhimin Zhang and Wen Su. Estimating the Gerber–Shiu function in a Lévy risk model by Laguerre series expansion. Journal of Computational and Applied Mathematics, 346(??):133–149, January 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718304230.

Zhang:2015:ESE

[ZSL15] G. L. Zhang, M. H. Song, and M. Z. Liu. Exponential stability of the exact solutions and the numerical solutions for a class of linear impulsive delay differential equations. Journal of Computational and Applied Mathematics, 285(??):32–44, September 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715000461.

Zhang:2018:SCP

[ZSL18] Wei Zhang, M. H. Song, and M. Z. Liu. Strong convergence of the partially truncated Euler– Maruyama method for a class of stochastic differential delay equations. Journal of Computational and Applied Mathematics, 335(??):114– 128, June 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042717305903.

Zhang:2015:AAM

[ZST15] Hong Zhang, Adrian Sandu, and Paul Tranquilli. Application of approximate matrix factorization to high order linearly implicit Runge– Kutta methods. Journal of Computational and Applied Mathematics, 286(??):196–210, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271500151X.

Zhang:2016:SIC

[ZSW16] Chunfang Zhang, Yimin Shi, and Min Wu. Statistical inference for competing risks model in step-stress partially accelerated life tests with progressively Type-I hybrid censored Weibull life data. Journal of Computational and Applied Mathematics, 297(??):65– 74, May 1, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715005476.

Zhang:2017:GAM

[ZSZ17] Fode Zhang, Yimin Shi, and Chunfang Zhang. Geometry of an accelerated model with censored data. Journal of Computational and Applied Mathematics, 317(??):137– 145, June 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716305921.

Zhang:2018:RFO

[ZT18] Tie Zhang and Can Tong.

A remark on the fractional order differential equations. Journal of Computational and Applied Mathematics, 340(??):375–379, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/

                                        science/article/pii/
                                        S0377042718301328.

Zhang:2018:GBS

[ZTGZ18] Li Zhang, Jieqing Tan, Xianyu Ge, and Guo Zheng. Generalized Bsplines’ geometric iterative fitting method with mutually different weights. Journal of Computational and Applied Mathematics, 329(??):331–343, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730287X.

Zaman:2017:ENM

[ZuI17] Sakhi Zaman and Siraj ul Islam. Efficient numerical methods for Bessel type of oscillatory integrals. Journal of Computational and Applied Mathematics, 315(??):161–174, May 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042716305210.

Zhang:2016:TBI

[ZW16] Kaili Zhang and Yiju Wang. An H-tensor based iterative scheme for identifying the positive definiteness of multivariate homogeneous forms.

Journal of Computational and Applied Mathematics, 305(??):1–10, October 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716301522.

Zhang:2018:AIH

[ZWGL18] Chao Zhang, Li-Lian Wang, Dongqin Gu, and Wenjie Liu. On approximate inverse of Hermite and Laguerre collocation differentiation matrices and new collocation schemes in unbounded domains. Journal of Computational and Applied Mathematics, 344 (??):553–571, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718303388.

Zong:2015:EMS

[ZWH15a] Xiaofeng Zong, Fuke Wu, and Chengming Huang. Exponential mean square stability of the theta approximations for neutral stochastic differential delay equations. Journal of Computational and Applied Mathematics, 286(??):172–185, October 1, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001624.

Zong:2015:TSS

[ZWH15b] Xiaofeng Zong, Fuke Wu, and Chengming Huang. Theta schemes for SDDEs with non-globally Lipschitz continuous coefficients. Journal of Computational and Applied Mathematics, 278(??):258–277, April 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042714004609.

Ziatdinov:2015:TPH

[ZwKN15] Rushan Ziatdinov, Tae wan Kim, and Rifkat I. Nabiyev. Two-point G^{1} Hermite interpolation in biangular coordinates. Journal of Computational and Applied Mathematics, 287(??):1–11, October 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715001107.

Zhou:2018:BQI

[ZWL18] Jian-Ping Zhou, Ren-Hong Wang, and Chong-Jun Li.

The bivariate quadratic C^{1} spline spaces with stable dimensions on the triangulations. Journal of Computational and Applied Mathematics, 329(??):364–373, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717303096.

Zhang:2019:GAM

[ZWT19] Chun-Hua Zhang, Xiang Wang, and Xiao-Bin Tang. Generalized AOR method for solving a class of generalized saddle point problems. Journal of Computational and Applied Mathematics, 350(??):69– 79, April 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718306022.

Zhang:2018:LCM

[ZWW^{+}18] Na Zhang, Yating Wang, Yuhe Wang, Bicheng Yan, and Qian Sun. A locally conservative Multiscale Finite Element Method for multiphase flow simulation through heterogeneous and fractured porous media. Journal of Computational and Applied Mathematics, 343(??):501–519, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302759.

Zheng:2019:SAD

[ZWWW19] Wei Zheng, Hongbin Wang, Hongrui Wang, and Shuhuan Wen. Stability analysis and dynamic output feedback controller design of T-S fuzzy systems with time-varying delays and external disturbances. Journal of Computational and Applied Mathematics, 358(??):111–135, October 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271930038X.

Zong:2018:CST

[ZWX18] Xiaofeng Zong, Fuke Wu, and Guiping Xu. Convergence and stability of two classes of theta-Milstein schemes for stochastic differential equations. Journal of Computational and Applied Mathematics, 336 (??):8–29, July 2018. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042717306532.

Zheng:2016:CPP

[ZWY16] Zhihao Zheng, Guozhao Wang, and Ping Yang. On control polygons of Pythagorean hodograph septic curves. Journal of Computational and Applied Mathematics, 296 (??):212–227, April 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0377042715004586.

Zhou:2018:SSU

[ZWZ^{+}18] Jun Zhou, Weiming Wang, Jie Zhang, Baocai Yin, and Xiuping Liu. 3D shape segmentation using multiple random walkers. Journal of Computational and Applied Mathematics, 329(??):353–363, February 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717301449.

Zhao:2016:PSP

[ZX16] Pan Zhao and Qingxian Xiao. Portfolio selection problem with value-at-risk constraints under nonextensive statistical mechanics. Journal of Computational and Applied Mathematics, 298(??):64– 71, May 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715006081.

Zhang:2019:FMM

[ZX19] Qingyang Zhang and Shuhuang Xiang. On fast multipole methods for Volterra integral equations with highly oscillatory kernels. Journal of Computational and Applied Mathematics, 348(??):535–554, March 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305545.

Zhang:2017:FFM

[ZXQH17] Ying Zhang, Yingtao Xu, Qiusheng Qiu, and Xiaowei He. A filled function method for minimizing control variation in constrained discrete-time optimal control problems. Journal of Computational and Applied Mathematics, 326(??):126–137, December 15, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302686.

Zhao:2018:SNM

[ZY18] Hong-Jie Zhao and Haijian Yang. Semismooth Newton methods with domain decomposition for American options. Journal of Computational and Applied Mathematics, 337(??):37–50, August 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718300141.

Zhang:2019:VGE

[ZY19a] Zhimin Zhang and Yaodi Yong. Valuing guaranteed equity-linked contracts by Laguerre series expansion. Journal of Computational and Applied Mathematics, 357(??):329–348, September 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719301141.

Zhou:2019:RAW

[ZY19b] Zhiyong Zhou and Jun Yu. Recovery analysis for weighted mixed l^{2}/l^{p} minimization with 0 < p ≤ 1. Journal of Computational and Applied Mathematics, 352(??):210–222, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307210.

Zhang:2019:MVE

[ZYF19] Baiju Zhang, Yan Yang, and Minfu Feng. Mixed virtual element methods for elastodynamics with weak symmetry. Journal of Computational and Applied Mathematics, 353(??):49– 71, June 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718307519.

Zygmunt:2017:CAA

[Zyg17] Mariusz Zygmunt. Circular arc approximation using polygons. Journal of Computational and Applied Mathematics, 322(??):81–85, October 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S037704271730153X.

Zhang:2018:TCO

[ZYL18] Wei Guo Zhang, Xing Yu, and Yong Jun Liu. Trade and currency options hedging model. Journal of Computational and Applied Mathematics, 343(??):328–340, December 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718302620.

Zhang:2019:NST

[ZYL19] Chao Zhang, Hanfeng Yao, and Huiyuan Li. New space-time spectral and structured spectral element methods for high order problems. Journal of Computational and Applied Mathematics, 351(??):153–166, May 1, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718305259.

Zhang:2016:MCV

[ZYXT16] Ying Zhang, Changjun Yu, Yingtao Xu, and Kok Lay Teo. Minimizing control variation in discrete-time optimal control problems. Journal of Computational and Applied Mathematics, 292(??):292–306, January 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003726.

Zhang:2019:MQI

[ZYY19] Shengliang Zhang, Hongqiang Yang, and Yu Yang. A multiquadric quasiinterpolations method for CEV option pricing model. Journal of Computational and Applied Mathematics, 347(??):1–11, February 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302279.

Zhang:2018:PCM

[ZYZ18] Jiansong Zhang, Danping Yang, and Jiang Zhu. Parallel characteristic mixed element method for saltwater intrusion problem. Journal of Computational and Applied Mathematics, 336(??):160– 174, July 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718300013.

Zhu:2015:CIM

[ZZ15] Mu-Zheng Zhu and Guo-Feng Zhang. A class of iteration methods based on the HSS for Toeplitz systems of weakly nonlinear equations. Journal of Computational and Applied Mathematics, 290 (??):433–444, December 15, 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042715003258.

Zhao:2016:INC

[ZZ16] Xuan-Yi Zhao and Chun-Gang Zhu. Injectivity of NURBS curves. Journal of Computational and Applied Mathematics, 302(??):129–138, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300267.

Zhu:2018:NCS

[ZZ18] Huijian Zhu and Caibin Zeng. A novel chaotification scheme for fractional system and its application. Journal of Computational and Applied Mathematics, 339(??):275–284, ???? 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717304983.

Zeng:2019:MLI

[ZZ19] Min-Li Zeng and Guo-Feng Zhang. An MHSSlike iteration method for two-by-two linear systems with application to FDE optimization problems. Journal of Computational and Applied Mathematics, 352(??):368–381, May 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718307209.

Zhao:2019:LMB

[ZZG^{+}19] Yong-Liang Zhao, Pei-Yong Zhu, Xian-Ming Gu, Xi-Le Zhao, and Jianxiong Cao. A limitedmemory block bi-diagonal Toeplitz preconditioner for block lower triangular Toeplitz system from time-space fractional diffusion equation. Journal of Computational and Applied Mathematics, 362(??):99–115, December 15, 2019. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL https://

www.sciencedirect.com/
science/article/pii/
S0377042719302547.

Zhou:2017:SSR

[ZZHS17] Weien Zhou, Jingjing Zhang, Jialin Hong, and Songhe Song. Stochastic symplectic Runge– Kutta methods for the strong approximation of Hamiltonian systems with additive noise. Journal of Computational and Applied Mathematics, 325(??):134–148, December 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042717302285.

Zhang:2016:OSC

[ZZLY16] Rongpei Zhang, Jiang Zhu, Abimael F. D. Loula, and Xijun Yu. Operator splitting combined with positivity-preserving discontinuous Galerkin method for the chemotaxis model. Journal of Computational and Applied Mathematics, 302(??):312–326, August 15, 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042716300656.

Zhang:2016:CCN

[ZZW16] Qifeng Zhang, Chengjian Zhang, and Li Wang. The compact and Crank– Nicolson ADI schemes for two-dimensional semilinear multidelay parabolic equations. Journal of Computational and Applied Mathematics, 306 (??):217–230, November 2016. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0377042716301893.

Zhou:2018:SCB

[ZZW18] Sanyu Zhou, Jingjing Zhu, and Defa Wang. Simultaneous confidence bands for a percentile hyper-plane with covariates constrained in a restricted range. Journal of Computational and Applied Mathematics, 344 (??):301–312, December 15, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042718302978.

Zheng:2015:BAS

[ZZZ15] Zhong Zheng, Guo-Feng Zhang, and Mu-Zheng Zhu. A block alternating splitting iteration method for a class of block twoby-two complex linear systems. Journal of Computational and Applied Mathematics, 288(??):203–214, November 2015. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

                                        www.sciencedirect.com/
                                        science/article/pii/
                                        S0377042715002150.

Zhang:2017:CDG

[ZZZ^{+}17] Jiansong Zhang, Jiang Zhu, Rongpei Zhang, Danping Yang, and Abimael F. D. Loula. A combined discontinuous Galerkin finite element method for miscible displacement problem. Journal of Computational and Applied Mathematics, 309(??):44–55, January 1, 2017. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S037704271630293X.

Zhang:2018:IFE

[ZZZ18] Yamiao Zhang, Jiazhong Zhang, and Lianning Zhu. Iterative finite element variational multiscale method for the incompressible Navier– Stokes equations. Journal of Computational and Applied Mathematics, 340(??):53–70, October 1, 2018. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http://

www.sciencedirect.com/
science/article/pii/
S0377042718301031.